A Complete Bibliography of Publications in Journal of Fish Diseases (2010–2019)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

29 August 2024 Version 1.06

Title word cross-reference

−20 [HPAL10]. −80 [SG12]. 2 [HST^{+}13]. ^{N438A} [CGW^{+}19]. _{1} [BSD^{+}18d]. _{18} [LZZN10]. _{4} [LZZN10]. α [LLZW14, TQW^{+}14]. α_{s2} [SJDC^{+}13]. β [CKTT10, CPB^{+}15, CMR^{+}16, HBH^{+}19, JSAH^{+}18, KBE^{+}14, PLCC11, SCS^{+}12, TQW^{+}14, UN10, UKS^{+}18, YSHG18]. γ [TQW^{+}14]. μ [BVL^{+}18, GVLE16]. × [AZSSZZ13, BSK^{+}10, DBS12, FCS^{+}18, LTC^{+}14, RHSM^{+}18b, SSF^{+}17, SL15, SRT^{+}14, ZDWH18]. ε [RBL^{+}11].

-1 [JSAH^{+}18, YSHG18]. -6-fluoro [SRD^{+}16]. -D-glucans [KBE^{+}14]. -G [CGW^{+}19]. -glucan [UKS^{+}18]. -haemoglobin [UN10]. -hydroxybutyrate [SCS^{+}12]. -like [BRB^{+}12, ZGY^{+}18]. -proteobacteria [CMR^{+}16].

. [ATP^{+}17, SPKS17, YFA^{+}17].

1 [CKTT10, EDP^{+}17, GVLE16, GCCLRP^{+}13, KYK^{+}12, KAST17, LASRA^{+}19, MMNL14, NJK^{+}17, SHKF16, UN10, VCS^{+}19].

1-chloropiperidin-4-yl [SRD^{+}16]. 12 [DJR^{+}19, MMNL14]. 16S [DPMK12, LWWS14, LGD^{+}17, RB15]. 16S-RFLP [LGD^{+}17].

2 [AHJS^{+}18, DKBS17, GM10, NSS^{+}17, PMW^{+}10, RAM^{+}18]. 23S [DPMK12]. 2L [LCC^{+}15].

3 [CWW^{+}18, IFF^{+}15, PFP^{+}13]. 3- [SRD^{+}16]. 3-glucan [YSHG18]. 3-year [KZJ^{+}17]. 3/1 [JSAH^{+}18]. 35-year-old [WGL17]. 36 [APG^{+}11, HO13].

5 [VSS17]. 5’-monophosphate [VSS17]. 55L [HLC^{+}15]. 5p [FBK^{+}16].

6 [GCNB16]. 6- [KBE^{+}14]. 6-glucan [JSAH^{+}18]. 610 [RGD^{+}12]. 64L [CCLL16].

73 [YK12].

8 [LRD^{+}19]. 89 [RSIP^{+}17, RSPA^{+}18].

90 [RSIP^{+}17, RSPA^{+}18, SHO^{+}17]. 97L [HLC^{+}15].

A-I [LSC^{+}14]. A-layer [GLK^{+}16]. A. [CSA^{+}15]. A1 [ZWG^{+}17]. abalone [JYZ^{+}14]. ABC [OJRL10]. ABC-transporter [OJRL10]. abdominal [GHB^{+}11]. abdominalis [LYD^{+}15]. Abilgaard [SAD^{+}12]. ability [CD17, PST^{+}16, SWM^{+}15]. abiotic [SHV^{+}16]. abnormalities [FPK^{+}17, MPS^{+}18]. abrupt [LTC17]. abscessus [PBW^{+}18]. absence [GCS^{+}18, JLJ16]. abundance [GBV^{+}16, GVB^{+}17a, GAVC16, JMF^{+}18, LDÁM14]. Acanthopagrus [Dig13, Dig15, LLDL13, Pol15]. acarne [RVB13]. Accelerated [AGB^{+}19]. accuracy [LTG^{+}18, LMG^{+}19]. accurate [KPSP^{+}19]. achromogenes [BBP18, SWM^{+}15]. Achyranthes [CR12]. Acid [ASHZ^{+}12, ACYC10, AMRHM18, BCBS^{+}13, DTF^{+}16, GRT12, JKL^{+}13, KSM^{+}17, MMNL14, PSBG^{+}11, PCK^{+}15, SOGR^{+}19, SMFM12, TNK12]. acid-specific [ACYC10]. acidification [ASW16, CBM^{+}17]. acidovorans [ARFG13]. acids [MLF^{+}10]. Acipenser [LGK^{+}14, MMD^{+}12, RHSM^{+}18b, RFA^{+}14, SAD^{+}12]. Acipenserid [DKBS17]. acquisition [BLF^{+}16, BPMI^{+}16, BAMYAH19]. act [PJL^{+}12]. Actinochlamydia [SPV^{+}18]. Actinochlamydiaceae [SPV^{+}18]. activation [RCGGM15]. active [LBF^{+}14]. activities [BSD^{+}18a, BA14, HDC18, YLX^{+}19b]. activity [BSSB18, BSD^{+}18d, BYAS^{+}19, BLFC12, PMH^{+}11, PSBG^{+}11, RBH^{+}17, SJDC^{+}13, SLPB18, SBW11, TFG19, UN10, YSK13, ZWG^{+}17]. acts [PBC15]. aculeatus [ASW16, PJZB19, WFH^{+}18]. Acute [HCCM12, PRT^{+}14, PÅD18, Bin14, BPK^{+}17, CCLGRM^{+}17, CZB^{+}11, DD18, DJR^{+}19, KAF^{+}19, KTN^{+}16, KYB^{+}15, MKY^{+}15, NPGA^{+}19, TTL^{+}16]. acute-phase [DJR^{+}19]. acutely [PRSK10]. acutus [KJB11]. Acyclovir [TKS^{+}18]. adapt [HGH^{+}18]. adaptation [MLH^{+}15]. additional [MJ17]. additives [AA10]. adenocarcinoma [RHSM^{+}18a]. adenoma [SBHC18]. adenosine [BSD^{+}18a, VSS17]. adenosinergic [BSD^{+}18b]. adherence [JSJR^{+}19, PDLW17, VMBGIC19]. Adhesion [HRNW15, HGW^{+}19, YZW^{+}19]. adipose [ASSP^{+}15, BHW^{+}11]. adjuvant [PVIM18]. adjuvanted [CLS^{+}15]. administered [FHD13, GSS^{+}10, HJN^{+}19, TNK12]. administration [CMH^{+}17, DPCdR14, ESV^{+}19b, FGJ^{+}10, FHZ^{+}16, LWJ^{+}16, LXHL11, OKS^{+}13, PPI^{+}13, dOAPLá^{+}17]. adomavirus [PWW19]. adult [BHT^{+}17, BVT^{+}18, FSV^{+}14, HKD^{+}10, JHM^{+}18, KCD^{+}14, MIO^{+}18, PPE^{+}18, RBAT^{+}17, RAM^{+}17, SSK17]. adulterant [PNBK13]. aduncum [MGH^{+}14]. Advice [TJR^{+}18]. aeglefinus [JHM^{+}18]. Aequidens [YCNW18]. aer [FMA^{+}19]. Aerococcus [CG11]. aerocystitis [CBH^{+}15, dSRSHGá^{+}19]. aerolysin [FMA^{+}19, MBB12]. aerolysin-like [MBB12]. aeromonads [PAZ18]. Aeromonas [AA10, BSV^{+}18, BSB^{+}18, BKS^{+}17, BKD17, BHF13, BLFRCQ10, BVK^{+}16, BE19, BBP18, BGCKF12, CHH^{+}14, CMSHS16, CLZ^{+}12, CD17, CYS17, ÇOÇ^{+}16, CCMA^{+}17, CTK^{+}10, DBS12, DLCF13, DTJ^{+}17, DSJA18, EDP^{+}17, EDV^{+}18, FCC^{+}13, FMA^{+}19, GDN^{+}16, GLK^{+}16, HKK^{+}11a, JAR^{+}11, JSAH^{+}18, JSJR^{+}19, KEY^{+}15, KBJ^{+}13, KBD13, LPA^{+}10, LL12, LLZW14, LCZ^{+}18, LGC^{+}19, LCR^{+}10, LHW^{+}15b, MSC^{+}17, MLEM16, MP18, NB11, NLC^{+}16, NDA10, OSAA18, PSCFB^{+}18, PFJ^{+}10, RKK^{+}19, RVC18, RPM^{+}19, SSHEM11, SWM^{+}15, SSC17, SJSA^{+}14, VSS17, VSDS^{+}19a, VDD^{+}18, VC17, WMD^{+}18, WYZ^{+}19, YFA^{+}17, YZW^{+}19, YK12, ZCJ^{+}18, ZWL^{+}16]. aeruginosa [BSSB18, BSD^{+}18c]. aestuarianus [ZQB^{+}11]. aetiological [CCLGRM^{+}17, QXW14]. aetiology [LGS^{+}14]. affect [FCJH19, HHO14, JOF^{+}19, SFR^{+}17]. affected [ACN12, CSS^{+}16, DPTFC10, HOHG13, JR19, OCLI^{+}19, RHS^{+}12, SJT14]. affecting [JFWH19, RAM^{+}18, RSF^{+}17, SVM^{+}18]. aflatoxin [BSD^{+}18d]. Africa [HGPD11, HOHG13, HWO^{+}17, MCCN14]. African [BHA^{+}11, HOHG13, MWS17, OAO^{+}12, RSR^{+}18, WTM14]. after [BHO^{+}19, DPCdR14, ELI^{+}17, FHZ^{+}16, GFM^{+}16, GMN12, HKBD15b, JMF^{+}18, JSAH^{+}18, MID^{+}19b, MID^{+}19a, NSE^{+}17, TQW^{+}14, TMC^{+}12, VCN^{+}10, YYW^{+}14]. against [ASHZ^{+}12, ANB^{+}19, BBF^{+}15, BYAS^{+}19, BVK^{+}16, BE19, Buc12, BLFC12, CCLGRM^{+}17, CHH^{+}14, CHV^{+}12, CCH^{+}19, CGR17, CCLL16, CGW^{+}19, CKTT10, CKvGJB17, ÇOÇ^{+}16, CSCCA14, CSSWK10, DD18, DLP^{+}13, ESA^{+}10, EDV^{+}18, ERÁpSB14, FLCC^{+}14, FHD13, GCNB16, HDC18, HLL^{+}12, HCL^{+}14, HJN^{+}19, HLC^{+}15, HBH^{+}19, JST^{+}19, JAR^{+}11, JOF^{+}19, KSS^{+}15, LPA^{+}10, LKS^{+}13, LRC^{+}18, LXHL11, LHW^{+}12, LLL^{+}18a, MMS^{+}16, MTK^{+}10, MID^{+}19b, MA16, NPGA^{+}19, NNC^{+}18, OMB^{+}19, PZL^{+}14, PSBG^{+}11, RPMRLJ^{+}19, RSGB^{+}14, RBH^{+}17, SHS14, SJDC^{+}13, SMR10, SGB^{+}16, SLKE10, SSSP18, SNM^{+}16, SWS^{+}16, SRD^{+}16, SRB12, TFG19, VDO^{+}18b, VDO^{+}18a, WBF^{+}10, WCR^{+}18, WM19, WHR^{+}18, WSA^{+}18, XZZ16, YLX^{+}19b, ZWG^{+}18]. agalactiae [DDBC^{+}15, AAZS^{+}15, AZSI^{+}12, AZSSZZ13, ATP^{+}17, BSD^{+}18b, BMMC^{+}15, BFFA^{+}12, BDB^{+}18, CWL^{+}17, CZL^{+}19, DZC^{+}16, HZX^{+}13, ICVV16, LLH^{+}14a, LLH^{+}14b, LLX^{+}19, LLL^{+}18a, MoSMS18, NYRS12, SFL^{+}16, SPS^{+}19, VMBGIC19, VJWC^{+}18, XXL^{+}19, ZKL^{+}19]. Agassiz [RHSM^{+}18a]. AGD [ACN12, CHWM^{+}19]. Age [JHW17]. agent [ACM^{+}14, CCLGRM^{+}17, CPP14, GTT^{+}17, GCGR13, HWM^{+}13, JFWH19, KEY^{+}15, MFS^{+}17, MSS^{+}17a, NLF11, RGD^{+}13, SSF^{+}17, SWD^{+}14, TVK^{+}19, WCR^{+}18, WEPB^{+}18, ZWG^{+}17]. agent-based [GCGR13]. agents [HH13a, HBS^{+}14, SSC17]. agreement [JWO^{+}10]. AHPND [BPK^{+}17, ITKH18]. AIP1 [SRSF10]. air [LWMT10]. Aires [Lek12]. akaara [GHH^{+}11]. al [JCN^{+}14, KRFT14]. AL09 [YK12]. AL09-73 [YK12]. Alaska [PPE^{+}18]. Albendazole [ANB^{+}19]. albigutta [DBHC17]. Albufera [BSC^{+}14]. Algamune^{T M} [YSHG18]. Algeria [KCD^{+}14]. alginolyticus [CLWJ13, CLW^{+}11, HGX^{+}19, LSC^{+}14, PKLH19, YLX^{+}19a, YLS^{+}19]. alive [CBM^{+}18]. Alix [SRSF10]. Alix/ [SRSF10]. alleles [LLZW14]. allicin [NDA10]. allis [WBE^{+}17, WEPB^{+}18]. Allium [MSCH14]. allogynogenetic [LJHG14, ZWLG10, ZYW^{+}10]. along [FSM^{+}16]. Alosa [WBE^{+}17, WEPB^{+}18]. alpha [LSA^{+}15]. alpha-2-macroglobulin [LSA^{+}15]. alphavirus [Ano10c, BTC^{+}17, CGR17, GFW^{+}10, GFM^{+}11, GFR^{+}12, GRF14, HFT^{+}12, HST^{+}13, HJT^{+}16, JTW^{+}10, JJB15, MIO^{+}18, RSP^{+}17, RSM^{+}18, SBH^{+}16, SHV^{+}16, TJB^{+}15]. alpinus [SWM^{+}15, WKK12]. also [GCM^{+}18b]. alter [PPF^{+}12]. alterations [LvD18]. altered [SSK17]. alternative [CB13]. alternifolia [BSD^{+}18c, VGI^{+}16]. altivelis [FNISE12, KSS^{+}15, ZCC18, CONN10, FNISE12, KSS^{+}15, ZCC18]. Aluminium [MGF17]. alvei [OSC^{+}14]. Amago [IOO15]. Amazon [MdOO16, OMB^{+}19, PdSG^{+}12, dSdAdS^{+}19]. amberjack [MKH^{+}11, NMM^{+}18, dSRSHGá^{+}19]. ambigua [HBB14]. Ameiurus [BPF^{+}18, PHR14]. amended [GDN^{+}16]. America [EKT^{+}11]. American [CG11, DKBS17, GL18, KPKP10, KWD^{+}16, LML12, SQC^{+}14]. American-HPR0 [GL18]. americanus [CG11, SQC^{+}14]. amine [FHZ^{+}16]. Amino [SOGR^{+}19]. ammonia [FSO^{+}10, SRB12]. Ammonium [RHS^{+}12, MoSMS18]. amoeba [DT16, LTC16, LTC17]. amoebae [FMM^{+}19]. Amoebic [CONN10, TCC10a, ACN12, BPCS19, DYML^{+}18, MLER^{+}18, MCCN14, NBJ10, WNO^{+}18]. Amoebozoa [CPP14]. among [ADR11, OEG^{+}13, PPSW19, SMJ^{+}12, SCD19, SVY^{+}14]. amongst [KBTMH^{+}16]. Amphiprion [BRMG13, SBK^{+}15]. amplification [HJOK11, HGG^{+}11, JYZ^{+}14, KTN^{+}16, LLWL15, MSI^{+}15, OBR^{+}14, WSX^{+}18, YWW^{+}19]. ampullary [CBC^{+}13]. Amyloid [CBFM16, RTM^{+}15]. Amyloid-like [CBFM16]. Amyloodinium [MSS^{+}17b, YB17]. anadromous [áKÁGA16]. Anaemia [CCCTA^{+}14, LLLG18, ADF14, ADMF15, ATM^{+}10, EYF^{+}19, GL18, GCS^{+}18, HWM^{+}13, IOO15, MMWK19, MAC^{+}15, MTHB13, OBFR12, SG12, VDMJ^{+}13, XQS^{+}15]. anal [CAJD^{+}18, FPK^{+}17]. analyses [NRMS14, RHVG^{+}19b, STIG17, YEK^{+}17]. Analysing [MLH^{+}15]. Analysis [FCL^{+}19, YMH^{+}11, ATM^{+}10, ACAS14, ASB^{+}12, BC11, CIL^{+}17, CGP^{+}19, CMEV17a, CMS^{+}17, CMEV17b, ECP^{+}10, FGJ^{+}10, FMA^{+}19, GPL^{+}12, IBFL19, JCN^{+}13, JCN^{+}14, JKH^{+}17, KKM^{+}19, KRFT14, LGD^{+}17, LCV^{+}17, LMG^{+}19, LITCAH19b, LKT^{+}13, MFS^{+}17, MBT^{+}17, MMH10, NSE^{+}17, NSC^{+}18, NMM^{+}18, PHvdH12, PKLH19, PSVK^{+}18, PJU^{+}13, RKK^{+}19, SHKF16, SOA^{+}18, TCFáS19, TDT^{+}16, VC17, WMM^{+}16, WMD^{+}18, WZA10, XJL^{+}19, YAHWU11, ZJR^{+}18, ZCJ^{+}18]. analytical [BHA^{+}17]. anatomy [FPK^{+}17, LGS^{+}14]. anchovy [RGMGG^{+}18]. Ancyrocephalus [BGRB14]. and/or [HKBD15a, HKBD15b]. angasi [BWP^{+}19b]. AngHV [KYK^{+}12, NJK^{+}17]. AngHV-1 [KYK^{+}12, NJK^{+}17]. angio [ASW16]. angio-architecture [ASW16]. angiolipoma [RhAS^{+}15]. anglerfish [FYO^{+}11, LRAP17]. Anguilla [ARFG13, BSC^{+}14, BMD^{+}14, BND^{+}17, BDM^{+}18, CSA^{+}15, DMB^{+}16, EMA17, FDST16, GFM^{+}16, KYK^{+}12, KPK^{+}18, KLD^{+}18, KAST17, LML12, MBD16, MFES18, MWS17, NJK^{+}17, OMO^{+}15, PWHW18b, SWE^{+}19, WCW^{+}16b]. anguillae [MWS17, NJK^{+}17]. anguillarum [ACLP^{+}19, BVK^{+}16, BFC^{+}15, CCH^{+}19, CKTT10, CDG^{+}14, FMB^{+}11, HRL15, JQK14, RGD^{+}12, RRO^{+}13, SCS^{+}12, TMC^{+}12, WM19, ZXW^{+}12]. anguillicaudatus [ZWL^{+}16]. Anguillicola [BMD^{+}14, BND^{+}17]. Anguillicoloides [FDST16, LML12, LFP^{+}11a, LFP^{+}11b]. anguillid [KYK^{+}12, vBVLR^{+}16, MFES18, KAST17]. anguilliseptica [ARFG13]. angulospora [SBP^{+}19]. animal [RAM^{+}19]. animals [LTG^{+}18, LTM^{+}19, SAS^{+}19, THW^{+}12]. anisakiasis [KHS11]. Anisakidae [MGH^{+}14, SDZN19]. Anisakis [MGH^{+}14]. anisitsi [Lek12]. annectens [RSR^{+}18]. anomalies [ACA^{+}16, BC11, dALF^{+}17]. anomaly [AAV^{+}16]. anterior [PAM^{+}19, PRSK10]. Anthony [Rob11]. Anthropogenic [KAST17, MMMD16]. anti [BSB^{+}18, DJQ^{+}12, MDF^{+}17, NPGA^{+}19]. anti-apoptotic [DJQ^{+}12]. anti-inflammatory [BSB^{+}18, MDF^{+}17]. Anti-PirA-like [NPGA^{+}19]. Antibacterial [LWJ^{+}16, LLL^{+}12, SHS14, ZZQ^{+}19]. Antibiotic [CDW17, CYS17, CIL^{+}17, FCL^{+}19, HKB^{+}16, NB11, RHBMM13, SK15, SFF^{+}17, ZCJ^{+}18, ZDWH18]. antibiotic-resistant [RHBMM13]. antibiotics [AAZS^{+}15, LRS^{+}17, SHS14]. antibodies [CLC^{+}17, ERÁpSB14, HLC^{+}15, LCR^{+}10, ZWG^{+}18]. Antibody [GSEM13, SSHEM11, Ano12c, BWZ^{+}17, BRB^{+}12, CGR17, CCLL16, ELI^{+}17, EDV^{+}18, KSS^{+}15, KJW^{+}17, MF12, PSK^{+}13, TCC10a]. antibody-based [PSK^{+}13]. Antibody-coated [SSHEM11]. Antifungal [BAH^{+}16]. antigen [CR12, DMD^{+}17, HCL^{+}14, MSC^{+}17, SGB^{+}16, TQW^{+}14, ZXW^{+}12]. antigenic [JKC^{+}11]. Antigens [MLEM16, áKÁGA16, KREM14, LMAKN16, RBAT^{+}17]. Antimicrobial [BYAS^{+}19, CMH^{+}17, DBV^{+}13, NSB^{+}18, DMP10, PMK12, RHTC19b, SW11, UN10, ZN10]. antioomycete [ZN10]. antioxidant [BSD^{+}18c, BA14]. antioxidative [LMC^{+}16]. Antiparasitic [UN10, HBS^{+}14, MDF^{+}17, ZN10]. antiparasitics [GOM^{+}19]. Antiviral [HDC18, RBH^{+}17, SJDC^{+}13, SVL^{+}17, YLX^{+}19b]. aorta [DKG^{+}17]. Aphanomyces [SMKBP17, VDHV^{+}16, VDH19]. Apicomplexa [MdOO16, dSdAdS^{+}19]. aplasia [FFP^{+}14]. apolipoprotein [LSC^{+}14]. Apoptosis [DRG^{+}17, CXW^{+}19, DJQ^{+}12, RBL^{+}11, WCL^{+}19]. apoptotic [DJQ^{+}12]. Apostichopus [LQL^{+}10]. apparatus [YK12]. applicability [CPdPP16]. Application [CDG^{+}14, LRS^{+}17, MMBS10, CLWJ13, CCCTA^{+}14, dSCBdS^{+}16, FSG^{+}12, GHW^{+}15, LML^{+}13, MoSMS18, NMIY14, PPK^{+}13, SCK13, WCR^{+}18, ZCC^{+}15]. Applications [Spe19a]. applied [GOM^{+}19, MSCH14]. approach [JMD^{+}18, LRMN17, LHW^{+}15b, OLVJ^{+}13, WWE^{+}13]. approaches [CRBLC^{+}17, ZXW^{+}12]. apta [YLX^{+}19a, ZLN^{+}17]. apta-sorbent [YLX^{+}19a, ZLN^{+}17]. aptamer [YGL^{+}18, YLX^{+}19a, ZLN^{+}17]. aptamer-based [YLX^{+}19a]. aptamers [HMP^{+}12, YLS^{+}19]. Aquaculture [Spe19a, BTN^{+}18, BMBS12, DVC^{+}17, GJMH17, GDW^{+}15, JJM^{+}17, JSAB^{+}19, KOS^{+}10, MSBB^{+}16, MBS^{+}17, MSCH14, OT12, PSVK^{+}18, QZZ^{+}18b, RAM^{+}18, SMR16, SBK^{+}15, TJA^{+}13, WCR^{+}18, ZCQF15]. Aquaflor(R) [DBS12]. AquaPathogen [EKT^{+}11]. aquareovirus [PWW19]. aquaria [SMJ^{+}14]. aquarium [DSJA18, KJB11, KZJ^{+}17, ONGQ^{+}14, PBW^{+}18, TIK^{+}14a]. aquarium-held [KJB11]. aquarium-maintained [TIK^{+}14a]. aquatic [EKT^{+}11, EMA17, LRS^{+}17, LTG^{+}18, LTM^{+}19, PJL^{+}12, THW^{+}12]. aqueous [YLX^{+}19b]. AQUI [CAS^{+}18]. AQUI-S(R) [CAS^{+}18]. Arabian [CIA^{+}18]. arabicum [CIA^{+}18]. Arborescent [OOO^{+}11]. archamoebae [CPdPP16]. architecture [ASW16]. archived [TVK^{+}19]. Arctic [KOB^{+}17, SWM^{+}15, SVL^{+}17, WKK12]. area [JR19, SMR16]. areas [DMK^{+}19]. arenaria [MMD16]. argenteus [LJC^{+}19]. Argentina [CCMA^{+}17, TB12]. argentine [TB12]. argus [DIH^{+}16, LL12]. Argyrosomus [EKS^{+}13, TAQ^{+}10]. aromatic [PVIM18]. arowana [KVC^{+}16]. array [LFH^{+}11]. array-based [LFH^{+}11]. artedi [WKW^{+}11]. Artemia [TNK12, WBF^{+}10]. arteries [BOH^{+}12, DRdS^{+}10]. Arteriosclerosis [DKG^{+}17]. arteriosus [DKG^{+}17, YP17]. arthromitus [DPTFC10]. Arthrospira [RBH^{+}17]. artificially [BLFRCQ10]. arugamensis [KCL14]. AsaP1 [SWM^{+}15]. AsaP1-toxoid [SWM^{+}15]. Ascanius [BT14, RSP^{+}17, RSM^{+}18, TB12]. ascidian [KOS^{+}10, PZL^{+}14]. ascidians [KOS^{+}10]. Ascomycota [ŘKH16]. Asia [DMK^{+}19, SDM^{+}19]. Asian [CCW^{+}15, CCW^{+}15, GKCC^{+}12, KNN19]. ASK [BAMYAH19, CCCTA^{+}14, SDR^{+}15]. asp [MAOS^{+}18]. aspartate [ZCJ^{+}18]. aspects [GJVT^{+}11, PMBM13, PMP^{+}18, SVY^{+}14]. aspera [CR12]. aspiration [ALG15]. Aspius [MAOS^{+}18]. assay [AS17, ASS19, BWZ^{+}17, CDF^{+}18, CCCTA^{+}14, DPMK12, DSLR16, FSO^{+}10, GM10, GDN^{+}16, JKS^{+}13, KBJ^{+}13, LFH^{+}11, LLWL15, LXX^{+}17, MMBS10, MSI^{+}15, NMM^{+}18, SMF^{+}10, SFL^{+}16, YWW^{+}19, YLX^{+}19a, ZFZ^{+}11, ZLN^{+}17, vBVLR^{+}16]. assays [CHAB10, DCL^{+}17, EAM^{+}13, GJK^{+}11, LPCC12, NMIY14, SK15]. assembling [LRD^{+}19]. assembly [FFLC17, KREM14, PLCC11]. assess [BFC^{+}15, JKH^{+}17, SMM^{+}13]. assessed [GDN^{+}16]. Assessing [GMC^{+}16b, YLL^{+}18]. Assessment [DBS12, MBS^{+}19a, AAV^{+}16, DCG18, MMH^{+}18, NJK^{+}17, PNBK13, RCD^{+}17, SCF^{+}13, SGB^{+}16]. assessments [WWE^{+}13]. assignment [LWWS14, LGD^{+}17]. assisted [CMEV17a, CMEV17b]. Asso [EKS^{+}13, TAQ^{+}10, VPP^{+}12]. associated [ACLP^{+}19, AHB^{+}17, BKG^{+}18, BB11, BA13, BK12, BMA13, BVT^{+}18, BVL^{+}18, CBC^{+}13, CDM^{+}16, CKTT10, CSA^{+}15, CG11, CMR^{+}16, DSMC17, DML^{+}16, DYML^{+}18, DT16, FCL^{+}19, GDW^{+}15, HBB14, HSA17, HBB^{+}15, HSHC^{+}16, JYZ^{+}14, KCD^{+}14, KOB^{+}17, KRD^{+}17, LGK^{+}14, LFP^{+}11a, LFP^{+}11b, LLZW14, LXL^{+}18, LDEH17, LFL19, LLL^{+}13, MLMR16, MLER^{+}18, MRSC13, MSR^{+}10, MKH^{+}11, MTLP^{+}18b, MCO11, MMW^{+}15, OAO^{+}12, ONGQ^{+}14, ONA10, PAM^{+}19, PCK^{+}15, QXW14, RCGGM15, RKPG16, RSPA^{+}18, RTM^{+}15, STF^{+}15, SBM^{+}11, SDZN19, SWE^{+}19, SPV^{+}18, SKGE13, SRT^{+}14, TGJCH15, VJWC^{+}18, VFB^{+}12, VME^{+}19, WNLF^{+}12, XQS^{+}15, XLZ^{+}17, YYF^{+}10, ZZC^{+}17, ZJDL18, ZSZF16, ZxYJ^{+}19]. Association [CPB^{+}15, MTG^{+}10, CHF^{+}12, FCC^{+}13, FFLC17, MCO11, SL15]. associations [KNN19]. astaci [SMKBP17, VDHV^{+}16]. Astacus [OSC^{+}14]. Astronotus [RHSM^{+}18a]. asymmetric [YGL^{+}18]. asymptomatic [LJGRI^{+}10, WIC^{+}19]. ater [SDZN19]. atherinoides [HAP^{+}10, LRB^{+}17, MGN^{+}16]. atherinoidi [LRB^{+}17]. atherosclerosis [DBHC17]. Atlantic [JCN^{+}14, VCS^{+}19, ACN12, ACA^{+}16, AAV^{+}16, AFA^{+}12, ASSP^{+}15, Ano10c, ACM^{+}14, BSWS11, BRL^{+}13, BHO^{+}19, BSSH17, BOH^{+}12, BA13, BBP18, BMT^{+}15, BTC^{+}17, BHA^{+}17, BNS13, BDD^{+}18, BSBH10, CBLN13, CJB^{+}15, CNN^{+}10, CGR17, CHWM^{+}19, CZB^{+}11, CRAG^{+}12, DTF^{+}16, DKG^{+}17, DLLB^{+}18, DMB^{+}19, DRG^{+}17, DYML^{+}18, ESBE18, ECP^{+}10, EWRH13, ELI^{+}17, EDV^{+}18, ESV^{+}19b, FMM^{+}19, FvdMF^{+}18, FSM^{+}16, FFP^{+}14, FHFF15, FMH^{+}15, FGR^{+}12, FSV^{+}18, FSB^{+}18c, GL18, GFO^{+}13, GFS^{+}18, GMC^{+}16a, GVB^{+}17b, GIW^{+}11, GTT^{+}17, GCM^{+}18b, GSS^{+}10, GGH^{+}10, GCMG^{+}11, GMC^{+}16b, GFW^{+}10, GBSF12, GFR^{+}12, GRT12, GHB^{+}11, GCGR13, GGGM14, HHO14, HLHS11, HKD^{+}10, HSA17, HST^{+}13, IAR^{+}10, JCN^{+}13, JKO^{+}13, JWO^{+}10, JTW^{+}10, JMF^{+}18, JBO^{+}15, JOF^{+}19, JNS^{+}19, JR19, JHGR13, JJSJ15, KOB^{+}17, KAF^{+}19, KHG^{+}17, KSA^{+}10, KRFT14, KBD13, KWD^{+}16, LGK^{+}14, LDÁM14]. Atlantic [LAN^{+}14, LMG^{+}19, LLLG18, LDA^{+}19, LRAP17, MVM^{+}18, MKI^{+}15, MIO^{+}18, MMWK19, MSMM19, MSG^{+}18, MLMR16, MLER^{+}18, MSBB^{+}16, MLF^{+}10, MMD16, MMD^{+}12, MRSC13, MAC^{+}15, MID^{+}19b, MID^{+}19a, MSR^{+}10, MGH^{+}14, MMH10, NCNB18, NB10, ONGQ^{+}14, ONA10, PSZ^{+}14, PdLC19, PBMB11, PHVM13, PPK^{+}13, PPI^{+}13, PG16, PÅD18, PJU^{+}13, RB15, RGMGG^{+}18, RMR14, RBR19, RRG^{+}19, RSIP^{+}17, RSPA^{+}18, RBWR^{+}15, SMR10, SMM^{+}13, SFR^{+}17, SAS^{+}19, SFU^{+}13, SG12, SQC^{+}14, SAD^{+}12, STES14, SWWN^{+}16, SWS^{+}16, SCZM19, SSA13, SSL18, SNK^{+}19, TWNB^{+}12, TJB^{+}15, TCC10a, TB12, TES11, TCM^{+}19, UMMS17, USM17, UKS^{+}18, VSDS^{+}19a, VDMJ^{+}13, VMBGE17, WDM^{+}15, WMM^{+}16, WHR^{+}18, WNLF^{+}12, WNSAL12, WNAF^{+}13, WNMK^{+}16, WNAJ^{+}16, WNGS^{+}17, WOD16, WNO^{+}18, YP17, ZCCN13, ZCC^{+}15, ZNHC11a, ZNHC11b, ZCP12]. atlanticus [BCW^{+}15]. ATP [PLCC11]. atresia [CZB^{+}11]. attachment [PMBM13, TGJCH15]. Attempted [GM11]. Attenuated [CSSWK10, CGW^{+}19, GPL^{+}12, LLX^{+}19, MBS^{+}19a, ZXW^{+}12]. attenuation [MSG^{+}18]. attributes [BFC^{+}15]. Atypical [HKK^{+}11a, BVK^{+}16, DKBS17, GLW^{+}18, GGH^{+}10, MLWC11, RVC18, VDD^{+}18, YFA^{+}17]. audit [EWRH13]. aurata [BC11, CVLJ^{+}10, CVGR^{+}13, CCME16, DdRTM13, DPCdR14, LGS^{+}14, MSS^{+}17b, RHVG^{+}19b, TPD^{+}19]. auratus [KKOY19, KJW^{+}17, LJHG14, NSS^{+}17, WXL16, ZWLG10, BBL^{+}16, GAEK^{+}14, LCR^{+}10, RMAM12, WIC^{+}19, YNL11, ZWLG10]. aureamaculosa [HSGD11]. aureus [OAO^{+}12, OGI^{+}18, SL15]. auriga [LJGRI^{+}10]. Austin [ABF15]. Australia [BFFA^{+}12, BDB^{+}18, HSGD11, MSS^{+}17a, NB11, OLC11]. Australian [CSA^{+}15, HO13, JFWH19, KMGK^{+}18, MMG^{+}11, MCO11, NSC^{+}15, RWTB17, SFF^{+}17]. australis [CSA^{+}15, Dig13, Dig15, Pol15]. Austria [LUP^{+}18]. autogenous [FLC^{+}17, RPMRLJ^{+}19]. Autophagy [BCLM^{+}19]. Autophagy-related [BCLM^{+}19]. Avenant [DAO19]. Avenant-Oldewage [DAO19]. Avermectins [OMB^{+}19]. avidus [LPC^{+}18]. avirulent [FNISE12]. avium [JK12]. axelrodi [LLG^{+}15, ŘKH17]. axillary [RVB13]. ayu [CONN10, FNISE12, KSS^{+}15, ZCC18]. Azumiobodo [PZL^{+}14].

B [JCN^{+}14, BSD^{+}18d]. B. [OSAA18]. Bacescu [FŠK^{+}19]. Bacillus [OSAA18]. Back [BWP^{+}19a, RHTC19a, PBBG19]. bacteria [DSSB18, EEMCEGSP16, PSBG^{+}11, RHBMM13, STIG17, THW^{+}12, TDLT^{+}16, WZQ^{+}16, YkSKY13, ZJDL18]. Bacterial [LST^{+}12, RGD^{+}12, SWH^{+}14, BSSH17, BLFC12, BSBH10, CHV^{+}12, DSA^{+}12, DAB^{+}12, GCNB16, GPFGAG19, GDW^{+}15, JAR^{+}11, JSAH^{+}18, KOB^{+}17, LFH^{+}11, LFL19, MFS^{+}17, MDC^{+}18, NFA15, PBC15, PHW^{+}16, QWTL17, RSH11, RHB^{+}15, VCN^{+}10, WMM17, ZZA^{+}14]. bacterially [SSV^{+}10, TMM^{+}15]. bacterin [PVIM18, SLKE10, WGL17]. Bacteriological [MCSM17, AdOG^{+}17, EAM^{+}13]. Bacteriophage [SPCR^{+}19, EDP^{+}17, HY16, LLL^{+}18a]. Bacteriophage-based [SPCR^{+}19]. Bacteriovorax [CHH^{+}14]. bacterium [CSS^{+}16, CHH^{+}14, LDEH17, SPV^{+}18]. baculovirus [RDN^{+}12]. Balantiocheilus [RTM^{+}15]. ballan [BT14, RSM^{+}18, SRM^{+}18, BVK^{+}16, RSP^{+}17]. Baltic [AHB^{+}17, BMGS17, BK12, GJVT^{+}11, OBD10]. Bangweulu [HHC^{+}18]. barb [SMD^{+}12]. barbel [ZPM^{+}18]. Barbonymus [ARR^{+}18]. Barbus [ZPM^{+}18]. barcoo [KRD^{+}17, LLH^{+}14b, KRD^{+}17, LLH^{+}14b]. barfin [KKK^{+}18]. barotrauma [HBB14, HFKF17, PSK^{+}12]. barracuda [GMP^{+}11, GMP^{+}11]. barramundi [BO15, JHW17, TGJCH15]. barrier [BSSB18, JSAH^{+}18]. base [MMNL14]. based [AS17, ÇOÇ^{+}16, DPMK12, DIVC16, ELI^{+}17, FSM^{+}16, GCGR13, JKS^{+}13, LFH^{+}11, MPBM^{+}16, PSK^{+}13, RB15, SPCR^{+}19, TCFáS19, VLVS^{+}16, YLX^{+}19a, ZWG^{+}18, ZCJ^{+}18, BMBS12, GCNB18]. baseline [GBV^{+}16]. Basilewsky [TIK^{+}14a, TIK^{+}14b, TQW^{+}14]. Basin [DAO19, SBW^{+}18, FDL^{+}17, MKKM15]. Bass [CGAS19, BCBS^{+}13, BVG^{+}10, CKTT10, DBS12, FPK^{+}17, JFWH19, LMC^{+}16, NOST^{+}16, PBC15, PST^{+}16, PBBG19, RGD^{+}12, SBW^{+}18, SBM^{+}11, SDJH16, TMC^{+}12, TNK12, VTM^{+}14, ZLX^{+}19]. batfish [LZGZ16]. bath [BCT^{+}17, FHD13, GVB^{+}17b, GVB^{+}17a, RHS^{+}12, SCR^{+}15, SOA^{+}18, TNK12]. bath-challenged [SOA^{+}18]. batrachus [BNC^{+}11]. Baumann [SNM^{+}16]. Bay [PBBG19, NRF^{+}18, SKB10]. Bayesian [LMG^{+}19]. be [DZV10, MBB12, PMH^{+}11]. bean [MTLP^{+}18b, CHF^{+}12, GM11]. behaviour [AHAMR^{+}19, BDD^{+}18, DLLB^{+}18, FHZ^{+}16, HGH^{+}18, OWM^{+}17, SSK17]. Behavioural [FSB18a]. being [PSZ^{+}14]. Belar [DT16]. bellied [EDSL11, LYD^{+}15]. Bench [EAM^{+}13]. Bench-top [EAM^{+}13]. benefits [GCGR13]. Bengal [PRWG17]. benign [VPCM19]. Bennett [WA14]. benzisoxazole [SRD^{+}16]. benzoate [GSS^{+}10, HLHS11, IPW^{+}13, JHGR13, KWGS19, LML12, PPI^{+}13, SMR10, SMM^{+}13, Sam10, WWE^{+}13]. bergylta [BVK^{+}16, BT14, RSP^{+}17, RSM^{+}18, SRM^{+}18]. Bering [WMF14]. Bermuda [BGC^{+}17]. best [JMD^{+}18, MMH^{+}18]. bester [RHSM^{+}18b]. Beta [SJSA^{+}14, LSA^{+}15, NYRS12]. beta-2 [LSA^{+}15]. Beta-glucan [SJSA^{+}14]. Betanodavirus [KCD^{+}14, LNC^{+}17, MPJ^{+}15, PST^{+}16, SOA^{+}18, TDS^{+}19, Bin14, CGG^{+}10, DTM^{+}10, HGB^{+}11, HLP^{+}16, KTP^{+}15, KKK^{+}18, KKM^{+}19, NOST^{+}16, SODB16, SOGR^{+}19, VTM^{+}14, Wen16, WCL^{+}12, FCM^{+}19]. Betanodavirus-associated [KCD^{+}14]. betanodaviruses [BBC^{+}10, BCL^{+}19, VGP^{+}18]. Betta [RhAS^{+}15]. between [BLFRCQ10, FNISE12, JWO^{+}10, JLJ16, KHG^{+}17, LMC^{+}16, LZZN10, LLS^{+}11, MSC^{+}17, MMH10, OLT^{+}10, SL15, VC17, WKK12]. bias [CXK16, MLH^{+}15]. bidyana [FHD13]. Bidyanus [FHD13]. bigeye [HHM^{+}14]. binding [vGJHS^{+}11, YHY^{+}16, ZGY^{+}18, WCL^{+}19]. bioassay [HH13b, HBS^{+}14]. bioassays [HH13a, MMH^{+}18, SMM^{+}13]. Biochemical [SSV^{+}17, DPTCF10, TAC15]. biochemistry [BHA^{+}17, LWMT10]. Biocontrol [GPFGAG19, CHH^{+}14]. bioenergetics [BSV^{+}18]. biofilm [ASS19, CD17, DSSB18, DSSB19, LFD^{+}17, LITCAH19b, SW11]. biofilm-forming [CD17]. Biography [HR15, PBW14]. Biological [LQL^{+}10, MSS^{+}17a, SDI^{+}18]. biologically [RHB^{+}15]. Biology [Spe19a, LYKN18]. bioluminescence [BKD17]. biomarker [BMT^{+}15, FBK^{+}16]. biomarkers [BMGS17]. biomass [MMG^{+}11]. biomolecular [GMB^{+}17]. biophysical [SCF^{+}13]. biopolymer [BYAS^{+}19]. biopsy [CMS^{+}19, TMWV14]. biosecurity [NSC^{+}15, TJR^{+}18]. biotype [FLVA10, SON^{+}18]. biphasic [MLWC11]. bivalve [ACYC10, VGP^{+}18]. Bivalvulida [LJHG14, LRB^{+}17]. Bizerte [YB17]. BKD [BSSH17]. black [BCT^{+}17, BPF^{+}18, DCMM13, DBM^{+}14, GPWF19, HKK^{+}11a, LLDL13, PSK^{+}13, PSK^{+}12, SHW^{+}11, SHS14, SRSF10, SHAM14, SMS^{+}19, XWP^{+}18, ZxYJ^{+}19]. blackeye [HSGD11]. blacktip [PC17]. bladder [BMD^{+}14, BCG^{+}17, ŘKH18]. Bleeker [BRMG13, BCJ13, DCV^{+}14, LYD^{+}15, LLDL13, RTM^{+}15, WZL^{+}11]. Bloch [BFFA^{+}12, BO15, CLWJ13, CBH^{+}15, CCM^{+}11, DDBC^{+}15, GKCC^{+}12, GHW^{+}15, HSGD11, HHM^{+}14, JHW17, LJHG14, LLL^{+}13, TGJCH15, WXL16, YLW^{+}14, ZWLG10]. blochii [AZSI^{+}12, CWL^{+}17, WKW13]. Blood [BSSB18, CMS^{+}19, Lek12, AI11, DPTCF10, WXL16, YSV^{+}13]. blood-free [YSV^{+}13]. blooded [SAS^{+}19]. bloom [KPM^{+}11]. blotched [CDM^{+}16]. blotting [LCR^{+}10]. blue [BRL^{+}13, FSM^{+}16, HBH^{+}19, KEY^{+}15, RGMGG^{+}18, XKBC12]. bluefin [CZB^{+}11, PSZ^{+}14]. Bluegill [LEM^{+}19, BMA13, DWL^{+}17, LDEH17, LFL19, LEM^{+}19]. Body [PPK^{+}13, LL11, LGS^{+}14, SSL18]. Bonamia [BLSG^{+}16, BWP^{+}19b]. bonasus [BB11, VPCM19]. bone [ARL^{+}17]. Bonnaterre [JK12]. bonnethead [CBC^{+}13, MCWL13]. Book [Rob10]. Boone [CMEV17a, CMS^{+}17, CMEV17b, CLW^{+}13, HMV^{+}17, HXZ^{+}16, SSV^{+}17, STG^{+}10, SNM^{+}16, TMM^{+}15, TDLT^{+}16, TDT^{+}16]. border [NSC^{+}15]. botanically [OWM^{+}17]. both [CGW^{+}19, CBM^{+}18]. Boulenger [SHAM14]. bovis [LRC^{+}18]. Boxhall [HBS^{+}14]. Boxshall [GAVC16, MSMM19, MGM^{+}18]. Brachydanio [Bin14]. brachypomus [FCS^{+}18]. brachyurus [RBF14]. brackish [NOV^{+}11]. brain [BSSB18, CLL^{+}10, OKS^{+}13, SON^{+}18, WWZ^{+}18, WZW^{+}18, WKW13, Wen16, XJL^{+}19, YCNW18]. branchial [BSD^{+}18b]. branchialis [BSWS11]. Branchinella [SHW^{+}11]. Branchiomonas [WNGS^{+}17]. Branchiomycosis [PdSG^{+}12]. branchiophilum [GDW^{+}15]. Brandt [RFA^{+}14]. Bravo [GAVC16, HBS^{+}14, MSMM19, MGM^{+}18]. Brazil [CMSHS16, CdAPdC^{+}19, FCS^{+}18, MdOO16]. Brazilian [ARL^{+}17, CLS^{+}15, FLC^{+}17, MDC^{+}18, PdSG^{+}12, SPKS17]. breakdown [BSSB18]. bream [ACSB^{+}19, BC11, DJQ^{+}12, DdRTM13, DPCdR14, Dig13, Dig15, JKH^{+}17, JJV^{+}15, JLJ16, KPT^{+}11, KHC^{+}18, KLW10, LGS^{+}14, MSS^{+}17b, Pol15, RHVG^{+}19b, TPD^{+}19]. bred [ADR11, XLZ^{+}17]. Breeding [VST^{+}19, SSO^{+}14]. brevifilum [SFBW^{+}17]. brevirostris [JMHK18]. brevirostrum [LGK^{+}14]. brill [LPVGL^{+}10]. British [BPPKJ10, BHT^{+}17, MMB^{+}15, SMR10, SMM^{+}13, TVK^{+}19]. broadly [KWD^{+}16]. Bromodeoxyuridine [EPCSBP18]. bronze [RBF14]. brood [ZZQ^{+}19]. broodfish [WNSAL12]. broodstock [LMG^{+}19, LPCC12, MMH10, RDN^{+}12, SBP^{+}19]. brook [DGS^{+}14]. Broth [YVM^{+}14, GCW17]. brown [AGFG19, BSPS^{+}18, KHG^{+}17, KFK^{+}16, LSC^{+}14, LSA^{+}15, OSC^{+}14, PGF^{+}19, PHR14, SVV^{+}17, Wen16, YB17]. brown-marbled [LSC^{+}14, LSA^{+}15, Wen16]. Brunswick [BSSH17, EWRH13, GVB^{+}17b]. bryosalmonae [BSPS^{+}18, BPPKJ10, JZK^{+}14, KAESEM13, LUP^{+}18, MJ17, Mor12]. bubble [RPJ^{+}18]. Buchanan [Bin14, CXK16]. Buenos [Lek12]. bulbus [DKG^{+}17, YP17]. Bulgaria [OSC^{+}14]. bullhead [BPF^{+}18, PHR14]. burbot [PFJ^{+}10]. Burchell [DAO19, HGPD11, HOHG13, LvD18, OOO^{+}11, OAO^{+}12, SRT^{+}14, TMWV14]. burdens [GVB^{+}17b]. Burm.f [HDC18]. burst [CHAB10, UKS^{+}18]. butterfly [BCJ13]. buttnerae [OMB^{+}19]. bv [ZWL^{+}16]. by-products [LKS^{+}13, MMNL14].

C [JCN^{+}14, OTN^{+}12, OKS^{+}13, GMN12, HPAL10, Sam10, SG12]. C. [ZWLG10]. C6 [GCNB16, APG^{+}11, HO13]. C6-6 [GCNB16]. C6/ [APG^{+}11, HO13]. Ca [AFTEML19, WNGS^{+}17]. Ca/P [AFTEML19]. cactaceae [BAH^{+}16]. cage [AZSI^{+}12, GBV^{+}16, LHW^{+}15a, SPS^{+}17, SZZSFN^{+}14, TWYC15, XQS^{+}15, ZxYJ^{+}19]. cage-cultured [AZSI^{+}12, SZZSFN^{+}14, ZxYJ^{+}19]. cage-farmed [SPS^{+}17]. cage-level [GBV^{+}16]. cage-reared [LHW^{+}15a, XQS^{+}15]. cainii [ABF15]. calcarifer [BO15, GKCC^{+}12, JHW17, KNN19, SDM^{+}19, TGJCH15]. California [CGAS19, GCCLRP^{+}13, KHS11]. californiensis [VCLW16]. Caligidae [FSB18a, KJB11, NLC^{+}16]. Caligus [BNS13, FVAVMGE18, GAVC16, GOM^{+}19, HLHS11, HBS^{+}14, MSMM19, MGM^{+}18, STIG17, VMBGE17, YAHWU11]. Callinectes [FSM^{+}16]. campbellii [NDA^{+}19, WBF^{+}10]. Can [BMK17, DZV10, LPA^{+}10]. Canada [BSSH17, BPPKJ10, EWRH13, FSV^{+}18, FSB^{+}18c, GL18, GVB^{+}17b, KHLL16, LGK^{+}14, LLLG18, LDA^{+}19, MMB^{+}15, MMD16, NRF^{+}18, SMM^{+}13, SCD19, SGQ^{+}15, SQC^{+}14, TVK^{+}19]. Canadian [VC17]. canadum [CLL^{+}10, HCL^{+}14, HSHC^{+}16, PSS17, SPS^{+}17, TWYC15]. Cancer [SWH^{+}14]. candidate [LMC^{+}16, SCK13]. candidates [ABF15, BLFC12, JKC^{+}11]. Candidatus [DPTFC10, MSR^{+}10, SPV^{+}18]. Cantor [LL12]. CAP18 [CMH^{+}17]. capensis [MVWS13]. Capillariidae [MWRSK17]. capillata [PÅD18]. Capsalidae [HSHC^{+}16, TGJCH15]. capsid [CGG^{+}10, FMT^{+}14, SCK13, SOGR^{+}19]. capsomers [LRD^{+}19]. Capsular [ZKL^{+}19]. capsulated [CLT^{+}14]. captive [ADR11, BDB^{+}18, CHF^{+}12, HSGD11, JK12, LYD^{+}15, MCWL13, RTM^{+}15, SBM^{+}11]. captive-bred [ADR11]. captive-raised [SBM^{+}11]. captivity [HFKF17, MS16, RBWR^{+}15]. capture [WNLF^{+}12]. capturing [OLVJ^{+}13]. carapo [VMCM14]. Carassius [BBL^{+}16, FCC^{+}13, FSC^{+}16, GAEK^{+}14, KKOY19, KJW^{+}17, LJHG14, LCR^{+}10, LLL^{+}13, NSS^{+}17, RMAM12, WXL16, WIC^{+}19, YNL11, YLW^{+}14, ZWLG10]. Carcharhinus [PC17, RBF14]. Carcharias [LL11, HKF^{+}16, RBF14]. Carcharodon [HKF^{+}16, RBF14]. carcinoma [SHAM14, TIK^{+}14b, VPCM19]. cardiac [GRT12, YLW^{+}14]. cardinal [LLG^{+}15, ŘKH17]. Cardiomyopathy [GFS^{+}18, RMR14, SNK^{+}19, WNLF^{+}12]. cardiorespiratory [PG16]. Caribbean [PRWG17, SBK^{+}13]. Carp [ABVS18, MBS^{+}17, ARR^{+}18, ATJS^{+}18, AI11, BSK^{+}10, BMK17, CLC^{+}17, CR12, CCT^{+}11, EBM^{+}18, FBS13, FSV^{+}14, GSEM13, GSBEM15, GFS^{+}14, GCWB11, JDRM17, JAR^{+}11, JSAH^{+}18, KKOY19, KJW^{+}17, KPM^{+}11, LWJ^{+}16, LLL^{+}12, LLZW14, LJHG14, LXHL11, LLL^{+}13, MPBM^{+}16, MSB^{+}19, MSS^{+}17a, MGN^{+}16, MTA^{+}15, MTAB15, OSAA18, OCLI^{+}19, PTM^{+}19, PFP^{+}13, SGB^{+}16, SBD11, SKN^{+}16, SEM18, SRB12, SJSA^{+}14, TIK^{+}14b, VCN^{+}10, WXL16, WCR^{+}18, XWP^{+}18, YFA^{+}17, YLW^{+}14, YkSKY13, YYW^{+}14, ZWG^{+}18, ZWLG10, ZYW^{+}10]. carpet [CIA^{+}18, DABO^{+}17]. carpio [AI11, BLFRCQ10, BDM^{+}18, CLC^{+}17, CR12, CCT^{+}11, CLHGK17, FSV^{+}14, GJM^{+}15, GSEM13, HDC18, JDRM17, JAR^{+}11, JSAH^{+}18, KPM^{+}11, LWJ^{+}16, LLZW14, MBS^{+}17, MSS^{+}17a, MTA^{+}15, MTAB15, OSAA18, OCLI^{+}19, PFP^{+}13, RHZA^{+}16, SBD11, SRB12, SJSA^{+}14, VCN^{+}10, WPH^{+}13, WZW^{+}18, YFA^{+}17]. Carrier [FLCC^{+}14, PFJ^{+}10, SBH^{+}16]. Cartilage [YP17]. Caryophyllidea [SBD11]. Case [BSSH17, CLHGK17, LGK^{+}14, AHJS^{+}18, ALG15, ATP^{+}16, ARFG13, BMK17, BO15, CIA^{+}18, DBHC17, DML^{+}16, FKB^{+}14, JDRM17, JR19, LL12, MNF^{+}15, OLVJ^{+}13, OOO^{+}11, TIK^{+}14a]. casein [PMH^{+}11, SJDC^{+}13]. cases [DSJA18]. caspase [LBF^{+}14, RCGGM15]. caspase-1 [RCGGM15]. caspase-3-positive [LBF^{+}14]. cataract [ALG15, SFR^{+}17]. cataracts [JHRW17]. Catarina [CdAPdC^{+}19]. catchment [DIH^{+}16, HHC^{+}18]. catecholamine [DSSB19]. catfish [BNC^{+}11, BSV^{+}18, BSSB18, BSD^{+}18b, BSD^{+}18a, BSD^{+}18d, BAD^{+}12, BBF^{+}15, BPF^{+}18, CCMA^{+}17, CTK^{+}10, DD15, DAO19, EAN^{+}18, HGPD11, HOHG13, KMGK^{+}18, LTC^{+}14, LLWL15, MP18, MAOS^{+}18, OOO^{+}11, OAO^{+}12, PRWG17, PRSK10, PMK12, RKPG16, SMV^{+}14, SLPB18, SZZSFN^{+}14, SMFM12, SRT^{+}14, TMWV14, VST^{+}19, XKBC12, YSK13, YK13]. Catostomus [MGN^{+}16]. caudal [ASW16, PWHW18b]. caught [CNN^{+}10, HFKF17, MIO^{+}18, RBWR^{+}15, RSM^{+}18, SDZN19]. causal [ACM^{+}14, SSF^{+}17]. causative [CPP14, KEY^{+}15, MFS^{+}17, RGD^{+}13, WEPB^{+}18, ZWG^{+}17]. cause [HB12, LCZ^{+}18]. caused [APP^{+}17, AHTCA^{+}17, CONN10, DDBC^{+}15, DTJ^{+}17, EEMCEGSP16, GVC^{+}19, KAF^{+}19, LLH^{+}14a, MKY^{+}15, MGM^{+}16, MAOS^{+}18, MCCN14, NBJ10, NDA10, PHVM13, ŘKH18, RSP^{+}17, RRG^{+}19, TSK^{+}19, ZZA^{+}14, dSdAdS^{+}19]. causes [LDÁM14, PZL^{+}14, QZZ^{+}18b, SSK17]. causing [BMMC^{+}15, BPF^{+}18, DD18, KYB^{+}15, LPC^{+}18, PRY^{+}18, QWTL17, SPP^{+}18, TOK^{+}16, WZQ^{+}16, ZWL^{+}16]. caution [KZJ^{+}17]. Cavernous [SMD^{+}12]. caviae [BSV^{+}18, BSB^{+}18]. CD3 [RBL^{+}11]. cefotaxime [GTGCM19]. cel [XXL^{+}19]. cel-EIIB [XXL^{+}19]. Cell [DLP^{+}13, AI11, AGB^{+}19, APG^{+}11, BNC^{+}11, BAMYAH19, BKG^{+}18, CHWM^{+}19, CLL^{+}10, CD17, CGCAH12, DMD^{+}17, EYF^{+}19, GHH^{+}11, GM10, GSBEM15, GBSF12, GHW^{+}15, HGW^{+}19, HO13, HCW^{+}16, IFF^{+}15, JMF^{+}18, JJS^{+}12, KLW10, LJC^{+}19, LSC^{+}14, MVM^{+}18, MB17, MTHB13, NMV^{+}19, NSE^{+}17, NYRS12, PWHW18b, PWW19, PMW^{+}10, PPF^{+}12, RPMRLJ^{+}19, RBL^{+}11, SVL^{+}17, SHAM14, SSG^{+}19, SDR^{+}15, TQW^{+}14, TDT^{+}16, VBL^{+}15, VBL^{+}16, VKK^{+}19, WBF^{+}10, WWS^{+}11, WWZ^{+}18, WZW^{+}18, WCL^{+}19, WKW13, Wen16, WCW^{+}16b, WBD^{+}10b, XWP^{+}18, YNL11, YCNW18]. cell-free [MB17]. cell-infecting [WCW^{+}16b]. cells [AHPR^{+}16, AJLC^{+}16, BCLM^{+}19, BGD^{+}17, BDM^{+}18, CCCTA^{+}14, CXW^{+}19, DPC^{+}10, DPADO16, GIW^{+}11, GCM^{+}18b, IFF^{+}15, LDÁM14, Lek12, LASRA^{+}19, LBF^{+}14, OMO^{+}15, PMH^{+}11, PSR^{+}11, PPK^{+}13, RSB13, RCGGM15, SHKF16, SJT14, WXL16, WBD^{+}10b, XJL^{+}19]. cells-infecting [OMO^{+}15]. Cellular [ASA10, CRS^{+}13, YHY^{+}16]. Celtic [RBWR^{+}15]. central [SMJ^{+}14, GFB^{+}13]. centres [PSZ^{+}14]. centrifugation [DLCC^{+}13, MTHB13]. centroura [CBFM16]. Cephalotheca [ŘKH16]. cephalus [MVWS13]. Cerastoderma [RAVC18]. Ceratomyxa [HB12]. cerebralis [FCJH19, MKKM15]. cernuus [MGM^{+}16]. Cestoda [BB11, BCJ13, SBD11]. cestode [WFH^{+}18]. ceti [CLS^{+}15, GMC^{+}16b, MHNY18]. CEV [MSB^{+}19, RAM^{+}18]. cf [MdOO16]. chain [BHT^{+}17, LSA^{+}15, NYRS12, TSS18, THW^{+}12, YEK^{+}17]. chalcogramma [WMF14]. Challenge [BVK^{+}16, CCLGRM^{+}17, CAJD^{+}18, CRS^{+}13, CTK^{+}10, GMC^{+}16b, GFM^{+}11, GBSF12, JAR^{+}11, JSAH^{+}18, KBTMH^{+}16, MRN^{+}19, MAC^{+}15, MID^{+}19b, MID^{+}19a, NSE^{+}17, PYSS19, SGB^{+}16, SPCR^{+}19, SSC17, TTL^{+}16, ZWL^{+}14]. challenged [LF16, NNCLGC^{+}11, RCGGM15, RSIP^{+}17, RSPA^{+}18, SOA^{+}18, SCS^{+}12, TSS18, ZJR^{+}18]. Change [OKS^{+}13, TTL^{+}16]. Changes [CHWM^{+}19, DCZ^{+}17, GTGCM19, VCN^{+}10, YK13, ACLP^{+}19, BCBS^{+}13, BHO^{+}19, CSA^{+}15, DTF^{+}16, DPTCF10, DAO19, FvdMF^{+}18, GFM^{+}16, JKL^{+}13, LDÁM14, LF16, MAOS^{+}18, PAM^{+}19, PCK^{+}15, SSV^{+}17, SOGR^{+}19, TQW^{+}14]. changing [JR19]. Channa [CLZ^{+}12, DIH^{+}16, LGC^{+}19]. channel [BBF^{+}15, LLWL15, MP18, PRSK10, PMK12, RKPG16, SZZSFN^{+}14, SMFM12, YSK13, YK13]. Chanodichthys [TIK^{+}14b]. chaperone [MFHHG18]. chaperones [RAS^{+}10]. Chapin [STF^{+}15]. char [SWM^{+}15]. Characidae [Lek12]. characteristics [CD17, HKK^{+}11b, LQL^{+}10, MB17, OLVJ^{+}13, TWYC15, WZA10]. Characterization [ACM^{+}14, CMSHS16, CD17, CCME16, CLS^{+}15, EDP^{+}17, GLW^{+}18, HBVB16, KYB^{+}15, LLL^{+}18a, PFJ^{+}10, PRY^{+}18, RB15, VKK^{+}19, Wen16, WNLF^{+}12, YCNW18, BNC^{+}11, BPF^{+}18, BCL^{+}19, BCM^{+}17, BDM^{+}18, CLZ^{+}12, CCLL16, CLL^{+}10, DTF^{+}16, DLCF13, DPADO16, DLPR15, DKBS17, DMK^{+}19, ERÁpSB14, FDL^{+}17, FMA^{+}19, GJM^{+}15, GBO^{+}18, GHH^{+}11, GFS^{+}14, GGGM14, HJT^{+}16, HY16, HLC^{+}15, HHX^{+}15, HCW^{+}16, IGLG^{+}17, KTP^{+}15, KLW10, LJC^{+}19, LCC^{+}15, LRB^{+}17, MFHHG18, NYRS12, NSC^{+}15, OLVJ^{+}13, OFA11, OGI^{+}18, PWHW18b, PKPG15, PSBG^{+}11, PAZ18, PLC^{+}19b, RGMGG^{+}18, RVC18, SRSF10, SDWA18, SDZN19, SPS^{+}17, SMV^{+}14, SMS^{+}19, SSO^{+}14, SMC^{+}11, TPD^{+}19, VMBGIC19, VGP^{+}18, WWS^{+}11, WWZ^{+}18, WZW^{+}18, WCL^{+}19, YNL11, YK12, YMH^{+}11, YLS^{+}19, ZQB^{+}11, ZB12, ZWG^{+}17]. characterize [BWM^{+}13, TAC15]. characterized [GTT^{+}17]. charr [SVL^{+}17, WKK12]. Chasmistes [JMHK18]. che [FMA^{+}19]. chelonae [CBW19]. Chemical [OBD10, SHS14]. chemicals [TFG19]. Chemotactic [YSK13]. chemotherapeutants [MMH^{+}18]. Chen [SS17]. Cherax [ABF15]. Chesapeake [SKB10]. Chile [AAUHW15, BNS13, CRAG^{+}12, GGH^{+}10, HBS^{+}14, MSMM19, PBMB11, VDMJ^{+}13, YAHWU11]. Chilean [AHTCA^{+}17, CIL^{+}17, CCCTA^{+}14, FCL^{+}19, HBB^{+}15, HKB^{+}16, IGLG^{+}17, ISCF^{+}19, LCV^{+}17]. Chilean-farmed [AHTCA^{+}17]. chilensis [IGLG^{+}17]. Chilodonella [GJMH17]. chilodonelliasis [BO15]. Chilodonellidae [GJMH17]. Chiloscyllium [CIA^{+}18]. Chimaera [DRdS^{+}10]. chimeric [EBM^{+}18]. China [LZGZ16, LHW^{+}15a, LXL^{+}18, CWW^{+}18, HHX^{+}15, LYL^{+}18, LQY11, LL12, LLH^{+}14b, LJHG14, LXX^{+}17, WZL^{+}11, XLZ^{+}17, YLSW14, ZYW^{+}10, ZSZF16, ZLX^{+}19, ZCC18, ZDWH18]. chinensis [WLX^{+}10]. Chinese [DML^{+}16, DCZ^{+}17, DCL^{+}17, FGJ^{+}10, SCS^{+}12, WLX^{+}10, YMH^{+}11, ZB12]. Chinook [BHT^{+}17, BMH^{+}19, DWP^{+}19, DPD^{+}18, GMC^{+}16a, GBO^{+}18, GTH11, LST^{+}12, LF16, LF18, LWMT10, MPS^{+}18, PMPG^{+}16, STF^{+}15]. chitosan [BYAS^{+}19, LLDL13, RSGB^{+}14]. Chlamydiales [SPV^{+}18]. chloramine [BA14]. chloramine-T [BA14]. chloride [CHWM^{+}19, MMD^{+}12, MGF17, RHS^{+}12]. chloride-induced [MGF17]. chlorine [DD18, MWRSK17]. chloropiperidin [SRD^{+}16]. Cholangioma [vD14]. cholerae [ZZC^{+}17]. cholinergic [BSD^{+}18b]. chondracanthid [ABC^{+}10]. Chondracanthus [ABC^{+}10]. Chondrichthyes [DRdS^{+}10]. chorion [SWS^{+}16]. chromatography [KKT^{+}18b, KKT^{+}18a]. chromogenes [LHW^{+}12]. chromogenic [CPEH15]. Chronic [KPT^{+}11, LLH^{+}14a, MKY^{+}15, SJT14]. Chryseobacterium [IAR^{+}10, PSVK^{+}18]. chrysops [DBS12]. chuatsi [TIK^{+}14a, TQW^{+}14]. chum [DPADO16, KBE^{+}14, MUM^{+}19]. cichlid [LFSZ18, RHST^{+}16, RhMM^{+}15, RhMV^{+}15]. cichlids [dASLS^{+}19]. ciliate [AJLK^{+}18, vGJKR^{+}17, YYW^{+}14]. ciliates [GJMH17]. Ciliophora [SPP^{+}18]. circle [ABM^{+}15]. citrus [LKS^{+}13]. CJ [ZWL^{+}14]. clam [ACYC10, DABO^{+}17, FBVC16, MMD16, PDdC^{+}14, RPJ^{+}18]. clams [MMMD16]. Clarias [BNC^{+}11, DAO19, HGPD11, HOHG13, LvD18, OOO^{+}11, OAO^{+}12, SRT^{+}14, TMWV14]. clarithromycin [CDW17]. clarkii [AMAR18, DXX^{+}15, JPG^{+}19, MKKM15, WXWD15]. class [ATM^{+}10, CPB^{+}15, LMG^{+}19, LLZW14, NB11, RBM^{+}16]. classes [AH16]. claudin [SJSA^{+}14]. Clavochlamydia [MSR^{+}10]. clay [BBF^{+}15]. Cleaner [Spe19a, Spe19b, BW18, GCGR13, GDN^{+}16, WDM^{+}15]. cleansing [Lek12]. clearance [CR12]. Cleidodiscus [BCT^{+}17]. Clinacanthus [HDC18]. Clinical [BHW^{+}11, CGP^{+}19, HBB14, JJB15, RMR14, SHW^{+}11, SMS^{+}10, BSSH17, BHA^{+}17, BSBH10, CPdPP16, CRAG^{+}12, HWM^{+}13, KTP^{+}15, QXW14, RHZA^{+}16, RHST^{+}16, RHSM^{+}18b, TSS18]. clinically [SG12, TWNB^{+}12]. Clinostomum [FŠK^{+}19]. clioquinol [LPC^{+}18]. clipping [ASSP^{+}15]. Clonal [MFS^{+}17, CHB^{+}17, TALA11]. clonality [RDW18]. clone [BMMC^{+}15]. Cloning [CLWJ13, JKH^{+}17, NMIY14, HCW^{+}16, NRMS14, NSE^{+}17, SHKF16]. close [LDEH17]. closure [ASSP^{+}15]. clownfish [BRMG13, SBK^{+}15]. clpV [LXZ^{+}19]. club [PPF^{+}12]. Clupea [GJVT^{+}11, GGW^{+}11, HLL^{+}12, HCEH16, MMBS10]. clusters [BKG^{+}18]. Cnidaria [MJ17, RKPG16]. Cnideria [PÅD18]. Co [ASHZ^{+}12, ATP^{+}16, ARFG13, MBOC19, VDMJ^{+}13, DYML^{+}18, LGJ19, WNAJ^{+}16]. Co-Infection [ARFG13, ATP^{+}16, MBOC19, VDMJ^{+}13, LGJ19]. co-infections [DYML^{+}18, WNAJ^{+}16]. coast [BMGS17, EEMCEGSP16, FSM^{+}16, GMP^{+}11, JKO^{+}13, LGK^{+}14, MDC^{+}18, RDN^{+}12, SDJH16]. coastal [CNN^{+}10, HGH^{+}16, RSF^{+}17]. coat [ZLN^{+}17]. coated [MTK^{+}10, SSHEM11]. cobia [CLL^{+}10, HCL^{+}14, HSHC^{+}16, PSS17, SPS^{+}17, TWYC15]. Cobitis [FŠK^{+}19]. cocaine [GFM^{+}16]. cocci [BHA^{+}11]. coccidiosis [LFL19, dSdAdS^{+}19]. cockle [DIVC16]. cockle-based [DIVC16]. cockles [DIVC16, RAVC18]. cocktail [SPCR^{+}19]. cod [BSWS11, BMGS17, BK12, FvdMF^{+}18, GIW^{+}11, GGGM14, HLHS11, HKD^{+}10, HFKF17, MVM^{+}18, MBB12, OBD10, PG16, RBWR^{+}15, Sam10, SJT14, ZCP12]. Codon [SCK13, MLH^{+}15]. Codon-optimized [SCK13]. coexistence [LVADB17]. coggygria [BE19]. cognate [ZWY^{+}11]. cohabitant [MID^{+}19a]. cohabitation [BWS^{+}18, GFM^{+}11, MID^{+}19b]. cohabited [Tre12]. Coho [AHTCA^{+}17, VMBGE17]. coinfection [ATP^{+}17, MP18]. coinfections [PTM^{+}19]. coioides [FHZ^{+}16, NNC^{+}18, WCL^{+}12]. Cold [KSS17, CHV^{+}12, SAS^{+}19, VBL^{+}16]. cold-blooded [SAS^{+}19]. Cold-water [KSS17]. coldwater [GCNB16, MBS^{+}19a, RHB^{+}15]. coli [SNM^{+}16]. collected [BOH^{+}12, BA13, BMGS17, DBV^{+}13, GOM^{+}19, MJ17, PPSW19, SMJ^{+}14, WWE^{+}13]. collecting [LST^{+}12, MWRSK17]. collection [ZCCN13]. colloidal [LXX^{+}17]. Colombia [BMMC^{+}15]. colonization [BK12, SFL^{+}16]. colonized [RHBMM13]. colonizing [LF18, SWWN^{+}16]. Colorimetric [ASS19]. Colossoma [ANB^{+}19, FCS^{+}18, OMB^{+}19, PdSG^{+}12, VDO^{+}18b, VDO^{+}18a]. Columbia [BPPKJ10, MMB^{+}15, SMM^{+}13, TVK^{+}19, BHT^{+}17, SMR10]. column [JHM^{+}18, MLF^{+}10]. columnare [BBF^{+}15, BLF^{+}16, BMA13, DBS12, DBV^{+}13, DLPR15, DSLR16, FDL^{+}17, GLW^{+}18, GCW17, LWWS14, LGD^{+}17, LGS19, LFD^{+}17, LZZN10, LLL^{+}12, OFA11, SL15, SLPB18, TQW^{+}14, WSA^{+}18]. columnaris [MA16]. columns [OM10]. combination [QZZ^{+}18b]. combinations [HCL^{+}14]. Combinatorial [PPI^{+}13]. Combined [MDF^{+}17, AGB^{+}19, TCFáS19, WSX^{+}18]. comes [BSW17]. Comment [KRFT14, JCN^{+}14]. Comments [Pol15, FYO^{+}11, MWS17, Dig15]. commercial [ABF15, CdAPdC^{+}19, MSBB^{+}16, TFG19]. commercially [LXX^{+}17, WGL17]. commersonii [MGN^{+}16]. Common [SBH^{+}16, ABVS18, BWS^{+}18, CLC^{+}17, CR12, GSEM13, JST^{+}19, JAR^{+}11, LWJ^{+}16, LLZW14, LJGRI^{+}10, MBS^{+}17, OSAA18, OCLI^{+}19, PTM^{+}19, PFP^{+}13, SGB^{+}16, SGQ^{+}15, SEM18, SSK17, SRB12, SJSA^{+}14, THW^{+}12, VCN^{+}10, VDD^{+}18, YFA^{+}17]. community [JSAB^{+}19]. companion [RBR19]. Comparative [BBP18, BFC^{+}15, FDST16, GPL^{+}12, IBFL19, LWZ^{+}18, MID^{+}19b, MNG^{+}18, OEG^{+}13, RCD^{+}17, RSIP^{+}17, VTM^{+}14, WZA10, XJL^{+}19, XKB12, YEK^{+}17, ABT^{+}14, BSPS^{+}18, DPTFC10, FLVA10, GFM^{+}11, MID^{+}19a, SOA^{+}18, SMS^{+}18, KPT^{+}11]. comparing [CMH^{+}17, EWRH13]. Comparison [GRG^{+}16, HOHG13, RSPA^{+}18, TAC15, BLF^{+}16, DZC^{+}16, EAM^{+}13, LLLG18]. Competition [YHY^{+}16]. complanatum [FŠK^{+}19]. Complementary [TQW^{+}14]. Complete [CGG^{+}10, KKM^{+}19, YMH^{+}11, ZxYJ^{+}19]. complex [CPB^{+}15, KDP^{+}17, TALA11, PdLC19]. Complex-related [PdLC19]. complexes [LC17, LRD^{+}19]. component [NDA10]. components [ASA10, BLF^{+}16]. composition [IBFL19, JBO^{+}15, KFK^{+}16, LDÁM14, LWZ^{+}18]. compositions [DTF^{+}16]. compounds [AHH16, AH16, MoSMS18, PPI^{+}13]. Comprehensive [HKB^{+}16]. compression [LGS^{+}14]. computed [LTC^{+}14]. computer [CMEV17a, CMEV17b]. computer-assisted [CMEV17a, CMEV17b]. concentration [GFM^{+}16, HH13a, MSC^{+}17, OM10, YVM^{+}14]. concentrations [FCJH19, GTGCM19, HHO14, HGH^{+}18, RHB^{+}15]. Concurrence [ACA^{+}16]. concurrent [LGK^{+}14]. condition [PBM^{+}15, SSJ^{+}19, SSL18]. conditions [CDG^{+}14, HBS^{+}14, HKG^{+}15, KCL14, MGM^{+}18, TAQ^{+}10]. conducting [WWE^{+}13]. confinement [PBC15]. confirm [KF10a]. confirmation [CHB^{+}17, GPWF19]. confirmed [JDRM17]. confirms [KDA^{+}18]. confocal [BCM^{+}17, CGCAH12]. conger [IGLG^{+}17]. Congo [HHC^{+}18]. conjoined [ADR11, BQP12, LTC^{+}14]. conjugated [RSGB^{+}14]. connection [MTG^{+}10]. consensus [TAC15]. consequence [AFTEML19]. consequences [LGS^{+}14]. conserved [YEK^{+}17]. consideration [SSL18]. Constance [BGRB14]. constant [TQW^{+}14]. constitutes [SPTDEE^{+}15]. Construction [EAN^{+}18, SWM^{+}15, YK12, ZWL^{+}14]. contact [SVY^{+}14]. containing [SNM^{+}16]. contaminated [BSD^{+}18d, WBD^{+}10b]. content [BCBS^{+}13]. continuous [MVM^{+}18, PWHW18b, SDR^{+}15, VKK^{+}19, YCNW18]. Contracaecum [DMB^{+}16]. contractile [LTC16]. contribute [FBS13]. contributes [YZW^{+}19]. control [ASA10, BNS13, DBS12, FLC^{+}17, GVB^{+}17b, HLHS11, JR19, MSS^{+}17a, NFA15, SKGE13, TDJ^{+}10, VGI^{+}16]. Controlled [SPLV^{+}15, KBTMH^{+}16, MMBS10, NRMS14]. controlling [BAH^{+}16, GCGR13, HGG^{+}11, MSMM19, SMFM12]. controls [BMBS12]. coordinated [JR19]. Copaifera [VDO^{+}18b]. Cope [JMHK18, LRB^{+}17]. copepod [ABC^{+}10, SAD^{+}12]. Copepoda [FSB18a, BSWS11, KJB11, MAOS^{+}18, PMBM13]. copepodid [BRL^{+}13]. copepodid-density [BRL^{+}13]. copepodids [BRL^{+}13, DLLB^{+}18, FSB18a]. copper [DBS12, FCJH19]. core [TMWV14]. Coregonus [SMS^{+}18, WKW^{+}11]. corkwing [SRM^{+}18]. Coronary [DTF^{+}16, BOH^{+}12, DRdS^{+}10]. correction [GM11]. correlate [PdLC19]. correlated [HZP^{+}12]. correlates [Dal18, NN11]. correlation [FPK^{+}17, JSJR^{+}19, LLS^{+}11, LWMT10, MSC^{+}17]. corresponds [GGW^{+}11].

Corrigendum [Ano12a, Ano12b, Ano17a, Ano17b]. cosmetic [VME^{+}19]. costae [KCD^{+}14]. Cotinus [BE19]. count [GBV^{+}16, GVB^{+}17a, Gis18]. counts [EWRH13]. County [MNM^{+}19]. coupled [YGL^{+}18]. course [RBM^{+}16]. Cover [BWP^{+}19a, ESV^{+}19a, FSB18b, GGMG18a, GCM^{+}18a, LITCAH19a, LLL^{+}18b, MBS^{+}19b, MTLP^{+}18a, PWHW18a, PLC^{+}19a, QZZ^{+}18a, RHTC19a, RHVG^{+}19a, VSDS^{+}19b]. cownose [BB11, VPCM19]. cowtail [SDZN19]. CpG [PFC^{+}13]. Cpn60 [CG11]. crab [DML^{+}16, DCZ^{+}17, DCL^{+}17, FGJ^{+}10, FSM^{+}16, LQY11, MRR^{+}14, RNM^{+}12, SCS^{+}12, ZB12]. crabs [SWH^{+}14]. crappie [GPWF19, SMS^{+}19]. crappies [BCT^{+}17]. Crassostrea [CBM^{+}18, GVLE16, GCCLRP^{+}13, HPY^{+}13]. crassus [BMD^{+}14, BND^{+}17, FDST16, LML12, LFP^{+}11a, LFP^{+}11b]. Crayfish [KPKP10, AMAR18, DXX^{+}15, JPG^{+}19, OSC^{+}14, SMKBP17, VDHV^{+}16, VDH19]. crimson [CLWJ13]. critical [WBD10a]. croaker [HZJ^{+}19, XQS^{+}15, ZZA^{+}14, ZWG^{+}17]. croakers [MPM16]. crocea [HZJ^{+}19, XQS^{+}15, ZZA^{+}14, ZWG^{+}17]. Cross [AAUHW15, GRF14, JLJ16, MBS^{+}19a]. Cross-neutralization [GRF14]. cross-protection [JLJ16, MBS^{+}19a]. Cross-sectional [AAUHW15]. crossback [KVC^{+}16]. crossbreeds [PFP^{+}13]. crowding [Bin14]. Crp [VSS17]. crucian [BSK^{+}10, KKOY19, KJW^{+}17, WXL16]. crumenophthalmus [HHM^{+}14]. Crustacea [BSWS11]. crustaceans [LXX^{+}17]. cryotherapy [BCW^{+}15]. Cryptocaryon [LMAKN16]. cryptocaryonosis [LMAKN16]. CT [BWM^{+}13]. Ctenochaetus [WA14]. Ctenolabrus [BT14]. Ctenopharyngodon [GCWB11, LLL^{+}12, YYW^{+}14]. Ctenopharyngon [LXHL11]. cucumber [LQL^{+}10]. cues [FSB18a]. cultivable [STIG17]. cultivated [VFB^{+}12]. cultivation [APG^{+}11, KAESEM13, KSAB^{+}14, Sla14]. cultivation-independent [Sla14]. culture [ABF15, AGB^{+}19, AdOG^{+}17, EAM^{+}13, GM10, GBSF12, GDN^{+}16, HO13, JJS^{+}12, LLX^{+}19, MB17, MLWC11, Mor12, ONA10, SMS^{+}10, STS^{+}15, YSV^{+}13]. culture-associated [ONA10]. cultured [AOAH13, AAZS^{+}15, AZSI^{+}12, ACLP^{+}19, BMMC^{+}15, BAD^{+}12, CVLJ^{+}10, CVGR^{+}13, CWW^{+}18, CCT^{+}11, CLW^{+}13, CSA^{+}15, CWL^{+}17, CPP14, CCM^{+}11, CCMA^{+}17, DMP10, EKS^{+}13, HHX^{+}15, KPT^{+}11, KPK^{+}18, KLD^{+}18, LGK^{+}14, LFSZ18, LYL^{+}18, LL12, LPVGL^{+}10, LLL^{+}13, MLRS^{+}17, ME12, MKMY10, PAZ18, PLC^{+}19b, QWTL17, RHSM^{+}18b, SMD^{+}12, SPKS17, SDJH16, SSJ^{+}19, SZZSFN^{+}14, SKPS16, TWYC15, WZL^{+}11, WZQ^{+}16, XJL^{+}15, YYF^{+}10, ZQB^{+}11, ZxYJ^{+}19, dALF^{+}17]. cultures [CHB^{+}17, PDdC^{+}14]. culturing [KDP^{+}17]. cuneata [LPVGL^{+}10]. curcumin [MDF^{+}17]. Current [GJMH17, LRMN17, MYWC18, WBD^{+}10b, GFS^{+}18, JJM^{+}17, KSS17, TALA11, MMH^{+}18, ZGY^{+}18]. curtus [HLHS11]. Cururu [MdOO16]. curve [TCFáS19]. cut [SEK^{+}16]. cut-off [SEK^{+}16]. Cutaneous [LLG^{+}15, MBD16, AGFG19, KHS11, LDÁM14, VPCM19]. cutoff [GCW17]. cutthroat [MKKM15]. Cuvier [BQP12, BHW^{+}11, CHF^{+}12, GOM^{+}19, PdSG^{+}12, SBK^{+}15, ZWY^{+}11].

Cyanea [PÅD18]. cyanobacterial [KPM^{+}11]. cyanuric [PCK^{+}15]. cycle [AJLK^{+}18, LNC^{+}17, RRC^{+}17]. cycles [JR19]. cyclic [VSS17]. cyclophilin [YSK13]. cyclopteri [ATP^{+}16, PB18]. Cyclopterus [ATP^{+}16, AGR^{+}16, CCH^{+}19, ESBE18, EWH^{+}19, ELI^{+}17, EDV^{+}18, ESV^{+}19b, GVC^{+}19, JHRW17, MLRS^{+}17, PAM^{+}19, PB18, SBP^{+}19, SFBW^{+}17, TDS^{+}19]. CyHV [BBL^{+}16, AHJS^{+}18, IFF^{+}15, NSS^{+}17, PFP^{+}13, RAM^{+}18]. CyHV-2 [BBL^{+}16, AHJS^{+}18, NSS^{+}17, RAM^{+}18]. CyHV-3 [IFF^{+}15, PFP^{+}13]. Cynoglossus [XJL^{+}15, ZQB^{+}11]. Cynoscion [CCM^{+}11]. cyprinacea [RHS^{+}12]. cyprini [GSBEM15, WBD^{+}10b]. Cyprinid [LLL^{+}13, MSS^{+}17a, ASB^{+}12, CLC^{+}17, CCW^{+}15, EBM^{+}18, MLH^{+}15, MKY^{+}15, MBT^{+}17, NSS^{+}17, PFP^{+}13, RAM^{+}18, SEM18, WXL16, WSX^{+}18, WIC^{+}19, BBL^{+}16, FBAS16, FCC^{+}13, FSC^{+}16, GSEM13, GAEK^{+}14, HDC18, IFF^{+}15, KJW^{+}17, TKS^{+}18]. Cyprinus [AI11, BLFRCQ10, BDM^{+}18, CLC^{+}17, CR12, CCT^{+}11, CLHGK17, FSV^{+}14, GJM^{+}15, GSEM13, HDC18, JDRM17, JAR^{+}11, JSAH^{+}18, KPM^{+}11, LWJ^{+}16, LLZW14, MBS^{+}17, MSS^{+}17a, MTA^{+}15, MTAB15, OSAA18, OCLI^{+}19, PFP^{+}13, RHZA^{+}16, SBD11, SRB12, SJSA^{+}14, VCN^{+}10, WPH^{+}13, WZW^{+}18, YFA^{+}17]. Cyrilia [MdOO16]. cyst [SdMVS^{+}16]. Cystic [RhMM^{+}15, WBE^{+}17]. cysticola [WNGS^{+}17]. Cystidicolidae [SBK^{+}13]. cytokine [AMA11]. cytokines [NSE^{+}17]. cytologic [RSR^{+}18]. cytometry [CHB^{+}17]. cytotoxicity [JSJR^{+}19]. Czech [PFP^{+}13].

D [ASHZ^{+}12, JCN^{+}14, KBE^{+}14]. dab [SBH^{+}16, VDD^{+}18, YYW^{+}14]. damage [MBB12]. damselae [BBP18, CDF^{+}18, CRS^{+}13, HCL^{+}14, KPSP^{+}19, OJRL10, SPS^{+}17, BBP18, HHM^{+}14, SPS^{+}17]. Danio [AMRHM18, CXK16, CKvGJB17, EDP^{+}17, KWGS19, LHW^{+}12, LHW^{+}15b, MLR^{+}12, MWRSK17, RAM^{+}19, SZM^{+}15, SSK17, ZJR^{+}18]. Dasyatis [CBFM16]. data [Gis18, HBS^{+}14, MPJ^{+}15, SMM^{+}13, SK15, SEK^{+}16]. database [EKT^{+}11, JSK^{+}10]. dead [CBM^{+}18]. deaminase [BSD^{+}18a]. Decapterus [BHW^{+}11]. decipiens [BK12]. decreased [MCO11]. decreases [CCHL13, USM17]. decussatus [DABO^{+}17]. defence [DLP^{+}13, RRO^{+}13]. deficiency [WPH^{+}13]. deficient [HCCM12]. definition [BSSH17, MNF^{+}15]. definitions [OLVJ^{+}13]. deformation [TPD^{+}19]. deformations [ASW16]. deformities [CLHGK17, Dig13, Dig15, FHFF15, GHB^{+}11, MLF^{+}10, Pol15, SMS^{+}18]. deformity [GCWB11, JHM^{+}18, NWMK16]. degenerated [VCOD17]. degenerative [ARL^{+}17]. Delftia [ARFG13]. delica [TFG19]. delivered [PBM^{+}15]. delivery [LLDL13]. delousing [HH13a, JR19]. Delta [BAD^{+}12]. Deltamethrin [HBS^{+}14]. Deltistes [JMHK18]. demonstrate [WKW^{+}11]. demonstrates [GCM^{+}18b, LLS^{+}11, PPE^{+}18]. demonstrating [JKS^{+}13]. demonstration [CBLN13]. Dendronereis [HVVV15]. denitrification [JSAB^{+}19]. density [BRL^{+}13, DLCC^{+}13]. densovirus [OLC11]. deoxynivalenol [RHB^{+}15]. dependency [JHW17]. dependent [ASSP^{+}15, CGG^{+}10, HZJ^{+}19, HLP^{+}16]. depletion [DdRTM13, DdRCP13, DPCdR14]. deposition [CHF^{+}12]. depositions [LAN^{+}14]. deprivation [LDÁM14]. derivative [PVIM18]. derived [GHH^{+}11, KLW10, LJC^{+}19, OWM^{+}17, PMW^{+}10, SSG^{+}19, WWS^{+}11, YSHG18, YNL11, YCNW18]. dermal [RKPG16]. dermatitis [PRT^{+}14, RBS^{+}16]. dermatopathy [KPT^{+}11, SJT14]. Dermocystidium [FCS^{+}18, KHLL16, LLG^{+}15]. Dermocystidium-like [KHLL16]. description [BLdA^{+}18, BVT^{+}18, CTD^{+}12, GKCC^{+}12, GGH^{+}10, MdOO16, SZM^{+}15, dSRSHGá^{+}19]. descriptive [WMM^{+}16]. Design [LTG^{+}18, VCOD17]. Desmarest [MPM16]. Desmozoon [MRSC13]. destructive [BMT^{+}15, CMS^{+}19, LJGRI^{+}10]. detect [CDG^{+}14, THW^{+}12]. detected [CCW^{+}15, SRM^{+}18]. detecting [AHITC18, BRB^{+}12, HWM^{+}13, WCW16a]. Detection [ACSB^{+}19, BKS^{+}17, BLSG^{+}16, BPF^{+}18, CPEH15, FCM^{+}19, FCC^{+}13, FGR^{+}12, GFB^{+}13, GBSF12, GCCLRP^{+}13, JKC^{+}11, JYZ^{+}14, JHGR13, KTN^{+}16, KJW^{+}17, LGC^{+}19, LJGRI^{+}10, MCW^{+}18, NJK^{+}17, NFF^{+}17, NSC^{+}15, OMO^{+}15, PBW^{+}18, SZZSFN^{+}14, SSO^{+}14, VGP^{+}18, WXL16, ARR^{+}18, AGB^{+}19, BBC^{+}10, BBL^{+}16, CLC^{+}17, CDF^{+}18, CWW^{+}18, CRAG^{+}12, CZL^{+}19, DPMK12, DTM^{+}10, DCL^{+}17, FSC^{+}16, FSG^{+}12, GMB^{+}17, GJK^{+}11, GM10, HJOK11, HGB^{+}11, JNN^{+}11, KJW^{+}12, KBJ^{+}13, KMGK^{+}18, LFH^{+}11, LLWL15, LXX^{+}17, LCR^{+}10, MSI^{+}15, MTHB13, MTAB15, MNM^{+}19, NLKA16, NMM^{+}18, NOST^{+}16, OCLI^{+}19, PHvdH12, PSK^{+}13, RHZA^{+}16, RPJ^{+}18, SMF^{+}10, SSHEM11, SPKS17, SKN^{+}16, Sla14, SEM18, TSS18, TSK^{+}19, WWZ^{+}18, WSX^{+}18, XJL^{+}15, YTL10, YWW^{+}19, YLX^{+}19a, ZCCN13, ZCC^{+}15, ZWG^{+}18, ZHFC11, ZLN^{+}17, vBVLR^{+}16]. detections [HST^{+}13, TVK^{+}19]. detects [ISCF^{+}19]. determinants [NB11]. Determination [OM10, DSJA18, RHTC19b]. determinations [YVM^{+}14]. determined [BHT^{+}17]. developed [NMV^{+}19]. Developing [WCW16a]. Development [Ano12c, BRB^{+}12, BMBS12, CDF^{+}18, CCCTA^{+}14, CLT^{+}14, ÇOÇ^{+}16, CPdPP16, DPMK12, DCL^{+}17, EKT^{+}11, FSG^{+}12, GJK^{+}11, HRL15, HLC^{+}15, JKS^{+}13, KJW^{+}12, KBJ^{+}13, LML^{+}13, LLWL15, LXX^{+}17, LLX^{+}19, MVM^{+}18, MSI^{+}15, MID^{+}19a, MF12, NLKA16, OFA11, PWHW18b, PSK^{+}13, RSGB^{+}14, SCF^{+}13, SVL^{+}17, SKN^{+}16, SFL^{+}16, TSS18, THW^{+}12, VBL^{+}15, YWW^{+}19, YLX^{+}19a, ZCC^{+}15, ZWG^{+}18, vBVLR^{+}16, BPCS19, BSK^{+}10, CRBLC^{+}17, CLS^{+}15, DMK^{+}19, EDN18, FLC^{+}17, HH13b, HFKF17, JWO^{+}10, JJV^{+}15, KBE^{+}14, LCOZ10, NMM^{+}18, OLVJ^{+}13, PWW19, SVV^{+}17, SNK^{+}19, dALF^{+}17]. DHA [BCBS^{+}13]. dhakensis [CMSHS16]. diagnose [NMIY14]. diagnosing [KKT^{+}18a]. diagnosis [ATM^{+}10, FBL^{+}14, FMB^{+}11, KKT^{+}18b, VCOD17, WXL16, YYF^{+}10]. Diagnostic [JSJR^{+}19, BTC^{+}17, CPdPP16, EAM^{+}13, FMT^{+}14, HH13b, HAS^{+}19, JWO^{+}10, JSJR^{+}18, LTG^{+}18, LMG^{+}19, LPCC12, MNG^{+}18, NJK^{+}17, VDH19, BWP^{+}19b, RHSM^{+}18b, RHSM^{+}18a]. diagnostics [LML^{+}13]. Dicauda [LRB^{+}17]. dicentrarchi [AHITC18, IGLG^{+}17, KAF^{+}19, LITCAH19b]. Dicentrarchus [CKTT10, FCM^{+}19, FPK^{+}17, LMC^{+}16, NOST^{+}16, PBC15, PST^{+}16, RRC^{+}17, RGD^{+}12, TMC^{+}12, TNK12, VTM^{+}14]. Dichelesthium [SAD^{+}12]. dichotomy [VC17]. diclina [SWWN^{+}16]. Dicologoglossa [LPVGL^{+}10]. did [OBFR12]. Didymodiclinus [PMP^{+}18]. Didymozoidae [PMP^{+}18]. diet [BSD^{+}18d, HOHG13]. Dietary [MLF^{+}10, AFTEML19, AMA11, BCBS^{+}13, DTF^{+}16, DMB^{+}19, LKS^{+}13, PCK^{+}15, SFR^{+}17]. differ [BDD^{+}18]. Difference [LZZN10]. Differences [JOF^{+}19, TDLT^{+}16, BLF^{+}16, RSPA^{+}18, SHO^{+}17, SFU^{+}13]. Different [ZXW^{+}12, BRL^{+}13, CRO^{+}18, CPdPP16, CRS^{+}13, DDBC^{+}15, DMK^{+}19, ELI^{+}17, GFS^{+}14, GRG^{+}16, HKK^{+}11b, HH13a, HZX^{+}13, JAR^{+}11, JR19, JPG^{+}19, JKH^{+}17, JHGR13, JBGP11, JJSJ15, JSJR^{+}18, JSJR^{+}19, KEY^{+}15, MAC^{+}15, PST^{+}16, RHTC19b, SSV^{+}10, SWWN^{+}16, TDS^{+}19, VTM^{+}14, WBF^{+}10, ZYW^{+}10]. Differential [CZL^{+}19, OLVJ^{+}13, PMW^{+}10, RSB13, VBL^{+}16, WCL^{+}12, ZJR^{+}18, CLW^{+}11, CLT^{+}14, CDG^{+}14, HB12, NSE^{+}17]. differentially [LSA^{+}15]. differentiated [CIA^{+}18, SJSA^{+}14]. differentiates [GLK^{+}16]. differentiation [JSJR^{+}18, RMSH09, SEM18]. differs [KBTMH^{+}16]. diffusion [SK15]. digenean [DLP^{+}13, DAO19]. digestion [RRG^{+}19]. digestive [BT14, DPTFC10]. Diggles [Dig15, Pol15]. dilation [LWMT10]. dimensional [KTK17, LLL^{+}12]. Dinophyceae [YB17]. dioctes [DD15]. Diodon [SBK^{+}13]. dioica [AA10]. dip [NCNB18]. Diphyllobothrium [KFK^{+}16, KKKA17]. Diplodus [KPT^{+}11]. diploid [AAV^{+}16, LWZ^{+}18, SFR^{+}17]. diplomonad [MLWC11]. dips [TJR^{+}18]. dipstick [WSX^{+}18]. Direct [DXX^{+}15, GCNB18, vGJHS^{+}11, LCR^{+}10, SSHEM11]. disc [SK15]. discus [AFTEML19, RKK^{+}19, RKK^{+}19]. Disease [AHTCA^{+}17, SWD^{+}14, ATJS^{+}18, ACN12, AAUHW15, AFTEML19, Ano10c, BSW17, BMMC^{+}15, BTN^{+}18, BHF13, BPCS19, BSK^{+}10, BSSH17, BPK^{+}17, BMT^{+}15, BSBH10, CCLGRM^{+}17, CVGR^{+}13, CJB^{+}15, CHV^{+}12, CPP14, CCME16, CRAG^{+}12, CONN10, DD18, DMB^{+}19, DBHC17, DML^{+}16, DTJ^{+}17, DYML^{+}18, DSJA18, DT16, FFLC17, GCNB16, GTT^{+}17, GCM^{+}18b, GKO^{+}14, GDW^{+}15, GMN12, GBSF12, HZP^{+}12, HHX^{+}15, JWO^{+}10, JTW^{+}10, JJB15, JJM^{+}17, KRD^{+}17, KTP^{+}15, KAF^{+}19, KOI^{+}14, KTN^{+}16, KYB^{+}15, LGK^{+}14, LKS^{+}13, LLS^{+}11, LRMN17, MBS^{+}19a, MNF^{+}15, MKI^{+}15, MLER^{+}18, MTG^{+}10, MR11, MA16, MCCN14, NPGA^{+}19, NBJ10, NDA10, OLVJ^{+}13, OSAA18, OCLI^{+}19, OLT^{+}10, PHvdH12, PRT^{+}14, PGF^{+}19, PRY^{+}18, QPB^{+}16, RSP^{+}17, RRG^{+}19, RPJ^{+}18, RHB^{+}15, SHW^{+}11, SHS14, SMS^{+}10, STS^{+}15, SDM^{+}19, SSF^{+}17, SQC^{+}14, SJK^{+}14, SVY^{+}14, SKJ13, TOK^{+}16, TWNB^{+}12, TCC10a, VFB^{+}12, WMM^{+}16]. disease [WNO^{+}18, XKB12, ZYW^{+}10, ZWG^{+}17, ZDWH18]. disease-causing [DD18]. diseased [AI11, BAD^{+}12, CMSHS16, CLZ^{+}12, GCMG^{+}11, KOS^{+}10, LPVGL^{+}10, OSC^{+}14, PKPG15, PAZ18, RKK^{+}19, SMV^{+}14, SG12, WNCR12]. Diseases [LYD^{+}15, Tre12, BMBS12, EEMCEGSP16, HAS^{+}19, JBO^{+}15, KKT^{+}18a, LTG^{+}18, LTM^{+}19, MKY^{+}15, NFA15, OLVJ^{+}13, Spe19b]. disinfectant [MoSMS18]. disinfectants [JST^{+}19, SHS14]. Disinfection [ITKH18, BMK17, DD18, MUM^{+}19, OBD10]. disorders [KOB^{+}17]. dispersal [SCF^{+}13, SDI^{+}18]. displaying [WNAJ^{+}16]. displays [BSB^{+}18]. disposal [DAB^{+}12]. Disruption [OWM^{+}17, ZCQF15]. Dissection [LNC^{+}17]. Disseminated [DSMC17, ŘKH17, SCZM19, DIVC16]. dissolved [MTLP^{+}18b]. distances [SMR16]. distinct [EDN18, IBFL19, LLC^{+}11]. distinguishing [FNISE12]. Distribution [AGFG19, BMH^{+}19, LUP^{+}18, MMMD16, AOAH13, CHWM^{+}19, DPC^{+}10, EKED^{+}17, GFR^{+}12, LBF^{+}14, MRR^{+}14, MJ17, NRF^{+}18, NHM^{+}10, PPE^{+}18, SHO^{+}17, SSV^{+}17, TDJ^{+}10]. disturbances [PHVM13]. divergence [FMA^{+}19, MBS^{+}17]. divergent [BVL^{+}18]. diverse [HGX^{+}19, LF18, ME12]. Diversity [CHV^{+}12, ASB^{+}12, BAD^{+}12, BFC^{+}15, DMP10, DLPR15, ISCF^{+}19, MPBM^{+}16, MWS17, PdLC19, PPSW19, TCM^{+}19, WMD^{+}18, WNMK^{+}16]. DlFBL [PBC15]. DNA [ASHZ^{+}12, CLGP15, CGR17, Dal18, DCMM13, EPCSBP18, GCCLRP^{+}13, HLL^{+}12, KKT^{+}18b, KKT^{+}18a, LLDL13, LFH^{+}11, RSGB^{+}14, TQW^{+}14, WCL^{+}19, YGL^{+}18, YEK^{+}17, ZWL^{+}14]. DNA-binding [WCL^{+}19]. DnaK [PHW^{+}16, SNM^{+}16]. Do [FBS13, NBJ10, PPF^{+}12, FCJH19, RHB^{+}15]. docking [SRD^{+}16]. docosahexaenoic [BCBS^{+}13]. doctor [VME^{+}19]. Does [HBH^{+}19, CRBLC^{+}17, BPCS19]. dollar [ALG15]. domain [ZGY^{+}18]. dopamine [DSSB19]. dorsal [KLW10]. dosage [dOAPLá^{+}17]. dose [CLGP15, HH13b, NSC^{+}18, PPI^{+}13, Sam10, WWE^{+}13]. dosing [RHTC19b]. dot [LCR^{+}10]. double [MKI^{+}15]. double-virus-infected [MKI^{+}15]. doubling [MLWC11]. Down [RBM^{+}16]. Down-regulation [RBM^{+}16]. Dr [Rob11]. Dr. [Cip11, EW17, Rob14a]. Drasche [KOS^{+}10, PZL^{+}14]. Dried [ZTF^{+}10]. drivers [JMD^{+}18]. drop [GKCC^{+}12, SDM^{+}19, ZDWH18]. drug [dSCBdS^{+}16, RKK^{+}19]. drugs [PZL^{+}14, ZN10]. dry [SG12]. dry-stored [SG12]. dsRNA [DPADO16, SSV^{+}10, TMM^{+}15]. dual [VLVS^{+}16]. dual-labelled [VLVS^{+}16]. due [ABM^{+}15, JJB15, ŘKH17]. Dujardin [KWGS19, MKK^{+}11]. Dumbo [DAO19]. dumerili [MKH^{+}11, dSRSHGá^{+}19]. duplex [FGR^{+}12]. duration [HLHS11]. during [BMT^{+}15, CHWM^{+}19, CG11, DMB^{+}19, DYML^{+}18, EDN18, FvdMF^{+}18, GTGCM19, HH13a, JKL^{+}13, KHG^{+}17, LGJ19, LXZ^{+}19, OJRL10, PTM^{+}19, RHS^{+}12, RCGGM15, SMS^{+}10, SMJ^{+}12, SHKF16, VMBGE17, XJL^{+}19]. dusky [PMP^{+}18, RBS^{+}16, RSF^{+}17]. dwarf [RWTB17]. dyed [Góm12]. Dynamics [KBTMH^{+}16, MSMM19, HHO14, HKG^{+}15, JPG^{+}19, SVY^{+}14, SSL18, YLL^{+}18]. dysfunction [LGJ19]. dysgalactiae [CLLF14, NLF11, NMIY14, TWYC15]. dysregulation [CCHL13].

E. [ZDWH18]. Early [WFH^{+}18, WOD16, BPK^{+}17, HGG^{+}11, ICVV16, vGJHS^{+}11, LF18, LSC^{+}14, MNG^{+}18, MTA^{+}15, MTAB15, MPS^{+}18, RRO^{+}13, SMJ^{+}12, SFU^{+}13, YK13, dALF^{+}17]. Early-stage [WOD16]. East [Dig13, Dig15, GMP^{+}11, RBR19, CCW^{+}15, LZGZ16, Pol15]. eastern [CBM^{+}17, CBM^{+}18, PdSG^{+}12, VPS^{+}14, WMF14, LDA^{+}19]. echinostomatid [MGM^{+}16]. Ecological [PMP^{+}18]. economic [KNN19]. ecosystem [JMD^{+}18]. ectoparasites [STIG17]. ectoparasitic [MUM^{+}19, TGJCH15]. Ectopic [BKG^{+}18]. ectopically [TPD^{+}19]. edema [ABVS18, MBS^{+}17, MSB^{+}19]. edible [DdRTM13, KOS^{+}10, PZL^{+}14, SWH^{+}14]. Editorial [RW10, RW11, Rob13, Rob14b, Spe15, Woo12]. edule [RAVC18]. edulis [BRL^{+}13, BLSG^{+}16]. Edward [MMMD16]. Edwards [CG11, GMP^{+}11, NRMS14, SQC^{+}14, ZB12]. Edwardsiella [ACAS14, ACLP^{+}19, BAD^{+}12, CDM^{+}16, CGAS19, CTD^{+}12, COB^{+}16, CTK^{+}10, DMK^{+}19, EAN^{+}18, EA18, GRG^{+}16, GPWF19, KTK17, KMGK^{+}18, KREM14, LKS^{+}13, LML^{+}13, PRWG17, PRSK10, PMK12, QXW14, RSR^{+}18, SMV^{+}14, SRT^{+}14, VST^{+}19, WZA10, YK13, ZXW^{+}12]. Edwardsiellosis [QXW14, LML^{+}13]. edwardsii [ZJDL18]. eel [ARFG13, BMD^{+}14, BND^{+}17, BCG^{+}17, GFM^{+}16, IGLG^{+}17, KYK^{+}12, KPK^{+}18, KLD^{+}18, MBD16, MFES18, MWS17, MSS^{+}12, NJK^{+}17, OMO^{+}15, PWHW18b, PWW19, SWE^{+}19, WCW^{+}16b, WTM14]. eelpout [AHB^{+}17]. eels [BSC^{+}14, CSA^{+}15, CHF^{+}12, EMA17, EA18, FDST16, KAST17, LML12, LFP^{+}11a, LFP^{+}11b, WCW^{+}16b]. Effect [AMA11, CMEV17a, CAS^{+}18, DSSB18, DSSB19, GVB^{+}17a, HKBD15a, JQK14, KSM^{+}17, KBE^{+}14, KCL14, MMD^{+}12, RRG^{+}19, SSSP18, SK15, SDR^{+}15, SCS^{+}12, SW11, AHH16, AH16, AMRHM18, ACA^{+}16, AAV^{+}16, BND^{+}17, BLP10, CGR17, CMH^{+}17, DSA^{+}12, ESA^{+}10, FFP^{+}14, GBV^{+}16, GHB^{+}11, HFKF17, IPW^{+}13, LAN^{+}14, MTK^{+}10, OLC11, PVIM18, SSC17, YSHG18, YFA^{+}17, SOGR^{+}19]. effective [ANB^{+}19, CB13, SMFM12]. effectively [RPMRLJ^{+}19]. effectiveness [GVB^{+}17b, GVB^{+}17a, NNC^{+}18, TJR^{+}18]. Effects [BPCS19, JBGP11, JJV^{+}15, KFK^{+}16, LKS^{+}13, LPC^{+}18, PG16, PMPG^{+}16, TTL^{+}16, TMC^{+}12, VDO^{+}18b, VMCM14, BSD^{+}18d, BHA^{+}17, EDP^{+}17, ELI^{+}17, HBB14, JST^{+}19, LLL^{+}18a, MDF^{+}17, MMMD16, NCNB18, PNBK13, PMBM13, PPI^{+}13, SHS14, UKS^{+}18, VCLW16, VBL^{+}15, WXWD15, XKB12]. Efficacy [BNS13, FHD13, HLL^{+}12, HJN^{+}19, MSCH14, RSH11, SSV^{+}10, SGB^{+}16, SKGE13, WGL17, ZPM^{+}18, BAH^{+}16, CDW17, DD18, ESA^{+}10, HLHS11, HCL^{+}14, KWGS19, LML12, LWJ^{+}16, MWRSK17, MA16, NCNB18, OMB^{+}19, PZL^{+}14, RNM^{+}12, SMR10, SSSP18, TNK12, VDO^{+}18a]. efficient [EYF^{+}19]. efforts [JR19]. EGCG [WCR^{+}18]. egg [AHH16, BAH^{+}16, FHFF15, GM11, MGM^{+}18]. eggs [KBE^{+}14, MWRSK17, OBD10, SWWN^{+}16, SMFM12, TES11]. Egusa [ZCC18]. Egyptian [NFF^{+}17]. EHP [SSV^{+}17]. EHP-infected [SSV^{+}17]. Eigenmann [Lek12]. eight [JCN^{+}13, JCN^{+}14, KRFT14]. EIIB [XXL^{+}19]. Eimeriidae [dSdAdS^{+}19]. ELASA [YLX^{+}19a, ZLN^{+}17]. Elasmobranchii [MdOO16]. elasmobranchs [KJB11]. electric [MSS^{+}12]. electricus [MSS^{+}12]. electrophoresis [LLL^{+}12, TAC15]. Electrophorus [MSS^{+}12]. Eleginops [GOM^{+}19]. element [CYS17]. Elevation [Bin14]. ELISA [BWZ^{+}17, LMG^{+}19, ZWG^{+}18]. Ellidaár [áKÁGA16]. Ellis [Rob11]. elongation [PLCC11]. elongatoides [FŠK^{+}19]. EM-90-like [RSIP^{+}17, RSPA^{+}18, SHO^{+}17]. EM90 [MID^{+}19b, MID^{+}19a]. EM90-like [MID^{+}19a, MID^{+}19b]. emaciation [KOI^{+}14]. emamectin [GSS^{+}10, HLHS11, IPW^{+}13, JHGR13, KWGS19, LML12, PPI^{+}13, SMR10, SMM^{+}13, Sam10, WWE^{+}13]. emarginatus [VDO^{+}18a]. embedded [FBL^{+}14, FGR^{+}12]. emerald [HAP^{+}10, LRB^{+}17, MGN^{+}16]. Emergence [AHB^{+}17, JHGR13]. Emerging [BK12, EKT^{+}11, NSC^{+}15, OLVJ^{+}13, PRT^{+}14, VJWC^{+}18]. emphasis [VGI^{+}16, WBD10a]. employ [SWWN^{+}16]. employing [MFES18]. empty [CBM^{+}18]. EMS [BPK^{+}17, ITKH18]. EMS/ [ITKH18]. EMS/AHPND [BPK^{+}17]. Encephalomyelitis [MKH^{+}11, YYF^{+}10]. encephalopathy [DVC^{+}17, GFB^{+}13, NOST^{+}16, VPP^{+}12]. encoded [CCLL16, CGW^{+}19, LCC^{+}15]. endangered [VPP^{+}12]. endocarditis [WEPB^{+}18]. endogenous [BMBS12]. Endolimax [CPP14, CPdPP16]. endoparasitic [Buc12]. endothelial [OMO^{+}15, WCW^{+}16b]. England [TDJ^{+}10, TWJP10]. Enhanced [MTHB13, NNC^{+}18]. Enhancement [SRB12, LSC^{+}14]. enhances [CR12, ZCJ^{+}18]. enhancing [BRB^{+}12]. enolase [BMT^{+}15]. enrichment [YGL^{+}18]. enteric [DPC^{+}10, LXL^{+}18, XLZ^{+}17]. entering [CDM^{+}13, VCR^{+}11]. enteritis [MTLP^{+}18b, WZQ^{+}16]. Enterobacter [GCNB16]. Enterobacteriaceae [KMGK^{+}18]. enterobacterial [TAC15]. enterocytes [RRO^{+}13]. Enterocytozoon [KS19, SSV^{+}17, TSK^{+}19]. enteromyxosis [LBF^{+}14]. Enteromyxum [ERÁpSB14, SSJ^{+}19]. Enterospora [ACSB^{+}19]. enterotoxin [MBB12]. envelope [LC17]. environment [JFWH19]. environmental [DIVC16, GFM^{+}16, HSA17, KOB^{+}17, SMR16]. environments [EMA17, MKKM15, PJL^{+}12]. enzyme [BWZ^{+}17, BA14, NMIY14, YLX^{+}19a, ZLN^{+}17]. enzyme-linked [BWZ^{+}17, NMIY14, YLX^{+}19a, ZLN^{+}17]. enzymes [BSV^{+}18, BSD^{+}18a]. epaulette [JK12]. EPC [WBD^{+}10b]. Ephippidae [DTM^{+}10]. epicarditis [DKG^{+}17]. Epidemiological [GJVT^{+}11, SEK^{+}16, DIVC16, GCW17, JHRW17, MSBB^{+}16, MFS^{+}17, SMS^{+}10]. epidemiology [EEMCEGSP16, JJM^{+}17, NBJ10]. Epidermal [CKJ^{+}14, PJU^{+}13]. epidermoid [SdMVS^{+}16]. Epinephalus [OKS^{+}13]. Epinephelus [BFFA^{+}12, CBH^{+}15, DDBC^{+}15, FHZ^{+}16, GHH^{+}11, GHW^{+}15, KCD^{+}14, LXL^{+}18, LSC^{+}14, LSA^{+}15, LMM^{+}15, LMM^{+}16, NNC^{+}18, NDA^{+}19, OTN^{+}12, PMP^{+}18, RBS^{+}16, RSF^{+}17, SSF^{+}17, Wen16, WCL^{+}12, XLZ^{+}17, ZDWH18]. epithelial [AHPR^{+}16, BCLM^{+}19, BKG^{+}18, GCM^{+}18b, RBL^{+}11, VBL^{+}16]. Epitheliocystis [CMR^{+}16, MSR^{+}10, SN14, SPV^{+}18, SVV^{+}17]. Epithelioma [GSBEM15, WBD^{+}10b]. Epizootic [CCM^{+}11, BO15, CCT^{+}11, CLW^{+}13, CdAPdC^{+}19, DWL^{+}17, HHC^{+}18, LFL19]. Epizootics [BDB^{+}18, HGG^{+}11]. epizootiologic [FLVA10]. erectus [BCW^{+}15, QWTL17, SBHC18, WZQ^{+}16]. Ergens [BGRB14]. Eriocheir [DML^{+}16, DCZ^{+}17, FGJ^{+}10, HGW^{+}19, SCS^{+}12, ZB12, DCL^{+}17]. eriocheiris [DXX^{+}15, HGW^{+}19, LXX^{+}17]. ERM [WGL17]. erosions [ONA10]. Erratum [Ano10a, Ano10b, Ano11, Ano14, RMSH10, TCC10b].

Erysipelothrix [CSA^{+}15, PRY^{+}18]. erythrocytes [VSDS^{+}19a]. Erythromycin [JML19, DdRCP13, DPCdR14, LLX^{+}19]. erythromycin-resistant [LLX^{+}19]. erythropolis [PBMB11]. erythropterus [CLWJ13, TIK^{+}14b]. escaped [MKI^{+}15]. Escherichia [SNM^{+}16]. Esox [Ano12c, KF10b, MF12, Řeh13]. Essential [VGI^{+}16, BSD^{+}18c, OJRL10, VDO^{+}18a, ZKL^{+}19]. establish [OM10]. established [AZSSZZ13]. Establishment [BNC^{+}11, CLL^{+}10, GHH^{+}11, GHW^{+}15, IFF^{+}15, KLW10, KAESEM13, LJC^{+}19, WWS^{+}11, WWZ^{+}18, WZW^{+}18, XWP^{+}18, YNL11, dOAPLá^{+}17]. estimate [KKKA17]. estimates [GBV^{+}16, GVB^{+}17a, MMG^{+}11]. Estimating [HDWM15]. estuary [DCV^{+}14]. ethanolic [BAH^{+}16]. Etroplus [MRN^{+}19]. EU [VCR^{+}11]. Euglena [YSHG18]. Eurasian [BGRB14]. europaeus [DABO^{+}17]. Europe [BLL^{+}17, BMMP^{+}14, COB^{+}16, DKBS17, SBD11, SMJ^{+}14, VCS^{+}19]. European [ARFG13, BMD^{+}14, BND^{+}17, BLSG^{+}16, CGP^{+}19, CKTT10, EA18, FCM^{+}19, FPK^{+}17, FDST16, GFM^{+}16, KAST17, LMC^{+}16, MBD16, MFES18, MAOS^{+}18, MNM^{+}18, NJK^{+}17, PBM^{+}15, PST^{+}16, RRC^{+}17, RPM^{+}19, SWE^{+}19, TMC^{+}12, TNK12, Tre12, VTM^{+}14, VC17]. eutropha [HBH^{+}19]. eutrophic [HWO^{+}17]. evaluate [LTG^{+}18, RSF^{+}17]. evaluated [MSC^{+}17]. Evaluating [GVB^{+}17b, HCL^{+}14]. Evaluation [ATM^{+}10, CBLN13, HKBD15b, RAM^{+}19, ASS19, DSSB18, DSSB19, EAN^{+}18, FDST16, HLHS11, KWGS19, LTC^{+}14, LFP^{+}11a, LFP^{+}11b, LWJ^{+}16, LLWL15, MID^{+}19b, MNG^{+}18, PVIM18, PPK^{+}13, YDO^{+}16, ZCCN13]. event [BGC^{+}17, BVT^{+}18, BVL^{+}18, DABO^{+}17, FSC^{+}16, KHLL16, LEM^{+}19, MMW^{+}15]. events [CdAPdC^{+}19, GVB^{+}17a, PTM^{+}19]. Evidence [AHAMR^{+}19, CZB^{+}11, RDN^{+}12, RGD^{+}13, VME^{+}19, YTKS19, ZN10, APP^{+}17, CSS^{+}16, CRO^{+}18, COB^{+}16, DJQ^{+}12, DCMM13, GB18, GFW^{+}10, HSC^{+}16, IAR^{+}10, vGJHS^{+}11, KYK^{+}12, KSA^{+}10, MBS^{+}17, MDC^{+}18, MUM^{+}19, ISCF^{+}19]. evolution [SJ11, YGL^{+}18]. Examination [MTA^{+}15, HKF^{+}16, JMHK18, MMMD16, MFES18]. examining [JSJR^{+}19]. excised [GGW^{+}11]. Exclusion [AJLC^{+}16]. exercise [GMC^{+}16b]. exerts [BSD^{+}18d]. exhibit [JGI^{+}10]. exilis [RKPG16]. exitiosa [BLSG^{+}16]. exogenous [Góm12]. Exophiala [CBH^{+}15, ŘKH17, ŘKH18, SBP^{+}19]. Exophthalmia [HFKF17]. exopolysaccharides [RBH^{+}17]. exotoxin [RGD^{+}13]. expansion [TALA11]. experiencing [GJM^{+}15, NFF^{+}17, RHZA^{+}16]. Experimental [ABT^{+}14, CAJD^{+}18, CTK^{+}10, DCMM13, FSV^{+}18, ICVV16, vGJHS^{+}11, KS19, KF10a, LSA^{+}15, MUM^{+}19, SON^{+}18, WKW^{+}11, ZCP12, AHHR^{+}10, BSW17, Buc12, CJB^{+}15, EDP^{+}17, GFM^{+}11, GBSF12, GRF14, GRP^{+}15, JAR^{+}11, KSA^{+}10, LPA^{+}10, LTG^{+}18, MGM^{+}18, MMMD16, MID^{+}19b, PBM^{+}15, PNBK13, ŘKH18, RHB^{+}15, SGB^{+}16, TDS^{+}19, WIC^{+}19, YDO^{+}16]. experimentally [ACN12, Ano12c, BSV^{+}18, BSSB18, BSD^{+}18b, BSD^{+}18c, BSB^{+}18, FCM^{+}19, HGB^{+}11, JML19, KSM^{+}17, LF16, MF12, MTA^{+}15, MTAB15, RSIP^{+}17, SSV^{+}17, SFL^{+}16, TJB^{+}15, TSS18, VTM^{+}14]. experiments [FBAS16, KAESEM13]. exploited [LXX^{+}17]. exponential [YGL^{+}18]. exposed [Ano12c, CTK^{+}10, FBK^{+}16, FSO^{+}10, HSGD11, KSA^{+}10, KPM^{+}11, MF12, SCR^{+}15]. exposure [BSW17, CAS^{+}18, DDBC^{+}15, GFM^{+}16, GGW^{+}11, HCCM12, HKBD15b, HGG^{+}11, IPW^{+}13, MMMD16, PÅD18, TMC^{+}12, VCN^{+}10, VMCM14]. exposures [WNO^{+}18]. expressed [JKC^{+}11, LSA^{+}15, OJRL10, SSV^{+}10, TMM^{+}15]. expresses [CGW^{+}19]. expressing [EBM^{+}18, ZXW^{+}12]. Expression [ACYC10, PRSK10, PMK12, ZWY^{+}11, AMA11, BPK^{+}17, CLWJ13, CKJ^{+}14, CG11, CDG^{+}14, DMD^{+}17, EDN18, FBVC16, IPW^{+}13, JKH^{+}17, JJS^{+}12, LCV^{+}17, NSE^{+}17, NMIY14, NYRS12, RBM^{+}16, RHVG^{+}19b, RSPA^{+}18, SHKF16, SCK13, SWD^{+}14, SJSA^{+}14, TvWO^{+}16, UKS^{+}18, WCL^{+}12, YK12, YYW^{+}14]. extant [GRP^{+}15]. external [AJLK^{+}18, DPD^{+}18, HCEH16, JMHK18, PDLW17]. extra [WCW16a]. extracellular [BGG11]. extract [BAH^{+}16, HDC18, MSCH14]. Extraction [LRD^{+}19]. extracts [YLX^{+}19b]. Extreme [CBM^{+}17, NNCLGC^{+}11]. extremely [SVL^{+}17]. Exuberant [HSHC^{+}16]. eye [PAM^{+}19, SM14].

F [BWM^{+}13, PBC15]. F-fluorodeoxyglucose [BWM^{+}13]. F-type [PBC15]. F. [GDW^{+}15]. Fabaceae [VDO^{+}18a]. Fabricius [DCMM13, DBM^{+}14, FSO^{+}10, JJS^{+}12, RDN^{+}12, STS^{+}15]. facilities [VPP^{+}12]. facing [GAVC16]. factor [BRB^{+}12, KREM14, LSC^{+}14, PLCC11, SSJ^{+}19, YAHWU11, ZKL^{+}19]. factor-Tu [PLCC11]. Factors [CHAB10, HGG^{+}11, TES16, BHF13, DSSB18, DSSB19, DIVC16, FMB^{+}11, JSJR^{+}19, KAS^{+}18, MMMD16, RMS^{+}13, ZZC^{+}17, ZJDL18]. facultative [AHAMR^{+}19]. faeces [GCNB18, GBSF12]. fairy [SHW^{+}11, SHS14]. fallowing [OM10]. families [JAR^{+}11]. family [CSS^{+}16, MNG^{+}18, UN10]. fan [ZJDL18]. fantail [CDM^{+}16]. fario [BSPS^{+}18]. farm [AZSSZZ13, AdOG^{+}17, BPK^{+}17, BVG^{+}10, EWRH13, KTP^{+}15, KDA^{+}18, LLH^{+}14b, RRC^{+}17, SMM^{+}13]. farm-raised [AdOG^{+}17, KTP^{+}15]. farmed [ATP^{+}16, AGR^{+}16, AGFG19, ATP^{+}17, AHTCA^{+}17, BVK^{+}16, BHO^{+}19, BO15, BMH^{+}19, BGCKF12, CGP^{+}19, CRAG^{+}12, DTM^{+}10, DWP^{+}19, DPD^{+}18, FKB^{+}14, FMH^{+}15, FSB^{+}18c, FCS^{+}18, GRB^{+}17, GBO^{+}18, GGH^{+}10, GFR^{+}12, GCGR13, HKK^{+}11b, HKD^{+}10, HPY^{+}13, IAR^{+}10, JCS11, JHRW17, JHGR13, KAS^{+}18, KOB^{+}17, KYK^{+}12, KAST17, MKI^{+}15, MSMM19, MLMR16, MLER^{+}18, MMB^{+}15, MCSM17, MRSC13, MSR^{+}10, MHNY18, MKH^{+}11, MGH^{+}14, MCW^{+}18, MNM^{+}18, MNM^{+}19, NOV^{+}11, NMM^{+}18, OGI^{+}18, PdLC19, PRT^{+}14, PGF^{+}19, QXW14, QPB^{+}16, ŘKH16, RFA^{+}14, RMSH09, SMR10, SCD19, SDWA18, SDM^{+}19, SPS^{+}17, SKJ13, SSA13, SNK^{+}19, Tre12, VDMJ^{+}13, WMM^{+}16, WNAJ^{+}16, YLL^{+}18, YAHWU11, YP17, ZCCN13, ZNHC11a, ZNHC11b, ZZA^{+}14, ZSZF16]. farming [BNS13, HSA17, JJB15, JR19, TAQ^{+}10]. farms [BPK^{+}17, BAH^{+}18, CLS^{+}15, CdAPdC^{+}19, EWRH13, FLC^{+}17, Gis18, GFW^{+}10, ISCF^{+}19, MSBB^{+}16, MSK^{+}12, MCO11, NB11, NRF^{+}18, NFF^{+}17, OM10, SVM^{+}18, SJK^{+}14, SVY^{+}14, WDM^{+}15]. Faroe [Gis18]. fasciatus [JJV^{+}15, OKS^{+}13]. fast [MLWC11, Sla14]. fat [RRG^{+}19, SHV^{+}16]. fatal [FSM^{+}16]. fate [BSWS11]. fathead [AHHR^{+}10, CXW^{+}19, JBGP11, MGN^{+}16, PPF^{+}12, WBD^{+}10b]. fatty [DTF^{+}16, MLF^{+}10]. fauna [HSA17]. favoured [EMA17]. Fe [PLCC11]. Fe-S [PLCC11]. features [BOH^{+}12, DMD^{+}17, FFLC17, RAM^{+}17, VFB^{+}12]. fecundity [USM17]. fed [BSD^{+}18d, TMC^{+}12]. feed [AA10, ESA^{+}10, JKL^{+}13, KSM^{+}17, NFA15, PPI^{+}13, PFC^{+}13, RHB^{+}15, XKB12, dOAPLá^{+}17]. feeding [JSAH^{+}18, SJSA^{+}14]. feeds [CBW19, NPGA^{+}19, PNBK13]. female [DBS12, RAM^{+}17, ZDWH18]. fenbendazole [FHD13]. Fenneropenaeus [NRMS14, WLX^{+}10]. fennicus [VDH19]. feral [AAUHW15]. fermented [LKS^{+}13]. Ferreira [HSGD11]. ferritin [NRMS14]. fibroblast [KLW10, SSG^{+}19, VBL^{+}16]. fibroblastoid [VKK^{+}19]. fibroblasts [DJQ^{+}12]. fibromas [BCW^{+}15]. fibrosarcoma [DD15]. Field [MMD16, OTN^{+}12, BSBH10, BWP^{+}19b, FSG^{+}12, HH13b, HBB^{+}15, HKB^{+}16, JMS13, LTG^{+}18, MRR^{+}14, RNM^{+}12, TAC15, WWE^{+}13]. field-collected [WWE^{+}13]. fighting [RhAS^{+}15]. filaments [MGM^{+}16]. Filifolinone [PVIM18]. fillets [TWNB^{+}12, TRS19]. Filtration [BRL^{+}13]. fin [ASSP^{+}15, BNC^{+}11, CLL^{+}10, FPK^{+}17, GGW^{+}11, KLW10, LJC^{+}19, MSC^{+}17, PWHW18b, SSG^{+}19, VBL^{+}16, YNL11]. fin-derived [SSG^{+}19]. findings [STS^{+}15]. fine [DPC^{+}10]. finfish [WBD10a]. fingerlings [PYSS19, SSSP18]. fingerprinting [CYS17]. Finland [IAR^{+}10]. Finstad [JCN^{+}14]. First [ARR^{+}18, BLdA^{+}18, BRMG13, BQP12, BBL^{+}16, BMMP^{+}14, BMA13, BVT^{+}18, CTD^{+}12, CWW^{+}18, CSA^{+}15, DML^{+}16, EPA^{+}10, GRB^{+}17, GGH^{+}10, HSC^{+}16, HHC^{+}18, HKF^{+}16, IGLG^{+}17, KMGK^{+}18, KPK^{+}18, KDP^{+}17, LFSZ18, LYL^{+}18, LL12, LPVGL^{+}10, MdOO16, MFES18, MVWS13, MDC^{+}18, MHNY18, MCCN14, MNM^{+}19, OGI^{+}18, OCLI^{+}19, PSVK^{+}18, RHZA^{+}16, RBF14, RPJ^{+}18, SPP^{+}18, SFBW^{+}17, SPS^{+}17, SMV^{+}14, SdMVS^{+}16, SWE^{+}19, VDD^{+}18, WBE^{+}17, ZxYJ^{+}19, ZCC18, dSRSHGá^{+}19, APP^{+}17, GL18, GKCC^{+}12, HST^{+}13, JDRM17, LUP^{+}18, TIK^{+}14a]. Fish [Buc12, DPADO16, KPSP^{+}19, Loh13, Spe19a, ADMF15, ACAS14, AJLK^{+}18, AAUHW15, AFTEML19, AHITC18, AHAMR^{+}19, BSD^{+}18c, BSB^{+}18, BKS^{+}17, BHF13, BPMI^{+}16, Bin14, BFJ11, BCL^{+}19, BMBS12, BGC^{+}17, BWNS^{+}11, BW18, BMMP^{+}14, BGD^{+}17, BFFA^{+}12, BNS13, BWM^{+}13, BMA13, CDM^{+}13, CHH^{+}14, CDF^{+}18, CMSHS16, CD17, ÇOÇ^{+}16, CGCAH12, dSCBdS^{+}16, Dal18, DTM^{+}10, DBV^{+}13, DZC^{+}16, DMD^{+}17, DLCF13, DSA^{+}12, DAB^{+}12, DMK^{+}19, DSJA18, EKT^{+}11, EBM^{+}18, ERÁpSB14, EPCSBP18, FBS13, FBAS16, FBK^{+}16, FLVA10, FMA^{+}19, FMB^{+}11, FCS^{+}18, GCNB18, GFB^{+}13, GJK^{+}11, GJMH17, GOM^{+}19, GM10, GMN12, GWM12, GCGR13, GDN^{+}16, HCCM12, HSA17, HGG^{+}11, HY16, HGB^{+}11, HAS^{+}19, HWO^{+}17, ISCF^{+}19, JDRM17, JSJR^{+}19, KHC^{+}18, KOI^{+}14, KFK^{+}16, KZJ^{+}17, LFSZ18, LITCAH19b, LFH^{+}11, LL12, LLH^{+}14b, LHW^{+}15a, LDEH17, LRMN17, MYWC18, MGM^{+}18, MFHHG18, ME12, MKK^{+}11]. fish [MLWC11, MR11, MJ17, MTHB13, MoSMS18, MSK^{+}12, MMNL14, NN11, NFA15, NFF^{+}17, NMIY14, NSC^{+}15, NHM^{+}10, OT12, OJRL10, PPSW19, PKPG15, PKLH19, PSCFB^{+}18, PNBK13, PRY^{+}18, PBW^{+}18, RHS^{+}12, RHST^{+}16, RhMM^{+}15, RhAS^{+}15, RhMV^{+}15, RAM^{+}19, RRC^{+}17, RRO^{+}13, RAS^{+}10, RCD^{+}17, RHBMM13, RE14, SJDC^{+}13, SMF^{+}10, SSHEM11, SZM^{+}15, SCK13, SWD^{+}14, SMJ^{+}14, Spe19b, SMS^{+}18, SKB10, SKJ13, SN14, SSO^{+}14, TIK^{+}14a, TSS18, TWJP10, TAQ^{+}10, TQW^{+}14, TMWV14, VMCM14, VCR^{+}11, VC17, VME^{+}19, WDM^{+}15, WMM17, WMD^{+}18, WYZ^{+}19, WKW13, WNCR12, ZZC^{+}17, ZHFC11, ZGY^{+}18, ZXW^{+}12]. fish-farming [BNS13, TAQ^{+}10]. fish-myxozoan [EPCSBP18]. fish-pathogenic [CHH^{+}14, FLVA10]. fish-to-fish [HGG^{+}11]. fishable [MMG^{+}11]. fishery [MMG^{+}11, TJR^{+}18]. fishes [ARL^{+}17, AAZS^{+}15, FDL^{+}17, KPT^{+}11, MA16, PAZ18, RWTB17, RBR19]. fishpathogens.eu [MPJ^{+}15]. fishpathogens.eu/noda [MPJ^{+}15]. five [BOH^{+}12, BBP18, THW^{+}12]. fixation [CBLN13, CMS^{+}17]. fixed [FGR^{+}12, WWE^{+}13]. fixed-dose [WWE^{+}13]. fjord [KAST17, SVY^{+}14]. flagellar [YK12]. flat [BLSG^{+}16]. flavescens [KHLL16, OEG^{+}13, PHvdH12, SGQ^{+}15]. Flavobacteria [ATJS^{+}18, LF18]. Flavobacterium [BBF^{+}15, BLF^{+}16, BMA13, BLFC12, CHV^{+}12, CAJD^{+}18, DBS12, DBV^{+}13, DMP10, DJR^{+}19, DLPR15, DSLR16, FDL^{+}17, FNISE12, GLW^{+}18, GCW17, GPL^{+}12, GDW^{+}15, HKBD15a, HKBD15b, HJN^{+}19, HRNW15, JDH^{+}18, JML19, KSS^{+}15, LLC^{+}11, LWWS14, LGD^{+}17, LGS19, LFD^{+}17, LZZN10, LLL^{+}12, LF16, LPCC12, MBS^{+}19a, MBOC19, NN11, NSB^{+}18, NOV^{+}11, OFA11, OBD^{+}11, PDLW17, PLCC11, PFJ^{+}10, RHB^{+}15, SDWA18, dASLS^{+}19, SL15, SLPB18, SEK^{+}16, SWE^{+}19, SBW11, SW11, TQW^{+}14, WSA^{+}18]. FLIV [JLJ16]. Florfenicol [DdRTM13, FHZ^{+}16, JML19, RHTC19b, RSH11, SKGE13, YVM^{+}14, dOAPLá^{+}17]. Florida [BA13]. flounder [DBHC17, HKK^{+}11b, JKC^{+}11, JLJ16, KKK^{+}18, LKS^{+}13, LPC^{+}18, SSJ^{+}19, TOK^{+}16, WWS^{+}11]. flow [BRL^{+}13, CHB^{+}17, PSK^{+}13, WSX^{+}18]. flow-through [PSK^{+}13]. flower [RHST^{+}16, RhMM^{+}15, RhMV^{+}15]. fluorescence [AS17, DXX^{+}15, KTK17, LGC^{+}19]. fluorescence-based [AS17]. fluorescens [GPFGAG19]. fluoro [SRD^{+}16]. fluorodeoxyglucose [BWM^{+}13]. fluviatilis [BGRB14, BGCKF12, CGP^{+}19, PBM^{+}15, RPM^{+}19, SKPS16, VPP^{+}12]. focal [BHO^{+}19]. folifer [BCJ13]. follicles [CZB^{+}11]. follow [BMGS17, PPI^{+}13]. follow-up [BMGS17, PPI^{+}13]. following [ASSP^{+}15, BSW17, CRS^{+}13, ECP^{+}10, FGJ^{+}10, GVB^{+}17a, GMC^{+}16b, GBSF12, HCCM12, IPW^{+}13, JAJC^{+}18, LWJ^{+}16, MBS^{+}19a, MMG^{+}11, PSK^{+}12, PFC^{+}13, TWNB^{+}12]. follows [KDA^{+}18]. fomites [PJL^{+}12]. fontinalis [DGS^{+}14]. Food [LDÁM14, SNM^{+}16]. foodborne [TOK^{+}16]. foraging [BND^{+}17]. foreign [LL11]. fork [DD15]. fork-tailed [DD15]. formalin [FGR^{+}12]. formalin-fixed [FGR^{+}12]. formation [DSSB18, DSSB19, LFD^{+}17, LITCAH19b, LTC17, SFR^{+}17, SW11]. formerly [MJ17]. forming [CD17]. formosus [KVC^{+}16]. forms [SQC^{+}14]. Forsskål [DTM^{+}10, LQY11, LSC^{+}14, LSA^{+}15, LMM^{+}15, LMM^{+}16, SGR^{+}18, Wen16]. Forster [ABC^{+}10, CB13]. fortis [WZQ^{+}16]. fortuitum [PBW^{+}18]. fosteri [SFF^{+}17]. found [AHTCA^{+}17]. Fouquet [vGJHS^{+}11]. four [KKT^{+}18a, NWMK16, XKB12]. Fourier [WNCR12]. fractions [LLC^{+}11]. fracture [FvdMF^{+}18]. France [BVL^{+}18, BBL^{+}16]. franciscana [TNK12, WBF^{+}10]. Francisella [ATP^{+}17, BFJ11, CDM^{+}13, GIW^{+}11, KBTMH^{+}16, PLC^{+}19b, SPKS17, SKGE13]. francisellosis [AdOG^{+}17, KBTMH^{+}16, RPMRLJ^{+}19, RBWR^{+}15]. Frank [Cip11]. Fraser [BLP10]. Fredericella [KAESEM13]. free [HKF^{+}16, MB17, MPJ^{+}15, YSV^{+}13]. free-swimming [HKF^{+}16]. freedom [JKS^{+}13]. Freeman [MRSC13]. French [DTM^{+}10]. frequency [LRS^{+}17]. fresh [JZK^{+}14, MSR^{+}10, NOV^{+}11, WNGS^{+}17]. freshwater [ADMF15, ACA^{+}16, AAV^{+}16, Ano10c, ACM^{+}14, BAD^{+}12, BMMP^{+}14, CJB^{+}15, DXX^{+}15, FMT^{+}14, GB18, GTT^{+}17, GGH^{+}10, GJMH17, JTW^{+}10, LFSZ18, LL12, LYKN18, MRR^{+}14, MdOO16, MSK^{+}12, NHM^{+}10, PPSW19, PKPG15, RNM^{+}12, RWTB17, VPP^{+}12, WMM^{+}16, WOD16, WNO^{+}18]. freshwater- [GTT^{+}17]. freshwaters [SMJ^{+}14]. FRET [SMF^{+}10]. frog [DPADO16]. Front [MBS^{+}19b, RHVG^{+}19a]. frozen [TRS19]. fry [CJB^{+}15, DSLR16, GCM^{+}18b, HKBD15b, KEY^{+}15, PDLW17, SCR^{+}15, WKK12]. FTA [PSK^{+}13]. Full [NRMS14, NDA^{+}19]. Full-length [NRMS14]. function [BMD^{+}14, JBGP11]. functional [WCL^{+}19, ZZQ^{+}19]. Fundy [NRF^{+}18]. funebris [BCG^{+}17, CHF^{+}12]. fungal [ŘKH16]. fungicide [MMMD16]. fungus [SMFM12]. furcatus [LTC^{+}14, XKBC12]. furnieri [MPM16]. Further [KWGS19]. furunculosis [GGH^{+}10]. Fusarium [KDP^{+}17]. fuscoguttatus [SSF^{+}17, LSC^{+}14, LSA^{+}15, LMM^{+}15, LMM^{+}16, NDA^{+}19, Wen16, ZDWH18]. fusions [DWP^{+}19]. future [ZGY^{+}18].

g [LSC^{+}14, CGW^{+}19, LZZN10]. Gadus [BSWS11, BMGS17, BK12, FvdMF^{+}18, GIW^{+}11, GGGM14, HLHS11, HKD^{+}10, HFKF17, MVM^{+}18, MBB12, OBD10, PG16, RBWR^{+}15, Sam10, ZCP12]. Gaimard [CCMA^{+}17, PC17, WCW^{+}16b]. Galicia [DIVC16, RAVC18]. gamete [LST^{+}12]. gametes [KSA^{+}10]. GAPDH [ZXW^{+}12]. gariepinus [DAO19, HGPD11, HOHG13, LvD18, OOO^{+}11, OAO^{+}12, SRT^{+}14, TMWV14]. garlic [MSCH14, NDA10]. Garman [AFA^{+}12]. Garra [VME^{+}19]. garvieae [ABT^{+}14, CLT^{+}14, DPMK12, FLC^{+}17, HY16, PSBG^{+}11, TAC15]. Gas [SBHC18, RPJ^{+}18]. Gasterosteus [ASW16, PJZB19, WFH^{+}18]. Gastric [LWMT10, dSdAdS^{+}19, BB11, LL11]. gastroenteritis [DPTFC10, DPTCF10, RBL^{+}11]. gastrointestinal [ABF15, BLFC12]. GBS [BFFA^{+}12]. GC [ZCJ^{+}18]. GC-MS-based [ZCJ^{+}18]. GCRV [ZWG^{+}18]. gel [TAC15]. Gene [RHVG^{+}19b, AMA11, CLWJ13, CCLL16, CGW^{+}19, CGG^{+}10, CKJ^{+}14, CKTT10, dJDAVRGM^{+}18, EDN18, FMA^{+}19, JJS^{+}12, LCV^{+}17, LCC^{+}15, LXZ^{+}19, MPBM^{+}16, MNG^{+}18, NSE^{+}17, OBR^{+}14, QZZ^{+}18b, RB15, RSPA^{+}18, SWD^{+}14, TMM^{+}15, UKS^{+}18, YYW^{+}14, ZWY^{+}11, ZZQ^{+}19]. gene-encoded [CGW^{+}19]. generated [SK15, SEK^{+}16]. Generation [CHB^{+}17, DWL^{+}17, LMM^{+}16]. genes [BCLM^{+}19, BGC^{+}17, CIL^{+}17, COB^{+}16, CDG^{+}14, FVAVMGE18, FCL^{+}19, JKH^{+}17, LWWS14, LZZN10, LMM^{+}15, LMM^{+}16, NRMS14, OJRL10, SSV^{+}10, SJSA^{+}14, TvWO^{+}16, YK12]. Genetic [DMP10, HJT^{+}16, KTP^{+}15, MPBM^{+}16, SDZN19, TOK^{+}16, TCM^{+}19, TWYC15, WM19, WNAF^{+}13, WSA^{+}18, ASB^{+}12, BFC^{+}15, COB^{+}16, DZV10, DLPR15, ISCF^{+}19, KNN19, LDEH17, NWMK16, RGMGG^{+}18, SDWA18, SMC^{+}11, VCS^{+}19, VST^{+}19, ZFZ^{+}11]. genetically [JAR^{+}11]. genogroup [ISCF^{+}19]. genogroups [ABVS18, KWD^{+}16]. genome [BFC^{+}15, IBFL19, KKM^{+}19, RKK^{+}19, ZxYJ^{+}19]. genomes [BHF13, YEK^{+}17]. Genomic [DZC^{+}16, RVC18, GBO^{+}18, JGI^{+}10, LCV^{+}17, LLLG18, TCS19, YMH^{+}11]. genomovar [BMA13, GLW^{+}18, LWWS14, LGD^{+}17, SL15]. genotoxicity [BA14]. genotype [Ano12c, CVLJ^{+}10, CCW^{+}15, CWW^{+}18, GJVT^{+}11, GVC^{+}19, KF10a, KKK^{+}18, LRD^{+}19, MF12, VCS^{+}19, VBL^{+}16, WKW^{+}11, ZWG^{+}18]. genotypes [HB12, MMWK19, RPGCIG17, TDS^{+}19, VLVS^{+}16, VDHV^{+}16]. Genotypic [DMK^{+}19, PPSW19, WMD^{+}18]. genotypical [PKPG15]. Genotyping [CLLF14, FSB^{+}18c, BBC^{+}10, FNISE12, LGS19]. genus [NSC^{+}15]. Genypterus [IGLG^{+}17]. geographic [AGFG19, GRG^{+}16, SHO^{+}17]. Geographical [GFR^{+}12, CRS^{+}13, DMK^{+}19]. geranyl [PVIM18]. Germany [AHJS^{+}18, BGRB14, KAST17]. GFP [RGD^{+}12]. GFP-labelled [RGD^{+}12]. Ghanaian [VJWC^{+}18]. Ghar [YMBE13]. giant [BFFA^{+}12, DCMM13, DBM^{+}14, FMT^{+}14, GHW^{+}15, LYKN18]. gibbosus [LDEH17]. gibel [LJHG14, YLW^{+}14, ZWLG10, ZYW^{+}10]. gibelio [KJW^{+}17, LJHG14, LLL^{+}13, WXL16, YLW^{+}14, ZWLG10]. GICA [LXX^{+}17]. gigas [GVLE16, GCCLRP^{+}13, HPY^{+}13]. Gilbert [DBHC17, WWS^{+}11]. Gill [BGRB14, MRSC13, MCO11, VCLW16, WNGS^{+}17, ACN12, AHPR^{+}16, ABC^{+}10, BCLM^{+}19, BPCS19, BLP10, CONN10, DYML^{+}18, DT16, FGR^{+}12, GGMG18b, GTT^{+}17, GCM^{+}18b, GDW^{+}15, GFS^{+}14, LDA^{+}19, MLER^{+}18, MR11, MGM^{+}16, MCCN14, NBJ10, OSAA18, PGF^{+}19, QPB^{+}16, TCC10a, WNO^{+}18]. Gill-associated [MCO11]. gills [CBLN13, FMM^{+}19, HKBD15a, vGJHS^{+}11, KSM^{+}17, KVC^{+}16, MAOS^{+}18, WNMK^{+}16]. gilthead [ACSB^{+}19, BC11, CVLJ^{+}10, CVGR^{+}13, CCME16, DPCdR14, RHVG^{+}19b, TPD^{+}19, LGS^{+}14]. ginbuna [NSS^{+}17]. Ginsburg [DSMC17, LKT^{+}13]. Girard [DXX^{+}15, PSK^{+}12]. GIS [AGFG19]. GIT [ABF15]. GIV [LCC^{+}15]. GIV-2L [LCC^{+}15]. gives [SWS^{+}16]. gland [SBHC18]. glanis [MAOS^{+}18]. glassfish [Góm12]. gleaned [MJ17]. Global [SJ11]. globalis [SAS^{+}19]. glomerular [DBHC17]. glomerulonephritis [LWMT10]. glucan [JSAH^{+}18, SJSA^{+}14, UKS^{+}18, YSHG18]. glucans [KBE^{+}14]. glucidic [LDÁM14]. Glugea [ZCC18]. glutaraldehyde [MoSMS18]. glycogenation [TIK^{+}14b]. Glycolic [ASHZ^{+}12]. Glycophorin [AI11]. glycoprotein [ÇOÇ^{+}16, EBM^{+}18, HLL^{+}12, IPW^{+}13, MPBM^{+}16]. glycoprotein-based [ÇOÇ^{+}16]. glycosylation [FE11]. Gnathia [HSGD11]. gnotobiotic [RGD^{+}12]. gnotobiotically [WBF^{+}10]. goby [GFB^{+}19, WZL^{+}11]. goes [BSW17]. gold [KVC^{+}16, LXX^{+}17, SSHEM11]. golden [AZSI^{+}12, CWL^{+}17, HBB14, KHC^{+}18, LYL^{+}18, SM14, SWD^{+}14]. goldfish [AHJS^{+}18, BSK^{+}10, BBL^{+}16, GAEK^{+}14, LCR^{+}10, RMAM12, WIC^{+}19, YNL11]. goldlined [SGR^{+}18]. goldring [WA14]. goldsinny [BT14]. goldsmidi [ABC^{+}10]. gonads [VMCM14]. gondii [SZM^{+}15]. Goode [CHF^{+}12]. Gorbunova [ME12]. gorbuscha [BPPKJ10, SMJ^{+}12]. gourami [RWTB17]. Goussia [dSdAdS^{+}19]. governs [FE11]. gracilis [YSHG18]. grade [VPCM19]. gradient [DLCC^{+}13]. Gram [BHA^{+}11, THW^{+}12]. Gram-negative [THW^{+}12]. Gram-positive [BHA^{+}11]. granular [RVC18]. granulation [HSHC^{+}16]. granulomas [AJLC^{+}16, JK12, ZZA^{+}14, ZWG^{+}17]. Granulomatous [MTLP^{+}18b, VPS^{+}14, CPP14, NHM^{+}10, dSRSHGá^{+}19]. grass [GCWB11, LLL^{+}12, LXHL11, WCR^{+}18, YYW^{+}14, ZWG^{+}18]. Great [FDL^{+}17, KF10b, GMP^{+}11, RBF14, EKED^{+}17, KF10a, OEG^{+}13, WKW^{+}11]. greater [MKH^{+}11, dSRSHGá^{+}19]. Greek [BCL^{+}19]. Green [TCFáS19, BCG^{+}17, KTK17]. Grenada [HHM^{+}14]. grey [BK12, PMBM13]. GroEL [KREM14]. gross [DPTCF10, LFP^{+}11a, LFP^{+}11b]. grounds [BK12]. group [BVL^{+}18, CDD^{+}18, EMA17, GAC10, VBL^{+}15]. grouper [BFFA^{+}12, CBH^{+}15, CCLL16, CGG^{+}10, DDBC^{+}15, FHZ^{+}16, GHH^{+}11, GHW^{+}15, HLC^{+}15, HHX^{+}15, KKK^{+}18, LCC^{+}15, LSC^{+}14, LSA^{+}15, MRN^{+}19, MSI^{+}15, NNC^{+}18, NDA^{+}19, OTN^{+}12, OKS^{+}13, PMP^{+}18, RBS^{+}16, RSF^{+}17, Wen16, WCL^{+}12, XJL^{+}19, YLX^{+}19b, ZLN^{+}17, ZDWH18]. groupers [KCD^{+}14, LXL^{+}18, SSF^{+}17, XLZ^{+}17]. groups [RB15]. grow [KDA^{+}18]. grow-out [KDA^{+}18]. grown [WBF^{+}10]. Growth [MB17, NN11, ACA^{+}16, CAS^{+}18, CDG^{+}14, DSSB18, DSSB19, GBO^{+}18, HMP^{+}12, LKS^{+}13, LWZ^{+}18, MLWC11, MPS^{+}18, OLT^{+}10, PMW^{+}10, PMH^{+}11, RHB^{+}15, STF^{+}15, SLPB18, SFU^{+}13, VDHV^{+}16, WKW13]. grunter [KRD^{+}17, LLH^{+}14b]. guamaensis [dSdAdS^{+}19]. Guangxi [LYL^{+}18]. Gubanov [CIA^{+}18]. gudgeon [MKK^{+}11]. gueldenstaedtii [RFA^{+}14]. Guichenot [IGLG^{+}17]. guidelines [HHO14, LTM^{+}19]. Guilford [SGQ^{+}15]. guitarfish [AFA^{+}12]. Guitart [ABM^{+}15]. Gulf [BB11, DBHC17, GCCLRP^{+}13, PJZB19]. Günther [Dig13, Dig15, IOO15, Pol15, RBF14, SRT^{+}14, XJL^{+}15, YCNW18, ZQB^{+}11]. guppies [BCM^{+}17, LCOZ10]. guppy [SPLV^{+}15]. Gut [JKL^{+}13, DCZ^{+}17, RBL^{+}11]. guttulatus [BQP12]. Gymnocephalus [MGM^{+}16]. Gymnocorymbus [SHAM14]. Gymnothorax [BCG^{+}17, CHF^{+}12]. Gymnotus [DMD^{+}17, VMCM14]. Gymnura [BCJ13]. gyrA [HBB^{+}15]. gyrodactylid [YEK^{+}17]. Gyrodactylus [GAC10, HHO14, HKG^{+}15, KHG^{+}17, LP17, PHVM13, PB18, WKK12, YEK^{+}17].

H [ZB12]. H. [NRMS14]. H6 [AJLK^{+}18]. Haan [ZSZF16]. habitats [OMO^{+}15, YkSKY13, ZYW^{+}10]. haddock [JHM^{+}18]. haem [OJRL10]. haemal [FvdMF^{+}18]. haemangioma [SMD^{+}12]. Haematological [KPM^{+}11]. Haematopoietic [LLL^{+}13, ASHZ^{+}12, EBM^{+}18, JSK^{+}10, MNM^{+}19, PMW^{+}10, PMH^{+}11, PBM^{+}15, SJDC^{+}13, YHY^{+}16]. haemic [MMMD16, MMD16]. haemocyte [TDLT^{+}16, TDT^{+}16]. haemocytes [FBVC16]. haemoglobin [UN10]. Haemogregarinidae [MdOO16]. haemolysin [DJQ^{+}12]. haemolytic [MFS^{+}17, OAO^{+}12]. haemorrhage [PWW19]. haemorrhagic [AHHR^{+}10, AHPR^{+}16, Ano12c, DPADO16, EKED^{+}17, GJVT^{+}11, GFB^{+}19, GGW^{+}11, GVC^{+}19, HLL^{+}12, HGG^{+}11, HPAL10, JKS^{+}13, KKK15, KF10a, KF10b, LVADB17, MMBS10, MF12, MMW^{+}15, OEG^{+}13, PBM^{+}15, PJL^{+}12, RMS^{+}13, SHJ^{+}12, SJ11, VLVS^{+}16, VBL^{+}15, VBL^{+}16, WDM^{+}15, WKW^{+}11, ZLX^{+}19]. Hafnia [OSC^{+}14]. half [XJL^{+}15]. half-smooth [XJL^{+}15]. halibut [BBP18, ONA10, RSH11]. Halieutaea [LZGZ16]. Halocynthia [KOS^{+}10, PZL^{+}14]. Hamilton [Bin14, CXK16, Góm12, KWGS19, KREM14, LHW^{+}12, LHW^{+}15b, MLR^{+}12, SMD^{+}12, SM14, WCL^{+}12]. Hamilton-Buchanan [Bin14, CXK16]. hand [SWS^{+}16]. handling [GAVC16, KZJ^{+}17]. handy [MPJ^{+}15]. haptor [TGJCH15]. harengus [GJVT^{+}11, MMBS10]. Harry [LFSZ18]. harveyi [DJQ^{+}12, HBVB16, MKMY10, NNC^{+}18, QWTL17, SSF^{+}17, SNM^{+}16, ZWY^{+}11, ZDWH18]. hatch [DPD^{+}18]. hatched [SBM^{+}11]. hatcheries [TES16]. hatchery [AZSSZZ13, BRMG13, CONN10, DABO^{+}17, FBS13, KDA^{+}18, LF18, PDdC^{+}14, SBP^{+}19, WBE^{+}17, XLZ^{+}17]. hatchery-bred [XLZ^{+}17]. hatchery-incubated [LF18]. hatchery-reared [BRMG13, CONN10, SBP^{+}19, WBE^{+}17]. hatching [VCLW16]. Hawaiian [WA14]. Haworth [BAH^{+}16]. Hb [UN10]. head [AFTEML19, CWW^{+}18, CHAB10, CSCCA14, Lek12, LRD^{+}19, NNCLGC^{+}11, STG^{+}10]. head-kidney [CSCCA14]. healing [SSG^{+}19]. health [BMGS17, HBB14, OT12, RAS^{+}10, TMWV14]. healthy [AI11, DPTFC10, GCCLRP^{+}13, HOHG13, KSM^{+}17, OSC^{+}14, TWNB^{+}12]. heart [FMH^{+}15, GRT12, GHW^{+}15, WHR^{+}18]. hearti [YLW^{+}14]. hearts [BOH^{+}12]. Heat [LRC^{+}18, RAS^{+}10, DJQ^{+}12, MWRSK17, SNM^{+}16, WBF^{+}10, ZWY^{+}11]. Heat-inactivated [LRC^{+}18]. heat-shock [SNM^{+}16, ZWY^{+}11]. heavily [JMHK18]. heavy [QWTL17, QZZ^{+}18b]. Heckel [AFTEML19, VME^{+}19]. held [KJB11, Sam10, WNGS^{+}17]. helminth [BGD^{+}17, ZPM^{+}18]. helminth-infected [ZPM^{+}18]. helminths [DPC^{+}10, DCG18]. hematopoietic [CGW^{+}19, HSC^{+}16, KWD^{+}16, LGJ19, MBOC19].

Hemiarius [DD15]. Hemigrammus [ŘKH18]. Hemigymnus [HSGD11]. Hemilepidotus [GM11]. Hemiscyllium [JK12]. hemocytes [HGW^{+}19]. Henle [CDM^{+}16]. Henneguya [RKPG16]. henneguyosis [RKPG16]. hepatic [BSV^{+}18, LvD18, SMJ^{+}12]. hepatitis [FKB^{+}14, KAS^{+}18]. hepatoblastoma [SCZM19]. hepatocellular [TIK^{+}14b]. hepatocytes [TIK^{+}14b]. hepatopancreas [CMEV17a, CMS^{+}17, CMEV17b]. hepatopancreatic [BPK^{+}17, CCLGRM^{+}17, DD18, DML^{+}16, KTN^{+}16, KYB^{+}15, MRR^{+}14, NPGA^{+}19, YTL10]. hepatopenaei [KS19, SSV^{+}17, TSK^{+}19]. Hepatospora [DCL^{+}17]. herbarum [dSRSHGá^{+}19]. Herbst [MRR^{+}14, RNM^{+}12]. heritage [GVLE16]. herpes [SGB^{+}16, SEM18]. herpesviral [LGK^{+}14]. Herpesvirus [BBL^{+}16, ASB^{+}12, BSK^{+}10, BMK17, BVL^{+}18, CLC^{+}17, CCW^{+}15, CCT^{+}11, DKBS17, EYF^{+}19, FBS13, FBAS16, FCC^{+}13, FSC^{+}16, GVLE16, GSEM13, GAEK^{+}14, HDC18, HPY^{+}13, IFF^{+}15, KYK^{+}12, KKOY19, KJW^{+}17, KAST17, LLL^{+}13, MLH^{+}15, MSS^{+}17a, MKY^{+}15, MTA^{+}15, MTAB15, MBT^{+}17, NSS^{+}17, PTM^{+}19, PFP^{+}13, RHZA^{+}16, RBH^{+}17, SMC^{+}11, TDJ^{+}10, TWJP10, TKS^{+}18, WXL16, WSX^{+}18, WIC^{+}19, YkSKY13, vBVLR^{+}16, MFES18, NJK^{+}17, GM10]. herpesvirus-1 [KYK^{+}12, MFES18]. herpesvirus-2 [NSS^{+}17, GM10]. herpesvirus-3 [GAEK^{+}14, MKY^{+}15]. herring [GJVT^{+}11, GGW^{+}11, HLL^{+}12, HCEH16, HGH^{+}16, MMBS10, WKW^{+}11]. Heterogeneity [BGCKF12, LWWS14]. heterologous [MBS^{+}19a, SLKE10]. heterophil [RSR^{+}18]. Heterophyidae [BCM^{+}17]. Heterosporis [KSAB^{+}14]. Hetrick [Cip11]. Hexacanalis [BCJ13]. Hexamitiasis [OP14]. Hfq [MFHHG18]. Hfq-sRNA-chaperone [MFHHG18]. HI [RGD^{+}12]. HI-610 [RGD^{+}12]. hidden [RBR19]. High [KKK^{+}18, RNM^{+}12, TCFáS19, BHT^{+}17, BLF^{+}16, BCBS^{+}13, DAB^{+}12, HGH^{+}18, HMW11, LITCAH19b, LZZN10, NFF^{+}17, OFA11, VCN^{+}10, VPCM19, WMD^{+}18, WZA10, WKW^{+}11, ZDWH18]. high- [BLF^{+}16, WZA10]. high-grade [VPCM19]. High-throughput [TCFáS19, BHT^{+}17, LITCAH19b]. higher [FHFF15]. highly [GCNB18, KF10b, LQL^{+}10, MAC^{+}15, YGL^{+}18]. Hilgendorf [HKK^{+}11a]. Hippocampus [BQP12, BCW^{+}15, DCV^{+}14, DSMC17, LYD^{+}15, LKT^{+}13, OBR^{+}14, QWTL17, SBHC18, WZQ^{+}16]. Hippoglossus [BBP18, ONA10, RSH11]. hirame [BMMP^{+}14, HMP^{+}12, SDJH16, YMH^{+}11]. histidine [FMA^{+}19, SFR^{+}17]. Histochemical [BDM^{+}18, DMD^{+}17]. histochemistry [DPC^{+}10, MBD16]. Histocompatibility [PdLC19, CPB^{+}15]. histologic [CSA^{+}15]. Histological [ADR11, GFM^{+}16, Góm12, LTC^{+}14, BCBS^{+}13, DPTCF10, HKF^{+}16, PPK^{+}13, PMBM13, RAM^{+}17, SVV^{+}17, TPD^{+}19]. Histology [ECP^{+}10, LvD18, CMEV17b, JBO^{+}15]. Histone [WCL^{+}19]. Histone-like [WCL^{+}19]. Histopathological [DCG18, DAO19, MKMY10, MAOS^{+}18, STS^{+}15, ABT^{+}14, MCSM17, MPM16, NB10, PMP^{+}18, QXW14, RHZA^{+}16, RHST^{+}16, RHSM^{+}18b].

Histopathology [JMHK18, ONA10, TGJCH15, ATM^{+}10, MBD16, SDZN19, YLW^{+}14]. historical [WMM17]. history [ARFG13]. histrix [MdOO16]. HN48 [LLL^{+}18a]. Hofer [MKKM15]. Hoffman [LRB^{+}17]. Holbrook [ALG15]. Hole [AFTEML19]. Hole-in-the-head [AFTEML19]. Holmberg [VGI^{+}16]. Holmes [SPTDEE^{+}15]. Holocene [ARL^{+}17]. Holocephali [DRdS^{+}10]. Homarus [CG11, SQC^{+}14]. homeostasis [TvWO^{+}16]. homogenate [SG12]. homogeneity [COB^{+}16]. hook [ABM^{+}15]. hooking [DLLB^{+}18]. Horizontal [NRF^{+}18, KS19, SHV^{+}16]. horizontally [WNGS^{+}17]. hormonal [GFM^{+}16]. hormones [DSSB19]. horn [RHST^{+}16, RhMM^{+}15, RhMV^{+}15]. horse [BCW^{+}15, ONGQ^{+}14, RGMGG^{+}18]. Host [ADMF15, ABC^{+}10, JFWH19, BCJ13, BLP10, EPCSBP18, FSB18a, FYO^{+}11, GSEM13, GAEK^{+}14, GGGM14, IBFL19, JPG^{+}19, KOB^{+}17, KF10a, LHW^{+}15a, LRB^{+}17, MMS^{+}16, MWS17, OWM^{+}17, PJZB19, RGD^{+}12, SPP^{+}18, SPTDEE^{+}15, SP15]. host-associated [KOB^{+}17]. host-related [FSB18a]. host-seeking [OWM^{+}17]. Hosts [SMKBP17, GRG^{+}16, HB12, KHG^{+}17, MLH^{+}15, ZYW^{+}10]. hounds [MS16]. hoyamushi [PZL^{+}14]. HPR0 [GL18]. HPV [YTL10]. Hsieh [SS17]. HSP70 [PHW^{+}16, SRB12]. human [VME^{+}19]. humic [YFA^{+}17]. humic-rich [YFA^{+}17]. humoral [CRS^{+}13]. humphead [HCW^{+}16, ZWY^{+}11]. Hungary [ABVS18, KPKP10]. Huso [RHSM^{+}18b]. hutCD [OJRL10]. hybrid [AZSSZZ13, BVG^{+}10, FCS^{+}18, ISCF^{+}19, LTC^{+}14, RHST^{+}16, RHSM^{+}18b, RhMM^{+}15, RhMV^{+}15, SSF^{+}17, SL15, SRT^{+}14, ZDWH18]. hybridization [ACSB^{+}19, CPEH15, DXX^{+}15, DCL^{+}17, LZZN10, MTA^{+}15]. hybrids [BSK^{+}10, KHG^{+}17]. hydrodomous [MRR^{+}14, RNM^{+}12]. Hydrogen [FVAVMGE18, ACN12, HKBD15a, HKBD15b, MGM^{+}18]. hydrolysates [SJDC^{+}13]. hydrophila [AA10, BLFRCQ10, BE19, ÇOÇ^{+}16, CTK^{+}10, DBS12, EDP^{+}17, JAR^{+}11, JSAH^{+}18, KEY^{+}15, LPA^{+}10, LLZW14, LCR^{+}10, LHW^{+}15b, NDA10, OSAA18, SJSA^{+}14, WMD^{+}18, WYZ^{+}19, YK12, ZCJ^{+}18]. hydroxybutyrate [HBH^{+}19, SCS^{+}12]. hymenocallidicola [ŘKH18]. Hypanus [SCZM19]. hyperinflation [BCG^{+}17]. hyperlipidaemia [CHF^{+}12]. hyperplasia [ONA10]. hyperplastic [BHW^{+}11]. hyperthermia [WXWD15]. hyphal [VDHV^{+}16]. Hyphessobrycon [Lek12]. hypochlorite [HHO14]. hypodermal [YHY^{+}16]. hypophthalmus [BAD^{+}12, CTK^{+}10, SMV^{+}14, SZZSFN^{+}14, SPV^{+}18, VST^{+}19]. hypoxic [EMA17]. Hysterothylacium [MGH^{+}14, RGMGG^{+}18]. hystrix [SBK^{+}13].

I-based [TCFáS19]. ice [DJR^{+}19]. ice-structuring [DJR^{+}19]. Iceland [GVC^{+}19, áKÁGA16]. Ichthyobacterium [MFS^{+}17]. Ichthyobodo [CKJ^{+}14, MUM^{+}19]. ichthyoenteri [BT14]. ichthyofauna [RAM^{+}18]. ichthyophoniasis [HCEH16]. Ichthyophonus [AJLC^{+}16, CPEH15, HGH^{+}16, WMF14]. Ichthyophonus-induced [AJLC^{+}16]. Ichthyophonus-infected [WMF14]. ichthyophthiriasis [AMRHM18, VGI^{+}16]. Ichthyophthirius [AJLK^{+}18, AMRHM18, BSD^{+}18a, CKvGJB17, vGJHS^{+}11, vGJKR^{+}17, XKBC12, XZZ16, YYW^{+}14]. IcHV [GM10]. IcHV-2 [GM10]. icmF [WYZ^{+}19]. ictaluri [BAD^{+}12, CTK^{+}10, GRG^{+}16, KMGK^{+}18, PRWG17, PRSK10, PMK12, SRT^{+}14, VST^{+}19, YK13]. Ictalurid [GM10]. Ictalurus [BBF^{+}15, LTC^{+}14, MP18, PRSK10, RKPG16, XKBC12, YSK13, YK13]. idella [GCWB11, LLL^{+}12, YYW^{+}14]. idellus [LXHL11]. identical [MNM^{+}18]. Identification [BGC^{+}17, CHH^{+}14, CCW^{+}15, CXW^{+}19, CHF^{+}12, DBM^{+}14, HXZ^{+}16, HHX^{+}15, LLC^{+}11, LCV^{+}17, LCC^{+}15, LLL^{+}12, LMM^{+}16, MFHHG18, PKLH19, PSBG^{+}11, SRSF10, TCS19, WCL^{+}19, WNCR12, YGL^{+}18, ZWG^{+}17, ZWL^{+}16, BLL^{+}17, dJDAVRGM^{+}18, GLW^{+}18, IGLG^{+}17, JSJR^{+}18, KPSP^{+}19, LDA^{+}19, LML^{+}13, LYL^{+}18, LFH^{+}11, OBR^{+}14, OGI^{+}18, PSCFB^{+}18, SSF^{+}17, SMC^{+}11, TCFáS19, ZSZF16, ZLX^{+}19, ZFZ^{+}11]. identifies [GSEM13, GLK^{+}16, LSA^{+}15, WDM^{+}15, WMM^{+}16]. identifying [BTN^{+}18, JSJR^{+}19]. identity [LDEH17]. iELISA [ZWG^{+}18]. IgD [AJLC^{+}16]. IgD- [AJLC^{+}16]. IgM [AJLC^{+}16, vGJHS^{+}11]. IgM-positive [AJLC^{+}16]. IgT [AJLC^{+}16, vGJHS^{+}11]. IgT- [AJLC^{+}16]. IgY [NPGA^{+}19]. IHHNV [YHY^{+}16]. IHN [HAS^{+}19]. IHNV [CGW^{+}19, HSC^{+}16, JSK^{+}10, PMH^{+}11, PFJ^{+}10]. II [BMA13, CPB^{+}15, GJVT^{+}11, LLZW14, PMH^{+}11, YYW^{+}14, ZWG^{+}18]. II-DAB [YYW^{+}14]. III [EAN^{+}18]. IIIA [SWWN^{+}16]. illegally [ARFG13]. Image [ESV^{+}19a, FSB18b, GGMG18a, GCM^{+}18a, LITCAH19a, LLL^{+}18b, MTLP^{+}18a, PWHW18a, PLC^{+}19a, QZZ^{+}18a, VSDS^{+}19b, CMEV17a, CMS^{+}17, CMEV17b]. imaging [BKD17, BWM^{+}13, FDST16, RHSM^{+}18b, RHSM^{+}18a]. imbalance [AFTEML19]. Immersion [MAC^{+}15, GOM^{+}19, JAJC^{+}18, MBS^{+}19a, MRN^{+}19, MSCH14, PBM^{+}15, SPLV^{+}15, SCR^{+}15, TTL^{+}16]. immersion-bath [SCR^{+}15]. Immune [HZX^{+}13, LHW^{+}12, TMM^{+}15, AJLC^{+}16, BSD^{+}18b, BSB^{+}18, BAMYAH19, Buc12, CLGP15, CCHL13, CKvGJB17, CCME16, CRS^{+}13, CSCCA14, DMB^{+}16, HKBD15b, HXZ^{+}16, HMW11, JAJC^{+}18, JMF^{+}18, JKH^{+}17, vGJKR^{+}17, JJS^{+}12, LMC^{+}16, LSC^{+}14, LMM^{+}15, LMM^{+}16, LHW^{+}15b, NNC^{+}18, PHW^{+}16, RSB13, RBAT^{+}17, RCGGM15, RSPA^{+}18, SWD^{+}14, XKB12, ZWL^{+}14]. immune-related [JJS^{+}12]. immune-relevant [SWD^{+}14]. Immunity [CGW^{+}19, CR12, JJV^{+}15, LKS^{+}13, SWM^{+}15]. Immunization [PLCC11, OTN^{+}12]. immunizes [NPGA^{+}19]. immunoassay [PSK^{+}13, SSHEM11]. immunoblotting [LLL^{+}12]. immunochromatographic [LXX^{+}17]. immunocytochemistry [ECP^{+}10]. immunoenzymatic [NOST^{+}16]. immunogenic [KREM14, LLC^{+}11, LLL^{+}12, NMIY14]. immunoglobulin [ESV^{+}19b, LSA^{+}15, NPGA^{+}19]. Immunohistochemical [ERÁpSB14, FBL^{+}14, YDO^{+}16, ZHFC11, AOAH13, ABT^{+}14, BDM^{+}18, CKJ^{+}14, RHSM^{+}18a, RAM^{+}17, SMS^{+}19]. immunohistochemistry [GCM^{+}18b, MBD16]. immunological [EAM^{+}13, FMT^{+}14, SPCR^{+}19, SSSP18, WCW16a, YSHG18]. Immunomodulatory [ESA^{+}10, WXWD15]. Immunophysiology [SAD^{+}12]. immunoproteomic [CRBLC^{+}17]. immunosorbent [BWZ^{+}17, NMIY14]. immunostimulants [NFA15]. Immunostimulation [IPW^{+}13]. immunostimulatory [PPI^{+}13]. Impact [AJLK^{+}18, BPK^{+}17, DD18, JCN^{+}13, JCN^{+}14, STG^{+}10, SFU^{+}13, DTF^{+}16, GMC^{+}16b, STES14, SSL18, TJA^{+}13, KRFT14]. Impacts [BA14, BGRB14]. impair [RHB^{+}15]. Impairment [BSV^{+}18]. impairs [LP17]. impedes [BMD^{+}14]. implementation [BMBS12]. implicated [BHF13]. Implications [DMK^{+}19, RBL^{+}11, SSK17]. importance [KPSP^{+}19, KNN19]. important [DS10, WLX^{+}10]. Imported [RHBMM13, GJK^{+}11, NSC^{+}15, PRWG17, RKK^{+}19, SBD11]. imports [WTM14]. impoundments [HWO^{+}17]. improved [GCM^{+}18b, SKN^{+}16, VST^{+}19]. Improvement [HAS^{+}19, TJR^{+}18]. improves [DMB^{+}19, FMM^{+}19]. in-vitro [SHS14]. Inactivated [WHR^{+}18, LRC^{+}18, MMNL14, RPMRLJ^{+}19, SGB^{+}16]. Inactivation [YkSKY13, DSA^{+}12, DAB^{+}12, HPAL10, MoSMS18, NCNB18]. inaequilabiatus [DMD^{+}17]. Incidence [NB11]. including [SFF^{+}17]. inclusion [PFC^{+}13]. Incomati [DAO19]. increase [FMH^{+}15, SJK^{+}14]. Increased [DMB^{+}19, CKTT10, FHFF15, NYRS12, OBFR12]. increases [CCHL13]. Increasing [FCJH19, USM17]. incubated [LF18]. incubation [FFP^{+}14, FHFF15, SK15]. independent [Sla14]. India [GJM^{+}15, GMP^{+}11, HMV^{+}17, RDN^{+}12, SPP^{+}18, SSV^{+}17, SM14, TSK^{+}19]. Indian [BNC^{+}11, Góm12, GMP^{+}11, JGI^{+}10, NRMS14, SVM^{+}18]. indica [AA10]. indicate [CKJ^{+}14, KZJ^{+}17]. indicates [MMBS10]. Indications [JNS^{+}19]. indicators [PSZ^{+}14]. indices [KPM^{+}11, LFP^{+}11a, LFP^{+}11b]. indicus [NRMS14]. Indies [HHM^{+}14]. indirect [ZWG^{+}18]. individual [DIVC16, HWM^{+}13]. Indo [CDM^{+}13]. Indo-Pacific [CDM^{+}13]. Indonesia [SMC^{+}11]. indoor [SBK^{+}15]. induce [CZB^{+}11, LPA^{+}10, OBFR12, SWM^{+}15]. induced [ACN12, AJLC^{+}16, BRMG13, CGW^{+}19, DAO19, FLCC^{+}14, Góm12, GHB^{+}11, JCN^{+}13, JCN^{+}14, KRFT14, MMD^{+}12, MLEM16, MGF17, XQS^{+}15, ZZQ^{+}19]. induces [CLGP15, CXW^{+}19, ESV^{+}19b, LF16, SON^{+}18]. Induction [DJQ^{+}12, HY16, FSV^{+}18, MMS^{+}16, WBF^{+}10]. industry [VJWC^{+}18]. infect [PST^{+}16]. infected [AOAH13, BSV^{+}18, BSSB18, BSD^{+}18b, BSD^{+}18c, BSB^{+}18, BSD^{+}18a, BLFRCQ10, BRB^{+}12, Bin14, BCJ13, BTC^{+}17, CLC^{+}17, DCG18, DPADO16, FCM^{+}19, FFLC17, GIW^{+}11, GCM^{+}18b, GMN12, HDWM15, HGB^{+}11, HZX^{+}13, JML19, JJV^{+}15, KSM^{+}17, KAST17, LRD^{+}19, LMM^{+}15, LMM^{+}16, MKI^{+}15, MKMY10, MTA^{+}15, MTAB15, NB10, NHM^{+}10, PB18, PRSK10, PJU^{+}13, SP15, SWD^{+}14, SFL^{+}16, TJB^{+}15, UMMS17, VTM^{+}14, WMF14, ZWY^{+}11, ZPM^{+}18, SSV^{+}17]. infecting [BLL^{+}17, BNS13, FDL^{+}17, IBFL19, LQL^{+}10, LZGZ16, LJHG14, OMO^{+}15, RPGCIG17, SAS^{+}19, SGQ^{+}15, TB12, USM17, WMD^{+}18, WCW^{+}16b]. Infection [ARFG13, BKD17, BLP10, CGAS19, DDBC^{+}15, NNCLGC^{+}11, ZWLG10, ASA10, AMAR18, AHHR^{+}10, AHPR^{+}16, ATP^{+}16, AZSI^{+}12, ABT^{+}14, AA10, BSPS^{+}18, BMD^{+}14, BBF^{+}15, BLdA^{+}18, BAMYAH19, BE19, BHO^{+}19, BFFA^{+}12, BDB^{+}18, BVG^{+}10, BVT^{+}18, CBLN13, CDM^{+}13, CBC^{+}13, CHWM^{+}19, CAJD^{+}18, CKvGJB17, CG11, CB13, CDD^{+}18, CCM^{+}11, CCMA^{+}17, DBS12, DIH^{+}16, DLP^{+}13, DMB^{+}16, DMD^{+}17, DJR^{+}19, DCZ^{+}17, EDP^{+}17, ESA^{+}10, ECP^{+}10, FBAS16, FSV^{+}14, FŠK^{+}19, FBVC16, FCC^{+}13, FSO^{+}10, FCS^{+}18, GSEM13, GRP^{+}15, GDN^{+}16, HKK^{+}11a, HPY^{+}13, ICVV16, JKL^{+}13, JHW17, JFWH19, JMF^{+}18, JKH^{+}17, KKK15, KSS^{+}15, KRD^{+}17, KPK^{+}18, KHC^{+}18, KLD^{+}18, KKOY19, KBE^{+}14, KJW^{+}17, KDA^{+}18, KKKA17, KDP^{+}17, LGK^{+}14, LLG^{+}15, LP17, LMG^{+}19, LC17, LCOZ10, LL12, LLH^{+}14b, LXL^{+}18, LGJ19, LCR^{+}10, LSA^{+}15, LYKN18, LHW^{+}15b, LXZ^{+}19, MBOC19, MSC^{+}17, MLRS^{+}17, MFES18, MSR^{+}10, MKH^{+}11]. infection [MUM^{+}19, MTA^{+}15, MCW^{+}18, MKKM15, NOV^{+}11, NYRS12, OJRL10, PBC15, PRWG17, PFP^{+}13, PYSS19, PPI^{+}13, PPF^{+}12, PG16, PMK12, PFC^{+}13, PMPG^{+}16, QWTL17, QZZ^{+}18b, RHS^{+}12, ŘKH17, RBM^{+}16, RFA^{+}14, RSR^{+}18, RPM^{+}19, SHKF16, SZM^{+}15, SBP^{+}19, SSV^{+}10, SGB^{+}16, SHJ^{+}12, SDJH16, SPLV^{+}15, SS17, SSJ^{+}19, SNM^{+}16, SAD^{+}12, SWWN^{+}16, SODB16, SOA^{+}18, SVY^{+}14, SON^{+}18, SRT^{+}14, SNK^{+}19, TIK^{+}14a, TTL^{+}16, TDS^{+}19, VDMJ^{+}13, VBL^{+}15, VBL^{+}16, WBF^{+}10, WLX^{+}10, WKW^{+}11, WIC^{+}19, Wen16, WFH^{+}18, WXWD15, XJL^{+}19, YFA^{+}17, YDO^{+}16, YK13, YYW^{+}14, YLX^{+}19b, ZNHC11a, ZNHC11b, ZKL^{+}19, ZLN^{+}17]. Infections [LHW^{+}15a, STF^{+}15, SBM^{+}11, BFJ11, CMH^{+}17, CLS^{+}15, DS10, DYML^{+}18, FYO^{+}11, GFW^{+}10, GRF14, HCL^{+}14, JQK14, JZK^{+}14, JJSJ15, KFK^{+}16, LML12, LST^{+}12, LPCC12, LLL^{+}18a, NMIY14, RAM^{+}19, RSH11, SWS^{+}16, SVV^{+}17, TSK^{+}19, UMMS17, WMF14, WBD10a, WNAJ^{+}16, ZHFC11, PTM^{+}19]. Infectious [ADF14, BAH^{+}18, CCCTA^{+}14, MNM^{+}18, MMNL14, RWTB17, TVK^{+}19, ADMF15, ATM^{+}10, ASHZ^{+}12, AOAH13, BMK17, CGW^{+}19, DSJA18, EYF^{+}19, ECP^{+}10, EBM^{+}18, EEMCEGSP16, FE11, GL18, GTT^{+}17, GCS^{+}18, HSC^{+}16, HWM^{+}13, HMV^{+}17, IOO15, JDRM17, JOF^{+}19, JSK^{+}10, JMS13, JJSJ15, KDA^{+}18, KBD13, KWD^{+}16, LTG^{+}18, LTM^{+}19, LLLG18, LASRA^{+}19, LGJ19, LJGRI^{+}10, LVADB17, LRMN17, MBOC19, MMWK19, MAC^{+}15, MR11, MTHB13, MNM^{+}19, MMH10, NB10, OBFR12, PMW^{+}10, PMH^{+}11, PBM^{+}15, RMSH09, SJDC^{+}13, SWD^{+}14, SG12, SSO^{+}14, TIK^{+}14a, VDMJ^{+}13, VCOD17, YLSW14, YHY^{+}16]. infective [FSB18a, GGGM14]. infective-stage [FSB18a]. Infectivity [CGCAH12, SCR^{+}15, FE11, GTH11, OKS^{+}13, SG12]. infects [RAVC18, VSDS^{+}19a]. inference [RRC^{+}17]. inferred [FMA^{+}19, MFS^{+}17, STIG17]. infestation [FDST16, GAVC16, HSHC^{+}16, MGM^{+}18, SFU^{+}13, VMBGE17]. infestations [GCGR13, HLHS11, JCN^{+}13, JCN^{+}14, KRFT14, PBBG19, SMR10, TAQ^{+}10]. infesting [YB17]. inflammation [BKG^{+}18, GRT12, MLF^{+}10, NHM^{+}10, WHR^{+}18]. inflammatory [BSB^{+}18, BSD^{+}18a, BSD^{+}18d, ESV^{+}19b, GIW^{+}11, MDF^{+}17, RCGGM15, SJSA^{+}14]. Influence [HH13a, JSAB^{+}19, SLPB18, SSL18, WBF^{+}10, BSD^{+}18a, DIVC16, IBFL19, MPS^{+}18]. influenced [BSPS^{+}18]. influences [CRBLC^{+}17, JMF^{+}18]. influencing [CHAB10, TES16]. Information [Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, AGFG19, MJ17]. informing [SMR16, SDI^{+}18]. infrared [WNCR12]. Ingestion [SNM^{+}16]. inhabiting [FDL^{+}17]. Inhibition [GSBEM15, HMP^{+}12, BSD^{+}18c, DRG^{+}17, HLP^{+}16]. inhibitors [AS17, ASS19, UKS^{+}18]. Inhibitory [SBW11, BLFC12, PSBG^{+}11, YVM^{+}14, YLX^{+}19b]. inhibits [BSV^{+}18, TKS^{+}18]. iniae [BYAS^{+}19, CZL^{+}19, EPA^{+}10, HJOK11, KNN19, OGI^{+}18, SLKE10, WZL13, ZFZ^{+}11]. initial [SZM^{+}15]. injectable [HJN^{+}19]. injection [CMH^{+}17, ELI^{+}17, GSS^{+}10, GMC^{+}16b, MID^{+}19b, PYSS19, RPMRLJ^{+}19, SPLV^{+}15]. injury [FBK^{+}16]. inkyongense [dASLS^{+}19]. innate [JKH^{+}17, LKS^{+}13, STS^{+}15, YSHG18]. innesi [SMS^{+}19]. inoculation [SDR^{+}15]. inorganic [MMNL14]. inornatus [SBW^{+}18]. insight [dALF^{+}17]. Insights [COB^{+}16, SP15]. integrate [JMD^{+}18]. integrity [FMM^{+}19, FE11]. integron [NB11]. intensity [BND^{+}17, BWP^{+}19b]. Intensive [TAQ^{+}10, JR19, LLH^{+}14b, STS^{+}15]. intensively [MP18, SKPS16, SMS^{+}18]. interacting [SRSF10]. interaction [GIW^{+}11, IBFL19, MKI^{+}15, RGD^{+}12, SP15, WNMK^{+}16]. Interactions [LMC^{+}16, GAEK^{+}14, JFWH19, Tre12]. interactome [FMA^{+}19]. interface [BCJ13]. interferes [VDH19]. interferons [NSE^{+}17]. intergenic [TAC15]. interleukin [CKTT10]. interleukin-1 [CKTT10]. internal [DPMK12, LCZ^{+}18]. internalization [TDT^{+}16]. internalize [LASRA^{+}19]. interpretation [GVB^{+}17a]. intestinal [BGD^{+}17, BWS^{+}18, DPC^{+}10, DMB^{+}16, DAO19, ERÁpSB14, JSAH^{+}18, LFL19, MBB12, MKK^{+}11, SJSA^{+}14, VMBGIC19]. intestine [BKG^{+}18, BDM^{+}18]. intestines [YkSKY13]. Intimal [DRdS^{+}10]. intra [GSS^{+}10]. intra-peritoneal [GSS^{+}10]. intracellular [AHAMR^{+}19, MB17, YVM^{+}14]. intracoelomic [PYSS19]. Intragastric [PYSS19]. Intragenomic [LWWS14]. Intramuscular [ESV^{+}19b, ELI^{+}17, FGJ^{+}10]. intranuclear [ACSB^{+}19, XLZ^{+}17]. intraperitoneal [ELI^{+}17, LWJ^{+}16, MID^{+}19b, MID^{+}19a, RPMRLJ^{+}19, SPLV^{+}15]. Intraspecific [BAD^{+}12]. Introduced [BMD^{+}14]. intubation [JSAH^{+}18, SJSA^{+}14]. Invasion [RhAS^{+}15, HGW^{+}19, KVC^{+}16, VMBGIC19]. invasions [WTM14]. invasive [BND^{+}17, GCNB18, KPKP10, MKK^{+}11, RHST^{+}16, SBD11]. invertebrates [BCL^{+}19, HCCM12]. investigate [AAUHW15]. Investigating [GAEK^{+}14, JBO^{+}15]. Investigation [DWL^{+}17, DYML^{+}18, GFB^{+}19, SS17, SFF^{+}17, VCR^{+}11, YLSW14, BSW17, HZJ^{+}19, KOI^{+}14, SSJ^{+}19]. Investigations [NLC^{+}16, SVV^{+}17]. investment [PPF^{+}12]. involved [CIL^{+}17, DLP^{+}13, GSEM13, LCV^{+}17, LSC^{+}14, MBB12, PKLH19, RMS^{+}13, WCL^{+}19]. Involvement [BSD^{+}18b, BSD^{+}18c, LTC16, NB10]. involving [CDD^{+}18, DABO^{+}17, LEM^{+}19]. iodine [MoSMS18]. iodixanol [DLCC^{+}13]. IPN [CRAG^{+}12]. IPNV [BAH^{+}18, CGW^{+}19, ECP^{+}10, JJSJ15, KBD13, LJGRI^{+}10, NSE^{+}17, NB10, PFJ^{+}10]. Iran [RHZA^{+}16]. Ireland [JCN^{+}14, BW18, GFW^{+}10, GFR^{+}12, JCN^{+}13, JKO^{+}13, KRFT14, MSR^{+}10, RMSH09, SBH^{+}16, TCM^{+}19]. iridoviral [LGK^{+}14]. Iridoviridae [MNG^{+}18]. Iridovirus [WBD10a, CCLL16, GHH^{+}11, HLC^{+}15, JLJ16, KHC^{+}18, LCC^{+}15, RWTB17, SCK13, WZL^{+}11, YLX^{+}19b]. iridovirus-type [KHC^{+}18]. iridoviruses [BLL^{+}17]. Irish [GM11]. Iron [BPMI^{+}16, BLF^{+}16, BAMYAH19, CRO^{+}18, OJRL10, VMBGE17, WZL13]. iron-acquisition [BAMYAH19]. irradiation [CB13]. irritans [LMAKN16]. ISA [CRAG^{+}12]. ISAV [AGB^{+}19, GL18, MMWK19, NSE^{+}17]. Isbrücker [ADR11]. ISKNV [TIK^{+}14a]. Island [HSGD11, MMMD16]. islandica [ATP^{+}16]. islands [LCV^{+}17, Gis18]. isogeneic [NSS^{+}17]. isolate [CHH^{+}14, HO13, HHX^{+}15, IOO15, JSK^{+}10, MID^{+}19a, ZLX^{+}19]. isolated [BT14, CVLJ^{+}10, CMSHS16, COB^{+}16, CLLF14, DMP10, DLPR15, DSJA18, GJVT^{+}11, GBO^{+}18, GCMG^{+}11, HKK^{+}11b, HY16, IGLG^{+}17, KTP^{+}15, KKM^{+}19, KSAB^{+}14, LEM^{+}19, LYL^{+}18, MNM^{+}18, NMIY14, NMM^{+}18, PKPG15, PSBG^{+}11, PAZ18, PLC^{+}19b, RKK^{+}19, RVC18, SDWA18, dASLS^{+}19, SKB10, TWYC15, YkSKY13, ZZC^{+}17]. isolates [ACAS14, AAZS^{+}15, BAD^{+}12, BGC^{+}17, BBP18, BFC^{+}15, CD17, CYS17, DBV^{+}13, DLCF13, DZV10, DMK^{+}19, FSB^{+}18c, FNISE12, GPFGAG19, GRG^{+}16, HBB^{+}15, HKB^{+}16, HXZ^{+}16, JDH^{+}18, JGI^{+}10, JMS13, JSJR^{+}18, KWD^{+}16, LLLG18, MPBM^{+}16, MNM^{+}18, NSB^{+}18, PSCFB^{+}18, PRY^{+}18, RB15, RSIP^{+}17, RSPA^{+}18, RMSH09, SHO^{+}17, SLKE10, TJB^{+}15, ZWL^{+}16]. Isolation [ABF15, BLFC12, CLZ^{+}12, DKBS17, FDL^{+}17, GJM^{+}15, GFS^{+}14, JDRM17, LDA^{+}19, OBR^{+}14, PBMB11, RSM^{+}18, SSF^{+}17, SMC^{+}11, WCW^{+}16b, ZB12, ZSZF16, ZLX^{+}19, ATM^{+}10, BMMP^{+}14, DSMC17, DABO^{+}17, LPVGL^{+}10, MHNY18, MNM^{+}19, SPS^{+}17, SMV^{+}14, SWE^{+}19, VDD^{+}18, ZxYJ^{+}19, EA18]. isopods [HSGD11]. isothermal [HJOK11, JYZ^{+}14, KTN^{+}16, LLWL15, MSI^{+}15, YWW^{+}19]. Issue [Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l]. Isurus [ABM^{+}15]. Itagaki [LML12]. Italy [APP^{+}17, VPS^{+}14, BVG^{+}10, BVT^{+}18, CGP^{+}19, FCC^{+}13, FSC^{+}16, QPB^{+}16]. items [PJL^{+}12]. IV [DJR^{+}19, GVC^{+}19, MSG^{+}18]. IVb [AHHR^{+}10, AHPR^{+}16, Ano12c, GMN12, GRP^{+}15, HPAL10, KF10a, MF12, OEG^{+}13, VBL^{+}15, VBL^{+}16, WKW^{+}11]. ivermectin [ANB^{+}19].

Jackson [KRFT14]. jandaei [DTJ^{+}17]. Japan [CONN10, MHNY18, MKMY10, TIK^{+}14a, ZYW^{+}10]. Japanese [FYO^{+}11, KYK^{+}12, KPK^{+}18, OMO^{+}15, WCW^{+}16b]. japonensis [SBD11]. japonica [KYK^{+}12, KPK^{+}18, KLD^{+}18, OMO^{+}15]. japonicus [ESA^{+}10, LQL^{+}10, MTK^{+}10, XJL^{+}19]. Jasus [ZJDL18]. Jaundice [OAO^{+}12, MFS^{+}17, GMC^{+}16a]. jaw [ACA^{+}16, WBE^{+}17]. JEECV [OMO^{+}15]. jellyfish [MLMR16, PÅD18]. jirovecii [AMAR18]. Jordan [DBHC17, FYO^{+}11, WWS^{+}11]. jordani [GM11]. juvenile [ASW16, AAV^{+}16, BCT^{+}17, BBP18, HSGD11, HKD^{+}10, KF10b, KPM^{+}11, LKS^{+}13, MUM^{+}19, MPM16, PST^{+}16, PSS17, STF^{+}15, SSF^{+}17, XLZ^{+}17, VCLW16]. juveniles [EEMCEGSP16, OP14, PBM^{+}15, RBWR^{+}15, SOA^{+}18].

Kaolinitic [BBF^{+}15]. Kaup [CVLJ^{+}10, CTD^{+}12, CPP14, CRS^{+}13, CSCCA14, FBL^{+}14, dALF^{+}17]. Kellogg [TNK12, WBF^{+}10]. Kempkes [ADR11]. Kenya [MNM^{+}18, MNM^{+}19]. keta [KBE^{+}14]. key [KFK^{+}16, LXZ^{+}19, WMM17, ZZQ^{+}19]. Khawia [SBD11]. KHV [BSK^{+}10, BWZ^{+}17, KKOY19, MTA^{+}15, PFP^{+}13, RBM^{+}16, SMC^{+}11, ZWL^{+}14]. KHV-CJ [ZWL^{+}14]. KHVD [BSK^{+}10]. kidney [BSW17, BLdA^{+}18, BSSH17, BLP10, BSBH10, CHAB10, CSCCA14, GHH^{+}11, KSM^{+}17, Lek12, PRSK10, RhMM^{+}15, RWTB17, SWD^{+}14, SSO^{+}14, TIK^{+}14a, WWS^{+}11, XWP^{+}18]. kidney-derived [WWS^{+}11]. kill [MP18]. killer [BRB^{+}12, LSC^{+}14]. killing [MWRSK17, TDLT^{+}16]. kills [LDEH17]. kinase [FMA^{+}19, HCW^{+}16, PMH^{+}11]. kind [BSPS^{+}18]. Kinetic [FGJ^{+}10, TDT^{+}16]. kinetics [SS17, TNK12]. Kingdom [NSB^{+}18]. kingfish [NWMK16, STIG17]. Kinkelin [HMDB13]. kisutch [AHTCA^{+}17]. Klamath [JMHK18]. Kner [EDSL11]. knowledge [GFS^{+}18, JJM^{+}17, KSS17, LRMN17, MYWC18, ZGY^{+}18]. known [MS16, VLVS^{+}16]. Koi [CCT^{+}11, PTM^{+}19, TDJ^{+}10, ATJS^{+}18, BSK^{+}10, BMK17, CLC^{+}17, CCW^{+}15, CLHGK17, CDD^{+}18, EYF^{+}19, FBS13, FSV^{+}14, GJM^{+}15, HDC18, KKOY19, MTA^{+}15, MTAB15, RHZA^{+}16, RBH^{+}17, SGB^{+}16, SEM18, SMC^{+}11, TWJP10, WZW^{+}18, YkSKY13]. koi-herpesvirus [YkSKY13]. Korea [HKK^{+}11a, HKK^{+}11b, HPY^{+}13, KYK^{+}12, KLD^{+}18, RKK^{+}19, SDJH16]. Korean [KOS^{+}10]. koreensis [SM14]. Krabbe [BK12, MGH^{+}14]. Krkosek [JCN^{+}14]. Krøyer [AHH16, DLLB^{+}18, IPW^{+}13, JR19, PJZB19]. Kruger [HGPD11, HOHG13]. kuda [DCV^{+}14, LYD^{+}15]. Kudo [RKPG16]. Kudoa [ATP^{+}16, CB13, MSBB^{+}16, SS17, TOK^{+}16]. kuruma [ESA^{+}10, MTK^{+}10]. Kuwahara [LML12]. kynurenic [KSM^{+}17].

L [ASHZ^{+}12, JCN^{+}14, ZWLG10, LFD^{+}17, ZCJ^{+}18]. L-aspartate [ZCJ^{+}18].

L-Lactic-Co-Glycolic [ASHZ^{+}12]. L-rhamnose [LFD^{+}17]. L. [ACN12, ATP^{+}16, AGFG19, AZSSZZ13, ACA^{+}16, AAV^{+}16, ARFG13, ASSP^{+}15, Ano10c, AI11, ACM^{+}14, ATP^{+}17, AdOG^{+}17, BNC^{+}11, BSWS11, BSC^{+}14, BMD^{+}14, BGRB14, BLFRCQ10, BC11, BT14, BSSH17, BMGS17, BMT^{+}15, BHA^{+}17, BNS13, BK12, BSBH10, BGCKF12, CBLN13, CBC^{+}13, CVLJ^{+}10, CVGR^{+}13, CJB^{+}15, CR12, CLL^{+}10, CCT^{+}11, CLHGK17, CKTT10, CCHL13, CCME16, CZB^{+}11, CRAG^{+}12, CLLF14, DTF^{+}16, DKG^{+}17, DLLB^{+}18, DMB^{+}19, DLP^{+}13, DdRTM13, DPCdR14, DTJ^{+}17, DRdS^{+}10, EPA^{+}10, ECP^{+}10, EWH^{+}19, ELI^{+}17, EDV^{+}18, ESV^{+}19b, EMA17, FLCC^{+}14, FSV^{+}14, FKB^{+}14, FMM^{+}19, FCC^{+}13, FSC^{+}16, FFP^{+}14, FHFF15, FMH^{+}15, GJVT^{+}11, GFO^{+}13, GFS^{+}18, GMG^{+}19, GMC^{+}16a, GJM^{+}15, GIW^{+}11, GSS^{+}10, GGH^{+}10, GCMG^{+}11, GMC^{+}16b, GSEM13, GBSF12, GFR^{+}12, GRT12, GHB^{+}11, GCGR13, GGGM14, HHO14, HLHS11, HKD^{+}10, HSA17, HST^{+}13, HCL^{+}14, HJN^{+}19, HMW11, HFKF17, HKF^{+}16]. L. [IPW^{+}13, IAR^{+}10, JCN^{+}13, JKO^{+}13, JQK14, JWO^{+}10, JTW^{+}10, JAR^{+}11, JBO^{+}15, JOF^{+}19, JNS^{+}19, JCS11, JHRW17, JHGR13, JJSJ15, KAS^{+}18, KTP^{+}15, KSA^{+}10, KPM^{+}11, KRFT14, KAST17, KBD13, KWD^{+}16, LGK^{+}14, LDÁM14, LAN^{+}14, LLH^{+}14a, LGS^{+}14, LCR^{+}10, LJGRI^{+}10, LPVGL^{+}10, LBF^{+}14, MBB12, MKI^{+}15, MIO^{+}18, MMWK19, MCWL13, MLMR16, MLRS^{+}17, MLER^{+}18, MSBB^{+}16, MLF^{+}10, MMBS10, MBS^{+}17, MRSC13, MAC^{+}15, MVWS13, MID^{+}19b, MID^{+}19a, MSS^{+}12, MSR^{+}10, MGH^{+}14, MGM^{+}16, MTA^{+}15, MTAB15, MMH10, NLF11, NJK^{+}17, NYRS12, NB10, OBR^{+}14, OSC^{+}14, ONGQ^{+}14, OCLI^{+}19, ONA10, OBD10, PAM^{+}19, PBC15, PBM^{+}15, PSZ^{+}14, PBMB11, PGF^{+}19, PHVM13, PFP^{+}13, PYSS19, PB18, PPK^{+}13, PPI^{+}13, PFJ^{+}10, PG16, PÅD18, PJU^{+}13, QXW14, RHZA^{+}16, Řeh13, RB15, RGD^{+}12, RMAM12, RBF14, RMR14, RSH11, RSIP^{+}17, RBWR^{+}15, SMR10, SMM^{+}13, SFR^{+}17]. L. [Sam10, SBD11, SFBW^{+}17, SWM^{+}15, SPKS17, SLKE10, SL15, SBH^{+}16, SFU^{+}13, SG12, SWH^{+}14, STES14, SWWN^{+}16, SWS^{+}16, SKGE13, SBK^{+}13, SKPS16, SKJ13, SSA13, SRB12, SNK^{+}19, SJSA^{+}14, TWNB^{+}12, TJB^{+}15, TCC10a, TES11, TCM^{+}19, TMC^{+}12, TNK12, Tre12, TWYC15, VDMJ^{+}13, VCN^{+}10, VTM^{+}14, VMCM14, VFB^{+}12, WDM^{+}15, WMM^{+}16, WPH^{+}13, WZW^{+}18, WIC^{+}19, WNLF^{+}12, WNSAL12, WNAF^{+}13, WNMK^{+}16, WNAJ^{+}16, WKK12, WEPB^{+}18, YSHG18, YFA^{+}17, YNL11, YMBE13, YP17, ZCCN13, ZNHC11a, ZNHC11b, ZCP12, ZWLG10, ZPM^{+}18]. labelled [KTK17, VLVS^{+}16, RGD^{+}12]. labelling [EPCSBP18]. Labeo [KREM14, SSSP18]. Laboratory [ZCCN13, CBW19, KCL14, KAESEM13, MMD16, SSK17]. labracis [RRC^{+}17]. labrax [CKTT10, FCM^{+}19, FPK^{+}17, LMC^{+}16, NOST^{+}16, PBC15, PST^{+}16, RRC^{+}17, RGD^{+}12, TMC^{+}12, TNK12, VTM^{+}14]. Labridae [MMW^{+}15, Tre12, WDM^{+}15]. Labrus [BVK^{+}16, BT14, RSP^{+}17, RSM^{+}18, SRM^{+}18]. Lacépède [AZSI^{+}12, CGAS19, CLZ^{+}12, CWL^{+}17, LYD^{+}15, MGN^{+}16, WKW13]. Lack [KSA^{+}10, SL15]. Lactic [ASHZ^{+}12, PSBG^{+}11]. lactis [WEPB^{+}18]. lactococcosis [CRBLC^{+}17, CPB^{+}15, FLC^{+}17]. Lactococcus [ABT^{+}14, CLT^{+}14, DPMK12, FLC^{+}17, HY16, PSBG^{+}11, TAC15, WEPB^{+}18]. Lagoon [YMBE13, YB17]. Lagos [AAUHW15]. Lake [BSC^{+}14, JMHK18, MKKM15, NFF^{+}17, ARR^{+}18, CMS^{+}19, CMR^{+}16, HWO^{+}17, JST^{+}19, MCW^{+}18, NMV^{+}19, PYSS19, TSS18, TRS19, WWZ^{+}18, WKW^{+}11, YTKS19, YLL^{+}18, YWW^{+}19, BGRB14, BWB^{+}17, HAP^{+}10, HWO^{+}17, MCW^{+}18]. lakes [AAUHW15, EKED^{+}17, FDL^{+}17, KF10a, KF10b, OEG^{+}13, WKW^{+}11]. lalandei [STIG17]. lalandi [KHS11, NWMK16, RBAT^{+}17, STIG17, SS17]. LamB [YZW^{+}19]. laminin [MBD16]. LAMP [HJOK11, YWW^{+}19]. Lamproglena [MAOS^{+}18]. lanceolatus [BFFA^{+}12, CBH^{+}15, DDBC^{+}15, GHW^{+}15, SSF^{+}17, ZDWH18]. Landsberg [ZYW^{+}10]. langsdorfii [KKOY19, NSS^{+}17]. large [HZJ^{+}19, SPP^{+}18, XQS^{+}15, ZZA^{+}14, ZWG^{+}17]. Largemouth [CGAS19, PBBG19, ZLX^{+}19]. Larimichthys [XQS^{+}15, ZZA^{+}14, ZWG^{+}17]. larvae [BCBS^{+}13, BRMG13, BA13, CVGR^{+}13, DMB^{+}16, DMD^{+}17, DABO^{+}17, EEMCEGSP16, HBH^{+}19, HPY^{+}13, KS19, KTK17, LMC^{+}16, MBB12, MRN^{+}19, MGF17, NRF^{+}18, NDA^{+}19, RGD^{+}12, RGMGG^{+}18, RSH11, RCD^{+}17, SPP^{+}18, SCS^{+}12, TNK12, WCL^{+}12, ZJR^{+}18]. larval [BT14, EDN18, MVM^{+}18, RRO^{+}13, RBAT^{+}17]. laser [BCM^{+}17, WNLF^{+}12]. late [CCLL16, LCC^{+}15, MLER^{+}18]. late-stage [MLER^{+}18]. latency [IFF^{+}15]. latent [ATM^{+}10, LMG^{+}19]. Lateolabrax [SDJH16, XJL^{+}19]. lateral [FPK^{+}17, WSX^{+}18]. lateral-line [FPK^{+}17]. Lates [BO15, GKCC^{+}12, JHW17, KNN19, SDM^{+}19, TGJCH15]. Latris [ABC^{+}10, CB13]. Laurentian [FDL^{+}17, KF10b]. layer [GLK^{+}16]. LE141 [GPFGAG19]. LE89 [GPFGAG19]. lead [CCHL13, GBV^{+}16, MSS^{+}12]. leads [ASW16, EDV^{+}18, MSG^{+}18, OP14]. leaves [ZTF^{+}10]. Lecanicephalidea [BCJ13]. lectin [ACYC10, PBC15]. Leech [PBBG19, KCL14]. leei [ERÁpSB14, SSJ^{+}19]. leg [TTL^{+}16]. length [NRMS14]. lentiginosus [AFA^{+}12]. Lepeophtheirus [AHH16, AH16, BRL^{+}13, BDD^{+}18, DLLB^{+}18, EDN18, FSB18a, GCGR13, HH13b, IPW^{+}13, JCN^{+}13, JCN^{+}14, JR19, JHGR13, KJB11, KRFT14, LGJ19, NRF^{+}18, NLC^{+}16, NBJ10, OWM^{+}17, PJZB19, PPI^{+}13, PFC^{+}13, SMR10, SMM^{+}13, SMR16, SFU^{+}13, SSA13, USM17, WWE^{+}13]. lepeophtherii [MRSC13]. lepidosarcomas [SMS^{+}19]. Lepidotrema [FHD13]. Lepomis [ALG15, BMA13, DWL^{+}17, LEM^{+}19, LDEH17, LFL19]. Lernaea [RHS^{+}12]. Lernaeidae [MAOS^{+}18]. Lernaeocera [BSWS11]. Lernaeopodidae [PMBM13]. lesion [HKF^{+}16]. Lesional [ONGQ^{+}14]. Lesions [BB11, AGFG19, BOH^{+}12, BMA13, CBFM16, Góm12, GHB^{+}11, LLS^{+}11, LWMT10, MPM16, PSS17, RhMM^{+}15, RVB13, RSF^{+}17, RBF14, SMJ^{+}12, SM14, SON^{+}18, VPS^{+}14, VPCM19, WNLF^{+}12]. Lesson [LYD^{+}15]. Lesueur [GPWF19, LML12, LTC^{+}14, PHR14, SCZM19]. Letaba [HGPD11]. lethal [FCM^{+}19, MTAB15, SRB12]. lethality [DIVC16]. lethostigma [WWS^{+}11]. leucocytes [CHAB10, CSCCA14]. levamisole [ANB^{+}19, OMB^{+}19]. level [GBV^{+}16, KAS^{+}18, MTLP^{+}18b, PSCFB^{+}18]. Levels [áKÁGA16, DIVC16, GDN^{+}16, HGG^{+}11, JKO^{+}13, MMH10, RHTC19b, YAHWU11]. LF [RSIP^{+}17, RSPA^{+}18]. LF-89-like [RSIP^{+}17, RSPA^{+}18]. Libyan [RBS^{+}16, RSF^{+}17]. Lice [BT13, Kan18, AH16, EWRH13, FVAVMGE18, GBV^{+}16, GVB^{+}17b, GVB^{+}17a, Gis18, GCGR13, HGH^{+}18, HH13b, JCN^{+}13, JKO^{+}13, JCN^{+}14, JMD^{+}18, JR19, JHGR13, KRFT14, LGJ19, MSMM19, MMH^{+}18, NRF^{+}18, NBJ10, PdLC19, PJU^{+}13, PFC^{+}13, SMR10, SMM^{+}13, SCF^{+}13, SDI^{+}18, SSA13, SSL18, TJA^{+}13, UMMS17, VDMJ^{+}13, WWE^{+}13, WOD16, YAHWU11, VCLW16]. lice-induced [JCN^{+}13, JCN^{+}14, KRFT14]. lids [BHW^{+}11]. Life [RRC^{+}17, AJLK^{+}18, JHGR13, LF18, LNC^{+}17]. ligands [YGL^{+}18]. light [BRL^{+}13, HPAL10, LSA^{+}15]. Ligurian [VPS^{+}14]. like [BRB^{+}12, BMH^{+}19, CCLGRM^{+}17, CBFM16, dJDAVRGM^{+}18, GMB^{+}17, GBO^{+}18, KHLL16, LLS^{+}11, MBB12, MRR^{+}14, MTG^{+}10, MID^{+}19b, NSE^{+}17, PRSK10, RAVC18, RPJ^{+}18, RSIP^{+}17, RSPA^{+}18, SHO^{+}17, SHKF16, SWD^{+}14, WCL^{+}19, ZGY^{+}18, CKJ^{+}14, MID^{+}19a, NPGA^{+}19]. Limanda [SBH^{+}16, VDD^{+}18]. Limited [PSCFB^{+}18]. Lindau [HDC18]. line [ADR11, APG^{+}11, BNC^{+}11, BAMYAH19, FPK^{+}17, GHH^{+}11, GSBEM15, GHW^{+}15, IFF^{+}15, KLW10, LJC^{+}19, MVM^{+}18, NSE^{+}17, PWHW18b, PWW19, PMW^{+}10, SVL^{+}17, SSG^{+}19, SDR^{+}15, VBL^{+}15, VKK^{+}19, WWS^{+}11, WWZ^{+}18, WZW^{+}18, WKW13, Wen16, WBD^{+}10b, XWP^{+}18, YNL11, YCNW18]. lineage [MBS^{+}17, SMC^{+}11]. lineages [WBD^{+}10b]. lineata [ABC^{+}10, CB13]. lined [BCW^{+}15, QWTL17, SBHC18]. lines [CLL^{+}10, CGCAH12, EYF^{+}19, MTHB13, NMV^{+}19, VBL^{+}16]. link [DPD^{+}18, MLF^{+}10]. linked [BWZ^{+}17, DCMM13, DBM^{+}14, FMA^{+}19, KHLL16, NMIY14, PMH^{+}11, TvWO^{+}16, YLX^{+}19a, ZLN^{+}17]. links [LC17]. Linn [ZKL^{+}19]. Linnaeus [ASW16, BAH^{+}16, FCM^{+}19, FPK^{+}17, GAVC16, HSHC^{+}16, LLZW14, LDEH17, MBD16, MSS^{+}17b, PJZB19, TDS^{+}19, WSA^{+}18]. Linnhe [SMR16, SDI^{+}18]. lion [PÅD18]. lipid [CHF^{+}12, DMB^{+}19, HOHG13]. lipidosis [DBHC17, SKPS16]. Lipoid [DBHC17]. lipomatosis [RHST^{+}16]. lipopolysaccharide [JBGP11]. Liposarcoma [RHST^{+}16, SBK^{+}15]. liposome [MTK^{+}10]. liposome-coated [MTK^{+}10]. Liquid [BCW^{+}15, SHV^{+}16]. listed [HAS^{+}19]. Listonella [RRO^{+}13]. Litopenaeus [HXZ^{+}16, BTN^{+}18, CLW^{+}13, HMV^{+}17, JQK14, KS19, RPGCIG17, SSV^{+}17, SVM^{+}18, TMM^{+}15, TDLT^{+}16, TDT^{+}16]. Little [GB18, MS16]. little-known [MS16]. litulon [FYO^{+}11]. live [CBW19, CBM^{+}18, GWM12, MBS^{+}19a, TWJP10, WKK12]. live-attenuated [MBS^{+}19a]. Liver [DMD^{+}17, NB10, PSZ^{+}14, CMS^{+}19, DBHC17, FBK^{+}16, JCS11, KSM^{+}17, RhMV^{+}15, RMAM12, TMWV14]. living [OMO^{+}15]. liza [PMBM13]. Lizard [HSGD11]. lizardfish [KSAB^{+}14]. loach [ZWL^{+}16]. load [DD18, OBFR12, USM17, VCN^{+}10]. loads [PdLC19]. lobster [CG11, SQC^{+}14, ZJDL18]. local [ASB^{+}12, ESV^{+}19b]. localization [ACYC10, PSS17]. located [HGW^{+}19]. location [BNS13]. locations [JCN^{+}13, JCN^{+}14, KRFT14]. Loch [SMR16, SDI^{+}18]. loci [PdLC19]. locus [MSG^{+}18]. Lom [ZYW^{+}10]. Loma [MVM^{+}18, PG16]. Long [DIVC16, MSBB^{+}16, KHG^{+}17]. Long-term [DIVC16, MSBB^{+}16, KHG^{+}17]. longfin [ABM^{+}15, MWS17, WTM14]. longiceps [GMP^{+}11]. longitudinal [GFW^{+}10, JWO^{+}10]. longsnout [DSMC17, LKT^{+}13]. loop [HJOK11, JYZ^{+}14, KTN^{+}16, LLWL15, MSI^{+}15]. loop-mediated [HJOK11, JYZ^{+}14, KTN^{+}16, LLWL15, MSI^{+}15]. loopmediated [YWW^{+}19]. Lophiiformes [LZGZ16]. Lophius [FYO^{+}11]. lordosis [LGS^{+}14]. lords [GM11]. Loskop [HWO^{+}17]. loss [TJB^{+}15]. Lost [JMHK18]. Lota [PFJ^{+}10]. Lough [MFES18]. louse [AHH16, BRL^{+}13, BDD^{+}18, EDN18, FSB18a, HBS^{+}14, KJB11, NLC^{+}16, OWM^{+}17, SFU^{+}13, USM17]. Low [HHO14, HZP^{+}12, JLJ16, AAZS^{+}15, BLF^{+}16, EDV^{+}18, GAVC16, HGG^{+}11, LRS^{+}17, LZZN10, MMBS10, MAC^{+}15, MTHB13, MSK^{+}12, SVL^{+}17, WZA10]. low-frequency [LRS^{+}17]. low-speed [MTHB13]. low-virulence [BLF^{+}16]. low-virulent [WZA10]. Lowe [KCD^{+}14, RBS^{+}16, RSF^{+}17]. Lower [HGPD11, ACA^{+}16, OP14]. LPS [UKS^{+}18, ZZQ^{+}19]. LPS-induced [ZZQ^{+}19]. LR [FBK^{+}16]. LS [DJR^{+}19]. LS-12 [DJR^{+}19]. lucioperca [BAH^{+}16]. lucius [Řeh13]. lugubris [PBBG19]. Luminal [RRO^{+}13]. luminescent [HBVB16]. Lumpfish [TDS^{+}19, ATP^{+}16, CCH^{+}19, ESBE18, ELI^{+}17, EDV^{+}18, ESV^{+}19b, GVC^{+}19, JHRW17, MLRS^{+}17, PAM^{+}19, PB18, RVC18, SBP^{+}19, SFBW^{+}17]. lumpsucker [AGR^{+}16]. lumpsuckers [EWH^{+}19]. lumpus [ATP^{+}16, AGR^{+}16, CCH^{+}19, ESBE18, EWH^{+}19, ELI^{+}17, EDV^{+}18, ESV^{+}19b, GVC^{+}19, JHRW17, MLRS^{+}17, PAM^{+}19, PB18, SBP^{+}19, SFBW^{+}17, TDS^{+}19]. lungfish [RSR^{+}18, SFF^{+}17]. lupin [AA10]. Lupinus [AA10]. luteus [dJDAVRGM^{+}18]. Lutjanus [CLWJ13, HCW^{+}16, ZWY^{+}11]. luxatus [JMHK18]. lymphocystis [CVGR^{+}13, CCME16, HHX^{+}15, JKC^{+}11, PHvdH12]. Lymphocystivirus [CVLJ^{+}10]. lymphocyte [HCW^{+}16]. Lymphoid [JJS^{+}12, SCZM19]. lymphoma [KVC^{+}16]. Lymphosarcoma [MCWL13]. lysogenic [HY16]. lysozyme [LSC^{+}14].

M [ESV^{+}19b, PMW^{+}10]. M. [FMT^{+}14, JCN^{+}14, PBW^{+}18, ZYW^{+}10]. macarellus [BHW^{+}11]. Maccullochella [SJT14]. mackerel [BHW^{+}11, ONGQ^{+}14, RGMGG^{+}18]. maclovinus [GOM^{+}19]. Macquaria [HBB14, JFWH19]. Macrobrachium [FMT^{+}14, HO13, LYKN18, RSGB^{+}14, SPP^{+}18, WCW16a, ZSZF16]. macrocephalus [SRT^{+}14]. macrochirus [BMA13, DWL^{+}17, LEM^{+}19, LDEH17, LFL19]. macroglobulin [LSA^{+}15]. macrophages [DRG^{+}17, UKS^{+}18, WYZ^{+}19]. macropinocytosis [LASRA^{+}19]. macropomum [OMB^{+}19, ANB^{+}19, FCS^{+}18, PdSG^{+}12, VDO^{+}18b, VDO^{+}18a]. macroscopic [LWMT10]. macrostomum [BCM^{+}17]. macrostomus [IOO15]. maculata [CLZ^{+}12, LGC^{+}19]. maculatus [SDJH16]. Magath [RGMGG^{+}18]. magellanicus [BRL^{+}13]. mahseer [SM14]. main [RMS^{+}13]. Maine [PJZB19]. mainland [CWW^{+}18, ZCC18]. maintained [TIK^{+}14a]. maintenance [KAESEM13]. major [CPB^{+}15, SSO^{+}14, PdLC19, KLW10]. make [DPADO16]. mako [ABM^{+}15]. Malacosporea [MJ17]. Malaysia [ARR^{+}18, AAZS^{+}15, AZSI^{+}12, SZZSFN^{+}14, SSO^{+}14]. MALDI [PSVK^{+}18, PSCFB^{+}18]. MALDI-TOF [PSVK^{+}18, PSCFB^{+}18]. male [CXK16, DBS12, RhAS^{+}15, VMCM14, ZDWH18]. malformation [WBE^{+}17]. Malformations [MGM^{+}16]. malformed [RHVG^{+}19b]. Malmberg [HHO14, WKK12]. Maltoporin [YZW^{+}19]. mammalian [CGCAH12]. Man [FMT^{+}14, RSGB^{+}14]. management [ALG15, BPK^{+}17, CHF^{+}12, JMD^{+}18, MSCH14, SMR16, SFF^{+}17]. mandarin [KHC^{+}18, SWD^{+}14, TIK^{+}14a, TQW^{+}14]. Manday [ABM^{+}15]. mandelii [LDEH17]. mane [PÅD18]. Mangifera [AA10]. mango [AA10]. manifestations [JJB15]. manifesting [AAZS^{+}15]. Manila [ACYC10, FBVC16]. mannan [TMC^{+}12]. marbled [LSC^{+}14, LSA^{+}15, PWHW18b, PWW19, WZL^{+}11, Wen16, WCW^{+}16b]. marcescens [CBC^{+}13]. marginatus [ALG15, KCD^{+}14, PMP^{+}18, PDdC^{+}14, RBS^{+}16, RSF^{+}17]. mariculture [YLX^{+}19a]. marine [ADMF15, ACYC10, BWNS^{+}11, GFB^{+}13, GFM^{+}11, GFR^{+}12, HSA17, HY16, JCN^{+}13, JCN^{+}14, JJB15, KPT^{+}11, KRFT14, KCL14, LTC16, LTC17, LHW^{+}15a, MR11, RCD^{+}17, SFU^{+}13, WDM^{+}15]. marinum [APP^{+}17, BVG^{+}10, CBW19, CSSWK10, PBW^{+}18, SMF^{+}10, Sla14, SMJ^{+}14]. marisflavi [LQL^{+}10]. maritimum [BMH^{+}19, CSCCA14, FSG^{+}12, FSV^{+}18, FSB^{+}18c, MMS^{+}16, VFB^{+}12]. Mark [GRB^{+}17, GMB^{+}17, CSS^{+}16, MTG^{+}10]. markedly [LF18]. marker [DCMM13, DBM^{+}14, MNG^{+}18]. markers [TDH^{+}10, ZFZ^{+}11]. marmorata [PWHW18b, WCW^{+}16b]. marmoratus [WZL^{+}11]. marron [ABF15]. Marsupenaeus [ESA^{+}10, MTK^{+}10]. Marteiliosis [GKO^{+}14]. Masenia [DAO19]. masou [IOO15]. masquinongy [Ano12c, KF10b, MF12]. Mass [XQS^{+}15, AHB^{+}17, BRMG13, FSC^{+}16, GJM^{+}15, LLC^{+}11, LXL^{+}18, MMG^{+}11, RHZA^{+}16]. masses [VPCM19]. materials [HH13a]. matters [PPK^{+}13]. mauve [MLMR16]. Mawsonascaris [SDZN19]. maximus [HMW11, JCS11, LBF^{+}14, MBB12, QXW14, SODB16, WM19, dOAPLá^{+}17]. may [LC17, MBB12, PMH^{+}11, WLX^{+}10]. Mayer [FŠK^{+}19]. MBV [RDN^{+}12]. McCulloch [KRD^{+}17, LLH^{+}14b]. MCP [MNG^{+}18]. MCP43 [FMT^{+}14]. MCsialec [ACYC10]. MDA5 [NSE^{+}17]. meagre [EKS^{+}13]. meal [MTLP^{+}18b]. means [CJB^{+}15]. measure [RHS^{+}12]. measured [JBO^{+}15]. measures [NWMK16, TJR^{+}18]. mechanism [WYZ^{+}19, YLX^{+}19b]. Mechanisms [ZCQF15, BAMYAH19, DLP^{+}13, JBO^{+}15]. Medaka [KTK17]. media [MB17, RSF^{+}17, YSV^{+}13]. mediate [LC17]. mediated [HJOK11, JOF^{+}19, JYZ^{+}14, KTN^{+}16, LLWL15, MSI^{+}15, XXL^{+}19]. medicated [TNK12, dOAPLá^{+}17]. medicinal [AHH16]. Mediterranean [EKS^{+}13, GFB^{+}13, PMP^{+}18, RGMGG^{+}18]. medium [CLW^{+}11, CLT^{+}14, KAESEM13]. megalocytivirus [JKH^{+}17, JJV^{+}15, KHC^{+}18, MNG^{+}18, NSC^{+}15]. Megalocytivirus-infected [JJV^{+}15]. megalocytiviruses [GJK^{+}11]. Megalops [BCW^{+}15]. Mekong [BAD^{+}12]. Melaleuca [BSD^{+}18c, VGI^{+}16]. melamine [PNBK13, PCK^{+}15]. melanin [LAN^{+}14]. Melanized [BHO^{+}19]. Melanogrammus [JHM^{+}18]. melanoma [RhAS^{+}15]. melanomacrophage [PSZ^{+}14]. melanops [PSK^{+}12]. melanopterus [PC17, RTM^{+}15]. melanotic [RVB13]. melapterus [HSGD11]. melas [BPF^{+}18]. Melh [YMBE13]. melleni [RBAT^{+}17]. melops [SRM^{+}18]. melting [TCFáS19]. member [UN10]. members [GAC10]. membrane [CLWJ13, HGW^{+}19, JSAB^{+}19, KREM14, NNC^{+}18, YHY^{+}16]. membrane-denitrification [JSAB^{+}19]. membranes [AI11]. membras [GJVT^{+}11]. Memoriam [EW17]. mercenariae [RAVC18]. mercenariae-like [RAVC18]. mercury [VMCM14]. merguiensis [OLC11, OLC11]. mesopotamicus [CMSHS16, VGI^{+}16]. metabolic [OP14, TvWO^{+}16, UKS^{+}18]. metabolite [FHZ^{+}16]. Metabolomic [KOI^{+}14, HZJ^{+}19, LRMN17]. metabolomics [ZCJ^{+}18]. Metabronema [SBK^{+}13]. metacercariae [FŠK^{+}19, MGM^{+}16]. Metanophrys [SPP^{+}18]. Metastatic [PC17]. metazoan [MS16]. Methacarn [FMM^{+}19]. methaemoglobinemia [MMD^{+}12]. method [KKT^{+}18a, KKKA17, MWRSK17, Mor12, PSR^{+}11, VLVS^{+}16]. methodology [PPK^{+}13]. Methods [JSJR^{+}18, AGB^{+}19, CBLN13, CMS^{+}17, EA18, FMT^{+}14, FCM^{+}19, JSJR^{+}19, KJW^{+}17, MTAB15, NJK^{+}17, NCNB18, PHvdH12, VDH19, WCW16a, XZZ16, ZCCN13, ZCC^{+}15]. metritis [BA13]. Mexican [NNCLGC^{+}11]. Mexico [BB11, dJDAVRGM^{+}18, EEMCEGSP16, OGI^{+}18, OCLI^{+}19, RPGCIG17]. meyeni [CDM^{+}16]. MGB [LGC^{+}19]. MH [YYW^{+}14]. MHC [LLZW14, RBM^{+}16]. MI03 [KF10b]. Miamiensis [LPC^{+}18]. MIC [SEK^{+}16, YVM^{+}14]. Michigan [BWB^{+}17, LST^{+}12]. miconazole [SSSP18]. microalga [JDRM17]. microbe [LKS^{+}13]. microbial [JSAB^{+}19]. microbiome [DCZ^{+}17]. microbiota [JKL^{+}13, KOB^{+}17, STIG17]. Micrococcus [dJDAVRGM^{+}18]. Microcotylidae [TAQ^{+}10]. microcystin [FBK^{+}16]. microcystin-LR [FBK^{+}16]. microcystins [KPM^{+}11]. microdilution [GCW17, YVM^{+}14]. microdissection [WNLF^{+}12]. Microencapsulation [GCNB16]. microflora [BT14]. microparasites [BHT^{+}17]. Micropogonias [MPM16]. microporidian [XLZ^{+}17]. Micropterus [CGAS19, PBBG19, ZLX^{+}19]. microsatellite [DCMM13]. Microscopic [SMS^{+}19, BCM^{+}17, LWMT10]. microscopy [CGCAH12]. Microsporida [FYO^{+}11]. Microsporidia [ATP^{+}16, KSAB^{+}14, PG16]. microsporidian [ACSB^{+}19, DML^{+}16, DCL^{+}17, MVM^{+}18, MRSC13, MKH^{+}11, ZCC18]. Microsporidians [LRAP17]. microsporidiosis [XLZ^{+}17]. middle [SMS^{+}10]. Midichloriaceae [CSS^{+}16]. migrating [BHT^{+}17, JCN^{+}13, JCN^{+}14, KRFT14]. migratory [HGH^{+}18]. Mikrogeophagus [LFSZ18]. Miller [RAM^{+}17]. Milne [NRMS14, ZB12, CG11, SQC^{+}14]. mimic [LRB^{+}17]. mimicus [ZZC^{+}17]. Minchinia [RAVC18]. mineralization [PC17]. mineralized [TPD^{+}19]. minibicornis [BLP10]. Minimal [TRS19, OM10]. minimum [YVM^{+}14]. minnow [AHHR^{+}10, CXW^{+}19, JBGP11, MGN^{+}16, WBD^{+}10b]. minnows [PPF^{+}12]. minor [PJZB19]. minus [DT16]. miR [FBK^{+}16]. miR-122-5p [FBK^{+}16]. Misgurnus [ZWL^{+}16]. Mississippi [LEM^{+}19]. Mitchell [FHD13, PHvdH12, SJT14, SGQ^{+}15]. Mitchill [Ano12c, BB11, CBFM16, DGS^{+}14, GRP^{+}15, KF10b, KHLL16, MMD^{+}12, MF12, OEG^{+}13, SAD^{+}12, VBL^{+}16]. mitochondrial [CXW^{+}19, YEK^{+}17]. mitochondrial-targeting [CXW^{+}19]. mitten [DML^{+}16, DCZ^{+}17, DCL^{+}17, FGJ^{+}10, SCS^{+}12, ZB12]. Mixed [FSV^{+}14, ESA^{+}10, KBD13]. model [CAJD^{+}18, CKvGJB17, EPCSBP18, GFM^{+}11, KSAB^{+}14, MSC^{+}17, NSS^{+}17, RAM^{+}19, SCF^{+}13, SZM^{+}15, SSG^{+}19, VMBGIC19, dOAPLá^{+}17]. Modelling [SSA13, Gis18, GCGR13, RHTC19b, SDI^{+}18]. models [BVK^{+}16, Buc12, PYSS19]. modified [SLKE10]. modulated [KPM^{+}11]. modulates [FVAVMGE18, GRT12, SPCR^{+}19]. Modulation [BSB^{+}18, PBC15]. Moebius [OBR^{+}14]. Molecular [BLL^{+}17, BCM^{+}17, CSS^{+}16, DTM^{+}10, EEMCEGSP16, FMA^{+}19, HCW^{+}16, KYK^{+}12, KDA^{+}18, LYKN18, NSE^{+}17, NYRS12, PLC^{+}19b, PPE^{+}18, RMSH09, SHKF16, SPKS17, SRD^{+}16, YYF^{+}10, ZQB^{+}11, BCL^{+}19, DPTFC10, EAM^{+}13, GAC10, IAR^{+}10, LLC^{+}11, LRB^{+}17, MFS^{+}17, MFES18, OBR^{+}14, PHvdH12, PAZ18, RHZA^{+}16, RCGGM15, SCD19, STIG17, SSO^{+}14, VDH19, VGP^{+}18, YLW^{+}14, ZCC^{+}15]. molluscan [MMG^{+}11]. molluscs [VGP^{+}18]. monitored [BKD17]. Monitoring [SNK^{+}19, AGB^{+}19, EWRH13, SFL^{+}16]. monoclonal [CCLL16, HLC^{+}15, KJW^{+}17, LCR^{+}10, PSK^{+}13]. monodon [RDN^{+}12, CWW^{+}18, DCMM13, DBM^{+}14, FSO^{+}10, JJS^{+}12, MCO11, PSK^{+}13, RDN^{+}12, SMS^{+}10, STS^{+}15, SRSF10]. Monogenea [HSHC^{+}16, TAQ^{+}10, YEK^{+}17]. monogenean [FHD13, HKG^{+}15, MWS17, PHVM13, TGJCH15, VDO^{+}18b]. monogeneans [ANB^{+}19, VDO^{+}18a]. monophosphate [VSS17]. monstrosa [DRdS^{+}10]. moray [BCG^{+}17, CHF^{+}12]. morbidity [HCCM12]. Moreau [LPVGL^{+}10]. morhua [BSWS11, BMGS17, BK12, FvdMF^{+}18, GIW^{+}11, GGGM14, HLHS11, HKD^{+}10, HFKF17, MVM^{+}18, MBB12, OBD10, PG16, RBWR^{+}15, Sam10, ZCP12]. moribund [OBR^{+}14, SWE^{+}19]. moringa [CHF^{+}12]. Moritella [BGG11, ESBE18]. Morone [DBS12, SBM^{+}11]. Moronidae [RRC^{+}17]. morphogenesis [GFS^{+}14, HFT^{+}12, MBT^{+}17]. Morphological [BOH^{+}12, PMBM13, RGMGG^{+}18, TPD^{+}19, VMBGIC19, WNMK^{+}16, BC11, SDZN19]. Morphology [NHM^{+}10, YMBE13, CHWM^{+}19, KSM^{+}17, KOS^{+}10, YLW^{+}14]. morphometric [BC11]. morphometry [SMS^{+}18]. Morphopathological [VFB^{+}12]. morphotypes [DSLR16]. mortalities [AHB^{+}17, GJM^{+}15, KCD^{+}14, MBB12, NFF^{+}17, RHZA^{+}16, SBM^{+}11, VJWC^{+}18, ZSZF16]. Mortality [ABM^{+}15, CGP^{+}19, DABO^{+}17, SDM^{+}19, SKJ13, SRT^{+}14, TJB^{+}15, BPF^{+}18, BRMG13, BGC^{+}17, BPK^{+}17, BVT^{+}18, BVL^{+}18, CRAG^{+}12, DTJ^{+}17, FSC^{+}16, HCCM12, HBB14, HB12, JCN^{+}13, JQK14, JJV^{+}15, KHLL16, KAF^{+}19, KRFT14, LEM^{+}19, LXL^{+}18, LF16, MLMR16, MMG^{+}11, MMW^{+}15, NNCLGC^{+}11, PTM^{+}19, RHB^{+}15, SPP^{+}18, SKGE13, TMC^{+}12, XQS^{+}15, JCN^{+}14]. MOS [TMC^{+}12]. mosquito [APG^{+}11]. mossambica [MWS17]. mossambicus [AZSSZZ13, NMV^{+}19, TvWO^{+}16, vD14]. most [VCOD17]. motile [DSJA18, PAZ18, TALA11]. motility [DSSB19]. moulting [TTL^{+}16]. mounds [ARL^{+}17]. mouthrot [FSV^{+}18]. moved [GMN12]. movements [GWM12, TWJP10, WMM^{+}16]. Mozambique [DAO19, TvWO^{+}16, vD14]. mRNA [IPW^{+}13, RBM^{+}16, SJSA^{+}14]. MS [PSVK^{+}18, ZCJ^{+}18]. Mucoadhesive [dSCBdS^{+}16]. mucosa [VMBGIC19]. Mucosal [RBL^{+}11, JSAH^{+}18, PDLW17]. mucous [BGD^{+}17, DPC^{+}10, LDÁM14, PSR^{+}11, PPK^{+}13]. mucus [FMM^{+}19, GBSF12, JSJR^{+}19, MMS^{+}16, PJU^{+}13, SLPB18, VCN^{+}10]. mud [LQY11]. Müeller [CDM^{+}16]. Mugil [MVWS13, PMBM13]. mugilid [DCG18]. Müller [HLHS11]. mullet [PMBM13, VPS^{+}14]. mullets [APP^{+}17]. multi [LCZ^{+}18, RKK^{+}19]. multi-drug-resistant [RKK^{+}19]. multi-organ [LCZ^{+}18]. Multicentric [DD15]. multifactorial [GTT^{+}17]. multifiliis [AJLK^{+}18, AMRHM18, BSD^{+}18a, CKvGJB17, vGJHS^{+}11, vGJKR^{+}17, XKBC12, XZZ16, YYW^{+}14]. Multilocus [ACAS14, ISCF^{+}19, NMM^{+}18, RDW18, WMD^{+}18, VJWC^{+}18]. Multiple [OBFR12, TSK^{+}19, VPCM19, AH16, JQK14, ZDWH18]. Multiplex [LGS19, HRL15, LML^{+}13, LFH^{+}11, NMM^{+}18, TSK^{+}19, THW^{+}12, VLVS^{+}16]. multiplication [TKS^{+}18]. multispecies [SFF^{+}17]. multivalent [ZXW^{+}12]. Multivalvulida [TOK^{+}16]. multiyear [GM11]. Muraena [CHF^{+}12]. Murray [FHD13, SJT14]. muscle [BCBS^{+}13, BHO^{+}19, FGJ^{+}10, GRT12, LAN^{+}14, LWZ^{+}18, WHR^{+}18, ZDWH18]. muskellunge [Ano12c, KF10b, MF12]. mussel [BRL^{+}13, HBH^{+}19]. mussels [GKO^{+}14]. Mustelus [MS16]. mutagenesis [MSG^{+}18]. mutants [EAN^{+}18, GTGCM19, HBVB16, OFA11, VSS17, WBF^{+}10]. Mya [MMD16]. Mycobacteria [KZJ^{+}17, SKB10, MSK^{+}12, NHM^{+}10]. Mycobacterial [DIH^{+}16, FSV^{+}14, PSS17, ZHFC11]. Mycobacteriosis [APP^{+}17, AFA^{+}12, ACM^{+}14, CDW17, CSSWK10, EDSL11, FFLC17, LRC^{+}18, ONGQ^{+}14, SFF^{+}17, VPS^{+}14, ZCP12]. Mycobacterium [APP^{+}17, ACM^{+}14, BVG^{+}10, CBW19, CSSWK10, JK12, LRC^{+}18, PBW^{+}18, RFA^{+}14, SMF^{+}10, Sla14, SMJ^{+}14, ZNHC11a, ZNHC11b]. mycoides [OSAA18]. mycosis [ŘKH18]. myeloperoxidase [BSSB18, YK13]. Myers [LFSZ18]. mykiss [ASA10, ASHZ^{+}12, AOAH13, AHHR^{+}10, AJLC^{+}16, ABT^{+}14, AA10, AMA11, BSW17, BSPS^{+}18, BCLM^{+}19, BHA^{+}11, BE19, BA14, BLFC12, CSS^{+}16, CHAB10, CMH^{+}17, CPB^{+}15, CLS^{+}15, DS10, DPTFC10, DPTCF10, DMP10, DJR^{+}19, DdRTM13, DdRCP13, DPADO16, DSJA18, DT16, GCNB16, GTH11, GPFGAG19, GDW^{+}15, GMC^{+}16b, HKBD15a, HKBD15b, HJN^{+}19, JKL^{+}13, JDH^{+}18, JML19, vGJHS^{+}11, vGJKR^{+}17, KSM^{+}17, LPA^{+}10, LLS^{+}11, MBOC19, MNF^{+}15, MSC^{+}17, MCSM17, MLEM16, MTLP^{+}18b, NOV^{+}11, NDA10, OP14, OBFR12, OBD^{+}11, OSAA18, OLT^{+}10, PDLW17, PSBG^{+}11, PLCC11, QPB^{+}16, RSB13, ŘKH16, RCGGM15, RHB^{+}15, SHJ^{+}12, SPCR^{+}19, SCR^{+}15, SKJ13, SON^{+}18, VSS17, XKB12, YDO^{+}16]. Myliobatiformes [BB11]. myocardial [WNLF^{+}12]. myocarditis [BWNS^{+}11, JNS^{+}19, SRM^{+}18, SNK^{+}19, TB12, TCM^{+}19, WNSAL12, WNAF^{+}13, WNAJ^{+}16]. myonecrosis [HMV^{+}17, YLSW14]. myopathy [BMT^{+}15]. Mytilus [BRL^{+}13]. myxidiosis [MBD16]. Myxobolidae [LRB^{+}17, RKPG16, SGQ^{+}15]. Myxobolus [FCJH19, HAP^{+}10, ME12, MKKM15, SBW^{+}18, SGQ^{+}15, YLW^{+}14, YYF^{+}10, ZWLG10, ZYW^{+}10]. myxoma [KHS11]. myxosporea [YLW^{+}14, CB13, ERÁpSB14, JNN^{+}11, LJHG14, LXL^{+}18, RRC^{+}17, SGQ^{+}15, SSJ^{+}19, YMBE13]. myxosporean [KOI^{+}14]. Myxozoa [MJ17, ATP^{+}16, CB13, ERÁpSB14, JNN^{+}11, KAESEM13, LXL^{+}18, MSBB^{+}16, ME12, SSJ^{+}19, TOK^{+}16, YMBE13]. myxozoan [EPCSBP18, SBW^{+}18]. myxozoans [DCG18]. Myxus [MVWS13]. Myzobdella [PBBG19].

n [LJHG14, RRC^{+}17, SDZN19, YMBE13, ZWLG10, dSdAdS^{+}19]. naïve [BMK17]. namaycush [CMR^{+}16]. nanobubbles [ITKH18]. Nanoemulsion [VDO^{+}18a, VDO^{+}18b]. nanoencapsulated [BSD^{+}18c]. nanoparticles [ASHZ^{+}12, dSCBdS^{+}16, LLDL13, SSHEM11]. Nanophyetus [STF^{+}15]. National [HGPD11, HOHG13, NSC^{+}15]. native [ÇOÇ^{+}16, GOM^{+}19, MKKM15, RBR19, SLPB18]. native-subunit [ÇOÇ^{+}16]. nattectin [LSA^{+}15]. nattereri [EDSL11]. Natural [ATP^{+}17, BFFA^{+}12, BLdA^{+}18, BRB^{+}12, BHO^{+}19, CCME16, CCMA^{+}17, DMB^{+}16, EA18, GRF14, KHC^{+}18, KSA^{+}10, LSC^{+}14, MDF^{+}17, OMO^{+}15, SBH^{+}16, SMJ^{+}14]. naturally [AOAH13, BSD^{+}18a, DMB^{+}19, FFLC17, HGB^{+}11, SSV^{+}17]. nauplii [AHH16, TNK12]. NB2011 [ZWG^{+}17]. Neagh [MFES18]. nearby [MKI^{+}15]. nebulosus [PHR14]. necator [CKJ^{+}14]. necrosis [ASHZ^{+}12, AOAH13, BLdA^{+}18, BRMG13, Bin14, BPK^{+}17, BAH^{+}18, CCLGRM^{+}17, CGW^{+}19, CGG^{+}10, DD18, DML^{+}16, ECP^{+}10, EBM^{+}18, FE11, GFS^{+}14, HSC^{+}16, HKD^{+}10, JHW17, JFWH19, JOF^{+}19, JSK^{+}10, JMS13, JJSJ15, KKK^{+}18, KTN^{+}16, KYB^{+}15, KDA^{+}18, KBD13, KWD^{+}16, LASRA^{+}19, LYL^{+}18, LHW^{+}15a, LCZ^{+}18, LGJ19, LJGRI^{+}10, LVADB17, LLL^{+}13, MBOC19, MRN^{+}19, MSI^{+}15, MNM^{+}18, MNM^{+}19, MMH10, MMNL14, NPGA^{+}19, NB10, OTN^{+}12, OKS^{+}13, OOO^{+}11, PMW^{+}10, PMH^{+}11, PBM^{+}15, RWTB17, RMSH09, SJDC^{+}13, SWD^{+}14, SSO^{+}14, TIK^{+}14a, TDS^{+}19, VCOD17, XJL^{+}19, XJL^{+}15, YHY^{+}16, ZJDL18, ZLN^{+}17, ZDWH18]. necrosis-induced [BRMG13]. necrotic [DCV^{+}14]. needle [TMWV14]. negative [THW^{+}12]. negatively [BGRB14]. Nematoda [MWRSK17, MGH^{+}14, MPM16, RGMGG^{+}18, SDZN19, SBK^{+}13]. Nematode [LZGZ16, BA13, DMD^{+}17, FDST16, LFP^{+}11a, LFP^{+}11b, MWRSK17]. nematodes [Buc12]. Nemeczek [ZWLG10]. Neobenedenia [HSHC^{+}16, MSCH14, RBAT^{+}17, TGJCH15]. Neoceratodus [SFF^{+}17]. Neoechinorhynchidae [OMB^{+}19]. Neoechinorhynchus [OMB^{+}19]. neomycin [ZCJ^{+}18]. neon [SMS^{+}19]. Neoparamoeba [CBLN13, CAS^{+}18, CONN10, FGR^{+}12, LTC16, LTC17, MCCN14, NBJ10, WNO^{+}18]. Neoplasia [GMP^{+}11, SGR^{+}18, BCW^{+}15, DIVC16, MMMD16, MMD16, RTM^{+}15, SCZM19]. neoplasm [BWS^{+}18]. neoplastic [HKF^{+}16]. Neotropical [VDO^{+}18b]. Nephroblastoma [RHSM^{+}18b]. nephrocalcinosis [MVWS13]. nephrolith [LKT^{+}13]. nephrolithiasis [LKT^{+}13]. Nereididae [HVVV15]. nerka [BLP10, GMC^{+}16a, LGJ19, TVK^{+}19]. nervous [BRMG13, Bin14, CGG^{+}10, FYO^{+}11, HKD^{+}10, JHW17, KKK^{+}18, LYL^{+}18, LHW^{+}15a, MRN^{+}19, MSI^{+}15, OTN^{+}12, OKS^{+}13, TDS^{+}19, XJL^{+}19, ZLN^{+}17, JFWH19]. nested [AOAH13]. nested-PCR [AOAH13]. net [TJR^{+}18]. Netherlands [HSC^{+}16]. nettle [AA10]. network [BSV^{+}18]. networks [SVY^{+}14]. neural [FvdMF^{+}18, SSK17, WKW13]. neurobehavioural [SSK17]. neurological [SON^{+}18]. neurophila [CB13]. neurophilia [CXK16, SSK17]. neurophilus [SGQ^{+}15]. neutralization [CLC^{+}17, GRF14]. neutralizing [Ano12c, MF12]. neutrophil [JBGP11]. newborn [BQP12]. newly [AZSSZZ13]. next [DWL^{+}17, LMM^{+}16]. next-generation [DWL^{+}17, LMM^{+}16]. Nigella [MA16]. nigromaculatus [BCT^{+}17, GPWF19, SMS^{+}19]. Niimi [LML12]. Nile [ACLP^{+}19, ATP^{+}17, AdOG^{+}17, CCHL13, CLLF14, DTJ^{+}17, FLCC^{+}14, LLH^{+}14a, LCZ^{+}18, LLL^{+}18a, MCW^{+}18, NLF11, NYRS12, PYSS19, RHTC19b, RPMRLJ^{+}19, SPKS17, SLKE10, SKGE13, SPS^{+}19, WSA^{+}18, XXL^{+}19, YSHG18, ZKL^{+}19]. niloticus [AZSSZZ13, ACLP^{+}19, ATP^{+}17, AdOG^{+}17, CCHL13, CLLF14, CZL^{+}19, DTJ^{+}17, FLCC^{+}14, FKB^{+}14, KAS^{+}18, KTP^{+}15, LLH^{+}14a, LCZ^{+}18, LLX^{+}19, LLL^{+}18a, MCW^{+}18, NLF11, NYRS12, PYSS19, RHTC19b, RPMRLJ^{+}19, SPKS17, SLKE10, SL15, SKGE13, SPS^{+}19, WSA^{+}18, YSHG18, ZKL^{+}19]. nipponense [ZSZF16]. nitrate [SSSP18]. nitric [YK13]. nitrite [MMD^{+}12]. nitrite-induced [MMD^{+}12]. nitrogen [BCW^{+}15]. NKEF [BRB^{+}12]. nkomatiensis [DAO19]. NNV [OTN^{+}12, OKS^{+}13]. noatunensis [CDM^{+}13, KBTMH^{+}16, PLC^{+}19b, SKGE13, ATP^{+}17, GIW^{+}11, SPKS17]. noble [OSC^{+}14]. nocardia [KPK^{+}18, CXW^{+}19, CCM^{+}11, DSMC17, KLD^{+}18, WCL^{+}19]. nocardiosis [DSMC17, EKS^{+}13, MYWC18]. noctiluca [MLMR16]. NOD [ZGY^{+}18]. noda [MPJ^{+}15]. nodavirus [FMT^{+}14, HO13, IBFL19, LYKN18, RSGB^{+}14, WCW16a]. Nodular [QPB^{+}16, DT16, LvD18, PGF^{+}19]. Non [CSCCA14, HO13, LMM^{+}15, BMT^{+}15, CMS^{+}19, DZC^{+}16, DLCF13, DSLR16, FCM^{+}19, GCNB18, GCM^{+}18b, HRL15, HBVB16, LJGRI^{+}10, MSS^{+}17a, MTAB15, RBR19, TALA11, TDLT^{+}16, ZZC^{+}17]. non-destructive [BMT^{+}15, CMS^{+}19, LJGRI^{+}10]. Non-immune-related [LMM^{+}15]. non-invasive [GCNB18]. non-lethal [FCM^{+}19, MTAB15]. non-luminescent [HBVB16]. non-motile [TALA11]. non-native [RBR19]. non-O139 [ZZC^{+}17]. non-pathogenic [TDLT^{+}16]. Non-permissive [HO13]. non-probed [HRL15]. non-respiratory [GCM^{+}18b]. non-rhizoid [DSLR16]. non-salmonid [DLCF13]. Non-specific [CSCCA14]. non-target [MSS^{+}17a]. non-virulent [DZC^{+}16]. norepinephrine [DSSB19]. Normandy [BVL^{+}18]. North [GL18, EEMCEGSP16, RPGCIG17, RBR19, SBH^{+}16, Tre12, DKBS17, EKT^{+}11, KPKP10, KWD^{+}16, VCS^{+}19, VDD^{+}18, WMF14]. north-east [RBR19]. north-west [SBH^{+}16]. north-western [EEMCEGSP16, RPGCIG17]. Northeast [STF^{+}15]. northern [BB11, BFFA^{+}12, FSM^{+}16, JMS13, KAST17, MIO^{+}18, MMWK19, NJK^{+}17, Řeh13, DIH^{+}16, VCS^{+}19]. Northwest [GCS^{+}18]. northwestern [BOH^{+}12]. Norway [ATP^{+}16, AGR^{+}16, Ano10c, GFO^{+}13, GGMG18b, GMG^{+}19, HKD^{+}10, HST^{+}13, HJT^{+}16, JWO^{+}10, JTW^{+}10, JNN^{+}11, JMS13, KAF^{+}19, MIO^{+}18, MMWK19, NOV^{+}11, SNK^{+}19, TJB^{+}15]. Norwegian [ATM^{+}10, HBS^{+}14, JJB15, MJ17, SJK^{+}14, SVY^{+}14, SKJ13, TES16, WNAF^{+}13]. notata [OBR^{+}14]. Notropis [LRB^{+}17, MGN^{+}16, HAP^{+}10]. nov [CPP14, SPV^{+}18, VDH19]. Nova [SCD19, DSMC17]. Novel [PSR^{+}11, TALA11, BTC^{+}17, BVL^{+}18, CCW^{+}15, CAJD^{+}18, DPMK12, DLLB^{+}18, DXX^{+}15, DKBS17, FFLC17, FNISE12, GFS^{+}14, HHX^{+}15, JKS^{+}13, LDA^{+}19, MFES18, PWW19, PKLH19, PPK^{+}13, RAM^{+}18, RRO^{+}13, RSF^{+}17, SPTDEE^{+}15, VLVS^{+}16, WZL13, WZQ^{+}16, WCW^{+}16b, YSV^{+}13, YLX^{+}19a]. novemaculeata [JFWH19]. NsHLP [WCL^{+}19]. nuclear [DMD^{+}17, TIK^{+}14b]. nucleophila [ACSB^{+}19]. Nucleospora [ATP^{+}16]. Nucleotide [ZGY^{+}18, CPB^{+}15, FCL^{+}19, IBFL19]. Nucleotide-binding [ZGY^{+}18]. number [ACAS14, GLK^{+}16, HZP^{+}12, RWTB17]. numbers [HZP^{+}12, PFC^{+}13, TES16]. nutans [HDC18]. nutrient [BCLM^{+}19]. Nutrition [OT12, JHRW17]. nutritional [HWO^{+}17]. NW [DIVC16, RAVC18]. Nyeri [MNM^{+}19]. NZ [BMH^{+}19]. NZ-RLO [BMH^{+}19].

O1 [SON^{+}18, ZZC^{+}17]. O1/ [ZZC^{+}17]. O139 [ZZC^{+}17]. O2b [AHTCA^{+}17]. obese [TvWO^{+}16]. Obituary [WR10]. object [GAVC16]. oblongum [SAD^{+}12]. observation [RSR^{+}18]. observations [ADR11, BQP12, BMA13, SHW^{+}11]. observe [GTH11]. obtained [AGFG19, JMS13, OGI^{+}18]. Obtainment [PVIM18]. Occurrence [BCL^{+}19, LKT^{+}13, PGF^{+}19, RAM^{+}18, HMV^{+}17, RSB13, RSF^{+}17]. occurring [ADR11, DMB^{+}19]. Ocean [ASW16, CBM^{+}17, BOH^{+}12, CNN^{+}10, STF^{+}15, VCS^{+}19]. ocellaris [SBK^{+}15]. ocellatum [JK12, MSS^{+}17b, YB17]. ocellatus [RHSM^{+}18a]. Ochroconis [SAS^{+}19]. Ocular [EDSL11, PAM^{+}19, PSS17, CHF^{+}12]. oedema [PTM^{+}19, SEM18]. off [BA13, BMGS17, SBH^{+}16, SEK^{+}16]. officinalis [VDO^{+}18b, ZTF^{+}10]. offshore [CNN^{+}10]. Ogcocephalidae [LZGZ16]. oil [BSD^{+}18c, CLS^{+}15, ELI^{+}17, GMP^{+}11, MA16, VDO^{+}18a]. oil-adjuvanted [CLS^{+}15]. oil-based [ELI^{+}17]. oils [VGI^{+}16]. old [WGL17]. Oldewage [DAO19]. oleoresin [VDO^{+}18b]. olfactory [LP17]. Olifants [HGPD11, HOHG13, HWO^{+}17]. oligochaetes [ME12]. oligomerization [ZGY^{+}18]. oligosaccharides [TMC^{+}12]. olivaceus [HKK^{+}11b, JKC^{+}11, LKS^{+}13, LPC^{+}18, LWZ^{+}18, SSJ^{+}19]. olive [HKK^{+}11b, JKC^{+}11, LKS^{+}13, LPC^{+}18, SMD^{+}12, SSJ^{+}19]. olseni [FBVC16]. omp85 [KREM14]. OmpU [CLWJ13]. Onchorynchus [DMP10]. Oncorhynchus [ASA10, ASHZ^{+}12, AOAH13, AHHR^{+}10, AJLC^{+}16, ABT^{+}14, AHTCA^{+}17, AA10, AMA11, BSW17, BSPS^{+}18, BCLM^{+}19, BHT^{+}17, BHA^{+}11, BE19, BA14, BPPKJ10, BLP10, BMH^{+}19, BLFC12, CSS^{+}16, CHAB10, CMH^{+}17, CPB^{+}15, CLS^{+}15, DS10, DWP^{+}19, DPD^{+}18, DPTFC10, DPTCF10, DJR^{+}19, DdRTM13, DdRCP13, DPADO16, DSJA18, DT16, GMC^{+}16a, GCNB16, GBO^{+}18, GTH11, GPFGAG19, GDW^{+}15, GMC^{+}16b, HKBD15a, HKBD15b, HJN^{+}19, IOO15, JKL^{+}13, JDH^{+}18, JML19, vGJHS^{+}11, vGJKR^{+}17, KSM^{+}17, KBE^{+}14, LPA^{+}10, LLS^{+}11, LST^{+}12, LF16, LF18, LGJ19, LWMT10, MBOC19, MNF^{+}15, MSC^{+}17, MCSM17, MLEM16, MTLP^{+}18b, MKKM15, NOV^{+}11, NDA10, OP14, OBFR12, OBD^{+}11, OSAA18, OLT^{+}10, PDLW17, PSBG^{+}11, PLCC11, PMPG^{+}16, PPE^{+}18, QPB^{+}16, RSB13, ŘKH16, RCGGM15, RHB^{+}15, SMJ^{+}12, STF^{+}15, SHJ^{+}12, SPCR^{+}19, SCR^{+}15, SKJ13, SON^{+}18, TVK^{+}19, VSS17, XKB12, YDO^{+}16]. online [MPJ^{+}15]. onset [GMN12]. Ontario [KHLL16]. ontogeny [BND^{+}17, FPK^{+}17]. oomycetes [SSSP18]. oophoritis [BA13]. opacity [PAM^{+}19]. opercular [BC11, TPD^{+}19]. operon [DS10]. Ophiocephalus [LL12]. Oplegnathus [JJV^{+}15]. optimal [RHTC19b, dOAPLá^{+}17]. Optimization [CMS^{+}17]. Optimized [LGD^{+}17, SCK13]. Opuntia [BAH^{+}16]. Oral [ASHZ^{+}12, SHJ^{+}12, CMH^{+}17, DPCdR14, FHZ^{+}16, LLDL13, LXHL11, NFA15, SCK13]. orally [FHD13]. orange [FHZ^{+}16, NNC^{+}18, WCL^{+}12]. orange-spotted [FHZ^{+}16, NNC^{+}18, WCL^{+}12]. orbicularis [DTM^{+}10]. ordalii [AHAMR^{+}19]. order [CSS^{+}16]. Oregon [JMHK18]. Oreochromis [AZSSZZ13, ACLP^{+}19, ATP^{+}17, AdOG^{+}17, BMMC^{+}15, CCHL13, CLLF14, CZL^{+}19, DLPR15, DSLR16, DTJ^{+}17, FLCC^{+}14, FKB^{+}14, ICVV16, KAS^{+}18, KTP^{+}15, KBTMH^{+}16, LLH^{+}14a, LCZ^{+}18, LLX^{+}19, LLL^{+}18a, MCW^{+}18, NMV^{+}19, NLF11, NYRS12, OGI^{+}18, PYSS19, PLC^{+}19b, RHTC19b, RPMRLJ^{+}19, SPKS17, SLKE10, SL15, SKGE13, SPS^{+}19, TvWO^{+}16, VMBGIC19, WSA^{+}18, YSHG18, ZKL^{+}19, vD14]. ORF25 [ZWL^{+}14]. organ [JJS^{+}12, LCZ^{+}18, OOO^{+}11, PMBM13]. organic [DD18, JKL^{+}13, SHS14]. organism [BMH^{+}19, KHLL16, LLS^{+}11, MTG^{+}10, VDHV^{+}16]. organisms [GMB^{+}17, GBO^{+}18, RPJ^{+}18]. organs [CSS^{+}16, LCZ^{+}18, ZCCN13]. oriental [ZSZF16]. orientalis [ATP^{+}17, CDM^{+}13, KBTMH^{+}16, PLC^{+}19b, SPKS17, SKGE13]. origin [BDD^{+}18, WDM^{+}15]. originating [NJK^{+}17]. origins [CRS^{+}13, GRG^{+}16]. ornamental [CDM^{+}13, CLHGK17, GJK^{+}11, LFSZ18, NSC^{+}15, NHM^{+}10, PRY^{+}18, RHBMM13, dASLS^{+}19, SSO^{+}14, VCR^{+}11, ZZC^{+}17]. orocutaneous [DD15]. Ortholinea [RRC^{+}17]. Ortholineidae [RRC^{+}17]. orthoreovirus [GB18, GMG^{+}19, MIO^{+}18, PPE^{+}18, BHO^{+}19, VCS^{+}19, WHR^{+}18]. orthoreovirus-3 [GMG^{+}19]. Oscar [RHSM^{+}18a]. OsHV [GVLE16, GCCLRP^{+}13]. OsHV-1 [GVLE16, GCCLRP^{+}13]. osmoregulation [LTC16]. Osmoregulatory [PHVM13, LGJ19]. Osteichthyes [CZB^{+}11, PMP^{+}18]. osteosarcoma [RHSM^{+}18a]. Ostrea [BLSG^{+}16, BWP^{+}19b]. Ostreid [GVLE16, HPY^{+}13, BVL^{+}18]. other [BFFA^{+}12, NB11]. Outbreak [BCT^{+}17, CGAS19, GVC^{+}19, LLH^{+}14b, LXL^{+}18, SDJH16, XLZ^{+}17, AZSI^{+}12, ACLP^{+}19, BGRB14, BFFA^{+}12, CGP^{+}19, CCME16, CCMA^{+}17, DYML^{+}18, EMA17, GFB^{+}19, HHC^{+}18, KTP^{+}15, KAF^{+}19, MSS^{+}17b, RBWR^{+}15, SFF^{+}17, TWNB^{+}12, VPP^{+}12, VFB^{+}12, WDM^{+}15, WCL^{+}12]. Outbreaks [HKD^{+}10, KRD^{+}17, WZL^{+}11, CWL^{+}17, CRAG^{+}12, FLC^{+}17, HZP^{+}12, HOHG13, JJB15, MHNY18, RDW18, STG^{+}10, SJK^{+}14]. outcome [TDT^{+}16]. Outer [KREM14, CLWJ13, NNC^{+}18]. ova [SWS^{+}16]. Ovarian [RAM^{+}17, CZB^{+}11, SdMVS^{+}16]. ovatus [LYL^{+}18]. overexpressing [SNM^{+}16]. Overview [GL18, BW18]. ovulatory [GM11]. oxidase [BSD^{+}18c, BSD^{+}18a, BSD^{+}18d]. Oxidative [BCBS^{+}13, MLER^{+}18, BSD^{+}18a, BSD^{+}18d, PKLH19]. oxide [YK13]. oxolinic [TNK12]. Oxyeleotris [WZL^{+}11]. oxygen [MTLP^{+}18b]. oxyrinchus [MMD^{+}12, SAD^{+}12, MMD^{+}12, SAD^{+}12]. oxysporum [KDP^{+}17]. oxytetracycline [FLCC^{+}14, FGJ^{+}10, YVM^{+}14]. oyster [BLSG^{+}16, BVL^{+}18, CBM^{+}17, CBM^{+}18, CdAPdC^{+}19, GCCLRP^{+}13, HPY^{+}13, MDC^{+}18]. oysters [BVT^{+}18, CBM^{+}18, GVLE16]. ozonation [CB13]. ozone [ITKH18].

P [AFTEML19, IPW^{+}13, UN10]. P-1 [UN10]. P-glycoprotein [IPW^{+}13]. p53 [XJL^{+}19]. Pacific [STF^{+}15, BTN^{+}18, BVT^{+}18, BVL^{+}18, CDM^{+}13, CLW^{+}13, GVLE16, GGW^{+}11, GCCLRP^{+}13, GCS^{+}18, HLL^{+}12, HCEH16, HGH^{+}16, HPY^{+}13, JQK14, PPE^{+}18, RPGCIG17, WMF14]. pacinii [BT14]. pacu [CMSHS16, VGI^{+}16]. Pagellus [RVB13]. Pagrus [EPA^{+}10, KLW10, LJGRI^{+}10]. pagurus [SWH^{+}14]. pAh [EDP^{+}17]. pAh-1 [EDP^{+}17]. pair [CZL^{+}19, VCOD17]. Palaemon [SPTDEE^{+}15]. Pallas [WMF14]. pallasii [GGW^{+}11, HLL^{+}12, HCEH16]. Pampus [LJC^{+}19]. pancerii [TAQ^{+}10]. Pancreas [JWO^{+}10, AAUHW15, Ano10c, BMT^{+}15, CJB^{+}15, DMB^{+}19, GBSF12, JTW^{+}10, JJB15, JJM^{+}17, RSP^{+}17, RRG^{+}19, SJK^{+}14, SVY^{+}14, SKJ13, TWNB^{+}12]. pancreatic [AOAH13, BAH^{+}18, CGW^{+}19, ECP^{+}10, FE11, JOF^{+}19, JMS13, JJSJ15, KDA^{+}18, KBD13, LASRA^{+}19, LJGRI^{+}10, LVADB17, MNM^{+}18, MMH10, MMNL14, NB10, RMSH09, VCOD17]. pangasiae [SPV^{+}18]. Pangasianodon [BAD^{+}12, CTK^{+}10, SMV^{+}14, SPV^{+}18, VST^{+}19]. Pangasius [PRWG17, SZZSFN^{+}14]. Pansteatitis [HWO^{+}17, HOHG13, TvWO^{+}16]. pansteatitis-affected [HOHG13]. pansteatitis-suffering [TvWO^{+}16]. Paphies [RPJ^{+}18]. papillary [ONA10]. papillomatosis [DWL^{+}17]. papulosum [GSBEM15, WBD^{+}10b]. Parabrachiella [PMBM13]. Paracheirodon [LLG^{+}15, ŘKH17, SMS^{+}19]. paraffin [FBL^{+}14, FGR^{+}12]. paraffin-embedded [FBL^{+}14, FGR^{+}12]. parahaemolyticus [CCLGRM^{+}17, DD18, HXZ^{+}16, ITKH18, KTN^{+}16, KYB^{+}15, LLDL13, LSA^{+}15, LMM^{+}15, LMM^{+}16, ZJR^{+}18]. parahaemolyticus-infected [LMM^{+}15, LMM^{+}16]. Paralichthys [DBHC17, HKK^{+}11b, JKC^{+}11, LKS^{+}13, LPC^{+}18, LWZ^{+}18, SSJ^{+}19, WWS^{+}11]. parameters [SPCR^{+}19, VST^{+}19, WM19, WSA^{+}18, XKB12]. Paramoeba [CHB^{+}17, HSA17, WNMK^{+}16]. Parasite [MWS17, USM17, ABC^{+}10, BMD^{+}14, BND^{+}17, BCJ13, CBM^{+}17, DAO19, EPCSBP18, FHD13, HHO14, HGH^{+}16, LRB^{+}17, ME12, MKK^{+}11, MLWC11, MPM16, PZL^{+}14, PMP^{+}18, RAVC18, RRC^{+}17, RKPG16, SBW^{+}18, SBD11, SGQ^{+}15, SSK17, TOK^{+}16, WTM14, WNO^{+}18, YMBE13, YYW^{+}14, ZCC18, MJ17]. Parasites [SMJ^{+}12, BCM^{+}17, ERÁpSB14, LZGZ16, MS16, MGM^{+}18, VDO^{+}18b]. Parasitic [BA13, FŠK^{+}19, AJLK^{+}18, BDD^{+}18, EDN18, GJMH17, vGJKR^{+}17, LTC^{+}14, LTC16, LTC17, SAD^{+}12]. parasitica [RDW18, SSSP18, STES14, SWWN^{+}16, TFG19]. parasiticides [XZZ16]. parasitized [JMHK18]. parasitizing [GAC10]. Parasitosis [SBK^{+}13]. Paratelphusa [MRR^{+}14, RNM^{+}12]. parauberis [HKK^{+}11b, LFSZ18, MP18, NLKA16]. parietalis [WEPB^{+}18]. Park [HGPD11, HOHG13]. parr [STES14]. part [BTN^{+}18, JMS13]. partial [GFS^{+}14, RB15, ZB12]. parva [MKK^{+}11]. parvalbumin [LSA^{+}15]. Parvicapsula [BLP10, JNN^{+}11]. parvo [MRR^{+}14]. parvo-like [MRR^{+}14]. parvovirus [YTL10]. passage [OBFR12]. passages [STES14]. passively [NPGA^{+}19]. Pasteurella [EWH^{+}19, RB15]. Pasteurellosis [AGR^{+}16]. Pastinachus [SDZN19]. pathogen [AHITC18, AHAMR^{+}19, BPMI^{+}16, BMA13, DXX^{+}15, DCL^{+}17, DABO^{+}17, EAN^{+}18, EKT^{+}11, FMA^{+}19, FMB^{+}11, KJB11, KPKP10, LITCAH19b, MB17, MSG^{+}18, MFHHG18, Mor12, PKLH19, ŘKH16, RCGGM15, SHW^{+}11, SP15, SMKBP17, VC17, XJL^{+}15, YSV^{+}13, YVM^{+}14, YLX^{+}19a, ZQB^{+}11, ZZA^{+}14]. pathogen-associated [RCGGM15]. Pathogenesis [MGN^{+}16, MKY^{+}15, ADF14, ICVV16, KOI^{+}14, MAC^{+}15, RSIP^{+}17]. Pathogenic [BHA^{+}11, ACAS14, BTN^{+}18, CHH^{+}14, FLVA10, HWM^{+}13, HBH^{+}19, KF10b, KKM^{+}19, LQL^{+}10, LML^{+}13, MS16, MoSMS18, NMM^{+}18, STIG17, SCS^{+}12, TDLT^{+}16, WZQ^{+}16, YLS^{+}19, ZDWH18]. Pathogenicity [EWH^{+}19, TES11, BBP18, FMA^{+}19, HXZ^{+}16, JGI^{+}10, JSJR^{+}19, JLJ16, LQL^{+}10, LQY11, NN11, PLC^{+}19b, ŘKH18, RSPA^{+}18, SSF^{+}17, VTM^{+}14]. pathogens [ATJS^{+}18, CGW^{+}19, DSA^{+}12, DAB^{+}12, DYML^{+}18, EKT^{+}11, LFH^{+}11, MDC^{+}18, RAM^{+}18, SHS14, VME^{+}19, WMM17, ZCQF15]. Pathologic [ACLP^{+}19, FFLC17]. pathological [LAN^{+}14, LF16, PAZ18]. Pathology [Loh13, MLMR16, SQC^{+}14, BWB^{+}17, BTC^{+}17, DDBC^{+}15, FYO^{+}11, GKCC^{+}12, LFP^{+}11a, LFP^{+}11b, MRSC13, SRT^{+}14, WPH^{+}13, WA14]. Pathophysiological [PCK^{+}15]. pathway [JNS^{+}19, WTM14, XJL^{+}19]. pathways [LMM^{+}15, LMM^{+}16]. pattern [AAZS^{+}15, DBV^{+}13, ZWY^{+}11]. patterns [AGFG19, DIVC16, EDN18, ONGQ^{+}14, PHR14, RCGGM15, VDMJ^{+}13, VMBGE17]. paucus [ABM^{+}15]. PCR [KKT^{+}18b, ATM^{+}10, AOAH13, AdOG^{+}17, AHITC18, BKS^{+}17, BBC^{+}10, BMBS12, CDF^{+}18, CYS17, CDG^{+}14, CZL^{+}19, DPMK12, DCL^{+}17, ECP^{+}10, FSM^{+}16, FSO^{+}10, FGR^{+}12, FSG^{+}12, GJK^{+}11, GTH11, GBSF12, HSA17, HRL15, HGB^{+}11, JKS^{+}13, JNN^{+}11, KJW^{+}12, KBJ^{+}13, KKT^{+}18a, LGS19, LML^{+}13, LGC^{+}19, LLS^{+}11, MMBS10, MCW^{+}18, NLKA16, NMM^{+}18, SKN^{+}16, SS17, SFL^{+}16, TSK^{+}19, TCFáS19, TAC15, YTL10, YGL^{+}18, ZCC^{+}15, ZHFC11, ZFZ^{+}11, vBVLR^{+}16]. PCR-based [BMBS12, FSM^{+}16]. PCR-DNA [KKT^{+}18a]. PCRs [MTAB15]. PD [Ano10c, BMT^{+}15, JWO^{+}10, JTW^{+}10, SVY^{+}14, TWNB^{+}12]. pear [BAH^{+}16]. pearlspot [MRN^{+}19]. peelii [SJT14, SJT14]. Pelagia [MLMR16]. peled [SMS^{+}18, SMS^{+}18]. Pelletan [HMDB13]. pellets [SNM^{+}16]. pelvic [FPK^{+}17]. penaeid [BPK^{+}17, CNN^{+}10, HVVV15, SP15, TDLT^{+}16, TDT^{+}16, YTL10, PHW^{+}16]. Penaeus [CMEV17a, CMS^{+}17, CMEV17b, CWW^{+}18, DCMM13, DBM^{+}14, EEMCEGSP16, FSO^{+}10, HXZ^{+}16, JJS^{+}12, MCO11, NNCLGC^{+}11, OLC11, PSK^{+}13, RDN^{+}12, SMS^{+}10, STS^{+}15, SRSF10, STG^{+}10, SNM^{+}16, TSK^{+}19, TTL^{+}16]. Peninsular [SSO^{+}14]. Pennsylvania [SBW^{+}18]. peptide [CMH^{+}17, UN10, ZN10]. peptides [PMK12]. Peracetic [SMFM12]. Perca [BGRB14, BGCKF12, CGP^{+}19, KHLL16, OEG^{+}13, PHvdH12, PBM^{+}15, RPM^{+}19, SGQ^{+}15, SKPS16]. percae [BGRB14]. perch [BGRB14, BGCKF12, CGP^{+}19, FHD13, HBB14, KHLL16, OEG^{+}13, PHvdH12, PBM^{+}15, RPM^{+}19, SGQ^{+}15, SKPS16]. Perciformes [EKS^{+}13, LXL^{+}18]. peregrinum [PBW^{+}18]. perennis [AA10]. performance [BWP^{+}19b, LKS^{+}13, LWZ^{+}18, PSCFB^{+}18, PG16]. period [OM10, SMS^{+}10]. peripheral [WXL16]. peritoneal [DRG^{+}17, GSS^{+}10]. Perkinsus [CdAPdC^{+}19, FBVC16]. permanganate [DBS12]. permissive [HO13]. Peroxide [MGM^{+}18, ACN12, FVAVMGE18, HKBD15a, HKBD15b]. Perry [BCW^{+}15, SBHC18, WZQ^{+}16]. Persistence [BMMC^{+}15, GMN12, HCEH16, WIC^{+}19, GFW^{+}10]. Persistent [JJSJ15, PWW19, Wen16]. perspective [dSCBdS^{+}16]. perspectives [Dal18, HWO^{+}17, ZGY^{+}18]. Peru [MCSM17]. perurans [CONN10, CBLN13, CAS^{+}18, CHB^{+}17, FGR^{+}12, HSA17, LTC16, LTC17, MCCN14, NBJ10, WNMK^{+}16, WNO^{+}18]. pet [KZJ^{+}17, BWM^{+}13]. pet-shop [KZJ^{+}17]. PET/ [BWM^{+}13]. petechial [PWW19]. Peters [AZSSZZ13, HVVV15, LCOZ10, MWS17, SPLV^{+}15, TvWO^{+}16, vD14]. pH [DSA^{+}12, DAB^{+}12, MMNL14]. Phacoemulsification [ALG15]. Phaeochromocytoma [TIK^{+}14b]. Phaeophleospora [ŘKH18]. phages [CHV^{+}12]. pharmaceutical [AH16]. pharmacodynamic [RHTC19b]. Pharmacokinetic [RHTC19b, dOAPLá^{+}17, LWJ^{+}16, Sam10]. Pharmacokinetics [DdRCP13, DPCdR14, FHZ^{+}16, GSS^{+}10]. phase [DJR^{+}19]. PHB [HBH^{+}19]. Phenotype [VSS17, DPD^{+}18, HBVB16]. phenotypes [HRNW15, TALA11]. Phenotypic [DLPR15, IAR^{+}10, PAZ18, DMK^{+}19, RVC18, SHO^{+}17, WZA10, ZQB^{+}11]. Phenotypical [PKPG15]. philippica [YLX^{+}19b]. philippinarum [FBVC16]. Philometra [SBM^{+}11]. philometrid [MPM16]. Philometridae [MPM16]. philppinarum [ACYC10]. phocae [CGCAH12, GCMG^{+}11]. Phoma [dSRSHGá^{+}19]. phospholipase [ZWG^{+}17]. phosphotransfer [BSV^{+}18]. photobacteriosis [KPSP^{+}19]. Photobacterium [BBP18, CDF^{+}18, CRS^{+}13, HHM^{+}14, HCL^{+}14, KPSP^{+}19, OJRL10, SPS^{+}17]. Phyllopharyngea [GJMH17]. Phyllopteryx [LYD^{+}15]. Phylogenetic [PHvdH12, VC17, CGP^{+}19, NRMS14, NSE^{+}17]. phylogeny [BMH^{+}19, GAC10, RRC^{+}17, SCD19, YLW^{+}14, YMBE13]. physical [SDI^{+}18]. Physiological [CRO^{+}18, MSS^{+}17b]. physiology [BLP10, GTGCM19]. Piaractus [CMSHS16, FCS^{+}18, VGI^{+}16]. Picornavirus [LEM^{+}19]. Pierre [HMDB13]. pigment [Góm12]. pike [Řeh13]. pikeperch [BAH^{+}16]. Pimephales [JBGP11, MGN^{+}16, PPF^{+}12, WBD^{+}10b]. Pimphales [AHHR^{+}10]. pink [BPPKJ10, SMJ^{+}12]. PirA [NPGA^{+}19, CCLGRM^{+}17, dJDAVRGM^{+}18]. pirA- [dJDAVRGM^{+}18]. PirA-like [CCLGRM^{+}17]. piranha [EDSL11]. pirB [dJDAVRGM^{+}18]. pirB-like [dJDAVRGM^{+}18]. Pisces [ANB^{+}19, MGM^{+}16, VDO^{+}18a]. Piscichlamydia [WNGS^{+}17]. piscicida [ACAS14, CDM^{+}16, CGAS19, CDF^{+}18, CRS^{+}13, EAN^{+}18, EA18, GPWF19, HCL^{+}14, KPSP^{+}19, OJRL10]. piscicida-associated [CDM^{+}16]. piscicola [IAR^{+}10]. piscidin [ZN10]. Piscine [GFO^{+}13, GMG^{+}19, GMC^{+}16a, MMB^{+}15, BBC^{+}10, BWNS^{+}11, GB18, JNS^{+}19, KBTMH^{+}16, MIO^{+}18, PPE^{+}18, TB12, TCM^{+}19, WNSAL12, WNAF^{+}13, BHO^{+}19, SRM^{+}18, SNK^{+}19, VCS^{+}19, WHR^{+}18]. pisciphila [ŘKH17, ŘKH18]. Piscirickettsia [CRO^{+}18, CIL^{+}17, DRG^{+}17, FCL^{+}19, HBB^{+}15, HKB^{+}16, ISCF^{+}19, LCV^{+}17, MB17, MSG^{+}18, MLRS^{+}17, MFHHG18, MID^{+}19b, OM10, PVIM18, RGD^{+}13, RSIP^{+}17, RSPA^{+}18, RE14, SHO^{+}17, SHKF16, SDR^{+}15, SCR^{+}15, YSV^{+}13, YVM^{+}14]. Piscirickettsiosis [RE14, MID^{+}19a, RGD^{+}13, RSIP^{+}17]. piscium [CPP14, CPdPP16]. PKD [MJ17]. PKD-parasite [MJ17]. placement [GAC10]. Placopecten [BRL^{+}13]. plague [KPKP10, SMKBP17, VDHV^{+}16, VDH19]. plant [ESA^{+}10, NFA15]. plant-mixed [ESA^{+}10]. plaque [CCCTA^{+}14]. plasma [FBVC16, FBK^{+}16, MMS^{+}16]. plasmid [BLFRCQ10]. Platax [DTM^{+}10]. platensis [RBH^{+}17]. platform [JJS^{+}12, MPJ^{+}15]. Platyhelminthes [BB11, YEK^{+}17]. play [BPCS19, WLX^{+}10]. plays [PJZB19, ZZQ^{+}19]. plecoglossi [ZCC18]. plecoglossicida [HZJ^{+}19, LXZ^{+}19, ZZA^{+}14, ZWG^{+}17]. Plecoglossus [CONN10, FNISE12, KSS^{+}15, ZCC18]. plerocerci [BB11]. Pleuronectiformes [YMBE13]. PLGA [ASHZ^{+}12]. ploidy [FFP^{+}14, LAN^{+}14]. PMCV [BWNS^{+}11]. Poecilia [ADR11, BCM^{+}17, LCOZ10, SPLV^{+}15]. Poeser [ADR11]. Poland [NJK^{+}17]. policies [JMD^{+}18]. Polish [BMGS17, MBS^{+}17]. Pollock [Dig15, WMF14]. polluted [HWO^{+}17]. poly [HBH^{+}19, SCS^{+}12, ASHZ^{+}12, OTN^{+}12, OKS^{+}13]. poly- [HBH^{+}19, SCS^{+}12]. polyclonal [ERÁpSB14]. Polycystic [JCS11, RhMV^{+}15, RMAM12]. Polydora [CBM^{+}18]. polydorid [CBM^{+}17]. polymerase [BHT^{+}17, CGG^{+}10, HLP^{+}16, TSS18, THW^{+}12, WSX^{+}18, YEK^{+}17]. polymicrobic [CDD^{+}18]. Polymorphism [HBB^{+}15, CPB^{+}15]. polymorphisms [FCL^{+}19]. Polynesia [DTM^{+}10]. polyomavirus [WCW^{+}16b]. Polyphasic [DLCF13]. polyphenol [MDF^{+}17]. polysaccharide [ZKL^{+}19]. polystyrene [HRNW15]. polyvalent [HJN^{+}19]. pomfret [CWL^{+}17, LJC^{+}19]. Pomoxis [BCT^{+}17, GPWF19, SMS^{+}19]. pompano [AZSI^{+}12, LYL^{+}18, WKW13]. pond [DML^{+}16, HMV^{+}17, SMS^{+}18, TSK^{+}19]. pond-reared [DML^{+}16, HMV^{+}17, SMS^{+}18, TSK^{+}19]. ponds [FBS13, MSK^{+}12, RHS^{+}12]. pool [LTM^{+}19]. pooled [HWM^{+}13]. pooling [LTM^{+}19]. Poorly [CIA^{+}18]. population [BDD^{+}18, DIVC16, EKS^{+}13, HHO14, ISCF^{+}19, LKT^{+}13, LUP^{+}18, MFES18, MVWS13, ONGQ^{+}14, SSL18, VPS^{+}14, YLL^{+}18]. populations [AAUHW15, BGCKF12, DBM^{+}14]. porcupinefish [SBK^{+}13]. pORF72 [KJW^{+}17]. porgy [EPA^{+}10]. Porpostoma [OBR^{+}14]. Portugal [BLSG^{+}16]. Portuguese [RRC^{+}17]. Positive [MSC^{+}17, AJLC^{+}16, BHA^{+}11, LLS^{+}11, LBF^{+}14, STS^{+}15]. possible [IFF^{+}15, MLF^{+}10, Tre12, YLSW14]. post [DPD^{+}18, ECP^{+}10, EEMCEGSP16, GM11, GDN^{+}16, HBB14, KS19, KAF^{+}19, NB10, RSIP^{+}17, RSPA^{+}18, SPCR^{+}19, SJSA^{+}14]. post-challenge [SPCR^{+}19]. post-hatch [DPD^{+}18]. post-larvae [EEMCEGSP16, KS19]. post-ovulatory [GM11]. post-release [HBB14]. post-smolt [NB10, RSIP^{+}17, RSPA^{+}18]. post-smolts [ECP^{+}10, KAF^{+}19]. post-stocked [GDN^{+}16]. pot [LYD^{+}15]. pot-bellied [LYD^{+}15]. Potamotrygon [MdOO16]. Potamotrygonidae [MdOO16]. potassium [DBS12, HCCM12]. potassium-deficient [HCCM12]. Potential [EKED^{+}17, LLDL13, MKI^{+}15, MSB^{+}19, STIG17, WCR^{+}18, ZJDL18, BE19, CLWJ13, CHH^{+}14, CHV^{+}12, GWM12, LMAKN16, MSS^{+}17a, PJL^{+}12, PSK^{+}12, PMPG^{+}16, RAS^{+}10, SHV^{+}16, SSL18, WDM^{+}15, WMM^{+}16, WXL16, WTM14, WMF14, XZZ16, YTKS19]. potentially [MS16].

Potomac [DIH^{+}16, PHR14]. pouch [ZZQ^{+}19]. povidone [MoSMS18]. Pox [WNGS^{+}17]. poxvirus [GGMG18b, GTT^{+}17, GCM^{+}18b, LDA^{+}19]. PP ExoU [ZWG^{+}17]. practical [BMBS12, HHO14]. practices [JR19, MMH^{+}18, RMS^{+}13]. pRAS1 [BLFRCQ10]. prawn [FMT^{+}14, IBFL19, LYKN18, MCO11, NRMS14, ZSZF16]. prawn-infecting [IBFL19]. praziquantel [BCT^{+}17, FHD13, OMB^{+}19, ZPM^{+}18]. pre [BHA^{+}17, GBV^{+}16, GDN^{+}16]. pre- [GDN^{+}16]. pre-analytical [BHA^{+}17]. pre-treatment [GBV^{+}16]. preadult [AH16]. precipitation [BTC^{+}17]. precision [SK15]. predatory [TIK^{+}14b]. predict [GCGR13]. Preliminary [ACN12, BVK^{+}16]. premises [NSC^{+}15]. preparation [WCW16a]. preparations [JBGP11, SGB^{+}16]. prerecruit [SWH^{+}14]. Presence [BSC^{+}14, GIW^{+}11, MGH^{+}14, VCS^{+}19, AAUHW15, DPTFC10, DPC^{+}10, MMW^{+}15, RDN^{+}12, VCR^{+}11, YLSW14]. present [ABVS18, BHT^{+}17]. presentation [CGP^{+}19]. preserves [FMM^{+}19]. pressure [SVY^{+}14]. presumptive [LGK^{+}14]. pretty [ŘKH18]. Prevalence [BWNS^{+}11, HKK^{+}11b, MIO^{+}18, MMWK19, MPS^{+}18, SHO^{+}17, SMS^{+}18, WNSAL12, BND^{+}17, BWB^{+}17, BWP^{+}19b, CXK16, DSJA18, FSM^{+}16, FFP^{+}14, FHFF15, HDWM15, KKK^{+}18, LUP^{+}18, LQY11, MMBS10, MSK^{+}12, PHR14, SBW^{+}18, SMJ^{+}14, ZDWH18]. prevalent [BAH^{+}18, GMG^{+}19]. prevent [AA10, BMK17, CB13, RHS^{+}12]. prevented [SJSA^{+}14]. prevention [FMB^{+}11, MUM^{+}19]. preventive [SSSP18]. prevents [NDA10]. previously [GLK^{+}16]. prickly [BAH^{+}16]. primary [RSB13]. primer [CZL^{+}19, GTH11, LGD^{+}17]. primer/probe [GTH11]. primers [VCOD17, YEK^{+}17]. Priming [PHW^{+}16]. Prince [MMMD16]. prior [GGW^{+}11]. Pro [RCGGM15, BSD^{+}18d]. Pro-inflammatory [RCGGM15, BSD^{+}18d]. pro-oxidative [BSD^{+}18d]. probe [GTH11, LGC^{+}19, SMF^{+}10]. probed [HRL15]. probes [BBC^{+}10, VLVS^{+}16]. probiotic [ABF15, BLFC12, DSSB18, GCNB16, LKS^{+}13]. Probiotics [NFA15, ASA10, LMC^{+}16]. problems [BK12]. Procambarus [AMAR18, DXX^{+}15, JPG^{+}19, WXWD15]. procedure [AHITC18, HAS^{+}19, LJGRI^{+}10, VCOD17]. procedures [HWM^{+}13]. process [LTC16]. processes [ARL^{+}17]. produced [KPM^{+}11, SBK^{+}15, WGL17]. produces [CRO^{+}18]. producing [RVC18]. product [WGL17]. Production [CCLL16, FMT^{+}14, KAS^{+}18, BPMI^{+}16, COB^{+}16, DPADO16, ESV^{+}19b, HDWM15, JR19, LAN^{+}14, MCO11, OLC11]. Production-level [KAS^{+}18]. products [ABF15, BGG11, LKS^{+}13, MMNL14, NFA15, TFG19]. profile [BSB^{+}18]. profiles [PRSK10, PMK12]. profiling [FBVC16, HKB^{+}16, LHW^{+}15b, WCL^{+}12]. progenitor [WKW13]. progeny [WNSAL12]. programme [GCS^{+}18]. programmes [BTN^{+}18]. progression [VPCM19]. proinflammatory [NSE^{+}17]. projections [RSR^{+}18]. Proliferation [JPG^{+}19, EPCSBP18, RBL^{+}11]. proliferative [BSW17, DMD^{+}17, RBF14]. promelas [AHHR^{+}10, JBGP11, MGN^{+}16, PPF^{+}12, WBD^{+}10b]. propagation [EYF^{+}19, GSEM13, NMV^{+}19]. properties [BGG11, GCMG^{+}11, HOHG13, NWMK16, ZWL^{+}16]. property [BSD^{+}18c]. proposed [TMWV14]. Prospective [GFW^{+}10, JWO^{+}10]. prospects [TDJ^{+}10]. protect [GCNB16, HBH^{+}19, LLDL13]. protected [ARFG13]. Protection [LXHL11, SLKE10, AHPR^{+}16, CHV^{+}12, Dal18, JOF^{+}19, JLJ16, LPA^{+}10, MBS^{+}19a, MSC^{+}17, SSV^{+}10, WBF^{+}10]. Protective [BGD^{+}17, CGR17, MA16, EDP^{+}17, ESA^{+}10, HCL^{+}14, JJV^{+}15, LLL^{+}18a, WCR^{+}18, YFA^{+}17]. protects [BBF^{+}15, CCLGRM^{+}17, CSSWK10, LRC^{+}18, RPMRLJ^{+}19, SNM^{+}16, SRB12, WHR^{+}18]. Protein [FBVC16, BRB^{+}12, CLWJ13, CXW^{+}19, DJQ^{+}12, DMB^{+}19, DJR^{+}19, FMT^{+}14, KTK17, KREM14, LC17, MTK^{+}10, NRMS14, NNC^{+}18, NMIY14, PLCC11, RRG^{+}19, SCK13, SNM^{+}16, SOGR^{+}19, VSS17, WCW16a, WCL^{+}19, XXL^{+}19, YZW^{+}19, ZWG^{+}18, ZWG^{+}17, ZLN^{+}17]. protein-based [ZWG^{+}18]. protein-labelled [KTK17]. protein-OmpU [CLWJ13]. protein-to-lipid [DMB^{+}19]. proteins [GSEM13, GAEK^{+}14, HLC^{+}15, JKC^{+}11, KREM14, LLC^{+}11, LLL^{+}12, LMAKN16, PLCC11, RAS^{+}10, SRSF10, SRD^{+}16, WBF^{+}10]. proteobacteria [CMR^{+}16, KMGK^{+}18]. proteolytic [SLPB18]. Proteomic [PJU^{+}13, GPL^{+}12, TCS19]. protocol [LWWS14, SEK^{+}16, YVM^{+}14]. Protopterus [RSR^{+}18]. Prototheca [JDRM17]. protothecosis [JDRM17]. Protozoa [SPP^{+}18]. protozoan [LPC^{+}18, PZL^{+}14]. protozoans [MUM^{+}19]. Province [LQY11]. Provisional [GCW17]. Prussian [LLL^{+}13]. PRV [BHO^{+}19, GFO^{+}13, GB18, VCS^{+}19]. PRV-1 [VCS^{+}19]. PSD [MNF^{+}15]. Pseudambassis [Góm12]. pseudobranchicola [JNN^{+}11]. Pseudocapillaria [KWGS19, MWRSK17, MKK^{+}11]. Pseudocyst [LTC17]. Pseudodactylogyrus [MWS17]. pseudodispar [ME12]. Pseudoloma [CXK16, SSK17]. Pseudomonas [AJLK^{+}18, ARFG13, BSSB18, BSD^{+}18c, GPFGAG19, HZJ^{+}19, LDEH17, LXZ^{+}19, SM14, SBW11, ZZA^{+}14, ZWG^{+}17]. pseudomycoides [OSAA18]. Pseudoplatystoma [FLC^{+}17]. Pseudorasbora [MKK^{+}11]. Pseudosciaena [HZJ^{+}19]. Pseudoterranova [BK12]. pseudotuberculosis [VPS^{+}14]. psychrophilum [SBW11, BLFC12, CHV^{+}12, CAJD^{+}18, DMP10, DJR^{+}19, FNISE12, GPL^{+}12, HKBD15a, HKBD15b, HJN^{+}19, HRNW15, JDH^{+}18, JML19, KSS^{+}15, LLC^{+}11, LPCC12, MBS^{+}19a, MBOC19, NN11, NSB^{+}18, NOV^{+}11, OBD^{+}11, PDLW17, PLCC11, PFJ^{+}10, RHB^{+}15, SDWA18, SEK^{+}16, SWE^{+}19, SW11]. Pterodon [VDO^{+}18a]. PTS [XXL^{+}19]. puffer [KOI^{+}14, MKMY10]. Puffy [MNF^{+}15]. pulchella [MAOS^{+}18]. pulcher [ŘKH18]. pulsed [TAC15]. pulsed-field [TAC15]. Pulteney [PDdC^{+}14]. pumpkinseed [LDEH17]. punctatus [BBF^{+}15, LTC^{+}14, MP18, PRSK10, RKPG16, YSK13, YK13]. puntazzo [KPT^{+}11]. Puntius [SMD^{+}12]. Purification [DLCC^{+}13, CCCTA^{+}14]. Purified [RHB^{+}15]. Purinergic [BSB^{+}18]. purposes [DAB^{+}12]. purse [LLZW14]. Putative [LSC^{+}14, GCNB16, GSEM13, YK12]. putitora [SM14]. putrefaciens [CDD^{+}18, EMA17, MP18, PKPG15]. Pygidiopsis [BCM^{+}17]. Pygocentrus [EDSL11].

qPCR [AGB^{+}19, GCNB18, GM10, GDN^{+}16, OBD^{+}11, VLVS^{+}16, VCOD17]. qPCR-based [GCNB18]. qRT [ECP^{+}10]. qRT-PCR [ECP^{+}10]. Quality [TWNB^{+}12, MoSMS18]. Quantification [OBD^{+}11, BKS^{+}17, CDF^{+}18, FSO^{+}10, GCNB18, LGC^{+}19, PPK^{+}13, TCFáS19, VCOD17]. quantifying [PSR^{+}11]. quantitation [NLKA16]. Quantitative [LLS^{+}11, NSC^{+}18, NWMK16, NOST^{+}16, SSJ^{+}19, BHT^{+}17, CMEV17b, EAM^{+}13, FGR^{+}12, FSG^{+}12, JBO^{+}15, LGC^{+}19, SS17, SFL^{+}16, TSS18, VST^{+}19]. quarantine [NSC^{+}15, RHS^{+}12]. quaternary [MoSMS18]. Queensland [BFFA^{+}12, BDB^{+}18, CBH^{+}15, DDBC^{+}15, Dig13, Dig15, OLC11, Pol15]. quelen [CCMA^{+}17]. quenching [DSSB18]. quinolones [HBB^{+}15]. quinqueradiata [MFS^{+}17, YYF^{+}10]. Quinsam [BPPKJ10]. Quorum [NDA^{+}19, DSSB18, HBVB16, ZCQF15]. Quoy [CCMA^{+}17, PC17, WCW^{+}16b].

r [FMT^{+}14]. r-MCP43 [FMT^{+}14]. rabbitfish [DRdS^{+}10]. Rachycentron [CLL^{+}10, HCL^{+}14, HSHC^{+}16, PSS17, SPS^{+}17, TWYC15]. Radiological [FvdMF^{+}18, RHST^{+}16]. Rafinesque [AHHR^{+}10, BBF^{+}15, BMA13, DWL^{+}17, HAP^{+}10, JBGP11, LTC^{+}14, LL11, LRB^{+}17, LDEH17, MGN^{+}16, PPF^{+}12, PRSK10, RKPG16, YSK13, YK13]. Rainbow [DJR^{+}19, vGJKR^{+}17, ASA10, ASHZ^{+}12, AOAH13, AHHR^{+}10, AHPR^{+}16, AJLC^{+}16, ABT^{+}14, AA10, AMA11, BSW17, BSPS^{+}18, BCLM^{+}19, BKD17, BHA^{+}11, BE19, BA14, BLFC12, BAH^{+}18, CSS^{+}16, CGW^{+}19, CHAB10, CKJ^{+}14, CMH^{+}17, CAJD^{+}18, CPB^{+}15, CLS^{+}15, DS10, DPTFC10, DPTCF10, DMP10, DSSB18, DSSB19, DdRTM13, DdRCP13, DPADO16, DSJA18, DT16, GRB^{+}17, GMB^{+}17, GCNB16, GPFGAG19, GDW^{+}15, GMC^{+}16b, HKBD15a, HKBD15b, HJN^{+}19, JKL^{+}13, JAJC^{+}18, JDH^{+}18, JML19, vGJHS^{+}11, KSM^{+}17, LPA^{+}10, LP17, LLS^{+}11, MBOC19, MNF^{+}15, MSC^{+}17, MCSM17, MHNY18, Mor12, MTLP^{+}18b, MNM^{+}18, MNM^{+}19, NB11, NOV^{+}11, NDA10, OP14, OLVJ^{+}13, OBFR12, OBD^{+}11, OSAA18, OLT^{+}10, PDLW17, PMW^{+}10, PMH^{+}11, PSBG^{+}11, PLCC11, QPB^{+}16, RSB13, ŘKH16, RCGGM15, RBL^{+}11, RHB^{+}15, SHJ^{+}12, SPCR^{+}19, SCR^{+}15, SON^{+}18, SSC17, VSS17, XKB12, YDO^{+}16]. raised [AdOG^{+}17, KTP^{+}15, SBM^{+}11]. Ralstonia [HBH^{+}19]. ram [LFSZ18]. ramirezi [LFSZ18]. ranavirus [GJM^{+}15, KRD^{+}17]. ranaviruses [VCR^{+}11, WBD10a]. ranched [SFU^{+}13]. Randomized [BSBH10]. ranga [Góm12]. range [BMBS12, BGD^{+}17, KF10a]. ranges [LHW^{+}15a]. Ranzani [BCG^{+}17, CHF^{+}12]. RAPD [DBM^{+}14]. RAPD-SCAR [DBM^{+}14]. Raphidascarididae [RGMGG^{+}18]. Rapid [ASSP^{+}15, CCHL13, HJOK11, KKT^{+}18b, LFH^{+}11, SEM18, WSX^{+}18, WKK12, ZFZ^{+}11, ZLN^{+}17, CDF^{+}18, CZB^{+}11, HH13b, KPSP^{+}19, KKT^{+}18a, LDÁM14, LXX^{+}17, MSI^{+}15, YWW^{+}19, YLX^{+}19a]. rare [GKO^{+}14]. rate [VDHV^{+}16]. rates [MPS^{+}18, OP14]. Rathbun [FSM^{+}16]. ratio [DMB^{+}19]. Ratzeburg [RFA^{+}14]. raw [TWNB^{+}12]. ray [BB11, BCJ13, VPCM19]. rays [BDB^{+}18]. razor [PDdC^{+}14]. rDNA [ZFZ^{+}11]. re [BSW17, PFC^{+}13]. re-exposure [BSW17]. re-infection [PFC^{+}13]. reaction [BHT^{+}17, BTC^{+}17, TSS18, THW^{+}12, YEK^{+}17]. reactions [ESV^{+}19b, GGGM14]. reactor [JSAB^{+}19]. Real [HGB^{+}11, JNN^{+}11, ATM^{+}10, AdOG^{+}17, BKS^{+}17, BKD17, BMBS12, CDF^{+}18, CDG^{+}14, DCL^{+}17, FSO^{+}10, FGR^{+}12, FSG^{+}12, GJK^{+}11, GTH11, GBSF12, HRL15, JKS^{+}13, KJW^{+}12, KBJ^{+}13, LGC^{+}19, MMBS10, MSI^{+}15, NLKA16, TCFáS19, YTL10, ZHFC11, vBVLR^{+}16]. Real-time [HGB^{+}11, JNN^{+}11, ATM^{+}10, AdOG^{+}17, BKS^{+}17, BKD17, BMBS12, CDF^{+}18, CDG^{+}14, DCL^{+}17, FSO^{+}10, FGR^{+}12, FSG^{+}12, GJK^{+}11, GTH11, GBSF12, HRL15, JKS^{+}13, KJW^{+}12, KBJ^{+}13, LGC^{+}19, MMBS10, MSI^{+}15, NLKA16, TCFáS19, YTL10, ZHFC11, vBVLR^{+}16]. reared [APP^{+}17, BRMG13, CONN10, DML^{+}16, GTT^{+}17, HMV^{+}17, JWO^{+}10, LHW^{+}15a, MSS^{+}17b, OBR^{+}14, PSS17, PFP^{+}13, RBWR^{+}15, SBP^{+}19, SMS^{+}18, TSK^{+}19, WBE^{+}17, XQS^{+}15]. rearing [MUM^{+}19, MPS^{+}18]. Reassortant [SODB16, SOA^{+}18, SOGR^{+}19]. recaptured [MKI^{+}15]. recently [GTT^{+}17]. receptor [NYRS12, VSS17]. receptors [NSE^{+}17, PRSK10, SHKF16, TQW^{+}14, ZGY^{+}18]. recirculating [BTN^{+}18]. recirculation [ACM^{+}14, GDW^{+}15, JSAB^{+}19]. recognized [GCM^{+}18b]. recognizing [YLS^{+}19]. Recombinant [CCLGRM^{+}17, CGW^{+}19, EBM^{+}18, FMT^{+}14, vGJKR^{+}17, LMAKN16, MTK^{+}10, WCW16a, YSK13, ZWG^{+}18]. recombinase [WSX^{+}18]. Recommendations [BTN^{+}18]. recompressed [PSK^{+}12]. record [BMA13, HHC^{+}18, LRB^{+}17, MCCN14, SPP^{+}18, SFBW^{+}17, ZCC18]. recovered [DLCF13]. Recovery [BC11, GPWF19, OSC^{+}14, OSAA18, PSK^{+}12, SM14, EA18, MGM^{+}18]. recruit [CBM^{+}18]. recruits [WOD16]. recurrent [VPCM19]. red [AMAR18, AZSSZZ13, AI11, CSS^{+}16, DLPR15, DSLR16, EPA^{+}10, EDSL11, GM11, ICVV16, IGLG^{+}17, JKH^{+}17, KKK^{+}18, KHC^{+}18, KLW10, LLZW14, MTG^{+}10, PCK^{+}15, RPMRLJ^{+}19, XJL^{+}19, GRB^{+}17, GMB^{+}17]. red-bellied [EDSL11]. red-mark [MTG^{+}10]. red-spotted [KKK^{+}18, XJL^{+}19]. redbanded [LJGRI^{+}10]. redescription [ZYW^{+}10]. Redfin [PBM^{+}15]. redspotted [CGG^{+}10, MRN^{+}19, ZLN^{+}17]. reduce [WNO^{+}18]. reduced [STF^{+}15]. reduces [CBM^{+}17, PFC^{+}13, RHB^{+}15]. reducing [SSC17]. reduction [LRS^{+}17, QWTL17]. reductions [QZZ^{+}18b]. reef [BGC^{+}17, CDM^{+}13, PC17]. refeeding [CMEV17a]. references [HBS^{+}14]. refined [NNC^{+}18]. regalis [CCM^{+}11]. Regan [RhAS^{+}15]. regardless [MTLP^{+}18b]. regime [LAN^{+}14]. regimen [RHTC19b]. regimes [BRL^{+}13]. Region [AAUHW15, MdOO16, DPMK12, EKED^{+}17, dSdAdS^{+}19, BNS13]. regions [HZP^{+}12, TQW^{+}14, CCW^{+}15]. regius [EKS^{+}13, TAQ^{+}10]. regulation [HGX^{+}19, RBM^{+}16, SJSA^{+}14, VMBGE17, XXL^{+}19]. reidi [DSMC17, LKT^{+}13]. reinforced [MKI^{+}15]. reinhardtii [CSA^{+}15]. related [BCLM^{+}19, COB^{+}16, FVAVMGE18, FSB18a, HST^{+}13, JBO^{+}15, JJS^{+}12, LMM^{+}15, LMM^{+}16, PdLC19, SKJ13, TCC10a, WCW^{+}16b]. relation [JHRW17, RAS^{+}10]. relationship [LWMT10, SAD^{+}12]. Relationships [OLT^{+}10, MKKM15]. Relative [SMJ^{+}14, CDG^{+}14, FSO^{+}10, FMH^{+}15, GMC^{+}16b, SSJ^{+}19]. relatively [GRP^{+}15]. release [HBB14]. relevant [PSVK^{+}18, RHB^{+}15, SWD^{+}14]. Reliability [HWM^{+}13]. reliable [Sla14]. remarkable [WGL17]. remedy [BE19]. removal [LL11]. renal [CBFM16, RHSM^{+}18a]. Renibacterium [BPMI^{+}16, BAMYAH19, EAM^{+}13, HDWM15, áKÁGA16, LMG^{+}19, PFJ^{+}10, PMPG^{+}16, STF^{+}15]. reovirus [DPADO16, FSM^{+}16, GFO^{+}13, GMC^{+}16a, LXHL11, MMB^{+}15, WCR^{+}18, WNSAL12, ZCCN13, ZCC^{+}15, ZWG^{+}18, ZB12, ZSZF16]. repair [FvdMF^{+}18]. repeat [ACAS14, HZP^{+}12]. Repeated [WNO^{+}18, CAS^{+}18, JAJC^{+}18]. repeats [HZP^{+}12, JGI^{+}10]. repellents [OWM^{+}17]. repetitive [CYS17, TAC15]. repetitive-element [CYS17]. replication [AGB^{+}19, GSBEM15, GGW^{+}11, HLP^{+}16, JQK14, RNM^{+}12, SOGR^{+}19]. Reply [Dig15]. report [ALG15, ATP^{+}16, BMK17, BRMG13, BO15, BHW^{+}11, CIA^{+}18, CSA^{+}15, DBHC17, EPA^{+}10, FKB^{+}14, GL18, GRB^{+}17, JDRM17, KPK^{+}18, KDP^{+}17, LGK^{+}14, LFSZ18, MdOO16, MVWS13, OOO^{+}11, SPS^{+}17, SdMVS^{+}16, WBE^{+}17, ZxYJ^{+}19]. reported [ZWLG10]. reports [RBF14, SWE^{+}19]. representing [KWD^{+}16]. reproduction [SSA13]. Republic [PFP^{+}13]. required [HGW^{+}19, KZJ^{+}17, NDA^{+}19]. rerio [AMRHM18, Bin14, CXK16, CKvGJB17, EDP^{+}17, KWGS19, LHW^{+}12, LHW^{+}15b, MLR^{+}12, MWRSK17, RAM^{+}19, SZM^{+}15, SSK17, ZJR^{+}18]. research [MPJ^{+}15, SSK17, TMWV14]. reservoir [WMF14]. reservoirs [EA18, GB18, MSK^{+}12]. residence [SMJ^{+}12]. resident [áKÁGA16]. residual [TNK12]. residue [DdRCP13, DPCdR14]. residues [FGJ^{+}10]. Resistance [JAR^{+}11, KNN19, MRN^{+}19, AAZS^{+}15, BSW17, CIL^{+}17, CKTT10, CYS17, CPB^{+}15, DMP10, DBM^{+}14, FCL^{+}19, HBS^{+}14, HBB^{+}15, JOF^{+}19, LKS^{+}13, LLZW14, NB11, OLT^{+}10, SMM^{+}13, SWD^{+}14, SPS^{+}19, TCC10a, VST^{+}19, WM19, WSA^{+}18, ZYW^{+}10, ZDWH18]. resistant [GTGCM19, GRP^{+}15, LLX^{+}19, LSA^{+}15, OFA11, PBM^{+}15, RKK^{+}19, RHBMM13, ZCJ^{+}18]. resisting [WXWD15]. respectively [MAOS^{+}18]. respiratory [CHAB10, GCM^{+}18b, UKS^{+}18]. Response [JCN^{+}14, JSAH^{+}18, MGM^{+}18, ABC^{+}10, BSD^{+}18b, BSB^{+}18, BCBS^{+}13, BAMYAH19, CLGP15, CGR17, CKJ^{+}14, CKvGJB17, CCME16, DPC^{+}10, DMB^{+}16, DJR^{+}19, DCZ^{+}17, ELI^{+}17, EDV^{+}18, FBVC16, GAVC16, HGH^{+}18, HSGD11, HKBD15b, HMW11, HGX^{+}19, HSHC^{+}16, JAJC^{+}18, vGJKR^{+}17, KBD13, LSC^{+}14, LMM^{+}16, LHW^{+}12, LHW^{+}15b, MWS17, NSC^{+}18, NYRS12, PKLH19, PMK12, RBAT^{+}17, RCGGM15, RSPA^{+}18, SSSP18, SPS^{+}19, SJSA^{+}14, TCC10a, TMC^{+}12]. response-related [LMM^{+}16]. responses [Ano12c, BLF^{+}16, BGD^{+}17, Buc12, CRS^{+}13, CSCCA14, FSB18a, HXZ^{+}16, HZX^{+}13, JMF^{+}18, JPG^{+}19, LMC^{+}16, MF12, MSS^{+}17b, NNC^{+}18, PPF^{+}12, QZZ^{+}18b, STS^{+}15, SVL^{+}17, TMM^{+}15, YSHG18, ZZQ^{+}19]. responsible [MHNY18]. restocking [VPP^{+}12]. Restricted [PMH^{+}11]. restriction [BCLM^{+}19, BLF^{+}16, RHB^{+}15]. Result [TDS^{+}19]. results [FBAS16, GRF14, HBS^{+}14, JWO^{+}10, KZJ^{+}17]. retained [ABM^{+}15]. retention [BDD^{+}18]. reticulata [LCOZ10, SPLV^{+}15]. retifera [CHF^{+}12]. retinopathy [DVC^{+}17, GFB^{+}13, NOST^{+}16, VPP^{+}12]. retrieved [VGP^{+}18]. Retrospective [CXK16, NB10]. returning [JKO^{+}13, LST^{+}12, MIO^{+}18, MMWK19]. revealed [WMD^{+}18]. reveals [BLF^{+}16, EDN18, EPCSBP18, IBFL19, MAC^{+}15, MLWC11, RDW18, RSPA^{+}18, WGL17, XJL^{+}19, ZCJ^{+}18]. reverse [LGD^{+}17, TSS18, YWW^{+}19]. Revie [JCN^{+}14]. review [DVC^{+}17, GFS^{+}18, MR11, Rob10, RAS^{+}10, RE14, SAS^{+}19, SMKBP17, WMM17, WBD10a, Dal18]. revision [ZWLG10]. RFLP [LGD^{+}17]. RGNNV [SOA^{+}18, ZLN^{+}17]. Rhabdosargus [SGR^{+}18]. rhabdovirus [AHB^{+}17, BPF^{+}18, BMMP^{+}14, CGP^{+}19, HMP^{+}12, SJDC^{+}13, SDJH16, YMH^{+}11]. Rhamdia [CCMA^{+}17]. rhamnose [LFD^{+}17]. Rhinobatos [AFA^{+}12]. Rhinoptera [BB11, VPCM19]. Rhinopteridae [BB11]. rhizoid [DSLR16]. Rhizoprionodon [BA13]. Rhodococcus [PBMB11]. Rhogostoma [DT16]. rhombus [LPVGL^{+}10]. rhusiopathiae [CSA^{+}15]. rhusiopathiae- [CSA^{+}15]. ribavirin [HLP^{+}16]. rice [MRR^{+}14, RNM^{+}12]. rice-field [MRR^{+}14, RNM^{+}12]. rich [HBH^{+}19, YFA^{+}17]. Richardson [BA13, CSA^{+}15, HBB14, KSAB^{+}14, MKKM15, XQS^{+}15, ZZA^{+}14]. rickettsia [BMH^{+}19, GBO^{+}18, LLS^{+}11, MTG^{+}10, RPJ^{+}18]. rickettsia-like [BMH^{+}19, GBO^{+}18, LLS^{+}11, MTG^{+}10, RPJ^{+}18]. Rickettsiales [CSS^{+}16, GMB^{+}17]. Rickettsiales-like [GMB^{+}17]. rifampicin [GPL^{+}12, OFA11]. rifampicin-attenuated [GPL^{+}12]. rifampicin-resistant [OFA11]. RIG [NSE^{+}17]. RIG-I-like [NSE^{+}17]. Risk [YAHWU11, JMD^{+}18, KAS^{+}18, MKKM15, NSC^{+}15, PPF^{+}12, TRS19, WMM^{+}16]. Risso [MKH^{+}11, RVB13, dSRSHGá^{+}19]. ritteri [SPTDEE^{+}15]. River [BPPKJ10, HOHG13, JMHK18, PHR14, SBW^{+}18, ARR^{+}18, MKI^{+}15, ZSZF16, BLP10, DIH^{+}16, HWO^{+}17, áKÁGA16, LEM^{+}19]. Rivers [HGPD11, LUP^{+}18, MJ17]. rivulatus [YCNW18]. RLO [BMH^{+}19]. RLRs [NSE^{+}17]. RMS [GRB^{+}17]. RNA [CGG^{+}10, EDN18, GMN12, HLP^{+}16, HMP^{+}12, KKT^{+}18b, PPE^{+}18, WNSAL12]. RNA-dependent [HLP^{+}16, CGG^{+}10]. RNAi [GSBEM15]. rock [JJV^{+}15, JLJ16]. rockfish [HKK^{+}11a, PSK^{+}12, ZxYJ^{+}19]. Rodlet [SJT14, BDM^{+}18]. Rodney [Rob14a]. rogercresseyi [BNS13, FVAVMGE18, GAVC16, GOM^{+}19, HBS^{+}14, MSMM19, MGM^{+}18, VMBGE17, YAHWU11]. rohita [KREM14, SSSP18]. role [BCLM^{+}19, BPCS19, DJQ^{+}12, HGX^{+}19, HOHG13, KSS^{+}15, NFA15, NLC^{+}16, NBJ10, PJZB19, RAS^{+}10, STIG17, TWJP10, WYZ^{+}19, XJL^{+}19].

Roles [MoSMS18, WLX^{+}10, ZZQ^{+}19]. roretzi [KOS^{+}10, PZL^{+}14]. rosenbergii [FMT^{+}14, HO13, LYKN18, RSGB^{+}14, SPP^{+}18, WCW16a]. Rosmarinus [ZTF^{+}10]. rostrata [LML12]. rotiferianus [ZxYJ^{+}19]. rotundus [ZWLG10]. rough [HRNW15]. roughtail [CBFM16]. round [GFB^{+}19]. route [SHJ^{+}12]. routes [BKD17, CMH^{+}17, DDBC^{+}15, FYO^{+}11, KSA^{+}10]. rPCR [LMG^{+}19]. rpoB [RB15]. rpoS [HGX^{+}19]. rRNA [DPMK12, LWWS14, OBR^{+}14, RB15]. RT [AGB^{+}19, GBSF12, HGB^{+}11, JKS^{+}13, KKT^{+}18b, LMG^{+}19, SKN^{+}16, VLVS^{+}16, VCOD17, YWW^{+}19]. RT-LAMP [YWW^{+}19]. RT-PCR [GBSF12, HGB^{+}11, JKS^{+}13, SKN^{+}16]. RT-PCR-DNA [KKT^{+}18b]. RT-qPCR [AGB^{+}19, VLVS^{+}16, VCOD17]. RT-rPCR [LMG^{+}19]. RTG [PMW^{+}10]. RTG-2 [PMW^{+}10]. RTGE [DPTCF10, RBL^{+}11]. rubripes [KOI^{+}14, MKMY10]. ruckeri [ASA10, AHTCA^{+}17, CMH^{+}17, DS10, DSSB18, DSSB19, GTGCM19, GCNB18, GTH11, JKL^{+}13, KSM^{+}17, KJW^{+}12, NCNB18, SON^{+}18, TALA11, TDH^{+}10, WNCR12]. Ruditapes [ACYC10, DABO^{+}17, FBVC16]. Rudolphi [FŠK^{+}19, MGH^{+}14]. rudolphii [DMB^{+}16]. rufa [VME^{+}19]. ruffe [MGM^{+}16]. runts [MGH^{+}14]. rupestris [BT14]. Russia [DKBS17]. Russian [RFA^{+}14]. ruthenus [RHSM^{+}18b].

s [ZWLG10, PLCC11]. S. [NMIY14, SWWN^{+}16]. sabinus [SCZM19]. sac [KEY^{+}15]. sacculitis [LWMT10]. Saddleback [Dig13, FPK^{+}17, Dig15, Pol15]. Safety [LML12]. salar [ACN12, ACA^{+}16, AAV^{+}16, ASSP^{+}15, Ano10c, ACM^{+}14, BSSH17, BMT^{+}15, BTC^{+}17, BHA^{+}17, BNS13, BSBH10, CBLN13, CJB^{+}15, CRAG^{+}12, DTF^{+}16, DKG^{+}17, DLLB^{+}18, DMB^{+}19, DRG^{+}17, ESBE18, ECP^{+}10, FMM^{+}19, FFP^{+}14, FHFF15, FMH^{+}15, GFO^{+}13, GFS^{+}18, GMC^{+}16a, GSS^{+}10, GGH^{+}10, GAVC16, GCMG^{+}11, GMC^{+}16b, GBSF12, GFR^{+}12, GRT12, GHB^{+}11, GCGR13, HHO14, HSA17, HST^{+}13, IPW^{+}13, IAR^{+}10, JCN^{+}13, JKO^{+}13, JCN^{+}14, JWO^{+}10, JTW^{+}10, JBO^{+}15, JOF^{+}19, JNS^{+}19, JHGR13, JJSJ15, KAF^{+}19, KSA^{+}10, KRFT14, KBD13, KWD^{+}16, LDÁM14, LAN^{+}14, LMG^{+}19, MKI^{+}15, MIO^{+}18, MMWK19, MSMM19, MSG^{+}18, MLMR16, MLER^{+}18, MSBB^{+}16, MLF^{+}10, MRSC13, MAC^{+}15, MID^{+}19b, MID^{+}19a, MSR^{+}10, MGH^{+}14, MMH10, NCNB18, NB10, PdLC19, PBMB11, PHVM13, PPK^{+}13, PPI^{+}13, PÅD18, PJU^{+}13, RB15, RMR14, RSIP^{+}17, SMR10, SMM^{+}13, SHKF16, SFR^{+}17, SAS^{+}19, SFU^{+}13]. salar [SG12, STES14, SWWN^{+}16, SWS^{+}16, SKJ13, SSA13, SNK^{+}19, TWNB^{+}12, TJB^{+}15, TCC10a, TES11, TCM^{+}19, Tre12, USM17, VSDS^{+}19a, VDMJ^{+}13, WDM^{+}15, WMM^{+}16, WNLF^{+}12, WNSAL12, WNAF^{+}13, WNMK^{+}16, WNAJ^{+}16, YP17, ZCCN13, ZNHC11a, ZNHC11b]. Salaria [VPP^{+}12]. salaris [HHO14, HKG^{+}15, KHG^{+}17, PHVM13, WKK12]. salinities [NNCLGC^{+}11]. salinity [CCHL13, KCL14, LTC17, PST^{+}16, TTL^{+}16]. Salmincola [VCLW16, STF^{+}15]. Salmo [ACN12, AGFG19, ACA^{+}16, AAV^{+}16, ASSP^{+}15, Ano10c, ACM^{+}14, BSPS^{+}18, BSSH17, BMT^{+}15, BTC^{+}17, BHA^{+}17, BNS13, BSBH10, CBLN13, CJB^{+}15, CRAG^{+}12, DTF^{+}16, DKG^{+}17, DLLB^{+}18, DMB^{+}19, DRG^{+}17, ESBE18, ECP^{+}10, FMM^{+}19, FFP^{+}14, FHFF15, FMH^{+}15, GFO^{+}13, GFS^{+}18, GMG^{+}19, GMC^{+}16a, GSS^{+}10, GGH^{+}10, GAVC16, GCMG^{+}11, GMC^{+}16b, GBSF12, GFR^{+}12, GRT12, GHB^{+}11, GCGR13, HHO14, HSA17, HST^{+}13, IPW^{+}13, IAR^{+}10, JCN^{+}13, JKO^{+}13, JCN^{+}14, JWO^{+}10, JTW^{+}10, JBO^{+}15, JOF^{+}19, JNS^{+}19, JHGR13, JJSJ15, KAF^{+}19, KSA^{+}10, KRFT14, KBD13, KWD^{+}16, LDÁM14, LAN^{+}14, LMG^{+}19, MKI^{+}15, MIO^{+}18, MMWK19, MSMM19, MSG^{+}18, MLMR16, MLER^{+}18, MSBB^{+}16, MLF^{+}10, MRSC13, MAC^{+}15, MID^{+}19b, MID^{+}19a, MSR^{+}10, MGH^{+}14, MMH10, NCNB18, NB10, OSC^{+}14, PdLC19, PBMB11, PGF^{+}19, PHVM13, PPK^{+}13, PPI^{+}13, PÅD18, PJU^{+}13, RB15, RMR14, RSIP^{+}17, SMR10]. Salmo [SMM^{+}13, SHKF16, SFR^{+}17, SAS^{+}19, SFU^{+}13, SG12, STES14, SWWN^{+}16, SWS^{+}16, SKJ13, SSA13, SNK^{+}19, TWNB^{+}12, TJB^{+}15, TCC10a, TES11, TCM^{+}19, Tre12, USM17, VSDS^{+}19a, VDMJ^{+}13, WDM^{+}15, WMM^{+}16, WNLF^{+}12, WNSAL12, WNAF^{+}13, WNMK^{+}16, WNAJ^{+}16, YP17, ZCCN13, ZNHC11a, ZNHC11b]. salmoides [CGAS19, PBBG19, ZLX^{+}19]. Salmon [CCCTA^{+}14, GTT^{+}17, GCM^{+}18b, LASRA^{+}19, TJA^{+}13, AHH16, AH16, ADF14, ADMF15, ATM^{+}10, ACN12, AAUHW15, ACA^{+}16, AAV^{+}16, ASSP^{+}15, Ano10c, ACM^{+}14, AHTCA^{+}17, BHT^{+}17, BHO^{+}19, BMT^{+}15, BTC^{+}17, BHA^{+}17, BPPKJ10, BLP10, BNS13, BMH^{+}19, BDD^{+}18, BSBH10, CBLN13, CJB^{+}15, CGR17, CHWM^{+}19, CRAG^{+}12, DTF^{+}16, DKG^{+}17, DWP^{+}19, DPD^{+}18, DLLB^{+}18, DMB^{+}19, DRG^{+}17, DPADO16, DYML^{+}18, EYF^{+}19, EDN18, ESBE18, ECP^{+}10, EWRH13, FMM^{+}19, FSB18a, FFP^{+}14, FHFF15, FMH^{+}15, FSV^{+}18, FSB^{+}18c, GL18, GFO^{+}13, GFS^{+}18, GGMG18b, GVB^{+}17b, GBO^{+}18, Gis18, GTH11, GSS^{+}10, GGH^{+}10, GCMG^{+}11, GMC^{+}16b, GFW^{+}10, GBSF12, GFR^{+}12, GRT12, GHB^{+}11, GCGR13, GCS^{+}18, HHO14, HWM^{+}13, HGH^{+}18, HSA17, HST^{+}13, IAR^{+}10, IOO15, JCN^{+}13, JKO^{+}13, JCN^{+}14, JWO^{+}10, JTW^{+}10, JJB15, JMF^{+}18, JBO^{+}15, JOF^{+}19, JNS^{+}19, JR19, JHGR13, JJSJ15, KOB^{+}17, KBE^{+}14]. salmon [KAF^{+}19, KHG^{+}17, KSA^{+}10, KRFT14, KBD13, KWD^{+}16, LDÁM14, LAN^{+}14, LMG^{+}19, LDA^{+}19, LST^{+}12, LF16, LF18, LGJ19, LWMT10, MKI^{+}15, MIO^{+}18, MMWK19, MSMM19, MSG^{+}18, MLMR16, MLER^{+}18, MSBB^{+}16, MLF^{+}10, MRSC13, MAC^{+}15, MID^{+}19b, MID^{+}19a, MSR^{+}10, MUM^{+}19, MGH^{+}14, MTHB13, MPS^{+}18, MMH10, NRF^{+}18, NCNB18, NB10, NLC^{+}16, NBJ10, OWM^{+}17, OM10, OBFR12, PdLC19, PBMB11, PHVM13, PPK^{+}13, PPI^{+}13, PÅD18, PJU^{+}13, PMPG^{+}16, PPE^{+}18, RB15, RMR14, RRG^{+}19, RSIP^{+}17, RSPA^{+}18, SMR10, SMJ^{+}12, SMM^{+}13, SFR^{+}17, SAS^{+}19, STF^{+}15, SFU^{+}13, SG12, STES14, SWWN^{+}16, SWS^{+}16, SJK^{+}14, SVY^{+}14, SSA13, SSL18, SNK^{+}19, TWNB^{+}12, TJB^{+}15, TCC10a, TVK^{+}19, TES11, TES16, TCM^{+}19, Tre12, UMMS17, USM17, UKS^{+}18, VSDS^{+}19a, VDMJ^{+}13, VMBGE17, WDM^{+}15, WMM^{+}16, WHR^{+}18, WNLF^{+}12, WNSAL12, WNAF^{+}13, WNMK^{+}16, WNAJ^{+}16, WNGS^{+}17]. salmon [WOD16, WNO^{+}18, YP17, ZCCN13, ZCC^{+}15, ZNHC11a, ZNHC11b, BSSH17, GMC^{+}16a, LLLG18, WNGS^{+}17]. salmonicida [BBP18, MLEM16, NLC^{+}16, SWM^{+}15, VSS17, BKS^{+}17, BKD17, BLFRCQ10, BVK^{+}16, DLCF13, EDV^{+}18, GDN^{+}16, GLK^{+}16, HKK^{+}11a, KBJ^{+}13, KBD13, MSC^{+}17, MLEM16, NLC^{+}16, PFJ^{+}10, RVC18, RPM^{+}19, SSHEM11, SSC17, VSS17, VSDS^{+}19a, VDD^{+}18, VC17, YFA^{+}17]. salmonicola [MSR^{+}10]. Salmonid [Ano10c, JTW^{+}10, BKG^{+}18, BTC^{+}17, CGR17, DLCF13, EBM^{+}18, FGR^{+}12, GB18, GFW^{+}10, GFM^{+}11, GFR^{+}12, GRF14, HDWM15, HFT^{+}12, HST^{+}13, HJT^{+}16, HB12, JJB15, JJM^{+}17, MIO^{+}18, MR11, NSE^{+}17, PSR^{+}11, RSP^{+}17, RSM^{+}18, SJDC^{+}13, SBH^{+}16, SHV^{+}16, SKJ13, TJB^{+}15, WKW^{+}11, WNCR12, YSV^{+}13, YVM^{+}14]. salmonids [BSPS^{+}18, BRB^{+}12, CLGP15, CHV^{+}12, GGMG18b, GAC10, JNN^{+}11, áKÁGA16, KKKA17, MMB^{+}15, MJ17, RMSH09, SDWA18, TJA^{+}13, VCS^{+}19, YAHWU11]. salmoninarum [BPMI^{+}16, BAMYAH19, EAM^{+}13, HDWM15, áKÁGA16, LMG^{+}19, PFJ^{+}10, PMPG^{+}16, STF^{+}15, TCFáS19, TCS19, YDO^{+}16]. salmoninarum-infected [HDWM15]. salmoniphilum [ACM^{+}14, RFA^{+}14, ZNHC11a, ZNHC11b]. salmonis [FSB18a, WNGS^{+}17, AHH16, AH16, BRL^{+}13, BDD^{+}18, CRO^{+}18, CIL^{+}17, DLLB^{+}18, DRG^{+}17, EDN18, FCL^{+}19, GAC10, GCGR13, HH13b, HBB^{+}15, HKB^{+}16, IPW^{+}13, ISCF^{+}19, JCN^{+}13, JCN^{+}14, JR19, JHGR13, KRFT14, LCV^{+}17, LLG^{+}15, LP17, LGJ19, MB17, MSG^{+}18, MLRS^{+}17, MFHHG18, MID^{+}19b, MUM^{+}19, NRF^{+}18, NLC^{+}16, NBJ10, OWM^{+}17, OM10, PVIM18, PJZB19, PPI^{+}13, PFC^{+}13, RGD^{+}13, RSIP^{+}17, RSPA^{+}18, RE14, SHO^{+}17, SMR10, SMM^{+}13, SMR16, SHKF16, SFU^{+}13, SDR^{+}15, SCR^{+}15, SSA13, USM17, WWE^{+}13, YSV^{+}13, YVM^{+}14]. salts [SHS14]. saltwater [HCCM12]. Salvelinus [CMR^{+}16, DGS^{+}14, SWM^{+}15, WKK12]. sample [ZCCN13]. samples [CPdPP16, FSG^{+}12, HSA17, LTM^{+}19, TSS18, TVK^{+}19]. sampling [FCM^{+}19, MTAB15]. sand [LL11]. Sander [BAH^{+}16, GRP^{+}15, VBL^{+}16, VKK^{+}19]. sandwich [ZLN^{+}17]. sanguinarine [LWJ^{+}16]. sanguineus [HCW^{+}16, ZWY^{+}11]. Santa [CdAPdC^{+}19]. Santos [DAO19]. sapidus [FSM^{+}16]. Saprolegnia [AS17, ASS19, KBE^{+}14, PPF^{+}12, RDW18, SCD19, SSSP18, STES14, SWWN^{+}16, SWS^{+}16, TFG19, TES11, TES16]. saprolegniasis [BAH^{+}16]. saprolegniosis [AGFG19, GPFGAG19]. sarana [SMD^{+}12]. sarba [SGR^{+}18]. sarcoma [CIA^{+}18]. sardine [GMP^{+}11]. Sardinella [GMP^{+}11]. Sardinia [APP^{+}17]. Saskatchewan [SGQ^{+}15]. sativa [MA16]. sativum [MSCH14]. saurida [KSAB^{+}14]. Sauvage [BAD^{+}12, CTK^{+}10, SMV^{+}14, SZZSFN^{+}14]. SAV [Ano10c, JTW^{+}10, KSA^{+}10, NSE^{+}17, RRG^{+}19, SBH^{+}16]. SAV2 [HST^{+}13]. SAV3 [BTC^{+}17, CGR17, JMF^{+}18]. saxatilis [DBS12, SBM^{+}11]. scad [BHW^{+}11, HHM^{+}14]. scale [ASB^{+}12, GKCC^{+}12, SPP^{+}18, SDM^{+}19, ZDWH18]. scallop [BRL^{+}13]. scanning [BCM^{+}17]. SCAR [DBM^{+}14]. scherzeri [KHC^{+}18, SWD^{+}14]. Schistocephalus [WFH^{+}18]. Schlegel [CONN10, FNISE12, GHH^{+}11, HKK^{+}11b, JKC^{+}11, JJV^{+}15, KSS^{+}15, KYK^{+}12, KLW10, LKS^{+}13, MKK^{+}11, MKMY10, NSS^{+}17, OMO^{+}15, SSJ^{+}19, YYF^{+}10, ZCC18]. schlegeli [HKK^{+}11a]. schlegelii [LLDL13, ZxYJ^{+}19]. Schlei [KAST17]. Schneider [CCM^{+}11]. schubertii [CLZ^{+}12, LL12, LGC^{+}19]. Schultz [LLG^{+}15]. schwanenfeldii [ARR^{+}18]. schwannoma [MLR^{+}12]. Sciaenidae [EKS^{+}13]. Sciaenocotyle [TAQ^{+}10]. Scleropages [KVC^{+}16]. Scombridae [CZB^{+}11]. scophthalmi [ERÁpSB14]. Scophthalmus [HMW11, LPVGL^{+}10, LBF^{+}14, MBB12, QXW14, SODB16, WM19, dOAPLá^{+}17]. Scopthalmus [JCS11]. Scortum [KRD^{+}17, LLH^{+}14b]. Scotia [SCD19]. Scotland [GFR^{+}12, GWM12, MMH10, SMR16, SDI^{+}18, SWE^{+}19]. Scottish [MMBS10, MRSC13, WMM^{+}16, WMM17]. screen [LPCC12]. screening [AHH16, AH16, AS17, ABF15, BLFC12, CMS^{+}19, GSEM13, XZZ16]. scuticociliate [OBR^{+}14]. scuticociliates [QZZ^{+}18b]. Scuticociliatida [SPP^{+}18]. scuticociliatosis [LPC^{+}18, MDF^{+}17]. scuticociliatosis-causing [LPC^{+}18]. Scyborskaya [PB18]. Scylla [LQY11]. Scyphozoa [PÅD18]. Sea [BT13, EWRH13, HGH^{+}18, JKO^{+}13, Kan18, PBC15, PdLC19, VPS^{+}14, ACSB^{+}19, BRL^{+}13, BC11, BCBS^{+}13, BCW^{+}15, CD17, CKTT10, CB13, DJQ^{+}12, DdRTM13, DPCdR14, FVAVMGE18, FPK^{+}17, GFO^{+}13, GBV^{+}16, GVB^{+}17b, GVB^{+}17a, Gis18, GCGR13, HH13a, HH13b, HBS^{+}14, JMD^{+}18, JWO^{+}10, JKH^{+}17, JHGR13, KPT^{+}11, KJB11, KHC^{+}18, KLW10, LMC^{+}16, LQL^{+}10, LGS^{+}14, MIO^{+}18, MSMM19, MMH^{+}18, MSS^{+}17b, NRF^{+}18, NOST^{+}16, PST^{+}16, PJU^{+}13, PFC^{+}13, RGD^{+}12, RHVG^{+}19b, SMR10, SMM^{+}13, SCF^{+}13, SDI^{+}18, Sam10, SDJH16, SFU^{+}13, SJK^{+}14, SSL18, TPD^{+}19, TMC^{+}12, TNK12, VDMJ^{+}13, VTM^{+}14, WWE^{+}13, WOD16, YAHWU11, AHB^{+}17, GFB^{+}13, GVLE16, LZGZ16, LHW^{+}15a, LXL^{+}18, PMP^{+}18, RBWR^{+}15, VDD^{+}18, WMF14]. sea-caught [MIO^{+}18]. sea-ranched [SFU^{+}13]. sea-reared [JWO^{+}10]. sea-trout [GFO^{+}13]. seabass [FCM^{+}19, GKCC^{+}12, KNN19, RRC^{+}17]. seabream [CVLJ^{+}10, CVGR^{+}13, CCME16, LLDL13, LJGRI^{+}10, RVB13, SGR^{+}18]. seadragon [LYD^{+}15]. Seahorse [QZZ^{+}18b, DSMC17, LYD^{+}15, LKT^{+}13, QWTL17, SBHC18, ZZQ^{+}19]. seahorses [BQP12, DCV^{+}14, OBR^{+}14, WZQ^{+}16]. seal [BK12]. Search [CIL^{+}17]. Seasonal [EA18, LFL19, SJK^{+}14]. Seasonality [YB17, YMBE13]. seatrout [GMG^{+}19]. seawater [Ano10c, GTT^{+}17, HRL15, JTW^{+}10, JMF^{+}18, KOB^{+}17, SMJ^{+}12]. seawater-farmed [KOB^{+}17]. seawater-reared [GTT^{+}17]. sebae [BRMG13]. Sebastes [HKK^{+}11a, PSK^{+}12, ZxYJ^{+}19]. Secondary [JAJC^{+}18, ATJS^{+}18, HFKF17, RSB13]. secreted [ZWG^{+}17]. secretion [EAN^{+}18, MSG^{+}18, RGD^{+}13]. secretory [CXW^{+}19]. sectional [AAUHW15]. seeds [MA16]. seeking [OWM^{+}17]. segment [PAM^{+}19]. Selar [HHM^{+}14]. selected [AAZS^{+}15, BMGS17, DAB^{+}12, EAM^{+}13, EEMCEGSP16, NJK^{+}17, SPCR^{+}19, TvWO^{+}16]. Selection [SPS^{+}19, YLS^{+}19, GTGCM19]. Selective [BTC^{+}17, CLW^{+}11, CLT^{+}14, SMF^{+}10]. selenium [WPH^{+}13]. semi [CMEV17b, STS^{+}15]. semi-intensive [STS^{+}15]. semi-quantitative [CMEV17b]. semilaevis [XJL^{+}15, ZQB^{+}11]. seminatural [HKG^{+}15]. senegalensis [CVLJ^{+}10, CTD^{+}12, CPP14, CPdPP16, CRS^{+}13, CSCCA14, FBL^{+}14, LVADB17, MMS^{+}16, SOA^{+}18, VFB^{+}12, YB17, dALF^{+}17]. Senegalese [CVLJ^{+}10, CTD^{+}12, CRS^{+}13, CSCCA14, FBL^{+}14, LVADB17, MMS^{+}16, YB17, dALF^{+}17]. senile [ARL^{+}17]. sensing [HBVB16, NDA^{+}19, ZCQF15]. sensitive [GCNB18, HJOK11, MMBS10, MSI^{+}15, SMF^{+}10, WOD16, YWW^{+}19, ZLN^{+}17]. Sensitivity [AdOG^{+}17, MMH^{+}18, MTAB15, PFP^{+}13, HH13b, ZCJ^{+}18]. sensory [FSB18a]. septemfasciatus [OTN^{+}12]. septempunctata [TOK^{+}16]. septicaemia [AHHR^{+}10, AHPR^{+}16, Ano12c, CBC^{+}13, CDM^{+}16, CSA^{+}15, DPADO16, EKED^{+}17, GJVT^{+}11, GFB^{+}19, GGW^{+}11, GVC^{+}19, HLL^{+}12, HGG^{+}11, HPAL10, JKS^{+}13, KKK15, KF10a, KF10b, LVADB17, MMBS10, MF12, MMW^{+}15, NLF11, OEG^{+}13, PBM^{+}15, PJL^{+}12, RMS^{+}13, SHJ^{+}12, SWH^{+}14, SJ11, VLVS^{+}16, VBL^{+}15, VBL^{+}16, WDM^{+}15, WKW^{+}11, ZLX^{+}19]. septicaemic [OCLI^{+}19]. septum [FFP^{+}14]. sequence [FMA^{+}19, ISCF^{+}19, JSK^{+}10, KKM^{+}19, MBS^{+}17, NMM^{+}18, RDW18, VJWC^{+}18, WMD^{+}18, YMH^{+}11]. sequences [OBR^{+}14, RB15, TQW^{+}14]. sequencing [BFC^{+}15, CGG^{+}10, DWL^{+}17, EDN18, LMM^{+}16]. sequential [MBT^{+}17]. sera [GRF14, LLL^{+}12]. Serbian [RAM^{+}18]. Seriola [KHS11, MFS^{+}17, MKH^{+}11, NWMK16, RBAT^{+}17, STIG17, SS17, YYF^{+}10, dSRSHGá^{+}19]. seriolae [CXW^{+}19, KLD^{+}18, STIG17, WCL^{+}19]. seriolicida [MFS^{+}17]. sero [LMAKN16]. sero-surveillance [LMAKN16]. Serodiagnosis [KKK15]. Serological [SDWA18, IAR^{+}10]. serotype [AOAH13, AHTCA^{+}17, BAH^{+}18, RB15]. serotypes [HKK^{+}11b]. serovars [FLVA10]. Serranidae [LXL^{+}18, PMP^{+}18]. Serrasalmidae [ANB^{+}19, OMB^{+}19, VDO^{+}18b, VDO^{+}18a]. serrata [LQY11]. Serratia [CBC^{+}13]. Serum [BSD^{+}18a, BMT^{+}15, BSD^{+}18d, CLC^{+}17, LWMT10, NN11, PBC15, YK13]. set [GTH11]. seven [CGCAH12, MTAB15, PMK12]. sevenband [OTN^{+}12, OKS^{+}13]. several [PPE^{+}18]. Severe [KHS11, CZB^{+}11, VFB^{+}12]. severity [LLS^{+}11]. sex [CXK16, FVAVMGE18]. sex-related [FVAVMGE18]. SGIV [GHH^{+}11]. SGPV [GGMG18b]. shad [WBE^{+}17, WEPB^{+}18]. sharing [MPJ^{+}15]. shark [ABM^{+}15, BA13, CBC^{+}13, CIA^{+}18, HKF^{+}16, JK12, LL11, MCWL13, RBF14]. sharkminnows [RTM^{+}15]. sharks [BOH^{+}12, PC17]. sharp [OAO^{+}12]. sharp-tooth [OAO^{+}12]. sharpnose [BA13]. sharpsnout [KPT^{+}11]. sharptooth [HGPD11, HOHG13, TMWV14]. shasta [HB12]. shed [RRO^{+}13]. shell [ARL^{+}17, DABO^{+}17, MMD16, SQC^{+}14]. shelled [MMMD16]. shellfish [BFJ11, KKK^{+}18, KKM^{+}19, RAS^{+}10]. shells [CBM^{+}17, CBM^{+}18]. Shewanella [CDD^{+}18, EMA17, JSJR^{+}18, LQL^{+}10, MP18, PPSW19, PKPG15]. shift [IFF^{+}15]. shiner [HAP^{+}10, LRB^{+}17, MGN^{+}16]. SHK [LASRA^{+}19, SHKF16]. SHK-1 [LASRA^{+}19, SHKF16]. shoaling [SSK17]. shock [DJQ^{+}12, RAS^{+}10, SNM^{+}16, WBF^{+}10, ZWY^{+}11]. shop [KZJ^{+}17]. short [LTC17]. short-term [LTC17]. shortnose [LGK^{+}14, JMHK18]. showed [KBD13]. shows [CXK16, VDO^{+}18a]. shrimp [APG^{+}11, BPK^{+}17, CCLGRM^{+}17, CNN^{+}10, CMEV17a, CMS^{+}17, CMEV17b, CLW^{+}13, DCMM13, DBM^{+}14, ESA^{+}10, EEMCEGSP16, HVVV15, JQK14, JJS^{+}12, KS19, KKT^{+}18b, KKT^{+}18a, LRD^{+}19, MRR^{+}14, MTK^{+}10, NPGA^{+}19, NSC^{+}18, PSK^{+}13, PHW^{+}16, RPGCIG17, SHS14, SPTDEE^{+}15, SRSF10, SSV^{+}17, SSV^{+}10, STG^{+}10, SP15, SVM^{+}18, TMM^{+}15, TTL^{+}16, TDLT^{+}16, TDT^{+}16, WLX^{+}10, YTL10, YLSW14, YHY^{+}16]. shrimps [BTN^{+}18, SHW^{+}11]. shrinking [LUP^{+}18]. shrivelling [JYZ^{+}14]. shuberti [LCZ^{+}18]. sialic [ACYC10]. Siamese [RhAS^{+}15]. Siberian [DKBS17]. Sickeningly [LFD^{+}17]. side [ELI^{+}17]. siderophore [BPMI^{+}16]. siderophores [CRO^{+}18]. sign [ZDWH18]. signalling [BSB^{+}18, LMM^{+}15, LMM^{+}16, XJL^{+}19]. significance [SMS^{+}10]. significant [MBS^{+}17]. significantly [BMD^{+}14]. signs [HBB14, HCEH16, SON^{+}18, XJL^{+}15]. Silurus [MAOS^{+}18]. silus [TB12]. Silva [KCL14]. silver [BSV^{+}18, BSSB18, BSD^{+}18b, BSD^{+}18a, BSD^{+}18d, CCMA^{+}17, FHD13, KKOY19, KJW^{+}17, LJC^{+}19, WXL16]. Simple [LCR^{+}10, MSI^{+}15, YWW^{+}19]. simplex [MGH^{+}14]. simplified [KKKA17]. Simulated [SMR16]. simulation [SRD^{+}16]. Simultaneous [CRAG^{+}12, TSK^{+}19]. simultaneously [KKT^{+}18a]. sinensis [DML^{+}16, DCZ^{+}17, DCL^{+}17, FGJ^{+}10, HGW^{+}19, SPP^{+}18, SCS^{+}12, ZB12]. Singapore [CWL^{+}17, GHH^{+}11]. Single [HH13b, BMMC^{+}15, CLGP15, CPB^{+}15, CZL^{+}19, FCL^{+}19, GM11, LWJ^{+}16, NCNB18, Sam10, VCOD17]. single- [GM11]. Single-dose [HH13b, Sam10]. Siniperca [KHC^{+}18, SWD^{+}14, TIK^{+}14a, TQW^{+}14]. Siphonostomatoida [KJB11]. sirindhornae [SHW^{+}11]. site [ESV^{+}19b, GVLE16, PPK^{+}13]. sites [Ano10c, HDWM15, HSA17, JTW^{+}10]. situ [ACSB^{+}19, CPEH15, DXX^{+}15, DCL^{+}17, vGJHS^{+}11, MTA^{+}15]. SJNNV [SOA^{+}18]. Skeletal [AAV^{+}16, ACA^{+}16, BHO^{+}19, BMT^{+}15, GRT12, RHVG^{+}19b, WHR^{+}18]. skeleton [DPD^{+}18]. Skin [LHW^{+}15b, WA14, CSS^{+}16, HMW11, JBO^{+}15, LJHG14, LHW^{+}12, MNF^{+}15, OLVJ^{+}13, ONA10, PWW19, PPK^{+}13, RSF^{+}17, SSF^{+}17, VCN^{+}10, VDD^{+}18, VKK^{+}19, XKB12, ZxYJ^{+}19]. skin11f [VKK^{+}19]. skyensis [RB15]. sleepy [ATJS^{+}18, WZL^{+}11]. SLICE(R) [PPI^{+}13, SMM^{+}13]. Slovakia [MKK^{+}11]. Slovene [JZK^{+}14]. Slovenia [GRB^{+}17]. small [WCW16a]. smallmouth [SBW^{+}18]. Smit [HSGD11]. smoked [TWNB^{+}12]. smolt [JOF^{+}19, KDA^{+}18, LAN^{+}14, MPS^{+}18, NB10, RSIP^{+}17, RSPA^{+}18, SFR^{+}17]. smolts [ECP^{+}10, FSV^{+}18, JCN^{+}13, JCN^{+}14, KAF^{+}19, KRFT14, LGJ19, MLF^{+}10, MID^{+}19a, PBMB11, SFU^{+}13]. smooth [HRNW15, MS16, XJL^{+}15]. smooth-hounds [MS16]. snakehead [CHH^{+}14, CLZ^{+}12, DIH^{+}16, LL12]. snapper [CLWJ13, HCW^{+}16, ZWY^{+}11]. SNPs [FCL^{+}19]. snubnose [WKW13]. sobria [CCMA^{+}17, FCC^{+}13, ZWL^{+}16]. social [RSF^{+}17]. sockeye [BLP10, LGJ19, GMC^{+}16a, TVK^{+}19]. sodium [HHO14]. soft [CIA^{+}18, KOS^{+}10, MMMD16, MMD16, PZL^{+}14]. soft-shell [MMD16]. soft-shelled [MMMD16]. sole [CVLJ^{+}10, CTD^{+}12, CPP14, CRS^{+}13, CSCCA14, FBL^{+}14, LVADB17, LPVGL^{+}10, MMS^{+}16, VFB^{+}12, XJL^{+}15, YB17, ZQB^{+}11, dALF^{+}17]. Solea [CVLJ^{+}10, CTD^{+}12, CPP14, CPdPP16, CRS^{+}13, CSCCA14, FBL^{+}14, LVADB17, MMS^{+}16, SOA^{+}18, VFB^{+}12, YMBE13, YB17, dALF^{+}17]. soleae [BVT^{+}18, LPVGL^{+}10, YMBE13]. Soleidae [YMBE13]. Solen [PDdC^{+}14]. solid [YSV^{+}13]. solidus [WFH^{+}18]. Sommerville [MRSC13]. sonophoresis [LRS^{+}17]. sorbent [YLX^{+}19a, ZLN^{+}17]. Sordariomycetes [ŘKH16]. source [OJRL10]. sources [CRO^{+}18]. South [HGPD11, HOHG13, HWO^{+}17, BHT^{+}17, GMP^{+}11, BHA^{+}11, CCW^{+}15, Dig13, Dig15, GFB^{+}13, LZGZ16, LHW^{+}15a, LXL^{+}18, MCCN14, Pol15]. south-east [GMP^{+}11, CCW^{+}15, Pol15]. south-western [BHT^{+}17]. southeast [RDN^{+}12, SDM^{+}19]. southern [BLSG^{+}16, BMGS17, CNN^{+}10, KYB^{+}15, MSMM19, SRT^{+}14, VDMJ^{+}13, WWS^{+}11, WZL^{+}11, YAHWU11, PRWG17]. soya [MTLP^{+}18b]. sp [BMMC^{+}15, CHH^{+}14, CPP14, DLPR15, DSLR16, EWH^{+}19, FCS^{+}18, FLC^{+}17, HGH^{+}16, HSHC^{+}16, LCOZ10, LJHG14, MdOO16, MSCH14, MAOS^{+}18, PMBM13, PMP^{+}18, PRY^{+}18, PLC^{+}19b, RRC^{+}17, SBM^{+}11, SDZN19, SPV^{+}18, SBK^{+}13, SBW11, TGJCH15, VMBGIC19, VDH19, YMBE13, ZWLG10, dSdAdS^{+}19, AOAH13, BAH^{+}18, VSS17]. spaC [PRY^{+}18]. spaC-type [PRY^{+}18]. spacer [DPMK12]. Spain [BSC^{+}14, DIVC16, RAVC18, DMP10, HGB^{+}11]. Spanish [RGMGG^{+}18]. spartansii [LF16]. Sparus [BC11, CVLJ^{+}10, CVGR^{+}13, CCME16, DdRTM13, DPCdR14, LGS^{+}14, MSS^{+}17b, RHVG^{+}19b, TPD^{+}19]. spat [BVL^{+}18]. Spatial [SBW^{+}18, PHR14]. Spatio [ASB^{+}12]. Spatio-temporal [ASB^{+}12]. spawning [BPPKJ10, BK12]. SPB [WKW13]. SPE0313 [HGW^{+}19]. special [VGI^{+}16]. species [ADMF15, ARFG13, BLL^{+}17, BWNS^{+}11, BOH^{+}12, DBV^{+}13, DCG18, DMK^{+}19, DSJA18, FBS13, FBAS16, GCNB16, GFB^{+}13, GOM^{+}19, GFS^{+}14, HY16, HZX^{+}13, IBFL19, JSJR^{+}19, KEY^{+}15, KBTMH^{+}16, KPKP10, KFK^{+}16, KDP^{+}17, LXL^{+}18, ME12, MSB^{+}19, MSS^{+}17a, MKK^{+}11, MPM16, MMW^{+}15, NSS^{+}17, NMM^{+}18, PSVK^{+}18, PSCFB^{+}18, SAS^{+}19, SN14, VMBGE17, VPP^{+}12, VDH19, ZZC^{+}17]. species-specific [IBFL19, VMBGE17]. specific [ACYC10, CLC^{+}17, CLGP15, CPB^{+}15, CSCCA14, IBFL19, KSS^{+}15, LCR^{+}10, MMBS10, SSV^{+}10, SKN^{+}16, TMM^{+}15, VMBGE17, YGL^{+}18, ZFZ^{+}11]. specifically [MLEM16, YLS^{+}19]. specificity [AdOG^{+}17]. spectroscopy [WNCR12]. speed [MTHB13]. sperm [VMCM14]. Sphaerospora [LXL^{+}18]. Sphyraena [GMP^{+}11]. Sphyrna [CBC^{+}13, MCWL13]. Spinal [GCWB11, CLHGK17, MLF^{+}10, MPS^{+}18]. spined [KFK^{+}16, PJZB19, WFH^{+}18]. spines [FvdMF^{+}18]. spiny [ZJDL18]. Spironucleus [MLWC11]. Spiroplasma [DXX^{+}15, HGW^{+}19, LXX^{+}17]. spirosulcatus [YYF^{+}10]. spleen [BSB^{+}18, BLdA^{+}18, GIW^{+}11, Lek12, LMM^{+}15, LMM^{+}16, RWTB17, SWD^{+}14, SSO^{+}14, TIK^{+}14a]. splendens [RhAS^{+}15]. splendidus [BT14, MBB12]. Splenic [AFA^{+}12, SKPS16, XJL^{+}15]. Spondylosis [BCG^{+}17]. Spontaneous [CGAS19, MLR^{+}12, RVB13, SMS^{+}19, dSRSHGá^{+}19]. spore [TES16]. spot [CNN^{+}10, CLW^{+}13, DLCC^{+}13, DZV10, DCZ^{+}17, DCMM13, DBM^{+}14, ESA^{+}10, FSO^{+}10, HVVV15, HZP^{+}12, JQK14, JGI^{+}10, JJS^{+}12, LQY11, MTK^{+}10, NSC^{+}18, PSK^{+}13, RNM^{+}12, RPGCIG17, RDN^{+}12, SMS^{+}10, STS^{+}15, SPTDEE^{+}15, SSV^{+}10, SP15, SVM^{+}18, SRD^{+}16, TMM^{+}15, TSK^{+}19, TTL^{+}16, TDT^{+}16, XKB12, YHY^{+}16]. spotted [FHZ^{+}16, KKK^{+}18, NNC^{+}18, SDJH16, WCL^{+}12, XJL^{+}19]. spp [BTN^{+}18, BHF13, BGC^{+}17, BBP18, CD17, CdAPdC^{+}19, GJMH17, HVVV15, ICVV16, JSJR^{+}18, KBTMH^{+}16, KFK^{+}16, KKKA17, MS16, NB11, PPSW19, PPE^{+}18, TES11, TES16, XLZ^{+}17]. Spraguea [FYO^{+}11]. Spraguidae [FYO^{+}11]. spread [RHS^{+}12, SJ11]. spreading [KAST17, TWJP10]. spring [EBM^{+}18, GSBEM15, LDEH17, LFL19, MPBM^{+}16, MGN^{+}16, OCLI^{+}19, SKN^{+}16, XWP^{+}18]. Sproston [GAC10, MWS17]. Squamous [SHAM14]. S(R) [CAS^{+}18]. sRNA [MFHHG18]. sRNAs [PKLH19]. ssDNA [YLS^{+}19]. ssp [BBP18, HCL^{+}14]. SSU [OBR^{+}14]. SSVR1601 [ZxYJ^{+}19]. stability [SG12, WGL17]. staff [EWRH13]. stage [FSB18a, MLER^{+}18, WOD16, YK13]. stages [AJLK^{+}18, EDN18, HGG^{+}11, JHGR13, LF18, MTA^{+}15, WFH^{+}18, dALF^{+}17]. standard [CDG^{+}14, GCW17, LWWS14, SEK^{+}16]. standardized [OLVJ^{+}13]. standards [LTG^{+}18]. Staphylococcus [LHW^{+}12, OAO^{+}12]. starry [LZGZ16]. starvation [CMEV17a, CZB^{+}11]. stasis [GM11]. state [FLCC^{+}14, GFS^{+}18, JJM^{+}17]. states [SPKS17, SSO^{+}14, CDM^{+}13, GCS^{+}18, PTM^{+}19]. Statistical [Gis18]. status [BSD^{+}18a, BCBS^{+}13, DSJA18, GJMH17, JOF^{+}19, KSS17, PFJ^{+}10, SMS^{+}10]. Steatitis [HGPD11]. steelhead [GTH11]. Steindachner [CSA^{+}15, KCD^{+}14, SWD^{+}14, SL15]. stellata [LZGZ16]. stellatus [dSdAdS^{+}19]. step [HRL15]. stickleback [ASW16, KFK^{+}16, WFH^{+}18, PJZB19]. stimulates [LFD^{+}17]. stimulation [TQW^{+}14]. stinger [MLMR16]. stinging [AA10]. stingray [CBFM16, CDM^{+}16, SDZN19, SCZM19, MdOO16]. stock [BGRB14]. stocked [GDN^{+}16, MP18]. stocking [KAST17]. stocks [AHJS^{+}18, OEG^{+}13]. storage [HPAL10]. stored [SG12]. strain [BVT^{+}18, CMSHS16, DKBS17, GCCLRP^{+}13, GGGM14, ITKH18, KF10b, KTN^{+}16, LZZN10, OEG^{+}13, RVC18, SVM^{+}18, SOGR^{+}19, TB12, VDD^{+}18, WKW^{+}11, YMH^{+}11, ZxYJ^{+}19]. strains [BLF^{+}16, CIL^{+}17, CLLF14, CRS^{+}13, dJDAVRGM^{+}18, FDL^{+}17, FCL^{+}19, GCMG^{+}11, GRF14, GRP^{+}15, LCV^{+}17, MBS^{+}19a, MNG^{+}18, OFA11, PKPG15, PFP^{+}13, SWM^{+}15, SOA^{+}18, TFG19, VTM^{+}14, VC17, WZA10, XKBC12, ZZC^{+}17]. strategies [SWWN^{+}16, ZCQF15]. strategy [FLC^{+}17, GCNB18, JMD^{+}18, LTC17, RRO^{+}13]. strawberry [LLS^{+}11, MTG^{+}10]. Streptocephalus [SHW^{+}11]. streptococcosis [FLCC^{+}14, LLH^{+}14a, ZTF^{+}10]. Streptococcus [AAZS^{+}15, AZSI^{+}12, AZSSZZ13, ATP^{+}17, BSD^{+}18b, BMMC^{+}15, BYAS^{+}19, BFFA^{+}12, BDB^{+}18, CWL^{+}17, CGCAH12, CLLF14, CZL^{+}19, DDBC^{+}15, DZC^{+}16, EPA^{+}10, GCMG^{+}11, HKK^{+}11b, HJOK11, HZX^{+}13, ICVV16, KNN19, LFSZ18, LLH^{+}14a, LLH^{+}14b, LLX^{+}19, LLL^{+}18a, MP18, MoSMS18, NLF11, NLKA16, NMIY14, NYRS12, OGI^{+}18, SLKE10, SFL^{+}16, SPS^{+}19, TWYC15, VMBGIC19, VJWC^{+}18, WZL13, XXL^{+}19, ZKL^{+}19, ZFZ^{+}11]. Stress [GAVC16, CZB^{+}11, DSSB19, MLER^{+}18, PBC15, PKLH19, TMC^{+}12]. stressed [CDG^{+}14]. stresses [HGX^{+}19]. stricta [BAH^{+}16]. strigosus [WA14]. strings [AHH16]. striped [ABC^{+}10, BVG^{+}10, CB13, SBM^{+}11, SMV^{+}14, VST^{+}19]. Strong [HLP^{+}16]. structural [LC17, SSV^{+}10]. structure [DPC^{+}10, ISCF^{+}19, LWZ^{+}18, MFS^{+}17, TPD^{+}19, VMCM14]. structures [DLP^{+}13, RHVG^{+}19b, RVC18, SMR16]. structuring [DJR^{+}19]. studied [GAC10]. Studies [DSA^{+}12, DAB^{+}12, ELI^{+}17, HMV^{+}17, LQY11, UKS^{+}18, CKJ^{+}14, CLHGK17, CTK^{+}10, GFW^{+}10, GRF14, KF10a, LTG^{+}18, LRMN17, MMS^{+}16, MMD16, MKMY10, RBAT^{+}17, SRD^{+}16, WKW^{+}11]. Study [DLLB^{+}18, JMS13, LBF^{+}14, AHJS^{+}18, AOAH13, AAUHW15, ABT^{+}14, BSPS^{+}18, BCM^{+}17, BMGS17, CXK16, CKvGJB17, CGCAH12, DPTFC10, DPTCF10, DIVC16, DZV10, FLVA10, GFM^{+}11, JWO^{+}10, JR19, JHRW17, KPT^{+}11, LVADB17, MVM^{+}18, MCSM17, MID^{+}19a, NB10, QXW14, RHZA^{+}16, RHST^{+}16, RHSM^{+}18b, RHSM^{+}18a, Sam10, SWM^{+}15, SVL^{+}17, SVV^{+}17, VTM^{+}14, VBL^{+}15]. sturgeon [BLL^{+}17, DKBS17, LGK^{+}14, MMD^{+}12, RHSM^{+}18b, RFA^{+}14, SAD^{+}12]. subclinical [BSSH17, LMG^{+}19]. subclinically [Bin14]. subinhibitory [GTGCM19]. sublethal [WNO^{+}18]. subpopulations [TDLT^{+}16, TDT^{+}16]. subsequent [IPW^{+}13]. subsp [ATP^{+}17, SPKS17, YFA^{+}17]. subsp. [CDM^{+}13, CDF^{+}18, CRS^{+}13, DABO^{+}17, KPSP^{+}19, KBTMH^{+}16, MLEM16, NLC^{+}16, OJRL10, PLC^{+}19b, SWM^{+}15, SPS^{+}17, SKGE13]. subspecies [GLK^{+}16, VC17]. substances [YFA^{+}17]. subtractive [LZZN10]. subtype [GL18, GRF14, HST^{+}13, JJB15, RSP^{+}17, RSM^{+}18, TJB^{+}15]. subtypes [GFM^{+}11, GFR^{+}12, GLK^{+}16]. subunit [ÇOÇ^{+}16, LSA^{+}15, LXHL11]. success [ACN12]. succinicans [GDW^{+}15]. sucker [MGN^{+}16, JMHK18]. suckers [BWB^{+}17]. Sueur [WKW^{+}11]. SufB [PLCC11]. suffering [ATJS^{+}18, GRT12, GFS^{+}14, TvWO^{+}16]. suggest [GB18, TALA11]. suggests [WXL16]. suicidal [CLGP15]. suitability [NJK^{+}17]. suitable [JKS^{+}13, KHG^{+}17, VLVS^{+}16]. sulfate [ZCJ^{+}18]. sulfate-resistant [ZCJ^{+}18]. sulfurea [ŘKH16]. sulphate [DBS12]. sultana [KAESEM13]. summary [JJM^{+}17]. sunfish [ALG15, LDEH17]. sunshine [DBS12]. Superior [HAP^{+}10]. supernatants [GM10]. supplementation [JKL^{+}13, KSM^{+}17, LKS^{+}13]. supplements [AMA11, NFA15]. support [BMBS12]. supports [GCS^{+}18]. suppression [LZZN10]. suratensis [MRN^{+}19]. surf [RPJ^{+}18]. Surface [GCMG^{+}11, CD17, NMIY14, OBD10]. surfaces [HRNW15, PDLW17]. surfactant [AJLK^{+}18]. surgeonfish [WA14]. Surgical [LL11, GM11]. surveillance [BTN^{+}18, BMBS12, GWM12, GCS^{+}18, HAS^{+}19, KOS^{+}10, LTM^{+}19, LMAKN16, MMBS10]. surveilling [HRL15]. Survey [MMG^{+}11, BHT^{+}17, GGMG18b, HSA17, JZK^{+}14, KZJ^{+}17, LUP^{+}18, MSBB^{+}16, MJ17, WDM^{+}15]. survival [DMB^{+}19, FFP^{+}14, LTC17, LGS^{+}14, MMS^{+}16, SPCR^{+}19, SFU^{+}13, WYZ^{+}19]. survived [WIC^{+}19, WCL^{+}12]. Susceptibility [BSK^{+}10, BDD^{+}18, GOM^{+}19, IOO15, NSS^{+}17, RWTB17, XKBC12, CLL^{+}10, CBM^{+}17, DBV^{+}13, DCMM13, GCW17, GMC^{+}16b, GRP^{+}15, GHW^{+}15, HKB^{+}16, JKH^{+}17, KKOY19, KBE^{+}14, ME12, MSS^{+}17a, NSB^{+}18, OEG^{+}13, PWHW18b, PFJ^{+}10, RHTC19b, WWS^{+}11, WKW^{+}11, WKW13, Wen16, XWP^{+}18, XKB12, YNL11, YCNW18]. susceptible [GHH^{+}11, IFF^{+}15, KWD^{+}16, RSP^{+}17, TDS^{+}19, UMMS17]. Suspected [KVC^{+}16, MSS^{+}12, LXL^{+}18]. Susquehanna [SBW^{+}18]. swabs [ZCCN13]. swamp [AMAR18]. swamps [HHC^{+}18]. Sweet [LFD^{+}17]. sweetfish [KSS^{+}15]. Swim [ŘKH18, BMD^{+}14, BCG^{+}17]. swimbladder [FDST16]. swimming [DSSB19, GMC^{+}16b, HKF^{+}16]. switch [CKJ^{+}14]. SYBR [TCFáS19]. Symphodus [SRM^{+}18]. Symphysodon [AFTEML19, RKK^{+}19]. Syncytial [FKB^{+}14, KAS^{+}18]. syndrome [APG^{+}11, BPK^{+}17, CSS^{+}16, CNN^{+}10, CLW^{+}13, DLCC^{+}13, DZV10, DCZ^{+}17, DCMM13, DBM^{+}14, ESA^{+}10, FSO^{+}10, FPK^{+}17, GFS^{+}18, GKCC^{+}12, HVVV15, HZP^{+}12, HHC^{+}18, JQK14, JYZ^{+}14, JGI^{+}10, JJS^{+}12, KEY^{+}15, KOS^{+}10, LQY11, LWMT10, MTK^{+}10, MTG^{+}10, NSC^{+}18, OOO^{+}11, OAO^{+}12, PZL^{+}14, PSK^{+}13, RNM^{+}12, RPGCIG17, RDN^{+}12, RMR14, SPTDEE^{+}15, SSV^{+}10, SP15, SVM^{+}18, SRD^{+}16, SNK^{+}19, TMM^{+}15, TSK^{+}19, TTL^{+}16, TDT^{+}16, WNLF^{+}12, YHY^{+}16, ZWL^{+}16, GRB^{+}17, GMB^{+}17, GMC^{+}16a]. syndrome-associated [JYZ^{+}14]. syndrome/acute [BPK^{+}17]. synergism [ZN10]. Synergistic [LGJ19]. Syngnathidae [FFLC17]. synthase [PLCC11]. synthesis [SRB12]. system [AGFG19, ACM^{+}14, CBC^{+}13, DPTFC10, DLLB^{+}18, EAN^{+}18, FYO^{+}11, JJS^{+}12, JSAB^{+}19, áKÁGA16, KDP^{+}17, LP17, MSG^{+}18, PHW^{+}16, XXL^{+}19]. systematic [GCS^{+}18, YGL^{+}18]. Systemic [CGAS19, CDD^{+}18, EKS^{+}13, NOV^{+}11, RSR^{+}18, RPM^{+}19, CPP14, DLP^{+}13, LCOZ10, PRY^{+}18, TCC10a]. systems [BSD^{+}18b, BTN^{+}18, BGD^{+}17, GDW^{+}15, LRS^{+}17, RSB13].

T [BA14, NYRS12, TQW^{+}14]. T-cell [NYRS12, TQW^{+}14]. T. [LUP^{+}18]. taeniolatus [LYD^{+}15]. Taeniura [CDM^{+}16]. tail [DCV^{+}14, MSC^{+}17, YNL11, ZJDL18]. tailed [DD15]. Taiwan [CCT^{+}11, CLW^{+}13, PLC^{+}19b]. Takahashi [ZCC18]. Takifugu [KOI^{+}14, MKMY10]. tambaqui [PdSG^{+}12]. tandem [ACAS14, HZP^{+}12, JGI^{+}10]. tank [GAVC16, PSS17]. tank-reared [PSS17]. tannic [AMRHM18]. tapetis [VDD^{+}18]. TaqMan [FSO^{+}10, LGC^{+}19, BMBS12, FGR^{+}12, JKS^{+}13, ZHFC11]. Taqman-based [JKS^{+}13]. tarda [CTD^{+}12, COB^{+}16, KTK17, KREM14, LKS^{+}13, LML^{+}13, QXW14, RSR^{+}18, SMV^{+}14, WZA10, ZXW^{+}12]. target [MSS^{+}17a]. targeted [GWM12, MPJ^{+}15]. targeting [CXW^{+}19, HLP^{+}16, vGJKR^{+}17]. tarpon [BCW^{+}15]. Tasmania [ABC^{+}10]. Tasmanian [SN14, ZCCN13, ZCC^{+}15]. Tasoh [ARR^{+}18]. Taura [APG^{+}11]. taurus [LL11]. Technical [BHA^{+}17]. technique [FNISE12, MTHB13, PSVK^{+}18, TMWV14]. techniques [CPdPP16, MFES18, TCS19]. Telangiectatic [RHSM^{+}18a]. teleost [MYWC18, SdMVS^{+}16, ZGY^{+}18]. Teleostei [Lek12, RRC^{+}17]. teleosts [BDM^{+}18, SCD19]. Temminck [CONN10, FNISE12, GHH^{+}11, HKK^{+}11b, JKC^{+}11, JJV^{+}15, KSS^{+}15, KYK^{+}12, KLW10, LKS^{+}13, MKK^{+}11, MKMY10, NSS^{+}17, OMO^{+}15, SSJ^{+}19, YYF^{+}10, ZCC18]. Temperature [VCLW16, AAV^{+}16, ASSP^{+}15, BSPS^{+}18, BPCS19, Bin14, CRBLC^{+}17, DSA^{+}12, EDV^{+}18, FFP^{+}14, GHB^{+}11, HMW11, HZJ^{+}19, IFF^{+}15, JBO^{+}15, JJV^{+}15, KCL14, MPS^{+}18, PMPG^{+}16, SFR^{+}17, SK15, SJK^{+}14, VBL^{+}15]. temperature-dependent [ASSP^{+}15, HZJ^{+}19]. temperature-influenced [BSPS^{+}18]. temperature-related [JBO^{+}15]. temperatures [FHFF15, GMN12, JPG^{+}19, RHTC19b, SVL^{+}17, VBL^{+}16]. Temporal [PHR14, ASB^{+}12, DIVC16, SBW^{+}18, SVV^{+}17]. Temporary [AHPR^{+}16]. ten [VTM^{+}14]. Tenacibaculosis [MMS^{+}16, FBL^{+}14]. Tenacibaculum [AHITC18, BMH^{+}19, BVT^{+}18, CSCCA14, FSG^{+}12, FSV^{+}18, FSB^{+}18c, IGLG^{+}17, KAF^{+}19, LITCAH19b, LPVGL^{+}10, MMS^{+}16, VFB^{+}12]. tench [DLP^{+}13]. tentative [VDH19]. tequila [RAM^{+}17]. teratoma [RAM^{+}17]. term [DIVC16, KHG^{+}17, LTC17, MSBB^{+}16]. ternetzi [SHAM14]. terraenovae [BA13]. territorial [BCL^{+}19]. test [BTC^{+}17, CLC^{+}17, DLLB^{+}18, HWM^{+}13, JWO^{+}10, SEK^{+}16, ZWL^{+}14]. Testate [DT16]. Testicular [Řeh13]. testing [BVK^{+}16, EAM^{+}13, GCW17, HH13b, LTM^{+}19, PPE^{+}18, WGL17]. tests [LTG^{+}18, OTN^{+}12, TAC15]. testudineus [LJHG14]. tetra [BE19, LLG^{+}15, Lek12, ŘKH17, ŘKH18, SHAM14, SMS^{+}19]. Tetracapsuloides [BSPS^{+}18, BPPKJ10, JZK^{+}14, KAESEM13, MJ17, Mor12]. Tetradecylthioacetic [GRT12]. Tetrahymena [LCOZ10, SPLV^{+}15, XZZ16]. Tetramicra [SFBW^{+}17]. tetraspanins [WLX^{+}10]. TH0426 [YZW^{+}19]. Th1 [KBD13, CKJ^{+}14]. Th1- [CKJ^{+}14]. Th1-/ [CKJ^{+}14]. Th1/Th2/Treg [KBD13]. Th2 [KBD13, CKJ^{+}14]. Th2-like [CKJ^{+}14]. Thailand [BPK^{+}17, DLPR15, KTP^{+}15, KYB^{+}15, SHW^{+}11, SHS14, STG^{+}10, SRT^{+}14]. thailandensis [SHW^{+}11]. their [AHH16, BCLM^{+}19, CPdPP16, DSSB18, DPC^{+}10, HGH^{+}18, HB12, JGI^{+}10, JSJR^{+}19, KBE^{+}14, NFA15, PPF^{+}12, RAS^{+}10, SWM^{+}15, TSK^{+}19, TQW^{+}14, WNMK^{+}16]. Thelohanellus [LJHG14]. them [JMD^{+}18]. theory [TALA11]. Theragra [WMF14]. therapeutic [BE19, MSCH14]. therapy [FLCC^{+}14]. thermoadapted [RPM^{+}19]. thermophila [XZZ16]. these [BSD^{+}18a]. thickening [DRdS^{+}10]. thicker [SWS^{+}16]. thicklip [HSGD11]. Thoracocharax [dSdAdS^{+}19]. Three [KTK17, WLX^{+}10, ABVS18, KKT^{+}18b, KFK^{+}16, MTHB13, OEG^{+}13, SGB^{+}16, SPKS17, SN14, WFH^{+}18, XKBC12, ZWG^{+}17, PJZB19]. Three-dimensional [KTK17]. three-spined [KFK^{+}16, WFH^{+}18, PJZB19]. threespine [ASW16]. thromboticans [WEPB^{+}18]. throughout [TWJP10]. throughput [BHT^{+}17, LITCAH19b, TCFáS19]. Thunberg [GVLE16, GCCLRP^{+}13, HPY^{+}13, OTN^{+}12, OKS^{+}13]. Thunnus [CZB^{+}11, PSZ^{+}14]. thymic [KVC^{+}16]. thynnus [CZB^{+}11, PSZ^{+}14]. thyrsites [MSBB^{+}16]. tiburo [CBC^{+}13, MCWL13]. tidal [PHR14]. tidepool [SPTDEE^{+}15]. tigecycline [CDW17]. tiger [DCMM13, DBM^{+}14, KOI^{+}14, LL11, MKMY10, NDA^{+}19, PSK^{+}13, SRSF10]. tilapia [ARR^{+}18, AZSSZZ13, ACLP^{+}19, ATP^{+}17, AdOG^{+}17, BMMC^{+}15, CMS^{+}19, CCHL13, CLLF14, CZL^{+}19, DLPR15, DSLR16, DTJ^{+}17, FLCC^{+}14, FKB^{+}14, HZX^{+}13, ICVV16, JST^{+}19, KAS^{+}18, KTP^{+}15, KBTMH^{+}16, LLH^{+}14a, LCZ^{+}18, LLX^{+}19, LLL^{+}18a, MCW^{+}18, MNM^{+}18, NMV^{+}19, NLF11, NYRS12, OGI^{+}18, PNBK13, PCK^{+}15, PYSS19, PLC^{+}19b, RHTC19b, RPMRLJ^{+}19, SPKS17, SLKE10, SL15, SKGE13, SFL^{+}16, SPS^{+}19, TSS18, TRS19, TvWO^{+}16, VMBGIC19, VJWC^{+}18, WWZ^{+}18, WSA^{+}18, XXL^{+}19, YSHG18, YTKS19, YLL^{+}18, YWW^{+}19, ZKL^{+}19, ZTF^{+}10, vD14, NFF^{+}17]. Tilesius [GM11]. TiLV [ARR^{+}18, MCW^{+}18]. Timah [ARR^{+}18]. Time [JMF^{+}18, ATM^{+}10, AGB^{+}19, AdOG^{+}17, BKS^{+}17, BKD17, BMBS12, CDF^{+}18, CDG^{+}14, DCL^{+}17, FSO^{+}10, FGR^{+}12, FSG^{+}12, GBV^{+}16, GJK^{+}11, GTH11, GBSF12, HRL15, HGB^{+}11, JKS^{+}13, JNN^{+}11, KJW^{+}12, KBJ^{+}13, LGC^{+}19, MMBS10, MSI^{+}15, MLWC11, NLKA16, SK15, TCFáS19, YTL10, ZHFC11, vBVLR^{+}16]. time-monitoring [AGB^{+}19]. timing [GVB^{+}17a]. Tinca [DLP^{+}13]. tips [DCV^{+}14]. Tissue [AOAH13, MRR^{+}14, BKS^{+}17, BTC^{+}17, CIA^{+}18, DdRTM13, HSHC^{+}16, LAN^{+}14, LRD^{+}19, LJC^{+}19, MB17, OKS^{+}13, PAM^{+}19, SSV^{+}17, SG12, SFL^{+}16, ZYW^{+}10]. tissues [CDF^{+}18, FBL^{+}14, FGR^{+}12, GMB^{+}17, GM10, GGW^{+}11, OBD^{+}11, RNM^{+}12, SSHEM11]. titration [CCCTA^{+}14]. titre [OKS^{+}13]. TLR1 [SHKF16]. TLR2 [ZZQ^{+}19]. TLR22 [SHKF16]. TLR5 [QZZ^{+}18b]. TLR5M [SHKF16]. TLR5S [SHKF16]. Todd [JCN^{+}14]. TOF [PSVK^{+}18, PSCFB^{+}18]. tolerance [CDW17, JHGR13, SW11, WWE^{+}13]. toll [PRSK10, SHKF16]. toll-like [PRSK10, SHKF16]. tomentosa [KWGS19, MWRSK17, MKK^{+}11]. tomographic [LTC^{+}14]. tongue [XJL^{+}15, ZQB^{+}11]. tool [HH13b, Sla14]. tooth [OAO^{+}12]. top [EAM^{+}13]. topmouth [MKK^{+}11]. Tor [SM14]. Tosh [EW17]. Toshio [EW17]. Toxicity [PNBK13, MGF17, PCK^{+}15, SRB12, ZPM^{+}18]. toxicology [GHW^{+}15]. toxicosis [DGS^{+}14, MSS^{+}12]. toxin [CCLGRM^{+}17, NPGA^{+}19]. toxoid [SWM^{+}15]. Toxoplasma [SZM^{+}15]. Trachinotus [AZSI^{+}12, CWL^{+}17, LYL^{+}18, WKW13]. Trachurus [ONGQ^{+}14]. tracing [KDA^{+}18]. tracking [EKT^{+}11]. tract [ABF15, BT14, BLFC12, MBB12]. Trade [RMS^{+}13, CDM^{+}13]. traded [AHJS^{+}18]. trafficked [ARFG13]. traits [KNN19]. transcribed [DPMK12]. transcription [FVAVMGE18, GRT12, TSS18, YWW^{+}19]. transcriptional [LHW^{+}15b, TQW^{+}14]. transcriptome [FFLC17, PKLH19, VMBGE17, XJL^{+}19, ZJR^{+}18]. Transcriptomic [CRBLC^{+}17]. transcriptomics [JBO^{+}15]. transfection [CLL^{+}10, WKW13]. transfer [BLFRCQ10, JMF^{+}18, RHS^{+}12]. transform [WNCR12]. translationally [NRMS14]. transmissible [GMC^{+}16a]. Transmission [AZSSZZ13, BWS^{+}18, CVGR^{+}13, CBW19, FBAS16, HVVV15, HKG^{+}15, SVY^{+}14, AHJS^{+}18, BMK17, FBS13, FYO^{+}11, HGG^{+}11, JNS^{+}19, KS19, KSA^{+}10, MMD16, NLC^{+}16, PMPG^{+}16, RMS^{+}13, SHJ^{+}12, STIG17, SHV^{+}16, SMKBP17, TRS19, WMM^{+}16, WKK12, YTKS19, YLL^{+}18]. transmit [WNGS^{+}17]. transmitting [CJB^{+}15]. transport [SMR16, SVY^{+}14]. transporter [OJRL10, WZL13]. transversum [FFP^{+}14]. treat [ACN12, AMRHM18, LML12]. treating [KWGS19]. Treatment [LCOZ10, PB18, TNK12, BCT^{+}17, BA14, CDW17, CB13, FVAVMGE18, GBV^{+}16, GVB^{+}17b, HHO14, JML19, KBE^{+}14, MGM^{+}18, MSCH14, RHS^{+}12, RSH11, SMM^{+}13, SFF^{+}17, VGI^{+}16, VME^{+}19, ZTF^{+}10]. treatments [BNS13, CDW17, GVB^{+}17a]. Treg [KBD13]. Trematoda [BCM^{+}17, PMP^{+}18]. triactinomyxon [FCJH19]. trial [BVK^{+}16, BSBH10, CCH^{+}19]. tributaries [BWB^{+}17]. tributary [MKKM15]. Trichodina [MUM^{+}19]. Trichosporon [AMAR18]. tricolour [RTM^{+}15]. trigger [Bin14]. triggers [SJK^{+}14]. triple [PPI^{+}13, SFF^{+}17]. triple-dose [PPI^{+}13]. triploid [ACA^{+}16, AAV^{+}16, GCWB11, LWZ^{+}18, SFR^{+}17]. triploidy [FHFF15, FMH^{+}15]. trophonts [vGJHS^{+}11]. trophozoites [WNMK^{+}16]. tropical [VMCM14]. tropism [ADF14, ADMF15, SFL^{+}16]. trout [ASA10, ASHZ^{+}12, AOAH13, AHHR^{+}10, AHPR^{+}16, AJLC^{+}16, AGFG19, ABT^{+}14, AA10, AMA11, BSW17, BSPS^{+}18, BCLM^{+}19, BKD17, BHA^{+}11, BE19, BA14, BLFC12, BAH^{+}18, CSS^{+}16, CGW^{+}19, CHAB10, CKJ^{+}14, CMH^{+}17, CAJD^{+}18, CPB^{+}15, CMR^{+}16, CLS^{+}15, DS10, DPTFC10, DPTCF10, DMP10, DSSB18, DSSB19, DJR^{+}19, DdRTM13, DdRCP13, DPADO16, DGS^{+}14, DSJA18, DT16, GRB^{+}17, GMB^{+}17, GFO^{+}13, GCNB16, GTH11, GPFGAG19, GDW^{+}15, GMC^{+}16b, HGH^{+}18, HKBD15a, HKBD15b, HJN^{+}19, IOO15, JKL^{+}13, JAJC^{+}18, JDH^{+}18, JML19, vGJHS^{+}11, vGJKR^{+}17, KSM^{+}17, KEY^{+}15, KHG^{+}17, KFK^{+}16, LPA^{+}10, LP17, LUP^{+}18, LLS^{+}11, MBOC19, MNF^{+}15, MSC^{+}17, MCSM17, MHNY18, Mor12, MTLP^{+}18b, MNM^{+}18, MNM^{+}19, MKKM15, NB11, NOV^{+}11, NDA10, OP14, OLVJ^{+}13, OBFR12, OBD^{+}11, OSC^{+}14, OSAA18, OLT^{+}10, PDLW17, PMW^{+}10, PMH^{+}11, PRT^{+}14, PSBG^{+}11, PGF^{+}19, PLCC11, PPE^{+}18, QPB^{+}16]. trout [RSB13, ŘKH16, RCGGM15, RBL^{+}11, RHB^{+}15, SHJ^{+}12, SPCR^{+}19, SCR^{+}15, SVV^{+}17, SON^{+}18, SSC17, VSS17, XKB12, YDO^{+}16]. trumpeter [ABC^{+}10, CB13]. trutta [BSPS^{+}18, AGFG19, GFO^{+}13, GMG^{+}19, OSC^{+}14, PGF^{+}19]. truttae [MUM^{+}19]. Trypanorhyncha [BB11]. tshawytscha [BHT^{+}17, BMH^{+}19, DWP^{+}19, DPD^{+}18, GMC^{+}16a, GBO^{+}18, GTH11, LST^{+}12, LF16, LF18, LWMT10, PMPG^{+}16, STF^{+}15]. TSV [APG^{+}11]. Tu [PLCC11]. tubiashii [DABO^{+}17]. tumour [NRMS14, PHR14, Řeh13]. Tumours [BWB^{+}17]. tuna [CZB^{+}11, PSZ^{+}14]. Tunic [KOS^{+}10, PZL^{+}14]. Tunisia [YMBE13, YB17]. Tunisian [CGG^{+}10]. turbot [COB^{+}16, HMW11, JCS11, LBF^{+}14, MBB12, MDF^{+}17, QXW14, SODB16, WM19, dOAPLá^{+}17]. Turkey [SDWA18]. Turkish [BAH^{+}18]. turpisrotundus [ZWLG10]. tuvira [VMCM14]. twin [LTC^{+}14]. twins [ADR11, BQP12]. Two [CdAPdC^{+}19, MS16, XZZ16, YSV^{+}13, BSPS^{+}18, BDM^{+}18, CRS^{+}13, DCG18, FLVA10, GPFGAG19, GFW^{+}10, HKK^{+}11b, HRL15, HLC^{+}15, HSHC^{+}16, LLL^{+}12, MFHHG18, RAM^{+}18, RB15, RTM^{+}15, SQC^{+}14, VDHV^{+}16]. two-dimensional [LLL^{+}12]. two-step [HRL15]. type [DJR^{+}19, EAN^{+}18, HBVB16, KHC^{+}18, LGS^{+}14, MSG^{+}18, PBC15, PMW^{+}10, PMH^{+}11, PRY^{+}18, VJWC^{+}18, ZWG^{+}17]. types [DLP^{+}13, VCOD17, XKB12]. typing [GLK^{+}16, ISCF^{+}19, RDW18, TCS19, VLVS^{+}16]. Tyrrhenian [GFB^{+}13].

U [PMW^{+}10, PMH^{+}11]. U-type [PMH^{+}11]. Uganda [WMD^{+}18]. ulcer [SSF^{+}17, ZxYJ^{+}19]. ulcerations [VDD^{+}18]. Ulcerative [RBS^{+}16, HHC^{+}18, KOB^{+}17, KPT^{+}11, KRD^{+}17, SJT14, VFB^{+}12, ZWL^{+}16]. ultrasonographical [RHST^{+}16]. Ultrastructural [GMB^{+}17, HFT^{+}12, MBT^{+}17, DCG18, SMS^{+}19]. ultrastructure [CHWM^{+}19, GFS^{+}14, LvD18]. Ultraviolet [CB13, HPAL10, MUM^{+}19]. unaffected [JOF^{+}19]. Uncovering [VMBGE17]. underlines [VC17]. underlying [JBO^{+}15]. undescribed [GLK^{+}16]. undosquamis [KSAB^{+}14]. UNESCO [GVLE16]. Unexpected [HSGD11, LHW^{+}15a]. United [CDM^{+}13, GCS^{+}18, NSB^{+}18, PTM^{+}19]. units [HZP^{+}12]. unusual [QXW14, RVC18]. unusually [MLWC11]. Updated [YLW^{+}14]. upon [TTL^{+}16]. upper [LEM^{+}19, SWS^{+}16, WBE^{+}17, HHC^{+}18, JMHK18]. uptake [RRO^{+}13, TDLT^{+}16]. Urtica [AA10]. Uruguay [PAZ18]. USA [PHR14, CGAS19, FSM^{+}16, MTG^{+}10, PBBG19]. usage [HBS^{+}14, MLH^{+}15]. Use [AA10, CMEV17b, GCGR13, SMM^{+}13, ABF15, BW18, CHV^{+}12, GTH11, JSJR^{+}18, LTM^{+}19, LLDL13, MVM^{+}18, RSF^{+}17, SVL^{+}17, SFF^{+}17, TMWV14, VBL^{+}15]. used [BNS13, DZV10, VME^{+}19, WDM^{+}15]. uses [CRO^{+}18]. Using [BWM^{+}13, SDI^{+}18, ATM^{+}10, ACN12, ASHZ^{+}12, AGFG19, BE19, CMEV17a, CHB^{+}17, DLLB^{+}18, DWL^{+}17, DXX^{+}15, FCM^{+}19, FSV^{+}18, GPFGAG19, GFM^{+}11, ITKH18, JYZ^{+}14, KTN^{+}16, KKT^{+}18b, KKT^{+}18a, KSA^{+}10, LZZN10, LFH^{+}11, LCR^{+}10, LJGRI^{+}10, MSI^{+}15, MID^{+}19a, MUM^{+}19, MTHB13, MTA^{+}15, OWM^{+}17, RBAT^{+}17, SLKE10, TCS19, VCOD17, WCW16a, WNLF^{+}12, XZZ16, YGL^{+}18, YkSKY13]. Utilization [ZYW^{+}10, BWM^{+}13, EKT^{+}11, OJRL10]. utilizing [ZFZ^{+}11]. Uttarakhand [SM14].

V. [ZZC^{+}17]. vaccinated [MSC^{+}17, PBMB11]. Vaccination [EDV^{+}18, FMH^{+}15, KBD13, ASHZ^{+}12, ESV^{+}19b, JAJC^{+}18, LAN^{+}14, MBS^{+}19a, MID^{+}19b, NCNB18]. vaccine [BVK^{+}16, BSBH10, CLWJ13, CLGP15, CCH^{+}19, CGR17, ÇOÇ^{+}16, CLS^{+}15, DMK^{+}19, ESV^{+}19b, FLC^{+}17, GHB^{+}11, HLL^{+}12, HJN^{+}19, JKC^{+}11, JOF^{+}19, vGJKR^{+}17, LPA^{+}10, LLDL13, LLX^{+}19, LXHL11, MBS^{+}19a, MSC^{+}17, RPMRLJ^{+}19, RSGB^{+}14, SCK13, SLKE10, WHR^{+}18, ZWL^{+}14]. vaccine-induced [GHB^{+}11]. vaccine-mediated [JOF^{+}19]. vaccines [Dal18, ELI^{+}17, NFA15, NNC^{+}18, ZXW^{+}12]. vacuoles [LTC16]. Vagococcus [TCFáS19, TCS19, YDO^{+}16]. Vahl [LZGZ16]. Valenciennes [BCW^{+}15, GMP^{+}11, GGW^{+}11, GCWB11, HLL^{+}12, HCEH16, KHS11, LLL^{+}12, LJGRI^{+}10, LXHL11, MVWS13, NWMK16, PMBM13, SS17, XKBC12, YYW^{+}14]. Validation [BWZ^{+}17, CLC^{+}17, GM10, LPCC12, BMBS12, ÇOÇ^{+}16, EAM^{+}13, JKS^{+}13, KJW^{+}12, KBJ^{+}13, TSS18, VCOD17, vBVLR^{+}16]. values [GCW17, SEK^{+}16]. valvularis [WEPB^{+}18]. vannamei [BTN^{+}18, CMEV17a, CMS^{+}17, CMEV17b, CLW^{+}13, EEMCEGSP16, HMV^{+}17, HXZ^{+}16, JQK14, KS19, NNCLGC^{+}11, RPGCIG17, SSV^{+}17, STG^{+}10, SNM^{+}16, SVM^{+}18, TMM^{+}15, TSK^{+}19, TTL^{+}16, TDLT^{+}16, TDT^{+}16]. vapA [GLK^{+}16]. Var [BVL^{+}18, GVLE16]. variability [SBW^{+}18, VCS^{+}19]. variable [ACAS14, HZP^{+}12, PPF^{+}12]. variant [CKTT10, LDA^{+}19, QXW14]. variants [BVL^{+}18, HBVB16, MFHHG18, TOK^{+}16]. Variation [GBV^{+}16, VDHV^{+}16, DZV10, LTC17, WNAF^{+}13]. variations [HBVB16, JJB15, JGI^{+}10]. various [BLL^{+}17]. vary [GRP^{+}15]. vascular [RSB13]. vector [MSB^{+}19, NLC^{+}16, SHV^{+}16]. vein [ASW16]. ventral [DKG^{+}17]. ventricosa [RPJ^{+}18]. verification [ŘKH18, SHW^{+}11]. veronii [ZWL^{+}16, CHH^{+}14, CYS17, DTJ^{+}17, FMA^{+}19, MP18, RKK^{+}19, YZW^{+}19]. Verrucous [RKPG16]. Vertebral [DWP^{+}19, JHM^{+}18, FHFF15, GHB^{+}11, LWZ^{+}18, LGS^{+}14, MLF^{+}10, dALF^{+}17]. vertical [JNS^{+}19, KSA^{+}10, NRF^{+}18, YTKS19]. VERV [GFB^{+}13]. VHS [GVC^{+}19, HLL^{+}12, HAS^{+}19]. VHSV [GMN12, GGW^{+}11, GRP^{+}15, MMBS10, ZLX^{+}19]. VHSV-IVb [GMN12]. via [KJW^{+}17, PBM^{+}15, RPMRLJ^{+}19, TRS19, DDBC^{+}15]. viability [CAS^{+}18, FCJH19, HPAL10, MGM^{+}18, VCLW16]. Vibrio [AHAMR^{+}19, BTN^{+}18, BVK^{+}16, BT14, BGC^{+}17, BBP18, BFC^{+}15, CLWJ13, CCLGRM^{+}17, CCH^{+}19, CLW^{+}11, CKTT10, CDG^{+}14, DD18, DJQ^{+}12, DABO^{+}17, FLVA10, FMB^{+}11, HRL15, HBVB16, HXZ^{+}16, HGX^{+}19, ITKH18, JQK14, KTN^{+}16, KYB^{+}15, LLDL13, LSC^{+}14, LSA^{+}15, LMM^{+}15, LMM^{+}16, MBB12, MKMY10, NNC^{+}18, NDA^{+}19, PKLH19, QWTL17, QZZ^{+}18b, RGD^{+}12, RRO^{+}13, RCD^{+}17, SSF^{+}17, SNM^{+}16, SCS^{+}12, TMC^{+}12, VDD^{+}18, WBF^{+}10, WZQ^{+}16, WM19, YGL^{+}18, YLX^{+}19a, YLS^{+}19, ZZC^{+}17, ZWY^{+}11, ZQB^{+}11, ZJR^{+}18, ZxYJ^{+}19, ZXW^{+}12, ZDWH18]. vibrioferrin [COB^{+}16]. Vibrionaceae [NMM^{+}18]. Vibrios [PDdC^{+}14, HBH^{+}19]. vibriosis [KSS17, TNK12]. Vibrogen [CCH^{+}19]. Vibrogen-2 [CCH^{+}19]. Victoria [MCW^{+}18]. Vietnam [BAD^{+}12]. Vietnamese [CTK^{+}10]. Viola [YLX^{+}19b]. viraemia [EBM^{+}18, GSBEM15, MPBM^{+}16, OCLI^{+}19, SKN^{+}16]. Viral [AHHR^{+}10, DVC^{+}17, GGW^{+}11, HPAL10, VPP^{+}12, WKW13, WNAJ^{+}16, AHPR^{+}16, Ano12c, BRMG13, Bin14, DSA^{+}12, DAB^{+}12, DPADO16, EKED^{+}17, GJVT^{+}11, GFB^{+}19, GFB^{+}13, GMN12, GFW^{+}10, GVC^{+}19, HLL^{+}12, HKD^{+}10, HGG^{+}11, JQK14, JKS^{+}13, KKK15, KF10a, KF10b, KOS^{+}10, LRD^{+}19, LFH^{+}11, LVADB17, LNC^{+}17, MMBS10, MDC^{+}18, MF12, MMW^{+}15, NOST^{+}16, OEG^{+}13, PBM^{+}15, PJL^{+}12, RAM^{+}18, RAM^{+}19, RMS^{+}13, SHJ^{+}12, SOGR^{+}19, SJ11, TDS^{+}19, VLVS^{+}16, VBL^{+}15, VBL^{+}16, WDM^{+}15, WMM17, WXL16, WKW^{+}11, Wen16, WNSAL12, XQS^{+}15, YNL11, ZLX^{+}19]. viral-induced [XQS^{+}15]. viremia [MGN^{+}16, XWP^{+}18]. Virginia [PBBG19]. virginica [CBM^{+}18]. viridans [CG11]. Virion [FE11]. Virucidal [JST^{+}19]. Virulence [BGG11, DSLR16, EBM^{+}18, JDH^{+}18, RPGCIG17, XXL^{+}19, ZZC^{+}17, BHF13, BLF^{+}16, BGC^{+}17, BFC^{+}15, COB^{+}16, CD17, CG11, CHB^{+}17, CDG^{+}14, DSSB18, DSSB19, EA18, FMB^{+}11, GTGCM19, HZJ^{+}19, HGX^{+}19, JMS13, JJSJ15, LCV^{+}17, LFD^{+}17, LLLG18, LZZN10, LXZ^{+}19, NDA^{+}19, RCD^{+}17, SL15, STES14, TDH^{+}10, VSS17, VDHV^{+}16, YZW^{+}19, ZJDL18, ZKL^{+}19, ZWL^{+}16, BAMYAH19]. Virulence-associated [ZZC^{+}17]. virulence-related [COB^{+}16]. virulent [DZC^{+}16, FNISE12, GPL^{+}12, MAC^{+}15, OFA11, RPM^{+}19, WZA10]. virus [ADMF15, ATM^{+}10, ARR^{+}18, ABVS18, ASHZ^{+}12, AOAH13, AHHR^{+}10, AHPR^{+}16, AAUHW15, Ano12c, APG^{+}11, BLdA^{+}18, BWNS^{+}11, BAH^{+}18, CVGR^{+}13, CCCTA^{+}14, CNN^{+}10, CWW^{+}18, CGW^{+}19, CLL^{+}10, CLW^{+}13, CGG^{+}10, CMS^{+}19, CCME16, DLCC^{+}13, DZV10, DCZ^{+}17, DPADO16, DCMM13, DBM^{+}14, EYF^{+}19, ESA^{+}10, ECP^{+}10, EKT^{+}11, EBM^{+}18, EKED^{+}17, FSO^{+}10, FE11, GJVT^{+}11, GL18, GFB^{+}13, GSBEM15, GGW^{+}11, GBSF12, GFS^{+}14, GVC^{+}19, GHW^{+}15, GCS^{+}18, HSC^{+}16, HMV^{+}17, HVVV15, HGG^{+}11, HZP^{+}12, HHX^{+}15, HPAL10, IBFL19, IOO15, JST^{+}19, JKC^{+}11, JQK14, JHW17, JFWH19, JNS^{+}19, JYZ^{+}14, JGI^{+}10, JSK^{+}10, JKS^{+}13, JJS^{+}12, JMS13, JJSJ15, KKK15, KF10a, KF10b, KKK^{+}18, KDA^{+}18, KBD13, KWD^{+}16, LC17, LASRA^{+}19, LYL^{+}18, LRD^{+}19, LQY11, LLWL15, LHW^{+}15a, LGJ19, LJGRI^{+}10, LVADB17, MLH^{+}15, MBOC19, MRR^{+}14, MKI^{+}15, MMWK19, MPBM^{+}16, MRN^{+}19, MMBS10, MBS^{+}17, MSB^{+}19, MTK^{+}10]. virus [MAC^{+}15, MFES18, MSI^{+}15, MF12, MGN^{+}16, MTHB13, MCW^{+}18, MNM^{+}18, MNM^{+}19, MMH10, MCO11, MMW^{+}15, MMNL14, NMV^{+}19, NNCLGC^{+}11, NSC^{+}18, NB10, NOST^{+}16, OTN^{+}12, OKS^{+}13, OMO^{+}15, OBFR12, OEG^{+}13, OCLI^{+}19, PTM^{+}19, PHvdH12, PWHW18b, PMW^{+}10, PMH^{+}11, PBM^{+}15, PSK^{+}13, PJL^{+}12, PYSS19, RNM^{+}12, RPGCIG17, RDN^{+}12, RMSH09, SJDC^{+}13, SPTDEE^{+}15, SSV^{+}10, SGB^{+}16, SHJ^{+}12, STG^{+}10, SP15, SKN^{+}16, SWD^{+}14, SZZSFN^{+}14, SVM^{+}18, SG12, SEM18, SJ11, SSO^{+}14, SRD^{+}16, TMM^{+}15, TIK^{+}14a, TSS18, TB12, TSK^{+}19, TRS19, TTL^{+}16, TCM^{+}19, TDT^{+}16, VLVS^{+}16, VCOD17, VBL^{+}15, VBL^{+}16, WWS^{+}11, WCW16a, WWZ^{+}18, WKW^{+}11, WCW^{+}16b, WNSAL12, WNAF^{+}13, XJL^{+}19, XWP^{+}18, YTKS19, YLSW14, YHY^{+}16, YLL^{+}18, YCNW18, YWW^{+}19, ZLX^{+}19, ZLN^{+}17, CCCTA^{+}14, LLLG18, NFF^{+}17, WNGS^{+}17, RWTB17, SRM^{+}18, SNK^{+}19]. virus-3 [SEM18]. virus-like [SWD^{+}14]. viruses [BSC^{+}14, CRAG^{+}12, DPADO16, KKT^{+}18b, NSC^{+}15, PPE^{+}18]. Visceral [ZZA^{+}14, ZWG^{+}17]. viscosa [BGG11, ESBE18]. visual [WSX^{+}18]. visualization [DXX^{+}15, FMM^{+}19, KTK17]. vitellogenic [CZB^{+}11]. vitreus [GRP^{+}15, VBL^{+}16, VKK^{+}19]. vitro [AMRHM18, ASS19, DSSB19, NMV^{+}19, OMB^{+}19, SSG^{+}19, SOGR^{+}19, TFG19, TKS^{+}18, VDO^{+}18a, VMBGIC19, YLX^{+}19b, AS17, APG^{+}11, CRBLC^{+}17, CHAB10, DSSB18, KSAB^{+}14, MLWC11, MBT^{+}17, Mor12, NSE^{+}17, PZL^{+}14, RBAT^{+}17, SHS14, STES14, SBW11, TDH^{+}10, VMCM14, XZZ16]. vivipara [BCM^{+}17]. viviparous [AHB^{+}17]. vivo [AMRHM18, DLLB^{+}18, DSSB19, GBO^{+}18, LLLG18, LXZ^{+}19, SOGR^{+}19, YLX^{+}19b, BKD17, BLFRCQ10, DSSB18, KAESEM13, LVADB17, MLEM16, PDLW17, PZL^{+}14, RBAT^{+}17]. VNTR [HZP^{+}12]. VNTRs [DZV10]. vollucellus [LRB^{+}17]. vortens [MLWC11]. VP2 [CGW^{+}19]. VP26 [SRD^{+}16]. VP28 [MTK^{+}10, SRD^{+}16, TMM^{+}15]. VP38 [ZWG^{+}18]. VP51B [LC17]. VREFT [GGW^{+}11]. vs [ZCCN13]. vs. [VPS^{+}14]. vulnificus [FLVA10, QZZ^{+}18b, YGL^{+}18]. vulvolacinata [SDZN19].

W [JCN^{+}14]. Wadden [GVLE16]. wageneri [GAC10]. wageneri-group [GAC10]. Waite [KRD^{+}17, LLH^{+}14b]. Walbaum [ASA10, ASHZ^{+}12, AOAH13, AHHR^{+}10, AJLC^{+}16, ABT^{+}14, AHTCA^{+}17, AA10, AMA11, BSW17, BCLM^{+}19, BHA^{+}11, BA14, BPPKJ10, BLP10, BLFC12, CSS^{+}16, CHAB10, CMH^{+}17, CPB^{+}15, CMR^{+}16, CLS^{+}15, DS10, DPTFC10, DPTCF10, DMP10, DJR^{+}19, DdRTM13, DdRCP13, DT16, GMC^{+}16a, GCNB16, GTH11, GPFGAG19, GDW^{+}15, GMC^{+}16b, HKBD15a, HKBD15b, JKL^{+}13, JDH^{+}18, JML19, vGJHS^{+}11, KSM^{+}17, KPT^{+}11, KBE^{+}14, LPA^{+}10, LLS^{+}11, LST^{+}12, LF16, LF18, LWMT10, MNF^{+}15, NOV^{+}11, NDA10, OP14, OBFR12, OBD^{+}11, OLT^{+}10, PSBG^{+}11, PLCC11, PMPG^{+}16, RSB13, RCGGM15, RHB^{+}15, SMJ^{+}12, STF^{+}15, SHJ^{+}12, SBM^{+}11, SCR^{+}15, SKJ13, XKB12, YDO^{+}16]. Wales [TDJ^{+}10, TWJP10]. Walker [LRB^{+}17]. walking [BNC^{+}11]. wall [BB11, WBF^{+}10]. Walleye [GRP^{+}15, VBL^{+}15, VBL^{+}16, VKK^{+}19, WMF14]. Wanted [CBM^{+}18]. Ward [RGMGG^{+}18]. warmer [GMN12]. was [GMC^{+}16a]. Washington [PPE^{+}18]. waste [DAB^{+}12, PJL^{+}12]. Water [SFR^{+}17, BMK17, CHV^{+}12, CD17, CB13, GDW^{+}15, GHB^{+}11, HH13a, HMW11, JJV^{+}15, KSS17, KPM^{+}11, MSR^{+}10, MUM^{+}19, MoSMS18, MTLP^{+}18b, MSK^{+}12, NOV^{+}11, OM10, PST^{+}16, RHTC19b, Sam10, SVY^{+}14, VCN^{+}10, WNGS^{+}17, YkSKY13]. waterborne [HKBD15b]. waters [BCL^{+}19, CNN^{+}10, EA18, JZK^{+}14, LRAP17, MMBS10, RBS^{+}16, RSF^{+}17, RGMGG^{+}18, RBR19, SBH^{+}16, SN14]. watershed [PHR14]. WE-skin11f [VKK^{+}19]. weakfish [CCM^{+}11]. Webb [RAM^{+}17]. websteri [CBM^{+}18]. wedge [LPVGL^{+}10]. weedy [LYD^{+}15]. weight [FMH^{+}15, TJB^{+}15]. weirs [LST^{+}12]. Weissella [CLS^{+}15, GMC^{+}16b, MHNY18]. weissellosis [MHNY18]. well [KKK^{+}18, PSZ^{+}14]. well-being [PSZ^{+}14]. West [HHM^{+}14, SBH^{+}16, PRWG17]. western [BHT^{+}17, BA13, EEMCEGSP16, PMP^{+}18, RPGCIG17, SDJH16, FSV^{+}18, FSB^{+}18c, RSR^{+}18]. whaler [RBF14]. which [CXW^{+}19, QZZ^{+}18b]. while [JOF^{+}19]. White [CNN^{+}10, CLW^{+}13, DCV^{+}14, BTN^{+}18, BWB^{+}17, BMT^{+}15, DLCC^{+}13, DZV10, DCMM13, DBM^{+}14, ESA^{+}10, EEMCEGSP16, FSO^{+}10, HVVV15, HZP^{+}12, HKF^{+}16, JQK14, JGI^{+}10, JJS^{+}12, LQY11, MTK^{+}10, MGN^{+}16, NRMS14, NSC^{+}18, PSK^{+}13, RNM^{+}12, RPGCIG17, RDN^{+}12, RBF14, SMS^{+}10, STS^{+}15, SPTDEE^{+}15, SSV^{+}10, SP15, SVM^{+}18, SRD^{+}16, TMM^{+}15, TSK^{+}19, TTL^{+}16, TDT^{+}16, XKB12, YHY^{+}16, DCZ^{+}17]. whiteleg [CMEV17a, CMS^{+}17, CMEV17b, KS19, SSV^{+}17, STG^{+}10, TMM^{+}15]. whitemouth [MPM16]. whiting [RGMGG^{+}18]. Whole [RPMRLJ^{+}19, RKK^{+}19, BHF13, KTK17]. Whole-genome [RKK^{+}19]. wickerhamii [JDRM17]. wide [KF10a, LHW^{+}15a]. widespread [PPE^{+}18]. wild [ARR^{+}18, AAUHW15, AGFG19, ADR11, BSC^{+}14, BFFA^{+}12, BGCKF12, CNN^{+}10, EMA17, EA18, FBS13, FBAS16, GFO^{+}13, GGMG18b, GMG^{+}19, GFB^{+}13, HBVB16, HGPD11, HFKF17, JKO^{+}13, JHM^{+}18, JHRW17, JNN^{+}11, KCD^{+}14, KMGK^{+}18, LHW^{+}15a, LDEH17, LFL19, MMWK19, MMB^{+}15, MMBS10, MVWS13, MCW^{+}18, PTM^{+}19, RDN^{+}12, RBS^{+}16, RSF^{+}17, RMSH09, RBWR^{+}15, RSM^{+}18, SBM^{+}11, SDZN19, SWE^{+}19, SN14, TJA^{+}13, VPS^{+}14, VCS^{+}19, WDM^{+}15, WMM17, ZLX^{+}19, dSRSHGá^{+}19]. wild-caught [HFKF17, RBWR^{+}15, RSM^{+}18]. wild-hatched [SBM^{+}11]. William [EW17]. wingei [ADR11]. Winter [MP18, KHG^{+}17]. within [BMBS12, DZV10, LLC^{+}11, YEK^{+}17]. Wootten [Rob14a]. world [GVLE16]. worldwide [DBV^{+}13, MNG^{+}18]. worm [BGRB14]. wound [ASSP^{+}15, SSG^{+}19]. wrasse [BVK^{+}16, BT14, MMW^{+}15, RSP^{+}17, RSM^{+}18, SRM^{+}18, Tre12, WDM^{+}15]. WSSV [DCZ^{+}17, DCMM13, HVVV15, JPG^{+}19, LC17, PSK^{+}13, RDN^{+}12, SSV^{+}10, WLX^{+}10, WXWD15, YHY^{+}16]. wulii [ZYW^{+}10].

X [EKT^{+}11, MFES18]. Xanthine [BSD^{+}18d, BSD^{+}18c, BSD^{+}18a]. xenobiotica [CBH^{+}15]. xenomas [LRAP17]. XI [BNS13].

YaeT [KREM14]. yasunagai [SS17]. Yasutake [EW17]. year [KZJ^{+}17, SBW^{+}18, WGL17]. yearling [LXL^{+}18, SFR^{+}17]. years [BMMC^{+}15]. yeast [SCK13, WBF^{+}10]. yellow [CWW^{+}18, GM11, HZJ^{+}19, KHLL16, LYD^{+}15, LRD^{+}19, NNCLGC^{+}11, OEG^{+}13, PHvdH12, SGQ^{+}15, STG^{+}10, XQS^{+}15, ZZA^{+}14, ZWG^{+}17]. yellowfin [Dig13, Dig15, Pol15]. Yellowstone [MKKM15]. yellowtail [KHS11, MFS^{+}17, NWMK16, RBAT^{+}17, STIG17, YYF^{+}10]. Yersinia [ASA10, AHTCA^{+}17, CMH^{+}17, DS10, DSSB18, DSSB19, GTGCM19, GCNB18, GTH11, JKL^{+}13, KSM^{+}17, KJW^{+}12, NCNB18, SON^{+}18, TALA11, TDH^{+}10, WNCR12]. Yersiniosis [GGGM14]. YHV [CWW^{+}18, LRD^{+}19]. YHV-3 [CWW^{+}18]. YHV-8 [LRD^{+}19]. Yin [GAC10, MWS17]. yl [SRD^{+}16]. York [GFB^{+}19]. young [SBW^{+}18].

Zambia [HHC^{+}18]. Zealand [BMH^{+}19, DWP^{+}19, GBO^{+}18, RPJ^{+}18]. zebra [Bin14]. Zebrafish [CKvGJB17, SSG^{+}19, AMRHM18, BLdA^{+}18, BWS^{+}18, CDW17, CBW19, CXK16, CSSWK10, KWGS19, KDP^{+}17, LRC^{+}18, LHW^{+}12, LHW^{+}15b, MLR^{+}12, MWRSK17, MGF17, RAM^{+}19, SZM^{+}15, SSK17, ZJR^{+}18, EDP^{+}17]. Zeuxapta [STIG17]. Zeylanicobdella [KCL14]. Zhejiang [LQY11]. Zinc [DGS^{+}14]. ZnuABC [DS10]. Zoarces [AHB^{+}17]. zonetail [BCJ13]. zonura [BCJ13]. Zoogoneticus [RAM^{+}17]. zoonotic [FŠK^{+}19]. Zschokkella [YMBE13].

References

Awad:2010:ULL

[AA10] E. Awad and B. Austin. Use of lupin, Lupinus perennis, mango, Mangifera indica, and stinging nettle, Urtica dioica, as feed additives to prevent Aeromonas hydrophila infection in rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 33(5):413–420, May 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Alegria:2015:CSS

[AAUHW15] R. Alegría, G. Arriagada, S. Urcelay, and C. Hamilton-West. Cross-sectional study to investigate the presence of salmon pancreas disease virus in wild and feral fish populations in 10 lakes, Los Lagos Region, Chile. Journal of Fish Diseases, 38(1):113–119, January 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Amoroso:2016:SAA

[AAV^{+}16] G. Amoroso, M. B. Adams, T. Ventura, C. G. Carter, and J. M. Cobcroft. Skeletal anomaly assessment in diploid and triploid juvenile Atlantic salmon (Salmo salar L.) and the effect of temperature in freshwater. Journal of Fish Diseases, 39(4):449–466, April 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Aisyhah:2015:SAI

[AAZS^{+}15] M. A. S. Aisyhah, M. N. A. Amal, M. Zamri-Saad, A. Siti-Zahrah, and N. N. Shaqinah. Streptococcus agalactiae isolates from cultured fishes in Malaysia manifesting low resistance pattern towards selected antibiotics. Journal of Fish Diseases, 38(12):1093–1098, December 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Andrews:2010:HRC

[ABC^{+}10] M. Andrews, S. Battaglene, J. Cobcroft, M. Adams, E. Noga, and B. Nowak. Host response to the chondracanthid copepod Chondracanthus goldsmidi, a gill parasite of the striped trumpeter, Latris lineata (Forster), in Tasmania. Journal of Fish Diseases, 33(3):211–220, March 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ambas:2015:ISP

[ABF15] I. Ambas, N. Buller, and R. Fotedar. Isolation and screening of probiotic candidates from marron, Cherax cainii (Austin, 2002) gastrointestinal tract (GIT) and commercial probiotic products for the use in marron culture. Journal of Fish Diseases, 38(5):467–476, May 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Adams:2015:MDR

[ABM^{+}15] D. H. Adams, J. D. Borucinska, K. Maillett, K. Whitburn, and T. E. Sander. Mortality due to a retained circle hook in a longfin mako shark Isurus paucus (Guitart–Manday). Journal of Fish Diseases, 38(7):621–628, July 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Avci:2014:ELG

[ABT^{+}14] H. Avci, S. S. Birincioglu, T. T. Tanrikul, E. T. Epikmen, N. Metin, and M. L. Avsever. Experimental Lactococcus garvieae infection in rainbow trout, Oncorhynchus mykiss, Walbaum 1792: a comparative histopathological and immunohistochemical study. Journal of Fish Diseases, 37(5): 481–495, May 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Adamek:2018:CEV

[ABVS18] M. Adamek, F. Baska, B. Vincze, and D. Steinhagen. Carp edema virus from three genogroups is present in common carp in Hungary. Journal of Fish Diseases, 41(3):463–468, March 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Amoroso:2016:CLJ

[ACA^{+}16] G. Amoroso, J. M. Cobcroft, M. B. Adams, T. Ventura, and C. G. Carter. Concurrence of lower jaw skeletal anomalies in triploid Atlantic salmon (Salmo salar L.) and the effect on growth in freshwater. Journal of Fish Diseases, 39(12): 1509–1521, December 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Abayneh:2014:MVN

[ACAS14] T. Abayneh, D. J. Colquhoun, D. Austin, and H. Sørum. Multilocus variable number tandem repeat analysis of Edwardsiella piscicida isolates pathogenic to fish. Journal of Fish Diseases, 37(11):941–948, November 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Armwood:2019:PCC

[ACLP^{+}19] Abigail R. Armwood, Alvin C. Camus, Adrián López-Porras, Cyndi Ware, Matt J. Griffin, and Esteban Soto. Pathologic changes in cultured Nile tilapia (Oreochromis niloticus) associated with an outbreak of Edwardsiella anguillarum. Journal of Fish Diseases, 42(10):1463–1469, October 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Aro:2014:CMS

[ACM^{+}14] L. Aro, K. Correa, A. Martínez, R. Ildefonso, and J. M. Yáñez. Characterization of Mycobacterium salmoniphilum as causal agent of mycobacteriosis in Atlantic salmon, Salmo salar L., from a freshwater recirculation system. Journal of Fish Diseases, 37(4):341–348, April 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Adams:2012:PSU

[ACN12] M. B. Adams, P. B. B. Crosbie, and B. F. Nowak. Preliminary success using hydrogen peroxide to treat Atlantic salmon, Salmo salar L., affected with experimentally induced amoebic gill disease (AGD). Journal of Fish Diseases, 35(11):839–848, November 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ahmed:2019:DIM

[ACSB^{+}19] Nahla Hossameldin Ahmed, Monica Caffara, Ariadna Sitjà-Bobadilla, Maria Letizia Fioravanti, Angelica Mazzone, Abbass Sayed Aboulezz, Asmaa Mohamed Metwally, Mosaab Adl-Eldin Omar, and Oswaldo R. Palenzuela. Detection of the intranuclear microsporidian Enterospora nucleophila in gilthead sea bream by in situ hybridization. Journal of Fish Diseases, 42(6):809–815, June 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Adhya:2010:ELM

[ACYC10] M. Adhya, K-S. Choi, Y. Yu, and M. Cho. Expression and localization of MCsialec, a sialic acid-specific lectin in the marine bivalve Manila clam, Ruditapes philppinarum. Journal of Fish Diseases, 33(11):889–899, November 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Aamelfot:2014:ISA

[ADF14] M. Aamelfot, O. B. Dale, and K. Falk. Infectious salmon anaemia — pathogenesis and tropism. Journal of Fish Diseases, 37(4):291–307, April 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Aamelfot:2015:HTI

[ADMF15] M. Aamelfot, O. B. Dale, A. McBeath, and K. Falk. Host tropism of infectious salmon anaemia virus in marine and freshwater fish species. Journal of Fish Diseases, 38(8):687– 694, August 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Assis:2017:SSR

[AdOG^{+}17] G. B. N. Assis, T. F. de Oliveira, I. A. Gardner, H. C. P. Figueiredo, and C. A. G. Leal. Sensitivity and specificity of real-time PCR and bacteriological culture for francisellosis in farm-raised Nile tilapia (Oreochromis niloticus L.). Journal of Fish Diseases, 40(6):785–795, June 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Arbuatti:2011:HOC

[ADR11] A. Arbuatti, L. Della Salda, and M. Romanucci. Histological observations on conjoined twins occurring among a captive-bred wild line of Poecilia wingei (Poeser, Kempkes & Isbrücker). Journal of Fish Diseases, 34(4):319–321, April 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anderson:2012:SMA

[AFA^{+}12] E. T. Anderson, S. Frasca, Jr., M. G. Asakawa, M. H. Fatzinger, J. Johnson, K. Marchetere, L. Goodale, G. R. Risatti, and C. A. Harms. Splenic mycobacteriosis in an Atlantic guitarfish, Rhinobatos lentiginosus Garman. Journal of Fish Diseases, 35(7):541–544, July 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Amesberger-Freitag:2019:HHD

[AFTEML19] Alessandra Amesberger-Freitag, Alexander Tichy, Mansour El-Matbouli, and Eva Lewisch. Hole-in-the-head disease in discus fish, Symphysodon (Heckel, 1840): Is it a consequence of a dietary Ca/P imbalance? Journal of Fish Diseases, 42 (8):1133–1142, August 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Arseneau:2019:AIR

[AGB^{+}19] Jean-René Arseneau, Chantal Gautreau, Linda Boston, Michel L. Goguen, and Mark Laflamme. Accelerated ISAV replication detection by cell culture methods combined with time-monitoring RT-qPCR. Journal of Fish Diseases, 42(2): 257–267, February 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Aller-Gancedo:2019:DPS

[AGFG19] José-Miguel Aller-Gancedo and Juan-Miguel Fregeneda-Grandes. Distribution patterns of saprolegniosis cutaneous lesions in wild and farmed brown trout (Salmo trutta L.) obtained using a geographic information system (GIS). Journal of Fish Diseases, 42(10):1419–1424, October 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Alarcon:2016:PLC

[AGR^{+}16] M. Alarcón, S. Gulla, M. V. Røsæg, A. Rønneseth, H. Wergeland, T. T. Poppe, H. Nilsen, and D. J. Colquhoun. Pasteurellosis in lumpsucker Cyclopterus lumpus, farmed in Norway. Journal of Fish Diseases, 39(4):489–495, April 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Aaen:2016:SMCb

[AH16] S. M. Aaen and T. E. Horsberg. A screening of multiple classes of pharmaceutical compounds for effect on preadult salmon lice Lepeophtheirus salmonis. Journal of Fish Diseases, 39(10):1213–1223, October 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Avendano-Herrera:2019:EFI

[AHAMR^{+}19] Ruben Avendaño-Herrera, Eloisa Arias-Muñoz, Verónica Rojas, Alicia E. Toranzo, Matías Poblete-Morales, Claudio Córdova, and Rute Irgang. Evidence for the facultative intracellular behaviour of the fish pathogen Vibrio ordalii. Journal of Fish Diseases, 42(10):1447–1455, October 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Axen:2017:ENR

[AHB^{+}17] C. Axén, M. Hakhverdyan, T. S. Boutrup, E. Blomkvist, F. Ljunghager, A. Alfjorden, Å. Hagström, N. J. Olesen, M. Juremalm, M. Leijon, and J-F. Valarcher. Emergence of a new rhabdovirus associated with mass mortalities in eelpout (Zoarces viviparous) in the Baltic Sea. Journal of Fish Diseases, 40(2):219–229, February 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Aaen:2016:SMCa

[AHH16] S. M. Aaen, L. A. Hamre, and T. E. Horsberg. A screening of medicinal compounds for their effect on egg strings and nauplii of the salmon louse Lepeophtheirus salmonis (Krøyer). Journal of Fish Diseases, 39(10):1201–1212, October 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Al-Hussinee:2010:VHS

[AHHR^{+}10] L. Al-Hussinee, P. Huber, S. Russell, V. LePage, A. Reid, K. M. Young, E. Nagy, R. M. W. Stevenson, and J. S. Lumsden. Viral haemorrhagic septicaemia virus IVb experimental infection of rainbow trout, Oncorhynchus mykiss (Walbaum), and fathead minnow, Pimphales promelas (Rafinesque). Journal of Fish Diseases, 33(4):347–360, April 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Avendano-Herrera:2018:PPD

[AHITC18] R. Avendaño-Herrera, R. Irgang, and D. Tapia-Cammas. PCR procedure for detecting the fish pathogen Tenacibaculum dicentrarchi. Journal of Fish Diseases, 41(4):715–719, April 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Adamek:2018:CTT

[AHJS^{+}18] M. Adamek, J. Hellmann, V. Jung-Schroers, F. Teitge, and D. Steinhagen. CyHV-2 transmission in traded goldfish stocks in Germany — a case study. Journal of Fish Diseases, 41(2):401–404, February 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Al-Hussinee:2016:TPR

[AHPR^{+}16] L. Al-Hussinee, Ph Pham, S. Russell, L. Tubbs, C. Tafalla, Nc Bols, B. Dixon, and Js Lumsden. Temporary protection of rainbow trout gill epithelial cells from infection with viral haemorrhagic septicaemia virus IVb. Journal of Fish Diseases, 39(9):1099–1112, September 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Avendano-Herrera:2017:DCY

[AHTCA^{+}17] R. Avendaño-Herrera, D. Tapia-Cammas, A. Aedo, P. Saldivia, C. Ortega, and R. Irgang. Disease caused by Yersinia ruckeri serotype O2b found in Chilean-farmed Coho salmon, Oncorhynchus kisutch (Walbaum, 1792). Journal of Fish Diseases, 40(2):279–285, February 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Aoki:2011:GRB

[AI11] T. Aoki and T. Inoue. Glycophorin in red blood cell membranes of healthy and diseased carp, Cyprinus carpio L. Journal of Fish Diseases, 34(7):573–576, July 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Al-Jubury:2016:EII

[AJLC^{+}16] A. Al-Jubury, S. LaPatra, N. D. Christensen, S. Zuo, C. Tafalla, and K. Buchmann. Exclusion of IgD-, IgT- and IgM-positive immune cells in Ichthyophonus-induced granulomas in rainbow trout Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 39(11):1399–1402, November 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Al-Jubury:2018:IPH

[AJLK^{+}18] A. Al-Jubury, C. Lu, P. W. Kania, L. von Gersdorff Jørgensen, Y. Liu, I. de Bruijn, J. Raaijmakers, and K. Buchmann. Impact of Pseudomonas H6 surfactant on all external life cycle stages of the fish parasitic ciliate Ichthyophthirius multifiliis. Journal of Fish Diseases, 41(7): 1147–1152, July 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kristmundsson:2016:LRS

[áKÁGA16] á Kristmundsson, F. Árnason, S. Gudmundsdóttir, and T. Antonsson. Levels of Renibacterium salmoninarum antigens in resident and anadromous salmonids in the River Ellidaár system in Iceland. Journal of Fish Diseases, 39(6):681– 692, June 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Adamovicz:2015:PAC

[ALG15] L. Adamovicz, G. Lewbart, and B. Gilger. Phacoemulsification and aspiration for cataract management in a dollar sunfish, Lepomis marginatus (Holbrook) — a case report. Journal of Fish Diseases, 38(12):1089–1092, December 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Awad:2011:EDS

[AMA11] E. Awad, W. J. Mitchell, and B. Austin. Effect of dietary supplements on cytokine gene expression in rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 34(8):629–634, August 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Abdallah:2018:TJI

[AMAR18] Ebtsam Sayed Hassan Abdallah, Mahmoud Mostafa Mahmoud, and Ismail Ramadan Abdel-Rahim. Trichosporon jirovecii infection of red swamp crayfish (Procambarus clarkii). Journal of Fish Diseases, 41(11):1719–1732, November 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Alavinia:2018:VVE

[AMRHM18] Seyed Jalil Alavinia, Seyed Saeed Mirzargar, Hooman Rahmati-Holasoo, and Hosseinali Ebrahimzadeh Mousavi. The in vitro and in vivo effect of tannic acid on Ichthyophthirius multifiliis in zebrafish (Danio rerio) to treat ichthyophthiriasis. Journal of Fish Diseases, 41(12):1793–1802, December 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Alves:2019:ALI

[ANB^{+}19] Carliane Maria Guimarães Alves, Joziele Neves Nogueira, Isaac Belo Barriga, Joilson Rodrigues dos Santos, Gracienhe Gomes Santos, and Marcos Tavares-Dias. Albendazole, levamisole and ivermectin are effective against monogeneans of Colossoma macropomum (Pisces: Serrasalmidae). Journal of Fish Diseases, 42(3):405–412, March 2019. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2010:Ea

[Ano10a] Anonymous. Erratum. Journal of Fish Diseases, 33(5):459, May 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2010:Eb

[Ano10b] Anonymous. Erratum. Journal of Fish Diseases, 33(7):623, July 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2010:SAS

[Ano10c] Anonymous. Salmonid alphavirus (SAV) and pancreas disease (PD) in Atlantic salmon, Salmo salar L., in freshwater and seawater sites in Norway from 2006 to 2008. Journal of Fish Diseases, 33(8):705, August 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2011:E

[Ano11] Anonymous. Erratum. Journal of Fish Diseases, 34(4):331, April 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2012:Ca

[Ano12a] Anonymous. Corrigendum. Journal of Fish Diseases, 35(8): 619, August 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2012:Cb

[Ano12b] Anonymous. Corrigendum. Journal of Fish Diseases, 35(9): 709, September 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2012:DNA

[Ano12c] Anonymous. Development of neutralizing antibody responses in muskellunge, Esox masquinongy (Mitchill), experimentally exposed to viral haemorrhagic septicaemia virus (genotype IVb). Journal of Fish Diseases, 35(5):393, May 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2014:E

[Ano14] Anonymous. Erratum. Journal of Fish Diseases, 37(1):75, January 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2017:Ca

[Ano17a] Anonymous. Corrigendum. Journal of Fish Diseases, 40 (10):1487, October 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2017:Cb

[Ano17b] Anonymous. Corrigendum. Journal of Fish Diseases, 40(12): 1941, December 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2018:IIa

[Ano18a] Anonymous. Issue information. Journal of Fish Diseases, 41 (1):ii–iv, January 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2018:IIb

[Ano18b] Anonymous. Issue information. Journal of Fish Diseases, 41 (2):ii–iv, February 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2018:IIc

[Ano18c] Anonymous. Issue information. Journal of Fish Diseases, 41(3):i–iv, March 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2018:IId

[Ano18d] Anonymous. Issue information. Journal of Fish Diseases, 41(4):i–iv, April 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2018:IIe

[Ano18e] Anonymous. Issue information. Journal of Fish Diseases, 41(5):i–iv, May 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2018:IIf

[Ano18f] Anonymous. Issue information. Journal of Fish Diseases, 41(6):ii–v, June 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2018:IIg

[Ano18g] Anonymous. Issue information. Journal of Fish Diseases, 41(7):ii–v, July 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2018:IIh

[Ano18h] Anonymous. Issue information. Journal of Fish Diseases, 41(8):ii–v, August 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2018:IIi

[Ano18i] Anonymous. Issue information. Journal of Fish Diseases, 41 (9):ii–v, September 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2018:IIj

[Ano18j] Anonymous. Issue information. Journal of Fish Diseases, 41 (10):ii–v, October 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2018:IIk

[Ano18k] Anonymous. Issue information. Journal of Fish Diseases, 41 (11):i–iv, November 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2018:IIl

[Ano18l] Anonymous. Issue information. Journal of Fish Diseases, 41 (12):ii–v, December 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2019:IIa

[Ano19a] Anonymous. Issue information. Journal of Fish Diseases, 42 (1):iii–vi, January 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2019:IIb

[Ano19b] Anonymous. Issue information. Journal of Fish Diseases, 42 (2):i–iv, February 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2019:IIc

[Ano19c] Anonymous. Issue information. Journal of Fish Diseases, 42(3):i–iv, March 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2019:IId

[Ano19d] Anonymous. Issue information. Journal of Fish Diseases, 42(4):ii–v, April 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2019:IIe

[Ano19e] Anonymous. Issue information. Journal of Fish Diseases, 42(5):ii–v, May 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2019:IIf

[Ano19f] Anonymous. Issue information. Journal of Fish Diseases, 42(6):i–iv, June 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2019:IIg

[Ano19g] Anonymous. Issue information. Journal of Fish Diseases, 42(7):i–iv, July 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2019:IIh

[Ano19h] Anonymous. Issue information. Journal of Fish Diseases, 42 (8):iii–vi, August 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2019:IIi

[Ano19i] Anonymous. Issue information. Journal of Fish Diseases, 42 (9):i–iv, September 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2019:IIj

[Ano19j] Anonymous. Issue information. Journal of Fish Diseases, 42 (10):i–iv, October 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2019:IIk

[Ano19k] Anonymous. Issue information. Journal of Fish Diseases, 42 (11):ii–v, November 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2019:IIl

[Ano19l] Anonymous. Issue information. Journal of Fish Diseases, 42 (12):ii–v, December 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ahmadi:2013:TDI

[AOAH13] N. Ahmadi, A. Oryan, M. Akhlaghi, and A. Hosseini. Tissue distribution of infectious pancreatic necrosis virus serotype Sp in naturally infected cultured rainbow trout, Oncorhynchus mykiss (Walbaum): an immunohistochemical and nested-PCR study. Journal of Fish Diseases, 36(7):629– 637, July 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Arunrut:2011:VCS

[APG^{+}11] N. Arunrut, J. Phromjai, W. Gangnonngiw, N. Kanthong, S. Sriurairatana, and W. Kiatpathomchai. In vitro cultivation of shrimp Taura syndrome virus (TSV) in a C6/ 36 mosquito cell line. Journal of Fish Diseases, 34(10):805–810, October 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Antuofermo:2017:MCM

[APP^{+}17] E. Antuofermo, A. Pais, M. Polinas, T. Cubeddu, M. Righetti, M. A. Sanna, and M. Prearo. Mycobacteriosis caused by Mycobacterium marinum in reared mullets: first evidence from Sardinia (Italy). Journal of Fish Diseases, 40(3):327–337, March 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Andree:2013:CIP

[ARFG13] K. B. Andree, C. J. Rodgers, D. Furones, and E. Gisbert. Co-infection with Pseudomonas anguilliseptica and Delftia acidovorans in the European eel, Anguilla anguilla (L.): a case history of an illegally trafficked protected species. Journal of Fish Diseases, 36(7):647–656, July 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Aguilera:2017:BDP

[ARL^{+}17] O. Aguilera, I. Rocha, M. S. Lopes, I. Lima, R. T. Lopes, A. S. Machado, R. B. Guimarães, M. A. C. Crapez, M. C. Tenório, and A. Nepomuceno. The bone degenerative processes in senile fishes from Holocene Brazilian shell mounds. Journal of Fish Diseases, 40(12):1869–1881, December 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Abdullah:2018:FDT

[ARR^{+}18] Azila Abdullah, Rimatulhana Ramly, Mohammad Syafiq Mohammad Ridzwan, Fahmi Sudirwan, Adnan Abas, Kamisa Ahmad, Munira Murni, and Beng Chu Kua. First detection of tilapia lake virus (TiLV) in wild river carp (Barbonymus schwanenfeldii) at Timah Tasoh Lake, Malaysia. Journal of Fish Diseases, 41(9):1459–1462, September 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ali:2017:FBA

[AS17] S. E. Ali and I. Skaar. A fluorescence-based assay for in vitro screening of Saprolegnia inhibitors. Journal of Fish Diseases, 40(10):1333–1339, October 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Abbass:2010:CCP

[ASA10] A. Abbass, S. M. Sharifuzzaman, and B. Austin. Cellular components of probiotics control Yersinia ruckeri infection in rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 33(1):31–37, January 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Avarre:2012:STA

[ASB^{+}12] J-C. Avarre, A. Santika, A. Bentenni, Z. Zainun, J-P. Madeira, M. Maskur, L. Bigarré, and D. Caruso. Spatiotemporal analysis of cyprinid herpesvirus 3 genetic diversity at a local scale. Journal of Fish Diseases, 35(10):767–774, October 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Adomako:2012:ODV

[ASHZ^{+}12] M. Adomako, S. St-Hilaire, Y. Zheng, J. Eley, R. D. Marcum, W. Sealey, B. C. Donahower, S. LaPatra, and P. P. Sheridan. Oral DNA vaccination of rainbow trout, Oncorhynchus mykiss (Walbaum), against infectious haematopoietic necrosis virus using PLGA [Poly(D,L-Lactic-Co-Glycolic Acid)] nanoparticles. Journal of Fish Diseases, 35(3):203–214, March 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ali:2019:CAV

[ASS19] Shimaa E. Ali, Mwansa M. Songe, and Ida Skaar. Colorimetric assay for the in vitro evaluation of Saprolegnia biofilm inhibitors. Journal of Fish Diseases, 42(8):1119–1124, August 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Andrews:2015:RTD

[ASSP^{+}15] M. Andrews, M. Stormoen, H. Schmidt-Posthaus, T. Wahli, and P. J. Midtlyng. Rapid temperature-dependent wound closure following adipose fin clipping of Atlantic salmon Salmo salar L. Journal of Fish Diseases, 38(6):523–531, June 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ahnelt:2016:OAL

[ASW16] H. Ahnelt, F. M. Schade, and M. Wegner. Ocean acidification leads to deformations of caudal vein angio-architecture in juvenile threespine stickleback, Gasterosteus aculeatus Linnaeus. Journal of Fish Diseases, 39(8):1001–1005, August 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Adamek:2018:FSP

[ATJS^{+}18] Mikolaj Adamek, Felix Teitge, Verena Jung-Schroers, Max Heling, David Gela, Veronika Piackova, Martin Kocour, and Dieter Steinhagen. Flavobacteria as secondary pathogens in carp suffering from koi sleepy disease. Journal of Fish Diseases, 41(11):1631–1642, November 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Abayneh:2010:EHR

[ATM^{+}10] T. Abayneh, N. Toft, A. B. Mikalsen, E. Brun, and M. Sandberg. Evaluation of histopathology, real-time PCR and virus isolation for diagnosis of infectious salmon anaemia in Norwegian salmon using latent class analysis. Journal of Fish Diseases, 33(6):529–532, June 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Alarcon:2016:CIN

[ATP^{+}16] M. Alarcón, E. Thoen, T. T. Poppe, G. Bornø, S. N. Mohammad, and H. Hansen. Co-infection of Nucleospora cyclopteri (Microsporidia) and Kudoa islandica (Myxozoa) in farmed lumpfish, Cyclopterus lumpus L., in Norway: a case report. Journal of Fish Diseases, 39(4):411–418, April 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Assis:2017:NCS

[ATP^{+}17] G. B. N. Assis, G. C. Tavares, F. L. Pereira, H. C. P. Figueiredo, and C. A. G. Leal. Natural coinfection by Streptococcus agalactiae and Francisella noatunensis subsp . orientalis in farmed Nile tilapia (Oreochromis niloticus L.). Journal of Fish Diseases, 40(1):51–63, January 2017. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Amal:2012:OSA

[AZSI^{+}12] M. N. A. Amal, M. Zamri-Saad, A. R. Iftikhar, A. Siti-Zahrah, S. Aziel, and S. Fahmi. An outbreak of Streptococcus agalactiae infection in cage-cultured golden pompano, Trachinotus blochii (Lacépède), in Malaysia. Journal of Fish Diseases, 35(11):849–852, November 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Amal:2013:TSA

[AZSSZZ13] M. N. A. Amal, M. Zamri-Saad, A. Siti-Zahrah, and A. R. Zulkafli. Transmission of Streptococcus agalactiae from a hatchery into a newly established red hybrid tilapia, Oreochromis niloticus (L.) × Oreochromis mossambicus (Peters), farm. Journal of Fish Diseases, 36(8):735–739, August 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Borucinska:2013:PMO

[BA13] J. D. Borucinska and D. H. Adams. Parasitic metritis and oophoritis associated with nematode larvae in an Atlantic sharpnose shark, Rhizoprionodon terraenovae (Richardson), collected in the western Atlantic off Florida. Journal of Fish Diseases, 36(9):811–817, September 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Boran:2014:ICT

[BA14] H. Boran and I. Altinok. Impacts of chloramine-T treatment on antioxidant enzyme activities and genotoxicity in rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 37(5):431–441, May 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bartie:2012:IDE

[BAD^{+}12] K. L. Bartie, F. W. Austin, A. Diab, C. Dickson, T. T. Dung, M. Giacomini, and M. Crumlish. Intraspecific diversity of Edwardsiella ictaluri isolates from diseased freshwater catfish, Pangasianodon hypophthalmus (Sauvage), cultured in the Mekong Delta, Vietnam. Journal of Fish Diseases, 35 (9):671–682, September 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

BenKhemis:2016:AEC

[BAH^{+}16] I. Ben Khemis, N. Besbes Aridh, N. Hamza, M. M’Hetli, and S. Sadok. Antifungal efficacy of the cactaceae Opuntia stricta (Haworth) prickly pear ethanolic extract in controlling pikeperch Sander lucioperca (Linnaeus) egg saprolegniasis. Journal of Fish Diseases, 39(3):377–383, March 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Buyukekiz:2018:IPN

[BAH^{+}18] A. G. Büyükekiz, S. Altun, E. F. Hansen, İ. B. Satıcıoğlu, M. Duman, T. Markussen, and E. Rimstad. Infectious pancreatic necrosis virus (IPNV) serotype Sp is prevalent in Turkish rainbow trout farms. Journal of Fish Diseases, 41 (1):95–104, January 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bethke:2019:RSI

[BAMYAH19] Jorn Bethke, Eloisa Arias-Muñoz, Alejandro Yáñez, and Ruben Avendaño-Herrera. Renibacterium salmoninarum iron-acquisition mechanisms and ASK cell line infection: Virulence and immune response. Journal of Fish Diseases, 42 (9):1283–1291, September 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Borucinska:2011:LAP

[BB11] J. D. Borucinska and S. A. Bullard. Lesions associated with plerocerci (Platyhelminthes: Cestoda: Trypanorhyncha) in the gastric wall of a cownose ray, Rhinoptera bonasus (Mitchill), (Myliobatiformes: Rhinopteridae) from the northern Gulf of Mexico. Journal of Fish Diseases, 34(2): 149–157, February 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bigarre:2010:NPP

[BBC^{+}10] L. Bigarré, M. Baud, J. Cabon, K. Crenn, and J. Castric. New PCR probes for detection and genotyping of piscine betanodaviruses. Journal of Fish Diseases, 33(11):907–912, November 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Beck:2015:KCP

[BBF^{+}15] B. H. Beck, L. M. Barnett, B. D. Farmer, E. Peatman, and D. Carter. Kaolinitic clay protects against Flavobacterium columnare infection in channel catfish Ictalurus punctatus (Rafinesque). Journal of Fish Diseases, 38(3):241–248, March 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Boitard:2016:FDC

[BBL^{+}16] P-M. Boitard, M. Baud, S. Labrut, C. de Boisséson, M. Jamin, and L. Bigarré. First detection of Cyprinid Herpesvirus 2 (CyHV-2) in goldfish (Carassius auratus) in France. Journal of Fish Diseases, 39(6):673–680, June 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bowden:2018:CPV

[BBP18] T. J. Bowden, I. R. Bricknell, and B. M. Preziosi. Comparative pathogenicity of Vibrio spp., Photobacterium damselae ssp. damselae and five isolates of Aeromonas salmonicida ssp. achromogenes in juvenile Atlantic halibut (Hippoglossus hippoglossus). Journal of Fish Diseases, 41(1):79–86, January 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Beraldo:2011:ROA

[BC11] P. Beraldo and B. Canavese. Recovery of opercular anomalies in gilthead sea bream, Sparus aurata L.: morphological and morphometric analysis. Journal of Fish Diseases, 34(1):21– 30, January 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Betancor:2013:OSH

[BCBS^{+}13] M. B. Betancor, M. J. Caballero, T. Benítez-Santana, R. Saleh, J. Roo, E. Atalah, and M. Izquierdo. Oxidative status and histological changes in sea bass larvae muscle in response to high dietary content of docosahexaenoic acid DHA. Journal of Fish Diseases, 36(5):453–465, May 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Boylan:2017:SGM

[BCG^{+}17] S. M. Boylan, A. Camus, J. Gaskins, J. Oliverio, M. Parks, A. Davis, and J. Cassel. Spondylosis in a green moray eel, Gymnothorax funebris (Ranzani 1839), with swim bladder hyperinflation. Journal of Fish Diseases, 40(7):963–969, July 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Borucinska:2013:PHI

[BCJ13] J. D. Borucinska, J. J. Cielocha, and K. Jensen. The parasite–host interface in the zonetail butterfly ray, Gymnura zonura (Bleeker), infected with Hexacanalis folifer (Cestoda: Lecanicephalidea). Journal of Fish Diseases, 36(1):1–8, January 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bitchava:2019:OMC

[BCL^{+}19] Konstantina Bitchava, Taxiarchis Chassalevris, Eirini Lampou, Fotini Athanassopoulou, Vangelis Economou, and Chrysostomos I. Dovas. Occurrence and molecular characterization of betanodaviruses in fish and invertebrates of the Greek territorial waters. Journal of Fish Diseases, 42(12): 1773–1783, December 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Balmori-Cedeno:2019:ARG

[BCLM^{+}19] Jaramar Balmori-Cedeño, Juan-Ting Liu, Ehab Misk, Brandon N. Lillie, and John S. Lumsden. Autophagy-related genes in rainbow trout Oncorhynchus mykiss (Walbaum) gill epithelial cells and their role in nutrient restriction. Journal of Fish Diseases, 42(4):549–558, April 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Borges:2017:MCC

[BCM^{+}17] J. N. Borges, V. S. Costa, C. Mantovani, E. Barros, E. G. N. Santos, C. L. Mafra, and C. P. Santos. Molecular characterization and confocal laser scanning microscopic study of Pygidiopsis macrostomum (Trematoda: Heterophyidae) parasites of guppies Poecilia vivipara. Journal of Fish Diseases, 40(2):191–203, February 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bader:2017:OCJ

[BCT^{+}17] C. Bader, J. Jesudoss Chelladurai, K. Thompson, D. Starling, and M. T. Brewer. Outbreak of Cleidodiscus in juvenile black crappies (Pomoxis nigromaculatus) and bath treatment with praziquantel. Journal of Fish Diseases, 40 (11):1737–1739, November 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Boylan:2015:LNC

[BCW^{+}15] S. M. Boylan, A. Camus, T. Waltzek, L. Yarbrough, S. R. Miller, and S. Howard. Liquid nitrogen cryotherapy for fibromas in tarpon, Megalops atlanticus, Valenciennes 1847, and neoplasia in lined sea horse, Hippocampus erectus, Perry 1810. Journal of Fish Diseases, 38(7):681–685, July 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bowater:2018:ESA

[BDB^{+}18] R. O. Bowater, M. M. Dennis, D. Blyde, B. Stone, A. C. Barnes, J. Delamare-Deboutteville, M. A. Horton, M. White, K. Condon, and R. Jones. Epizootics of Streptococcus agalactiae infection in captive rays from Queensland, Australia. Journal of Fish Diseases, 41(2):223–232, February 2018. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bui:2018:SBR

[BDD^{+}18] S. Bui, S. Dalvin, T. Dempster, O. F. Skulstad, R. B. Edvardsen, A. Wargelius, and F. Oppedal. Susceptibility, behaviour, and retention of the parasitic salmon louse (Lepeophtheirus salmonis) differ with Atlantic salmon population origin. Journal of Fish Diseases, 41(3):431–442, March 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bosi:2018:HIC

[BDM^{+}18] G. Bosi, J. A. DePasquale, M. Manera, G. Castaldelli, L. Giari, and B. Sayyaf Dezfuli. Histochemical and immunohistochemical characterization of rodlet cells in the intestine of two teleosts, Anguilla anguilla and Cyprinus carpio. Journal of Fish Diseases, 41(3):475–485, March 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bilen:2019:NPT

[BE19] Soner Bilen and Housam Taher Abdul Ghani Elbeshti. A new potential therapeutic remedy against Aeromonas hydrophila infection in rainbow trout (Oncorhynchus mykiss) using tetra, Cotinus coggygria. Journal of Fish Diseases, 42(10):1369–1381, October 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Busschaert:2015:CGS

[BFC^{+}15] P. Busschaert, I. Frans, S. Crauwels, B. Zhu, K. Willems, P. Bossier, C. Michiels, K. Verstrepen, B. Lievens, and H. Rediers. Comparative genome sequencing to assess the genetic diversity and virulence attributes of 15 Vibrio anguillarum isolates. Journal of Fish Diseases, 38(9):795–807, September 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bowater:2012:NOS

[BFFA^{+}12] R. O. Bowater, J. Forbes-Faulkner, I. G. Anderson, K. Condon, B. Robinson, F. Kong, G. L. Gilbert, A. Reynolds, S. Hyland, G. McPherson, J. O’ Brien, and D. Blyde. Natural outbreak of Streptococcus agalactiae (GBS) infection in wild giant Queensland grouper, Epinephelus lanceolatus (Bloch), and other wild fish in northern Queensland, Australia. Journal of Fish Diseases, 35(3):173–186, March 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Birkbeck:2011:FIF

[BFJ11] T. H. Birkbeck, S. W. Feist, and D. W. Verner Jeffreys. Francisella infections in fish and shellfish. Journal of Fish Diseases, 34(3):173–187, March 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bluford:2017:IVG

[BGC^{+}17] J. Bluford, D. Gauthier, M. Colasanto, M. Rhodes, W. Vogelbein, and A. Haines. Identification of virulence genes in Vibrio spp. isolates from the 2009 Bermuda reef fish mortality event. Journal of Fish Diseases, 40(4):597–600, April 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Burr:2012:HAP

[BGCKF12] S. E. Burr, E. Goldschmidt-Clermont, P. Kuhnert, and J. Frey. Heterogeneity of Aeromonas populations in wild and farmed perch, Perca fluviatilis L. Journal of Fish Diseases, 35(8):607–613, August 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bosi:2017:PRI

[BGD^{+}17] G. Bosi, L. Giari, J. A. DePasquale, A. Carosi, M. Lorenzoni, and B. Sayyaf Dezfuli. Protective responses of intestinal mucous cells in a range of fish–helminth systems. Journal of Fish Diseases, 40(8):1001–1014, August 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bjornsdottir:2011:VPM

[BGG11] B. Bjornsdottir, T. Gudmundsdottir, and B. K. Gudmundsdottir. Virulence properties of Moritella viscosa extracellular products. Journal of Fish Diseases, 34(5):333–343, May 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Behrmann-Godel:2014:GWA

[BGRB14] J. Behrmann-Godel, S. Roch, and A. Brinker. Gill worm Ancyrocephalus percae (Ergens 1966) outbreak negatively impacts the Eurasian perch Perca fluviatilis L. stock of Lake Constance, Germany. Journal of Fish Diseases, 37(10):925– 930, October 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bekker:2011:PGP

[BHA^{+}11] A. Bekker, C. Hugo, J. Albertyn, C. E. Boucher, and R. R. Bragg. Pathogenic Gram-positive cocci in South African rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 34(6):483–487, June 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Braceland:2017:TPA

[BHA^{+}17] M. Braceland, K. Houston, A. Ashby, C. Matthews, H. Haining, H. Rodger, and P. D. Eckersall. Technical pre-analytical effects on the clinical biochemistry of Atlantic salmon (Salmo salar L.). Journal of Fish Diseases, 40(1):29–40, January 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Beaz-Hidalgo:2013:ASW

[BHF13] R. Beaz-Hidalgo and M. J. Figueras. Aeromonas spp. whole genomes and virulence factors implicated in fish disease. Journal of Fish Diseases, 36(4):371–388, April 2013. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bjorgen:2019:MFC

[BHO^{+}19] Håvard Bjørgen, Randi Haldorsen, Øyvind Oaland, Agnar Kvellestad, Dhamotharan Kannimuthu, Espen Rimstad, and Erling Olaf Koppang. Melanized focal changes in skeletal muscle in farmed Atlantic salmon after natural infection with Piscine orthoreovirus (PRV). Journal of Fish Diseases, 42 (6):935–945, June 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bass:2017:SMP

[BHT^{+}17] A. L. Bass, S. G. Hinch, A. K. Teffer, D. A. Patterson, and K. M. Miller. A survey of microparasites present in adult migrating Chinook salmon (Oncorhynchus tshawytscha) in south-western British Columbia determined by high-throughput quantitative polymerase chain reaction.

Journal of Fish Diseases, 40(4):453–477, April 2017. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Boylan:2011:CRH

[BHW^{+}11] S. M. Boylan, C. A. Harms, T. Waltzek, J. M. Law, M. Garner, J. Cassell, M. H. Fatzinger, and P. Govett. Clinical report: hyperplastic adipose lids in mackerel scad, Decapterus macarellus (Cuvier). Journal of Fish Diseases, 34(12):921– 925, December 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Binesh:2014:ETC

[Bin14] C. P. Binesh. Elevation of temperature and crowding trigger acute viral nervous necrosis in zebra fish, Brachydanio rerio (Hamilton-Buchanan), subclinically infected with betanodavirus. Journal of Fish Diseases, 37(3):279–282, March 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Buchmann:2012:EPD

[BK12] K. Buchmann and P. Kania. Emerging Pseudoterranova decipiens (Krabbe, 1878) problems in Baltic cod, Gadus morhua L., associated with grey seal colonization of spawning grounds. Journal of Fish Diseases, 35(11):861–866, November 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bartkova:2017:IRA

[BKD17] S. Bartkova, B. Kokotovic, and I. Dalsgaard. Infection routes of Aeromonas salmonicida in rainbow trout monitored in vivo by real-time bioluminescence imaging. Journal of Fish Diseases, 40(1):73–82, January 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bjorgen:2018:EEC

[BKG^{+}18] H. Bjørgen, E. O. Koppang, G. Gunnes, I. Hordvik, T. Moldal, M. Kaldhusdal, and O. B. Dale. Ectopic epithelial cell clusters in salmonid intestine are associated with inflammation. Journal of Fish Diseases, 41(7):1031–1040, July 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bartkova:2017:DQA

[BKS^{+}17] S. Bartkova, B. Kokotovic, H. F. Skall, N. Lorenzen, and I. Dalsgaard. Detection and quantification of Aeromonas salmonicida in fish tissue by real-time PCR. Journal of Fish Diseases, 40(2):231–242, February 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bermudez:2018:FDN

[BLdA^{+}18] Roberto Bermúdez, Ana Paula Losada, Ana Manuela de Azevedo, Jorge Guerra-Varela, David Pérez-Fernández, Laura Sánchez, Francesc Padrós, Barbara Nowak, and María Isabel Quiroga. First description of a natural infection with spleen and kidney necrosis virus in zebrafish. Journal of Fish Diseases, 41(8):1283–1294, August 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Beck:2016:CHL

[BLF^{+}16] B. H. Beck, C. Li, B. D. Farmer, L. M. Barnett, M. D. Lange, and E. Peatman. A comparison of high- and lowvirulence Flavobacterium columnare strains reveals differences in iron acquisition components and responses to iron restriction. Journal of Fish Diseases, 39(3):259–268, March 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Burbank:2012:IBP

[BLFC12] D. R. Burbank, S. E. LaPatra, G. Fornshell, and K. D. Cain. Isolation of bacterial probiotic candidates from the gastrointestinal tract of rainbow trout, Oncorhynchus mykiss (Walbaum), and screening for inhibitory activity against Flavobacterium psychrophilum. Journal of Fish Diseases, 35(11):809–816, November 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bello-Lopez:2010:VTP

[BLFRCQ10] J. M. Bello-López, E. Fernández-Rendón, and E. Curiel-Quesada. In vivo transfer of plasmid pRAS1 between Aeromonas salmonicida and Aeromonas hydrophila in artificially infected Cyprinus carpio L. Journal of Fish Diseases, 33(3):251–259, March 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bigarre:2017:MII

[BLL^{+}17] L. Bigarré, M. Lesne, A. Lautraite, V. Chesneau, A. Leroux, M. Jamin, P. M. Boitard, A. Toffan, M. Prearo, S. Labrut, and P. Daniel. Molecular identification of iridoviruses infecting various sturgeon species in Europe. Journal of Fish Diseases, 40(1):105–118, January 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bradford:2010:IGK

[BLP10] M. J. Bradford, J. Lovy, and D. A. Patterson. Infection of gill and kidney of Fraser River sockeye salmon, Oncorhynchus nerka (Walbaum), by Parvicapsula minibicornis and its effect on host physiology. Journal of Fish Diseases, 33(9): 769–779, September 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Batista:2016:DBE

[BLSG^{+}16] F. M. Batista, M. López-Sanmartín, A. Grade, J. I. Navas, and F. Ruano. Detection of Bonamia exitiosa in the European flat oyster Ostrea edulis in southern Portugal. Journal of Fish Diseases, 39(5):607–611, May 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bullard:2013:FRF

[BMA13] S. A. Bullard, H. Mohammed, and C. R. Arias. First record of the fish pathogen Flavobacterium columnare genomovar II from bluegill, Lepomis macrochirus (Rafinesque), with observations on associated lesions. Journal of Fish Diseases, 36 (4):447–451, April 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bland:2012:DVR

[BMBS12] F. Bland, R. McIntosh, N. Bain, and M. Snow. Development and validation of a range of endogenous controls to support the implementation of practical Taqman real-time PCR-based surveillance for fish diseases within aquaculture. Journal of Fish Diseases, 35(6):447–454, June 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Barry:2014:IPA

[BMD^{+}14] J. Barry, J. McLeish, J. A. Dodd, J. F. Turnbull, P. Boylan, and C. E. Adams. Introduced parasite Anguillicola crassus infection significantly impedes swim bladder function in the European eel Anguilla anguilla (L.). Journal of Fish Diseases, 37(10):921–924, October 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Borucinska:2017:FSS

[BMGS17] J. D. Borucinska, D. Morka, Z. Grabowski, and H. Smith. A follow-up study of selected biomarkers of health in cod Gadus morhua L. collected from the southern Baltic off the Polish coast. Journal of Fish Diseases, 40(12):1883–1894, December 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Brosnahan:2019:NZR

[BMH^{+}19] Cara L. Brosnahan, John S. Munday, Hye Jeong Ha, Mark Preece, and John B. Jones. New Zealand rickettsia-like organism (NZ-RLO) and Tenacibaculum maritimum: Distribution and phylogeny in farmed Chinook salmon (Oncorhynchus tshawytscha). Journal of Fish Diseases, 42(1): 85–95, January 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bergmann:2017:CWD

[BMK17] S. M. Bergmann, E. S. Monro, and J. Kempter. Can water disinfection prevent the transmission of infectious koi herpesvirus to naïve carp? — A case report. Journal of Fish Diseases, 40(7):885–893, July 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Barato:2015:PSC

[BMMC^{+}15] P. Barato, E. R. Martins, J. Melo-Cristino, C. A. Iregui, and M. Ramirez. Persistence of a single clone of Streptococcus agalactiae causing disease in tilapia (Oreochromis sp.) cultured in Colombia over 8 years. Journal of Fish Diseases, 38(12):1083–1087, December 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Borzym:2014:FIH

[BMMP^{+}14] E. Borzym, M. Matras, J. Maj-Paluch, M. Baud, C. De Boisséson, C. Talbi, N. J. Olesen, and L. Bigarré. First isolation of hirame rhabdovirus from freshwater fish in Europe. Journal of Fish Diseases, 37(5):423–430, May 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Braceland:2015:SEN

[BMT^{+}15] M. Braceland, M. F. McLoughlin, J. Tinsley, C. Wallace, D. Cockerill, M. McLaughlin, and P. D. Eckersall. Serum enolase: a non-destructive biomarker of white skeletal myopathy during pancreas disease (PD) in Atlantic salmon Salmo salar L. Journal of Fish Diseases, 38(9):821–831, September 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Babu:2011:ECF

[BNC^{+}11] V. S. Babu, K. S. N. Nambi, V. Chandra, V. P. Ishaq Ahmed, R. Bhonde, and A. S. Sahul Hameed. Establishment and characterization of a fin cell line from Indian walking catfish, Clarias batrachus (L.). Journal of Fish Diseases, 34(5):355– 364, May 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Barry:2017:EFO

[BND^{+}17] J. Barry, M. Newton, J. A. Dodd, D. Evans, J. Newton, and C. E. Adams. The effect of foraging and ontogeny on the prevalence and intensity of the invasive parasite Anguillicola crassus in the European eel Anguilla anguilla. Journal of Fish Diseases, 40(9):1213–1222, September 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bravo:2013:ETU

[BNS13] S. Bravo, M. Nuñez, and M. T. Silva. Efficacy of the treatments used for the control of Caligus rogercresseyi infecting Atlantic salmon, Salmo salar L., in a new fish-farming location in Region XI, Chile. Journal of Fish Diseases, 36(3): 221–228, March 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bowater:2015:ECF

[BO15] R. O. Bowater and P. J. O’Donoghue. An epizootic of chilodonelliasis in farmed barramundi Lates calcarifer (Bloch), a case report. Journal of Fish Diseases, 38(10): 931–936, October 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Borucinska:2012:MFC

[BOH^{+}12] J. D. Borucinska, O. A. Obasa, N. M. Haffey, J. P. Scott, L. N. Williams, S. M. Baker, S. J. Min, A. Kaplan, and R. Mudimala. Morphological features of coronary arteries and lesions in hearts from five species of sharks collected from the northwestern Atlantic Ocean. Journal of Fish Diseases, 35(10):741–754, October 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Benedicenti:2019:ETA

[BPCS19] Ottavia Benedicenti, Tom G. Pottinger, Catherine Collins, and Christopher J. Secombes. Effects of temperature on amoebic gill disease development: Does it play a role? Journal of Fish Diseases, 42(9):1241–1258, September 2019. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bedendo:2018:DCR

[BPF^{+}18] Giulia Bedendo, Valentina Panzarin, Andrea Fortin, Gianpiero Zamperin, Tobia Pretto, Alessandra Buratin, Rosita Quartesan, Matteo Sabbion, Cristian Salogni, Francesco Pascoli, and Anna Toffan. Detection and characterization of a rhabdovirus causing mortality in black bullhead catfish, Ameiurus melas. Journal of Fish Diseases, 41(7):1063–1075, July 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Boonyawiwat:2017:IFM

[BPK^{+}17] V. Boonyawiwat, T. Patanasatienkul, J. Kasornchandra, C. Poolkhet, S. Yaemkasem, L. Hammell, and J. Davidson. Impact of farm management on expression of early mortality syndrome/acute hepatopancreatic necrosis disease (EMS/AHPND) on penaeid shrimp farms in Thailand. Journal of Fish Diseases, 40(5):649–659, May 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bethke:2016:IAS

[BPMI^{+}16] J. Bethke, M. Poblete-Morales, R. Irgang, A. Yáñez, and R. Avendaño-Herrera. Iron acquisition and siderophore production in the fish pathogen Renibacterium salmoninarum. Journal of Fish Diseases, 39(11):1275–1283, November 2016.

CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Braden:2010:TBS

[BPPKJ10] L. M. Braden, G. Prosperi-Porta, E. Kim, and S. R. M. Jones. Tetracapsuloides bryosalmonae in spawning pink salmon, Oncorhynchus gorbuscha (Walbaum), in the Quinsam River, British Columbia, Canada. Journal of Fish Diseases, 33(7):617–621, July 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Blanco:2012:FOC

[BQP12] A. Blanco, P. Quintas, and M. Planas. First observations of conjoined twins in newborn seahorses, Hippocampus guttulatus Cuvier. Journal of Fish Diseases, 35(9):705–708, September 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bethke:2012:DNA

[BRB^{+}12] J. Bethke, V. Rojas, J. Berendsen, C. Cárdenas, F. Guzmán, J. A. Gallardo, and L. Mercado. Development of a new antibody for detecting natural killer enhancing factor (NKEF)- like protein in infected salmonids. Journal of Fish Diseases, 35(5):379–388, May 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bartsch:2013:FSL

[BRL^{+}13] A. Bartsch, S. M. C. Robinson, M. Liutkus, K. P. Ang, J. Webb, and C. M. Pearce. Filtration of sea louse, Lepeophtheirus salmonis, copepodids by the blue mussel, Mytilus edulis, and the Atlantic sea scallop, Placopecten magellanicus, under different flow, light and copepodid-density regimes. Journal of Fish Diseases, 36(3):361–370, March 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Binesh:2013:FRV

[BRMG13] C. P. Binesh, K. Renuka, N. Malaichami, and C. Greeshma. First report of viral nervous necrosis-induced mass mortality in hatchery-reared larvae of clownfish, Amphiprion sebae Bleeker. Journal of Fish Diseases, 36(12):1017–1020, December 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Baldissera:2018:PSD

[BSB^{+}18] M. D. Baldissera, C. F. Souza, N. B. Bottari, C. M. Verdi, R. C. V. Santos, B. S. Vizzotto, and B. Baldisserotto. Purinergic signalling displays an anti-inflammatory profile in the spleen of fish experimentally infected with Aeromonas caviae: Modulation of the immune response. Journal of Fish Diseases, 41(4):683–687, April 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Burnley:2010:RCF

[BSBH10] T. A. Burnley, H. Stryhn, H. J. Burnley, and K. L. Hammell. Randomized clinical field trial of a bacterial kidney disease vaccine in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 33(7):545–557, July 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bandin:2014:PVW

[BSC^{+}14] I. Bandín, S. Souto, J. M. Cutrín, C. López-Vázquez, J. G. Olveira, C. Esteve, E. Alcaide, and C. P. Dopazo. Presence of viruses in wild eels Anguilla anguilla L., from the Albufera Lake (Spain). Journal of Fish Diseases, 37(7):597–607, July 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Baldissera:2018:SAD

[BSD^{+}18a] M. D. Baldissera, C. F. Souza, P. H. Doleski, S. G. Monteiro, A. S. da Silva, and B. Baldisserotto. Serum adenosine deaminase and xanthine oxidase activities in silver catfish naturally infected with Ichthyophthirius multifiliis: the influence of these enzymes on inflammatory and oxidative status. Journal of Fish Diseases, 41(2):263–268, February 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Baldissera:2018:ICA

[BSD^{+}18b] M. D. Baldissera, C. F. Souza, P. H. Doleski, K. L. S. Moreira, M. L. da Veiga, M. I. U. M. da Rocha, R. C. V. Santos, and B. Baldisserotto. Involvement of cholinergic and adenosinergic systems on the branchial immune response of experimentally infected silver catfish with Streptococcus agalactiae. Journal of Fish Diseases, 41(1):27–32, January 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Baldissera:2018:IXO

[BSD^{+}18c] M. D. Baldissera, C. F. Souza, P. H. Doleski, R. C. V. Santos, R. P. Raffin, and B. Baldisserotto. Involvement of xanthine oxidase inhibition with the antioxidant property of nanoencapsulated Melaleuca alternifolia essential oil in fish experimentally infected with Pseudomonas aeruginosa. Journal of Fish Diseases, 41(5):791–796, May 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Baldissera:2018:XOA

[BSD^{+}18d] M. D. Baldissera, C. F. Souza, P. H. Doleski, C. C. Zeppenfeld, S. Descovi, A. S. Da Silva, and B. Baldisserotto. Xanthine oxidase activity exerts pro-oxidative and proinflammatory effects in serum of silver catfish fed with a diet contaminated with aflatoxin B_{1}. Journal of Fish Diseases, 41(7):1153–1158, July 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bergmann:2010:SKC

[BSK^{+}10] S. M. Bergmann, J. Sadowski, M. Kiełpiński, M. Bartłomiejczyk, D. Fichtner, R. Riebe, M. Lenk, and J. Kempter. Susceptibility of koi × crucian carp and koi × goldfish hybrids to koi herpesvirus (KHV) and the development of KHV disease (KHVD). Journal of Fish Diseases, 33(3):267–272, March 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bailey:2018:BTS

[BSPS^{+}18] C. Bailey, H. Schmidt-Posthaus, H. Segner, T. Wahli, and N. Strepparava. Are brown trout Salmo trutta fario and rainbow trout Oncorhynchus mykiss two of a kind? A comparative study of salmonids to temperature-influenced Tetracapsuloides bryosalmonae infection. Journal of Fish Diseases, 41(2):191–198, February 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Baldissera:2018:BBB

[BSSB18] M. D. Baldissera, C. F. Souza, R. C. V. Santos, and B. Baldisserotto. Blood–brain barrier breakdown and myeloperoxidase activity in silver catfish experimentally infected with Pseudomonas aeruginosa. Journal of Fish Diseases, 41(2): 209–213, February 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Boerlage:2017:CDC

[BSSH17] A. S. Boerlage, H. Stryhn, J. Sanchez, and K. L. Hammell. Case definition for clinical and subclinical bacterial kidney disease (BKD) in Atlantic Salmon (Salmo salar L.) in New Brunswick, Canada. Journal of Fish Diseases, 40(3):395– 409, March 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Baldissera:2018:ACI

[BSV^{+}18] M. D. Baldissera, C. F. Souza, C. M. Verdi, K. L. M. dos Santos, M. L. Da Veiga, M. I. U. M. da Rocha, R. C. V. Santos, B. S. Vizzotto, and B. Baldisserotto. Aeromonas caviae inhibits hepatic enzymes of the phosphotransfer network in experimentally infected silver catfish: Impairment on bioenergetics. Journal of Fish Diseases, 41(3):469–474, March 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bailey:2017:WGA

[BSW17] C. Bailey, H. Segner, and T. Wahli. What goes around comes around: an investigation of resistance to proliferative kidney disease in rainbow trout Oncorhynchus mykiss (Walbaum) following experimental re-exposure. Journal of Fish Diseases, 40(11):1599–1612, November 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Baily:2011:FLB

[BSWS11] J. E. Baily, J. L. Smith, R. Wootten, and C. Sommerville. The fate of Lernaeocera branchialis (L.) (Crustacea; Copepoda) in Atlantic cod, Gadus morhua L. Journal of Fish Diseases, 34(2):139–147, February 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Boxaspen:2013:SL

[BT13] Karin Boxaspen and Ole Torrissen. Sea lice 2012. Journal of Fish Diseases, 36(3):169, March 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Birkbeck:2014:VSV

[BT14] T. H. Birkbeck and J. W. Treasurer. Vibrio splendidus, Vibrio ichthyoenteri and Vibrio pacinii isolated from the digestive tract microflora of larval ballan wrasse, Labrus bergylta Ascanius, and goldsinny wrasse, Ctenolabrus rupestris (L.). Journal of Fish Diseases, 37(1):69–74, January 2014. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Braceland:2017:SPR

[BTC^{+}17] M. Braceland, J. Tinsley, D. Cockerill, R. Bickerdike, M. F. McLoughlin, and P. D. Eckersall. Selective precipitation reaction: a novel diagnostic test for tissue pathology in Atlantic salmon, Salmo salar, infected with salmonid alphavirus (SAV3). Journal of Fish Diseases, 40(8):1077–1087, August 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bauer:2018:RIP

[BTN^{+}18] Julia Bauer, Felix Teitge, Lisa Neffe, Mikolaj Adamek, Arne Jung, Christina Peppler, Dieter Steinhagen, and Verena Jung-Schroers. Recommendations for identifying pathogenic Vibrio spp. as part of disease surveillance programmes in recirculating aquaculture systems for Pacific white shrimps (Litopenaeus vannamei). Journal of Fish Diseases, 41(12): 1877–1897, December 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Buchmann:2012:FIR

[Buc12] K. Buchmann. Fish immune responses against endoparasitic nematodes — experimental models. Journal of Fish Diseases, 35(9):623–635, September 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bozzetta:2010:MMI

[BVG^{+}10] E. Bozzetta, K. Varello, I. Giorgi, M. L. Fioravanti, M. Pezzolato, R. G. Zanoni, and M. Prearo. Mycobacterium marinum infection in a hybrid striped bass farm in Italy. Journal of Fish Diseases, 33(9):781–785, September 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Biering:2016:CMA

[BVK^{+}16] E. Biering, Ø. Vaagnes, B. Krossøy, S. Gulla, and Dj Colquhoun. Challenge models for atypical Aeromonas salmonicida and Vibrio anguillarum in farmed Ballan wrasse (Labrus bergylta) and preliminary testing of a trial vaccine against atypical Aeromonas salmonicida. Journal of Fish Diseases, 39(10):1257–1261, October 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Burioli:2018:NDG

[BVL^{+}18] Erika A. V. Burioli, Katia Varello, Antonio Lavazza, Elena Bozzetta, Marino Prearo, and Maryline Houssin. A novel divergent group of Ostreid herpesvirus 1 μ Var variants associated with a mortality event in Pacific oyster spat in Normandy (France) in 2016. Journal of Fish Diseases, 41(11): 1759–1769, November 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Burioli:2018:FDM

[BVT^{+}18] E. A. V. Burioli, K. Varello, S. Trancart, E. Bozzetta, A. Gorla, M. Prearo, and M. Houssin. First description of a mortality event in adult Pacific oysters in Italy associated with infection by a Tenacibaculum soleae strain. Journal of Fish Diseases, 41(2):215–221, February 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bolton-Warberg:2018:OCF

[BW18] M. Bolton-Warberg. An overview of cleaner fish use in Ireland. Journal of Fish Diseases, 41(6):935–939, June 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Blazer:2017:TWS

[BWB^{+}17] V. S. Blazer, H. L. Walsh, R. P. Braham, C. M. Hahn, P. Mazik, and P. B. McIntyre. Tumours in white suckers from Lake Michigan tributaries: pathology and prevalence. Journal of Fish Diseases, 40(3):377–393, March 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Browning:2013:UPC

[BWM^{+}13] Z. S. Browning, A. A. Wilkes, D. S. MacKenzie, R. M. Patterson, and M. W. Lenox. Using PET/ CT imaging to characterize 18 F-fluorodeoxyglucose utilization in fish. Journal of Fish Diseases, 36(11):911–919, November 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bockerman:2011:PPM

[BWNS^{+}11] I. Böckerman, C. R. Wiik-Nielsen, H. Sindre, R. Johansen, and T. Tengs. Prevalence of piscine myocarditis virus (PMCV) in marine fish species. Journal of Fish Diseases, 34(12):955–957, December 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Buss:2019:BC

[BWP^{+}19a] Jessica J. Buss, Kathryn H. Wiltshire, Thomas A. A. Prowse, James O. Harris, and Marty R. Deveney. Back cover. Journal of Fish Diseases, 42(1):ii, January 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Buss:2019:BOA

[BWP^{+}19b] Jessica J. Buss, Kathryn H. Wiltshire, Thomas A. A. Prowse, James O. Harris, and Marty R. Deveney. Bonamia in Ostrea angasi: Diagnostic performance, field prevalence and intensity. Journal of Fish Diseases, 42(1):63–74, January 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Burns:2018:TCI

[BWS^{+}18] A. R. Burns, V. Watral, S. Sichel, S. Spagnoli, A. V. Banse, E. Mittge, T. J. Sharpton, K. Guillemin, and M. L. Kent. Transmission of a common intestinal neoplasm in zebrafish by cohabitation. Journal of Fish Diseases, 41(4):569–579, April 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bergmann:2017:VKA

[BWZ^{+}17] S. M. Bergmann, Q. Wang, W. Zeng, Y. Li, Y. Wang, M. Matras, M. Reichert, D. Fichtner, M. Lenk, T. Morin, N. J. Olesen, H. F. Skall, P-Y. Lee, S. Zheng, S. Monaghan, S. Reiche, W. Fuchs, M. Kotler, K. Way, G. Bräuer, K. Böttcher, A. Kappe, and J. Kielpinska. Validation of a KHV antibody enzyme-linked immunosorbent assay (ELISA). Journal of Fish Diseases, 40(11):1511–1527, November 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Beck:2019:AAB

[BYAS^{+}19] Benjamin H. Beck, Mediha Yildirim-Aksoy, Craig A. Shoemaker, Sidney Adam Fuller, and Eric Peatman. Antimicrobial activity of the biopolymer chitosan against Streptococcus iniae. Journal of Fish Diseases, 42(3):371–377, March 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chettri:2018:EAI

[CAJD^{+}18] Jiwan K. Chettri, Azmi Al-Jubury, Inger Dalsgaard, Peter M. H. Heegaard, and Kurt Buchmann. Experimental anal infection of rainbow trout with Flavobacterium psychrophilum: a novel challenge model. Journal of Fish Diseases, 41(12): 1917–1919, December 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chance:2018:ERE

[CAS^{+}18] R. J. Chance, Z. Allcock, C. J. Secombes, B. Collet, and C. Collins. Effect of repeated exposure to AQUI-S(R) on the viability and growth of Neoparamoeba perurans. Journal of Fish Diseases, 41(2):291–298, February 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cobcroft:2013:UIE

[CB13] J. M. Cobcroft and S. C. Battaglene. Ultraviolet irradiation is an effective alternative to ozonation as a sea water treatment to prevent Kudoa neurophila (Myxozoa: Myxosporea) infection of striped trumpeter, Latris lineata (Forster). Journal of Fish Diseases, 36(1):57–65, January 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Camus:2013:SMA

[CBC^{+}13] A. Camus, A. Berliner, T. Clauss, N. Hatcher, and D. Marancik. Serratia marcescens associated ampullary system infection and septicaemia in a bonnethead shark, Sphyrna tiburo (L.). Journal of Fish Diseases, 36(10):891–895, October 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Camus:2016:ALR

[CBFM16] A. Camus, C. Brown, C. Field, and A. McDermott. Amyloidlike renal lesions in a roughtail stingray Dasyatis centroura (Mitchill 1815). Journal of Fish Diseases, 39(5):613–617, May 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Camus:2015:EXA

[CBH^{+}15] A. Camus, A. Berliner, M. Hyatt, N. Hatcher, and T. Clauss. Exophiala xenobiotica aerocystitis in a Queensland grouper Epinephelus lanceolatus (Bloch). Journal of Fish Diseases, 38(2):221–225, February 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cadoret:2013:EFM

[CBLN13] K. Cadoret, A. R. Bridle, M. J. Leef, and B. F. Nowak. Evaluation of fixation methods for demonstration of Neoparamoeba perurans infection in Atlantic salmon, Salmo salar L., gills. Journal of Fish Diseases, 36(10):831–839, October 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Clements:2017:EOA

[CBM^{+}17] J. C. Clements, D. Bourque, J. McLaughlin, M. Stephenson, and L. A. Comeau. Extreme ocean acidification reduces the susceptibility of eastern oyster shells to a polydorid parasite. Journal of Fish Diseases, 40(11):1573–1585, November 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Clements:2018:WDA

[CBM^{+}18] J. C. Clements, D. Bourque, J. McLaughlin, M. Stephenson, and L. A. Comeau. Wanted dead or alive: Polydora websteri recruit to both live oysters and empty shells of the eastern oyster, Crassostrea virginica. Journal of Fish Diseases, 41 (5):855–858, May 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chang:2019:TMC

[CBW19] Carolyn T. Chang, Samuel Benedict, and Christopher M. Whipps. Transmission of Mycobacterium chelonae and Mycobacterium marinum in laboratory zebrafish through live feeds. Journal of Fish Diseases, 42(10):1425–1431, October 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Castillo-Cerda:2014:DPA

[CCCTA^{+}14] M. T. Castillo-Cerda, L. Cottet, D. Toro-Ascuy, E. Spencer, and M. Cortez-San Martín. Development of plaque assay for Chilean Infectious Salmon Anaemia Virus, application for virus purification and titration in salmon ASK cells. Journal of Fish Diseases, 37(11):989–995, November 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chakraborty:2019:VVT

[CCH^{+}19] Setu Chakraborty, Trung Cao, Ahmed Hossain, Hajarooba Gnanagobal, Ignacio Vasquez, Danny Boyce, and Javier Santander. Vibrogen-2 vaccine trial in lumpfish (Cyclopterus lumpus) against Vibrio anguillarum. Journal of Fish Diseases, 42(7):1057–1064, July 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Choi:2013:RDS

[CCHL13] K. Choi, W. G. Cope, C. A. Harms, and J. M. Law. Rapid decreases in salinity, but not increases, lead to immune dysregulation in Nile tilapia, Oreochromis niloticus (L.). Journal of Fish Diseases, 36(4):389–399, April 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Campa-Cordova:2017:RPL

[CCLGRM^{+}17] A. I. Campa-Córdova, A. F. León-Gallo, A. Romero-Maldonado, A. C. Ibarra-Serrano, S. Rosales-Mendoza, I. Hirono, and C. Angulo. Recombinant PirA-like toxin protects shrimp against challenge with Vibrio parahaemolyticus, the aetiological agent of acute hepatopancreatic necrosis disease. Journal of Fish Diseases, 40(11):1725–1729, November 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chen:2016:PCM

[CCLL16] Z-Y. Chen, P. P. Chiou, C-J. Liou, and Y-S. Lai. Production and characterization of a monoclonal antibody against a late gene encoded by grouper iridovirus 64L. Journal of Fish Diseases, 39(2):129–141, February 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cornwell:2011:ENI

[CCM^{+}11] E. R. Cornwell, M. J. Cinelli, D. M. McIntosh, G. S. Blank, G. A. Wooster, G. H. Groocock, R. G. Getchell, and P. R. Bowser. Epizootic Nocardia infection in cultured weakfish, Cynoscion regalis (Bloch and Schneider). Journal of Fish Diseases, 34(7):567–571, July 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Coscelli:2017:ASO

[CCMA^{+}17] G. A. Coscelli, C. Casabonne, E. Morón-Alcain, N. Arancegui, and F. A. Vigliano. Aeromonas sobria, an outbreak of natural infection in cultured silver catfish Rhamdia quelen (Quoy & Gaimard, 1824) in Argentina. Journal of Fish Diseases, 40(12):1929–1933, December 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cordero:2016:CGS

[CCME16] H. Cordero, A. Cuesta, J. Meseguer, and M. A. Esteban. Characterization of the gilthead seabream (Sparus aurata L.) immune response under a natural lymphocystis disease virus outbreak. Journal of Fish Diseases, 39(12):1467–1476, December 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cheng:2011:KHE

[CCT^{+}11] L. Cheng, C-Y. Chen, M-A. Tsai, P-C. Wang, J-P. Hsu, R-S. Chern, and S-C. Chen. Koi herpesvirus epizootic in cultured carp and koi, Cyprinus carpio L., in Taiwan. Journal of Fish Diseases, 34(7):547–554, July 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chen:2015:INC

[CCW^{+}15] J. Chen, D. Chee, Y. Wang, G. Y. Lim, S. M. Chong, Y. N. Lin, and T. Huangfu. Identification of a novel cyprinid herpesvirus 3 genotype detected in koi from the East Asian and South-East Asian Regions. Journal of Fish Diseases, 38(10): 915–923, October 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chenia:2017:CVC

[CD17] H. Y. Chenia and S. Duma. Characterization of virulence, cell surface characteristics and biofilm-forming ability of Aeromonas spp. isolates from fish and sea water. Journal of Fish Diseases, 40(3):339–350, March 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cunha:2019:TEP

[CdAPdC^{+}19] Ana C. Luz Cunha, Vitor de A. Pontinha, Maria Alcina M. de Castro, Simone Sühnel, Sthefanie C. Medeiros, Ângela M. Moura da Luz, Ricardo Harakava, Leonardo Tachibana, Danielle F. Mello, Naissa M. Danielli, Alcir L. Dafre, Aimê R. M. Magalhães, and José Luiz P. Mouriño. Two epizootic Perkinsus spp. events in commercial oyster farms at Santa Catarina, Brazil. Journal of Fish Diseases, 42(3):455–463, March 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cocchi:2018:SPI

[CDD^{+}18] Monia Cocchi, Gabrita De Zan, Tiziana Di Giusto, Silvia Deotto, Tobia Pretto, Amedeo Manfrin, Romina Brunetta, and Anna Toffan. Systemic polymicrobic infection involving Shewanella putrefaciens group in koi. Journal of Fish Diseases, 41(12):1929–1931, December 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Carraro:2018:DRT

[CDF^{+}18] R. Carraro, G. Dalla Rovere, S. Ferraresso, L. Carraro, R. Franch, A. Toffan, F. Pascoli, T. Patarnello, and L. Bargelloni. Development of a real-time PCR assay for rapid detection and quantification of Photobacterium damselae subsp. piscicida in fish tissues. Journal of Fish Diseases, 41(2):247–254, February 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Crisafi:2014:ARR

[CDG^{+}14] F. Crisafi, R. Denaro, M. Genovese, M. Yakimov, and L. Genovese. Application of relative real-time PCR to detect differential expression of virulence genes in Vibrio anguillarum under standard and stressed growth conditions. Journal of Fish Diseases, 37(7):629–640, July 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Camus:2013:FNS

[CDM^{+}13] A. C. Camus, J. A. Dill, A. J. McDermott, T. M. Clauss, A. L. Berliner, S. M. Boylan, and E. Soto. Francisella noatunensis subsp. orientalis infection in Indo-Pacific reef fish entering the United States through the ornamental fish trade. Journal of Fish Diseases, 36(7):681–684, July 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Camus:2016:EPA

[CDM^{+}16] A. Camus, J. Dill, A. McDermott, N. Hatcher, and M. Griffin. Edwardsiella piscicida-associated septicaemia in a blotched fantail stingray Taeniura meyeni (Müeller & Henle). Journal of Fish Diseases, 39(9):1125–1131, September 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chang:2017:ATZ

[CDW17] C. T. Chang, K. M. Doerr, and C. M. Whipps. Antibiotic treatment of zebrafish mycobacteriosis: tolerance and efficacy of treatments with tigecycline and clarithromycin. Journal of Fish Diseases, 40(10):1473–1485, October 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Clark:2011:AVE

[CG11] K. F. Clark and S. J. Greenwood. Aerococcus viridans expression of Cpn60 is associated with virulence during infection of the American lobster, Homarus americanus Milne Edwards. Journal of Fish Diseases, 34(11):831–843, November 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Camus:2019:SOS

[CGAS19] Al Camus, Matt Griffin, Abigail Armwood, and Esteban Soto. A spontaneous outbreak of systemic Edwardsiella piscicida infection in largemouth bass Micropterus salmoides (Lacépède, 1802) in California, USA. Journal of Fish Diseases, 42(5):759–763, May 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cortez-SanMartin:2012:ISS

[CGCAH12] M. Cortez-San Martín, A. González-Contreras, and R. Avendaño-Herrera. Infectivity study of Streptococcus phocae to seven fish and mammalian cell lines by confocal microscopy. Journal of Fish Diseases, 35(6):431–436, June 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cherif:2010:CST

[CGG^{+}10] N. Chérif, N. Gagné, D. Groman, F. Kibenge, T. Iwamoto, C. Yason, and S. Hammami. Complete sequencing of Tunisian redspotted grouper nervous necrosis virus betanodavirus capsid gene and RNA-dependent RNA polymerase gene. Journal of Fish Diseases, 33(3):231–240, March 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Caruso:2019:MOP

[CGP^{+}19] Claudio Caruso, Andrea Gustinelli, Paolo Pastorino, Pier Luigi Acutis, Riccardo Prato, Loretta Masoero, Simone Peletto, Maria Letizia Fioravanti, and Marino Prearo. Mortality outbreak by perch rhabdovirus in European perch (Perca fluviatilis) farmed in Italy: Clinical presentation and phylogenetic analysis. Journal of Fish Diseases, 42(5):773– 776, May 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chang:2017:PEA

[CGR17] C. J. Chang, J. Gu, and B. Robertsen. Protective effect and antibody response of DNA vaccine against salmonid alphavirus 3 (SAV3) in Atlantic salmon. Journal of Fish Diseases, 40(12):1775–1781, December 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chen:2019:IIR

[CGW^{+}19] Yaping Chen, Mengting Guo, Yanxue Wang, Xiaojing Hua, Shuai Gao, Yuting Wang, Dechuan Li, Wen Shi, Lijie Tang, Yijing Li, and Min Liu. Immunity induced by recombinant attenuated IHNV (infectious hematopoietic necrosis virus)- G^{N438A} expresses VP2 gene-encoded IPNV (infectious pancreatic necrosis virus) against both pathogens in rainbow trout. Journal of Fish Diseases, 42(5):631–642, May 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chettri:2010:FIV

[CHAB10] J. K. Chettri, L. Holten-Andersen, and K. Buchmann. Factors influencing in vitro respiratory burst assays with head kidney leucocytes from rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 33(7):593–602, July 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Collins:2017:GPP

[CHB^{+}17] C. Collins, M. Hall, D. Bruno, J. Sokolowska, L. Duncan, R. Yuecel, U. McCarthy, M. J. Fordyce, C. C. Pert, R. McIntosh, and Z. MacKay. Generation of Paramoeba perurans clonal cultures using flow cytometry and confirmation of virulence. Journal of Fish Diseases, 40(3):351–365, March 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Clode:2012:IMO

[CHF^{+}12] A. B. Clode, C. Harms, M. H. Fatzinger, F. Young, C. Colitz, and D. Wert. Identification and management of ocular lipid deposition in association with hyperlipidaemia in captive moray eels, Gymnothorax funebris Ranzani, Gymnothorax moringa (Cuvier) and Muraena retifera Goode and Bean. Journal of Fish Diseases, 35(9):683–693, September 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cao:2014:IBS

[CHH^{+}14] H. Cao, S. Hou, S. He, L. Lu, and X. Yang. Identification of a Bacteriovorax sp. isolate as a potential biocontrol bacterium against snakehead fish-pathogenic Aeromonas veronii. Journal of Fish Diseases, 37(3):283–289, March 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Castillo:2012:DFP

[CHV^{+}12] D. Castillo, G. Higuera, M. Villa, M. Middelboe, I. Dalsgaard, L. Madsen, and R. T. Espejo. Diversity of Flavobacterium psychrophilum and the potential use of its phages for protection against bacterial cold water disease in salmonids. Journal of Fish Diseases, 35(3):193–201, March 2012. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chang:2019:CDM

[CHWM^{+}19] Yao-Chung Chang, Harry Hamlin-Wright, Sean Monaghan, Tharangani Herath, Johanna Baily, Jorge del Pozo, Jamie Downes, Andrew Preston, Lynn Chalmers, Nilantha Jayasuriya, James E. Bron, Alexandra Adams, and Sophie Fridman. Changes in distribution, morphology and ultrastructure of chloride cell in Atlantic salmon during an AGD infection. Journal of Fish Diseases, 42(10):1433–1446, October 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Camus:2018:PDS

[CIA^{+}18] A. C. Camus, M. M. Ibrahiem, F. A. Alhizab, T. A. Aboellail, and A. M. Ibrahim. Poorly differentiated soft tissue sarcoma in an Arabian carpet shark Chiloscyllium arabicum (Gubanov): a case report. Journal of Fish Diseases, 41(1): 181–185, January 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cartes:2017:SAG

[CIL^{+}17] C. Cartes, A. Isla, F. Lagos, D. Castro, M. Muñoz, A. Yañez, D. Haussmann, and J. Figueroa. Search and analysis of genes involved in antibiotic resistance in Chilean strains of Piscirickettsia salmonis. Journal of Fish Diseases, 40(8): 1025–1039, August 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cipriano:2011:DFH

[Cip11] Rocco C. Cipriano. Dr. Frank Hetrick. Journal of Fish Diseases, 34(9):641, September 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cano:2015:EMT

[CJB^{+}15] I. Cano, C. Joiner, A. Bayley, G. Rimmer, K. Bateman, S. W. Feist, D. Stone, and R. Paley. An experimental means of transmitting pancreas disease in Atlantic salmon Salmo salar L. fry in freshwater. Journal of Fish Diseases, 38(3):271– 281, March 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chettri:2014:ERR

[CKJ^{+}14] J. K. Chettri, J. A. Kuhn, R. M. Jaafar, P. W. Kania, O. S. Møller, and K. Buchmann. Epidermal response of rainbow trout to Ichthyobodo necator: immunohistochemical and gene expression studies indicate a Th1-/ Th2-like switch. Journal of Fish Diseases, 37(9):771–783, September 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chistiakov:2010:VGI

[CKTT10] D. A. Chistiakov, F. V. Kabanov, O. D. Troepolskaya, and M. M. Tischenko. A variant of the interleukin-1β gene in European sea bass, Dicentrarchus labrax L., is associated with increased resistance against Vibrio anguillarum. Journal of Fish Diseases, 33(9):759–767, September 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Christoffersen:2017:ZDR

[CKvGJB17] T. B. Christoffersen, P. W. Kania, L. von Gersdorff Jørgensen, and K. Buchmann. Zebrafish Danio rerio as a model to study the immune response against infection with Ichthyophthirius multifiliis. Journal of Fish Diseases, 40(6):847–852, June 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cabon:2017:VSN

[CLC^{+}17] J. Cabon, L. Louboutin, J. Castric, S. Bergmann, G. Bovo, M. Matras, O. Haenen, N. J. Olesen, and T. Morin. Validation of a serum neutralization test for detection of antibodies specific to cyprinid herpesvirus 3 in infected common and koi carp (Cyprinus carpio). Journal of Fish Diseases, 40(5):687– 701, May 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Castells:2015:SDS

[CLGP15] M. L. Ogas Castells, J. L. La Torre, P. R. Grigera, and T. V. Poggio. A single dose of a suicidal DNA vaccine induces a specific immune response in salmonids. Journal of Fish Diseases, 38(6):581–587, June 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chin:2017:CSS

[CLHGK17] H. N. Chin, R. Loh, Y. C. Hong, and S. Gibson-Kueh. Case studies of spinal deformities in ornamental koi, Cyprinus carpio L. Journal of Fish Diseases, 40(1):65–71, January 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cheng:2010:ECV

[CLL^{+}10] T-C. Cheng, Y-S. Lai, I-Y. Lin, C-P. Wu, S-L. Chang, T-I. Chen, and M-S. Su. Establishment, characterization, virus susceptibility and transfection of cell lines from cobia, Rachycentron canadum (L.), brain and fin. Journal of Fish Diseases, 33(2):161–169, February 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Costa:2014:GSD

[CLLF14] F. A. A. Costa, C. A. G. Leal, R. C. Leite, and H. C. P. Figueiredo. Genotyping of Streptococcus dysgalactiae strains isolated from Nile tilapia, Oreochromis niloticus (L.). Journal of Fish Diseases, 37(5):463–469, May 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Costa:2015:CWC

[CLS^{+}15] F. A. A. Costa, C. A. G. Leal, N. D. Schuenker, R. C. Leite, and H. C. P. Figueiredo. Characterization of Weissella ceti infections in Brazilian rainbow trout, Oncorhynchus mykiss (Walbaum), farms and development of an oil-adjuvanted vaccine. Journal of Fish Diseases, 38(3):295–302, March 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chang:2014:DSD

[CLT^{+}14] C-I. Chang, C-F. Lee, J-M. Tsai, C-C. Wu, L-H. Chen, S-C. Chen, and K-J. Lin. Development of a selective and differential medium for capsulated Lactococcus garvieae. Journal of Fish Diseases, 37(8):719–728, August 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chang:2011:SDM

[CLW^{+}11] C-I. Chang, C-F. Lee, C-C. Wu, T. C. Cheng, J-M. Tsai, and K-J. Lin. A selective and differential medium for Vibrio alginolyticus. Journal of Fish Diseases, 34(3):227–234, March 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cheng:2013:WSS

[CLW^{+}13] L. Cheng, W-H. Lin, P-C. Wang, M-A. Tsai, J-P. Hsu, and S-C. Chen. White spot syndrome virus epizootic in cultured Pacific white shrimp Litopenaeus vannamei (Boone) in Taiwan. Journal of Fish Diseases, 36(12):977–985, December 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cai:2013:CEV

[CLWJ13] S. H. Cai, Y. S. Lu, Z. H. Wu, and J. C. Jian. Cloning, expression of Vibrio alginolyticus outer membrane protein-OmpU gene and its potential application as vaccine in crimson snapper, Lutjanus erythropterus Bloch. Journal of Fish Diseases, 36(8):695–702, August 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chen:2012:ICA

[CLZ^{+}12] Y. F. Chen, R. S. Liang, X. L. Zhuo, X. T. Wu, and J. X. Zou. Isolation and characterization of Aeromonas schubertii from diseased snakehead, Channa maculata (Lacépède). Journal of Fish Diseases, 35(6):421–430, June 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cervellione:2017:ESR

[CMEV17a] F. Cervellione, C. McGurk, T. Berger Eriksen, and W. Van den Broeck. Effect of starvation and refeeding on the hepatopancreas of whiteleg shrimp Penaeus vannamei (Boone) using computer-assisted image analysis. Journal of Fish Diseases, 40(11):1707–1715, November 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cervellione:2017:UCA

[CMEV17b] F. Cervellione, C. McGurk, T. Berger Eriksen, and W. Van den Broeck. Use of computer-assisted image analysis for semi-quantitative histology of the hepatopancreas in whiteleg shrimp Penaeus vannamei (Boone). Journal of Fish Diseases, 40(9):1223–1234, September 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chettri:2017:APC

[CMH^{+}17] J. K. Chettri, F. Mehrdana, E. B. Hansen, A. Ebbensgaard, M. T. Overgaard, A. H. Lauritsen, I. Dalsgaard, and K. Buchmann. Antimicrobial peptide CAP18 and its effect on Yersinia ruckeri infections in rainbow trout Oncorhynchus mykiss (Walbaum): comparing administration by injection and oral routes. Journal of Fish Diseases, 40 (1):97–104, January 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Contador:2016:ELT

[CMR^{+}16] E. Contador, P. Methner, I. Ryerse, P. Huber, B. N. Lillie, S. Frasca, Jr., and J. S. Lumsden. Epitheliocystis in lake trout Salvelinus namaycush (Walbaum) is associated with a β-proteobacteria. Journal of Fish Diseases, 39(3): 353–366, March 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cervellione:2017:OFM

[CMS^{+}17] F. Cervellione, C. McGurk, P. Silva, M. A. G. Owen, and W. Van den Broeck. Optimization of fixation methods for image analysis of the hepatopancreas in whiteleg shrimp, Penaeus vannamei (Boone). Journal of Fish Diseases, 40 (4):517–527, April 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chiamkunakorn:2019:BLB

[CMS^{+}19] Chutipong Chiamkunakorn, Vimbai Irene Machimbirike, Saengchan Senapin, Pongsak Khunrae, Ha Thanh Dong, and Triwit Rattanarojpong. Blood and liver biopsy for the non-destructive screening of tilapia lake virus. Journal of Fish Diseases, 42(11):1629–1636, November 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Carriero:2016:CNS

[CMSHS16] M. M. Carriero, A. A. Mendes Maia, R. L. Moro Sousa, and F. Henrique-Silva. Characterization of a new strain of Aeromonas dhakensis isolated from diseased pacu fish (Piaractus mesopotamicus) in Brazil. Journal of Fish Diseases, 39(11):1285–1295, November 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cavalli:2010:WSS

[CNN^{+}10] L. S. Cavalli, B. F. S. Nornberg, S. A. Netto, L. Poersch, L. A. Romano, L. F. Marins, and P. C. Abreu. White spot syndrome virus in wild penaeid shrimp caught in coastal and offshore waters in the southern Atlantic Ocean. Journal of Fish Diseases, 33(6):533–536, June 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Castro:2016:IVR

[COB^{+}16] N. Castro, C. R. Osorio, N. Buján, J. C. Fuentes, J. Rodríguez, M. Romero, C. Jiménez, A. E. Toranzo, and B. Magariños. Insights into the virulence-related genes of Edwardsiella tarda isolated from turbot in Europe: genetic homogeneity and evidence for vibrioferrin production. Journal of Fish Diseases, 39(5):565–576, May 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ciftci:2016:DVG

[ÇOÇ^{+}16] A. Çiftci, E. E. Onuk, G. Çiftci, A. Fındık, M. ü Söğüt, B. I. Didinen, A. Aksoy, K. üstünakın, T. Gülhan, F. Balta, and S. Altun. Development and validation of glycoproteinbased native-subunit vaccine for fish against Aeromonas hydrophila. Journal of Fish Diseases, 39(8):981–992, August 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Crosbie:2010:AGD

[CONN10] P. B. B. Crosbie, K. Ogawa, D. Nakano, and B. F. Nowak. Amoebic gill disease in hatchery-reared ayu, Plecoglossus altivelis (Temminck & Schlegel), in Japan is caused by Neoparamoeba perurans. Journal of Fish Diseases, 33(5):455–458, May 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Colussi:2015:ASM

[CPB^{+}15] S. Colussi, M. Prearo, S. A. Bertuzzi, T. Scanzio, S. Peletto, L. Favaro, P. Modesto, M. G. Maniaci, G. Ru, R. Desiato, and P. L. Acutis. Association of a specific major histocompatibility complex class IIβ single nucleotide polymorphism with resistance to lactococcosis in rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 38 (1):27–35, January 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Constenla:2016:DDD

[CPdPP16] M. Constenla, F. Padrós, R. del Pozo, and O. Palenzuela. Development of different diagnostic techniques for Endolimax piscium (archamoebae) and their applicability in Solea senegalensis clinical samples. Journal of Fish Diseases, 39(12):

1433–1443, December 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Conway:2015:DIC

[CPEH15] C. M. Conway, M. K. Purcell, D. G. Elliott, and P. K. Hershberger. Detection of Ichthyophonus by chromogenic in situ hybridization. Journal of Fish Diseases, 38(9):853–857, September 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Constenla:2014:EPS

[CPP14] M. Constenla, F. Padrós, and O. Palenzuela. Endolimax piscium sp. nov. (Amoebozoa), causative agent of systemic granulomatous disease of cultured sole, Solea senegalensis Kaup. Journal of Fish Diseases, 37(3):229–240, March 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chakrabarti:2012:AAE

[CR12] R. Chakrabarti and Y. V. Rao. Achyranthes aspera enhances immunity and antigen clearance in common carp, Cyprinus carpio L. Journal of Fish Diseases, 35(5):389–392, May 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cortez-SanMartin:2012:SDI

[CRAG^{+}12] M. Cortez-San Martín, A. Rivas-Aravena, S. Guajardo, M. T. Castillo, M. Jashes, A. M. Sandino, and E. Spencer. Simultaneous detection of the IPN and ISA viruses in outbreaks of clinical disease and mortality in farmed Atlantic salmon, Salmo salar L., in Chile. Journal of Fish Diseases, 35(6):461–465, June 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Castro:2017:HDT

[CRBLC^{+}17] R. Castro, M. Reguera-Brito, G. H. López-Campos, M. M. Blanco, M. Aguado-Urda, J. F. Fernández-Garayzábal, and A. Gibello. How does temperature influences the development of lactococcosis? Transcriptomic and immunoproteomic in vitro approaches. Journal of Fish Diseases, 40(10): 1285–1297, October 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Calquin:2018:PEP

[CRO^{+}18] P. Calquín, P. Ruiz, C. Oliver, P. Sánchez, R. Haro, H. Oliva, L. Vargas-Chacoff, R. Avendaño-Herrera, and A. J. Yáñez. Physiological evidence that Piscirickettsia salmonis produces siderophores and uses iron from different sources. Journal of Fish Diseases, 41(3):553–558, March 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Costas:2013:CHI

[CRS^{+}13] B. Costas, P. C. N. P. Rêgo, I. Simões, J. F. Marques, M. Castro-Cunha, and A. Afonso. Cellular and humoral immune responses of Senegalese sole, Solea senegalensis (Kaup), following challenge with two Photobacterium damselae subsp. piscicida strains from different geographical origins. Journal of Fish Diseases, 36(6):543–553, June 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chong:2015:FRE

[CSA^{+}15] R. S-M. Chong, M. W. Shinwari, M. J. Amigh, M. Aravena-Roman, and T. V. Riley. First report of Erysipelothrix rhusiopathiae- associated septicaemia and histologic changes in cultured Australian eels, Anguilla reinhardtii (Steindachner, 1867) and A. australis (Richardson, 1841). Journal of Fish Diseases, 38(9):839–847, September 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Costas:2014:NSI

[CSCCA14] B. Costas, I. Simões, M. Castro-Cunha, and A. Afonso. Nonspecific immune responses of Senegalese sole, Solea senegalensis (Kaup), head-kidney leucocytes against Tenacibaculum maritimum. Journal of Fish Diseases, 37(8):765–769, August 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cafiso:2016:MEB

[CSS^{+}16] A. Cafiso, D. Sassera, V. Serra, C. Bandi, U. McCarthy, and C. Bazzocchi. Molecular evidence for a bacterium of the family Midichloriaceae (order Rickettsiales) in skin and organs of the rainbow trout Oncorhynchus mykiss (Walbaum) affected by red mark syndrome. Journal of Fish Diseases, 39 (4):497–501, April 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cui:2010:AMM

[CSSWK10] Z. Cui, D. Samuel-Shaker, V. Watral, and M. L. Kent. Attenuated Mycobacterium marinum protects zebrafish against mycobacteriosis. Journal of Fish Diseases, 33(4):371–375, April 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Castro:2012:FDE

[CTD^{+}12] N. Castro, A. E. Toranzo, S. Devesa, A. González, S. Nuñez, and B. Magariños. First description of Edwardsiella tarda in Senegalese sole, Solea senegalensis (Kaup). Journal of Fish Diseases, 35(1):79–82, January 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Crumlish:2010:ECS

[CTK^{+}10] M. Crumlish, P. C. Thanh, J. Koesling, V. T. Tung, and K. Gravningen. Experimental challenge studies in Vietnamese catfish, Pangasianodon hypophthalmus (Sauvage), exposed to Edwardsiella ictaluri and Aeromonas hydrophila. Journal of Fish Diseases, 33(9):717–722, September 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cano:2013:TLD

[CVGR^{+}13] I. Cano, E. J. Valverde, E. Garcia-Rosado, M. C. Alonso, B. Lopez-Jimena, J. B. Ortiz-Delgado, J. J. Borrego, C. Sarasquete, and D. Castro. Transmission of lymphocystis disease virus to cultured gilthead seabream, Sparus aurata L., larvae. Journal of Fish Diseases, 36(6):569–576, June 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cano:2010:NGL

[CVLJ^{+}10] I. Cano, E. J. Valverde, B. Lopez-Jimena, M. C. Alonso, E. Garcia-Rosado, C. Sarasquete, J. J. Borrego, and D. Castro. A new genotype of Lymphocystivirus isolated from cultured gilthead seabream, Sparus aurata L., and Senegalese sole, Solea senegalensis (Kaup). Journal of Fish Diseases, 33(8):695–700, August 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chong:2017:SAO

[CWL^{+}17] S. M. Chong, W. K. Wong, W. Y. Lee, Z. B. Tan, Y. H. Tay, X. H. Teo, L. D. Chee, and C. J. Fernandez. Streptococcus agalactiae outbreaks in cultured golden pomfret, Trachinotus blochii (Lacépède), in Singapore. Journal of Fish Diseases, 40(7):971–974, July 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chen:2018:FDY

[CWW^{+}18] Jiayuan Chen, Wenchao Wang, Xiuhua Wang, Qingli Zhang, Yanbei Ren, Jipeng Song, Xuepeng Wang, Xuan Dong, and Jie Huang. First detection of yellow head virus genotype 3 (YHV-3) in cultured Penaeus monodon, mainland China. Journal of Fish Diseases, 41(9):1449–1451, September 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chow:2016:RSP

[CXK16] F. W. Chow, L. Xue, and M. L. Kent. Retrospective study of the prevalence of Pseudoloma neurophilia shows male sex bias in zebrafish Danio rerio (Hamilton-Buchanan). Journal of Fish Diseases, 39(3):367–370, March 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chen:2019:IMT

[CXW^{+}19] Jianlin Chen, Liqun Xia, Wenji Wang, Zhiwen Wang, Suying Hou, Caixia Xie, Jia Cai, and Yishan Lu. Identification of a mitochondrial-targeting secretory protein from Nocardia seriolae which induces apoptosis in fathead minnow cells. Journal of Fish Diseases, 42(11):1493–1507, November 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chung:2017:ARR

[CYS17] T-H. Chung, S-W. Yi, and G-W. Shin. Antibiotic resistance and repetitive-element PCR fingerprinting in Aeromonas veronii isolates. Journal of Fish Diseases, 40(6):821–829, June 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Corriero:2011:ESA

[CZB^{+}11] A. Corriero, R. Zupa, G. Bello, C. C. Mylonas, M. Deflorio, S. Genovese, G. Basilone, G. Buscaino, G. Buffa, C. Pousis, G. De Metrio, and N. Santamaria. Evidence that severe acute stress and starvation induce rapid atresia of ovarian vitellogenic follicles in Atlantic bluefin tuna, Thunnus thynnus (L.) (Osteichthyes: Scombridae). Journal of Fish Diseases, 34(11):853–860, November 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cui:2019:DPD

[CZL^{+}19] Miao Cui, Huijie Zhang, Jingjing Li, Ru Liu, Min Wu, Delin Xu, and Qizhong Zhang. Differential PCR detection of Streptococcus agalactiae and Streptococcus iniae in tilapia (Oreochromis niloticus) by a single primer pair. Journal of Fish Diseases, 42(8):1211–1216, August 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dixon:2012:SIS

[DAB^{+}12] P. F. Dixon, M. Algoët, A. Bayley, M. Dodge, C. Joiner, and E. Roberts. Studies on the inactivation of selected viral and bacterial fish pathogens at high pH for waste disposal purposes. Journal of Fish Diseases, 35(1):65–72, January 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dubert:2017:MEI

[DABO^{+}17] J. Dubert, J. A. Aranda-Burgos, J. Ojea, J. L. Barja, and S. Prado. Mortality event involving larvae of the carpet shell clam Ruditapes decussatus in a hatchery: isolation of the pathogen Vibrio tubiashii subsp. europaeus. Journal of Fish Diseases, 40(9):1185–1193, September 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dalmo:2018:DVF

[Dal18] R. A. Dalmo. DNA vaccines for fish: Review and perspectives on correlates of protection. Journal of Fish Diseases, 41(1):1–9, January 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

deAzevedo:2017:NIV

[dALF^{+}17] A. M. de Azevedo, A. P. Losada, I. Ferreiro, A. Riaza, S. Vázquez, and M. I. Quiroga. New insight on vertebral anomalies in cultured Senegalese sole (Solea senegalensis, Kaup) at early stages of development. Journal of Fish Diseases, 40(8):987–1000, August 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dumbo:2019:HCI

[DAO19] José Chissiua Dumbo and Annemariè Avenant-Oldewage. Histopathological changes induced by the digenean intestinal parasite Masenia nkomatiensis Dumbo, Dos Santos, & Avenant-Oldewage, 2019 of the catfish Clarias gariepinus (Burchell) from Incomati Basin, Mozambique. Journal of Fish Diseases, 42(10):1341–1350, October 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sebastiao:2019:FII

[dASLS^{+}19] Fernanda de Alexandre Sebastião, Benjamin R. LaFrentz, John P. Shelley, Brittany Stevens, David Marancik, Freeland Dunker, Drury Reavill, and Esteban Soto. Flavobacterium inkyongense isolated from ornamental cichlids. Journal of Fish Diseases, 42(9):1309–1313, September 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dill:2017:LLD

[DBHC17] J. Dill, C. Brown, K. Heym, and A. Camus. Lipoid liver disease, atherosclerosis and glomerular lipidosis in a Gulf flounder Paralichthys albigutta (Jordan & Gilbert 1882): a case report. Journal of Fish Diseases, 40(2):273–278, February 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dutta:2014:IRS

[DBM^{+}14] S. Dutta, S. Biswas, K. Mukherjee, U. Chakrabarty, A. Mallik, and N. Mandal. Identification of RAPD-SCAR marker linked to white spot syndrome virus resistance in populations of giant black tiger shrimp, Penaeus monodon Fabricius. Journal of Fish Diseases, 37(5):471–480, May 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Darwish:2012:AAC

[DBS12] A. M. Darwish, J. A. Bebak, and K. K. Schrader. Assessment of Aquaflor(R), copper sulphate and potassium permanganate for control of Aeromonas hydrophila and Flavobacterium columnare infection in sunshine bass, Morone chrysops female × Morone saxatilis male. Journal of Fish Diseases, 35(9):637–647, September 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Declercq:2013:ASP

[DBV^{+}13] A. M. Declercq, F. Boyen, W. Van den Broeck, P. Bossier, A. Karsi, F. Haesebrouck, and A. Decostere. Antimicrobial susceptibility pattern of Flavobacterium columnare isolates collected worldwide from 17 fish species. Journal of Fish Diseases, 36(1):45–55, January 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dezfuli:2018:HUA

[DCG18] B. Sayyaf Dezfuli, G. Castaldelli, and L. Giari. Histopathological and ultrastructural assessment of two mugilid species infected with myxozoans and helminths. Journal of Fish Diseases, 41(2):299–307, February 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ding:2017:DSH

[DCL^{+}17] Z. F. Ding, J. Q. Chen, J. Lin, X. S. Zhu, G. H. Xu, R. L. Wang, Q. G. Meng, and W. Wang. Development of in situ hybridization and real-time PCR assays for the detection of Hepatospora eriocheir, a microsporidian pathogen in the Chinese mitten crab Eriocheir sinensis. Journal of Fish Diseases, 40(7):919–927, July 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dutta:2013:EEW

[DCMM13] S. Dutta, U. Chakrabarty, A. Mallik, and N. Mandal. Experimental evidence for white spot syndrome virus (WSSV) susceptibility linked to a microsatellite DNA marker in giant black tiger shrimp, Penaeus monodon (Fabricius). Journal of Fish Diseases, 36(6):593–597, June 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Declercq:2014:WNT

[DCV^{+}14] A. M. Declercq, K. Chiers, W. Van den Broeck, A. Rekecki, S. Teerlinck, D. Adriaens, F. Haesebrouck, and A. Decostere. White necrotic tail tips in estuary seahorses, Hippocampus kuda, Bleeker. Journal of Fish Diseases, 37(5):501–504, May 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ding:2017:CGM

[DCZ^{+}17] Z. F. Ding, M. J. Cao, X. S. Zhu, G. H. Xu, and R. L. Wang. Changes in the gut microbiome of the Chinese mitten crab (Eriocheir sinensis) in response to White spot syndrome virus (WSSV) infection. Journal of Fish Diseases, 40 (11):1561–1571, November 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dennis:2015:MOF

[DD15] M. M. Dennis and B. K. Diggles. Multicentric orocutaneous fibrosarcoma in a fork-tailed catfish (Hemiarius dioctes). Journal of Fish Diseases, 38(12):1077–1081, December 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

DeBodt:2018:IOL

[DD18] Jana De Bodt and Tom Defoirdt. Impact of the organic load on the efficacy of chlorine disinfection against acute hepatopancreatic necrosis disease-causing Vibrio parahaemolyticus. Journal of Fish Diseases, 41(10):1609–1612, October 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Delamare-Deboutteville:2015:IPQ

[DDBC^{+}15] J. Delamare-Deboutteville, R. Bowater, K. Condon, A. Reynolds, A. Fisk, F. Aviles, and A. C. Barnes. Infection and pathology in Queensland grouper, Epinephelus lanceolatus, (Bloch), caused by exposure to Streptococcus agalactiae via different routes. Journal of Fish Diseases, 38(12):1021–1035, December 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

DiSalvo:2013:PRD

[DdRCP13] A. Di Salvo, G. della Rocca, P. Cagnardi, and R. M. Pellegrino. Pharmacokinetics and residue depletion of erythromycin in rainbow trout Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 36(12):1021–1029, December 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

DiSalvo:2013:FDE

[DdRTM13] A. Di Salvo, G. della Rocca, E. Terzetti, and J. Malvisi. Florfenicol depletion in edible tissue of rainbow trout, Oncorhynchus mykiss (Walbaum), and sea bream, Sparus aurata L. Journal of Fish Diseases, 36(8):685–693, August 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dreyfuss:2014:ZTB

[DGS^{+}14] J. Dreyfuss, J. Geyer, M. A. Stamper, A. Baldessari, and G. A. Lewbart. Zinc toxicosis in a brook trout, Salvelinus fontinalis Mitchill. Journal of Fish Diseases, 37(4):397–399, April 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Diggles:2013:SDY

[Dig13]  B. K. Diggles.  Saddleback deformities in yellowfin

bream, Acanthopagrus australis (Günther), from South East Queensland. Journal of Fish Diseases, 36(5):521–527, May 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Diggles:2015:RPC

[Dig15] B. K. Diggles. Reply to pollock (2014) Comments on ‘Saddleback deformities in yellowfin bream, Acanthopagrus australis (Günther), from South East Queensland’ by Diggles (2013). Journal of Fish Diseases, 38(3):331–333, March 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Densmore:2016:MIN

[DIH^{+}16] C. L. Densmore, L. R. Iwanowicz, A. P. Henderson, D. D. Iwanowicz, and J. S. Odenkirk. Mycobacterial infection in Northern snakehead (Channa argus) from the Potomac River catchment. Journal of Fish Diseases, 39(6):771–775, June 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Diaz:2016:LTE

[DIVC16] S. Díaz, D. Iglesias, A. Villalba, and M. J. Carballal. Longterm epidemiological study of disseminated neoplasia of cockles in Galicia (NW Spain): temporal patterns at individual and population levels, influence of environmental and cockle-based factors and lethality. Journal of Fish Diseases, 39(9):1027–1042, September 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Duran-Avelar:2018:GPG

[dJDAVRGM^{+}18] Ma. de Jesús Durán-Avelar, Alejandro Vázquez-Reyes, Ana Lourdes González-Mercado, José Francisco Zambrano-Zaragoza, Miriam Fabiola Ayón-Pérez, Juan Manuel Agraz-Cibrián, Jorge Gutiérrez-Franco, and Norberto Vibanco-Pérez. pirA- and pirB-like gene identification in Micrococcus luteus strains in Mexico. Journal of Fish Diseases, 41 (11):1667–1673, November 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Deane:2012:IAS

[DJQ^{+}12] E. E. Deane, A. Jia, Z. Qu, J-X. Chen, X-H. Zhang, and N. Y. S. Woo. Induction of apoptosis in sea bream fibroblasts by Vibrio harveyi haemolysin and evidence for an antiapoptotic role of heat shock protein 70. Journal of Fish Diseases, 35(4):287–302, April 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

DiNatale:2019:RTO

[DJR^{+}19] Adrian Di Natale, Maureen Jarau, Spencer K. Russell, Ryan A. Horricks, Brandon N. Lillie, Brian Dixon, and John S. Lumsden. Rainbow trout Oncorhynchus mykiss (Walbaum) type IV ice-structuring protein LS-12 in the acute-phase response to Flavobacterium psychrophilum infection. Journal of Fish Diseases, 42(7):975–984, July 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Doszpoly:2017:ICA

[DKBS17] A. Doszpoly, I. M. Kalabekov, R. Breyta, and I. S. Shchelkunov. Isolation and characterization of an atypical Siberian sturgeon herpesvirus strain in Russia: novel North American Acipenserid herpesvirus 2 strain in Europe? Journal of Fish Diseases, 40(10):1363–1372, October 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dalum:2017:AVA

[DKG^{+}17] A. S. Dalum, K. H. Kristthorsdottir, D. J. Griffiths, K. Bjørklund, and T. T. Poppe. Arteriosclerosis in the ventral aorta and epicarditis in the bulbus arteriosus of Atlantic salmon (Salmo salar L.). Journal of Fish Diseases, 40(6): 797–809, June 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dantas-Lima:2013:PWS

[DLCC^{+}13] J. J. Dantas-Lima, M. Corteel, M. Cornelissen, P. Bossier, P. Sorgeloos, and H. J. Nauwynck. Purification of white spot syndrome virus by iodixanol density gradient centrifugation. Journal of Fish Diseases, 36(10):841–851, October 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Diamanka:2013:PCA

[DLCF13] A. Diamanka, T. P. Loch, R. C. Cipriano, and M. Faisal. Polyphasic characterization of Aeromonas salmonicida isolates recovered from salmonid and non-salmonid fish. Journal of Fish Diseases, 36(11):949–963, November 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Delfosse:2018:SHB

[DLLB^{+}18] C. Delfosse, C. Lafont-Lecuelle, H. Barthélémy, C. Chabaud, E. Teruel, C. Bienboire-Frosini, and P. Pageat. Study of the hooking behaviour of Lepeophtheirus salmonis (Krøyer, 1837) copepodids on Atlantic salmon, Salmo salar L., using a novel in vivo test system. Journal of Fish Diseases, 41 (6):969–972, June 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dezfuli:2013:CTS

[DLP^{+}13] B. S. Dezfuli, A. Lui, F. Pironi, M. Manera, A. P. Shinn, and M. Lorenzoni. Cell types and structures involved in tench, Tinca tinca (L.), defence mechanisms against a systemic digenean infection. Journal of Fish Diseases, 36(6):577–585, June 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dong:2015:PCG

[DLPR15] H. T. Dong, B. LaFrentz, N. Pirarat, and C. Rodkhum. Phenotypic characterization and genetic diversity of Flavobacterium columnare isolated from red tilapia, Oreochromis sp., in Thailand. Journal of Fish Diseases, 38(10):901–913, October 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dezfuli:2016:AAI

[DMB^{+}16] B. S. Dezfuli, M. Manera, G. Bosi, J. A. DePasquale, S. D’Amelio, G. Castaldelli, and L. Giari. Anguilla anguilla intestinal immune response to natural infection with Contracaecum rudolphii A larvae. Journal of Fish Diseases, 39(10): 1187–1200, October 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dessen:2019:IDP

[DMB^{+}19] Jens-Erik Dessen, Turid Mørkøre, Jan Ivar Bildøy, Stig N. Johnsen, Lars Thomas Poppe, Bjarne Hatlen, Magny S. Thomassen, and Kjell-Arne Rørvik. Increased dietary protein-to-lipid ratio improves survival during naturally occurring pancreas disease in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 42(1):21–34, January 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dezfuli:2017:LFG

[DMD^{+}17] B. Sayyaf Dezfuli, M. Manera, J. A. DePasquale, F. Pironi, and L. Giari. Liver of the fish Gymnotus inaequilabiatus and nematode larvae infection: Histochemical features and expression of proliferative cell nuclear antigen. Journal of Fish Diseases, 40(12):1765–1774, December 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dubey:2019:GPC

[DMK^{+}19] Saurabh Dubey, Biswajit Maiti, Sung-Hyun Kim, Sangeetha Madambithara Sivadasan, Dhamotharan Kannimuthu, Pramod Kumar Pandey, Shivani Kallappa Girisha, Stephen Mutoloki, Shih-Chu Chen, Øystein Evensen, Indrani Karunasagar, and Hetron Mweemba Munang’andu. Genotypic and phenotypic characterization of Edwardsiella isolates from different fish species and geographical areas in Asia: Implications for vaccine development. Journal of Fish Diseases, 42(6):835–850, June 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ding:2016:FCH

[DML^{+}16] Z. Ding, Q. Meng, H. Liu, S. Yuan, F. Zhang, M. Sun, Y. Zhao, M. Shen, G. Zhou, J. Pan, H. Xue, and W. Wang. First case of hepatopancreatic necrosis disease in pondreared Chinese mitten crab, Eriocheir sinensis, associated with microsporidian. Journal of Fish Diseases, 39(9):1043– 1051, September 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

DelCerro:2010:GDA

[DMP10] A. Del Cerro, I. Márquez, and J. M. Prieto. Genetic diversity and antimicrobial resistance of Flavobacterium psychrophilum isolated from cultured rainbow trout, Onchorynchus mykiss (Walbaum), in Spain. Journal of Fish Diseases, 33(4):285–291, April 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

deOcenda:2017:PMF

[dOAPLá^{+}17] V-R. de Ocenda, S. Almeida-Prieto, A. Luzardo-álvarez, J. L. Barja, F. J. Otero-Espinar, and J. Blanco-Méndez. Pharmacokinetic model of florfenicol in turbot (Scophthalmus maximus): establishment of optimal dosage and administration in medicated feed. Journal of Fish Diseases, 40(3): 411–424, March 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Doherty:2016:FVM

[DPADO16] L. Doherty, S. J. Poynter, A. Aloufi, and S. J. DeWitte-Orr. Fish viruses make dsRNA in fish cells: characterization of dsRNA production in rainbow trout (Oncorhynchus mykiss) cells infected with viral haemorrhagic septicaemia virus, chum salmon reovirus and frog virus 3. Journal of Fish Diseases, 39(9):1133–1137, September 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dezfuli:2010:RIM

[DPC^{+}10] B. S. Dezfuli, F. Pironi, M. Campisi, A. P. Shinn, and L. Giari. The response of intestinal mucous cells to the presence of enteric helminths: their distribution, histochemistry and fine structure. Journal of Fish Diseases, 33(6):481–488, June 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

DiSalvo:2014:PRD

[DPCdR14] A. Di Salvo, R. M. Pellegrino, P. Cagnardi, and G. della Rocca. Pharmacokinetics and residue depletion of erythromycin in gilthead sea bream Sparus aurata L. after oral administration. Journal of Fish Diseases, 37(9):797–803, September 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

DeClercq:2018:EPS

[DPD^{+}18] A. De Clercq, M. R. Perrott, P. S. Davie, M. A. Preece, A. Huysseune, and P. E. Witten. The external phenotype– skeleton link in post-hatch farmed Chinook salmon (Oncorhynchus tshawytscha). Journal of Fish Diseases, 41(3): 511–527, March 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dang:2012:DNP

[DPMK12] H. T. Dang, H. K. Park, S. C. Myung, and W. Kim. Development of a novel PCR assay based on the 16S–23S rRNA internal transcribed spacer region for the detection of Lactococcus garvieae. Journal of Fish Diseases, 35(7):481–487, July 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Del-Pozo:2010:SGH

[DPTCF10] J. Del-Pozo, J. F. Turnbull, M. Crumlish, and H. W. Ferguson. A study of gross, histological and blood biochemical changes in rainbow trout, Oncorhynchus mykiss (Walbaum), with rainbow trout gastroenteritis (RTGE). Journal of Fish Diseases, 33(4):301–310, April 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Del-Pozo:2010:CMS

[DPTFC10] J. Del-Pozo, J. Turnbull, H. Ferguson, and M. Crumlish. A comparative molecular study of the presence of “ Candidatus arthromitus” in the digestive system of rainbow trout, Oncorhynchus mykiss (Walbaum), healthy and affected with rainbow trout gastroenteritis. Journal of Fish Diseases, 33 (3):241–250, March 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Duran:2010:ITC

[DRdS^{+}10] A. C. Durán, C. Rodríguez, L. Gil de Sola, B. Fernández, J. M. Arqué, and V. Sans-Coma. Intimal thickening of coronary arteries in the rabbitfish, Chimaera monstrosa L. (Chondrichthyes: Holocephali). Journal of Fish Diseases, 33 (8):675–682, August 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Diaz:2017:AIA

[DRG^{+}17] S. Díaz, M. E. Rojas, M. Galleguillos, C. Maturana, P. I. Smith, F. Cifuentes, I. Contreras, and P. A. Smith. Apoptosis inhibition of Atlantic salmon (Salmo salar) peritoneal macrophages by Piscirickettsia salmonis. Journal of Fish Diseases, 40(12):1895–1902, December 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dahiya:2010:ZOI

[DS10] I. Dahiya and R. M. W. Stevenson. The ZnuABC operon is important for Yersinia ruckeri infections of rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 33(4):331–340, April 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dixon:2012:SET

[DSA^{+}12] P. F. Dixon, D. A. Smail, M. Algoët, T. S. Hastings, A. Bayley, H. Byrne, M. Dodge, A. Garden, C. Joiner, E. Roberts, D. Verner-Jeffreys, and F. Thompson. Studies on the effect of temperature and pH on the inactivation of fish viral and bacterial pathogens. Journal of Fish Diseases, 35(1):51–64, January 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Costa:2016:MNN

[dSCBdS^{+}16] A. C. da S. Costa, H. M. Brandão, S. R. da Silva, A. R. Bentes-Sousa, J. A. P. Diniz, Jr., J. de J. Viana Pinheiro, M. de F. C. de Melo, J. O. C. Silva, Jr., E. R. Matos, and R. M. Ribeiro-Costa. Mucoadhesive nanoparticles: a new perspective for fish drug application. Journal of Fish Diseases, 39(4):503–506, April 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

daSilva:2019:GCT

[dSdAdS^{+}19] Marcelo F. da Silva, Rebeca S. de Araújo, Diehgo T. da Silva, Osimar Sanches, Igor Hamoy, and Edilson Matos. Gastric coccidiosis in Thoracocharax stellatus caused by Goussia guamaensis n. sp. (Apicomplexa: Eimeriidae) from the Amazon region. Journal of Fish Diseases, 42(6):905–912, June 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Duman:2018:DIS

[DSJA18] Muhammed Duman, Izzet Burcin Saticioglu, J. Michael Janda, and Soner Altun. The determination of the infectious status and prevalence of motile Aeromonas species isolated from disease cases in rainbow trout (Oncorhynchus mykiss) and aquarium fish. Journal of Fish Diseases, 41(12):1843– 1857, December 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dong:2016:VAR

[DSLR16] H. T. Dong, S. Senapin, B. LaFrentz, and C. Rodkhum. Virulence assay of rhizoid and non-rhizoid morphotypes of Flavobacterium columnare in red tilapia, Oreochromis sp., fry. Journal of Fish Diseases, 39(6):649–655, June 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dill:2017:DNA

[DSMC17] J. Dill, S. Sanchez, A. McDermott, and A. Camus. Disseminated nocardiosis associated with the isolation of Nocardia nova in a longsnout seahorse Hippocampus reidi (Ginsburg). Journal of Fish Diseases, 40(9):1235–1239, September 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

deSales-Ribeiro:2019:FDS

[dSRSHGá^{+}19] Carolina de Sales-Ribeiro, Andres Sanchez-Henao, Natalia García-álvarez, Fernando Real, Miguel A. Rivero, Antonio Fernández, and María José Caballero. First description of spontaneous granulomatous aerocystitis by Phoma herbarum in a wild greater amberjack (Seriola dumerili Risso, 1810). Journal of Fish Diseases, 42(9):1321–1325, September 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Delshad:2018:EQQ

[DSSB18] Somayeh Torabi Delshad, Siyavash Soltanian, Hassan Sharifiyazdi, and Peter Bossier. Effect of quorum quenching bacteria on growth, virulence factors and biofilm formation of Yersinia ruckeri in vitro and an in vivo evaluation of their probiotic effect in rainbow trout. Journal of Fish Diseases, 41(9):1429–1438, September 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Delshad:2019:ECS

[DSSB19] Somayeh Torabi Delshad, Siyavash Soltanian, Hassan Sharifiyazdi, and Peter Bossier. Effect of catecholamine stress hormones (dopamine and norepinephrine) on growth, swimming motility, biofilm formation and virulence factors of Yersinia ruckeri in vitro and an in vivo evaluation in rainbow trout. Journal of Fish Diseases, 42(4):477–487, April 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dykova:2016:TAR

[DT16] I. Dyková and T. Tyml. Testate amoeba Rhogostoma minus Belar, 1921, associated with nodular gill disease of rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 39(5):539–546, May 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dalum:2016:CCA

[DTF^{+}16] A. Dalum, R. Tangen, K. Falk, I. Hordvik, G. Rosenlund, B. Torstensen, and E. O. Koppang. Coronary changes in the Atlantic salmon Salmo salar L.: characterization and impact of dietary fatty acid compositions. Journal of Fish Diseases, 39(1):41–54, January 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dong:2017:AJA

[DTJ^{+}17] H. T. Dong, C. Techatanakitarnan, P. Jindakittikul, A. Thaiprayoon, S. Taengphu, W. Charoensapsri, P. Khunrae, T. Rattanarojpong, and S. Senapin. Aeromonas jandaei and Aeromonas veronii caused disease and mortality in Nile tilapia, Oreochromis niloticus (L.). Journal of Fish Diseases, 40(10):1395–1403, October 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

David:2010:MDB

[DTM^{+}10] R. David, C. Tréguier, C. Montagnani, C. Belliard, P. Levy, G. Nédélec, V. Joufoques, G. Remoissenet, Y. Gueguen, and N. Cochennec-Laureau. Molecular detection of betanodavirus from the farmed fish, Platax orbicularis (Forsskål) (Ephippidae), in French Polynesia. Journal of Fish Diseases, 33(5):451–454, May 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Doan:2017:VER

[DVC^{+}17] Q. K. Doan, M. Vandeputte, B. Chatain, T. Morin, and F. Allal. Viral encephalopathy and retinopathy in aquaculture: a review. Journal of Fish Diseases, 40(5):717–742, May 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dill:2017:IEP

[DWL^{+}17] J. A. Dill, S. M. Williams, J. H. Leary, T. F. F. Ng, and A. C. Camus. Investigation of epizootic papillomatosis in bluegill Lepomis macrochirus (Rafinesque 1810) using nextgeneration sequencing. Journal of Fish Diseases, 40(7):947– 952, July 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Davie:2019:VFF

[DWP^{+}19] Peter S. Davie, Seamus P. Walker, Matthew R. Perrott, Jane E. Symonds, Mark Preece, Bailey A. Lovett, and John S. Munday. Vertebral fusions in farmed Chinook salmon (Oncorhynchus tshawytscha) in New Zealand. Journal of Fish Diseases, 42(7):965–974, July 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ding:2015:DVN

[DXX^{+}15] Z. F. Ding, S. Y. Xia, H. Xue, J. Q. Tang, Q. Ren, W. Gu, Q. G. Meng, and W. Wang. Direct visualization of the novel pathogen, Spiroplasma eriocheiris, in the freshwater crayfish Procambarus clarkii (Girard) using fluorescence in situ hybridization. Journal of Fish Diseases, 38(9):787–794, September 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Downes:2018:ICI

[DYML^{+}18] J. K. Downes, T. Yatabe, M. Marcos-Lopez, H. D. Rodger, E. MacCarthy, I. O’Connor, E. Collins, and N. M. Ruane. Investigation of co-infections with pathogens associated with gill disease in Atlantic salmon during an amoebic gill disease outbreak. Journal of Fish Diseases, 41(8):1217–1227, August 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Delannoy:2016:GCV

[DZC^{+}16] C. M. J. Delannoy, R. N. Zadoks, M. Crumlish, D. Rodgers, F. A. Lainson, H. W. Ferguson, J. Turnbull, and M. C. Fontaine. Genomic comparison of virulent and non-virulent Streptococcus agalactiae in fish. Journal of Fish Diseases, 39 (1):13–29, January 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dieu:2010:CVU

[DZV10] B. T. M. Dieu, M. P. Zwart, and J. M. Vlak. Can VN-TRs be used to study genetic variation within white spot syndrome virus isolates? Journal of Fish Diseases, 33(8): 689–693, August 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Esteve:2018:SRE

[EA18] Consuelo Esteve and Elena Alcaide. Seasonal recovery of Edwardsiella piscicida from wild European eels and natural waters: Isolation methods, virulence and reservoirs. Journal of Fish Diseases, 41(11):1613–1623, November 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Elliott:2013:BTV

[EAM^{+}13] D. G. Elliott, L. J. Applegate, A. L. Murray, M. K. Purcell, and C. L. McKibben. Bench-top validation testing of selected immunological and molecular Renibacterium salmoninarum diagnostic assays by comparison with quantitative bacteriological culture. Journal of Fish Diseases, 36(9):779–809, September 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Edrees:2018:CET

[EAN^{+}18] A. Edrees, H. Abdelhamed, S. W. Nho, S. B. Park, A. Karsi, F. W. Austin, M. Essa, T. Pechan, and M. L. Lawrence. Construction and evaluation of type III secretion system mutants of the catfish pathogen Edwardsiella piscicida. Journal of Fish Diseases, 41(5):805–816, May 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Emmenegger:2018:VCR

[EBM^{+}18] E. J. Emmenegger, S. Biacchesi, E. Mérour, J. A. Glenn, A. D. Palmer, M. Brémont, and G. Kurath. Virulence of a chimeric recombinant infectious haematopoietic necrosis virus expressing the spring viraemia of carp virus glycoprotein in salmonid and cyprinid fish. Journal of Fish Diseases, 41(1):67–78, January 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ellis:2010:HIQ

[ECP^{+}10] A. E. Ellis, A. Cavaco, A. Petrie, K. Lockhart, M. Snow, and B. Collet. Histology, immunocytochemistry and qRT-PCR analysis of Atlantic salmon, Salmo salar L., post-smolts following infection with infectious pancreatic necrosis virus (IPNV). Journal of Fish Diseases, 33(10):803–818, October 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Eichner:2018:RSR

[EDN18] C. Eichner, M. Dondrup, and F. Nilsen. RNA sequencing reveals distinct gene expression patterns during the development of parasitic larval stages of the salmon louse (Lepeophtheirus salmonis). Journal of Fish Diseases, 41(6):1005– 1029, June 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Easwaran:2017:CBP

[EDP^{+}17] M. Easwaran, S. H. S. Dananjaya, S. C. Park, J. Lee, H-J. Shin, and M. De Zoysa. Characterization of bacteriophage pAh-1 and its protective effects on experimental infection of Aeromonas hydrophila in Zebrafish (Danio rerio). Journal of Fish Diseases, 40(6):841–846, June 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Evely:2011:OMR

[EDSL11] M. M. Evely, J. M. Donahue, S. F. Sells, and A. T. Loynachan. Ocular mycobacteriosis in a red-bellied piranha, Pygocentrus nattereri Kner. Journal of Fish Diseases, 34 (4):323–326, April 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Erkinharju:2018:VAL

[EDV^{+}18] T. Erkinharju, R. A. Dalmo, Ø. Vågsnes, I. Hordvik, and T. Seternes. Vaccination of Atlantic lumpfish (Cyclopterus lumpus L.) at a low temperature leads to a low antibody response against Aeromonas salmonicida. Journal of Fish Diseases, 41(4):613–623, April 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Enriquez-Espinoza:2016:MES

[EEMCEGSP16] T. Enríquez-Espinoza, F. Mendoza-Cano, T. Encinas-García, and A. Sánchez-Paz. Molecular epidemiology of selected infectious diseases caused by bacteria in juveniles and post-larvae of the white shrimp Penaeus vannamei from the north-western coast of Mexico. Journal of Fish Diseases, 39 (11):1393–1397, November 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Escobar:2017:PDV

[EKED^{+}17] L. E. Escobar, G. Kurath, J. Escobar-Dodero, M. E. Craft, and N. B. D. Phelps. Potential distribution of the viral haemorrhagic septicaemia virus in the Great Lakes region. Journal of Fish Diseases, 40(1):11–28, January 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Elkesh:2013:SNM

[EKS^{+}13] A. Elkesh, K. P. L. Kantham, A. P. Shinn, M. Crumlish, and R. H. Richards. Systemic nocardiosis in a Mediterranean population of cultured meagre, Argyrosomus regius Asso (Perciformes: Sciaenidae). Journal of Fish Diseases, 36 (2):141–149, February 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Emmenegger:2011:DAP

[EKT^{+}11] E. J. Emmenegger, E. Kentop, T. M. Thompson, S. Pittam, A. Ryan, D. Keon, J. A. Carlino, J. Ranson, R. B. Life, R. M. Troyer, K. A. Garver, and G. Kurath. Development of an aquatic pathogen database (AquaPathogen X) and its utilization in tracking emerging fish virus pathogens in North America. Journal of Fish Diseases, 34(8):579–587, August 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Erkinharju:2017:SAR

[ELI^{+}17] T. Erkinharju, M. R. Lundberg, E. Isdal, I. Hordvik, R. A. Dalmo, and T. Seternes. Studies on the antibody response and side effects after intramuscular and intraperitoneal injection of Atlantic lumpfish (Cyclopterus lumpus L.) with different oil-based vaccines. Journal of Fish Diseases, 40 (12):1805–1813, December 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Esteve:2017:OSP

[EMA17] C. Esteve, R. Merchán, and E. Alcaide. An outbreak of Shewanella putrefaciens group in wild eels Anguilla anguilla L. favoured by hypoxic aquatic environments. Journal of Fish Diseases, 40(7):929–939, July 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

ElAamri:2010:FRS

[EPA^{+}10] F. El Aamri, D. Padilla, F. Acosta, M. J. Caballero, J. Roo, J. Bravo, J. Vivas, and F. Real. First report of Streptococcus iniae in red porgy (Pagrus pagrus, L.). Journal of Fish Diseases, 33(11):901–905, November 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Estensoro:2018:BDL

[EPCSBP18] I. Estensoro, G. Pérez-Cordón, A. Sitjà-Bobadilla, and M. C. Piazzon. Bromodeoxyuridine DNA labelling reveals host and parasite proliferation in a fish-myxozoan model. Journal of Fish Diseases, 41(4):651–662, April 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Estensoro:2014:ICP

[ERÁpSB14] I. Estensoro, M. J. Redondo, P. Álvarez-pellitero, and A. Sitjà-Bobadilla. Immunohistochemical characterization of polyclonal antibodies against Enteromyxum leei and Enteromyxum scophthalmi (Myxozoa: Myxosporea), intestinal parasites of fish. Journal of Fish Diseases, 37(9):785–796, September 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

ElAsely:2010:IEP

[ESA^{+}10] Amel M. El Asely, A. A. Shaheen, Amany A. Abbass, R. Sudhakaran, N. T. H. Linh, T. Yoshida, Y. Tachikawa, S. Yoshida, and T. Itami. Immunomodulatory effect of plantmixed feed in kuruma shrimp, Marsupenaeus japonicus, and its protective efficacy against white spot syndrome virus infection. Journal of Fish Diseases, 33(10):859–863, October 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Einarsdottir:2018:MVL

[ESBE18] Thorbjorg Einarsdottir, Heida Sigurdardottir, Thorunn Soley Bjornsdottir, and Elisabet Einarsdottir. Moritella viscosa in lumpfish (Cyclopterus lumpus) and Atlantic salmon (Salmo salar). Journal of Fish Diseases, 41(11):1751–1758, November 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Erkinharju:2019:CI

[ESV^{+}19a] Toni Erkinharju, Guro Strandskog, Øyvind Vågnes, Ivar Hordvik, Roy A. Dalmo, and Tore Seternes. Cover image. Journal of Fish Diseases, 42(12):i, December 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Erkinharju:2019:IVA

[ESV^{+}19b] Toni Erkinharju, Guro Strandskog, Øyvind Vågnes, Ivar Hordvik, Roy A. Dalmo, and Tore Seternes. Intramuscular vaccination of Atlantic lumpfish (Cyclopterus lumpus L.) induces inflammatory reactions and local immunoglobulin M production at the vaccine administration site. Journal of Fish Diseases, 42(12):1731–1743, December 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Elliott:2017:MDW

[EW17] Diane Elliott and Jim Winton. In memoriam: [Dr. William Toshio (Tosh) Yasutake (1922–2016)]. Journal of Fish Diseases, 40(5):607–608, May 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ellul:2019:PPS

[EWH^{+}19] Rebecca Marie Ellul, Cecilie Walde, Gyri Teien Haugland, Heidrun Wergeland, and Anita Rønneseth. Pathogenicity of Pasteurella sp. in lumpsuckers (Cyclopterus lumpus L.). Journal of Fish Diseases, 42(1):35–46, January 2019. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Elmoslemany:2013:SLM

[EWRH13] A. Elmoslemany, S. K. Whyte, C. W. Revie, and K. L. Hammell. Sea lice monitoring on Atlantic salmon farms in New Brunswick, Canada: comparing audit and farm staff counts. Journal of Fish Diseases, 36(3):241–247, March 2013. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Eckart:2019:NCL

[EYF^{+}19] Valentin Eckart, Takuya Yamaguchi, Kati Franzke, Sven M. Bergmann, Pierre Boudinot, Edwige Quillet, Motokazu Kawanobe, Neila Alvarez de Haro, and Uwe Fischer. New cell lines for efficient propagation of koi herpesvirus and infectious salmon anaemia virus. Journal of Fish Diseases, 42 (2):181–187, February 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fabian:2016:TCH

[FBAS16] M. Fabian, A. Baumer, M. Adamek, and D. Steinhagen. Transmission of Cyprinid herpesvirus 3 by wild fish species — results from infection experiments. Journal of Fish Diseases, 39(5):625–628, May 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Florczyk:2016:MPB

[FBK^{+}16] M. Florczyk, P. Brzuzan, J. Krom, M. Woźny, and A. Łakomiak. miR-122-5p as a plasma biomarker of liver injury in fish exposed to microcystin-LR. Journal of Fish Diseases, 39(6):741–751, June 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Failde:2014:IDT

[FBL^{+}14] L. D. Faílde, R. Bermúdez, A. P. Losada, A. Riaza, Y. Santos, and M. I. Quiroga. Immunohistochemical diagnosis of tenacibaculosis in paraffin-embedded tissues of Senegalese sole Solea senegalensis Kaup, 1858. Journal of Fish Diseases, 37(11):959–968, November 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fabian:2013:DWF

[FBS13] M. Fabian, A. Baumer, and D. Steinhagen. Do wild fish species contribute to the transmission of koi herpesvirus to carp in hatchery ponds? Journal of Fish Diseases, 36(5): 505–514, May 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fernandez-Boo:2016:PEP

[FBVC16] S. Fernández-Boo, A. Villalba, and A. Cao. Protein expression profiling in haemocytes and plasma of the Manila clam Ruditapes philippinarum in response to infection with Perkinsus olseni. Journal of Fish Diseases, 39(11):1369– 1385, November 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fichi:2013:DCH

[FCC^{+}13] G. Fichi, G. Cardeti, C. Cocumelli, N. Vendramin, A. Toffan, C. Eleni, N. Siemoni, R. Fischetti, and F. Susini. Detection of Cyprinid herpesvirus 2 in association with an Aeromonas sobria infection of Carassius carassius (L.), in Italy. Journal of Fish Diseases, 36(10):823–830, October 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fetherman:2019:ICC

[FCJH19] Eric R. Fetherman, Pete Cadmus, Abbie L. Jefferson, and Marta K. Hura. Increasing copper concentrations do not affect Myxobolus cerebralis triactinomyxon viability. Journal of Fish Diseases, 42(9):1327–1331, September 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Figueroa:2019:ASN

[FCL^{+}19] Jaime Figueroa, Diana Castro, Fernando Lagos, Carlos Cartes, Adolfo Isla, Alejandro J. Yáñez, Rubén Avendaño-Herrera, and Denise Haussmann. Analysis of single nucleotide polymorphisms (SNPs) associated with antibiotic resistance genes in Chilean Piscirickettsia salmonis strains. Journal of Fish Diseases, 42(12):1645–1655, December 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ferreira:2019:DBE

[FCM^{+}19] Inês Almeida Ferreira, Janina Z. Costa, Valeria Macchia, Kimberly Dawn Thompson, and Teresa Baptista. Detection of Betanodavirus in experimentally infected European seabass (Dicentrarchus labrax, Linnaeus 1758) using nonlethal sampling methods. Journal of Fish Diseases, 42(8): 1097–1105, August 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fujimoto:2018:DSI

[FCS^{+}18] R. Y. Fujimoto, M. V. S. Couto, N. C. Sousa, D. G. Diniz, J. A. P. Diniz, R. R. Madi, M. L. Martins, and J. C. Eiras. Dermocystidium sp. infection in farmed hybrid fish Colossoma macropomum × Piaractus brachypomus in Brazil. Journal of Fish Diseases, 41(3):565–568, March 2018. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Faisal:2017:ICF

[FDL^{+}17] M. Faisal, A. Diamanka, T. P. Loch, B. R. LaFrentz, A. D. Winters, J. C. García, and B. S. Toguebaye. Isolation and characterization of Flavobacterium columnare strains infecting fishes inhabiting the Laurentian Great Lakes basin. Journal of Fish Diseases, 40(5):637–648, May 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Frisch:2016:CIE

[FDST16] K. Frisch, A. Davie, T. Schwarz, and Jf Turnbull. Comparative imaging of European eels (Anguilla anguilla) for the evaluation of swimbladder nematode (Anguillicoloides crassus) infestation. Journal of Fish Diseases, 39(6):635–647, June 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fridholm:2011:VGG

[FE11] H. Fridholm and E. Everitt. Virion glycosylation governs integrity and infectivity of infectious pancreatic necrosis virus. Journal of Fish Diseases, 34(9):663–675, September 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fogelson:2017:PFM

[FFLC17] S. B. Fogelson, M. D. Fast, J. Leary, and A. C. Camus. Pathologic features of mycobacteriosis in naturally infected Syngnathidae and novel transcriptome assembly in association with disease. Journal of Fish Diseases, 40(11):1681– 1694, November 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fraser:2014:EPI

[FFP^{+}14] T. W. K. Fraser, M. S. Fleming, T. T. Poppe, T. Hansen, and P. G. Fjelldal. The effect of ploidy and incubation temperature on survival and the prevalence of aplasia of the septum transversum in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 37(3):189–200, March 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Feng:2010:KAO

[FGJ^{+}10] Q. Feng, T. Gao, H. Ji, X. Jiang, T. Liang, W. Gu, G. Wu, G. Gao, and W. Wang. Kinetic analysis of oxytetracycline residues in Chinese mitten crab, Eriocheir sinensis, muscle following intramuscular administration. Journal of Fish Diseases, 33(8):639–647, August 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fringuelli:2012:DNP

[FGR^{+}12] E. Fringuelli, A. W. Gordon, H. Rodger, M. D. Welsh, and D. A. Graham. Detection of Neoparamoeba perurans by duplex quantitative Taqman real-time PCR in formalin-fixed, paraffin-embedded Atlantic salmonid gill tissues. Journal of Fish Diseases, 35(10):711–724, October 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Forwood:2013:EBO

[FHD13] J. M. Forwood, J. O. Harris, and M. R. Deveney. Efficacy of bath and orally administered praziquantel and fenbendazole against Lepidotrema bidyana Murray, a monogenean parasite of silver perch, Bidyanus bidyanus (Mitchell). Journal of Fish Diseases, 36(11):939–947, November 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fraser:2015:PVD

[FHFF15] T. W. K. Fraser, T. Hansen, M. S. Fleming, and P. G. Fjelldal. The prevalence of vertebral deformities is increased with higher egg incubation temperatures and triploidy in Atlantic salmon Salmo salar L. Journal of Fish Diseases, 38(1):75– 89, January 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Feng:2016:PFB

[FHZ^{+}16] J-B. Feng, D-R. Huang, M. Zhong, P. Liu, and J-D. Dong. Pharmacokinetics of florfenicol and behaviour of its metabolite florfenicol amine in orange-spotted grouper (Epinephelus coioides) after oral administration. Journal of Fish Diseases, 39(7):833–843, July 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ferguson:2014:SHF

[FKB^{+}14] H. W. Ferguson, R. Kabuusu, S. Beltran, E. Reyes, J. A. Lince, and J. del Pozo. Syncytial hepatitis of farmed tilapia, Oreochromis niloticus (L.): a case report. Journal of Fish Diseases, 37(6):583–589, June 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fukushima:2017:LGO

[FLC^{+}17] H. C. S. Fukushima, C. A. G. Leal, R. B. Cavalcante, H. C. P. Figueiredo, S. Arijo, M. A. Moriñigo, M. Ishikawa, R. C. Borra, and M. J. T. Ranzani-Paiva. Lactococcus garvieae outbreaks in Brazilian farms Lactococcosis in Pseudoplatystoma sp. — development of an autogenous vaccine as a control strategy. Journal of Fish Diseases, 40(2):263–272, February 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Faria:2014:CSI

[FLCC^{+}14] F. C. Faria, C. A. G. Leal, G. A. Carvalho-Castro, R. C. Leite, and H. C. P. Figueiredo. Carrier state induced by oxytetracycline therapy against streptococcosis in Nile tilapia, Oreochromis niloticus (L.). Journal of Fish Diseases, 37(9):853–857, September 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fouz:2010:CES

[FLVA10] B. Fouz, A. Llorens, E. Valiente, and C. Amaro. A comparative epizootiologic study of the two fish-pathogenic serovars of Vibrio vulnificus biotype 2. Journal of Fish Diseases, 33 (5):383–390, May 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Foysal:2019:MCI

[FMA^{+}19] Md Javed Foysal, Farhana Momtaz, Md Hazrat Ali, Muhammad A. B. Siddik, Md Reaz Chaklader, Md Mahbubur Rahman, Md Shamsul Haque Prodhan, and Anthony Cole. Molecular characterization and interactome analysis of aerolysin (aer) gene from fish pathogen Aeromonas veronii: the pathogenicity inferred from sequence divergence and linked to histidine kinase (che A). Journal of Fish Diseases, 42(4):465–475, April 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Frans:2011:VAF

[FMB^{+}11] I. Frans, C. W. Michiels, P. Bossier, K. A. Willems, B. Lievens, and H. Rediers. Vibrio anguillarum as a fish pathogen: virulence factors, diagnosis and prevention. Journal of Fish Diseases, 34(9):643–661, September 2011. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fraser:2015:VTI

[FMH^{+}15] T. W. K. Fraser, I. Mayer, T. Hansen, T. T. Poppe, J. E. Skjæraasen, E. O. Koppang, and P. G. Fjelldal. Vaccination and triploidy increase relative heart weight in farmed Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 38 (2):151–160, February 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fernandez:2019:MPM

[FMM^{+}19] Carolina Fernandez, Dario Mascolo, Sean J. Monaghan, Johanna L. Baily, Lynn Chalmers, Giuseppe Paladini, Alexandra Adams, James E. Bron, and Sophie Fridman. Methacarn preserves mucus integrity and improves visualization of amoebae in gills of Atlantic salmon (Salmo salar L.). Journal of Fish Diseases, 42(6):883–894, June 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Farook:2014:PRC

[FMT^{+}14] M. A. Farook, N. Madan, G. Taju, S. Abdul Majeed, K. S. N. Nambi, N. Sundar Raj, S. Vimal, and A. S. Sahul Hameed. Production of recombinant capsid protein of Macrobrachium rosenbergii nodavirus (r-MCP43) of giant freshwater prawn, M. rosenbergii (de Man) for immunological diagnostic methods. Journal of Fish Diseases, 37(8):703–710, August 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fujiwara-Nagata:2012:NGT

[FNISE12] E. Fujiwara-Nagata, J. Ikeda, K. Sugahara, and M. Eguchi. A novel genotyping technique for distinguishing between Flavobacterium psychrophilum isolates virulent and avirulent to ayu, Plecoglossus altivelis altivelis (Temminck & Schlegel). Journal of Fish Diseases, 35(7):471–480, July 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fragkoulis:2017:SSE

[FPK^{+}17] S. Fragkoulis, H. Paliogiannis, P. Kokkinias, K. Chiers, D. Adriaens, and G. Koumoundouros. Saddleback syndrome in European sea bass Dicentrarchus labrax (Linnaeus, 1758): anatomy, ontogeny and correlation with lateral-line, anal and pelvic fin abnormalities. Journal of Fish Diseases, 40(1):83– 95, January 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fields:2018:BRI

[FSB18a] D. M. Fields, A. B. Skiftesvik, and H. I. Browman. Behavioural responses of infective-stage copepodids of the salmon louse (Lepeophtheirus salmonis, Copepoda:Caligidae) to host-related sensory cues. Journal of Fish Diseases, 41 (6):875–884, June 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fields:2018:CI

[FSB18b] D. M. Fields, A. B. Skiftesvik, and H. I. Browman. Cover image. Journal of Fish Diseases, 41(6):i, June 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Frisch:2018:GTM

[FSB^{+}18c] K. Frisch, S. B. Småge, Ø. J. Brevik, H. Duesund, and A. Nylund. Genotyping of Tenacibaculum maritimum isolates from farmed Atlantic salmon in Western Canada. Journal of Fish Diseases, 41(1):131–137, January 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fichi:2016:NDC

[FSC^{+}16] G. Fichi, F. Susini, C. Cocumelli, A. Cersini, M. Salvadori, M. Guarducci, and G. Cardeti. New detection of Cyprinid herpesvirus 2 in mass mortality event of Carassius carassius (L.), in Italy. Journal of Fish Diseases, 39(12):1523–1527, December 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fringuelli:2012:DQR

[FSG^{+}12] E. Fringuelli, P. D. Savage, A. Gordon, E. J. Baxter, H. D. Rodger, and D. A. Graham. Development of a quantitative real-time PCR for the detection of Tenacibaculum maritimum and its application to field samples. Journal of Fish Diseases, 35(8):579–590, August 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fedorcak:2019:PIC

[FŠK^{+}19] Jakub Fedorčák, Łubomír Šmiga, Iuliia Kutsokon, Vladislav Kolarčik, Lenka Koščová, Mikuláš Oros, and Ján Koščo. Parasitic infection of Cobitis elongatoides Băcescu & Mayer, 1969 by zoonotic metacercariae Clinostomum complanatum (Rudolphi, 1814). Journal of Fish Diseases, 42(12):1677– 1685, December 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Flowers:2016:PBP

[FSM^{+}16] E. M. Flowers, K. Simmonds, G. A. Messick, L. Sullivan, and E. J. Schott. PCR-based prevalence of a fatal reovirus of the blue crab, Callinectes sapidus (Rathbun) along the northern Atlantic coast of the USA. Journal of Fish Diseases, 39 (6):705–714, June 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fouzi:2010:TRT

[FSO^{+}10] M. Fouzi, M. Shariff, A. R. Omar, F. Md Yusoff, and S. W. Tan. TaqMan real-time PCR assay for relative quantification of white spot syndrome virus infection in Penaeus monodon Fabricius exposed to ammonia. Journal of Fish Diseases, 33(12):931–938, December 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Favaro:2014:MMI

[FSV^{+}14] L. Favaro, T. Scanzio, K. Varello, M. Caffara, M. Righetti, E. Bozzetta, and M. Prearo. Mixed mycobacterial infection in an adult koi carp Cyprinus carpio L. Journal of Fish Diseases, 37(8):753–755, August 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Frisch:2018:EIM

[FSV^{+}18] K. Frisch, S. B. Småge, C. Vallestad, H. Duesund, Ø. J. Brevik, A. Klevan, R. H. Olsen, S. T. Sjaatil, D. Gauthier, B. Brudeseth, and A. Nylund. Experimental induction of mouthrot in Atlantic salmon smolts using Tenacibaculum maritimum from Western Canada. Journal of Fish Diseases, 41(8):1247–1258, August 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Farlora:2018:HPT

[FVAVMGE18] R. Farlora, F. Valdebenito-Aguayo, V. Valenzuela-Muñoz, and C. Gallardo-Escárate. Hydrogen peroxide treatment modulates the transcription of sex-related genes in the sea lice Caligus rogercresseyi. Journal of Fish Diseases, 41(6): 921–926, June 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fjelldal:2018:RCD

[FvdMF^{+}18] Per Gunnar Fjelldal, Terje van der Meeren, Thomas W. K. Fraser, Florian Sambraus, Laith Jawad, and Tom Johnny Hansen. Radiological changes during fracture and repair in neural and haemal spines of Atlantic cod (Gadus morhua). Journal of Fish Diseases, 41(12):1871–1875, December 2018.

CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Freeman:2011:SMS

[FYO^{+}11] M. A. Freeman, H. Yokoyama, A. Osada, T. Yoshida, A. Yamanobe, and K. Ogawa. Spraguea (Microsporida: Spraguidae) infections in the nervous system of the Japanese anglerfish, Lophius litulon (Jordan), with comments on transmission routes and host pathology. Journal of Fish Diseases, 34(6):445–452, June 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gilmore:2010:PGS

[GAC10] S. R. Gilmore, C. L. Abbott, and D. K. Cone. The placement of Gyrodactylus salmonis (Yin & Sproston) in the molecular phylogeny of studied members of the Gyrodactylus wagenerigroup parasitizing salmonids. Journal of Fish Diseases, 33 (6):461–467, June 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gotesman:2014:IIC

[GAEK^{+}14] M. Gotesman, A. Abd-Elfattah, J. Kattlun, H. Soliman, and M. El-Matbouli. Investigating the interactions of Cyprinid herpesvirus-3 with host proteins in goldfish Carassius auratus. Journal of Fish Diseases, 37(9):835–841, September 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gomez:2016:SRS

[GAVC16] M. P. González Gómez, S. L. Marín Arribas, and L. Vargas-Chacoff. Stress response of Salmo salar (Linnaeus 1758) facing low abundance infestation of Caligus rogercresseyi (Boxshall & Bravo 2000), an object in the tank, and handling. Journal of Fish Diseases, 39(7):853–865, July 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Garseth:2018:LES

[GB18] Åse Helen Garseth and Eirik Biering. Little evidence to suggest salmonid freshwater reservoirs of piscine orthoreovirus (PRV). Journal of Fish Diseases, 41(8):1313–1315, August 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gias:2018:VGG

[GBO^{+}18] E. Gias, C. L. Brosnahan, D. Orr, B. Binney, H. J. Ha, M. A. Preece, and B. Jones. In vivo growth and genomic characterization of rickettsia-like organisms isolated from farmed Chinook salmon (Oncorhynchus tshawytscha) in New Zealand. Journal of Fish Diseases, 41(8):1235–1245, August 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Graham:2012:DSP

[GBSF12] D. A. Graham, A. Brown, P. Savage, and P. Frost. Detection of salmon pancreas disease virus in the faeces and mucus of Atlantic salmon, Salmo salar L., by real-time RT-PCR and cell culture following experimental challenge. Journal of Fish Diseases, 35(12):949–951, December 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gautam:2016:VPT

[GBV^{+}16] R. Gautam, A. S. Boerlage, R. Vanderstichel, C. W. Revie, and K. L. Hammell. Variation in pre-treatment count lead time and its effect on baseline estimates of cage-level sea lice abundance. Journal of Fish Diseases, 39(11):1297–1303, November 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Grijalva-Chon:2013:DNO

[GCCLRP^{+}13] J. M. Grijalva-Chon, R. Castro-Longoria, J. Ramos-Paredes, T. L. Enríquez-Espinoza, and F. Mendoza-Cano. Detection of a new OsHV-1 DNA strain in the healthy Pacific oyster, Crassostrea gigas Thunberg, from the Gulf of California. Journal of Fish Diseases, 36(11):965–968, November 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Groner:2013:UAB

[GCGR13] M. L. Groner, R. Cox, G. Gettinby, and C. W. Revie. Use of agent-based modelling to predict benefits of cleaner fish in controlling sea lice, Lepeophtheirus salmonis, infestations on farmed Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 36(3):195–208, March 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gjessing:2018:CI

[GCM^{+}18a] M. C. Gjessing, D. H. Christensen, F. Manji, S. Mohammad, P. E. Petersen, B. Saure, C. Skjengen, S. C. Weli, and O. B. Dale. Cover image. Journal of Fish Diseases, 41(7):i, July 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gjessing:2018:SGP

[GCM^{+}18b] M. C. Gjessing, D. H. Christensen, F. Manji, S. Mohammad, P. E. Petersen, B. Saure, C. Skjengen, S. C. Weli, and O. B. Dale. Salmon gill poxvirus disease in Atlantic salmon fry as recognized by improved immunohistochemistry also demonstrates infected cells in non-respiratory epithelial cells. Journal of Fish Diseases, 41(7):1103–1110, July 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gonzalez-Contreras:2011:SPS

[GCMG^{+}11] A. González-Contreras, B. Magariños, M. Godoy, R. Irgang, A. E. Toranzo, and R. Avendaño-Herrera. Surface properties of Streptococcus phocae strains isolated from diseased Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 34 (3):203–215, March 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ghosh:2016:MPP

[GCNB16] B. Ghosh, K. D. Cain, B. F. Nowak, and A. R. Bridle. Microencapsulation of a putative probiotic Enterobacter species, C6-6, to protect rainbow trout, Oncorhynchus mykiss (Walbaum), against bacterial coldwater disease. Journal of Fish Diseases, 39(1):1–11, January 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ghosh:2018:HSN

[GCNB18] Bikramjit Ghosh, Philip B. B. Crosbie, Barbara F. Nowak, and Andrew R. Bridle. A highly sensitive, non-invasive qPCR-based strategy for direct quantification of Yersinia ruckeri in fish faeces. Journal of Fish Diseases, 41(9):1421– 1428, September 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gustafson:2018:SSP

[GCS^{+}18] L. L. Gustafson, L. H. Creekmore, K. R. Snekvik, J. A. Ferguson, J. V. Warg, M. Blair, T. R. Meyers, B. Stewart, K. I. Warheit, J. Kerwin, A. E. Goodwin, L. D. Rhodes, J. E. Whaley, M. K. Purcell, C. Bentz, D. Shasa, J. Bader, and J. R. Winton. A systematic surveillance programme for infectious salmon anaemia virus supports its absence in the Pacific Northwest of the United States. Journal of Fish Diseases, 41(2):337–346, February 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gieseker:2017:PEC

[GCW17] C. M. Gieseker, T. C. Crosby, and L. C. Woods I.I.I. Provisional epidemiological cutoff values for standard broth microdilution susceptibility testing of Flavobacterium columnare. Journal of Fish Diseases, 40(5):679–686, May 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Grimmett:2011:SDT

[GCWB11] S. G. Grimmett, H. J. Chalmers, J. C. Wolf, and P. R. Bowser. Spinal deformity in triploid grass carp Ctenopharyngodon idella (Valenciennes). Journal of Fish Diseases, 34(3): 217–225, March 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gulla:2016:ASI

[GDN^{+}16] S. Gulla, S. Duodu, A. Nilsen, I. Fossen, and D. J. Colquhoun. Aeromonas salmonicida infection levels in preand post-stocked cleaner fish assessed by culture and an amended qPCR assay. Journal of Fish Diseases, 39(7):867– 877, July 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Good:2015:FBF

[GDW^{+}15] C. Good, J. Davidson, G. D. Wiens, T. J. Welch, and S. Summerfelt. Flavobacterium branchiophilum and F. succinicans associated with bacterial gill disease in rainbow trout Oncorhynchus mykiss (Walbaum) in water recirculation aquaculture systems. Journal of Fish Diseases, 38(4):409–413, April 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Giacopello:2013:DVE

[GFB^{+}13] C. Giacopello, M. Foti, T. Bottari, V. Fisichella, and G. Barbera. Detection of viral encephalopathy and retinopathy virus (VERV) in wild marine fish species of the South Tyrrhenian Sea (Central Mediterranean). Journal of Fish Diseases, 36(9):819–821, September 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Getchell:2019:IRG

[GFB^{+}19] Rodman G. Getchell, Erika J. First, Steven M. Bogdanowicz, Jose A. Andrés, Adam T. Schulman, Jordan Kramer, Geofrey E. Eckerlin, John M. Farrell, and Hélène Marquis. Investigation of round goby viral haemorrhagic septicaemia outbreak in New York. Journal of Fish Diseases, 42(7):1023– 1033, July 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Graham:2011:CSM

[GFM^{+}11] D. A. Graham, P. Frost, K. McLaughlin, H. M. Rowley, I. Gabestad, A. Gordon, and M. F. McLoughlin. A comparative study of marine salmonid alphavirus subtypes 1–6 using an experimental cohabitation challenge model. Journal of Fish Diseases, 34(4):273–286, April 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gay:2016:HHC

[GFM^{+}16] F. Gay, I. Ferrandino, A. Monaco, M. Cerulo, G. Capasso, and A. Capaldo. Histological and hormonal changes in the European eel (Anguilla anguilla) after exposure to environmental cocaine concentration. Journal of Fish Diseases, 39 (3):295–308, March 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Garseth:2013:PRP

[GFO^{+}13] Å. H. Garseth, C. Fritsvold, M. Opheim, E. Skjerve, and E. Biering. Piscine reovirus (PRV) in wild Atlantic salmon, Salmo salar L., and sea-trout, Salmo trutta L., in Norway. Journal of Fish Diseases, 36(5):483–493, May 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Graham:2012:GDS

[GFR^{+}12] D. A. Graham, E. Fringuelli, H. M. Rowley, D. Cockerill, D. I. Cox, T. Turnbull, H. Rodger, D. Morris, and M. F. Mc Loughlin. Geographical distribution of salmonid alphavirus subtypes in marine farmed Atlantic salmon, Salmo salar L., in Scotland and Ireland. Journal of Fish Diseases, 35(10): 755–765, October 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Granzow:2014:IPC

[GFS^{+}14] H. Granzow, D. Fichtner, H. Schütze, M. Lenk, B. Dresenkamp, H. Nieper, and T. C. Mettenleiter. Isolation and partial characterization of a novel virus from different carp species suffering gill necrosis — ultrastructure and morphogenesis. Journal of Fish Diseases, 37(6):559–569, June 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Garseth:2018:CSA

[GFS^{+}18] Å. H. Garseth, C. Fritsvold, J. C. Svendsen, B. Bang Jensen, and A. B. Mikalsen. Cardiomyopathy syndrome in Atlantic salmon Salmo salar L.: a review of the current state of knowledge. Journal of Fish Diseases, 41(1):11–26, January 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Graham:2010:PLS

[GFW^{+}10] D. A. Graham, E. Fringuelli, C. Wilson, H. M. Rowley, A. Brown, H. Rodger, M. F. McLoughlin, C. McManus, E. Casey, L. J. McCarthy, and N. M. Ruane. Prospective longitudinal studies of salmonid alphavirus infections on two Atlantic salmon farms in Ireland; evidence for viral persistence. Journal of Fish Diseases, 33(2):123–135, February 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gudmundsdottir:2014:YAC

[GGGM14] B. K. Gudmundsdottir, S. Gudmundsdottir, S. Gudmundsdottir, and B. Magnadottir. Yersiniosis in Atlantic cod, Gadus morhua (L.), characterization of the infective strain and host reactions. Journal of Fish Diseases, 37(6):511–519, June 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Godoy:2010:FDA

[GGH^{+}10] M. Godoy, V. Gherardelli, A. Heisinger, J. Fernández, P. Olmos, L. Ovalle, P. Ilardi, and R. Avendaño-Herrera. First description of atypical furunculosis in freshwater farmed Atlantic salmon, Salmo salar L., in Chile. Journal of Fish Diseases, 33(5):441–449, May 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Garseth:2018:CI

[GGMG18a] Å. H. Garseth, M. C. Gjessing, T. Moldal, and A. G. Gjevre. Cover image. Journal of Fish Diseases, 41(1):i, January 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Garseth:2018:SSG

[GGMG18b] Å. H. Garseth, M. C. Gjessing, T. Moldal, and A. G. Gjevre. A survey of salmon gill poxvirus (SGPV) in wild salmonids in Norway. Journal of Fish Diseases, 41(1):139–145, January 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Grady:2011:VRE

[GGW^{+}11] C. A. Grady, J. L. Gregg, R. M. Wade, J. R. Winton, and P. K. Hershberger. Viral replication in excised fin tissues (VREFT) corresponds with prior exposure of Pacific herring, Clupea pallasii (Valenciennes), to viral haemorrhagic septicaemia virus (VHSV). Journal of Fish Diseases, 34(1):3–12, January 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Grini:2011:EWT

[GHB^{+}11] A. Grini, T. Hansen, A. Berg, A. Wargelius, and P. G. Fjelldal. The effect of water temperature on vertebral deformities and vaccine-induced abdominal lesions in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 34(7):531–546, July 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gong:2011:ECN

[GHH^{+}11] J. Gong, Y. Huang, X. Huang, Z. Ouyang, M. Guo, and Q. Qin. Establishment and characterization of a new cell line derived from kidney of grouper, Epinephelus akaara (Temminck & Schlegel), susceptible to Singapore grouper iridovirus (SGIV). Journal of Fish Diseases, 34(9):677–686, September 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Guo:2015:ENC

[GHW^{+}15] C. Y. Guo, Y. H. Huang, S. N. Wei, Z. L. Ouyang, Y. Yan, X. H. Huang, and Q. W. Qin. Establishment of a new cell line from the heart of giant grouper, Epinephelus lanceolatus (Bloch), and its application in toxicology and virus susceptibility. Journal of Fish Diseases, 38(2):175–186, February 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gislason:2018:SMS

[Gis18] H. Gislason. Statistical modelling of sea lice count data from salmon farms in the Faroe Islands. Journal of Fish Diseases, 41(6):973–993, June 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gjessing:2011:PII

[GIW^{+}11] M. C. Gjessing, M. Inami, S. C. Weli, T. Ellingsen, K. Falk, E. O. Koppang, and A. Kvellestad. Presence and interaction of inflammatory cells in the spleen of Atlantic cod, Gadus morhua L., infected with Francisella noatunensis. Journal of Fish Diseases, 34(9):687–699, September 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gias:2011:DRT

[GJK^{+}11] E. Gias, C. Johnston, S. Keeling, R. P. Spence, and W. L. McDonald. Development of real-time PCR assays for detection of megalocytiviruses in imported ornamental fish. Journal of Fish Diseases, 34(8):609–618, August 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

George:2015:ICR

[GJM^{+}15] M. R. George, K. R. John, M. M. Mansoor, R. Saravanakumar, P. Sundar, and V. Pradeep. Isolation and characterization of a ranavirus from koi, Cyprinus carpio L., experiencing mass mortalities in India. Journal of Fish Diseases, 38 (4):389–403, April 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gomes:2017:CSP

[GJMH17] G. Bastos Gomes, D. R. Jerry, T. L. Miller, and K. S. Hutson. Current status of parasitic ciliates Chilodonella spp. (Phyllopharyngea: Chilodonellidae) in freshwater fish aquaculture. Journal of Fish Diseases, 40(5):703–715, May 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gadd:2011:EAV

[GJVT^{+}11] T. Gadd, M. Jakava-Viljanen, H. Tapiovaara, P. Koski, and L. Sihvonen. Epidemiological aspects of viral haemorrhagic septicaemia virus genotype II isolated from Baltic herring, Clupea harengus membras L. Journal of Fish Diseases, 34 (7):517–529, July 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gibson-Kueh:2012:PSD

[GKCC^{+}12] S. Gibson-Kueh, D. Chee, J. Chen, Y. H. Wang, S. Tay, L. N. Leong, M. L. Ng, J. B. Jones, P. K. Nicholls, and H. W. Ferguson. The pathology of ‘scale drop syndrome’ in Asian seabass, Lates calcarifer Bloch, a first description. Journal of Fish Diseases, 35(1):19–27, January 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gombac:2014:MMR

[GKO^{+}14] M. Gombač, D. Kušar, M. Ocepek, M. Pogačnik, I. Arzul, Y. Couraleau, and V. Jenčič. Marteiliosis in mussels: a rare disease? Journal of Fish Diseases, 37(9):805–814, September 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gagne:2018:OIS

[GL18] N. Gagné and F. LeBlanc. Overview of infectious salmon anaemia virus (ISAV) in Atlantic Canada and first report of an ISAV North American-HPR0 subtype. Journal of Fish Diseases, 41(3):421–430, March 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gulla:2016:VLT

[GLK^{+}16] S. Gulla, V. Lund, A. B. Kristoffersen, H. Sørum, and D. J. Colquhoun. vapA (A-layer) typing differentiates Aeromonas salmonicida subspecies and identifies a number of previously undescribed subtypes. Journal of Fish Diseases, 39(3):329– 342, March 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Garcia:2018:CAF

[GLW^{+}18] J. C. García, B. R. LaFrentz, G. C. Waldbieser, F. S. Wong, and S. F. Chang. Characterization of atypical Flavobacterium columnare and identification of a new genomovar. Journal of Fish Diseases, 41(7):1159–1164, July 2018. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Goodwin:2010:VQA

[GM10] A. E. Goodwin and E. Marecaux. Validation of a qPCR assay for the detection of Ictalurid herpesvirus-2 (IcHV-2) in fish tissues and cell culture supernatants. Journal of Fish Diseases, 33(4):341–346, April 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Goertz:2011:ASC

[GM11] C. E. C. Goertz and D. M. Mulcahy. Attempted surgical correction of single- and multiyear post-ovulatory egg stasis in yellow and red Irish lords, Hemilepidotus jordani (Bean) and Hemilepidotus hemilepidotus (Tilesius). Journal of Fish Diseases, 34(1):75–79, January 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Galeotti:2017:UBD

[GMB^{+}17] M. Galeotti, M. Manzano, P. Beraldo, C. Bulfon, G. Rossi, D. Volpatti, and G. E. Magi. Ultrastructural and biomolecular detection of Rickettsiales-like organisms in tissues of rainbow trout with Red Mark Syndrome. Journal of Fish Diseases, 40(7):907–917, July 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Garver:2016:PRJ

[GMC^{+}16a] K. A. Garver, G. D. Marty, S. N. Cockburn, J. Richard, L. M. Hawley, A. Müller, R. L. Thompson, M. K. Purcell, and S. Saksida. Piscine reovirus, but not Jaundice Syndrome, was transmissible to Chinook Salmon, Oncorhynchus tshawytscha (Walbaum), Sockeye Salmon, Oncorhynchus nerka (Walbaum), and Atlantic Salmon, Salmo salar L.

Journal of Fish Diseases, 39(2):117–128, February 2016. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Good:2016:AIS

[GMC^{+}16b] C. Good, T. May, C. Crouse, S. Summerfelt, and T. J. Welch. Assessing the impact of swimming exercise and the relative susceptibility of rainbow trout Oncorhynchus mykiss (Walbaum) and Atlantic salmon Salmo salar L. following injection challenge with Weissella ceti. Journal of Fish Diseases, 39(11):1387–1391, November 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Garseth:2019:POP

[GMG^{+}19] Åse Helen Garseth, Torfinn Moldal, Siri Kristine Gåsnes, Monika Jankowska Hjortaas, Vegard Pedersen Sollien, and Anne-Gerd Gjevre. Piscine orthoreovirus-3 is prevalent in wild seatrout (Salmo trutta L.) in Norway. Journal of Fish Diseases, 42(3):391–396, March 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Goodwin:2012:PVR

[GMN12] A. E. Goodwin, G. E. Merry, and A. D. Noyes. Persistence of viral RNA in fish infected with VHSV-IVb at 15^{o}C and then moved to warmer temperatures after the onset of disease. Journal of Fish Diseases, 35(7):523–528, July 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gopalakrishnan:2011:NIO

[GMP^{+}11] A. Gopalakrishnan, B. A. V. Maran, V. Puvanendran, M. Rajkumar, T. Balasubramanian, and H. W. Ferguson. Neoplasia in the Indian oil sardine, Sardinella longiceps (Valenciennes), and the great barracuda, Sphyraena barracuda (Edwards), from the south-east coast of India. Journal of Fish Diseases, 34(11):881–885, November 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gomez:2012:HLI

[Góm12] Serafín Gómez. Histological lesions induced by exogenous pigment in dyed Indian glassfish, Pseudambassis ranga (Hamilton). Journal of Fish Diseases, 35(12):953–954, December 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gomez:2019:SCR

[GOM^{+}19] Margarita P. González Gómez, Loreto Ovalle, Mylena Menanteau, Claudia Spinetto, Renato Oyarzún, Mario Rivas, and Cristian Oyarzo. Susceptibility of Caligus rogercresseyi collected from the native fish species Eleginops maclovinus (Cuvier) to antiparasitics applied by immersion. Journal of Fish Diseases, 42(8):1143–1149, August 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gonzalez-Palacios:2019:BSR

[GPFGAG19] Concepción González-Palacios, Juan-Miguel Fregeneda-Grandes, and José-Miguel Aller-Gancedo. Biocontrol of saprolegniosis in rainbow trout (Oncorhynchus mykiss Walbaum) using two bacterial isolates (LE89 and LE141) of Pseudomonas fluorescens. Journal of Fish Diseases, 42(2): 269–275, February 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gliniewicz:2012:CPA

[GPL^{+}12] K. Gliniewicz, K. P. Plant, S. E. LaPatra, B. R. LaFrentz, K. Cain, K. R. Snekvik, and D. R. Call. Comparative proteomic analysis of virulent and rifampicin-attenuated Flavobacterium psychrophilum. Journal of Fish Diseases, 35 (7):529–539, July 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Griffin:2019:RCE

[GPWF19] Matt J. Griffin, B. Denise Petty, Cynthia Ware, and Susan B. Fogelson. Recovery and confirmation of Edwardsiella piscicida from a black crappie Pomoxis nigromaculatus (Lesueur, 1829). Journal of Fish Diseases, 42(10):1457–1461, October 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Galeotti:2017:FRR

[GRB^{+}17] M. Galeotti, P. Ronza, P. Beraldo, C. Bulfon, G. E. Magi, M. Manzano, and D. Volpatti. First report of Red Mark Syndrome (RMS) in farmed rainbow trout in Slovenia. Journal of Fish Diseases, 40(12):1935–1939, December 2017. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Graham:2014:CNS

[GRF14] D. A. Graham, H. R. Rowley, and P. Frost. Crossneutralization studies with salmonid alphavirus subtype 1– 6 strains: results with sera from experimental studies and natural infections. Journal of Fish Diseases, 37(8):683–691, August 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Griffin:2016:CEI

[GRG^{+}16] M. J. Griffin, S. R. Reichley, T. E. Greenway, S. M. Quiniou, C. Ware, D. X. Gao, P. S. Gaunt, R. P. E. Yanong, D. B. Pouder, J. P. Hawke, and E. Soto. Comparison of Edwardsiella ictaluri isolates from different hosts and geographic origins. Journal of Fish Diseases, 39(8):947–969, August 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Grice:2015:WSV

[GRP^{+}15] J. Grice, A. Reid, A. Peterson, K. Blackburn, L. Tubbs, S. Lord, P. Huber, R. Horricks, B. Dixon, N. C. Bols, and J. S. Lumsden. Walleye Sander vitreus (Mitchill) are relatively resistant to experimental infection with VHSV IVb and extant walleye strains vary in susceptibility. Journal of Fish Diseases, 38(10):859–872, October 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Grammes:2012:TAM

[GRT12] F. Grammes, K-A. Rørvik, and H. Takle. Tetradecylthioacetic acid modulates cardiac transcription in Atlantic salmon, Salmo salar L., suffering heart and skeletal muscle inflammation. Journal of Fish Diseases, 35(2):109–117, February 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gotesman:2015:ISV

[GSBEM15] M. Gotesman, H. Soliman, R. Besch, and M. El-Matbouli. Inhibition of spring viraemia of carp virus replication in an Epithelioma papulosum cyprini cell line by RNAi. Journal of Fish Diseases, 38(2):197–207, February 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gotesman:2013:ASI

[GSEM13] M. Gotesman, H. Soliman, and M. El-Matbouli. Antibody screening identifies 78 putative host proteins involved in Cyprinid herpesvirus 3 infection or propagation in common carp, Cyprinus carpio L. Journal of Fish Diseases, 36(8): 721–733, August 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Glover:2010:PEB

[GSS^{+}10] K. A. Glover, O. B. Samuelsen, O. T. Skilbrei, K. Boxaspen, and B. T. Lunestad. Pharmacokinetics of emamectin benzoate administered to Atlantic salmon, Salmo salar L., by intra-peritoneal injection. Journal of Fish Diseases, 33(2): 183–186, February 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Garcia-Torrico:2019:CPV

[GTGCM19] Ana Isabel García-Torrico, José A. Guijarro, Desirée Cascales, and Jessica Méndez. Changes in physiology and virulence during the selection of resistant Yersinia ruckeri mutants under subinhibitory cefotaxime concentrations. Journal of Fish Diseases, 42(12):1687–1696, December 2019. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Glenn:2011:URT

[GTH11] R. A. Glenn, P. W. Taylor, and K. C. Hanson. The use of a real-time PCR primer/probe set to observe infectivity of Yersinia ruckeri in Chinook salmon, Oncorhynchus tshawytscha (Walbaum), and steelhead trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 34(10):783– 791, October 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gjessing:2017:SGP

[GTT^{+}17] M. C. Gjessing, E. Thoen, T. Tengs, S. A. Skotheim, and O. B. Dale. Salmon gill poxvirus, a recently characterized infectious agent of multifactorial gill disease in freshwaterand seawater-reared Atlantic salmon. Journal of Fish Diseases, 40(10):1253–1265, October 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gautam:2017:ETC

[GVB^{+}17a] R. Gautam, R. Vanderstichel, A. S. Boerlage, C. W. Revie, and K. L. Hammell. Effect of timing of count events on estimates of sea lice abundance and interpretation of effectiveness following bath treatments. Journal of Fish Diseases, 40 (3):367–375, March 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gautam:2017:EBT

[GVB^{+}17b] R. Gautam, R. Vanderstichel, A. S. Boerlage, C. W. Revie, and K. L. Hammell. Evaluating bath treatment effectiveness in the control of sea lice burdens on Atlantic salmon in New Brunswick, Canada. Journal of Fish Diseases, 40(7):895– 905, July 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Guethmundsdottir:2019:OVH

[GVC^{+}19] Sigrídur Gudmundsdóttir, Niccoló Vendramin, Argelia Cuenca, Heida Sigurdardóttir, Arni Kristmundsson, Tine Moesgaard Iburg, and Niels Jørgen Olesen. Outbreak of viral haemorrhagic septicaemia (VHS) in lumpfish (Cyclopterus lumpus) in Iceland caused by VHS virus genotype IV. Journal of Fish Diseases, 42(1):47–62, January 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gittenberger:2016:OHO

[GVLE16] A. Gittenberger, M. A. Voorbergen-Laarman, and M. Y. Engelsma. Ostreid herpesvirus OsHV-1μVar in Pacific oysters Crassostrea gigas (Thunberg 1793) of the Wadden Sea, a UN-ESCO world heritage site. Journal of Fish Diseases, 39(1): 105–109, January 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Green:2012:PTS

[GWM12] D. M. Green, M. Werkman, and L. A. Munro. The potential for targeted surveillance of live fish movements in Scotland. Journal of Fish Diseases, 35(1):29–37, January 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Horner:2010:MNE

[HAP^{+}10] D. Horner, S. D. Atkinson, D. M. Pratt, S. Marcquenski, and J. L. Bartholomew. Myxobolus notropis from emerald shiner, Notropis atherinoides Rafinesque, in Lake Superior. Journal of Fish Diseases, 33(3):279–282, March 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hoferer:2019:IDP

[HAS^{+}19] Marc Hoferer, Valerij Akimkin, Julia Skrypski, Heike Schütze, and Reinhard Sting. Improvement of a diagnostic procedure in surveillance of the listed fish diseases IHN and VHS. Journal of Fish Diseases, 42(4):559–572, April 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hurst:2012:CSG

[HB12] C. N. Hurst and J. L. Bartholomew. Ceratomyxa shasta genotypes cause differential mortality in their salmonid hosts. Journal of Fish Diseases, 35(10):725–732, October 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hall:2014:CSB

[HBB14] K. C. Hall, M. K. Broadhurst, and P. A. Butcher. Clinical signs of barotrauma in golden perch, Macquaria ambigua (Richardson), and associated effects on post-release mortality and health. Journal of Fish Diseases, 37(3):251– 264, March 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Henriquez:2015:PGA

[HBB^{+}15] P. Henríquez, H. Bohle, F. Bustamante, P. Bustos, and M. Mancilla. Polymorphism in gyrA is associated to quinolones resistance in Chilean Piscirickettsia salmonis field isolates. Journal of Fish Diseases, 38(4):415–418, April 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hung:2019:DRE

[HBH^{+}19] Nguyen Van Hung, Peter Bossier, Nguyen Thi X. Hong, Christine Ludeseve, Linsey Garcia-Gonzalez, Nancy Nevejan, and Peter De Schryver. Does Ralstonia eutropha, rich in poly-β hydroxybutyrate (PHB), protect blue mussel larvae against pathogenic vibrios? Journal of Fish Diseases, 42 (6):777–787, June 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Helgesen:2014:DRS

[HBS^{+}14] K. O. Helgesen, S. Bravo, S. Sevatdal, J. Mendoza, and T. E. Horsberg. Deltamethrin resistance in the sea louse Caligus rogercresseyi (Boxhall and Bravo) in Chile: bioassay results and usage data for antiparasitic agents with references to Norwegian conditions. Journal of Fish Diseases, 37(10): 877–890, October 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hong:2016:CPV

[HBVB16] N. T. X. Hong, K. Baruah, D. Vanrompay, and P. Bossier. Characterization of phenotype variations of luminescent and non-luminescent variants of Vibrio harveyi wild type and quorum sensing mutants. Journal of Fish Diseases, 39(3): 317–327, March 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hadfield:2012:AMM

[HCCM12] C. A. Hadfield, L. A. Clayton, D. K. Cohrs, and D. S. Murphy. Acute morbidity and mortality in invertebrates and fish following exposure to potassium-deficient saltwater. Journal of Fish Diseases, 35(7):549–553, July 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hart:2016:PES

[HCEH16] L. M. Hart, C. M. Conway, D. G. Elliott, and P. K. Hershberger. Persistence of external signs in Pacific herring Clupea pallasii Valenciennes with ichthyophoniasis. Journal of Fish Diseases, 39(4):429–440, April 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ho:2014:EPE

[HCL^{+}14] L-P. Ho, C-J. Chang, H-C. Liu, H-L. Yang, and J. H-Y. Lin. Evaluating the protective efficacy of antigen combinations against Photobacterium damselae ssp. piscicida infections in cobia, Rachycentron canadum L. Journal of Fish Diseases, 37(1):51–62, January 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Huang:2016:MCC

[HCW^{+}16] Y. Huang, J. Cai, B. Wang, J-F. Tang, J-C. Jian, Z-H. Wu, Z. Gan, and Y-S. Lu. Molecular cloning and characterization of lymphocyte cell kinase from humphead snapper (Lutjanus sanguineus). Journal of Fish Diseases, 39(7):809–819, July 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Haetrakul:2018:AAC

[HDC18] T. Haetrakul, S. G. Dunbar, and N. Chansue. Antiviral activities of Clinacanthus nutans (Burm.f.) Lindau extract against Cyprinid herpesvirus 3 in koi (Cyprinus carpio koi). Journal of Fish Diseases, 41(4):581–587, April 2018. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hall:2015:EPR

[HDWM15] L. M. Hall, S. Duguid, I. S. Wallace, and A. G. Murray. Estimating the prevalence of Renibacterium salmoninaruminfected salmonid production sites. Journal of Fish Diseases, 38(2):231–235, February 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Humborstad:2017:EWC

[HFKF17] O-B. Humborstad, K. Ferter, H. Kryvi, and P. G. Fjelldal. Exophthalmia in wild-caught cod (Gadus morhua L.): development of a secondary barotrauma effect in captivity. Journal of Fish Diseases, 40(1):41–49, January 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Herath:2012:UMS

[HFT^{+}12] T. K. Herath, H. W. Ferguson, K. D. Thompson, A. Adams, and R. H. Richards. Ultrastructural morphogenesis of salmonid alphavirus 1. Journal of Fish Diseases, 35(11): 799–808, November 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hodneland:2011:RTR

[HGB^{+}11] K. Hodneland, R. García, J. A. Balbuena, C. Zarza, and B. Fouz. Real-time RT-PCR detection of betanodavirus in naturally and experimentally infected fish from Spain. Journal of Fish Diseases, 34(3):189–202, March 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hershberger:2011:FCE

[HGG^{+}11] P. K. Hershberger, J. L. Gregg, C. A. Grady, L. M. Hart, S. R. Roon, and J. R. Winton. Factors controlling the early stages of viral haemorrhagic septicaemia epizootics: low exposure levels, virus amplification and fish-to-fish transmission. Journal of Fish Diseases, 34(12):893–899, December 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hershberger:2016:PIS

[HGH^{+}16] P. K. Hershberger, J. L. Gregg, L. M. Hart, S. Moffitt, R. Brenner, K. Stick, E. Coonradt, E. O. Otis, J. J. Vollenweider, K. A. Garver, J. Lovy, and T. R. Meyers. The parasite Ichthyophonus sp. in Pacific herring from the coastal NE Pacific. Journal of Fish Diseases, 39(4):395–410, April 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Halttunen:2018:STA

[HGH^{+}18] E. Halttunen, K-Ø Gjelland, S. Hamel, R-M. Serra-Llinares, R. Nilsen, P. Arechavala-Lopez, J. Skardhamar, I. A. Johnsen, L. Asplin, Ø. Karlsen, P-A. Bjørn, and B. Finstad. Sea trout adapt their migratory behaviour in response to high salmon lice concentrations. Journal of Fish Diseases, 41(6):953–967, June 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Huchzermeyer:2011:SWS

[HGPD11] K. D. A. Huchzermeyer, D. Govender, D. J. Pienaar, and A. R. Deacon. Steatitis in wild sharptooth catfish, Clarias gariepinus (Burchell), in the Olifants and Lower Letaba Rivers in the Kruger National Park, South Africa. Journal of Fish Diseases, 34(7):489–498, July 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hao:2019:SLC

[HGW^{+}19] Wenjing Hao, Qi Gao, Jian Wang, Wei Gu, Wen Wang, and Qingguo Meng. SPE0313 located at cell membrane of Spiroplasma eriocheiris is required for adhesion and invasion Eriocheir sinensis hemocytes. Journal of Fish Diseases, 42(3):

423–430, March 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Huang:2019:RGR

[HGX^{+}19] Lixing Huang, Lina Guo, Xiaojin Xu, Yingxue Qin, Lingmin Zhao, Yongquan Su, and Qingpi Yan. The role of rpoS in the regulation of Vibrio alginolyticus virulence and the response to diverse stresses. Journal of Fish Diseases, 42(5):703–712, May 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Helgesen:2013:IDM

[HH13a] K. O. Helgesen and T. E. Horsberg. Influence of different materials on the concentration of delousing agents in sea water during bioassays. Journal of Fish Diseases, 36(5):529–532, May 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Helgesen:2013:SDF

[HH13b] K. O. Helgesen and T. E. Horsberg. Single-dose field bioassay for sensitivity testing in sea lice, Lepeophtheirus salmonis: development of a rapid diagnostic tool. Journal of Fish Diseases, 36(3):261–272, March 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Huchzermeyer:2018:FRE

[HHC^{+}18] C. F. Huchzermeyer, K. D. A. Huchzermeyer, K. W. Christison, B. M. Macey, P. A. Colly, B. M. Hang’ombe, and M. M. Songe. First record of epizootic ulcerative syndrome from the Upper Congo catchment: an outbreak in the Bangweulu swamps, Zambia. Journal of Fish Diseases, 41(1):87–94, January 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Herbers:2014:PDB

[HHM^{+}14] M. Herbers, H. Hariharan, V. Matthew, S. Oliveira, A. Rovira, R. N. Sharma, and H. W. Ferguson. Photobacterium damselae in bigeye scad, Selar crumenophthalmus (Bloch), from Grenada, West Indies. Journal of Fish Diseases, 37(7):675–677, July 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hagen:2014:LCS

[HHO14] A. G. Hagen, S. Hytterød, and K. Olstad. Low concentrations of sodium hypochlorite affect population dynamics in Gyrodactylus salaris (Malmberg, 1957): practical guidelines for the treatment of the Atlantic salmon, Salmo salar L. parasite. Journal of Fish Diseases, 37(12):1003–1011, December 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Huang:2015:ICN

[HHX^{+}15] X. Huang, Y. Huang, L. Xu, S. Wei, Z. Ouyang, J. Feng, and Q. Qin. Identification and characterization of a novel lymphocystis disease virus isolate from cultured grouper in China. Journal of Fish Diseases, 38(4):379–387, April 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hoare:2019:EPI

[HJN^{+}19] Rowena Hoare, Sung-Ju Jung, Thao P. H. Ngo, Kerry L. Bartie, Kim D. Thompson, and Alexandra Adams. Efficacy of a polyvalent injectable vaccine against Flavobacterium psychrophilum administered to rainbow trout (Oncorhynchus mykiss L.). Journal of Fish Diseases, 42(2):229–236, February 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Han:2011:RSD

[HJOK11] H-J. Han, S-J. Jung, M-J. Oh, and D-H. Kim. Rapid and sensitive detection of Streptococcus iniae by loop-mediated isothermal amplification (LAMP). Journal of Fish Diseases, 34(5):395–398, May 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hjortaas:2016:GCS

[HJT^{+}16] M. J. Hjortaas, B. Bang Jensen, T. Taksdal, A. B. Olsen, A. Lillehaug, E. Trettenes, and H. Sindre. Genetic characterization of salmonid alphavirus in Norway. Journal of Fish Diseases, 39(2):249–257, February 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Henriquez:2016:CAS

[HKB^{+}16] P. Henríquez, M. Kaiser, H. Bohle, P. Bustos, and M. Mancilla. Comprehensive antibiotic susceptibility profiling of Chilean Piscirickettsia salmonis field isolates. Journal of Fish Diseases, 39(4):441–448, April 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Henriksen:2015:EHP

[HKBD15a] M. M. M. Henriksen, P. W. Kania, K. Buchmann, and I. Dalsgaard. Effect of hydrogen peroxide and/or Flavobacterium psychrophilum on the gills of rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 38 (3):259–270, March 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Henriksen:2015:EIR

[HKBD15b] M. M. M. Henriksen, P. W. Kania, K. Buchmann, and I. Dalsgaard. Evaluation of the immune response in rainbow trout fry, Oncorhynchus mykiss (Walbaum), after waterborne exposure to Flavobacterium psychrophilum and/or hydrogen peroxide. Journal of Fish Diseases, 38(1):55–66, January 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hellberg:2010:OVN

[HKD^{+}10] H. Hellberg, A. Kvellestad, B. Dannevig, G. Bornø, I. Modahl, R. N. Haldorsen, F. Vik-Mo, K. Ottesen, E. M. Sætre, and H. Sindre. Outbreaks of viral nervous necrosis in juvenile and adult farmed Atlantic cod, Gadus morhua L., in Norway. Journal of Fish Diseases, 33(1):75–81, January 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Huveneers:2016:FHE

[HKF^{+}16] C. Huveneers, S. Klebe, A. Fox, B. Bruce, R. Robbins, J. D. Borucinska, R. Jones, and M. Z. Michael. First histological examination of a neoplastic lesion from a free-swimming white shark, Carcharodon carcharias L. Journal of Fish Diseases, 39(10):1269–1273, October 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hendrichsen:2015:TDM

[HKG^{+}15] D. K. Hendrichsen, R. Kristoffersen, K. Ø. Gjelland, R. Knudsen, S. Kusterle, A. H. Rikardsen, E. H. Henriksen, A. Smalås, and K. Olstad. Transmission dynamics of the monogenean Gyrodactylus salaris under seminatural conditions. Journal of Fish Diseases, 38(6):541–550, June 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Han:2011:AAS

[HKK^{+}11a] H-J. Han, D-Y. Kim, W-S. Kim, C-S. Kim, S-J. Jung, M-J. Oh, and D-H. Kim. Atypical Aeromonas salmonicida infection in the black rockfish, Sebastes schlegeli Hilgendorf, in Korea. Journal of Fish Diseases, 34(1):47–55, January 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Han:2011:PDC

[HKK^{+}11b] S. Y. Han, B. K. Kang, B. J. Kang, S. P. Shin, B. H. Soen, J. M. Kim, J. H. Kim, C. H. Choresca, Jr., J. E. Han, J. W. Jun, and S. C. Park. Prevalence and different characteristics of two serotypes of Streptococcus parauberis isolated from the farmed olive flounder, Paralichthys olivaceus (Temminck and Schlegel), in Korea. Journal of Fish Diseases, 34(10): 731–739, October 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hu:2015:DCT

[HLC^{+}15] S-L. Hu, C-J. Liou, Y-H. Cheng, J-C. Yiu, P. P. Chiou, and Y-S. Lai. Development and characterization of two monoclonal antibodies against grouper iridovirus 55L and 97L proteins. Journal of Fish Diseases, 38(3):249–258, March 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hamre:2011:EDE

[HLHS11] L. A. Hamre, B. T. Lunestad, R. Hannisdal, and O. B. Samuelsen. An evaluation of the duration of efficacy of emamectin benzoate in the control of Caligus curtus Müller infestations in Atlantic cod, Gadus morhua L. Journal of Fish Diseases, 34(6):453–457, June 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hart:2012:EGD

[HLL^{+}12] L. M. Hart, N. Lorenzen, S. E. LaPatra, C. A. Grady, S. E. Roon, J. O’Reilly, J. L. Gregg, and P. K. Hershberger. Efficacy of a glycoprotein DNA vaccine against viral haemorrhagic septicaemia (VHS) in Pacific herring, Clupea pallasii Valenciennes. Journal of Fish Diseases, 35(10):775–779, October 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Huang:2016:SIB

[HLP^{+}16] Y-C. Huang, T-S. Lin, C. Peng, N-L. Chan, and Y-S. Han. Strong inhibition of betanodavirus replication by ribavirin targeting RNA-dependent RNA polymerase. Journal of Fish Diseases, 39(5):619–623, May 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hedrick:2013:PKP

[HMDB13] Ronald Hedrick, Christian Michel, Michel Dorson, and Jean-Francois Bernardet. Pierre de Kinkelin Pelletan (1936–2013). Journal of Fish Diseases, 36(12):975–976, December 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hwang:2012:IHR

[HMP^{+}12] S. D. Hwang, N. Midorikawa, P. Punnarak, Y. Kikuchi, H. Kondo, I. Hirono, and T. Aoki. Inhibition of hirame rhabdovirus growth by RNA aptamers. Journal of Fish Diseases, 35(12):927–934, December 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hameed:2017:SOI

[HMV^{+}17] A. S. Sahul Hameed, S. Abdul Majeed, S. Vimal, N. Madan, T. Rajkumar, S. Santhoshkumar, and S. Sivakumar. Studies on the occurrence of infectious myonecrosis virus in pondreared Litopenaeus vannamei (Boone, 1931) in India. Journal of Fish Diseases, 40(12):1823–1830, December 2017. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Huang:2011:IRT

[HMW11] Z-H. Huang, A-J. Ma, and X-A. Wang. The immune response of turbot, Scophthalmus maximus (L.), skin to high water temperature. Journal of Fish Diseases, 34(8):619–627, August 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hayakijkosol:2013:NPC

[HO13] O. Hayakijkosol and L. Owens. Non-permissive C6/ 36 cell culture for the Australian isolate of Macrobrachium rosenbergii nodavirus. Journal of Fish Diseases, 36(4):401–409, April 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Huchzermeyer:2013:CLP

[HOHG13] K. D. A. Huchzermeyer, G. Osthoff, A. Hugo, and D. Govender. Comparison of the lipid properties of healthy and pansteatitis-affected African sharptooth catfish, Clarias gariepinus (Burchell), and the role of diet in pansteatitis outbreaks in the Olifants River in the Kruger National Park, South Africa. Journal of Fish Diseases, 36(11):897–909, November 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Huber:2010:VHS

[HPAL10] P. Huber, B. Petri, S. Allen, and J. S. Lumsden. Viral haemorrhagic septicaemia virus IVb inactivation by ultraviolet light, and storage viability at 4 and −20 ^{o}C. Journal of Fish Diseases, 33(4):377–380, April 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hwang:2013:OHI

[HPY^{+}13] J. Y. Hwang, J. J. Park, H. J. Yu, Y. B. Hur, I. Arzul, Y. Couraleau, and M. A. Park. Ostreid herpesvirus 1 infection in farmed Pacific oyster larvae Crassostrea gigas (Thunberg) in Korea. Journal of Fish Diseases, 36(11):969–972, November 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hardy:2015:B

[HR15] R. W. Hardy and H. Rodger. Biography. Journal of Fish Diseases, 38(7):595–597, July 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hickey:2015:DTS

[HRL15] M. E. Hickey, G. P. Richards, and J-L. Lee. Development of a two-step, non-probed multiplex real-time PCR for surveilling Vibrio anguillarum in seawater. Journal of Fish Diseases, 38 (6):551–559, June 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hogfors-Ronnholm:2015:ASR

[HRNW15] E. Högfors-Rönnholm, J. Norrgård, and T. Wiklund. Adhesion of smooth and rough phenotypes of Flavobacterium psychrophilum to polystyrene surfaces. Journal of Fish Diseases, 38(5):429–437, May 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hellebo:2017:PSP

[HSA17] A. Hellebø, A. Stene, and V. Aspehaug. PCR survey for Paramoeba perurans in fauna, environmental samples and fish associated with marine farming sites for Atlantic salmon (Salmo salar L.). Journal of Fish Diseases, 40(5):661–670, May 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Haenen:2016:FEI

[HSC^{+}16] O. L. M. Haenen, H. Schuetze, M. Cieslak, S. Oldenburg, M. A. H. Spierenburg, I. Roozenburg-Hengst, M. Voorbergen-Laarman, M. Y. Engelsma, and N. J. Olesen. First evidence of infectious hematopoietic necrosis virus (IHNV) in the Netherlands. Journal of Fish Diseases, 39(8): 971–979, August 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hayes:2011:URC

[HSGD11] P. M. Hayes, N. J. Smit, A. S. Grutter, and A. J. Davies. Unexpected response of a captive blackeye thicklip, Hemigymnus melapterus (Bloch), from Lizard Island, Australia, exposed to juvenile isopods Gnathia aureamaculosa Ferreira & Smit. Journal of Fish Diseases, 34(7):563–566, July 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hurley-Sanders:2016:EGT

[HSHC^{+}16] J. Hurley-Sanders, C. Harms, E. Christiansen, E. Clarke I.I.I, and J. Law. Exuberant granulation tissue response associated with Neobenedenia sp. (Monogenea: Capsalidae) infestation in two cobia, Rachycentron canadum (Linnaeus). Journal of Fish Diseases, 39(3):277–283, March 2016. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hjortaas:2013:FDS

[HST^{+}13] M. J. Hjortaas, H. R. Skjelstad, T. Taksdal, A. B. Olsen, R. Johansen, B. Bang-Jensen, I. Ørpetveit, and H. Sindre. The first detections of subtype 2–related salmonid alphavirus (SAV2) in Atlantic salmon, Salmo salar L., in Norway. Journal of Fish Diseases, 36(1):71–74, January 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Haryadi:2015:TWS

[HVVV15] D. Haryadi, J. A. J. Verreth, M. C. J. Verdegem, and J. M. Vlak. Transmission of white spot syndrome virus (WSSV) from Dendronereis spp. (Peters) (Nereididae) to penaeid shrimp. Journal of Fish Diseases, 38(5):419–428, May 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hall:2013:RIP

[HWM^{+}13] M. Hall, I. S. Wallace, L. A. Munro, E. S. Munro, R. McIntosh, P. Cook, C. E. T. Allan, and A. G. Murray. Reliability of individual and pooled test procedures for detecting the pathogenic agent for clinical infectious salmon anaemia. Journal of Fish Diseases, 36(8):741–745, August 2013. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Huchzermeyer:2017:PPO

[HWO^{+}17] K. D. A. Huchzermeyer, S. Woodborne, G. Osthoff, A. Hugo, A. C. Hoffman, H. Kaiser, J. C. A. Steyl, and J. G. Myburgh. Pansteatitis in polluted Olifants River impoundments: nutritional perspectives on fish in a eutrophic lake, Lake Loskop, South Africa. Journal of Fish Diseases, 40(11):1665–1680, November 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hong:2016:IPV

[HXZ^{+}16] X. P. Hong, D. Xu, Y. Zhuo, H. Q. Liu, and L. Q. Lu. Identification and pathogenicity of Vibrio parahaemolyticus isolates and immune responses of Penaeus (Litopenaeus) vannamei (Boone). Journal of Fish Diseases, 39(9):1085–1097, September 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hoai:2016:ICL

[HY16] T. D. Hoai and T. Yoshida. Induction and characterization of a lysogenic bacteriophage of Lactococcus garvieae isolated from marine fish species. Journal of Fish Diseases, 39 (7):799–808, July 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Huang:2019:MIT

[HZJ^{+}19] Lixing Huang, Yanfei Zuo, Qingling Jiang, Yongquan Su, Yingxue Qin, Xiaojin Xu, Lingmin Zhao, and Qingpi Yan. A metabolomic investigation into the temperature-dependent virulence of Pseudomonas plecoglossicida from large yellow croaker (Pseudosciaena crocea). Journal of Fish Diseases, 42 (3):431–446, March 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hoa:2012:LNR

[HZP^{+}12] T. T. T. Hoa, M. P. Zwart, N. T. Phuong, M. C. M. de Jong, and J. M. Vlak. Low numbers of repeat units in variable number of tandem repeats (VNTR) regions of white spot syndrome virus are correlated with disease outbreaks. Journal of Fish Diseases, 35(11):817–826, November 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Huang:2013:IRD

[HZX^{+}13] B. F. Huang, L. L. Zou, J. G. Xie, Z. C. Huang, Y. W. Li, and A. X. Li. Immune responses of different species of tilapia infected with Streptococcus agalactiae. Journal of Fish Diseases, 36(8):747–752, August 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ilardi:2010:PSM

[IAR^{+}10] P. Ilardi, J. Abad, P. Rintamäki, J-F. Bernardet, and R. Avendaño-Herrera. Phenotypic, serological and molecular evidence of Chryseobacterium piscicola in farmed Atlantic salmon, Salmo salar L., in Finland. Journal of Fish Diseases, 33(2):179–181, February 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ismail:2019:CGA

[IBFL19] Siti Noor Fatimah Binti Ismail, Syarul Nataqain Baharum, Shazrul Fazry, and Chen Fei Low. Comparative genome analysis reveals a distinct influence of nucleotide composition on virus–host species-specific interaction of prawn-infecting nodavirus. Journal of Fish Diseases, 42(12):1761–1772, December 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Iregui:2016:EEP

[ICVV16] C. A. Iregui, J. Comas, G. M. Vásquez, and N. Verján. Experimental early pathogenesis of Streptococcus agalactiae infection in red tilapia Oreochromis spp. Journal of Fish Diseases, 39(2):205–215, February 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Imajoh:2015:ENC

[IFF^{+}15] M. Imajoh, H. Fujioka, K. Furusawa, K. Tamura, K. Yamasaki, S. Kurihara, J. Yamane, K. Kawai, and S. Oshima. Establishment of a new cell line susceptible to Cyprinid herpesvirus 3 (CyHV-3) and possible latency of CyHV-3 by temperature shift in the cells. Journal of Fish Diseases, 38(6): 507–514, June 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Irgang:2017:FIC

[IGLG^{+}17] R. Irgang, R. González-Luna, J. Gutiérrez, M. Poblete-Morales, V. Rojas, D. Tapia-Cammas, and R. Avendaño-Herrera. First identification and characterization of Tenacibaculum dicentrarchi isolated from Chilean red conger eel (Genypterus chilensis, Guichenot 1848). Journal of Fish Diseases, 40(12):1915–1920, December 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ito:2015:SAT

[IOO15] T. Ito, N. Oseko, and M. Ototake. Susceptibility of Amago trout, Oncorhynchus masou macrostomus (Günther) to an isolate of infectious salmon anaemia virus. Journal of Fish Diseases, 38(2):237–240, February 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Igboeli:2013:ISS

[IPW^{+}13] O. O. Igboeli, S. L. Purcell, H. Wotton, J. Poley, J. F. Burka, and M. D. Fast. Immunostimulation of Salmo salar L., and its effect on Lepeophtheirus salmonis (Krøyer) P-glycoprotein mRNA expression following subsequent emamectin benzoate exposure. Journal of Fish Diseases, 36 (3):339–351, March 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Isla:2019:MST

[ISCF^{+}19] Adolfo Isla, Mónica Saldarriaga-Córdoba, Derie E. Fuentes, Romina Albornoz, Denise Haussmann, Jorge Mancilla-Schulz, Alexis Martínez, Jaime Figueroa, Ruben Avendaño-Herrera, and Alejandro Yáñez. Multilocus sequence typing detects new Piscirickettsia salmonis hybrid genogroup in Chilean fish farms: Evidence for genetic diversity and population structure. Journal of Fish Diseases, 42(5):721– 737, May 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Imaizumi:2018:DEA

[ITKH18] K. Imaizumi, S. Tinwongger, H. Kondo, and I. Hirono. Disinfection of an EMS/ AHPND strain of Vibrio parahaemolyticus using ozone nanobubbles. Journal of Fish Diseases, 41 (4):725–727, April 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jaafar:2018:SIR

[JAJC^{+}18] R. M. Jaafar, A. Al-Jubury, J. K. Chettri, I. Dalsgaard, P. W. Kania, and K. Buchmann. Secondary immune response of rainbow trout following repeated immersion vaccination. Journal of Fish Diseases, 41(1):117–123, January 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jeney:2011:RGD

[JAR^{+}11] G. Jeney, L. Ardó, A. Rónyai, M. Bercsényi, and Z. Jeney. Resistance of genetically different common carp, Cyprinus carpio L., families against experimental bacterial challenge with Aeromonas hydrophila. Journal of Fish Diseases, 34 (1):65–70, January 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jovanovic:2011:EDL

[JBGP11] B. Jovanović, E. Baran, F. W. Goetz, and D. Palić. Effects of different lipopolysaccharide preparations on neutrophil function in the fathead minnow, Pimephales promelas Rafinesque. Journal of Fish Diseases, 34(11):877–880, November 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jensen:2015:IUM

[JBO^{+}15] L. B. Jensen, S. Boltana, A. Obach, C. McGurk, R. Waagbø, and S. MacKenzie. Investigating the underlying mechanisms of temperature-related skin diseases in Atlantic salmon, Salmo salar L., as measured by quantitative histology, skin transcriptomics and composition. Journal of Fish Diseases, 38(11):977–992, November 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jackson:2013:ILS

[JCN^{+}13] D. Jackson, D. Cotter, J. Newell, S. McEvoy, P. O’Donohoe, F. Kane, T. McDermott, S. Kelly, and A. Drumm. Impact of Lepeophtheirus salmonis infestations on migrating Atlantic salmon, Salmo salar L., smolts at eight locations in Ireland with an analysis of lice-induced marine mortality. Journal of Fish Diseases, 36(3):273–281, March 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic). See response [JCN^{+}14].

Jackson:2014:RMK

[JCN^{+}14] D. Jackson, D. Cotter, J. Newell, P. O’Donohoe, F. Kane, T. McDermott, S. Kelly, and A. Drumm. Response to M. Krkošek, C. W. Revie, B. Finstad and C. D. Todd’s comment on Jackson et al. ‘Impact of Lepeophtheirus salmonis infestations on migrating Atlantic salmon, Salmo salar L., smolts at eight locations in Ireland with an analysis of lice-induced marine mortality’. Journal of Fish Diseases, 37(4):419–421, April 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic). See [JCN^{+}13].

Jeronimo:2011:PLF

[JCS11] D. Jeronimo, C. Cruz, and A. Saraiva. Polycystic liver in farmed turbot, Scopthalmus maximus (L.). Journal of Fish Diseases, 34(12):937–938, December 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jarau:2018:VFP

[JDH^{+}18] Maureen Jarau, Adrian Di Natale, Paul E. Huber, Janet I. MacInnes, and John S. Lumsden. Virulence of Flavobacterium psychrophilum isolates in rainbow trout Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 41(10):1505–1514, October 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jagielski:2017:IIM

[JDRM17] T. Jagielski, M. Dylag, U. Roesler, and J. Murugaiyan. Isolation of infectious microalga Prototheca wickerhamii from a carp (Cyprinus carpio) — a first confirmed case report of protothecosis in a fish. Journal of Fish Diseases, 40(10): 1417–1421, October 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jaramillo:2019:HAE

[JFWH19] Diana Jaramillo, Stewart Fielder, Richard J. Whittington, and Paul Hick. Host, agent and environment interactions affecting Nervous necrosis virus infection in Australian bass Macquaria novemaculeata. Journal of Fish Diseases, 42(2): 167–180, February 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

John:2010:IIW

[JGI^{+}10] K. R. John, M. R. George, T. Iyappan, A. J. Thangarani, and M. J. P. Jeyaseelan. Indian isolates of white spot syndrome virus exhibit variations in their pathogenicity and genomic tandem repeats. Journal of Fish Diseases, 33(9):749–758, September 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jones:2013:DEB

[JHGR13] P. G. Jones, K. L. Hammell, G. Gettinby, and C. W. Revie. Detection of emamectin benzoate tolerance emergence in different life stages of sea lice, Lepeophtheirus salmonis, on farmed Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 36(3):209–220, March 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jawad:2018:VCD

[JHM^{+}18] Laith A. Jawad, Tom J. Hansen, Sonnich Meier, Jon Egil Skjæraasen, and Per Gunnar Fjelldal. Vertebral column deformity in adult wild haddock (Melanogrammus aeglefinus). Journal of Fish Diseases, 41(12):1909–1915, December 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jonassen:2017:ESC

[JHRW17] T. Jonassen, M. Hamadi, S. C. Remø, and R. Waagbø. An epidemiological study of cataracts in wild and farmed lumpfish (Cyclopterus lumpus L.) and the relation to nutrition. Journal of Fish Diseases, 40(12):1903–1914, December 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jaramillo:2017:ADN

[JHW17] D. Jaramillo, P. Hick, and R. J. Whittington. Age dependency of nervous necrosis virus infection in barramundi Lates calcarifer (Bloch). Journal of Fish Diseases, 40(8): 1089–1101, August 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jansen:2015:CMP

[JJB15] M. D. Jansen, B. Bang Jensen, and E. Brun. Clinical manifestations of pancreas disease outbreaks in Norwegian marine salmon farming — variations due to salmonid alphavirus subtype. Journal of Fish Diseases, 38(4):343–353, April 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jansen:2017:EPD

[JJM^{+}17] M. D. Jansen, B. Bang Jensen, M. F. McLoughlin, H. D. Rodger, T. Taksdal, H. Sindre, D. A. Graham, and A. Lillehaug. The epidemiology of pancreas disease in salmonid aquaculture: a summary of the current state of knowledge. Journal of Fish Diseases, 40(1):141–155, January 2017. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jose:2012:LOC

[JJS^{+}12] S. Jose, P. Jayesh, N. S. Sudheer, G. Poulose, A. Mohandas, R. Philip, and I. S. Bright Singh. Lymphoid organ cell culture system from Penaeus monodon (Fabricius) as a platform for white spot syndrome virus and shrimp immune-related gene expression. Journal of Fish Diseases, 35(5):321–334, May 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Julin:2015:PII

[JJSJ15] K. Julin, L-H. Johansen, A-I. Sommer, and J. B. Jørgensen. Persistent infections with infectious pancreatic necrosis virus (IPNV) of different virulence in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 38(11):1005–1019, November 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jung:2015:EWT

[JJV^{+}15] M. H. Jung, S. J. Jung, T. N. Vinay, C. Nikapitiya, J. O. Kim, J. H. Lee, J. Lee, and M. J. Oh. Effects of water temperature on mortality in Megalocytivirus-infected rock bream Oplegnathus fasciatus (Temminck et Schlegel) and development of protective immunity. Journal of Fish Diseases, 38 (8):729–737, August 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Janse:2012:MAG

[JK12] M. Janse and M. J. L. Kik. Mycobacterium avium granulomas in a captive epaulette shark, Hemiscyllium ocellatum (Bonnaterre). Journal of Fish Diseases, 35(12):935–940, December 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jang:2011:DAP

[JKC^{+}11] H. B. Jang, Y. R. Kim, I. S. Cha, S. W. Noh, S. B. Park, M. Ohtani, J. Hikima, T. Aoki, and T. S. Jung. Detection of antigenic proteins expressed by lymphocystis virus as vaccine candidates in olive flounder, Paralichthys olivaceus (Temminck & Schlegel). Journal of Fish Diseases, 34(7):555– 562, July 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jin:2017:CEA

[JKH^{+}17] J. W. Jin, Y. C. Kim, S. Hong, M. S. Kim, J. B. Jeong, and H. D. Jeong. Cloning and expression analysis of innate immune genes from red sea bream to assess different susceptibility to megalocytivirus infection. Journal of Fish Diseases, 40(4):583–595, April 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jaafar:2013:GMC

[JKL^{+}13] R. M. Jaafar, P. W. Kania, A. H. Larsen, D. S. Nielsen, B. Fouz, C. Browdy, and K. Buchmann. Gut microbiota changes in rainbow trout, Oncorhynchus mykiss (Walbaum), during organic acid feed supplementation and Yersinia ruckeri infection. Journal of Fish Diseases, 36(6):599–606, June 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jackson:2013:SLL

[JKO^{+}13] D. Jackson, F. Kane, P. O’Donohoe, T. Mc Dermott, S. Kelly, A. Drumm, and J. Newell. Sea lice levels on wild Atlantic salmon, Salmo salar L., returning to the coast of Ireland. Journal of Fish Diseases, 36(3):293–298, March 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jonstrup:2013:DVN

[JKS^{+}13] S. P. Jonstrup, S. Kahns, H. F. Skall, T. S. Boutrup, and N. J. Olesen. Development and validation of a novel Taqmanbased real-time RT-PCR assay suitable for demonstrating freedom from viral haemorrhagic septicaemia virus. Journal of Fish Diseases, 36(1):9–23, January 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jung:2016:LPF

[JLJ16] M. H. Jung, J. Lee, and S. J. Jung. Low pathogenicity of flounder iridovirus (FLIV) and the absence of crossprotection between FLIV and rock bream iridovirus. Journal of Fish Diseases, 39(11):1325–1333, November 2016. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jackson:2018:DSL

[JMD^{+}18] D. Jackson, O. Moberg, E. M. Stenevik Djupevåg, F. Kane, and H. Hareide. The drivers of sea lice management policies and how best to integrate them into a risk management strategy: an ecosystem approach to sea lice management. Journal of Fish Diseases, 41(6):927–933, June 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jarungsriapisit:2018:TAS

[JMF^{+}18] J. Jarungsriapisit, L. J. Moore, I. U. Fiksdal, H. Bjørgen, A. Tangerås, B. Köllner, E. O. Koppang, and S. Patel. Time after seawater transfer influences immune cell abundance and responses to SAV3 infection in Atlantic salmon. Journal of Fish Diseases, 41(8):1269–1282, August 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Janik:2018:HEE

[JMHK18] Andrew J. Janik, Douglas F. Markle, Jerry R. Heidel, and Michael L. Kent. Histopathology and external examination of heavily parasitized Lost River Sucker Deltistes luxatus (Cope 1879) and Shortnose Sucker Chasmistes brevirostris (Cope 1879) from Upper Klamath Lake, Oregon. Journal of Fish Diseases, 41(11):1675–1687, November 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jarau:2019:EFT

[JML19] Maureen Jarau, Janet I. MacInnes, and John S. Lumsden. Erythromycin and florfenicol treatment of rainbow trout Oncorhynchus mykiss (Walbaum) experimentally infected with Flavobacterium psychrophilum. Journal of Fish Diseases, 42 (3):325–334, March 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Julin:2013:SVF

[JMS13] K. Julin, S. Mennen, and A-I. Sommer. Study of virulence in field isolates of infectious pancreatic necrosis virus obtained from the northern part of Norway. Journal of Fish Diseases, 36(2):89–102, February 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jorgensen:2011:RTP

[JNN^{+}11] A. Jørgensen, A. Nylund, V. Nikolaisen, S. Alexandersen, and E. Karlsbakk. Real-time PCR detection of Parvicapsula pseudobranchicola (Myxozoa: Myxosporea) in wild salmonids in Norway. Journal of Fish Diseases, 34(5):365– 371, May 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jensen:2019:IVT

[JNS^{+}19] Britt Bang Jensen, Stian Nylund, Julie Christine Svendsen, Paul-Martin R. Ski, and Harald Takle. Indications for a vertical transmission pathway of piscine myocarditis virus in Atlantic salmon (Salmo salar L.). Journal of Fish Diseases, 42(6):825–833, June 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jensen:2019:DSS

[JOF^{+}19] Ingvill Jensen, Mathias C. Overrein, Børge N. Fredriksen, Guro Strandskog, and Tore Seternes. Differences in smolt status affect the resistance of Atlantic salmon (Salmo salar L.) against infectious pancreatic necrosis, while vaccinemediated protection is unaffected. Journal of Fish Diseases, 42(9):1271–1282, September 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jiang:2019:PDW

[JPG^{+}19] Nan Jiang, Xiaoyi Pan, Zemao Gu, Wenzhi Liu, Kaige Si, Yong Zhou, Yongze Zhou, Liwen Zhai, Yuding Fan, and Lingbing Zeng. Proliferation dynamics of WSSV in crayfish, Procambarus clarkii, and the host responses at different temperatures. Journal of Fish Diseases, 42(4):497–510, April 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jang:2014:EMI

[JQK14] I. K. Jang, G. Qiao, and S-K. Kim. Effect of multiple infections with white spot syndrome virus and Vibrio anguillarum on Pacific white shrimp Litopenaeus vannamei (L.): mortality and viral replication. Journal of Fish Diseases, 37(10): 911–920, October 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jevne:2019:HSL

[JR19] Lone Sunniva Jevne and Kjell Inge Reitan. How are the salmon lice (Lepeophtheirus salmonis Krøyer, 1837) in Atlantic salmon farming affected by different control efforts: a case study of an intensive production area with coordinated production cycles and changing delousing practices in 2013– 2018. Journal of Fish Diseases, 42(11):1573–1586, November 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jung-Schroers:2019:IMD

[JSAB^{+}19] Verena Jung-Schroers, Mikolaj Adamek, Angela Boley, Anna Korshun, and Dieter Steinhagen. Influence of a membranedenitrification reactor on the microbial community of an aquaculture recirculation system. Journal of Fish Diseases, 42(1):141–146, January 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jung-Schroers:2018:RIM

[JSAH^{+}18] V. Jung-Schroers, M. Adamek, S. Harris, H. Syakuri, A. Jung, I. Irnazarow, and D. Steinhagen. Response of the intestinal mucosal barrier of carp (Cyprinus carpio) to a bacterial challenge by Aeromonas hydrophila intubation after feeding with β-1,3/1,6-glucan. Journal of Fish Diseases, 41 (7):1077–1092, July 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jung-Schroers:2018:MID

[JSJR^{+}18] V. Jung-Schroers, A. Jung, M. Ryll, J. Bauer, F. Teitge, and D. Steinhagen. Methods for identification and differentiation of different Shewanella spp. isolates for diagnostic use. Journal of Fish Diseases, 41(4):689–714, April 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jung-Schroers:2019:DMI

[JSJR^{+}19] Verena Jung-Schroers, Arne Jung, Martin Ryll, Julia Bauer, Felix Teitge, and Dieter Steinhagen. Diagnostic methods for identifying different Aeromonas species and examining their pathogenicity factors, their correlation to cytotoxicity and adherence to fish mucus. Journal of Fish Diseases, 42(2): 189–219, February 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jonstrup:2010:ISD

[JSK^{+}10] S. P. Jonstrup, H. Schuetze, G. Kurath, T. Gray, B. Bang Jensen, and N. J. Olesen. An isolate and sequence database of infectious haematopoietic necrosis virus (IHNV). Journal of Fish Diseases, 33(6):469–471, June 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jaemwimol:2019:VEC

[JST^{+}19] Phitchaya Jaemwimol, Kwanrawee Sirikanchana, Puntanat Tattiyapong, Skorn Mongkolsuk, and Win Surachetpong. Virucidal effects of common disinfectants against tilapia lake virus. Journal of Fish Diseases, 42(10):1383–1389, October 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jansen:2010:SAS

[JTW^{+}10] M. D. Jansen, T. Taksdal, M. A. Wasmuth, B. Gjerset, E. Brun, A. B. Olsen, O. Breck, and M. Sandberg. Salmonid alphavirus (SAV) and pancreas disease (PD) in Atlantic salmon, Salmo salar L., in freshwater and seawater sites in Norway from 2006 to 2008. Journal of Fish Diseases, 33 (5):391–402, May 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jansen:2010:PDP

[JWO^{+}10] M. D. Jansen, M. A. Wasmuth, A. B. Olsen, B. Gjerset, I. Modahl, O. Breck, R. N. Haldorsen, R. Hjelmeland, and T. Taksdal. Pancreas disease (PD) in sea-reared Atlantic salmon, Salmo salar L., in Norway; a prospective, longitudinal study of disease development and agreement between diagnostic test results. Journal of Fish Diseases, 33(9):723– 736, September 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jiang:2014:DAS

[JYZ^{+}14] J-Z. Jiang, R. Yang, H. Zhang, G-F. Liu, Z-N. Zhu, Y-L. Su, E-Y. Wu, and J-Y. Wang. Detection of abalone shrivelling syndrome-associated virus using loop-mediated isothermal amplification. Journal of Fish Diseases, 37(1):63–67, January 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jencic:2014:STB

[JZK^{+}14] V. Jenčič, U. Zajc, D. Kušar, M. Ocepek, and M. Pate. A survey on Tetracapsuloides bryosalmonae infections in Slovene fresh waters. Journal of Fish Diseases, 37(8):711–717, August 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kumar:2013:EML

[KAESEM13] G. Kumar, A. Abd-Elfattah, H. Soliman, and M. El-Matbouli. Establishment of medium for laboratory cultivation and maintenance of Fredericella sultana for in vivo experiments with Tetracapsuloides bryosalmonae (Myxozoa). Journal of Fish Diseases, 36(2):81–88, February 2013. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Klakegg:2019:OAD

[KAF^{+}19] Øystein Klakegg, Takele Abayneh, Aud Kari Fauske, Michael Fülberth, and Henning Sørum. An outbreak of acute disease and mortality in Atlantic salmon (Salmo salar) post-smolts in Norway caused by Tenacibaculum dicentrarchi. Journal of Fish Diseases, 42(6):789–807, June 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kane:2018:SL

[Kan18] F. Kane. Sea lice 2016. Journal of Fish Diseases, 41(6):859– 860, June 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kabuusu:2018:PLR

[KAS^{+}18] R. M. Kabuusu, A. T. Aire, D. F. Stroup, C. N. L. Macpherson, and H. W. Ferguson. Production-level risk factors for syncytial hepatitis in farmed tilapia (Oreochromis niloticus L.). Journal of Fish Diseases, 41(1):61–66, January 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kullmann:2017:ASA

[KAST17] B. Kullmann, M. Adamek, D. Steinhagen, and R. Thiel. Anthropogenic spreading of anguillid herpesvirus 1 by stocking of infected farmed European eels, Anguilla anguilla (L.), in the Schlei fjord in northern Germany. Journal of Fish Diseases, 40(11):1695–1706, November 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kumari:2013:VAS

[KBD13] J. Kumari, J. Bøgwald, and R. A. Dalmo. Vaccination of Atlantic salmon, Salmo salar L., with Aeromonas salmonicida and infectious pancreatic necrosis virus (IPNV) showed a mixed Th1/Th2/Treg response. Journal of Fish Diseases, 36(10):881–886, October 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kiseleva:2014:ETC

[KBE^{+}14] M. Kiseleva, L. Balabanova, L. Elyakova, V. Rasskazov, and T. Zvyagintseva. Effect of treatment of chum salmon Oncorhynchus keta (Walbaum) eggs with 1,3;1,6-β-D-glucans on their development and susceptibility to Saprolegnia infection. Journal of Fish Diseases, 37(1):3–10, January 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Keeling:2013:DVR

[KBJ^{+}13] S. E. Keeling, C. L. Brosnahan, C. Johnston, R. Wallis, N. Gudkovs, and W. L. McDonald. Development and validation of a real-time PCR assay for the detection of Aeromonas salmonicida. Journal of Fish Diseases, 36(5):495–503, May 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Klinger-Bowen:2016:DPF

[KBTMH^{+}16] R. Klinger-Bowen, C. Tamaru, K. McGovern-Hopkins, Y. Li, F. Sithole, M. T. Kearney, S. Francis, and E. Soto. Dynamics of piscine francisellosis differs amongst tilapia species (Oreochromis spp.) in a controlled challenge with Francisella noatunensis subsp. orientalis. Journal of Fish Diseases, 39 (11):1305–1312, November 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kara:2014:BAM

[KCD^{+}14] H. M. Kara, L. Chaoui, F. Derbal, R. Zaidi, C. de Boisséson, M. Baud, and L. Bigarré. Betanodavirus-associated mortalities of adult wild groupers Epinephelus marginatus (Lowe) and Epinephelus costae (Steindachner) in Algeria. Journal of Fish Diseases, 37(3):273–278, March 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kua:2014:EST

[KCL14] B. C. Kua, F. C. Choong, and Y. Y. Leaw. Effect of salinity and temperature on marine leech, Zeylanicobdella arugamensis (De Silva) under laboratory conditions. Journal of Fish Diseases, 37(3):201–207, March 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kristoffersen:2018:MTC

[KDA^{+}18] Anja B. Kristoffersen, Magnus Devold, Vidar Aspehaug, Ove Gjelstenli, Olav Breck, and Britt Bang Jensen. Molecular tracing confirms that infection with infectious pancreatic necrosis virus follows the smolt from hatchery to grow-out farm. Journal of Fish Diseases, 41(10):1601–1607, October 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kulatunga:2017:FRF

[KDP^{+}17] D. C. M. Kulatunga, S. H. S. Dananjaya, B. K. Park, C-H. Kim, J. Lee, and M. De Zoysa. First report of Fusarium oxysporum species complex infection in zebrafish culturing system. Journal of Fish Diseases, 40(4):485–494, April 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kayis:2015:AHC

[KEY^{+}15]  Ş Kayiş, A. Er, C. Yilmaz, A. Düzgün, Ö. Köse, and I. Z.

Kurtoğlu. Aeromonas hydrophila as a causative agent of blue sac fry syndrome in different trout species. Journal of Fish Diseases, 38(12):1069–1071, December 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kim:2010:ESC

[KF10a] R. Kim and M. Faisal. Experimental studies confirm the wide host range of the Great Lakes viral haemorrhagic septicaemia virus genotype IVb. Journal of Fish Diseases, 33(1):83–88, January 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kim:2010:LGL

[KF10b] R. K. Kim and M. Faisal. The Laurentian Great Lakes strain (MI03) of the viral haemorrhagic septicaemia virus is highly pathogenic for juvenile muskellunge, Esox masquinongy (Mitchill). Journal of Fish Diseases, 33(6):513–527, June 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kuhn:2016:EFS

[KFK^{+}16] J. A. Kuhn, A. Frainer, R. Knudsen, R. Kristoffersen, and P-A. Amundsen. Effects of fish species composition on Diphyllobothrium spp. infections in brown trout — is threespined stickleback a key species? Journal of Fish Diseases, 39(11):1313–1323, November 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kim:2018:NIR

[KHC^{+}18] K. I. Kim, S. D. Hwang, M. Y. Cho, S. H. Jung, Y. C. Kim, and H. D. Jeong. A natural infection by the red sea bream iridovirus-type Megalocytivirus in the golden mandarin fish Siniperca scherzeri. Journal of Fish Diseases, 41 (8):1229–1233, August 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Knudsen:2017:HBA

[KHG^{+}17] R. Knudsen, E. H. Henriksen, K. Ø. Gjelland, H. Hansen, D. K. Hendrichsen, R. Kristoffersen, and K. Olstad. Are hybrids between Atlantic salmon and brown trout suitable longterm hosts of Gyrodactylus salaris during winter? Journal of Fish Diseases, 40(10):1299–1307, October 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kirkbright:2016:DLO

[KHLL16] D. Kirkbright, P. Huber, B. N. Lillie, and J. S. Lumsden. Dermocystidium-like organism linked with a mortality event in yellow perch Perca flavescens (Mitchill) in Ontario, Canada. Journal of Fish Diseases, 39(5):597–601, May 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Keller:2011:SAC

[KHS11] M. Keller, S. Han, and K. Snekvik. Severe anisakiasis and cutaneous myxoma in a California yellowtail, Seriola lalandi Valenciennes. Journal of Fish Diseases, 34(8):635–639, August 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kik:2011:SLL

[KJB11] M. J. L. Kik, M. Janse, and G. W. Benz. The sea louse Lepeophtheirus acutus (Caligidae, Siphonostomatoida, Copepoda) as a pathogen of aquarium-held elasmobranchs. Journal of Fish Diseases, 34(10):793–799, October 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Keeling:2012:DVR

[KJW^{+}12] S. E. Keeling, C. Johnston, R. Wallis, C. L. Brosnahan, N. Gudkovs, and W. L. McDonald. Development and validation of real-time PCR for the detection of Yersinia ruckeri. Journal of Fish Diseases, 35(2):119–125, February 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kong:2017:DMC

[KJW^{+}17] S. Y. Kong, Y. S. Jiang, Q. Wang, J. F. Lu, D. Xu, and L. Q. Lu. Detection methods of Cyprinid herpesvirus 2 infection in silver crucian carp (Carassius auratus gibelio) via a pORF72 monoclonal antibody. Journal of Fish Diseases, 40(12):1791– 1798, December 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kang:2015:SVH

[KKK15] B. J. Kang, K. J. Kim, and T. J. Kim. Serodiagnosis of a viral haemorrhagic septicaemia virus infection. Journal of Fish Diseases, 38(2):227–229, February 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kim:2018:HPB

[KKK^{+}18] Y. C. Kim, W. J. Kwon, M. S. Kim, K. I. Kim, J. G. Min, and H. D. Jeong. High prevalence of betanodavirus barfin flounder nervous necrosis virus as well as red-spotted grouper nervous necrosis virus genotype in shellfish. Journal of Fish Diseases, 41(2):233–246, February 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kuhn:2017:SME

[KKKA17] J. A. Kuhn, R. Knudsen, R. Kristoffersen, and P-A. Amundsen. A simplified method to estimate Diphyllobothrium spp. infection in salmonids. Journal of Fish Diseases, 40(7):863– 871, July 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kim:2019:CGS

[KKM^{+}19] Young Chul Kim, Woo Ju Kwon, Joon Gyu Min, Kwang Il Kim, and Hyun Do Jeong. Complete genome sequence and pathogenic analysis of a new betanodavirus isolated from shellfish. Journal of Fish Diseases, 42(4):519–531, April 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kim:2019:SSC

[KKOY19] Hyoung Jun Kim, Se Ryun Kwon, Niels Jørgen Olesen, and Kei Yuasa. The susceptibility of silver crucian carp (Carassius auratus langsdorfii) to infection with koi herpesvirus (KHV). Journal of Fish Diseases, 42(10):1333–1340, October 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Koiwai:2018:RMS

[KKT^{+}18a] K. Koiwai, T. Kodera, J. Thawonsuwan, M. Kawase, H. Kondo, and I. Hirono. A rapid method for simultaneously diagnosing four shrimp diseases using PCR-DNA chromatography method. Journal of Fish Diseases, 41(2): 395–399, February 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Koiwai:2018:RDT

[KKT^{+}18b] Keiichiro Koiwai, Takuya Kodera, Jumroensri Thawonsuwan, Septya Riani, Mitsuo Kawase, Hidehiro Kondo, and Ikuo Hirono. Rapid diagnosis of three shrimp RNA viruses using RT-PCR-DNA chromatography. Journal of Fish Diseases, 41(8):1309–1312, August 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kim:2018:NSI

[KLD^{+}18] Jin Do Kim, Nam-Sil Lee, Jeung Wan Do, Myoung Sug Kim, Han Gill Seo, Miyoung Cho, Sung Hee Jung, and Hyun-Ja Han. Nocardia seriolae infection in the cultured eel Anguilla japonica in Korea. Journal of Fish Diseases, 41(11):1745– 1750, November 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ku:2010:ECF

[KLW10] C-C. Ku, C-H. Lu, and C-S. Wang. Establishment and characterization of a fibroblast cell line derived from the dorsal fin of red sea bream, Pagrus major (Temminck & Schlegel). Journal of Fish Diseases, 33(3):187–196, March 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kelly:2018:FDE

[KMGK^{+}18] E. Kelly, P. A. J. Martin, S. Gibson-Kueh, D. L. Morgan, B. C. Ebner, J. Donaldson, N. Buller, D. A. Crook, S. Brooks, A. M. Davis, M. P. Hammer, L. Foyle, S. Hair, and A. J. Lymbery. First detection of Edwardsiella ictaluri (Proteobacteria: Enterobacteriaceae) in wild Australian catfish. Journal of Fish Diseases, 41(2):199–208, February 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Khang:2019:RSI

[KNN19] Pham Van Khang, Vo Van Nha, and Nguyen Hong Nguyen. Resistance to Streptococcus iniae and its genetic associations with traits of economic importance in Asian seabass (Lates calcarifer). Journal of Fish Diseases, 42(12):1657–1666, December 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Karlsen:2017:EHA

[KOB^{+}17] C. Karlsen, K. F. Ottem, Øyvind Jakobsen Brevik, M. Davey, H. Sørum, and H. C. Winther-Larsen. The environmental and host-associated bacterial microbiota of Arctic seawater-farmed Atlantic salmon with ulcerative disorders. Journal of Fish Diseases, 40(11):1645–1663, November 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kodama:2014:MIP

[KOI^{+}14] H. Kodama, K. Otani, T. Iwasaki, S. Takenaka, Y. Horitani, and H. Togase. Metabolomic investigation of pathogenesis of myxosporean emaciation disease of tiger puffer fish Takifugu rubripes. Journal of Fish Diseases, 37(7):619–627, July 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kitamura:2010:TMV

[KOS^{+}10] S-I. Kitamura, S-I. Ohtake, J-Y. Song, S-J. Jung, M-J. Oh, B-D. Choi, K. Azumi, and E. Hirose. Tunic morphology and viral surveillance in diseased Korean ascidians: soft tunic syndrome in the edible ascidian, Halocynthia roretzi (Drasche), in aquaculture. Journal of Fish Diseases, 33(2): 153–160, February 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kim:2018:FRN

[KPK^{+}18] Bo Seong Kim, Ji Won Park, Gyoung Sik Kang, Ji Hye Jin, Heyong Jin Roh, Do Hyung Kim, Mu Kun Lee, and Min Do Huh. First report of nocardia infection in cultured Japanese eel, Anguilla japonica. Journal of Fish Diseases, 41(12):1921–1927, December 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kozubikova:2010:CPP

[KPKP10] E. Kozubíková, M. Puky, P. Kiszely, and A. Petrusek. Crayfish plague pathogen in invasive North American crayfish species in Hungary. Journal of Fish Diseases, 33(11):925– 929, November 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kopp:2011:HIM

[KPM^{+}11] R. Kopp, M. Palíková, J. Mareš, S. Navrátil, Z. Kubíček, and A. Ziková. Haematological indices are modulated in juvenile carp, Cyprinus carpio L., exposed to microcystins produced by cyanobacterial water bloom. Journal of Fish Diseases, 34 (2):103–114, February 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kazazic:2019:FPI

[KPSP^{+}19] Snježana Pečur Kazazić, Natalija Topić Popović, Ivančica Strunjak-Perović, Daniela Florio, Maria Fioravanti, Sanja Babić, and Rozelindra Čož-Rakovac. Fish photobacteriosis — the importance of rapid and accurate identification of Photobacterium damselae subsp. piscicida. Journal of Fish Diseases, 42(8):1201–1209, August 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Katharios:2011:CUD

[KPT^{+}11] P. Katharios, M. Papadaki, S. Ternengo, P. K. Kantham, C. Zeri, P. E. Petraki, and P. Divanach. Chronic ulcerative dermatopathy in cultured marine fishes. Comparative study in sharpsnout sea bream, Diplodus puntazzo (Walbaum). Journal of Fish Diseases, 34(6):459–474, June 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kayansamruaj:2017:OUD

[KRD^{+}17] P. Kayansamruaj, A. Rangsichol, H. T. Dong, C. Rodkhum, M. Maita, T. Katagiri, and N. Pirarat. Outbreaks of ulcerative disease associated with ranavirus infection in barcoo grunter, Scortum barcoo (McCulloch & Waite). Journal of Fish Diseases, 40(10):1341–1350, October 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kumar:2014:OMP

[KREM14] G. Kumar, G. Rathore, and M. El-Matbouli. Outer membrane protein assembly factor YaeT (omp85) and GroEL proteins of Edwardsiella tarda are immunogenic antigens for Labeo rohita (Hamilton). Journal of Fish Diseases, 37(12): 1055–1059, December 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Krkosek:2014:CJA

[KRFT14] M. Krkošek, C. W. Revie, B. Finstad, and C. D. Todd. Comment on Jackson et al. ‘Impact of Lepeophtheirus salmonis infestations on migrating Atlantic salmon, Salmo salar L., smolts at eight locations in Ireland with an analysis of liceinduced marine mortality’. Journal of Fish Diseases, 37(4): 415–417, April 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Karthikeyan:2019:EHT

[KS19] Kesavan Karthikeyan and Raja Sudhakaran. Experimental horizontal transmission of Enterocytozoon hepatopenaei in post-larvae of whiteleg shrimp, Litopenaeus vannamei. Journal of Fish Diseases, 42(3):397–404, March 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kongtorp:2010:LEV

[KSA^{+}10] R. T. Kongtorp, A. Stene, P. A. Andreassen, V. Aspehaug, D. A. Graham, T. M. Lyngstad, A. B. Olsen, R. S. Olsen, M. Sandberg, N. Santi, C. Wallace, and O. Breck. Lack of evidence for vertical transmission of SAV 3 using gametes of Atlantic salmon, Salmo salar L., exposed by natural and experimental routes. Journal of Fish Diseases, 33(11):879– 888, November 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kumar:2014:VCM

[KSAB^{+}14] G. Kumar, M. Saleh, A-A. S. Abdel-Baki, S. Al-Quraishy, and M. El-Matbouli. In vitro cultivation model for Heterosporis saurida (Microsporidia) isolated from lizardfish, Saurida undosquamis (Richardson). Journal of Fish Diseases, 37(5):443–449, May 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kaczorek:2017:EFS

[KSM^{+}17] E. Kaczorek, J. Szarek, M. Mikiewicz, E. Terech-Majewska, P. Schulz, J. Małaczewska, R. Wójcik, and A. K. Siwicki. Effect of feed supplementation with kynurenic acid on the morphology of the liver, kidney and gills in rainbow trout (Oncorhynchus mykiss Walbaum, 1792), healthy and experimentally infected with Yersinia ruckeri. Journal of Fish Diseases, 40(7):873–884, July 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kato:2015:RSA

[KSS^{+}15] G. Kato, K. Suzuki, T. Sakai, M. Kawakami, T. Takano, T. Matsuyama, and C. Nakayasu. The role of a specific antibody against Flavobacterium psychrophilum infection in ayu sweetfish, Plecoglossus altivelis altivelis (Temminck & Schlegel, 1846). Journal of Fish Diseases, 38(1):107–112, January 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kashulin:2017:CWV

[KSS17] A. Kashulin, N. Seredkina, and H. Sørum. Cold-water vibriosis. The current status of knowledge. Journal of Fish Diseases, 40(1):119–126, January 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kataoka:2017:TDV

[KTK17] C. Kataoka, H. Tomiyama, and S. Kashiwada. Threedimensional visualization of green fluorescence proteinlabelled Edwardsiella tarda in whole Medaka larvae. Journal of Fish Diseases, 40(4):479–484, April 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Koiwai:2016:DAH

[KTN^{+}16] K. Koiwai, S. Tinwongger, R. Nozaki, H. Kondo, and I. Hirono. Detection of acute hepatopancreatic necrosis disease strain of Vibrio parahaemolyticus using loop-mediated isothermal amplification. Journal of Fish Diseases, 39(5): 603–606, May 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Keawcharoen:2015:GCB

[KTP^{+}15] J. Keawcharoen, S. Techangamsuwan, A. Ponpornpisit, E. D. Lombardini, T. Patchimasiri, and N. Pirarat. Genetic characterization of a betanodavirus isolated from a clinical disease outbreak in farm-raised tilapia Oreochromis niloticus (L.) in Thailand. Journal of Fish Diseases, 38(1):49–54, January 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kasantikul:2016:STL

[KVC^{+}16] T. Kasantikul, A. Visessripong, S. Chatrungreungkul, E. D. Lombardini, and W. Sirimanapong. Suspected thymic lymphoma with invasion of the gills in a gold crossback arowana (Scleropages formosus). Journal of Fish Diseases, 39(6):783– 786, June 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kurath:2016:ASS

[KWD^{+}16] G. Kurath, J. R. Winton, O. B. Dale, M. K. Purcell, K. Falk, and R. A. Busch. Atlantic salmon, Salmo salar L. are broadly susceptible to isolates representing the North American genogroups of infectious hematopoietic necrosis virus. Journal of Fish Diseases, 39(1):55–67, January 2016. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kent:2019:FEE

[KWGS19] Michael L. Kent, Virginia Watral, Christopher A. Gaulke, and Thomas J. Sharpton. Further evaluation of the efficacy of emamectin benzoate for treating Pseudocapillaria tomentosa (Dujardin 1843) in zebrafish Danio rerio (Hamilton 1822). Journal of Fish Diseases, 42(10):1351–1357, October 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kongrueng:2015:CVP

[KYB^{+}15] J. Kongrueng, M. Yingkajorn, S. Bunpa, N. Sermwittayawong, K. Singkhamanan, and V. Vuddhakul. Characterization of Vibrio parahaemolyticus causing acute hepatopancreatic necrosis disease in southern Thailand. Journal of Fish Diseases, 38(11):957–966, November 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kim:2012:MEA

[KYK^{+}12] H. J. Kim, J-H. Yu, D-W. Kim, S. R. Kwon, and S-W. Park. Molecular evidence of anguillid herpesvirus-1 (AngHV-1) in the farmed Japanese eel, Anguilla japonica Temminck & Schlegel, in Korea. Journal of Fish Diseases, 35(4):315– 319, April 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kusar:2017:MAF

[KZJ^{+}17] D. Kušar, U. Zajc, V. Jenčič, M. Ocepek, J. Higgins, M. Žolnir-Dovč, and M. Pate. Mycobacteria in aquarium fish: results of a 3-year survey indicate caution required in handling pet-shop fish. Journal of Fish Diseases, 40(6):773– 784, June 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Larsen:2014:EVP

[LAN^{+}14] H. A. S. Larsen, L. Austbø, A. Nødtvedt, T. W. K. Fraser, E. Rimstad, P. G. Fjelldal, T. Hansen, and E. O. Koppang. The effect of vaccination, ploidy and smolt production regime on pathological melanin depositions in muscle tissue of Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 37 (4):327–340, April 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Levican-Asenjo:2019:SCS

[LASRA^{+}19] Jorge Levicán-Asenjo, Ricardo Soto-Rifo, Francisco Aguayo, Aldo Gaggero, and Oscar Leon. Salmon cells SHK-1 internalize infectious pancreatic necrosis virus by macropinocytosis. Journal of Fish Diseases, 42(7):1035–1046, July 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Losada:2014:SDA

[LBF^{+}14] A. P. Losada, R. Bermúdez, L. D. Faílde, M. V. Ruiz de Ocenda, and M. I. Quiroga. Study of the distribution of active caspase-3-positive cells in turbot, Scophthalmus maximus (L.), enteromyxosis. Journal of Fish Diseases, 37(1): 21–32, January 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lee:2017:WEP

[LC17] Y. J. Lee and L. L. Chen. WSSV envelope protein VP51B links structural protein complexes and may mediate virus infection. Journal of Fish Diseases, 40(4):571–581, April 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lin:2015:ICL

[LCC^{+}15] H-Y. Lin, C-F. Cheng, P. P. Chiou, C-J. Liou, J-C. Yiu, and Y-S. Lai. Identification and characterization of a late gene encoded by grouper iridovirus 2L (GIV-2L). Journal of Fish Diseases, 38(10):881–890, October 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Leibowitz:2010:TDS

[LCOZ10] M. Pimenta Leibowitz, J. Kumar Chettri, R. Ofir, and D. Zilberg. Treatment development for systemic Tetrahymena sp. infection in guppies, Poecilia reticulata Peters. Journal of Fish Diseases, 33(6):473–480, June 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Longyant:2010:SDD

[LCR^{+}10] S. Longyant, K. Chaiyasittrakul, S. Rukpratanporn, P. Chaivisuthangkura, and P. Sithigorngul. Simple and direct detection of Aeromonas hydrophila infection in the goldfish, Carassius auratus (L.), by dot blotting using specific monoclonal antibodies. Journal of Fish Diseases, 33(12):973–984, December 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lagos:2017:IGI

[LCV^{+}17] F. Lagos, C. Cartes, T. Vera, D. Haussmann, and J. Figueroa. Identification of genomic islands in Chilean Piscirickettsia salmonis strains and analysis of gene expression involved in virulence. Journal of Fish Diseases, 40(10): 1321–1331, October 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Liu:2018:ASC

[LCZ^{+}18] C. Liu, O. Q. Chang, D. F. Zhang, K. B. Li, F. Wang, M. H. Lin, C. B. Shi, L. Jiang, Q. Wang, and S. M. Bergmann. Aeromonas shuberti as a cause of multi-organ necrosis in internal organs of Nile tilapia, Oreochromis niloticus. Journal of Fish Diseases, 41(10):1529–1538, October 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

LeBlanc:2019:IIN

[LDA^{+}19] Francis LeBlanc, Delphine Ditlecadet, Jean-René Arseneau, Royce Steeves, Linda Boston, Pascal Boudreau, and Nellie Gagné. Isolation and identification of a novel salmon gill poxvirus variant from Atlantic salmon in Eastern Canada. Journal of Fish Diseases, 42(2):315–318, February 2019. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Landeira-Dabarca:2014:FDC

[LDÁM14] A. Landeira-Dabarca, M. Álvarez, and P. Molist. Food deprivation causes rapid changes in the abundance and glucidic composition of the cutaneous mucous cells of Atlantic salmon Salmo salar L. Journal of Fish Diseases, 37(10): 899–909, October 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lovy:2017:BCG

[LDEH17] J. Lovy, D. Dicarlo-Emery, and J. M. Hutcheson. A bacterium with close genetic identity to Pseudomonas mandelii associated with spring fish kills in wild bluegill Lepomis macrochirus Rafinesque and pumpkinseed sunfish Lepomis gibbosus (Linnaeus). Journal of Fish Diseases, 40(12):1757– 1764, December 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Leknes:2012:BCC

[Lek12] I. L. Leknes. Blood cleansing cells in head kidney and spleen in Buenos Aires tetra, Hyphessobrycon anisitsi (Eigenmann), (Characidae: Teleostei). Journal of Fish Diseases, 35(6): 455–460, June 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Leis:2019:BPI

[LEM^{+}19] Eric Leis, Sara Erickson, Rebekah McCann, Isaac Standish, Ryan Katona, Brian Brecka, and Wes Baumgartner. Bluegill Picornavirus isolated from a mortality event involving Bluegill (Lepomis macrochirus) in the upper Mississippi River. Journal of Fish Diseases, 42(9):1233–1240, September 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Loch:2016:FSI

[LF16] T. P. Loch and M. Faisal. Flavobacterium spartansii induces pathological changes and mortality in experimentally challenged Chinook salmon Oncorhynchus tshawytscha (Walbaum). Journal of Fish Diseases, 39(4):483–488, April 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Loch:2018:FCE

[LF18] T. P. Loch and M. Faisal. Flavobacteria colonizing the early life stages of hatchery-incubated Chinook salmon Oncorhynchus tshawytscha (Walbaum 1792) are markedly diverse. Journal of Fish Diseases, 41(5):829–845, May 2018.

CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lange:2017:SSR

[LFD^{+}17] M. D. Lange, B. D. Farmer, A. M. Declercq, E. Peatman, A. Decostere, and B. H. Beck. Sickeningly sweet: Lrhamnose stimulates Flavobacterium columnare biofilm formation and virulence. Journal of Fish Diseases, 40(11):1613– 1624, November 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lievens:2011:RDI

[LFH^{+}11] B. Lievens, I. Frans, C. Heusdens, A. Justé, S. P. Jonstrup, F. Lieffrig, and K. A. Willems. Rapid detection and identification of viral and bacterial fish pathogens using a DNA array-based multiplex assay. Journal of Fish Diseases, 34 (11):861–875, November 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lovy:2019:SIC

[LFL19] Jan Lovy, Sarah E. Friend, and Nicole L. Lewis. Seasonal intestinal coccidiosis in wild bluegill Lepomis macrochirus is associated with a spring bacterial epizootic. Journal of Fish Diseases, 42(12):1697–1711, December 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lefebvre:2011:EIGa

[LFP^{+}11a] F. Lefebvre, G. Fazio, A. P. Palstra, C. Székely, and A. J. Crivelli. An evaluation of indices of gross pathology associated with the nematode Anguillicoloides crassus in eels. Journal of Fish Diseases, 34(1):31–45, January 2011. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lefebvre:2011:EIGb

[LFP^{+}11b] F. Lefebvre, G. Fazio, A. P. Palstra, C. Székely, and A. J. Crivelli. An evaluation of indices of gross pathology associated with the nematode Anguillicoloides crassus in eels. Journal of Fish Diseases, 34(5):409, May 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lazado:2018:FRS

[LFSZ18] Carlo C. Lazado, Sophie Fridman, Tamar Sinai, and Dina Zilberg. First report of Streptococcus parauberis in a cultured freshwater ornamental fish, the ram cichlid Mikrogeophagus ramirezi (Myers & Harry, 1948). Journal of Fish Diseases, 41(1):161–164, January 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Liu:2019:DQA

[LGC^{+}19] Chun Liu, Yanming M. Guo, Jizhen Z. Cao, De-Feng Zhang, Ou-Qin Chang, Kaibin Li, Fang Wang, Cun-Bin Shi, Lan Jiang, Qing Wang, and Li Lin. Detection and quantification of Aeromonas schubertii in Channa maculata by TaqMan MGB probe fluorescence real-time quantitative PCR. Journal of Fish Diseases, 42(1):109–117, January 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

LaFrentz:2017:ORP

[LGD^{+}17] B. R. LaFrentz, J. C. García, H. T. Dong, G. C. Waldbieser, C. Rodkhum, F. S. Wong, and S. F. Chang. Optimized reverse primer for 16S-RFLP analysis and genomovar assignment of Flavobacterium columnare. Journal of Fish Diseases, 40(8):1103–1108, August 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Long:2019:SOD

[LGJ19] Amy Long, Kyle A. Garver, and Simon R. M. Jones. Synergistic osmoregulatory dysfunction during salmon lice (Lepeophtheirus salmonis) and infectious hematopoietic necrosis virus co-infection in sockeye salmon (Oncorhynchus nerka) smolts. Journal of Fish Diseases, 42(6):869–882, June 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

LaPatra:2014:CRC

[LGK^{+}14] S. E. LaPatra, J. M. Groff, I. Keith, W. E. Hogans, and D. Groman. Case report: concurrent herpesviral and presumptive iridoviral infection associated with disease in cultured shortnose sturgeon, Acipenser brevirostrum (L.), from the Atlantic coast of Canada. Journal of Fish Diseases, 37 (2):141–147, February 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Loizides:2014:NTL

[LGS^{+}14] M. Loizides, A. N. Georgiou, S. Somarakis, P. E. Witten, and G. Koumoundouros. A new type of lordosis and vertebral body compression in Gilthead sea bream, Sparus aurata L.: aetiology, anatomy and consequences for survival. Journal of Fish Diseases, 37(11):949–957, November 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

LaFrentz:2019:MPG

[LGS19] Benjamin R. LaFrentz, Julio C. García, and John P. Shelley. Multiplex PCR for genotyping Flavobacterium columnare. Journal of Fish Diseases, 42(11):1531–1542, November 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lu:2012:IRS

[LHW^{+}12] A-J. Lü, X-C. Hu, Y. Wang, Q-L. Ming, and H-X. Hu. Immune response in the skin of zebrafish, Danio rerio (Hamilton), against Staphylococcus chromogenes. Journal of Fish Diseases, 35(9):699–704, September 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Liu:2015:INN

[LHW^{+}15a] X. D. Liu, J. N. Huang, S. P. Weng, X. Q. Hu, W. J. Chen, Z. D. Qin, X. X. Dong, X. L. Liu, Y. Zhou, M. Asim, W. M. Wang, J. G. He, and L. Lin. Infections of nervous necrosis virus in wild and cage-reared marine fish from South China Sea with unexpected wide host ranges. Journal of Fish Diseases, 38(6):533–540, June 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lu:2015:SIR

[LHW^{+}15b] A-J. Lü, X-C. Hu, Y. Wang, A-H. Zhu, L-L. Shen, J. Tian, Z-Z. Feng, and Z-J. Feng. Skin immune response in the zebrafish, Danio rerio (Hamilton), to Aeromonas hydrophila infection: a transcriptional profiling approach. Journal of Fish Diseases, 38(2):137–150, February 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Levipan:2019:CI

[LITCAH19a] Héctor A. Levipan, Rute Irgang, Diana Tapia-Cammas, and Ruben Avendaño-Herrera. Cover image. Journal of Fish Diseases, 42(4):i, April 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Levipan:2019:HTA

[LITCAH19b] Héctor A. Levipan, Rute Irgang, Diana Tapia-Cammas, and Ruben Avendaño-Herrera. A high-throughput analysis of biofilm formation by the fish pathogen Tenacibaculum dicentrarchi. Journal of Fish Diseases, 42(4):617–621, April 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Li:2019:ECF

[LJC^{+}19] Jianhuan Li, Peng Jia, Xueji Chen, Mingyan Lai, Fanming Jin, Wei Liu, Meisheng Yi, and Kuntong Jia. Establishment and characterization of a fin tissue cell line derived from silver pomfret, Pampus argenteus. Journal of Fish Diseases, 42 (10):1391–1399, October 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lopez-Jimena:2010:DIP

[LJGRI^{+}10] B. Lopez-Jimena, E. Garcia-Rosado, C. Infante, I. Cano, M. Manchado, D. Castro, J. J. Borrego, and M. C. Alonso. Detection of infectious pancreatic necrosis virus (IPNV) from asymptomatic redbanded seabream, Pagrus auriga Valenciennes, and common seabream, Pagrus pagrus (L.), using a non-destructive procedure. Journal of Fish Diseases, 33 (4):311–319, April 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Liu:2014:TTS

[LJHG14] Y. Liu, L. Jia, M. J. Huang, and Z. M. Gu. Thelohanellus testudineus n. sp. (Myxosporea: Bivalvulida) infecting the skin of allogynogenetic gibel carp Carassius auratus gibelio (Bloch) in China. Journal of Fish Diseases, 37(6):535–542, June 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lee:2013:EDS

[LKS^{+}13] B-J. Lee, S-S. Kim, J-W. Song, D-H. Oh, J-H. Cha, J-B. Jeong, M-S. Heo, K-W. Kim, and K-J. Lee. Effects of dietary supplementation of citrus by-products fermented with a probiotic microbe on growth performance, innate immunity and disease resistance against Edwardsiella tarda in juvenile olive flounder, Paralichthys olivaceus (Temminck & Schlegel). Journal of Fish Diseases, 36(7):617–628, July 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lewisch:2013:ONP

[LKT^{+}13] E. Lewisch, M. Kucera, R. Tappert, R. Tessadri, M. Tappert, and F. Kanz. Occurrence of nephrolithiasis in a population of longsnout seahorse, Hippocampus reidi Ginsburg, and analysis of a nephrolith. Journal of Fish Diseases, 36(2):163–167, February 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lloyd:2011:SRG

[LL11] R. Lloyd and C. Lloyd. Surgical removal of a gastric foreign body in a sand tiger shark, Carcharias taurus Rafinesque. Journal of Fish Diseases, 34(12):951–953, December 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Liu:2012:FCA

[LL12] J. Y. Liu and A. H. Li. First case of Aeromonas schubertii infection in the freshwater cultured snakehead fish, Ophiocephalus argus (Cantor), in China. Journal of Fish Diseases, 35(5):335–342, May 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

LaFrentz:2011:IIP

[LLC^{+}11] B. R. LaFrentz, S. E. LaPatra, D. R. Call, G. D. Wiens, and K. D. Cain. Identification of immunogenic proteins within distinct molecular mass fractions of Flavobacterium psychrophilum. Journal of Fish Diseases, 34(11):823–830, November 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Li:2013:PUC

[LLDL13] L. Li, S-L. Lin, L. Deng, and Z-G. Liu. Potential use of chitosan nanoparticles for oral delivery of DNA vaccine in black seabream Acanthopagrus schlegelii Bleeker to protect from Vibrio parahaemolyticus. Journal of Fish Diseases, 36 (12):987–995, December 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Langenmayer:2015:CID

[LLG^{+}15] M. C. Langenmayer, E. Lewisch, M. Gotesman, W. Hoedt, M. Schneider, M. El-Matbouli, and W. Hermanns. Cutaneous infection with Dermocystidium salmonis in cardinal tetra, Paracheirodon axelrodi (Schultz, 1956). Journal of Fish Diseases, 38(5):503–506, May 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Li:2014:CSN

[LLH^{+}14a] Y. W. Li, L. Liu, P. R. Huang, W. Fang, Z. P. Luo, H. L. Peng, Y. X. Wang, and A. X. Li. Chronic streptococcosis in Nile tilapia, Oreochromis niloticus (L.), caused by Streptococcus agalactiae. Journal of Fish Diseases, 37(8):757–763, August 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Liu:2014:OSA

[LLH^{+}14b] L. Liu, Y. W. Li, R. Z. He, X. X. Xiao, X. Zhang, Y. L. Su, J. Wang, and A. X. Li. Outbreak of Streptococcus agalactiae infection in barcoo grunter, Scortum barcoo (McCulloch & Waite), in an intensive fish farm in China. Journal of Fish Diseases, 37(12):1067–1072, December 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Liu:2012:IIP

[LLL^{+}12] Z. X. Liu, G. Y. Liu, N. Li, F. S. Xiao, H. X. Xie, and P. Nie. Identification of immunogenic proteins of Flavobacterium columnare by two-dimensional electrophoresis immunoblotting with antibacterial sera from grass carp, Ctenopharyngodon idella (Valenciennes). Journal of Fish Diseases, 35 (4):255–263, April 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Luo:2013:HNC

[LLL^{+}13] Y. Z. Luo, L. Lin, Y. Liu, Z. X. Wu, Z. M. Gu, L. J. Li, and J. F. Yuan. Haematopoietic necrosis of cultured Prussian carp, Carassius gibelio (Bloch), associated with cyprinid herpesvirus 2. Journal of Fish Diseases, 36(12):1035–1039, December 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Luo:2018:CBH

[LLL^{+}18a] X. Luo, G. Liao, C. Liu, X. Jiang, M. Lin, C. Zhao, J. Tao, and Z. Huang. Characterization of bacteriophage HN48 and its protective effects in Nile tilapia Oreochromis niloticus against Streptococcus agalactiae infections. Journal of Fish Diseases, 41(10):1477–1484, October 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Luo:2018:CI

[LLL^{+}18b] X. Luo, G. Liao, C. Liu, X. Jiang, M. Lin, C. Zhao, J. Tao, and Z. Huang. Cover image. Journal of Fish Diseases, 41 (10):i, October 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

LeBlanc:2018:VVG

[LLLG18] Francis LeBlanc, Steven Leadbeater, Mark Laflamme, and Nellie Gagné. In vivo virulence and genomic comparison of infectious Salmon Anaemia Virus isolates from Atlantic Canada. Journal of Fish Diseases, 41(9):1373–1384, September 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lloyd:2011:QPD

[LLS^{+}11] S. J. Lloyd, S. E. LaPatra, K. R. Snekvik, K. D. Cain, and D. R. Call. Quantitative PCR demonstrates a positive correlation between a Rickettsia-like organism and severity of strawberry disease lesions in rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 34(9):701–709, September 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Liu:2015:DEL

[LLWL15] G. X. Liu, L. Lin, M. Wang, and X. Q. Liu. Development and evaluation of a loop-mediated isothermal amplification assay for the detection of channel catfish virus. Journal of Fish Diseases, 38(12):1073–1076, December 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Liu:2019:DAE

[LLX^{+}19] Ling Liu, Dan-Qi Lu, Jun Xu, Heng-Li Luo, and An-Xing Li. Development of attenuated erythromycin-resistant Streptococcus agalactiae vaccine for tilapia (Oreochromis niloticus) culture. Journal of Fish Diseases, 42(5):693–701, May 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Liu:2014:MCI

[LLZW14] J. Liu, Z-Z. Liu, X-J. Zhao, and C-H. Wang. MHC class IIα alleles associated with resistance to Aeromonas hydrophila in purse red common carp, Cyprinus carpio Linnaeus. Journal of Fish Diseases, 37(6):571–575, June 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lokanathan:2016:CIR

[LMAKN16] Y. Lokanathan, A. Mohd-Adnan, B-C. Kua, and S. Nathan. Cryptocaryon irritans recombinant proteins as potential antigens for sero-surveillance of cryptocaryonosis. Journal of Fish Diseases, 39(9):1069–1083, September 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lamari:2016:IBC

[LMC^{+}16] F. Lamari, A. Mahdhi, I. Chakroun, M. A. Esteban, D. Mazurais, B. Amina, and F-J. Gatesoupe. Interactions between candidate probiotics and the immune and antioxidative responses of European sea bass (Dicentrarchus labrax) larvae. Journal of Fish Diseases, 39(12):1421–1432, December 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Laurin:2019:BLC

[LMG^{+}19] Emilie Laurin, Diane Morrison, Ian A. Gardner, Ahmed Siah, James F. F. Powell, and Mykolas Kamaitis. Bayesian latent class analysis of ELISA and RT-rPCR diagnostic accuracy for subclinical Renibacterium salmoninarum infection in Atlantic salmon (Salmo salar) broodstock. Journal of Fish Diseases, 42(2):303–313, February 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Larrat:2012:SEE

[LML12] S. Larrat, J. Marvin, and S. Lair. Safety and efficacy of emamectin benzoate to treat Anguillicoloides crassus (Kuwahara, Niimi & Itagaki) infections in American eels, Anguilla rostrata (Lesueur). Journal of Fish Diseases, 35 (6):467–470, June 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Li:2013:DMP

[LML^{+}13] G. Y. Li, Z. L. Mo, J. Li, P. Xiao, B. Hao, and Y. H. Guo. Development of a multiplex PCR for the identification of pathogenic Edwardsiella tarda and application to edwardsiellosis diagnostics. Journal of Fish Diseases, 36(2):151–157, February 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Low:2015:NIR

[LMM^{+}15] C-F. Low, N. S. Mariana, A. Maha, H-Y. Chee, and M. Y. Fatimah. Non-immune-related genes and signalling pathways in spleen of Vibrio parahaemolyticus-infected Epinephelus fuscoguttatus (Forsskål). Journal of Fish Diseases, 38(8): 761–764, August 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Low:2016:IIR

[LMM^{+}16] C-F. Low, N. S. Mariana, A. Maha, H-Y. Chee, and M. Y. Fatimah. Identification of immune response-related genes and signalling pathways in spleen of Vibrio parahaemolyticus-infected Epinephelus fuscoguttatus (Forsskål) by nextgeneration sequencing. Journal of Fish Diseases, 39(3):389– 394, March 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Low:2017:BDV

[LNC^{+}17] C-F. Low, B. Syarul Nataqain, H-Y. Chee, M. Z. H. Rozaini, and M. Najiah. Betanodavirus: Dissection of the viral life cycle. Journal of Fish Diseases, 40(11):1489–1496, November 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Loh:2013:FP

[Loh13] Richmond Loh. Fish pathology. Journal of Fish Diseases, 36(6):607, June 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lari:2017:GSI

[LP17] E. Lari and G. G. Pyle. Gyrodactylus salmonis infection impairs the olfactory system of rainbow trout. Journal of Fish Diseases, 40(10):1279–1284, October 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

LaPatra:2010:EVA

[LPA^{+}10] S. E. LaPatra, K. P. Plant, S. Alcorn, V. Ostland, and J. Winton. An experimental vaccine against Aeromonas hydrophila can induce protection in rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 33 (2):143–151, February 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lee:2018:ECS

[LPC^{+}18] J-H. Lee, J-J. Park, J-H. Choi, S-Y. Kang, Y-J. Kang, and K. H. Park. Effects of clioquinol on the scuticociliatosiscausing protozoan Miamiensis avidus in olive flounder Paralichthys olivaceus. Journal of Fish Diseases, 41(3):451–462, March 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Long:2012:VDA

[LPCC12] A. Long, M. P. Polinski, D. R. Call, and K. D. Cain. Validation of diagnostic assays to screen broodstock for Flavobacterium psychrophilum infections. Journal of Fish Diseases, 35(6):407–419, June 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lopez:2010:FIT

[LPVGL^{+}10] J. R. López, M. Piñeiro-Vidal, N. García-Lamas, R. De La Herran, J. I. Navas, I. Hachero-Cruzado, and Y. Santos. First isolation of Tenacibaculum soleae from diseased cultured wedge sole, Dicologoglossa cuneata (Moreau), and brill, Scophthalmus rhombus (L.). Journal of Fish Diseases, 33 (3):273–278, March 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Li:2010:BCP

[LQL^{+}10] H. Li, G. Qiao, Q. Li, W. Zhou, K. M. Won, D-H. Xu, and S-I. Park. Biological characteristics and pathogenicity of a highly pathogenic Shewanella marisflavi infecting sea cucumber, Apostichopus japonicus. Journal of Fish Diseases, 33(11):865–877, November 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Liu:2011:SPP

[LQY11] W. Liu, D. Qian, and X. J. Yan. Studies on pathogenicity and prevalence of white spot syndrome virus in mud crab, Scylla serrata (Forsskål), in Zhejiang Province, China. Journal of Fish Diseases, 34(2):131–138, February 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Llarena-Reino:2017:MXA

[LRAP17] M. Llarena-Reino, E. Abollo, and S. Pascual. Microsporidians xenomas of anglerfish from NE Atlantic waters. Journal of Fish Diseases, 40(11):1587–1598, November 2017. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Loch:2017:NHR

[LRB^{+}17] T. P. Loch, T. G. Rosser, W. A. Baumgartner, T. Boontai, M. Faisal, and M. J. Griffin. New host record and molecular characterization of Dicauda atherinoidi Hoffman & Walker (Bivalvulida: Myxobolidae): a parasite of the emerald shiner Notropis atherinoides Rafinesque, 1818 and mimic shiner Notropis vollucellus Cope, 1865. Journal of Fish Diseases, 40(10):1405–1415, October 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lopez:2018:HIM

[LRC^{+}18] Vladimir López, María Angeles Risalde, Marinela Contreras, Lourdes Mateos-Hernández, Joaquin Vicente, Christian Gortázar, and José de la Fuente. Heat-inactivated Mycobacterium bovis protects zebrafish against mycobacteriosis. Journal of Fish Diseases, 41(10):1515–1528, October 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Li:2019:EAC

[LRD^{+}19] Chen Li, Yanbei Ren, Xuan Dong, Chongming Wang, and Jie Huang. Extraction of assembling complexes of viral capsomers from shrimp tissue infected with yellow head virus genotype 8 (YHV-8). Journal of Fish Diseases, 42(4):613– 616, April 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Low:2017:CKM

[LRMN17] C-F. Low, M. Z. H. Rozaini, N. Musa, and B. Syarul Nataqain. Current knowledge of metabolomic approach in infectious fish disease studies. Journal of Fish Diseases, 40 (10):1267–1277, October 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Labarca:2017:ALF

[LRS^{+}17] C. Cobo Labarca, J. Radinger, V. Schöning, R. Ariav, R. Jung, K. D. Thompson, W. Kloas, and K. Knopf. Application of low-frequency sonophoresis and reduction of antibiotics in the aquatic systems. Journal of Fish Diseases, 40(11):1635–1643, November 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Low:2015:EIB

[LSA^{+}15] C-F. Low, M. Nor Shamsudin, M. Abdullah, H-Y. Chee, and M. Aliyu-Paiko. Experimental infection of brown-marbled grouper, Epinephelus fuscoguttatus (Forsskål), with Vibrio parahaemolyticus identifies parvalbumin beta-2 subunit I, alpha-2-macroglobulin, nattectin and immunoglobulin light chain, differentially expressed in resistant grouper. Journal of Fish Diseases, 38(1):17–25, January 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Low:2014:PAN

[LSC^{+}14] C-F. Low, M. N. Shamsudin, H-Y. Chee, M. Aliyu-Paiko, and E. S. Idrus. Putative apolipoprotein A-I, natural killer cell enhancement factor and lysozyme g are involved in the early immune response of brown-marbled grouper, Epinephelus fuscoguttatus, Forsskål, to Vibrio alginolyticus. Journal of Fish Diseases, 37(8):693–701, August 2014. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Loch:2012:BIC

[LST^{+}12] T. P. Loch, K. Scribner, R. Tempelman, G. Whelan, and M. Faisal. Bacterial infections of Chinook salmon, Oncorhynchus tshawytscha (Walbaum), returning to gamete collecting weirs in Michigan. Journal of Fish Diseases, 35 (1):39–50, January 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Leach:2014:HCT

[LTC^{+}14] C. A. Leach, M. W. Thomas, A. J. Cooley, G. C. Waldbieser, B. G. Bosworth, E. L. Torrans, and L. H. Khoo. Histological and computed tomographic evaluation of a parasitic conjoined twin in hybrid catfish (Ictalurus punctatus [Rafinesque] × Ictalurus furcatus [Lesueur]). Journal of Fish Diseases, 37(6):521–526, June 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lima:2016:ICV

[LTC16] P. C. Lima, R. S. Taylor, and M. Cook. Involvement of contractile vacuoles in the osmoregulation process of the marine parasitic amoeba Neoparamoeba perurans. Journal of Fish Diseases, 39(5):629–633, May 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lima:2017:PFM

[LTC17] P. C. Lima, R. S. Taylor, and M. Cook. Pseudocyst formation in the marine parasitic amoeba Neoparamoeba perurans: a short-term survival strategy to abrupt salinity variation. Journal of Fish Diseases, 40(8):1109–1113, August 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Laurin:2018:DSE

[LTG^{+}18] E. Laurin, K. K. Thakur, I. A. Gardner, P. Hick, N. J. G. Moody, M. S. J. Crane, and I. Ernst. Design standards for experimental and field studies to evaluate diagnostic accuracy of tests for infectious diseases in aquatic animals. Journal of Fish Diseases, 41(5):729–749, May 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Laurin:2019:PPG

[LTM^{+}19] Emilie Laurin, Krishna Thakur, Peter G. Mohr, Paul Hick, Mark St. J. Crane, Ian A. Gardner, Nicholas J. G. Moody, Axel Colling, and Ingo Ernst. To pool or not to pool? Guidelines for pooling samples for use in surveillance testing of infectious diseases in aquatic animals. Journal of Fish Diseases, 42(11):1471–1491, November 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lewisch:2018:DPB

[LUP^{+}18] Eva Lewisch, Guenther Unfer, Kurt Pinter, Thomas Bechter, and Mansour El-Matbouli. Distribution and prevalence of T. bryosalmonae in Austria: a first survey of trout from rivers with a shrinking population. Journal of Fish Diseases, 41 (10):1549–1557, October 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lopez-Vazquez:2017:VSV

[LVADB17] C. López-Vázquez, M. C. Alonso, C. P. Dopazo, and I. Bandín. In vivo study of viral haemorrhagic septicaemia virus and infectious pancreatic necrosis virus coexistence in Senegalese sole (Solea senegalensis). Journal of Fish Diseases, 40(9):1129–1139, September 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lukhwareni:2018:HUH

[LvD18] Refilwe Lukhwareni and Cobus van Dyk. Histology and ultrastructure of hepatic nodular alterations in Clarias gariepinus (Burchell, 1822). Journal of Fish Diseases, 41(12):1859– 1870, December 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ling:2016:AEP

[LWJ^{+}16] F. Ling, Z-Q. Wu, C. Jiang, L. Liu, and G-X. Wang. Antibacterial efficacy and pharmacokinetic evaluation of sanguinarine in common carp (Cyprinus carpio) following a single intraperitoneal administration. Journal of Fish Diseases, 39(8):993–1000, August 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lumsden:2010:GDA

[LWMT10] J. S. Lumsden, B. Wybourne, M. Minamikawa, and L. Tubbs. Gastric dilation and air sacculitis in Chinook salmon, Oncorhynchus tshawytscha (Walbaum); correlation of macroscopic and microscopic lesions, and relationship of the syndrome to glomerulonephritis and serum biochemistry. Journal of Fish Diseases, 33(9):737–747, September 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

LaFrentz:2014:IHR

[LWWS14] B. R. LaFrentz, G. C. Waldbieser, T. J. Welch, and C. A. Shoemaker. Intragenomic heterogeneity in the 16S rRNA genes of Flavobacterium columnare and standard protocol for genomovar assignment. Journal of Fish Diseases, 37(7):657– 669, July 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Liu:2018:CPG

[LWZ^{+}18] Hui Liu, Zhihao Wu, Xiangping Zhu, Zongcheng Song, Jinwei Hu, Lijuan Wang, Jun Li, and Feng You. Comparative performance of growth, vertebral structure and muscle composition in diploid and triploid Paralichthys olivaceus. Journal of Fish Diseases, 41(10):1495–1504, October 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lu:2011:PGC

[LXHL11] L. Lu, H. Xu, Y. He, and J. Li. Protection of grass carp, Ctenopharyngon idellus (Valenciennes), through oral administration of a subunit vaccine against reovirus. Journal of Fish Diseases, 34(12):939–942, December 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Liu:2018:OMM

[LXL^{+}18] X. H. Liu, L. W. Xu, D. Luo, Y. L. Zhao, Q. Q. Zhang, G. F. Liu, and J. Y. Zhang. Outbreak of mass mortality of yearling groupers of Epinephelus (Perciformes, Serranidae) associated with the infection of a suspected new enteric Sphaerospora (Myxozoa: Myxosporea) species in South China Sea. Journal of Fish Diseases, 41(4):663–672, April 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Liu:2017:DCG

[LXX^{+}17] H. Liu, Y. Xiu, Y. Xu, M. Tang, S. Li, W. Gu, Q. Meng, and W. Wang. Development of a colloidal gold immunochromatographic assay (GICA) for the rapid detection of Spiroplasma eriocheiris in commercially exploited crustaceans from China. Journal of Fish Diseases, 40(12):1839–1847, December 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Luo:2019:GCK

[LXZ^{+}19] Gang Luo, Xiaojin Xu, Lingmin Zhao, Yingxue Qin, Lixing Huang, Yongquan Su, and Qingpi Yan. clpV is a key virulence gene during in vivo Pseudomonas plecoglossicida infection. Journal of Fish Diseases, 42(7):991–1000, July 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

LePage:2015:DCY

[LYD^{+}15] V. LePage, J. Young, C. J. Dutton, G. Crawshaw, J. A. Paré, M. Kummrow, D. J. McLelland, P. Huber, K. Young, S. Russell, L. Al-Hussinee, and J. S. Lumsden. Diseases of captive yellow seahorse Hippocampus kuda Bleeker, pot-bellied seahorse Hippocampus abdominalis Lesson and weedy seadragon Phyllopteryx taeniolatus (Lacépède). Journal of Fish Diseases, 38(5):439–450, May 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Low:2018:MBM

[LYKN18] Chen-Fei Low, Mohd Radzi Md Yusoff, Giva Kuppusamy, and Nur Farahiyah Ahmad Nadzri. Molecular biology of Macrobrachium rosenbergii nodavirus infection in giant freshwater prawn. Journal of Fish Diseases, 41(12):1771– 1781, December 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Li:2018:FIN

[LYL^{+}18] P. Li, Q. Yu, F. Li, X. Qin, D. Dong, B. Chen, and Q. Qin. First identification of the nervous necrosis virus isolated from cultured golden pompano (Trachinotus ovatus) in Guangxi, China. Journal of Fish Diseases, 41(7):1177–1180, July 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Li:2016:NPI

[LZGZ16] L. Li, W-T. Zhao, Y-N. Guo, and L-P. Zhang. Nematode parasites infecting the starry batfish Halieutaea stellata (Vahl) (Lophiiformes: Ogcocephalidae) from the East and South China Sea. Journal of Fish Diseases, 39(5):515–529, May 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Li:2010:DGB

[LZZN10] N. Li, J. Zhang, L. Q. Zhang, and P. Nie. Difference in genes between a high virulence strain G_{4} and a low virulence strain G_{18} of Flavobacterium columnare by using suppression subtractive hybridization. Journal of Fish Diseases, 33(5):403– 412, May 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mohammed:2016:PEN

[MA16] H. H. Mohammed and C. R. Arias. Protective efficacy of Nigella sativa seeds and oil against columnaris disease in fishes. Journal of Fish Diseases, 39(6):693–703, June 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

McBeath:2015:ICL

[MAC^{+}15] A. McBeath, M. Aamelfot, D. H. Christiansen, I. Matejusova, T. Markussen, M. Kaldhusdal, O. B. Dale, S. C. Weli, and K. Falk. Immersion challenge with low and highly virulent infectious salmon anaemia virus reveals different pathogenesis in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 38(1):3–15, January 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Molnar:2018:HCG

[MAOS^{+}18] K. Molnár, A. Avenant-Oldewage, B. Sellyei, Á. Varga, and C. Székely. Histopathological changes on the gills of asp (Aspius aspius) and European catfish (Silurus glanis) caused by Lamproglena pulchella and a Lamproglena sp. (Copepoda: Lernaeidae), respectively. Journal of Fish Diseases, 41(1): 33–39, January 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Makrinos:2017:GCI

[MB17] D. L. Makrinos and T. J. Bowden. Growth characteristics of the intracellular pathogen, Piscirickettsia salmonis, in tissue culture and cell-free media. Journal of Fish Diseases, 40(8): 1115–1127, August 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Macpherson:2012:ALE

[MBB12] H. L. Macpherson, Ø. Bergh, and T. H. Birkbeck. An aerolysin-like enterotoxin from Vibrio splendidus may be involved in intestinal tract damage and mortalities in turbot, Scophthalmus maximus (L.), and cod, Gadus morhua L., larvae. Journal of Fish Diseases, 35(2):153–167, February 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Manera:2016:CME

[MBD16] M. Manera, C. Borreca, and B. S. Dezfuli. Cutaneous myxidiosis in European eel, Anguilla anguilla (Linnaeus, 1758): histopathology, histochemistry and laminin immunohistochemistry. Journal of Fish Diseases, 39(7):845–851, July 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ma:2019:CIR

[MBOC19] Jie Ma, Timothy J. Bruce, Luke P. Oliver, and Kenneth D. Cain. Co-infection of rainbow trout (Oncorhynchus mykiss) with infectious hematopoietic necrosis virus and Flavobacterium psychrophilum. Journal of Fish Diseases, 42(7):1065– 1076, July 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Matras:2017:CEV

[MBS^{+}17] M. Matras, E. Borzym, D. Stone, K. Way, M. Stachnik, J. Maj-Paluch, M. Palusińska, and M. Reichert. Carp edema virus in Polish aquaculture — evidence of significant sequence divergence and a new lineage in common carp Cyprinus carpio (L.). Journal of Fish Diseases, 40(3):319–325, March 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ma:2019:ACP

[MBS^{+}19a] Jie Ma, Timothy J. Bruce, Ponnerassery S. Sudheesh, Christopher Knupp, Thomas P. Loch, Mohamed Faisal, and Kenneth D. Cain. Assessment of cross-protection to heterologous strains of Flavobacterium psychrophilum following vaccination with a live-attenuated coldwater disease immersion vaccine. Journal of Fish Diseases, 42(1):75–84, January 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ma:2019:FC

[MBS^{+}19b] Jie Ma, Timothy J. Bruce, Ponnerassery S. Sudheesh, Christopher Knupp, Thomas P. Loch, Mohamed Faisal, and Kenneth D. Cain. Front cover. Journal of Fish Diseases, 42(1):i, January 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Monaghan:2017:UAS

[MBT^{+}17] S. J. Monaghan, S. M. Bergmann, K. D. Thompson, L. Brown, T. Herath, J. del Pozo, and A. Adams. Ultrastructural analysis of sequential cyprinid herpesvirus 3 morphogenesis in vitro. Journal of Fish Diseases, 40(8):1041–1054, August 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mouton:2014:FRA

[MCCN14] A. Mouton, P. Crosbie, K. Cadoret, and B. Nowak. First record of amoebic gill disease caused by Neoparamoeba perurans in South Africa. Journal of Fish Diseases, 37(4):407– 409, April 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Munro:2011:GAV

[MCO11] J. Munro, R. Callinan, and L. Owens. Gill-associated virus and its association with decreased production of Penaeus monodon in Australian prawn farms. Journal of Fish Diseases, 34(1):13–20, January 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mateo:2017:BHS

[MCSM17] E. C. Mateo, T. Castro, V. Sierralta, and D. R. Mateo. Bacteriological and histopathological study of farmed rainbow trout, Oncorhynchus mykiss, in Peru. Journal of Fish Diseases, 40(7):941–946, July 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mugimba:2018:DTL

[MCW^{+}18] K. K. Mugimba, A. A. Chengula, S. Wamala, E. D. Mwega, C. J. Kasanga, D. K. Byarugaba, R. H. Mdegela, S. Tal, B. Bornstein, A. Dishon, S. Mutoloki, L. David, Ø. Evensen, and H. M. Munang’andu. Detection of tilapia lake virus (TiLV) infection by PCR in farmed and wild Nile tilapia (Oreochromis niloticus) from Lake Victoria. Journal of Fish Diseases, 41(8):1181–1189, August 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Manire:2013:LCB

[MCWL13] C. A. Manire, A. C. Clarke, D. Wert, and J. Landolfi. Lymphosarcoma in a captive bonnethead shark, Sphyrna tiburo (L.). Journal of Fish Diseases, 36(4):437–440, April 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mello:2018:FEV

[MDC^{+}18] D. F. Mello, N. M. Danielli, F. Curbani, V. A. Pontinha, S. Suhnel, M. A. M. Castro, S. C. Medeiros, N. C. Wendt, R. Trevisan, A. R. M. Magalhães, and A. L. Dafre. First evidence of viral and bacterial oyster pathogens in the Brazilian coast. Journal of Fish Diseases, 41(3):559–563, March 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mallo:2017:CAA

[MDF^{+}17] N. Mallo, A. P. DeFelipe, I. Folgueira, R. A. Sueiro, J. Lamas, and J. M. Leiro. Combined antiparasitic and anti-inflammatory effects of the natural polyphenol curcumin on turbot scuticociliatosis. Journal of Fish Diseases, 40(2): 205–217, February 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Magro:2016:FRD

[MdOO16] N. M. Magro, A. T. de Oliveira, and L. H. O’Dwyer. First report and description of a Cyrilia sp. (Apicomplexa: Haemogregarinidae) from a freshwater Cururu Stingray Potamotrygon cf. histrix (Elasmobranchii: Potamotrygonidae), from the Amazon Region, Brazil. Journal of Fish Diseases, 39(8):907–911, August 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Marton:2012:SDS

[ME12] S. Marton and E. Eszterbauer. The susceptibility of diverse species of cultured oligochaetes to the fish parasite Myxobolus pseudodispar Gorbunova (Myxozoa). Journal of Fish Diseases, 35(4):303–314, April 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Millard:2012:DNA

[MF12] E. V. Millard and M. Faisal. Development of neutralizing antibody responses in muskellunge, Esox masquinongy (Mitchill), experimentally exposed to viral haemorrhagic septicaemia virus (genotype IVb). Journal of Fish Diseases, 35(1):11–18, January 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

McConville:2018:FEL

[MFES18] James McConville, Elena Fringuelli, Derek Evans, and Paul Savage. First examination of the Lough Neagh European eel (Anguilla anguilla) population for eel virus European, eel virus European X and Anguillid Herpesvirus-1 infection by employing novel molecular techniques. Journal of Fish Diseases, 41(12):1783–1791, December 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Marshall:2018:ICT

[MFHHG18] S. H. Marshall, P. Flores-Herrera, F. A. Henríquez, and F. A. Gómez. Identification and characterization of two variants of the Hfq-sRNA-chaperone in the fish pathogen Piscirickettsia salmonis. Journal of Fish Diseases, 41(3):501–509, March 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Matsuyama:2017:CSI

[MFS^{+}17] T. Matsuyama, Y. Fukuda, T. Sakai, N. Tanimoto, M. Nakanishi, Y. Nakamura, T. Takano, and C. Nakayasu. Clonal structure in Ichthyobacterium seriolicida, the causative agent of bacterial haemolytic jaundice in yellowtail, Seriola quinqueradiata, inferred from molecular epidemiological analysis. Journal of Fish Diseases, 40(8):1065–1075, August 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Monaco:2017:ACI

[MGF17] A. Monaco, M. C. Grimaldi, and I. Ferrandino. Aluminium chloride-induced toxicity in zebrafish larvae. Journal of Fish Diseases, 40(5):629–635, May 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mo:2014:PAS

[MGH^{+}14] T. A. Mo, A. Gahr, H. Hansen, E. Hoel, Ø. Oaland, and T. T. Poppe. Presence of Anisakis simplex (Rudolphi, 1809 det. Krabbe, 1878) and Hysterothylacium aduncum (Rudolphi, 1802) (Nematoda; Anisakidae) in runts of farmed Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 37(2): 135–140, February 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Molnar:2016:MGF

[MGM^{+}16] K. Molnár, D. I. Gibson, G. Majoros, C. Székely, D. Sándor, and G. Cech. Malformations of the gill filaments of the ruffe Gymnocephalus cernuus (L.) (Pisces) caused by echinostomatid metacercariae. Journal of Fish Diseases, 39(11):1357– 1367, November 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Marin:2018:RCR

[MGM^{+}18] S. L. Marín, M. P. González, S. T. Madariaga, M. Mancilla, and J. Mancilla. Response of Caligus rogercresseyi (Boxshall & Bravo, 2000) to treatment with Hydrogen Peroxide: Recovery of parasites, fish infestation and egg viability under experimental conditions. Journal of Fish Diseases, 41(6): 861–873, June 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Misk:2016:PSV

[MGN^{+}16] E. Misk, K. Garver, E. Nagy, S. Isaac, L. Tubbs, P. Huber, L. Al-Hussinee, and J. S. Lumsden. Pathogenesis of spring viremia of carp virus in emerald shiner Notropis atherinoides Rafinesque, fathead minnow Pimephales promelas Rafinesque and white sucker Catostomus commersonii (Lacépède). Journal of Fish Diseases, 39(6):729–739, June 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mitomi:2018:FIW

[MHNY18] K. Mitomi, T. D. Hoai, I. Nishiki, and T. Yoshida. First isolation of Weissella ceti responsible for outbreaks of weissellosis in farmed rainbow trout in Japan. Journal of Fish Diseases, 41(5):847–850, May 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Meza:2019:DPA

[MID^{+}19a] Karla Meza, Makoto Inami, Alf S. Dalum, Ane M. Bjelland, Henning Sørum, and Marie Løvoll. Development of piscirickettsiosis in Atlantic salmon (Salmo salar L.) smolts after intraperitoneal and cohabitant challenge using an EM90-like isolate: a comparative study. Journal of Fish Diseases, 42 (7):1001–1011, July 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Meza:2019:CEE

[MID^{+}19b] Karla Meza, Makoto Inami, Alf S. Dalum, Hege Lund, Ane M. Bjelland, Henning Sørum, and Marie Løvoll. Comparative evaluation of experimental challenge by intraperitoneal injection and cohabitation of Atlantic salmon (Salmo salar L.) after vaccination against Piscirickettsia salmonis (EM90-like). Journal of Fish Diseases, 42(12):1713–1730, December 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Madhun:2018:PPO

[MIO^{+}18] A. S. Madhun, C. H. Isachsen, L. M. Omdal, A. C. B. Einen, S. Mæhle, V. Wennevik, E. Niemelä, T. Svåsand, and E. Karlsbakk. Prevalence of piscine orthoreovirus and salmonid alphavirus in sea-caught returning adult Atlantic salmon (Salmo salar L.) in northern Norway. Journal of Fish Diseases, 41(5):797–803, May 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mo:2017:SDP

[MJ17] T. A. Mo and A. Jørgensen. A survey of the distribution of the PKD-parasite Tetracapsuloides bryosalmonae (Cnidaria: Myxozoa: Malacosporea) in salmonids in Norwegian rivers — additional information gleaned from formerly collected fish. Journal of Fish Diseases, 40(5):621–627, May 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Miwa:2011:EAM

[MKH^{+}11] S. Miwa, T. Kamaishi, T. Hirae, T. Murase, and T. Nishioka. Encephalomyelitis associated with microsporidian infection in farmed greater amberjack, Seriola dumerili (Risso). Journal of Fish Diseases, 34(12):901–910, December 2011. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Madhun:2015:PDI

[MKI^{+}15] A. S. Madhun, E. Karlsbakk, C. H. Isachsen, L. M. Omdal, A. G. Eide Sørvik, Ø. Skaala, B. T. Barlaup, and K. A. Glover. Potential disease interaction reinforced: doublevirus-infected escaped farmed Atlantic salmon, Salmo salar L., recaptured in a nearby river. Journal of Fish Diseases, 38 (2):209–219, February 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mihok:2011:IPP

[MKK^{+}11] T. Mihok, P. Košuth, A. Kočišová, L. Pekárik, E. Bártová, and P. Major. The intestinal parasite Pseudocapillaria tomentosa (Dujardin, 1843) of the invasive fish species topmouth gudgeon, Pseudorasbora parva (Temminck and Schlegel), in Slovakia. Journal of Fish Diseases, 34(9):711– 714, September 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Murcia:2015:MCH

[MKKM15] S. Murcia, B. L. Kerans, T. M. Koel, and E. MacConnell. Myxobolus cerebralis (Hofer) infection risk in native cutthroat trout Oncorhynchus clarkii (Richardson) and its relationships to tributary environments in the Yellowstone Lake Basin. Journal of Fish Diseases, 38(7):637–652, July 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mohi:2010:HSV

[MKMY10] M. M. Mohi, M. Kuratani, T. Miyazaki, and T. Yoshida. Histopathological studies on Vibrio harveyi — infected tiger puffer, Takifugu rubripes (Temminck et Schlegel), cultured in Japan. Journal of Fish Diseases, 33(10):833–840, October 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Miwa:2015:PAC

[MKY^{+}15] S. Miwa, I. Kiryu, K. Yuasa, T. Ito, and T. Kaneko. Pathogenesis of acute and chronic diseases caused by cyprinid herpesvirus-3. Journal of Fish Diseases, 38(8):695–712, August 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Menanteau-Ledouble:2016:AAS

[MLEM16] S. Menanteau-Ledouble and M. El-Matbouli. Antigens of Aeromonas salmonicida subsp. salmonicida specifically induced in vivo in Oncorhynchus mykiss. Journal of Fish Diseases, 39(8):1015–1019, August 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Marcos-Lopez:2018:OSA

[MLER^{+}18] M. Marcos-López, C. Ruiz Espinosa, H. D. Rodger, I. O’Connor, E. MacCarthy, and M. A. Esteban. Oxidative stress is associated with late-stage amoebic gill disease in farmed Atlantic salmon (Salmo salar L.). Journal of Fish Diseases, 41(2):383–387, February 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Martens:2010:DFA

[MLF^{+}10] L. Gil Martens, E. J. Lock, P. G. Fjelldal, A. Wargelius, P. Araujo, B. E. Torstensen, P. E. Witten, T. Hansen, R. Waagbø, and R. Ørnsrud. Dietary fatty acids and inflammation in the vertebral column of Atlantic salmon, Salmo salar L., smolts: a possible link to spinal deformities. Journal of Fish Diseases, 33(12):957–972, December 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ma:2015:ACU

[MLH^{+}15] Y. P. Ma, Z. X. Liu, L. Hao, J. Y. Ma, Z. L. Liang, Y. G. Li, and H. Ke. Analysing codon usage bias of cyprinid herpesvirus 3 and adaptation of this virus to the hosts. Journal of Fish Diseases, 38(7):665–673, July 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Marcos-Lopez:2016:PMA

[MLMR16] M. Marcos-López, S. O. Mitchell, and H. D. Rodger. Pathology and mortality associated with the mauve stinger jellyfish Pelagia noctiluca in farmed Atlantic salmon Salmo salar L. Journal of Fish Diseases, 39(1):111–115, January 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Marino:2012:SSZ

[MLR^{+}12] F. Marino, G. Lanteri, G. Rapisarda, A. Perillo, and B. Macrì. Spontaneous schwannoma in zebrafish, Danio rerio (Hamilton). Journal of Fish Diseases, 35(3):239–242, March 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Marcos-Lopez:2017:PSI

[MLRS^{+}17] M. Marcos-López, N. M. Ruane, F. Scholz, M. Bolton-Warberg, S. O. Mitchell, S. Murphy O’Sullivan, A. Irwin Moore, and H. D. Rodger. Piscirickettsia salmonis infection in cultured lumpfish (Cyclopterus lumpus L.). Journal of Fish Diseases, 40(11):1625–1634, November 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Millet:2011:VCD

[MLWC11] C. O. M. Millet, D. Lloyd, C. Williams, and J. Cable. In vitro culture of the diplomonad fish parasite Spironucleus vortens reveals unusually fast doubling time and atypical biphasic growth. Journal of Fish Diseases, 34(1):71–73, January 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Marty:2015:PRW

[MMB^{+}15] G. D. Marty, D. B. Morrison, J. Bidulka, T. Joseph, and A. Siah. Piscine reovirus in wild and farmed salmonids in British Columbia, Canada: 1974–2013. Journal of Fish Diseases, 38(8):713–728, August 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Matejusova:2010:ASS

[MMBS10] I. Matejusova, P. McKay, F. Bland, and M. Snow. Application of a sensitive, specific and controlled real-time PCR assay to surveillance indicates a low prevalence of viral haemorrhagic septicaemia virus (VHSV) in wild herring, Clupea harengus L., in Scottish waters. Journal of Fish Diseases, 33 (10):841–847, October 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Matsche:2012:ECN

[MMD^{+}12] M. A. Matsche, E. Markin, E. Donaldson, A. Hengst, and A. Lazur. Effect of chloride on nitrite-induced methaemoglobinemia in Atlantic sturgeon, Acipenser oxyrinchus oxyrinchus (Mitchill). Journal of Fish Diseases, 35(12):873– 885, December 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mateo:2016:FLT

[MMD16] D. R. Mateo, G. S. MacCallum, and J. Davidson. Field and laboratory transmission studies of haemic neoplasia in the soft-shell clam, Mya arenaria, from Atlantic Canada. Journal of Fish Diseases, 39(8):913–927, August 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mayfield:2011:SEF

[MMG^{+}11] S. Mayfield, R. McGarvey, H. K. Gorfine, H. Peeters, P. Burch, and S. Sharma. Survey estimates of fishable biomass following a mass mortality in an Australian molluscan fishery. Journal of Fish Diseases, 34(4):287–302, April 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Munro:2010:ALI

[MMH10] E. S. Munro, C. P. Millar, and T. S. Hastings. An analysis of levels of infectious pancreatic necrosis virus in Atlantic salmon, Salmo salar L., broodstock in Scotland between 1990–2002. Journal of Fish Diseases, 33(2):171–177, February 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Marin:2018:SAS

[MMH^{+}18] S. L. Marín, J. Mancilla, M. A. Hausdorf, D. Bouchard, M. S. Tudor, and F. Kane. Sensitivity assessment of sea lice to chemotherapeutants: Current bioassays and best practices. Journal of Fish Diseases, 41(6):995–1003, June 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mateo:2016:DHN

[MMMD16] D. R. Mateo, G. S. MacCallum, S. E. McGladdery, and J. Davidson. Distribution of haemic neoplasia of soft-shelled clams in Prince Edward Island: an examination of anthropogenic factors and effects of experimental fungicide exposure. Journal of Fish Diseases, 39(5):585–596, May 2016.

CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Myrmel:2014:IPN

[MMNL14] M. Myrmel, I. Modahl, H. Nygaard, and K. M. Lie. Infectious pancreatic necrosis virus in fish by-products is inactivated with inorganic acid (pH 1) and base (pH 12). Journal of Fish Diseases, 37(4):349–355, April 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mabrok:2016:TIS

[MMS^{+}16] M. Mabrok, M. Machado, C. R. Serra, A. Afonso, L. M. P. Valente, and B. Costas. Tenacibaculosis induction in the Senegalese sole (Solea senegalensis) and studies of Tenacibaculum maritimum survival against host mucus and plasma. Journal of Fish Diseases, 39(12):1445–1455, December 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Munro:2015:MEW

[MMW^{+}15] E. S. Munro, R. E. McIntosh, S. J. Weir, P. A. Noguera, J. M. Sandilands, I. Matejusova, A. S. Mayes, and R. Smith. A mortality event in wrasse species (Labridae) associated with the presence of viral haemorrhagic septicaemia virus. Journal of Fish Diseases, 38(4):335–341, April 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Madhun:2019:PGI

[MMWK19] Abdullah Sami Madhun, Stig Mæhle, Vidar Wennevik, and Egil Karlsbakk. Prevalence and genotypes of infectious salmon anaemia virus (ISAV) in returning wild Atlantic salmon (Salmo salar L.) in northern Norway. Journal of Fish Diseases, 42(8):1217–1221, August 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Maddocks:2015:PSD

[MNF^{+}15] C. E. Maddocks, E. T. Nolan, S. W. Feist, M. Crumlish, R. H. Richards, and C. F. Williams. Puffy skin disease (PSD) in rainbow trout, Oncorhynchus mykiss (Walbaum): a case definition. Journal of Fish Diseases, 38(7):653–664, July 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mishra:2018:CEG

[MNG^{+}18] A. Mishra, G-H. Nam, J-A. Gim, H-E. Lee, A. Jo, D. Yoon, S. Oh, S. Kim, A. Kim, D-H. Kim, Y. C. Kim, H. D. Jeong, H-J. Cha, Y. H. Choi, and H-S. Kim. Comparative evaluation of MCP gene in worldwide strains of Megalocytivirus (Iridoviridae family) for early diagnostic marker. Journal of Fish Diseases, 41(1):105–116, January 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mulei:2018:IPN

[MNM^{+}18] I. R. Mulei, P. N. Nyaga, P. G. Mbuthia, R. M. Waruiru, L. W. Njagi, E. W. Mwihia, A. A. A. Gamil, Ø. Evensen, and S. Mutoloki. Infectious pancreatic necrosis virus isolated from farmed rainbow trout and tilapia in Kenya is identical to European isolates. Journal of Fish Diseases, 41(8): 1191–1200, August 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mulei:2019:FDI

[MNM^{+}19] Isaac R. Mulei, Phillip N. Nyaga, Paul G. Mbuthia, Robert M. Waruiru, Cheng Xu, Øystein Evensen, and Stephen Mutoloki. First detection and isolation of infectious haematopoietic necrosis virus from farmed rainbow trout in Nyeri County, Kenya. Journal of Fish Diseases, 42(5):751– 758, May 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Morris:2012:TVC

[Mor12] D. J. Morris. Towards an in vitro culture method for the rainbow trout pathogen Tetracapsuloides bryosalmonae. Journal of Fish Diseases, 35(12):941–944, December 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mon-on:2018:RWQ

[MoSMS18] N. Mon-on, W. Surachetpong, S. Mongkolsuk, and K. Sirikanchana. Roles of water quality and disinfectant application on inactivation of fish pathogenic Streptococcus agalactiae with povidone iodine, quaternary ammonium compounds and glutaraldehyde. Journal of Fish Diseases, 41(5):783–789, May 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mohammed:2018:WKI

[MP18] Haitham H. Mohammed and Eric Peatman. Winter kill in intensively stocked channel catfish (Ictalurus punctatus): Coinfection with Aeromonas veronii, Streptococcus parauberis and Shewanella putrefaciens. Journal of Fish Diseases, 41(9):1339–1347, September 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Maj-Paluch:2016:GDS

[MPBM^{+}16] J. Maj-Paluch, E. Borzym, M. Matras, M. Stachnik, and M. Reichert. Genetic diversity of spring viraemia of carp virus isolates based on the glycoprotein gene. Journal of Fish Diseases, 39(10):1247–1252, October 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mikkelsen:2015:FEN

[MPJ^{+}15] S. S. Mikkelsen, V. Panzarin, S. P. Jonstrup, L. Bigarré, M. Baud, T. Gray, P-M. Agapow, and N. J. Olesen. fishpathogens.eu/noda: a free and handy online platform for betanodavirus targeted research and data sharing. Journal of Fish Diseases, 38(8):755–760, August 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Montes:2016:NSP

[MPM16] M. M. Montes, S. E. Plaul, and S. R. Martorelli. A new species of philometrid parasite (Nematoda, Philometridae) and histopathological lesions in juvenile whitemouth croakers, Micropogonias furnieri (Desmarest). Journal of Fish Diseases, 39(9):1053–1059, September 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Munday:2018:PSA

[MPS^{+}18] J. S. Munday, M. R. Perrott, J. E. Symonds, S. P. Walker, M. A. Preece, and P. S. Davie. Prevalence of spinal abnormalities in Chinook salmon smolt and influence of early rearing temperature and growth rates. Journal of Fish Diseases, 41(7):1111–1116, July 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mitchell:2011:RIG

[MR11] S. O. Mitchell and H. D. Rodger. A review of infectious gill disease in marine salmonid fish. Journal of Fish Diseases, 34 (6):411–432, June 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Marappan:2019:RPL

[MRN^{+}19] Makesh Marappan, Aathithya Rengarajan, Venkata Satyanarayana Nallala, Krishna Sukumaran, Aritra Bera, Thirugnamurty Sivaramakrishnan, Govindarajan Thiagarajan, Muniyandi Kailasam, and Vijayan Koyadan Kizhakedath. Resistance of pearlspot larvae, Etroplus suratensis, to redspotted grouper nervous necrosis virus by immersion challenge. Journal of Fish Diseases, 42(2):249–256, February 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Madan:2014:TDH

[MRR^{+}14] N. Madan, T. Rajkumar, N. Sundar Raj, M. A. Farook, K. S. N. Nambi, S. Abdul Majeed, and A. S. Sahul Hameed. Tissue distribution of hepatopancreatic parvo-like virus of shrimp in freshwater rice-field crab, Paratelphusa hydrodomous (Herbst). Journal of Fish Diseases, 37(11):969– 980, November 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Matthews:2013:GPS

[MRSC13] C. G. G. Matthews, R. H. Richards, A. P. Shinn, and D. I. Cox. Gill pathology in Scottish farmed Atlantic salmon, Salmo salar L., associated with the microsporidian Desmozoon lepeophtherii Freeman et Sommerville, 2009. Journal of Fish Diseases, 36(10):861–869, October 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

MacKenzie:2016:TLK

[MS16] K. MacKenzie and L. E. Smith. Two little-known metazoan parasites potentially pathogenic to smooth-hounds, Mustelus spp., in captivity. Journal of Fish Diseases, 39(4):511–514, April 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Matras:2019:PVS

[MSB^{+}19] Marek Matras, Magdalena Stachnik, Ewa Borzym, Joanna Maj-Paluch, and Michal Reichert. Potential vector species of carp edema virus (CEV). Journal of Fish Diseases, 42 (7):959–964, July 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Marshall:2016:LTE

[MSBB^{+}16] W. L. Marshall, A. Sitjà-Bobadilla, H. M. Brown, T. MacWilliam, Z. Richmond, H. Lamson, D. B. Morrison, and L. O. B. Afonso. Long-term epidemiological survey of Kudoa thyrsites (Myxozoa) in Atlantic salmon (Salmo salar L.) from commercial aquaculture farms. Journal of Fish Diseases, 39 (8):929–946, August 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Marana:2017:PCB

[MSC^{+}17] M. H. Marana, J. Skov, J. K. Chettri, B. Krossøy, I. Dalsgaard, P. W. Kania, and K. Buchmann. Positive correlation between Aeromonas salmonicida vaccine antigen concentration and protection in vaccinated rainbow trout Oncorhynchus mykiss evaluated by a tail fin infection model. Journal of Fish Diseases, 40(4):507–516, April 2017. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Militz:2014:EGA

[MSCH14] T. A. Militz, P. C. Southgate, A. G. Carton, and K. S. Hutson. Efficacy of garlic (Allium sativum) extract applied as a therapeutic immersion treatment for Neobenedenia sp. management in aquaculture. Journal of Fish Diseases, 37(5):451– 461, May 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mancilla:2018:MTI

[MSG^{+}18] M. Mancilla, J. Saavedra, M. Grandón, E. Tapia, E. Navas, H. Grothusen, and P. Bustos. The mutagenesis of a type IV secretion system locus of Piscirickettsia salmonis leads to the attenuation of the pathogen in Atlantic salmon, Salmo salar. Journal of Fish Diseases, 41(4):625–634, April 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mekata:2015:DSR

[MSI^{+}15] T. Mekata, J. Satoh, M. Inada, S. Dinesh, P. Harsha, T. Itami, and R. Sudhakaran. Development of simple, rapid and sensitive detection assay for grouper nervous necrosis virus using real-time loop-mediated isothermal amplification. Journal of Fish Diseases, 38(10):873–879, October 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mrlik:2012:LPM

[MSK^{+}12] V. Mrlik, M. Slany, J. Kubecka, J. Seda, A. Necas, V. Babak, I. Slana, P. Kriz, and I. Pavlik. A low prevalence of mycobacteria in freshwater fish from water reservoirs, ponds and farms. Journal of Fish Diseases, 35(7):497–504, July 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mancilla-Schulz:2019:DCR

[MSMM19] Jorge Mancilla-Schulz, Sandra L. Marín, and Carlos Molinet. Dynamics of Caligus rogercresseyi (Boxshall & Bravo, 2000) in farmed Atlantic salmon (Salmo salar) in southern Chile: Are we controlling sea lice? Journal of Fish Diseases, 42 (3):357–369, March 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mitchell:2010:EAS

[MSR^{+}10] S. O. Mitchell, T. Steinum, H. Rodger, C. Holland, K. Falk, and D. J. Colquhoun. Epitheliocystis in Atlantic salmon, Salmo salar L., farmed in fresh water in Ireland is associated with ‘Candidatus Clavochlamydia salmonicola’ infection. Journal of Fish Diseases, 33(8):665–673, August 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Minter:2012:SLT

[MSS^{+}12] L. J. Minter, M. K. Stoskopf, M. Serrano, O. Burrus, and G. A. Lewbart. Suspected lead toxicosis in an electric eel, Electrophorus electricus (L.). Journal of Fish Diseases, 35 (8):603–606, August 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

McColl:2017:CHP

[MSS^{+}17a] K. A. McColl, A. Sunarto, J. Slater, K. Bell, M. Asmus, W. Fulton, K. Hall, P. Brown, D. Gilligan, J. Hoad, L. M. Williams, and M. St J. Crane. Cyprinid herpesvirus 3 as a potential biological control agent for carp (Cyprinus carpio) in Australia: susceptibility of non-target species. Journal of Fish Diseases, 40(9):1141–1153, September 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Moreira:2017:PRR

[MSS^{+}17b] M. Moreira, D. Schrama, F. Soares, T. Wulff, P. Pousão-Ferreira, and P. Rodrigues. Physiological responses of reared sea bream (Sparus aurata Linnaeus, 1758) to an Amyloodinium ocellatum outbreak. Journal of Fish Diseases, 40(11): 1545–1560, November 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Monaghan:2015:EEI

[MTA^{+}15] S. J. Monaghan, K. D. Thompson, A. Adams, J. Kempter, and S. M. Bergmann. Examination of the early infection stages of koi herpesvirus (KHV) in experimentally infected carp, Cyprinus carpio L. using in situ hybridization. Journal of Fish Diseases, 38(5):477–489, May 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Monaghan:2015:SSP

[MTAB15] S. J. Monaghan, K. D. Thompson, A. Adams, and S. M. Bergmann. Sensitivity of seven PCRs for early detection of koi herpesvirus in experimentally infected carp, Cyprinus carpio L., by lethal and non-lethal sampling methods. Journal of Fish Diseases, 38(3):303–319, March 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Metselaar:2010:ARM

[MTG^{+}10] M. Metselaar, K. D. Thompson, R. M. L. Gratacap, M. J. L. Kik, S. E. LaPatra, S. J. Lloyd, D. R. Call, P. D. Smith, and A. Adams. Association of red-mark syndrome with a Rickettsia-like organism and its connection with strawberry disease in the USA. Journal of Fish Diseases, 33(10):849–858, October 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Molloy:2013:EDI

[MTHB13] S. D. Molloy, E. Thomas, K. Hoyt, and D. A. Bouchard. Enhanced detection of infectious salmon anaemia virus using a low-speed centrifugation technique in three fish cell lines. Journal of Fish Diseases, 36(1):35–44, January 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mavichak:2010:ELC

[MTK^{+}10] R. Mavichak, T. Takano, H. Kondo, I. Hirono, S. Wada, K. Hatai, H. Inagawa, Y. Takahashi, T. Yoshimura, H. Kiyono, Y. Yuki, and T. Aoki. The effect of liposome-coated recombinant protein VP28 against white spot syndrome virus in kuruma shrimp, Marsupenaeus japonicus. Journal of Fish Diseases, 33(1):69–74, January 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mosberian-Tanha:2018:CI

[MTLP^{+}18a] P. Mosberian-Tanha, T. Landsverk, C. M. Press, L. T. Mydland, J. W. Schrama, and M. Øverland. Cover image. Journal of Fish Diseases, 41(2):i, February 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mosberian-Tanha:2018:GER

[MTLP^{+}18b] P. Mosberian-Tanha, T. Landsverk, C. M. Press, L. T. Mydland, J. W. Schrama, and M. Øverland. Granulomatous enteritis in rainbow trout (Oncorhynchus mykiss) associated with soya bean meal regardless of water dissolved oxygen level. Journal of Fish Diseases, 41(2):269–280, February 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mizuno:2019:EEP

[MUM^{+}19] Shinya Mizuno, Shigehiko Urawa, Mahito Miyamoto, Makoto Hatakeyama, Nobuhisa Koide, and Hiroshi Ueda. Experimental evidence on prevention of infection by the ectoparasitic protozoans Ichthyobodo salmonis and Trichodina truttae in juvenile chum salmon using ultraviolet disinfection of rearing water. Journal of Fish Diseases, 42(1):129–140, January 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

MacLeod:2018:DCC

[MVM^{+}18] Michael J. MacLeod, Nguyen T. K. Vo, Michael S. Mikhaeil, S. Richelle Monaghan, J. Andrew N. Alexander, Mandeep K. Saran, and Lucy E. J. Lee. Development of a continuous cell line from larval Atlantic cod (Gadus morhua) and its use in the study of the microsporidian, Loma morhua. Journal of Fish Diseases, 41(9):1359–1372, September 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

McHugh:2013:FRN

[MVWS13] K. J. McHugh, J. C. Van Dyk, O. L. F. Weyl, and N. J. Smit. First report of nephrocalcinosis in a wild population of Mugil cephalus L. and Myxus capensis (Valenciennes). Journal of Fish Diseases, 36(10):887–889, October 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Martins:2017:MCE

[MWRSK17] M. L. Martins, V. Watral, J. P. Rodrigues-Soares, and M. L. Kent. A method for collecting eggs of Pseudocapillaria tomentosa (Nematoda: Capillariidae) from zebrafish Danio rerio and efficacy of heat and chlorine for killing the nematode’s eggs. Journal of Fish Diseases, 40(2):169–182, February 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

McHugh:2017:PDA

[MWS17] K. J. McHugh, O. L. F. Weyl, and N. J. Smit. Parasite diversity of African longfin eel Anguilla mossambica Peters with comments on host response to the monogenean Pseudodactylogyrus anguillae (Yin and Sproston). Journal of Fish Diseases, 40(7):959–961, July 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Maekawa:2018:CKN

[MYWC18] S. Maekawa, T. Yoshida, P-C. Wang, and S-C. Chen. Current knowledge of nocardiosis in teleost fish. Journal of Fish Diseases, 41(3):413–419, March 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Noguera:2010:LIP

[NB10] P. A. Noguera and D. W. Bruno. Liver involvement in post-smolt Atlantic salmon, Salmo salar L., infected with infectious pancreatic necrosis virus (IPNV): a retrospective histopathological study. Journal of Fish Diseases, 33(10): 819–832, October 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ndi:2011:ICI

[NB11] O. L. Ndi and M. D. Barton. Incidence of class 1 integron and other antibiotic resistance determinants in Aeromonas spp. from rainbow trout farms in Australia. Journal of Fish Diseases, 34(8):589–599, August 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nowak:2010:DSL

[NBJ10] B. F. Nowak, J. Bryan, and S. R. M. Jones. Do salmon lice, Lepeophtheirus salmonis, have a role in the epidemiology of amoebic gill disease caused by Neoparamoeba perurans? Journal of Fish Diseases, 33(8):683–687, August 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nguyen:2018:EIM

[NCNB18] T. D. Nguyen, P. B. B. Crosbie, B. F. Nowak, and A. R. Bridle. The effects of inactivation methods of Yersinia ruckeri on the efficacy of single dip vaccination in Atlantic salmon (Salmo salar). Journal of Fish Diseases, 41(7):1173–1176, July 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nya:2010:GCA

[NDA10] E. J. Nya, Z. Dawood, and B. Austin. The garlic component, allicin, prevents disease caused by Aeromonas hydrophila in rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 33(4):293–300, April 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Noor:2019:QSR

[NDA^{+}19] Noorashikin Md. Noor, Tom Defoirdt, Norfarrah Alipiah, Murni Karim, Hassan Daud, and Ikhsan Natrah. Quorum sensing is required for full virulence of Vibrio campbellii towards tiger grouper (Epinephelus fuscoguttatus) larvae. Journal of Fish Diseases, 42(4):489–495, April 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Newaj-Fyzul:2015:PIP

[NFA15] A. Newaj-Fyzul and B. Austin. Probiotics, immunostimulants, plant products and oral vaccines, and their role as feed supplements in the control of bacterial fish diseases. Journal of Fish Diseases, 38(11):937–955, November 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nicholson:2017:DTL

[NFF^{+}17] P. Nicholson, M. A. Fathi, A. Fischer, C. Mohan, E. Schieck, N. Mishra, A. Heinimann, J. Frey, B. Wieland, and J. Jores. Detection of Tilapia Lake Virus in Egyptian fish farms experiencing high mortalities in 2015. Journal of Fish Diseases, 40(12):1925–1928, December 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Novotny:2010:MDG

[NHM^{+}10] L. Novotny, R. Halouzka, L. Matlova, O. Vavra, L. Bartosova, M. Slany, and I. Pavlik. Morphology and distribution of granulomatous inflammation in freshwater ornamental fish infected with mycobacteria. Journal of Fish Diseases, 33(12): 947–955, December 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nguyen:2017:DHA

[NJK^{+}17] Tuan Thuc Nguyen, Yeonhwa Jin, Jolanta Kiełpińska, Sven M. Bergmann, Matthias Lenk, and Remigiusz Panicz. Detection of Herpesvirus anguillae (AngHV-1) in European eel Anguilla anguilla (L.) originating from northern Poland — assessment of suitability of selected diagnostic methods. Journal of Fish Diseases, 40(11):1717–1723, November 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Novak:2016:IRS

[NLC^{+}16] C. W. Novak, D. L. Lewis, B. Collicutt, K. Verkaik, and D. E. Barker. Investigations on the role of the salmon louse, Lepeophtheirus salmonis (Caligidae), as a vector in the transmission of Aeromonas salmonicida subsp. salmonicida. Journal of Fish Diseases, 39(10):1165–1178, October 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Netto:2011:SDA

[NLF11] Lamartine N. Netto, C. A. G. Leal, and H. C. P. Figueiredo. Streptococcus dysgalactiae as an agent of septicaemia in Nile tilapia, Oreochromis niloticus (L.). Journal of Fish Diseases, 34(3):251–254, March 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nguyen:2016:DRT

[NLKA16] T. L. Nguyen, Y. J. Lim, D-H. Kim, and B. Austin. Development of real-time PCR for detection and quantitation of Streptococcus parauberis. Journal of Fish Diseases, 39(1): 31–39, January 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nishiki:2014:CES

[NMIY14] I. Nishiki, T. Minami, T. Itami, and T. Yoshida. Cloning and expression of a surface immunogenic protein in Streptococcus dysgalactiae isolated from fish and its application in enzyme-linked immunosorbent assays to diagnose S. dysgalactiae infections in fish. Journal of Fish Diseases, 37(12): 1031–1039, December 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nishiki:2018:MSA

[NMM^{+}18] Issei Nishiki, Takayuki Minami, Ayana Murakami, Truong Dinh Hoai, and Atushi Fujiwara. Multilocus sequence analysis of Vibrionaceae isolated from farmed amberjack and the development of a multiplex PCR assay for the detection of pathogenic species. Journal of Fish Diseases, 41(8):1295– 1301, August 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nanthini:2019:VPT

[NMV^{+}19] Ravi Nanthini, Seepoo Abdul Majeed, Sugumar Vimal, Gani Taju, Selvam Sivakumar, Shanmugam Santhosh Kumar, Devika Pillai, Kalasseril Girijan Sneha, Chakalaparambil Gokulan Rakesh, and Azeez Sait Sahul Hameed. In vitro propagation of tilapia lake virus in cell lines developed from Oreochromis mossambicus. Journal of Fish Diseases, 42(11): 1543–1552, November 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nagai:2011:GFP

[NN11] T. Nagai and T. Nakai. Growth of Flavobacterium psychrophilum in fish serum correlates with pathogenicity. Journal of Fish Diseases, 34(4):303–310, April 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nguyen:2018:EIR

[NNC^{+}18] Hai Trong Nguyen, Thuy Thi Thu Nguyen, Yan-Chun Chen, Hung Vu-Khac, Pei-Chi Wang, and Shih-Chu Chen. Enhanced immune responses and effectiveness of refined outer membrane protein vaccines against Vibrio harveyi in orangespotted grouper (Epinephelus coioides). Journal of Fish Diseases, 41(9):1349–1358, September 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Navarro-Nava:2011:IMP

[NNCLGC^{+}11] F. Navarro-Nava, R. Castro-Longoria, J. M. Grijalva-Chon, J. Ramos-Paredes, and J. De la Rosa-Vélez. Infection and mortality of Penaeus vannamei at extreme salinities when challenged with Mexican yellow head virus. Journal of Fish Diseases, 34(4):327–329, April 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nunez-Ortiz:2016:QID

[NOST^{+}16] N. Nuñez-Ortiz, V. Stocchi, A. Toffan, F. Pascoli, N. Sood, F. Buonocore, S. Picchietti, C. Papeschi, A. R. Taddei, K. D. Thompson, and G. Scapigliati. Quantitative immunoenzymatic detection of viral encephalopathy and retinopathy virus (betanodavirus) in sea bass Dicentrarchus labrax. Journal of Fish Diseases, 39(7):821–831, July 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nilsen:2011:SFP

[NOV^{+}11] H. Nilsen, A. B. Olsen, Ø. Vaagnes, H. Hellberg, K. Bottolfsen, H. Skjelstad, and D. J. Colquhoun. Systemic Flavobacterium psychrophilum infection in rainbow trout, Oncorhynchus mykiss (Walbaum), farmed in fresh and brackish water in Norway. Journal of Fish Diseases, 34 (5):403–408, May 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nakamura:2019:APL

[NPGA^{+}19] Rika Nakamura, Ivane R. Pedrosa-Gerasmio, Rod Russel R. Alenton, Reiko Nozaki, Hidehiro Kondo, and Ikuo Hirono. Anti-PirA-like toxin immunoglobulin (IgY) in feeds passively immunizes shrimp against acute hepatopancreatic necrosis disease. Journal of Fish Diseases, 42(8):1125–1132, August 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nelson:2018:HVD

[NRF^{+}18] E. J. Nelson, S. M. C. Robinson, N. Feindel, A. Sterling, A. Byrne, and K. Pee Ang. Horizontal and vertical distribution of sea lice larvae (Lepeophtheirus salmonis) in and around salmon farms in the Bay of Fundy, Canada. Journal of Fish Diseases, 41(6):885–899, June 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nayak:2014:FLC

[NRMS14] S. Nayak, N. Ramaiah, R. M. Meena, and R. A. Sreepada. Full-length cloning and phylogenetic analyses of translationally controlled tumour protein and ferritin genes from the Indian white prawn, Fenneropenaeus indicus (H. Milne Edwards). Journal of Fish Diseases, 37(2):77–87, February 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ngo:2018:ASF

[NSB^{+}18] T. P. H. Ngo, P. Smith, K. L. Bartie, K. D. Thompson, D. W. Verner-Jeffreys, R. Hoare, and A. Adams. Antimicrobial susceptibility of Flavobacterium psychrophilum isolates from the United Kingdom. Journal of Fish Diseases, 41(2): 309–320, February 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nolan:2015:DCV

[NSC^{+}15] D. Nolan, F. Stephens, M. Crockford, J. B. Jones, and M. Snow. Detection and characterization of viruses of the genus Megalocytivirus in ornamental fish imported into an Australian border quarantine premises: an emerging risk to national biosecurity. Journal of Fish Diseases, 38(2):187– 195, February 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ngo:2018:QAD

[NSC^{+}18] Thuy T. N. Ngo, Alistair M. Senior, Antica Culina, Eduardo S. A. Santos, Just M. Vlak, and Mark P. Zwart. Quantitative analysis of the dose–response of white spot syndrome virus in shrimp. Journal of Fish Diseases, 41(11):1733–1744, November 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nerbovik:2017:MCM

[NSE^{+}17] I-K. G. Nerbøvik, M. A. Solheim, H. Ø. Eggestøl, A. Rønneseth, R. A. Jakobsen, H. I. Wergeland, and G. T. Haugland. Molecular cloning of MDA5, phylogenetic analysis of RIG-I-like receptors (RLRs) and differential gene expression of RLRs, interferons and proinflammatory cytokines after in vitro challenge with IPNV, ISAV and SAV in the salmonid cell line TO. Journal of Fish Diseases, 40(11): 1529–1544, November 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nanjo:2017:SIG

[NSS^{+}17] A. Nanjo, T. Shibata, M. Saito, K. Yoshii, M. Tanaka, T. Nakanishi, H. Fukuda, T. Sakamoto, G. Kato, and M. Sano. Susceptibility of isogeneic ginbuna Carassius auratus langsdorfii Temminck et Schlegel to cyprinid herpesvirus-2 (CyHV-2) as a model species. Journal of Fish Diseases, 40 (2):157–168, February 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nguyen:2016:QGP

[NWMK16] N. H. Nguyen, P. Whatmore, A. Miller, and W. Knibb. Quantitative genetic properties of four measures of deformity in yellowtail kingfish Seriola lalandi Valenciennes, 1833. Journal of Fish Diseases, 39(2):217–228, February 2016. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nithikulworawong:2012:MCI

[NYRS12] N. Nithikulworawong, A. Yakupitiyage, Sk Rakshit, and P. Srisapoome. Molecular characterization and increased expression of the Nile tilapia, Oreochromis niloticus (L.), Tcell receptor beta chain in response to Streptococcus agalactiae infection. Journal of Fish Diseases, 35(5):343–358, May 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Oladele:2012:JSA

[OAO^{+}12] O. O. Oladele, O. L. Ajayi, O. O. Olude, O. O. Stephen, A. A. Adediji, I. O. Arasi, and U. G. Ntiwunka. Jaundice syndrome in African sharp-tooth catfish, Clarias gariepinus (Burchell), associated with haemolytic Staphylococcus aureus. Journal of Fish Diseases, 35(12):945–947, December 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Overton:2010:CSD

[OBD10] J. L. Overton, M. S. Bruun, and I. Dalsgaard. Chemical surface disinfection of eggs of Baltic cod, Gadus morhua L. Journal of Fish Diseases, 33(9):707–716, September 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Orieux:2011:QFP

[OBD^{+}11] N. Orieux, J-P. Bourdineaud, D-G. Douet, P. Daniel, and M. Le Hénaff. Quantification of Flavobacterium psychrophilum in rainbow trout, Oncorhynchus mykiss (Walbaum), tissues by qPCR. Journal of Fish Diseases, 34(11): 811–821, November 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Olsen:2012:MPI

[OBFR12] C. M. Olsen, S. Braaen, K. Falk, and E. Rimstad. Multiple passage of infectious salmon anaemia virus in rainbow trout, Oncorhynchus mykiss (Walbaum), did not induce increased virus load. Journal of Fish Diseases, 35(11):827–838, November 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ofelio:2014:IMI

[OBR^{+}14] C. Ofelio, A. Blanco, Á. Roura, J. Pintado, S. Pascual, and M. Planas. Isolation and molecular identification of the scuticociliate Porpostoma notata Moebius, 1888 from moribund reared Hippocampus hippocampus (L.) seahorses, by amplification of the SSU rRNA gene sequences. Journal of Fish Diseases, 37(12):1061–1065, December 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ortega:2019:FDS

[OCLI^{+}19] Cesar Ortega, Leticia Cañas-Lopez, Rute Irgang, Raúl Fajardo, Matías Poblete-Morales, Benjamin Valladares-Carranza, Diana Tapia-Cammas, and Ruben Avendaño-Herrera. First detection of spring viraemia of carp virus in common carp (Cyprinus carpio L.) affected by a septicaemic disease in Mexico. Journal of Fish Diseases, 42(5):667–675, May 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Olson:2013:CSA

[OEG^{+}13] W. Olson, E. Emmenegger, J. Glenn, J. Winton, and F. Goetz. Comparative susceptibility among three stocks of yellow perch, Perca flavescens (Mitchill), to viral haemorrhagic septicaemia virus strain IVb from the Great Lakes. Journal of Fish Diseases, 36(8):711–719, August 2013. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Olivares-Fuster:2011:DCR

[OFA11] O. Olivares-Fuster and C. R. Arias. Development and characterization of rifampicin-resistant mutants from high virulent strains of Flavobacterium columnare. Journal of Fish Diseases, 34(5):385–394, May 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ortega:2018:FIC

[OGI^{+}18] C. Ortega, I. García, R. Irgang, R. Fajardo, D. Tapia-Cammas, J. Acosta, and R. Avendaño-Herrera. First identification and characterization of Streptococcus iniae obtained from tilapia (Oreochromis aureus) farmed in Mexico. Journal of Fish Diseases, 41(5):773–782, May 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Osorio:2010:ATG

[OJRL10] C. R. Osorio, S. Juiz-Río, and M. L. Lemos. The ABCtransporter hutCD genes of Photobacterium damselae subsp. piscicida are essential for haem utilization as iron source and are expressed during infection in fish. Journal of Fish Diseases, 33(8):649–655, August 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Oh:2013:CIT

[OKS^{+}13] M-J. Oh, W-S. Kim, H. G. Seo, H. J. Gye, and T. Nishizawa. Change in infectivity titre of nervous necrosis virus (NNV) in brain tissue of sevenband grouper, Epinephalus fasciatus Thunberg, with Poly(I:C) administration. Journal of Fish Diseases, 36(2):159–162, February 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Owens:2011:EPM

[OLC11] L. Owens, K. La Fauce, and K. Claydon. The effect of Penaeus merguiensis densovirus on Penaeus merguiensis production in Queensland, Australia. Journal of Fish Diseases, 34(7):509–515, July 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Overturf:2010:RBG

[OLT^{+}10] K. Overturf, S. LaPatra, R. Towner, N. Campbell, and S. Narum. Relationships between growth and disease resistance in rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 33(4):321–329, April 2010. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Oidtmann:2013:DCE

[OLVJ^{+}13] B. Oidtmann, S. E. LaPatra, D. Verner-Jeffreys, M. Pond, E. J. Peeler, P. A. Noguera, D. W. Bruno, S. St-Hilaire, C. B. Schubiger, K. Snekvik, M. Crumlish, D. M. Green, M. Metselaar, H. Rodger, H. Schmidt-Posthaus, M. Galeotti, and S. W. Feist. Differential characterization of emerging skin diseases of rainbow trout — a standardized approach to capturing disease characteristics and development of case definitions. Journal of Fish Diseases, 36(11):921–937, November 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Olivares:2010:DMC

[OM10] J. Olivares and S. H. Marshall. Determination of minimal concentration of Piscirickettsia salmonis in water columns to establish a fallowing period in salmon farms. Journal of Fish Diseases, 33(3):261–266, March 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Oliveira:2019:APL

[OMB^{+}19] Lígia C. D. Oliveira, Cláudia Majolo, Franmir R. Brandão, Caio Francisco S. Farias, Maria Inês B. Oliveira, Welliton B. Santos, Patrícia C. Monteiro, Maria Juliete S. Rocha, Edsandra C. Chagas, and Marcos Tavares-Dias. Avermectins, praziquantel and levamisole have in vitro efficacy against Neoechinorhynchus buttnerae (Neoechinorhynchidae) in Colossoma macropomum: A Serrasalmidae from the Amazon. Journal of Fish Diseases, 42(5):765–772, May 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Okazaki:2015:DJE

[OMO^{+}15] S. Okazaki, H. Manabe, T. Omatsu, S. Tsuchiaka, T. Yamamoto, S. Chow, T. Shibuno, K. Watanabe, S. Ono, H. Kuwada, and T. Mizutani. Detection of Japanese eel endothelial cells-infecting virus (JEECV) in the Japanese eel Anguilla japonica (Temminck & Schlegel), living in natural habitats. Journal of Fish Diseases, 38(9):849–852, September 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ottesen:2010:HCA

[ONA10] O. H. Ottesen, E. J. Noga, and A. Amin. Histopathology of culture-associated skin erosions and papillary hyperplasia of Atlantic halibut, Hippoglossus hippoglossus (L.). Journal of Fish Diseases, 33(6):489–496, June 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ortega:2014:LPA

[ONGQ^{+}14] J. Ortega, A. Noguera, A. García-Quirós, D. Viana, L. Selva, L. de Juan, B. Romero, D. García-Parraga, J. L. Crespo, and J. M. Corpa. Lesional patterns associated with mycobacteriosis in an Atlantic horse mackerel, Trachurus trachurus (L.), aquarium population. Journal of Fish Diseases, 37(6):591–595, June 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Oladele:2011:AON

[OOO^{+}11] O. O. Oladele, B. E. Olufemi, G. A. Oladosu, O. L. Ajayi, A. A. Adediji, and I. O. Arasi. Arborescent organ necrosis syndrome in catfish, Clarias gariepinus (Burchell): a case report. Journal of Fish Diseases, 34(10):801–804, October 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ogut:2014:HLL

[OP14] H. Ogut and R. Parlak. Hexamitiasis leads to lower metabolic rates in rainbow trout Oncorhynchus mykiss (Walbaum) juveniles. Journal of Fish Diseases, 37(12):1013–1020, December 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Orozova:2018:RBM

[OSAA18] P. Orozova, I. Sirakov, D. A. Austin, and B. Austin. Recovery of Bacillus mycoides, B. pseudomycoides and Aeromonas hydrophila from common carp (Cyprinus carpio) and rainbow trout (Oncorhynchus mykiss) with gill disease. Journal of Fish Diseases, 41(1):125–129, January 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Orozova:2014:RHA

[OSC^{+}14] P. Orozova, I. Sirakov, V. Chikova, R. Popova, A. H. Al-Harbi, M. Crumlish, and B. Austin. Recovery of Hafnia alvei from diseased brown trout, Salmo trutta L., and healthy noble crayfish, Astacus astacus (L.), in Bulgaria. Journal of Fish Diseases, 37(10):891–898, October 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Oliva-Teles:2012:NHA

[OT12] A. Oliva-Teles. Nutrition and health of aquaculture fish. Journal of Fish Diseases, 35(2):83–108, February 2012. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Oh:2012:FTP

[OTN^{+}12] M-J. Oh, I. Takami, T. Nishizawa, W-S. Kim, C-S. Kim, S-R. Kim, and M-A. Park. Field tests of Poly(I:C) immunization with nervous necrosis virus (NNV) in sevenband grouper, Epinephelus septemfasciatus (Thunberg). Journal of Fish Diseases, 35(3):187–191, March 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

OShea:2017:DHS

[OWM^{+}17] B. O’Shea, S. Wadsworth, J. Pino Marambio, M. A. Birkett, J. A. Pickett, and A. J. Mordue (Luntz). Disruption of host-seeking behaviour by the salmon louse, Lepeophtheirus salmonis, using botanically derived repellents. Journal of Fish Diseases, 40(4):495–505, April 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Powell:2018:ALM

[PÅD18] M. D. Powell, Å. Åtland, and T. Dale. Acute lion’s mane jellyfish, Cyanea capillata (Cnideria: Scyphozoa), exposure to Atlantic salmon (Salmo salar L.). Journal of Fish Diseases, 41(5):751–759, May 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Paradis:2019:OTC

[PAM^{+}19] Helene Paradis, Raahyma Ahmad, James McDonald, Danny Boyce, and Robert L. Gendron. Ocular tissue changes associated with anterior segment opacity in lumpfish (Cyclopterus lumpus L.) eye. Journal of Fish Diseases, 42(10):1401–1408, October 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Perretta:2018:PMP

[PAZ18] Alejandro Perretta, Karina Antúnez, and Pablo Zunino. Phenotypic, molecular and pathological characterization of motile aeromonads isolated from diseased fishes cultured in Uruguay. Journal of Fish Diseases, 41(10):1559–1569, October 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pietrak:2018:TLC

[PB18] M. Pietrak and S. Backman. Treatment of lumpfish (Cyclopterus lumpus L.) infected with Gyrodactylus cyclopteri (Scyborskaya 1948). Journal of Fish Diseases, 41(4):721– 723, April 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pomposini:2019:LML

[PBBG19] Amanda Pomposini, Jonathan Blubaugh, Robert Chad Boyce, and David T. Gauthier. Leech (Myzobdella lugubris) infestations in largemouth bass (Micropterus salmoides) in Back Bay, Virginia, USA. Journal of Fish Diseases, 42(5): 739–749, May 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Parisi:2015:SBD

[PBC15] M. G. Parisi, G. Benenati, and M. Cammarata. Sea bass Dicentrarchus labrax (L.) bacterial infection and confinement stress acts on F-type lectin (DlFBL) serum modulation. Journal of Fish Diseases, 38(11):967–976, November 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pascoli:2015:RPJ

[PBM^{+}15] F. Pascoli, F. Borghesan, A. Manfrin, F. Quaglio, F. Gatti, and A. Toffan. Redfin perch juveniles, Perca fluviatilis L., are resistant to European viral haemorrhagic septicaemia virus and infectious haematopoietic necrosis virus delivered via immersion in experimental condition. Journal of Fish Diseases, 38(6):589–594, June 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Perdiguero:2011:IRE

[PBMB11] G. Muñoz Perdiguero, A. Barrientos, E. Madrid, and T. H. Birkbeck. Isolation of Rhodococcus erythropolis from vaccinated Atlantic salmon, Salmo salar L., smolts in Chile. Journal of Fish Diseases, 34(9):715–717, September 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pinkney:2014:B

[PBW14] A. E. Pinkney, P. C. Baumann, and J. C. Wolf. Biography. Journal of Fish Diseases, 37(10):859–861, October 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Puk:2018:DMM

[PBW^{+}18] K. Puk, T. Banach, A. Wawrzyniak, Ł. Adaszek, J. Zietek, S. Winiarczyk, and L. Guz. Detection of Mycobacterium marinum, M. peregrinum, M. fortuitum and M. abscessus in aquarium fish. Journal of Fish Diseases, 41(1):153–156, January 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Perpinan:2017:MMB

[PC17] D. Perpiñán and T. Costa. Metastatic mineralization in blacktip reef sharks, Carcharhinus melanopterus (Quoy & Gaimard 1824). Journal of Fish Diseases, 40(3):447–451, March 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Phromkunthong:2015:PCA

[PCK^{+}15] W. Phromkunthong, P. Choochuay, V. Kiron, N. Nuntapong, and M. Boonyaratpalin. Pathophysiological changes associated with dietary melamine and cyanuric acid toxicity in red tilapia. Journal of Fish Diseases, 38(2):161–173, February 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Prado:2014:VHC

[PDdC^{+}14] S. Prado, J. Dubert, F. da Costa, D. Martínez-Patiño, and J. L. Barja. Vibrios in hatchery cultures of the razor clam, Solen marginatus (Pulteney). Journal of Fish Diseases, 37 (3):209–217, March 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pawluk:2019:SLL

[PdLC19] Rebecca Jane Pawluk, Carlos Garcia de Leaniz, and Sofia Consuegra. Sea lice loads correlate with the diversity at the Major Histocompatibility Complex-related loci in farmed Atlantic salmon, Salmo salar. Journal of Fish Diseases, 42 (7):1091–1093, July 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Papadopoulou:2017:VAF

[PDLW17] A. Papadopoulou, I. Dalsgaard, A. Lindén, and T. Wiklund. In vivo adherence of Flavobacterium psychrophilum to mucosal external surfaces of rainbow trout (Oncorhynchus mykiss) fry. Journal of Fish Diseases, 40(10):1309–1320, October 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pereira:2012:BTC

[PdSG^{+}12] W. L. A. Pereira, A. J. S. de Souza, Á. M. Gabriel, A. M. C. Cardoso, S. G. B. Monger, I. C. A. Seligmann, A. C. A. Pereira, and D. K. S. Queiroz. ‘Branchiomycosis in tambaqui, Colossoma macropomum (Cuvier), from the eastern Brazilian Amazon’. Journal of Fish Diseases, 35(8):615–617, August 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Purcell:2013:CIF

[PFC^{+}13] S. L. Purcell, S. E. Friend, J. M. Covello, A. Donkin, D. B. Groman, J. Poley, and M. D. Fast. CpG inclusion in feed reduces sea lice, Lepeophtheirus salmonis, numbers following re-infection. Journal of Fish Diseases, 36(3):229–240, March 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Polinski:2010:CSC

[PFJ^{+}10] M. P. Polinski, T. R. Fehringer, K. A. Johnson, K. R. Snekvik, S. E. LaPatra, B. R. LaFrentz, S. C. Ireland, and K. D. Cain. Characterization of susceptibility and carrier status of burbot, Lota lota (L.), to IHNV, IPNV, Flavobacterium psychrophilum, Aeromonas salmonicida and Renibacterium salmoninarum. Journal of Fish Diseases, 33(7):559– 570, July 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Piackova:2013:SCC

[PFP^{+}13] V. Piačková, M. Flajšhans, D. Pokorová, S. Reschová, D. Gela, A. Čížek, and T. Veselý. Sensitivity of common carp, Cyprinus carpio L., strains and crossbreeds reared in the Czech Republic to infection by cyprinid herpesvirus 3 (CyHV-3; KHV). Journal of Fish Diseases, 36(1):75–80, January 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Powell:2016:ELM

[PG16] M. D. Powell and A. K. Gamperl. Effects of Loma morhua (Microsporidia) infection on the cardiorespiratory performance of Atlantic cod Gadus morhua (L.). Journal of Fish Diseases, 39(2):189–204, February 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Perolo:2019:ONG

[PGF^{+}19] Alberto Perolo, Andrea Gustinelli, Maria Letizia Fioravanti, Amedeo Manfrin, Manuela Dalla Pozza, Fernando Lunelli, Adelchi Accini, and Francesco Quaglio. Occurrence of nodular gill disease in farmed brown trout (Salmo trutta L.). Journal of Fish Diseases, 42(9):1315–1320, September 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pinkney:2014:TSP

[PHR14] A. E. Pinkney, J. C. Harshbarger, and M. A. Rutter. Temporal and spatial patterns in tumour prevalence in brown bullhead Ameiurus nebulosus (Lesueur) in the tidal Potomac River watershed (USA). Journal of Fish Diseases, 37(10): 863–876, October 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Palmer:2012:PAM

[PHvdH12] L. J. Palmer, N. S. Hogan, and M. R. van den Heuvel. Phylogenetic analysis and molecular methods for the detection of lymphocystis disease virus from yellow perch, Perca flavescens (Mitchell). Journal of Fish Diseases, 35(9):661– 670, September 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pettersen:2013:ODA

[PHVM13] R. A. Pettersen, S. Hytterød, L. A. Vøllestad, and T. A. Mo. Osmoregulatory disturbances in Atlantic salmon, Salmo salar L., caused by the monogenean Gyrodactylus salaris. Journal of Fish Diseases, 36(1):67–70, January 2013. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Phuoc:2016:PIS

[PHW^{+}16] L. H. Phuoc, B. Hu, M. Wille, N. T. Hien, V. H. Phuong, N. T. N. Tinh, N. H. Loc, P. Sorgeloos, and P. Bossier. Priming the immune system of Penaeid shrimp by bacterial HSP70 (DnaK). Journal of Fish Diseases, 39(5):555– 564, May 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pham:2012:PWI

[PJL^{+}12] P. H. Pham, J. Jung, J. S. Lumsden, B. Dixon, and N. C. Bols. The potential of waste items in aquatic environments to act as fomites for viral haemorrhagic septicaemia virus. Journal of Fish Diseases, 35(1):73–77, January 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Provan:2013:PAE

[PJU^{+}13] F. Provan, L. B. Jensen, K. E. Uleberg, E. Larssen, T. Rajalahti, J. Mullins, and A. Obach. Proteomic analysis of epidermal mucus from sea lice–infected Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 36(3):311–321, March 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pietrak:2019:TSS

[PJZB19] Michael Pietrak, Alexander Jensen, Gayle Barbin Zydlewski, and Ian Bricknell. The Three-spined Stickleback, Gasterosteus aculeatus Linnaeus 1758, plays a minor role as a host of Lepeophtheirus salmonis (Krøyer 1837) in the Gulf of Maine. Journal of Fish Diseases, 42(7):985–989, July 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Peng:2019:INS

[PKLH19] Tao Peng, Jie Kan, Jingsheng Lun, and Zhong Hu. Identification of novel sRNAs involved in oxidative stress response in the fish pathogen Vibrio alginolyticus by transcriptome analysis. Journal of Fish Diseases, 42(2):277–291, February 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pekala:2015:PGC

[PKPG15] A. Pekala, A. Kozińska, E. Paździor, and H. Głowacka. Phenotypical and genotypical characterization of Shewanella putrefaciens strains isolated from diseased freshwater fish. Journal of Fish Diseases, 38(3):283–293, March 2015. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pulpipat:2019:CI

[PLC^{+}19a] Theeraporn Pulpipat, Kuo-Hua Lin, Ying-Hao Chen, Pei-Chi Wang, and Shih-Chu Chen. Cover image. Journal of Fish Diseases, 42(5):i, May 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pulpipat:2019:MCP

[PLC^{+}19b] Theeraporn Pulpipat, Kuo-Hua Lin, Ying-Hao Chen, Pei-Chi Wang, and Shih-Chu Chen. Molecular characterization and pathogenicity of Francisella noatunensis subsp. orientalis isolated from cultured tilapia (Oreochromis sp.) in Taiwan. Journal of Fish Diseases, 42(5):643–655, May 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Plant:2011:IRT

[PLCC11] K. P. Plant, S. E. LaPatra, D. R. Call, and K. D. Cain. Immunization of rainbow trout, Oncorhynchus mykiss (Walbaum), with Flavobacterium psychrophilum proteins elongation factor-Tu, SufB Fe-S assembly protein and ATP synthaseβ. Journal of Fish Diseases, 34(3):247–250, March 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Plaul:2013:MAH

[PMBM13] S. E. Plaul, M. M. Montes, C. G. Barbeito, and S. R. Martorelli. Morphological aspects and histological effects of the attachment organ of Parabrachiella sp. (Copepoda: Lernaeopodidae) on the grey mullet, Mugil liza Valenciennes. Journal of Fish Diseases, 36(10):853–859, October 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Park:2011:RGU

[PMH^{+}11] J. W. Park, C. H. Moon, A. Harmache, A. R. Wargo, M. K. Purcell, M. Bremont, and G. Kurath. Restricted growth of U-type infectious haematopoietic necrosis virus (IHNV) in rainbow trout cells may be linked to casein kinase II activity. Journal of Fish Diseases, 34(2):115–129, February 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pridgeon:2012:EPS

[PMK12] J. W. Pridgeon, X. Mu, and P. H. Klesius. Expression profiles of seven channel catfish antimicrobial peptides in response to Edwardsiella ictaluri infection. Journal of Fish Diseases, 35 (3):227–237, March 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Polinas:2018:EHA

[PMP^{+}18] Marta Polinas, Salvatore Mele, Francesc Padrós, Paolo Merella, Elisabetta Antuofermo, Adam Gouraguine, and Olga Reñones. Ecological and histopathological aspects of Didymodiclinus sp. (Trematoda: Didymozoidae) parasite of the dusky grouper, Epinephelus marginatus (Osteichthyes: Serranidae), from the western Mediterranean Sea. Journal of Fish Diseases, 41(9):1385–1393, September 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Purcell:2016:ETR

[PMPG^{+}16] M. K. Purcell, C. L. McKibben, S. Pearman-Gillman, D. G. Elliott, and J. R. Winton. Effects of temperature on Renibacterium salmoninarum infection and transmission potential in Chinook salmon, Oncorhynchus tshawytscha (Walbaum). Journal of Fish Diseases, 39(7):787–798, July 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Park:2010:DGU

[PMW^{+}10] J. W. Park, C. H. Moon, A. R. Wargo, M. K. Purcell, and G. Kurath. Differential growth of U and M type infectious haematopoietic necrosis virus in a rainbow trout–derived cell line, RTG-2. Journal of Fish Diseases, 33(7):583–591, July 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Phromkunthong:2013:TMA

[PNBK13] W. Phromkunthong, N. Nuntapong, M. Boonyaratpalin, and V. Kiron. Toxicity of melamine, an adulterant in fish feeds: experimental assessment of its effects on tilapia. Journal of Fish Diseases, 36(6):555–568, June 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pollock:2015:CSD

[Pol15] B. R. Pollock. Comments on ‘Saddleback deformities in yellowfin bream, Acanthopagrus australis (Günther), from South-East Queensland’ by Diggles (2013). Journal of Fish Diseases, 38(3):329–330, March 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Purcell:2018:MTA

[PPE^{+}18] M. K. Purcell, R. L. Powers, J. Evered, J. Kerwin, T. R. Meyers, B. Stewart, and J. R. Winton. Molecular testing of adult Pacific salmon and trout (Oncorhynchus spp.) for several RNA viruses demonstrates widespread distribution of piscine orthoreovirus in Alaska and Washington. Journal of Fish Diseases, 41(2):347–355, February 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pollock:2012:DFM

[PPF^{+}12] R. J. Pollock, M. S. Pollock, M. C. O. Ferrari, S. G. W. Kaminskyj, and D. P. Chivers. Do fathead minnows, Pimephales promelas Rafinesque, alter their club cell investment in responses to variable risk of infection from Saprolegnia? Journal of Fish Diseases, 35(4):249–254, April 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Poley:2013:CEA

[PPI^{+}13] J. Poley, S. L. Purcell, O. O. Igboeli, A. Donkin, H. Wotton, and M. D. Fast. Combinatorial effects of administration of immunostimulatory compounds in feed and follow-up administration of triple-dose SLICE(R) (emamectin benzoate) on Atlantic salmon, Salmo salar L., infection with Lepeophtheirus salmonis. Journal of Fish Diseases, 36(3):299–309, March 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pittman:2013:BSM

[PPK^{+}13] K. Pittman, A. Pittman, S. Karlson, T. Cieplinska, P. Sourd, K. Redmond, B. Ravnøy, and E. Sweetman. Body site matters: an evaluation and application of a novel histological methodology on the quantification of mucous cells in the skin of Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 36(2):115–127, February 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pazdzior:2019:GDA

[PPSW19] Ewa Paździor, Agnieszka Pekala-Safińska, and Dariusz Wasyl. Genotypic diversity among Shewanella spp. collected from freshwater fish. Journal of Fish Diseases, 42(5):677– 684, May 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pridgeon:2010:EPT

[PRSK10] J. W. Pridgeon, R. Russo, C. A. Shoemaker, and P. H. Klesius. Expression profiles of toll-like receptors in anterior kidney of channel catfish, Ictalurus punctatus (Rafinesque), acutely infected by Edwardsiella ictaluri. Journal of Fish Diseases, 33(6):497–505, June 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Peeler:2014:ADF

[PRT^{+}14] E. J. Peeler, D. Ryder, M. A. Thrush, J. Mewett, J. Hulland, and S. W. Feist. Acute dermatitis in farmed trout: an emerging disease. Journal of Fish Diseases, 37(12):1021– 1029, December 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Phillips:2017:EII

[PRWG17] A. C. N. Phillips, S. R. Reichley, C. Ware, and M. J. Griffin. Edwardsiella ictaluri infection in Pangasius catfish imported from West Bengal into the Southern Caribbean. Journal of Fish Diseases, 40(6):743–756, June 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pomaranski:2018:CST

[PRY^{+}18] E. K. Pomaranski, S. R. Reichley, R. Yanong, J. Shelley, D. B. Pouder, J. C. Wolf, K. V. Kenelty, B. Van Bonn, F. Oliaro, B. Byrne, K. A. Clothier, M. J. Griffin, A. C. Camus, and E. Soto. Characterization of spaC-type Erysipelothrix sp. isolates causing systemic disease in ornamental fish. Journal of Fish Diseases, 41(1):49–60, January 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Perez-Sanchez:2011:ICL

[PSBG^{+}11] T. Pérez-Sánchez, J. L. Balcázar, Y. García, N. Halaihel, D. Vendrell, I. de Blas, D. L. Merrifield, and I. Ruiz-Zarzuela. Identification and characterization of lactic acid bacteria isolated from rainbow trout, Oncorhynchus mykiss (Walbaum), with inhibitory activity against Lactococcus garvieae. Journal of Fish Diseases, 34(7):499–507, July 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Perez-Sancho:2018:LPM

[PSCFB^{+}18] M. Pérez-Sancho, I. Cerdá, A. Fernández-Bravo, L. Domínguez, M. J. Figueras, J. F. Fernández-Garayzábal, and A. I. Vela. Limited performance of MALDI-TOF for identification of fish Aeromonas isolates at species level. Journal of Fish Diseases, 41(10):1485–1493, October 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pribyl:2012:RPB

[PSK^{+}12] A. L. Pribyl, C. B. Schreck, M. L. Kent, K. M. Kelley, and S. J. Parker. Recovery potential of black rockfish, Sebastes melanops Girard, recompressed following barotrauma. Journal of Fish Diseases, 35(4):275–286, April 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Patil:2013:DMA

[PSK^{+}13] R. Patil, K. M. Shankar, B. T. N. Kumar, A. Kulkarni, P. Patil, and N. Moger. Development of a monoclonal antibody-based flow-through immunoassay (FTA) for detection of white spot syndrome virus (WSSV) in black tiger shrimp Penaeus monodon. Journal of Fish Diseases, 36(9): 753–762, September 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pittman:2011:NMQ

[PSR^{+}11] K. Pittman, P. Sourd, B. Ravnøy, Ø. Espeland, I. U. Fiksdal, T. Oen, A. Pittman, K. Redmond, and J. Sweetman. Novel method for quantifying salmonid mucous cells. Journal of Fish Diseases, 34(12):931–936, December 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Phillips:2017:OLM

[PSS17] A. C. N. Phillips, R. Suepaul, and E. Soto. Ocular localization of mycobacterial lesions in tank-reared juvenile cobia, Rachycentron canadum. Journal of Fish Diseases, 40(12): 1799–1804, December 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pascoli:2016:BAI

[PST^{+}16] F. Pascoli, M. Serra, M. Toson, T. Pretto, and A. Toffan. Betanodavirus ability to infect juvenile European sea bass, Dicentrarchus labrax, at different water salinity. Journal of Fish Diseases, 39(9):1061–1068, September 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Perez-Sancho:2018:FAM

[PSVK^{+}18] M. Pérez-Sancho, A. I. Vela, M. Kostrzewa, L. Zamora, A. Casamayor, L. Domínguez, and J. F. Fernández-Garayzábal. First analysis by MALDI-TOF MS technique of Chryseobacterium species relevant to aquaculture. Journal of Fish Diseases, 41(2):389–393, February 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Passantino:2014:LMC

[PSZ^{+}14] L. Passantino, N. Santamaria, R. Zupa, C. Pousis, R. Garofalo, A. Cianciotta, E. Jirillo, F. Acone, and A. Corriero. Liver melanomacrophage centres as indicators of Atlantic bluefin tuna, Thunnus thynnus L. well-being. Journal of Fish Diseases, 37(3):241–250, March 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Padhi:2019:KHC

[PTM^{+}19] Soumesh K. Padhi, Isaiah Tolo, Margaret McEachran, Alexander Primus, Sunil K. Mor, and Nicholas B. D. Phelps. Koi herpesvirus and carp oedema virus: Infections and coinfections during mortality events of wild common carp in the United States. Journal of Fish Diseases, 42(11):1609–1621, November 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Parra:2018:OEA

[PVIM18] M. Parra, B. Valenzuela, M. Imarai, and B. Modak. Obtainment and evaluation of adjuvant effect of the aromatic geranyl derivative Filifolinone in bacterin of Piscirickettsia salmonis. Journal of Fish Diseases, 41(1):157–159, January 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pao:2018:CI

[PWHW18a] H. Y. Pao, C. Y. Wu, C. H. Huang, and C. M. Wen. Cover image. Journal of Fish Diseases, 41(9):i, September 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pao:2018:DCV

[PWHW18b] H. Y. Pao, C. Y. Wu, C. H. Huang, and C. M. Wen. Development, characterization and virus susceptibility of a continuous cell line from the caudal fin of marbled eel (Anguilla marmorata). Journal of Fish Diseases, 41(9):1331– 1338, September 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pao:2019:PDA

[PWW19] Hui-Yu Pao, Cheng-Yu Wu, and Chiu Ming Wen. Persistent development of adomavirus and aquareovirus in a novel cell line from marbled eel with petechial skin haemorrhage. Journal of Fish Diseases, 42(3):345–355, March 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pierezan:2019:III

[PYSS19] Felipe Pierezan, Susan Yun, Win Surachetpong, and Esteban Soto. Intragastric and intracoelomic injection challenge models of tilapia lake virus infection in Nile tilapia (Oreochromis niloticus L.) fingerlings. Journal of Fish Diseases, 42(9):1301–1307, September 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Park:2014:VVE

[PZL^{+}14] K. H. Park, S-R. Zeon, J-G. Lee, S-H. Choi, Y. K. Shin, and K-I. Park. In vitro and in vivo efficacy of drugs against the protozoan parasite Azumiobodo hoyamushi that causes soft tunic syndrome in the edible ascidian Halocynthia roretzi (Drasche). Journal of Fish Diseases, 37(4):309–317, April 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Quaglio:2016:NGD

[QPB^{+}16] F. Quaglio, A. Perolo, P. Bronzatti, A. Gustinelli, V. Menconi, G. Cavazza, M. Caffara, A. Manfrin, E. Gallo, and M. L. Fioravanti. Nodular gill disease in farmed rainbow trout (Oncorhynchus mykiss) in Italy. Journal of Fish Diseases, 39(9):1139–1142, September 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Qin:2017:BIV

[QWTL17] G. Qin, X. Wang, S. Tan, and Q. Lin. A bacterial infection by Vibrio harveyi causing heavy reduction of cultured lined seahorse Hippocampus erectus. Journal of Fish Diseases, 40 (4):601–605, April 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Qin:2014:EFT

[QXW14] L. Qin, J. Xu, and Y. G. Wang. Edwardsiellosis in farmed turbot, Scophthalmus maximus (L.), associated with an unusual variant of Edwardsiella tarda: a clinical, aetiological and histopathological study. Journal of Fish Diseases, 37 (2):103–111, February 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Qin:2018:CI

[QZZ^{+}18a] Geng Qin, Yuan Zhang, Bo Zhang, Xin Wang, Jianping Yin, and Qiang Lin. Cover image. Journal of Fish Diseases, 41 (12):i, December 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Qin:2018:SGT

[QZZ^{+}18b] Geng Qin, Yuan Zhang, Bo Zhang, Xin Wang, Jianping Yin, and Qiang Lin. Seahorse TLR5 gene responses to Vibrio vulnificus infection, which in combination with scuticociliates causes heavy reductions in seahorse aquaculture. Journal of Fish Diseases, 41(12):1933–1936, December 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Romanucci:2017:OTA

[RAM^{+}17] M. Romanucci, A. Arbuatti, M. Massimini, S. V. P. Defourny, and L. Della Salda. Ovarian teratoma in an adult female Zoogoneticus tequila (Webb & Miller 1998): histological and immunohistochemical features. Journal of Fish Diseases, 40(6):859–862, June 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Radosavljevic:2018:OTN

[RAM^{+}18] V. Radosavljevic, M. Adamek, V. Milicevic, J. Maksimovic-Zoric, and D. Steinhagen. Occurrence of two novel viral pathogens (CEV and CyHV-2) affecting Serbian cyprinid aquaculture and ichthyofauna. Journal of Fish Diseases, 41 (5):851–854, May 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rakus:2019:EZD

[RAM^{+}19] Krzysztof Rakus, Mikołaj Adamek, Miriam Mojżesz, Piotr Podlasz, Małgorzata Chmielewska-Krzesińska, Karolina Naumowicz, Natalia Kasica-Jarosz, Katarzyna Kłak, Sebastian Rakers, Keith Way, Dieter Steinhagen, and Magdalena Chadzińska. Evaluation of zebrafish (Danio rerio) as an animal model for the viral infections of fish. Journal of Fish Diseases, 42(6):923–934, June 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roberts:2010:HSP

[RAS^{+}10] R. J. Roberts, C. Agius, C. Saliba, P. Bossier, and Y. Y. Sung. Heat shock proteins (chaperones) in fish and shellfish and their potential role in relation to fish health: a review. Journal of Fish Diseases, 33(10):789–801, October 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ramilo:2018:MML

[RAVC18] A. Ramilo, E. Abollo, A. Villalba, and M. J. Carballal. A Minchinia mercenariae-like parasite infects cockles Cerastoderma edule in Galicia (NW Spain). Journal of Fish Diseases, 41(1):41–48, January 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Reid:2015:CTG

[RB15] H. I. Reid and T. H. Birkbeck. Characterization of two groups of Pasteurella skyensis isolates from Atlantic salmon, Salmo salar L., based on serotype and 16S rRNA and rpoB gene partial sequences. Journal of Fish Diseases, 38(4):405– 408, April 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Reyes-Becerril:2017:VVS

[RBAT^{+}17] M. Reyes-Becerril, E. Alamillo, A. Trasviña, I. Hirono, H. Kondo, W. Jirapongpairoj, F. Ascencio-Valle, and C. Angulo. In vivo and in vitro studies using larval and adult antigens from Neobenedenia melleni on immune response in yellowtail (Seriola lalandi). Journal of Fish Diseases, 40(11): 1497–1509, November 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Robbins:2014:FRP

[RBF14] Rachel Robbins, Barry Bruce, and Andrew Fox. First reports of proliferative lesions in the great white shark, Carcharodon carcharias L., and bronze whaler shark, Carcharhinus brachyurus Günther. Journal of Fish Diseases, 37(11):997– 1000, November 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Reichert:2017:AAE

[RBH^{+}17] M. Reichert, S. M. Bergmann, J. Hwang, R. Buchholz, and C. Lindenberger. Antiviral activity of exopolysaccharides from Arthrospira platensis against koi herpesvirus. Journal of Fish Diseases, 40(10):1441–1450, October 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ronza:2011:MCC

[RBL^{+}11] P. Ronza, R. Bermúdez, A. P. Losada, A. Robles, and M. I. Quiroga. Mucosal CD3ε + cell proliferation and gut epithelial apoptosis: implications in rainbow trout gastroenteritis (RTGE). Journal of Fish Diseases, 34(6):433–443, June 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Reichert:2016:RMC

[RBM^{+}16] M. Reichert, E. Borzym, M. Matras, J. Maj-Paluch, M. Stachnik, and M. Palusinska. Down-regulation of MHC class I mRNA expression in the course of KHV infection. Journal of Fish Diseases, 39(10):1253–1256, October 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rodriguez:2019:HCN

[RBR19] Helena Rodríguez, Rafael Bañón, and Andrea Ramilo. The hidden companion of non-native fishes in north-east Atlantic waters. Journal of Fish Diseases, 42(7):1013–1021, July 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rizgalla:2016:UDW

[RBS^{+}16] J. Rizgalla, J. E. Bron, A. P. Shinn, T. K. Herath, G. Paladini, and H. W. Ferguson. Ulcerative dermatitis in wild dusky grouper Epinephelus marginatus (Lowe) from Libyan waters. Journal of Fish Diseases, 39(12):1457–1466, December 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ruane:2015:OFW

[RBWR^{+}15] N. M. Ruane, M. Bolton-Warberg, H. D. Rodger, D. J. Colquhoun, M. Geary, S. J. McCleary, K. O’Halloran, K. Maher, D. O’Keeffe, L. Mirimin, K. Henshilwood, F. Geoghegan, and R. D. Fitzgerald. An outbreak of francisellosis in wild-caught Celtic Sea Atlantic cod, Gadus morhua L., juveniles reared in captivity. Journal of Fish Diseases, 38(1): 97–102, January 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ronneseth:2017:CAV

[RCD^{+}17] A. Rønneseth, D. Castillo, P. D’Alvise, Ø. Tønnesen, G. Haugland, T. Grotkjær, K. Engell-Sørensen, L. Nørremark, Ø. Bergh, H. I. Wergeland, and L. Gram. Comparative assessment of Vibrio virulence in marine fish larvae. Journal of Fish Diseases, 40(10):1373–1385, October 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rojas:2015:PIC

[RCGGM15] V. Rojas, H. Camus-Guerra, F. Guzmán, and L. Mercado. Pro-inflammatory caspase-1 activation during the immune response in cells from rainbow trout Oncorhynchus mykiss (Walbaum 1792) challenged with pathogen-associated molecular patterns. Journal of Fish Diseases, 38(11):993–1003, November 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Remany:2012:EPW

[RDN^{+}12] M. C. Remany, C. Daly, S. Nagaraj, A. K. Panda, K. Jaideep, and Y. C. T. Samraj. Evidence for the presence of white spot syndrome virus (WSSV) and monodon baculovirus (MBV) in wild Penaeus monodon (Fabricius) broodstock, in the southeast coast of India. Journal of Fish Diseases, 35(11):793–798, November 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ravasi:2018:MST

[RDW18] Damiana Ravasi, Sophie De Respinis, and Thomas Wahli. Multilocus sequence typing reveals clonality in Saprolegnia parasitica outbreaks. Journal of Fish Diseases, 41(11):1653– 1665, November 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rozas:2014:PPS

[RE14] M. Rozas and R. Enríquez. Piscirickettsiosis and Piscirickettsia salmonis in fish: a review. Journal of Fish Diseases, 37 (3):163–188, March 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rehulka:2013:TTN

[Řeh13] J. Řehulka. Testicular tumour in northern pike, Esox lucius L. Journal of Fish Diseases, 36(7):669–673, July 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Righetti:2014:MSI

[RFA^{+}14] M. Righetti, L. Favaro, E. Antuofermo, M. Caffara, S. Nuvoli, T. Scanzio, and M. Prearo. Mycobacterium salmoniphilum infection in a farmed Russian sturgeon, Acipenser gueldenstaedtii (Brandt & Ratzeburg). Journal of Fish Diseases, 37(7):671–674, July 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rekecki:2012:BHI

[RGD^{+}12] A. Rekecki, R. A. Y. S. A. Gunasekara, K. Dierckens, S. Laureau, N. Boon, H. Favoreel, M. Cornelissen, P. Sorgeloos, R. Ducatelle, P. Bossier, and W. Van den Broeck. Bacterial host interaction of GFP-labelled Vibrio anguillarum HI-610 with gnotobiotic sea bass, Dicentrarchus labrax (L.), larvae. Journal of Fish Diseases, 35(4):265–273, April 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rojas:2013:EES

[RGD^{+}13] M. E. Rojas, M. Galleguillos, S. Díaz, A. Machuca, A. Carbonero, and P. A. Smith. Evidence of exotoxin secretion of Piscirickettsia salmonis, the causative agent of piscirickettsiosis. Journal of Fish Diseases, 36(8):703–709, August 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roca-Gerones:2018:MGC

[RGMGG^{+}18] Xavier Roca-Geronès, Isabel Montoliu, Carla Godínez-González, Roser Fisa, and Shokoofeh Shamsi. Morphological and genetic characterization of Hysterothylacium Ward & Magath, 1917 (Nematoda: Raphidascarididae) larvae in horse mackerel, blue whiting and anchovy from Spanish Atlantic and Mediterranean waters. Journal of Fish Diseases, 41(10):1463–1475, October 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rahmati-holasoo:2015:IMA

[RhAS^{+}15] H. Rahmati-holasoo, M. Alishahi, S. Shokrpoor, A. Jangarannejad, and B. Mohammadian. Invasion of melanoma to angiolipoma in a male Siamese fighting fish, Betta splendens, Regan. Journal of Fish Diseases, 38(10):925–930, October 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ryerse:2015:PDF

[RHB^{+}15] I. A. Ryerse, J. M. Hooft, D. P. Bureau, M. A. Hayes, and J. S. Lumsden. Purified deoxynivalenol or feed restriction reduces mortality in rainbow trout, Oncorhynchus mykiss (Walbaum), with experimental bacterial coldwater disease but biologically relevant concentrations of deoxynivalenol do not impair the growth of Flavobacterium psychrophilum. Journal of Fish Diseases, 38(9):809–819, September 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rose:2013:IOF

[RHBMM13] S. Rose, R. Hill, L. E. Bermudez, and T. Miller-Morgan. Imported ornamental fish are colonized with antibiotic-resistant bacteria. Journal of Fish Diseases, 36(6):533–542, June 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rahmati-holasoo:2015:CLK

[RhMM^{+}15] H. Rahmati-holasoo, M. Masoudifard, H. Ebrahimzadeh Mousavi, S. Shokrpoor, A. Tavakkoli, and M. S. Farazandemehr. Cystic lesions in the kidney of flower horn fish, hybrid cichlid. Journal of Fish Diseases, 38(9):833–838, September 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rahmati-holasoo:2015:PLF

[RhMV^{+}15] H. Rahmati-holasoo, H. Ebrahimzadeh Mousavi, A. Vajhi, S. Shokrpoor, A. Tavakkoli, M. A. Mirdamadi, and S. Fayyaz. Polycystic liver in flower horn fish, hybrid cichlid. Journal of Fish Diseases, 38(3):325–328, March 2015.

CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Raghavendra:2012:ACB

[RHS^{+}12] A. Raghavendra, K. P. Hemaprasanth, R. Singh, N. Sridhar, V. Kumar, and M. R. Raghunath. Ammonium chloride bath treatment as a quarantine measure to prevent spread of Lernaea cyprinacea infection during transfer of fish from affected ponds. Journal of Fish Diseases, 35(3):243–247, March 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rahmati-Holasoo:2018:TOR

[RHSM^{+}18a] H. Rahmati-Holasoo, S. Shokrpoor, M. Masoudifard, S. Davudypoor, and M. Vaseghi. Telangiectatic osteosarcoma and renal adenocarcinoma in an Oscar (Astronotus ocellatus, Agassiz): Diagnostic imaging and immunohistochemical study. Journal of Fish Diseases, 41(7):1165–1172, July 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rahmati-Holasoo:2018:NBS

[RHSM^{+}18b] H. Rahmati-Holasoo, M. Soltani, M. Masoudifard, S. Shokrpoor, A. Taheri Mirghaed, and M. Ahmadpoor. Nephroblastoma in bester sturgeon, a cultured hybrid of Huso huso × Acipenser ruthenus: Diagnostic imaging, clinical and histopathological study. Journal of Fish Diseases, 41(7): 1093–1101, July 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rahmati-Holasoo:2016:LIL

[RHST^{+}16] H. Rahmati-Holasoo, S. Shokrpoor, A. Tavakkoli, A. Vajhi, and H. Ebrahimzadeh Mousavi. Liposarcoma or invasive lipomatosis in flower horn fish, hybrid cichlid: clinical, radiological, ultrasonographical and histopathological study. Journal of Fish Diseases, 39(3):309–315, March 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rairat:2019:BC

[RHTC19a] Tirawat Rairat, Chia-Yu Hsieh, Wipavee Thongpiam, and Chi-Chung Chou. Back cover. Journal of Fish Diseases, 42(8):ii, August 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rairat:2019:PPM

[RHTC19b] Tirawat Rairat, Chia-Yu Hsieh, Wipavee Thongpiam, and Chi-Chung Chou. Pharmacokinetic–pharmacodynamic modelling for the determination of optimal dosing regimen of florfenicol in Nile tilapia (Oreochromis niloticus) at different water temperatures and antimicrobial susceptibility levels. Journal of Fish Diseases, 42(8):1181–1190, August 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Riera-Heredia:2019:FC

[RHVG^{+}19a] Natàlia Riera-Heredia, Emilio J. Vélez, Joaquim Gutiérrez, Isabel Navarro, and Encarnación Capilla. Front cover. Journal of Fish Diseases, 42(8):i, August 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Riera-Heredia:2019:GEA

[RHVG^{+}19b] Natàlia Riera-Heredia, Emilio J. Vélez, Joaquim Gutiérrez, Isabel Navarro, and Encarnación Capilla. Gene expression analyses in malformed skeletal structures of gilthead sea bream (Sparus aurata). Journal of Fish Diseases, 42(8): 1169–1180, August 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rahmati-Holasoo:2016:FDK

[RHZA^{+}16] H. Rahmati-Holasoo, A. Zargar, S. Ahmadivand, S. Shokrpoor, S. Ezhari, and H. A. Ebrahimzadeh Mousavi. First detection of koi herpesvirus from koi, Cyprinus carpio L. experiencing mass mortalities in Iran: clinical, histopathological and molecular study. Journal of Fish Diseases, 39(10): 1153–1163, October 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rehulka:2016:CSA

[ŘKH16] J. Řehulka, A. Kubátová, and V. Hubka. Cephalotheca sulfurea (Ascomycota, Sordariomycetes), a new fungal pathogen of the farmed rainbow trout Oncorhynchus mykiss. Journal of Fish Diseases, 39(12):1413–1419, December 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rehulka:2017:DID

[ŘKH17] J. Řehulka, M. Kolařík, and V. Hubka. Disseminated infection due to Exophiala pisciphila in Cardinal tetra, Paracheirodon axelrodi. Journal of Fish Diseases, 40(8): 1015–1024, August 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rehulka:2018:SBM

[ŘKH18] J. Řehulka, A. Kubátová, and V. Hubka. Swim bladder mycosis in pretty tetra (Hemigrammus pulcher) caused by Exophiala pisciphila and Phaeophleospora hymenocallidicola, and experimental verification of pathogenicity. Journal of Fish Diseases, 41(3):487–500, March 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roh:2019:WGA

[RKK^{+}19] Heyong Jin Roh, Bo-Seong Kim, Ahran Kim, Nam Eun Kim, Yoonhang Lee, Won-Kyong Chun, Tho Diem Ho, and Do-Hyung Kim. Whole-genome analysis of multi-drug-resistant Aeromonas veronii isolated from diseased discus (Symphysodon discus) imported to Korea. Journal of Fish Diseases, 42(1):147–153, January 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rosser:2016:VDH

[RKPG16] T. G. Rosser, L. H. Khoo, L. M. Pote, and M. J. Griffin. Verrucous dermal henneguyosis associated with Henneguya exilis (Kudo, 1929) (Cnidaria: Myxobolidae), a parasite of the channel catfish Ictalurus punctatus (Rafinesque, 1818). Journal of Fish Diseases, 39(10):1263–1267, October 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rezaie:2012:PLG

[RMAM12] A. Rezaie, S. M. Mousavi, E. Ahmadmoradi, and F. Mohammadi. Polycystic liver in a goldfish, Carassius auratus (L.). Journal of Fish Diseases, 35(10):785–787, October 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rodger:2014:CCS

[RMR14] H. D. Rodger, S. J. McCleary, and N. M. Ruane. Clinical cardiomyopathy syndrome in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 37(10):935–939, October 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Reichert:2013:TPM

[RMS^{+}13] M. Reichert, M. Matras, H. F. Skall, N. J. Olesen, and S. Kahns. Trade practices are main factors involved in the transmission of viral haemorrhagic septicaemia. Journal of Fish Diseases, 36(2):103–114, February 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ruane:2009:MDI

[RMSH09] N. M. Ruane, L. J. McCarthy, D. Swords, and K. Henshilwood. Molecular differentiation of infectious pancreatic necrosis virus isolates from farmed and wild salmonids in Ireland. Journal of Fish Diseases, 32(12):979–987, December 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic). See erratum [TCC10b].

Ruane:2010:E

[RMSH10] N. M. Ruane, L. J. McCarthy, D. Swords, and K. Henshilwood. Erratum. Journal of Fish Diseases, 33(7):623, July 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Raj:2012:HEW

[RNM^{+}12] N. Sundar Raj, K. S. Nathiga Nambi, S. Abdul Majeed, G. Taju, S. Vimal, M. A. Farook, and A. S. Sahul Hameed. High efficacy of white spot syndrome virus replication in tissues of freshwater rice-field crab, Paratelphusa hydrodomous (Herbst). Journal of Fish Diseases, 35(12):917–925, December 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roberts:2010:BR

[Rob10] R. J. Roberts. Book review. Journal of Fish Diseases, 33(3): 283, March 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roberts:2011:DAE

[Rob11] R. J. Roberts. Dr Anthony Ellis (1947–2010). Journal of Fish Diseases, 34(8):577–578, August 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roberts:2013:E

[Rob13] Ronald J. Roberts. Editorial. Journal of Fish Diseases, 36 (12):973, December 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roberts:2014:DRW

[Rob14a] Ronald J. Roberts. Dr. Rodney Wootten (1946–2013). Journal of Fish Diseases, 37(1):1–2, January 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roberts:2014:E

[Rob14b] Ronald J. Roberts. Editorial. Journal of Fish Diseases, 37 (12):1001, December 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ramos-Paredes:2017:VGW

[RPGCIG17] J. Ramos-Paredes, J. M. Grijalva-Chon, and J. C. Ibarra-Gámez. Virulence and genotypes of white spot syndrome virus infecting Pacific white shrimp Litopenaeus vannamei in north-western Mexico. Journal of Fish Diseases, 40(3): 425–435, March 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ross:2018:FDG

[RPJ^{+}18] P. M. Ross, A. Pande, J. B. Jones, J. Cope, and G. Flowers. First detection of gas bubble disease and Rickettsialike organisms in Paphies ventricosa, a New Zealand surf clam. Journal of Fish Diseases, 41(1):187–190, January 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rupp:2019:SIE

[RPM^{+}19] Melanie Rupp, Paola Pilo, Barbara Müller, Ralph Knüsel, Beat von Siebenthal, Joachim Frey, Paul-Daniel Sindilariu, and Heike Schmidt-Posthaus. Systemic infection in European perch with thermoadapted virulent Aeromonas salmonicida (Perca fluviatilis). Journal of Fish Diseases, 42 (5):685–691, May 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ramirez-Paredes:2019:WCI

[RPMRLJ^{+}19] José Gustavo Ramírez-Paredes, Miguel Angel Mendoza-Roldan, Benjamin Lopez-Jimena, Khalid Shahin, Matthijs Metselaar, Kim D. Thompson, David J. Penman, Randolph H. Richards, and Alexandra Adams. Whole cell inactivated autogenous vaccine effectively protects red Nile tilapia (Oreochromis niloticus) against francisellosis via intraperitoneal injection. Journal of Fish Diseases, 42(8):1191–1200, August 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rangel:2017:LCI

[RRC^{+}17] L. F. Rangel, S. Rocha, G. Casal, R. Castro, R. Severino, C. Azevedo, F. Cavaleiro, and M. J. Santos. Life cycle inference and phylogeny of Ortholinea labracis n. sp. (Myxosporea: Ortholineidae), a parasite of the European seabass Dicentrarchus labrax (Teleostei: Moronidae), in a Portuguese fish farm. Journal of Fish Diseases, 40(2):243– 262, February 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rosaeg:2019:EPD

[RRG^{+}19] Magnus Vikan Røsæg, Espen Rimstad, Arne Guttvik, Bård Skjelstad, Eldar Åsgard Bendiksen, and Åse Helen Garseth. Effect of pancreas disease caused by SAV 2 on protein and fat digestion in Atlantic salmon. Journal of Fish Diseases, 42 (1):97–108, January 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rekecki:2013:LUV

[RRO^{+}13] A. Rekecki, E. Ringø, R. Olsen, R. Myklebust, K. Dierckens, Ø. Bergh, S. Laureau, M. Cornelissen, R. Ducatelle, A. Decostere, P. Bossier, and W. Van den Broeck. Luminal uptake of Vibrio (Listonella) anguillarum by shed enterocytes — a novel early defence strategy in larval fish. Journal of Fish Diseases, 36(4):419–426, April 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rasmussen:2013:DOI

[RSB13] K. J. Rasmussen, J. F. Steffensen, and K. Buchmann. Differential occurrence of immune cells in the primary and secondary vascular systems in rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 36(7):675– 679, July 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rizgalla:2017:NUS

[RSF^{+}17] J. Rizgalla, A. P. Shinn, H. W. Ferguson, G. Paladini, N. S. Jayasuriya, and J. E. Bron. A novel use of social media to evaluate the occurrence of skin lesions affecting wild dusky grouper, Epinephelus marginatus (Lowe, 1834), in Libyan coastal waters. Journal of Fish Diseases, 40(5):609–620, May 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ramya:2014:DCC

[RSGB^{+}14] V. L. Ramya, R. Sharma, P. Gireesh-Babu, S. R. Patchala, A. Rather, P. C. Nandanpawar, and S. Eswaran. Development of chitosan conjugated DNA vaccine against nodavirus in Macrobrachium rosenbergii (De Man, 1879). Journal of Fish Diseases, 37(9):815–824, September 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roiha:2011:EFT

[RSH11] I. S. Roiha, O. B. Samuelsen, and T. Harboe. Efficacy of florfenicol in the treatment of bacterial infections in halibut, Hippoglossus hippoglossus (L.), larvae. Journal of Fish Diseases, 34(12):927–930, December 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rozas-Serri:2017:CPP

[RSIP^{+}17] M. Rozas-Serri, R. Ildefonso, A. Peña, R. Enríquez, S. Barrientos, and L. Maldonado. Comparative pathogenesis of piscirickettsiosis in Atlantic salmon (Salmo salar L.) post-smolt experimentally challenged with LF-89-like and EM-90-like Piscirickettsia salmonis isolates. Journal of Fish Diseases, 40(10):1451–1472, October 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ruane:2018:ISA

[RSM^{+}18] Neil M. Ruane, David Swords, Teresa Morrissey, Michelle Geary, Cathy Hickey, Evelyn M. Collins, Fiona Geoghegan, and Fiona Swords. Isolation of salmonid alphavirus subtype 6 from wild-caught ballan wrasse, Labrus bergylta (Ascanius). Journal of Fish Diseases, 41(11):1643–1651, November 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rosaeg:2017:BWL

[RSP^{+}17] M. V. Røsæg, H. Sindre, D. Persson, O. Breck, D. Knappskog, A. B. Olsen, and T. Taksdal. Ballan wrasse (Labrus bergylta Ascanius) is not susceptible to pancreas disease caused by salmonid alphavirus subtype 2 and 3. Journal of Fish Diseases, 40(7):975–978, July 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rozas-Serri:2018:CGE

[RSPA^{+}18] M. Rozas-Serri, A. Peña, G. Arriagada, R. Enríquez, and L. Maldonado. Comparison of gene expression in post-smolt Atlantic salmon challenged by LF-89-like and EM-90-like Piscirickettsia salmonis isolates reveals differences in the immune response associated with pathogenicity. Journal of Fish Diseases, 41(3):539–552, March 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rousselet:2018:SET

[RSR^{+}18] Estelle Rousselet, Nicole I. Stacy, David S. Rotstein, Tom B. Waltzek, Matt J. Griffin, and Ruth Francis-Floyd. Systemic Edwardsiella tarda infection in a Western African lungfish (Protopterus annectens) with cytologic observation of heterophil projections. Journal of Fish Diseases, 41(9):1453– 1458, September 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Russell:2015:AAN

[RTM^{+}15] S. Russell, L. Tubbs, D. J. McLelland, V. LePage, K. M. Young, P. Huber, and J. S. Lumsden. Amyloid associated with neoplasia in two captive tricolour sharkminnows Balantiocheilus melanopterus Bleeker. Journal of Fish Diseases, 38(6):561–565, June 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ramos:2013:SML

[RVB13] P. Ramos, P. Victor, and S. Branco. Spontaneous melanotic lesions in axillary seabream, Pagellus acarne (Risso). Journal of Fish Diseases, 36(9):769–777, September 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rouleau:2018:GPC

[RVC18] F. D. Rouleau, A. T. Vincent, and S. J. Charette. Genomic and phenotypic characterization of an atypical Aeromonas salmonicida strain isolated from a lumpfish and producing unusual granular structures. Journal of Fish Diseases, 41 (4):673–681, April 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roberts:2010:E

[RW10] R. J. Roberts and Rodney Wootten. Editorial. Journal of Fish Diseases, 33(2):93, February 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roberts:2011:E

[RW11] Ronald J. Roberts and Rod Wootten. Editorial. Journal of Fish Diseases, 34(1):1–2, January 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rimmer:2017:SNA

[RWTB17] A. E. Rimmer, R. J. Whittington, A. Tweedie, and J. A. Becker. Susceptibility of a number of Australian freshwater fishes to dwarf gourami iridovirus (Infectious spleen and kidney necrosis virus). Journal of Fish Diseases, 40(3):293– 310, March 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sokolowski:2012:IAS

[SAD^{+}12] M. S. Sokolowski, B. A. Allam, K. J. Dunton, M. A. Clark, E. B. Kurtz, and M. D. Fast. Immunophysiology of Atlantic sturgeon, Acipenser oxyrinchus oxyrinchus (Mitchill), and the relationship to parasitic copepod, Dichelesthium oblongum (Abilgaard) infection. Journal of Fish Diseases, 35 (9):649–660, September 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Samuelsen:2010:SDP

[Sam10] O. B. Samuelsen. A single-dose pharmacokinetic study of emamectin benzoate in cod, Gadus morhua L., held in sea water at 9 ^{o}C. Journal of Fish Diseases, 33(2):137–142, February 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Samerpitak:2019:OGI

[SAS^{+}19] Kittipan Samerpitak, Anders Alfjorden, Seyedmojtaba Seyedmousavi, Antonius J. M. M. Rijs, Albertus Hendrikus Gerardus Gerrits van den Ende, and Gert Sybren de Hoog. Ochroconis globalis infecting Atlantic salmon (Salmo salar), with a review of Ochroconis species in cold-blooded animals. Journal of Fish Diseases, 42(6):947–957, June 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Scholz:2011:KJC

[SBD11] T. Scholz, T. T. Binh, and B. S. Dezfuli. Khawia japonensis (Cestoda: Caryophyllidea): another invasive parasite of carp, Cyprinus carpio L., imported to Europe. Journal of Fish Diseases, 34(12):943–949, December 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Simons:2016:CDL

[SBH^{+}16] J. Simons, D. W. Bruno, Y-M. Ho, W. Murray, and I. Matejusova. Common dab, Limanda limanda (L.), as a natural carrier of salmonid alphavirus (SAV) from waters off northwest Ireland. Journal of Fish Diseases, 39(4):507–510, April 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stilwell:2018:GGA

[SBHC18] J. M. Stilwell, S. M. Boylan, S. Howard, and A. C. Camus. Gas gland adenoma in a lined seahorse, Hippocampus erectus, Perry 1810. Journal of Fish Diseases, 41(1):171–174, January 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Soto:2013:PMS

[SBK^{+}13] E. Soto, L. Bogdanovic, R. C. Krecek, J. A. Jovonovich, M. Arauz, and R. M. Overstreet. Parasitosis of Metabronema sp. (Nematoda: Cystidicolidae) in Caribbean porcupinefish, Diodon hystrix (L.). Journal of Fish Diseases, 36(12):1031– 1034, December 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sharon:2015:LCA

[SBK^{+}15] G. Sharon, D. Benharroch, L. Kachko, N. Reis-Hevlin, and D. Zilberg. Liposarcoma in clownfish, Amphiprion ocellaris Cuvier, produced in indoor aquaculture. Journal of Fish Diseases, 38(6):575–580, June 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Seguin:2011:IPS

[SBM^{+}11] G. Séguin, F. Bouchard, L. N. Measures, C. F. Uhland, and S. Lair. Infections with Philometra sp. associated with mortalities in wild-hatched captive-raised striped bass, Morone saxatilis (Walbaum). Journal of Fish Diseases, 34(6):475– 481, June 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Saraiva:2019:EAI

[SBP^{+}19] Marcia Saraiva, Max J. Beckmann, Sara Pflaum, Marianne Pearson, Daniel Carcajona, James W. Treasurer, and Pieter van West. Exophiala angulospora infection in hatcheryreared lumpfish (Cyclopterus lumpus) broodstock. Journal of Fish Diseases, 42(3):335–343, March 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Strom-Bestor:2011:IAP

[SBW11] M. Ström-Bestor and T. Wiklund. Inhibitory activity of Pseudomonas sp. on Flavobacterium psychrophilum, in vitro. Journal of Fish Diseases, 34(4):255–264, April 2011. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Schall:2018:STV

[SBW^{+}18] Megan K. Schall, Vicki S. Blazer, Heather L. Walsh, Geoffrey D. Smith, Timothy Wertz, and Tyler Wagner. Spatial and temporal variability of myxozoan parasite, Myxobolus inornatus, prevalence in young of the year smallmouth bass in the Susquehanna River Basin, Pennsylvania. Journal of Fish Diseases, 41(11):1689–1700, November 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sarowar:2019:SMP

[SCD19] Mohammad Nasif Sarowar, Roland Cusack, and James Duston. Saprolegnia molecular phylogeny among farmed teleosts in Nova Scotia, Canada. Journal of Fish Diseases, 42(12): 1745–1760, December 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Salama:2013:DAB

[SCF^{+}13] N. K. G. Salama, C. M. Collins, J. G. Fraser, J. Dunn, C. C. Pert, A. G. Murray, and B. Rabe. Development and assessment of a biophysical dispersal model for sea lice. Journal of Fish Diseases, 36(3):323–337, March 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Seo:2013:COE

[SCK13] J. Y. Seo, H. J. Chung, and T. J. Kim. Codon-optimized expression of fish iridovirus capsid protein in yeast and its application as an oral vaccine candidate. Journal of Fish Diseases, 36(9):763–768, September 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Smith:2015:IPS

[SCR^{+}15] P. A. Smith, J. R. Contreras, M. E. Rojas, A. Guajardo, S. Díaz, and A. Carbonero. Infectivity of Piscirickettsia salmonis in immersion-bath exposed rainbow trout Oncorhynchus mykiss (Walbaum) fry. Journal of Fish Diseases, 38(8):765–770, August 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sui:2012:EPH

[SCS^{+}12] L. Sui, J. Cai, H. Sun, M. Wille, and P. Bossier. Effect of poly-β-hydroxybutyrate on Chinese mitten crab, Eriocheir sinensis, larvae challenged with pathogenic Vibrio anguillarum. Journal of Fish Diseases, 35(5):359–364, May 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stilwell:2019:DLN

[SCZM19] Justin M. Stilwell, Alvin C. Camus, Trevor T. Zachariah, and Rita McManamon. Disseminated lymphoid neoplasia and hepatoblastoma in an Atlantic stingray, Hypanus sabinus (Lesueur 1824). Journal of Fish Diseases, 42(2):319–323, February 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Salama:2018:UBP

[SDI^{+}18] N. K. G. Salama, A. C. Dale, V. V. Ivanov, P. F. Cook, C. C. Pert, C. M. Collins, and B. Rabe. Using biological– physical modelling for informing sea lice dispersal in Loch Linnhe, Scotland. Journal of Fish Diseases, 41(6):901–919, June 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Seo:2016:OHR

[SDJH16] H-G. Seo, J. W. Do, S. H. Jung, and H-J. Han. Outbreak of hirame rhabdovirus infection in cultured spotted sea bass Lateolabrax maculatus on the western coast of Korea. Journal of Fish Diseases, 39(10):1239–1246, October 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Senapin:2019:MSD

[SDM^{+}19] Saengchan Senapin, Ha Thanh Dong, Watcharachai Meemetta, Warachin Gangnonngiw, Pakakul Sangsuriya, Rapeepun Vanichviriyakit, Molruedee Sonthi, and Bunlung Nuangsaeng. Mortality from scale drop disease in farmed Lates calcarifer in Southeast Asia. Journal of Fish Diseases, 42(1):119–127, January 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Silveira:2016:FRO

[SdMVS^{+}16] E. L. Silveira, A. de M. Viott, R. L. Santos, A. A. Macêdo, J. M. R. Aranha, and M. S. Menezes. First report of ovarian epidermoid cyst in a teleost. Journal of Fish Diseases, 39 (3):385–388, March 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Smith:2015:EPS

[SDR^{+}15] P. A. Smith, F. E. Díaz, M. E. Rojas, S. Díaz, M. Galleguillos, and A. Carbonero. Effect of Piscirickettsia salmonis inoculation on the ASK continuous cell line. Journal of Fish Diseases, 38(3):321–324, March 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Saticioglu:2018:SGC

[SDWA18] I. B. Saticioglu, M. Duman, T. Wiklund, and S. Altun. Serological and genetic characterization of Flavobacterium psychrophilum isolated from farmed salmonids in Turkey. Journal of Fish Diseases, 41(12):1899–1908, December 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shamsi:2019:GMC

[SDZN19] Shokoofeh Shamsi, Mai Dang, Xiaocheng Zhu, and Barbara Nowak. Genetic and morphological characterization of Mawsonascaris vulvolacinata n. sp. (Nematoda: Anisakidae) and associated histopathology in a wild caught cowtail stingray, Pastinachus ater. Journal of Fish Diseases, 42(7):1047–1056, July 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Smith:2016:ECV

[SEK^{+}16] P. Smith, R. Endris, G. Kronvall, V. Thomas, D. Verner-Jeffreys, C. Wilhelm, and I. Dalsgaard. Epidemiological cutoff values for Flavobacterium psychrophilum MIC data generated by a standard test protocol. Journal of Fish Diseases, 39(2):143–154, February 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Soliman:2018:RDD

[SEM18] H. Soliman and M. El-Matbouli. Rapid detection and differentiation of carp oedema virus and cyprinid herpes virus-3 in koi and common carp. Journal of Fish Diseases, 41(5):761– 772, May 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Scholz:2017:FRT

[SFBW^{+}17] F. Scholz, E. Fringuelli, M. Bolton-Warberg, M. Marcos-López, S. Mitchell, P. Prodhol, D. Moffet, P. Savage, S. Murphy O’Sullivan, I. O’Connor, E. McCarthy, and H. D. Rodger. First record of Tetramicra brevifilum in lumpfish (Cyclopterus lumpus, L.). Journal of Fish Diseases, 40(6): 757–771, June 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Strike:2017:IMO

[SFF^{+}17] T. B. Strike, Y. Feltrer, E. Flach, S. K. Macgregor, and S. Guillaume. Investigation and management of an outbreak of multispecies mycobacteriosis in Australian lungfish (Neoceratodus fosteri) including the use of triple antibiotic treatment. Journal of Fish Diseases, 40(4):557–570, April 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Su:2016:DQP

[SFL^{+}16] Y-L. Su, J. Feng, Y-W. Li, J-S. Bai, and A-X. Li. Development of a quantitative PCR assay for monitoring Streptococcus agalactiae colonization and tissue tropism in experimentally infected tilapia. Journal of Fish Diseases, 39(2): 229–238, February 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sambraus:2017:WTD

[SFR^{+}17] F. Sambraus, P. G. Fjelldal, S. C. Remø, E. M. Hevrøy, T. O. Nilsen, A. Thorsen, T. J. Hansen, and R. Waagbø. Water temperature and dietary histidine affect cataract formation in Atlantic salmon (Salmo salar L.) diploid and triploid yearling smolt. Journal of Fish Diseases, 40(9):1195–1212, September 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Skilbrei:2013:IES

[SFU^{+}13] O. T. Skilbrei, B. Finstad, K. Urdal, G. Bakke, F. Kroglund, and R. Strand. Impact of early salmon louse, Lepeophtheirus salmonis, infestation and differences in survival and marine growth of sea-ranched Atlantic salmon, Salmo salar L., smolts 1997–2009. Journal of Fish Diseases, 36(3):249– 260, March 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Smail:2012:SIS

[SG12] D. A. Smail and R. Grant. The stability of infectious salmon anaemia virus infectivity at −80^{o}C in tissue homogenate and dry-stored tissue from clinically diseased Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 35(10):789–792, October 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Schmid:2016:EAT

[SGB^{+}16] T. Schmid, L. Gaede, K. Böttcher, G. Bräuer, D. Fichtner, R. Beckmann, S. Speck, F. Becker, and U. Truyen. Efficacy assessment of three inactivated koi herpes virus antigen preparations against experimental challenge virus infection in common carp. Journal of Fish Diseases, 39(8):1007–1013, August 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Scott:2015:MNG

[SGQ^{+}15] S. J. Scott, M. J. Griffin, S. Quiniou, L. Khoo, and T. K. Bollinger. Myxobolus neurophilus Guilford 1963 (Myxosporea: Myxobolidae): a common parasite infecting yellow perch Perca flavescens (Mitchell, 1814) in Saskatchewan, Canada. Journal of Fish Diseases, 38(4):355–364, April 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Singaravel:2018:NGS

[SGR^{+}18] V. Singaravel, A. Gopalakrishnan, K. Raja, M. Rajkumar, and H. W. Ferguson. Neoplasia in goldlined seabream, Rhabdosargus sarba (Forsskål, 1775). Journal of Fish Diseases, 41 (2):405–411, February 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sharifpour:2014:SCC

[SHAM14] M. F. Sharifpour, R. Hobbenaghi, P. Alyali, and A. Morvaridi. Squamous cell carcinoma in a black tetra, Gymnocorymbus ternetzi (Boulenger). Journal of Fish Diseases, 37 (11):985–987, November 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Schonherz:2012:OTR

[SHJ^{+}12] A. A. Schönherz, M. H. H. Hansen, H. B. H. Jørgensen, P. Berg, N. Lorenzen, and K. Einer-Jensen. Oral transmission as a route of infection for viral haemorrhagic septicaemia virus in rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 35(6):395–406, June 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Salazar:2016:MCS

[SHKF16] C. Salazar, D. Haussmann, G. Kausel, and J. Figueroa. Molecular cloning of Salmo salar Toll-like receptors (TLR1, TLR22, TLR5M and TLR5S) and expression analysis in SHK-1 cells during Piscirickettsia salmonis infection. Journal of Fish Diseases, 39(2):239–248, February 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Saavedra:2017:PGD

[SHO^{+}17] J. Saavedra, N. Hernandez, A. Osses, A. Castillo, A. Cancino, H. Grothusen, E. Navas, P. Henriquez, H. Bohle, F. Bustamante, P. Bustos, and M. Mancilla. Prevalence, geographic distribution and phenotypic differences of Piscirickettsia salmonis EM-90-like isolates. Journal of Fish Diseases, 40(8):1055–1063, August 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Saejung:2014:VAE

[SHS14] C. Saejung, K. Hatai, and L. Sanoamuang. The in-vitro antibacterial effects of organic salts, chemical disinfectants and antibiotics against pathogens of black disease in fairy shrimp of Thailand. Journal of Fish Diseases, 37(1):33–41, January 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stene:2016:LFP

[SHV^{+}16] A. Stene, A. Hellebø, H. Viljugrein, S. E. Solevåg, M. Devold, and V. Aspehaug. Liquid fat, a potential abiotic vector for horizontal transmission of salmonid alphavirus? Journal of Fish Diseases, 39(5):531–537, May 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Saejung:2011:COB

[SHW^{+}11] C. Saejung, K. Hatai, S. Wada, O. Kurata, and L. Sanoamuang. Clinical observations of black disease in fairy shrimps, Streptocephalus sirindhornae and Branchinella thailandensis, from Thailand and pathogen verification. Journal of Fish Diseases, 34(12):911–920, December 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Studer:2011:GSE

[SJ11] J. Studer and D. A. Janies. Global spread and evolution of viral haemorrhagic septicaemia virus. Journal of Fish Diseases, 34(10):741–747, October 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Saint-Jean:2013:AAC

[SJDC^{+}13] S. Rodríguez Saint-Jean, A. De las Heras, W. Carrillo, I. Recio, J. B. Ortiz-Delgado, M. Ramos, J. A. Gomez-Ruiz, C. Sarasquete, and S. I. Pérez-Prieto. Antiviral activity of casein and α_{s2} casein hydrolysates against the infectious haematopoietic necrosis virus, a rhabdovirus from salmonid fish. Journal of Fish Diseases, 36(5):467–481, May 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stene:2014:SIS

[SJK^{+}14] A. Stene, B. Bang Jensen, Ø. Knutsen, A. Olsen, and H. Viljugrein. Seasonal increase in sea temperature triggers pancreas disease outbreaks in Norwegian salmon farms. Journal of Fish Diseases, 37(8):739–751, August 2014. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Syakuri:2014:BGF

[SJSA^{+}14] H. Syakuri, V. Jung-Schroers, M. Adamek, G. Brogden, I. Irnazarow, and D. Steinhagen. Beta-glucan feeding differentiated the regulation of mRNA expression of claudin genes and prevented an intestinal inflammatory response post Aeromonas hydrophila intubation in common carp, Cyprinus carpio L. Journal of Fish Diseases, 37(2):149–156, February 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Schultz:2014:RCM

[SJT14] A. G. Schultz, P. L. Jones, and T. Toop. Rodlet cells in Murray cod, Maccullochella peelii peelii (Mitchell), affected with chronic ulcerative dermatopathy. Journal of Fish Diseases, 37(3):219–228, March 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Smith:2015:EIT

[SK15] P. Smith and G. Kronvall. Effect of incubation temperature and time on the precision of data generated by antibiotic disc diffusion assays. Journal of Fish Diseases, 38(7):629– 636, July 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stine:2010:MIC

[SKB10] C. B. Stine, A. S. Kane, and A. M. Baya. Mycobacteria isolated from Chesapeake Bay fish. Journal of Fish Diseases, 33(1):39–46, January 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Soto:2013:EFC

[SKGE13] E. Soto, S. Kidd, P. S. Gaunt, and R. Endris. Efficacy of florfenicol for control of mortality associated with Francisella noatunensis subsp. orientalis in Nile tilapia, Oreochromis niloticus (L.). Journal of Fish Diseases, 36(4):411–418, April 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stormoen:2013:MRP

[SKJ13] M. Stormoen, A. B. Kristoffersen, and P. A. Jansen. Mortality related to pancreas disease in Norwegian farmed salmonid fish, Salmo salar L. and Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 36(7):639–645, July 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shimahara:2016:DIR

[SKN^{+}16] Y. Shimahara, J. Kurita, T. Nishioka, I. Kiryu, K. Yuasa, T. Sakai, N. Oseko, M. Sano, and P. Dixon. Development of an improved RT-PCR for specific detection of spring viraemia of carp virus. Journal of Fish Diseases, 39(3):269– 275, March 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stejskal:2016:SLI

[SKPS16] V. Stejskal, J. Kouřil, T. Policar, and Z. Svobodová. Splenic lipidosis in intensively cultured perch, Perca fluviatilis L. Journal of Fish Diseases, 39(1):87–93, January 2016. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shoemaker:2015:LAB

[SL15] C. A. Shoemaker and B. R. LaFrentz. Lack of association between Flavobacterium columnare genomovar and virulence in hybrid tilapia Oreochromis niloticus (L.) × Oreochromis aureus (Steindachner). Journal of Fish Diseases, 38(5):491– 498, May 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Slany:2014:NCI

[Sla14] M. Slany. A new cultivation-independent tool for fast and reliable detection of Mycobacterium marinum. Journal of Fish Diseases, 37(4):363–369, April 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shoemaker:2010:PAH

[SLKE10] C. A. Shoemaker, B. R. LaFrentz, P. H. Klesius, and J. J. Evans. Protection against heterologous Streptococcus iniae isolates using a modified bacterin vaccine in Nile tilapia, Oreochromis niloticus (L.). Journal of Fish Diseases, 33(7):537– 544, July 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shoemaker:2018:INC

[SLPB18] Craig A. Shoemaker, Benjamin R. LaFrentz, Eric Peatman, and Benjamin H. Beck. Influence of native catfish mucus on Flavobacterium columnare growth and proteolytic activity. Journal of Fish Diseases, 41(9):1395–1402, September 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shahi:2014:RPK

[SM14] N. Shahi and S. K. Mallik. Recovery of Pseudomonas koreensis from eye lesions in golden mahseer, Tor putitora (Hamilton, 1822) in Uttarakhand, India. Journal of Fish Diseases, 37(5):497–500, May 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sunarto:2011:ICK

[SMC^{+}11] A. Sunarto, K. A. McColl, M. St J. Crane, T. Sumiati, A. D. Hyatt, A. C. Barnes, and P. J. Walker. Isolation and characterization of koi herpesvirus (KHV) from Indonesia: identification of a new genetic lineage. Journal of Fish Diseases, 34(2):87–101, February 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sahoo:2012:CHC

[SMD^{+}12] P. K. Sahoo, B. R. Mohanty, A. Das, K. D. Mahapatra, and J. N. Saha. Cavernous haemangioma in cultured olive barb, Puntius sarana (Hamilton). Journal of Fish Diseases, 35(9): 695–697, September 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Salati:2010:SFP

[SMF^{+}10] F. Salati, M. Meloni, A. Fenza, G. Angelucci, A. Colorni, and G. Orrù. A sensitive FRET probe assay for the selective detection of Mycobacterium marinum in fish. Journal of Fish Diseases, 33(1):47–56, January 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Straus:2012:PAE

[SMFM12] D. L. Straus, T. Meinelt, B. D. Farmer, and A. J. Mitchell. Peracetic acid is effective for controlling fungus on channel catfish eggs. Journal of Fish Diseases, 35(7):505–511, July 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Saksida:2012:PHL

[SMJ^{+}12] S. M. Saksida, G. D. Marty, S. R. M. Jones, H. A. Manchester, C. L. Diamond, J. Bidulka, and S. St-Hilaire. Parasites and hepatic lesions among pink salmon, Oncorhynchus gorbuscha (Walbaum), during early seawater residence. Journal of Fish Diseases, 35(2):137–151, February 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Slany:2014:RPM

[SMJ^{+}14] M. Slany, J. Makovcova, P. Jezek, M. Bodnarova, and I. Pavlik. Relative prevalence of Mycobacterium marinum in fish collected from aquaria and natural freshwaters in central Europe. Journal of Fish Diseases, 37(6):527–533, June 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Svoboda:2017:HTC

[SMKBP17] J. Svoboda, A. Mrugała, E. Kozubíková-Balcarová, and A. Petrusek. Hosts and transmission of the crayfish plague pathogen Aphanomyces astaci: a review. Journal of Fish Diseases, 40(1):127–140, January 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Saksida:2013:UAS

[SMM^{+}13] S. M. Saksida, D. Morrison, P. McKenzie, B. Milligan, E. Downey, B. Boyce, and A. Eaves. Use of Atlantic salmon, Salmo salar L., farm treatment data and bioassays to assess for resistance of sea lice, Lepeophtheirus salmonis, to emamectin benzoate (SLICE(R)) in British Columbia, Canada. Journal of Fish Diseases, 36(5):515–520, May 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Saksida:2010:EEB

[SMR10] S. M. Saksida, D. Morrison, and C. W. Revie. The efficacy of emamectin benzoate against infestations of sea lice, Lepeophtheirus salmonis, on farmed Atlantic salmon, Salmo salar L., in British Columbia. Journal of Fish Diseases, 33 (11):913–917, November 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Salama:2016:SET

[SMR16] N. K. G. Salama, A. G. Murray, and B. Rabe. Simulated environmental transport distances of Lepeophtheirus salmonis in Loch Linnhe, Scotland, for informing aquaculture area management structures. Journal of Fish Diseases, 39(4):419– 428, April 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sahoo:2010:CWS

[SMS^{+}10] A. K. Sahoo, C. V. Mohan, K. M. Shankar, F. Corsin, J. F. Turnbull, P. C. Thakur, N. V. Hao, K. L. Morgan, and A. P. Padiyar. Clinical white spot disease status in Penaeus monodon during the middle of the culture period — its epidemiological significance. Journal of Fish Diseases, 33(7):609–615, July 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stejskal:2018:PDI

[SMS^{+}18] V. Stejskal, J. Matousek, R. Sebesta, M. Prokesova, T. Vanina, and P. Podhorec. Prevalence of deformities in intensively reared peled Coregonus peled and comparative morphometry with pond-reared fish. Journal of Fish Diseases, 41(2): 375–381, February 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stilwell:2019:MIU

[SMS^{+}19] Justin M. Stilwell, Rita McManamon, Ginger L. Sturgeon, Melinda S. Camus, and Alvin C. Camus. Microscopic, immunohistochemical and ultrastructural characterization of spontaneous lepidosarcomas in a neon tetra, Paracheirodon innesi, and a black crappie, Pomoxis nigromaculatus. Journal of Fish Diseases, 42(7):1077–1083, July 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shetty:2014:FIC

[SMV^{+}14] M. Shetty, B. Maiti, M. N. Venugopal, I. Karunasagar, and I. Karunasagar. First isolation and characterization of Edwardsiella tarda from diseased striped catfish, Pangasianodon hypophthalmus (Sauvage). Journal of Fish Diseases, 37(3):265–271, March 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stride:2014:ETW

[SN14] M. C. Stride and B. F. Nowak. Epitheliocystis in three wild fish species in Tasmanian waters. Journal of Fish Diseases, 37(2):157–162, February 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Svendsen:2019:MIP

[SNK^{+}19] Julie Christine Svendsen, Stian Nylund, Anja B. Kristoffersen, Harald Takle, Julia Fossberg Buhaug, and Britt Bang Jensen. Monitoring infection with Piscine myocarditis virus and development of cardiomyopathy syndrome in farmed Atlantic salmon (Salmo salar L.) in Norway. Journal of Fish Diseases, 42(4):511–518, April 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sinnasamy:2016:IFP

[SNM^{+}16] S. Sinnasamy, N. Mat Noordin, T. H. MacRae, M. Ikhwanuddin bin Abdullah, P. Bossier, M. E. bin Abdul Wahid, A. Noriaki, and Y. Y. Sung. Ingestion of food pellets containing Escherichia coli overexpressing the heat-shock protein DnaK protects Penaeus vannamei (Boone) against Vibrio harveyi (Baumann) infection. Journal of Fish Diseases, 39(5):577– 584, May 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Souto:2018:BIB

[SOA^{+}18] Sandra Souto, José G. Olveira, M. Carmen Alonso, Carlos P. Dopazo, and Isabel Bandín. Betanodavirus infection in bathchallenged Solea senegalensis juveniles: a comparative analysis of RGNNV, SJNNV and reassortant strains. Journal of Fish Diseases, 41(10):1571–1578, October 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Souto:2016:RBI

[SODB16] S. Souto, J. G. Olveira, C. P. Dopazo, and I. Bandín. Reassortant betanodavirus infection in turbot (Scophthalmus maximus). Journal of Fish Diseases, 39(11):1347–1356, November 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Souto:2019:AAC

[SOGR^{+}19] Sandra Souto, José G. Olveira, Esther García-Rosado, Carlos P. Dopazo, and Isabel Bandín. Amino acid changes in the capsid protein of a reassortant betanodavirus strain: Effect on viral replication in vitro and in vivo. Journal of Fish Diseases, 42(2):221–227, February 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Strom:2018:EIY

[SON^{+}18] H. K. Strøm, M. Ohtani, B. Nowak, T. S. Boutrup, B. Jones, M. K. Raida, and A. M. Bojesen. Experimental infection by Yersinia ruckeri O1 biotype 2 induces brain lesions and neurological signs in rainbow trout (Oncorhynchus mykiss). Journal of Fish Diseases, 41(3):529–537, March 2018. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shekhar:2015:RIH

[SP15] M. S. Shekhar and A. G. Ponniah. Recent insights into host– pathogen interaction in white spot syndrome virus infected penaeid shrimp. Journal of Fish Diseases, 38(7):599–612, July 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Schulz:2019:BBC

[SPCR^{+}19] Patrycja Schulz, Joanna Pajdak-Czaus, Stanisław Robak, Jarosław Dastych, and Andrzej K. Siwicki. Bacteriophagebased cocktail modulates selected immunological parameters and post-challenge survival of rainbow trout (Oncorhynchus mykiss). Journal of Fish Diseases, 42(8):1151–1160, August 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Speare:2015:E

[Spe15] Dave Speare. Editorial. Journal of Fish Diseases, 38(1):1, January 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Speare:2019:CFB

[Spe19a] David Speare. Cleaner fish biology and aquaculture applications. Journal of Fish Diseases, 42(2):157, February 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Speare:2019:CFD

[Spe19b] David John Speare. Cleaner fish diseases. Journal of Fish Diseases, 42(2):155–156, February 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sebastiao:2017:MDF

[SPKS17] F. A. Sebastião, F. Pilarski, M. T. Kearney, and E. Soto. Molecular detection of Francisella noatunensis subsp . orientalis in cultured Nile tilapia (Oreochromis niloticus L.) in three Brazilian states. Journal of Fish Diseases, 40(11): 1731–1735, November 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sharon:2015:CIP

[SPLV^{+}15] G. Sharon, M. Pimenta-Leibowitz, M. C. L. Vilchis, N. Isakov, and D. Zilberg. Controlled infection of Poecilia reticulata Peters (guppy) with Tetrahymena by immersion and intraperitoneal injection. Journal of Fish Diseases, 38 (1):67–74, January 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sahoo:2018:FRM

[SPP^{+}18] P. K. Sahoo, S. Pattanayak, A. Paul, M. K. Sahoo, P. Rajesh Kumar, D. Panda, and B. R. Pillai. First record of Metanophrys sinensis (Protozoa: Ciliophora: Scuticociliatida) from India causing large scale mortality in a new host Macrobrachium rosenbergii larvae. Journal of Fish Diseases, 41(8):1303–1307, August 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sharma:2017:FRI

[SPS^{+}17] S. R. Krupesha Sharma, M. A. Pradeep, N. Sadu, Praveen N. Dube, and K. K. Vijayan. First report of isolation and characterization of Photobacterium damselae subsp. damselae from cage-farmed cobia (Rachycentron canadum). Journal of Fish Diseases, 40(7):953–958, July 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Suebsong:2019:SRS

[SPS^{+}19] Wasana Suebsong, Supawadee Poompuang, Prapansak Srisapoome, Skorn Koonawootrittriron, Amorn Luengnaruemitchai, Harry Johansen, and Morten Rye. Selection response for Streptococcus agalactiae resistance in Nile tilapia Oreochromis niloticus. Journal of Fish Diseases, 42 (11):1553–1562, November 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sanchez-Paz:2015:TSP

[SPTDEE^{+}15] A. Sánchez-Paz, B. Terán-Díaz, T. Enríquez-Espinoza, T. Encinas-Garcia, I. Vázquez-Sánchez, and F. Mendoza-Cano. The tidepool shrimp, Palaemon ritteri Holmes, constitutes a novel host to the white spot syndrome virus. Journal of Fish Diseases, 38(7):613–620, July 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sood:2018:CAP

[SPV^{+}18] N. Sood, P. K. Pradhan, D. K. Verma, M. K. Yadav, Ravindra, A. K. Dev, T. R. Swaminathan, and N. K. Sood. Candidatus Actinochlamydia pangasiae sp. nov. (Chlamydiales, Actinochlamydiaceae), a bacterium associated with epitheliocystis in Pangasianodon hypophthalmus. Journal of Fish Diseases, 41(2):281–290, February 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Smolowitz:2014:PTF

[SQC^{+}14] R. Smolowitz, R. A. Quinn, R. J. Cawthorn, R. L. Summerfield, and A. Y. Chistoserdov. Pathology of two forms of shell disease of the American lobster Homarus americanus Milne Edwards in Atlantic Canada. Journal of Fish Diseases, 37 (6):577–581, June 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sung:2012:EHS

[SRB12] Y. Y. Sung, R. J. Roberts, and P. Bossier. Enhancement of Hsp70 synthesis protects common carp, Cyprinus carpio L., against lethal ammonia toxicity. Journal of Fish Diseases, 35(8):563–568, August 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sudharsana:2016:MDS

[SRD^{+}16] S. Sudharsana, C. B. Rajashekar Reddy, S. Dinesh, S. Rajasekhara Reddy, A. Mohanapriya, T. Itami, and R. Sudhakaran. Molecular docking and simulation studies of 3-(1-chloropiperidin-4-yl)-6-fluoro benzisoxazole 2 against VP26 and VP28 proteins of white spot syndrome virus. Journal of Fish Diseases, 39(10):1231–1238, October 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Scholz:2018:PMV

[SRM^{+}18] F. Scholz, N. M. Ruane, T. Morrissey, M. Marcos-López, S. Mitchell, I. O’Connor, L. Mirimin, E. MacCarthy, and H. D. Rodger. Piscine myocarditis virus detected in corkwing wrasse (Symphodus melops) and ballan wrasse (Labrus bergylta). Journal of Fish Diseases, 41(1):147–152, January 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sangsuriya:2010:ICA

[SRSF10] P. Sangsuriya, J. Rojtinnakorn, S. Senapin, and T. W. Flegel. Identification and characterization of Alix/ AIP1 interacting proteins from the black tiger shrimp, Penaeus monodon. Journal of Fish Diseases, 33(7):571–581, July 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Suanyuk:2014:MPH

[SRT^{+}14] N. Suanyuk, M. Rogge, R. Thune, M. Watthanaphiromsakul, N. Champhat, and W. Wiangkum. Mortality and pathology of hybrid catfish, Clarias macrocephalus (Günther) × Clarias gariepinus (Burchell), associated with Edwardsiella ictaluri infection in southern Thailand. Journal of Fish Diseases, 37 (4):385–395, April 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shin:2017:IKY

[SS17] S. P. Shin and S. Shirakashi. Investigation of Kudoa yasunagai (Hsieh & Chen, 1984) infection kinetics in Seriola lalandi (Valenciennes, 1833) by quantitative PCR. Journal of Fish Diseases, 40(2):287–291, February 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stormoen:2013:MSL

[SSA13] M. Stormoen, E. Skjerve, and A. Aunsmo. Modelling salmon lice, Lepeophtheirus salmonis, reproduction on farmed Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 36 (1):25–33, January 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Supattapone:2017:ERA

[SSC17] S. Y. Supattapone, S. Supattapone, and R. A. Cramer. The effect of reducing agents on challenge of rainbow trout with Aeromonas salmonicida. Journal of Fish Diseases, 40(3): 437–441, March 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shen:2017:IIP

[SSF^{+}17] G. M. Shen, C. Y. Shi, C. Fan, D. Jia, S. Q. Wang, G. S. Xie, G. Y. Li, Z. L. Mo, and J. Huang. Isolation, identification and pathogenicity of Vibrio harveyi, the causal agent of skin ulcer disease in juvenile hybrid groupers Epinephelus fuscoguttatus × Epinephelus lanceolatus. Journal of Fish Diseases, 40(10):1351–1362, October 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sivalingam:2019:ZFD

[SSG^{+}19] Nathiga Nambi Kalaiselvi Sivalingam, Abdul Majeed Seepoo, Taju Gani, Sivakumar Selvam, and Sahul Hameed Azeez Sait. Zebrafish fin-derived fibroblast cell line: a model for in vitro wound healing. Journal of Fish Diseases, 42(4):573– 584, April 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Saleh:2011:ACG

[SSHEM11] M. Saleh, H. Soliman, O. Haenen, and M. El-Matbouli. Antibody-coated gold nanoparticles immunoassay for direct detection of Aeromonas salmonicida in fish tissues. Journal of Fish Diseases, 34(11):845–852, November 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shin:2019:QIE

[SSJ^{+}19] Sang Phil Shin, Han Chang Sohn, Chang Nam Jin, Bong Jo Kang, and Jehee Lee. Quantitative investigation of Enteromyxum leei (Myxozoa: Myxosporea) infection and relative condition factor in cultured olive flounder Paralichthys olivaceus (Temminck and Schlegel). Journal of Fish Diseases, 42(2):159–165, February 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Spagnoli:2017:CNP

[SSK17] S. Spagnoli, J. Sanders, and M. L. Kent. The common neural parasite Pseudoloma neurophilia causes altered shoaling behaviour in adult laboratory zebrafish (Danio rerio) and its implications for neurobehavioural research. Journal of Fish Diseases, 40(3):443–446, March 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Susdorf:2018:IBC

[SSL18] R. Susdorf, N. K. G. Salama, and D. Lusseau. Influence of body condition on the population dynamics of Atlantic salmon with consideration of the potential impact of sea lice. Journal of Fish Diseases, 41(6):941–951, June 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Subramaniam:2014:DMC

[SSO^{+}14] K. Subramaniam, M. Shariff, A. R. Omar, M. Hair-Bejo, and B. L. Ong. Detection and molecular characterization of infectious spleen and kidney necrosis virus from major ornamental fish breeding states in Peninsular Malaysia. Journal of Fish Diseases, 37(7):609–618, July 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Singh:2018:EMN

[SSSP18] Mukta Singh, Ratan Kumar Saha, Himadri Saha, and Janmejay Parhi. Effect of miconazole nitrate on immunological response and its preventive efficacy in Labeo rohita fingerlings against oomycetes Saprolegnia parasitica. Journal of Fish Diseases, 41(10):1539–1548, October 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sarathi:2010:EBE

[SSV^{+}10] M. Sarathi, M. C. Simon, C. Venkatesan, J. Thomas, M. Ravi, N. Madan, S. Thiyagarajan, and A. S. Sahul Hameed. Efficacy of bacterially expressed dsRNA specific to different structural genes of white spot syndrome virus (WSSV) in protection of shrimp from WSSV infection. Journal of Fish Diseases, 33(7):603–607, July 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Santhoshkumar:2017:BCT

[SSV^{+}17] S. Santhoshkumar, S. Sivakumar, S. Vimal, S. Abdul Majeed, G. Taju, P. Haribabu, A. Uma, and A. S. Sahul Hameed. Biochemical changes and tissue distribution of Enterocytozoon hepatopenaei (EHP) in naturally and experimentally EHP-infected whiteleg shrimp, Litopenaeus vannamei (Boone, 1931), in India. Journal of Fish Diseases, 40 (4):529–539, April 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Songe:2014:VPI

[STES14] M. M. Songe, E. Thoen, Ø. Evensen, and I. Skaar. In vitro passages impact on virulence of Saprolegnia parasitica to Atlantic salmon, Salmo salar L. parr. Journal of Fish Diseases, 37(9):825–834, September 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sandell:2015:IRS

[STF^{+}15] T. A. Sandell, D. J. Teel, J. Fisher, B. Beckman, and K. C. Jacobson. Infections by Renibacterium salmoninarum and Nanophyetus salmincola Chapin are associated with reduced growth of juvenile Chinook salmon, Oncorhynchus tshawytscha (Walbaum), in the Northeast Pacific Ocean. Journal of Fish Diseases, 38(4):365–378, April 2015. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Senapin:2010:IYH

[STG^{+}10] S. Senapin, Y. Thaowbut, W. Gangnonngiw, N. Chuchird, S. Sriurairatana, and T. W. Flegel. Impact of yellow head virus outbreaks in the whiteleg shrimp, Penaeus vannamei (Boone), in Thailand. Journal of Fish Diseases, 33(5):421– 430, May 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sepulveda:2017:PRE

[STIG17] F. A. Sepúlveda, J. F. Torres, C. D. Infante, and M. T. González. Potential role of ectoparasites (Zeuxapta seriolae and Caligus lalandei) in the transmission of pathogenic bacteria in yellowtail kingfish Seriola lalandi, inferred from cultivable microbiota and molecular analyses. Journal of Fish Diseases, 40(7):979–985, July 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sahoo:2015:HFI

[STS^{+}15] A. K. Sahoo, P. C. Thakur, K. M. Shankar, C. V. Mohan, S. R. Krupesha Sharma, and F. Corsin. Histopathological findings on innate responses of white spot disease positive Penaeus monodon (Fabricius) under semi-intensive culture. Journal of Fish Diseases, 38(1):91–95, January 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Semple:2017:DUA

[SVL^{+}17] S. L. Semple, N. T. K. Vo, A. R. Li, P. H. Pham, N. C. Bols, and B. Dixon. Development and use of an Arctic charr cell line to study antiviral responses at extremely low temperatures. Journal of Fish Diseases, 40(10):1423–1439, October 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sivakumar:2018:NSW

[SVM^{+}18] S. Sivakumar, S. Vimal, S. Abdul Majeed, S. Santhosh Kumar, G. Taju, N. Madan, T. Rajkumar, S. Thamizhvanan, K. V. Shamsudheen, V. Scaria, S. Sivasubbu, and A. S. Sahul Hameed. A new strain of white spot syndrome virus affecting Litopenaeus vannamei in Indian shrimp farms. Journal of Fish Diseases, 41(7):1129–1146, July 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Soto:2017:ITD

[SVV^{+}17] M. Guevara Soto, B. Vidondo, L. Vaughan, J-F. Rubin, H. Segner, S. Samartin, and H. Schmidt-Posthaus. Investigations into the temporal development of epitheliocystis infections in brown trout: a histological study. Journal of Fish Diseases, 40(6):811–819, June 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stene:2014:TDP

[SVY^{+}14] A. Stene, H. Viljugrein, H. Yndestad, S. Tavornpanich, and E. Skjerve. Transmission dynamics of pancreas disease (PD) in a Norwegian fjord: aspects of water transport, contact networks and infection pressure among salmon farms. Journal of Fish Diseases, 37(2):123–134, February 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sundell:2011:EBF

[SW11] K. Sundell and T. Wiklund. Effect of biofilm formation on antimicrobial tolerance of Flavobacterium psychrophilum. Journal of Fish Diseases, 34(5):373–383, May 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shin:2014:DRI

[SWD^{+}14] G. W. Shin, S. L. White, H. U. Dahms, H. D. Jeong, and J. H. Kim. Disease resistance and immune-relevant gene expression in golden mandarin fish, Siniperca scherzeri Steindachner, infected with infectious spleen and kidney necrosis virus-like agent. Journal of Fish Diseases, 37(12):1041–1054, December 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Soares:2019:FIF

[SWE^{+}19] Sílvia M. C. Soares, Amanda Walker, Svenja A. Elwenn, Sion Bayliss, Alison Garden, Hannah E. B. Stagg, and Eann S. Munro. First isolation of Flavobacterium psychrophilum associated with reports of moribund wild European eel (Anguilla anguilla) in Scotland. Journal of Fish Diseases, 42 (11):1509–1521, November 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Smith:2014:BSP

[SWH^{+}14] A. L. Smith, M. M. A. Whitten, L. Hirschle, E. C. Pope, E. C. Wootton, C. L. Vogan, and A. F. Rowley. Bacterial septicaemia in prerecruit edible crabs, Cancer pagurus L. Journal of Fish Diseases, 37(8):729–737, August 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Schwenteit:2015:CAS

[SWM^{+}15] J. M. Schwenteit, B. Weber, D. L. Milton, U. T. Bornscheuer, and B. K. Gudmundsdottir. Construction of Aeromonas salmonicida subsp. achromogenes AsaP1-toxoid strains and study of their ability to induce immunity in Arctic char, Salvelinus alpinus L. Journal of Fish Diseases, 38(10):891– 900, October 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Songe:2016:TCG

[SWS^{+}16] M. M. Songe, A. Willems, M. N. Sarowar, K. Rajan, Ø. Evensen, K. Drynan, I. Skaar, and P. van West. A thicker chorion gives ova of Atlantic salmon (Salmo salar L.) the upper hand against Saprolegnia infections. Journal of Fish Diseases, 39(7):879–888, July 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Songe:2016:SDI

[SWWN^{+}16] M. M. Songe, A. Willems, J. Wiik-Nielsen, E. Thoen, Ø. Evensen, P. van West, and I. Skaar. Saprolegnia diclina IIIA and S. parasitica employ different infection strategies when colonizing eggs of Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 39(3):343–352, March 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sanders:2015:ZDR

[SZM^{+}15] J. L. Sanders, Y. Zhou, H. M. Moulton, Z. X. Moulton, R. McLeod, J. P. Dubey, L. M. Weiss, and M. L. Kent. The zebrafish, Danio rerio, as a model for Toxoplasma gondii: an initial description of infection in fish. Journal of Fish Diseases, 38(7):675–679, July 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Siti-Zahrah:2014:DCC

[SZZSFN^{+}14] A. Siti-Zahrah, M. Zamri-Saad, M. Firdaus-Nawi, M. K. Hazreen-Nita, and M. Nur-Nazifah. Detection of channel catfish virus in cage-cultured Pangasius hypophthalmus (Sauvage, 1878) in Malaysia. Journal of Fish Diseases, 37 (11):981–983, November 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ture:2015:CPF

[TAC15] M. Ture, I. Altinok, and E. Capkin. Comparison of pulsedfield gel electrophoresis and enterobacterial repetitive intergenic consensus PCR and biochemical tests to characterize Lactococcus garvieae. Journal of Fish Diseases, 38(1):37–47, January 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tinsley:2011:NNM

[TALA11] J. W. Tinsley, D. A. Austin, A. R. Lyndon, and B. Austin. Novel non-motile phenotypes of Yersinia ruckeri suggest expansion of the current clonal complex theory. Journal of Fish Diseases, 34(4):311–317, April 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ternengo:2010:IIS

[TAQ^{+}10] S. Ternengo, S. Agostini, Y. Quilichini, L. Euzet, and B. Marchand. Intensive infestations of Sciaenocotyle pancerii (Monogenea, Microcotylidae) on Argyrosomus regius (Asso) under fish-farming conditions. Journal of Fish Diseases, 33 (1):89–92, January 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tengs:2012:SPM

[TB12] T. Tengs and I. Böckerman. A strain of piscine myocarditis virus infecting Atlantic argentine, Argentina silus (Ascanius). Journal of Fish Diseases, 35(7):545–547, July 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Taylor:2010:AGD

[TCC10a] R. S. Taylor, P. B. Crosbie, and M. T. Cook. Amoebic gill disease resistance is not related to the systemic antibody response of Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 33(1):1–14, January 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Taylor:2010:E

[TCC10b] R. S. Taylor, P. B. Crosbie, and M. T. Cook. Erratum. Journal of Fish Diseases, 33(7):623, July 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic). See [RMSH09].

Torres-Corral:2019:HTI

[TCFáS19] Yolanda Torres-Corral, Clara Fernández-álvarez, and Ysabel Santos. High-throughput identification and quantification of Vagococcus salmoninarum by SYBR Green I-based real-time PCR combined with melting curve analysis. Journal of Fish Diseases, 42(10):1359–1368, October 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tighe:2019:GDP

[TCM^{+}19] Andrew J. Tighe, Jens Carlsson, Teresa Morrissey, Fiona Swords, and Neil M. Ruane. Genetic diversity of piscine myocarditis virus in Atlantic salmon Salmo salar L. in Ireland. Journal of Fish Diseases, 42(8):1161–1168, August 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Torres-Corral:2019:ITV

[TCS19] Yolanda Torres-Corral and Ysabel Santos. Identification and typing of Vagococcus salmoninarum using genomic and proteomic techniques. Journal of Fish Diseases, 42(4):597–612, April 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tobback:2010:VMV

[TDH^{+}10] E. Tobback, A. Decostere, K. Hermans, W. Van den Broeck, F. Haesebrouck, and K. Chiers. In vitro markers for virulence in Yersinia ruckeri. Journal of Fish Diseases, 33(3):197– 209, March 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Taylor:2010:KHD

[TDJ^{+}10] N. G. H. Taylor, P. F. Dixon, K. R. Jeffery, E. J. Peeler, K. L. Denham, and K. Way. Koi herpesvirus: distribution and prospects for control in England and Wales. Journal of Fish Diseases, 33(3):221–230, March 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tuan:2016:DUK

[TDLT^{+}16] V. V. Tuan, J. J. Dantas-Lima, K. V. Thuong, W. Li, K. Grauwet, P. Bossier, and H. J. Nauwynck. Differences in uptake and killing of pathogenic and non-pathogenic bacteria by haemocyte subpopulations of penaeid shrimp, Litopenaeus vannamei, (Boone). Journal of Fish Diseases, 39(2): 163–174, February 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Toffan:2019:LCL

[TDS^{+}19] Anna Toffan, Maria De Salvador, Felix Scholz, Tobia Pretto, Alessandra Buratin, Hamish D. Rodger, Marica Toson, Argelia Cuenca, and Niccolò Vendramin. Lumpfish (Cyclopterus lumpus, Linnaeus) is susceptible to viral nervous necrosis: Result of an experimental infection with different genotypes of betanodavirus. Journal of Fish Diseases, 42 (12):1667–1676, December 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tuan:2016:KAI

[TDT^{+}16] V. V. Tuan, G. M. A. De Gryse, K. V. Thuong, P. Bossier, and H. J. Nauwynck. Kinetic analysis of internalization of white spot syndrome virus by haemocyte subpopulations of penaeid shrimp, Litopenaeus vannamei (Boone), and the outcome for virus and cell. Journal of Fish Diseases, 39 (12):1477–1493, December 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Thoen:2011:PSS

[TES11] E. Thoen, Ø. Evensen, and I. Skaar. Pathogenicity of Saprolegnia spp. to Atlantic salmon, Salmo salar L., eggs. Journal of Fish Diseases, 34(8):601–608, August 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Thoen:2016:FIS

[TES16] E. Thoen, Ø. Evensen, and I. Skaar. Factors influencing Saprolegnia spp. spore numbers in Norwegian salmon hatcheries. Journal of Fish Diseases, 39(6):657–665, June 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tedesco:2019:VAC

[TFG19] Perla Tedesco, Maria Letizia Fioravanti, and Roberta Galuppi. In vitro activity of chemicals and commercial products against Saprolegnia parasitica and Saprolegnia delica strains. Journal of Fish Diseases, 42(2):237–248, February 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Trujillo-Gonzalez:2015:HAH

[TGJCH15] A. Trujillo-González, L. K. Johnson, C. C. Constantinoiu, and K. S. Hutson. Histopathology associated with haptor attachment of the ectoparasitic monogenean Neobenedenia sp. (Capsalidae) to barramundi, Lates calcarifer (Bloch). Journal of Fish Diseases, 38(12):1063–1067, December 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tsai:2012:DMP

[THW^{+}12] M-A. Tsai, P-Y. Ho, P-C. Wang, Y-J. E., L-L. Liaw, and S-C. Chen. Development of a multiplex polymerase chain reaction to detect five common Gram-negative bacteria of aquatic animals. Journal of Fish Diseases, 35(7):489–495, July 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tanaka:2014:FCI

[TIK^{+}14a] N. Tanaka, T. Izawa, M. Kuwamura, N. Higashiguchi, C. Kezuka, O. Kurata, S. Wada, and J. Yamate. The first case of infectious spleen and kidney necrosis virus (ISKNV) infection in aquarium-maintained mandarin fish, Siniperca chuatsi (Basilewsky), in Japan. Journal of Fish Diseases, 37 (4):401–405, April 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tanaka:2014:PHC

[TIK^{+}14b] N. Tanaka, T. Izawa, M. Kuwamura, N. Higashiguchi, C. Kezuka, and J. Yamate. Phaeochromocytoma and hepatocellular carcinoma with nuclear glycogenation of the hepatocytes in a predatory carp, Chanodichthys erythropterus (Basilewsky). Journal of Fish Diseases, 37(4):411–414, April 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Torrissen:2013:SLI

[TJA^{+}13] O. Torrissen, S. Jones, F. Asche, A. Guttormsen, O. T. Skilbrei, F. Nilsen, T. E. Horsberg, and D. Jackson. Salmon lice — impact on wild salmonids and salmon aquaculture. Journal of Fish Diseases, 36(3):171–194, March 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Taksdal:2015:MWL

[TJB^{+}15] T. Taksdal, B. Bang Jensen, I. Böckerman, M. F. McLoughlin, M. J. Hjortaas, A. Ramstad, and H. Sindre. Mortality and weight loss of Atlantic salmon, Salmo salar L., experimentally infected with salmonid alphavirus subtype 2 and subtype 3 isolates from Norway. Journal of Fish Diseases, 38(12):1047–1061, December 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tidbury:2018:EFN

[TJR^{+}18] Hannah J. Tidbury, Claire L. Joiner, Georgina S. E. Rimmer, Holly V. Potter, and Nick G. H. Taylor. The effectiveness of fishery net dips: Advice for the improvement of biosecurity measures. Journal of Fish Diseases, 41(11):1625–1630, November 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Troszok:2018:AIC

[TKS^{+}18] Agnieszka Troszok, Ludmiła Kolek, Joanna Szczygieł, Joanna Wawrzeczko, Ewa Borzym, Michał Reichert, Teresa Kamińska, Tomasz Ostrowski, Patrycja Jurecka, Mikołaj Adamek, Krzysztof Rakus, and Ilgiz Irnazarow. Acyclovir inhibits Cyprinid herpesvirus 3 multiplication in vitro. Journal of Fish Diseases, 41(11):1709–1718, November 2018. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Torrecillas:2012:EMS

[TMC^{+}12] S. Torrecillas, A. Makol, M. J. Caballero, D. Montero, A. K. S. Dhanasiri, J. Sweetman, and M. Izquierdo. Effects on mortality and stress response in European sea bass, Dicentrarchus labrax (L.), fed mannan oligosaccharides (MOS) after Vibrio anguillarum exposure. Journal of Fish Diseases, 35(8):591–602, August 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Taju:2015:IRW

[TMM^{+}15] G. Taju, N. Madan, S. Abdul Majeed, T. Raj Kumar, S. Thamizhvanan, S. K. Otta, and A. S. Sahul Hameed. Immune responses of whiteleg shrimp, Litopenaeus vannamei (Boone, 1931), to bacterially expressed dsRNA specific to VP28 gene of white spot syndrome virus. Journal of Fish Diseases, 38(5):451–465, May 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tresise:2014:PLN

[TMWV14] M. M. Tresise, M. L. L. Mokae, G. M. Wagenaar, and J. C. Van Dyk. A proposed liver needle core biopsy technique for the sharptooth catfish Clarias gariepinus (Burchell) for use in fish health research. Journal of Fish Diseases, 37(10): 931–934, October 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Touraki:2012:TVE

[TNK12] M. Touraki, I. Niopas, and V. Karagiannis. Treatment of vibriosis in European sea bass larvae, Dicentrarchus labrax L., with oxolinic acid administered by bath or through medicated nauplii of Artemia franciscana (Kellogg): efficacy and residual kinetics. Journal of Fish Diseases, 35(7):513–522, July 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Takeuchi:2016:GVK

[TOK^{+}16] F. Takeuchi, Y. Ogasawara, K. Kato, T. Sekizuka, T. Nozaki, Y. Sugita-Konishi, T. Ohnishi, and M. Kuroda. Genetic variants of Kudoa septempunctata (Myxozoa: Multivalvulida), a flounder parasite causing foodborne disease. Journal of Fish Diseases, 39(6):667–672, June 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Thuong:2019:MHC

[TPD^{+}19] Nguyen Phuc Thuong, Edina Prondvai, Barbara De Kegel, Tania De Wolf, Paul Eckhard Witten, and Dominique Adriaens. Morphological and histological characterization of an ectopically mineralized structure in a gilthead sea bream Sparus aurata with opercular deformation. Journal of Fish Diseases, 42(9):1259–1270, September 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tian:2014:CDS

[TQW^{+}14] J. Y. Tian, Z. T. Qi, N. Wu, M. X. Chang, and P. Nie. Complementary DNA sequences of the constant regions of Tcell antigen receptors α, β and γ in mandarin fish, Siniperca chuatsi Basilewsky, and their transcriptional changes after stimulation with Flavobacterium columnare. Journal of Fish Diseases, 37(2):89–101, February 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Treasurer:2012:DNE

[Tre12] J. W. Treasurer. Diseases of north European wrasse (Labridae) and possible interactions with cohabited farmed salmon, Salmo salar L. Journal of Fish Diseases, 35(8):555–562, August 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Thammatorn:2019:MRT

[TRS19] Worrayanee Thammatorn, Pattarasuda Rawiwan, and Win Surachetpong. Minimal risk of tilapia lake virus transmission via frozen tilapia fillets. Journal of Fish Diseases, 42(1):3–9, January 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Thamizhvanan:2019:MIC

[TSK^{+}19] S. Thamizhvanan, S. Sivakumar, S. Santhosh Kumar, D. Vinoth Kumar, S. Suryakodi, K. Balaji, T. Rajkumar, S. Vimal, S. Abdul Majeed, G. Taju, and A. S. Sahul Hameed. Multiple infections caused by white spot syndrome virus and Enterocytozoon hepatopenaei in pond-reared Penaeus vannamei in India and multiplex PCR for their simultaneous detection. Journal of Fish Diseases, 42(3):447– 454, March 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tattiyapong:2018:DVR

[TSS18] P. Tattiyapong, K. Sirikanchana, and W. Surachetpong. Development and validation of a reverse transcription quantitative polymerase chain reaction for tilapia lake virus detection in clinical samples and experimentally challenged fish. Journal of Fish Diseases, 41(2):255–261, February 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Thuong:2016:EAC

[TTL^{+}16] K. Van Thuong, V. Van Tuan, W. Li, P. Sorgeloos, P. Bossier, and H. Nauwynck. Effects of acute change in salinity and moulting on the infection of white leg shrimp (Penaeus vannamei) with white spot syndrome virus upon immersion challenge. Journal of Fish Diseases, 39(12):1403– 1412, December 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Thakur:2019:IAD

[TVK^{+}19] Krishna K. Thakur, Raphaël Vanderstichel, Karia Kaukinen, Omid Nekouei, Emilie Laurin, and Kristina M. Miller. Infectious agent detections in archived Sockeye salmon (Oncorhynchus nerka) samples from British Columbia, Canada (1985–94). Journal of Fish Diseases, 42(4):533–547, April 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Truter:2016:ESG

[TvWO^{+}16] J. C. Truter, J. H. van Wyk, P. J. Oberholster, A-M. Botha, and W. J. Luus-Powell. The expression of selected genes linked to metabolic homeostasis in obese pansteatitissuffering Mozambique tilapia, Oreochromis mossambicus (Peters). Journal of Fish Diseases, 39(1):69–85, January 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Taylor:2010:RLF

[TWJP10] N. G. H. Taylor, K. Way, K. R. Jeffery, and E. J. Peeler. The role of live fish movements in spreading koi herpesvirus throughout England and Wales. Journal of Fish Diseases, 33(12):1005–1007, December 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Taksdal:2012:QRS

[TWNB^{+}12] T. Taksdal, J. Wiik-Nielsen, S. Birkeland, P. Dalgaard, and T. Mørkøre. Quality of raw and smoked fillets from clinically healthy Atlantic salmon, Salmo salar L., following an outbreak of pancreas disease (PD). Journal of Fish Diseases, 35(12):897–906, December 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tsai:2015:GCS

[TWYC15] M-A. Tsai, P-C. Wang, T. Yoshida, and S-C. Chen. Genetic characteristics of Streptococcus dysgalactiae isolated from cage cultured cobia, Rachycentron canadum (L.). Journal of Fish Diseases, 38(12):1037–1046, December 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ulvestad:2018:SEL

[UKS^{+}18] J. S. Ulvestad, J. Kumari, T. Seternes, H. Chi, and R. A. Dalmo. Studies on the effects of LPS, β-glucan and metabolic inhibitors on the respiratory burst and gene expression in Atlantic salmon macrophages. Journal of Fish Diseases, 41 (7):1117–1127, July 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ugelvik:2017:ASI

[UMMS17] M. S. Ugelvik, T. Mo, A. Mennerat, and A. Skorping. Atlantic salmon infected with salmon lice are more susceptible to new lice infections. Journal of Fish Diseases, 40(3):311– 317, March 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ullal:2010:AAA

[UN10] A. J. Ullal and E. J. Noga. Antiparasitic activity of the antimicrobial peptide Hbβ P-1, a member of the β-haemoglobin peptide family. Journal of Fish Diseases, 33(8):657–664, August 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ugelvik:2017:PFD

[USM17] M. S. Ugelvik, A. Skorping, and A. Mennerat. Parasite fecundity decreases with increasing parasite load in the salmon louse Lepeophtheirus salmonis infecting Atlantic salmon Salmo salar. Journal of Fish Diseases, 40(5):671– 678, May 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Vo:2015:DWC

[VBL^{+}15] N. T. K. Vo, A. W. Bender, L. E. J. Lee, J. S. Lumsden, N. Lorenzen, B. Dixon, and N. C. Bols. Development of a walleye cell line and use to study the effects of temperature on infection by viral haemorrhagic septicaemia virus group IVb. Journal of Fish Diseases, 38(2):121–136, February 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Vo:2016:DVH

[VBL^{+}16] N. T. K. Vo, A. W. Bender, J. S. Lumsden, B. Dixon, and N. C. Bols. Differential viral haemorrhagic septicaemia virus genotype IVb infection in fin fibroblast and epithelial cell lines from walleye, Sander vitreus (Mitchill), at cold temperatures. Journal of Fish Diseases, 39(2):175–188, February 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

vanBeurden:2016:DVR

[vBVLR^{+}16] S. J. van Beurden, M. A. Voorbergen-Laarman, I. Roozenburg, J. van Tellingen, O. L. M. Haenen, and M. Y. Engelsma. Development and validation of a real-time PCR assay for the detection of anguillid herpesvirus 1. Journal of Fish Diseases, 39(1):95–104, January 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Vincent:2017:PAF

[VC17] A. T. Vincent and S. J. Charette. Phylogenetic analysis of the fish pathogen Aeromonas salmonicida underlines the dichotomy between European and Canadian strains for the salmonicida subspecies. Journal of Fish Diseases, 40(9): 1241–1247, September 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Vigil:2016:TEH

[VCLW16] E. M. Vigil, K. R. Christianson, J. M. Lepak, and P. J. Williams. Temperature effects on hatching and viability of Juvenile Gill Lice, Salmincola californiensis. Journal of Fish Diseases, 39(7):899–905, July 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

VanDerMarel:2010:CSM

[VCN^{+}10] M. Van Der Marel, N. Caspari, H. Neuhaus, W. Meyer, M-L. Enss, and D. Steinhagen. Changes in skin mucus of common carp, Cyprinus carpio L., after exposure to water with a high bacterial load. Journal of Fish Diseases, 33(5):431–439, May 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Vazquez:2017:DVR

[VCOD17] D. Vázquez, J. M. Cutrín, J. G. Olveira, and C. P. Dopazo. Design and validation of a RT-qPCR procedure for diagnosis and quantification of most types of infectious pancreatic necrosis virus using a single pair of degenerated primers. Journal of Fish Diseases, 40(9):1155–1167, September 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Vesely:2011:IOF

[VCR^{+}11] T. Vesely, K. Cinkova, S. Reschova, F. Gobbo, E. Ariel, M. Vicenova, D. Pokorova, P. Kulich, and G. Bovo. Investigation of ornamental fish entering the EU for the presence of ranaviruses. Journal of Fish Diseases, 34(2):159–166, February 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Vendramin:2019:PGV

[VCS^{+}19] Niccoló Vendramin, Argelia Cuenca, Juliane Sørensen, Anna L. F. Alencar, Debes H. Christiansen, Jan A. Jacobsen, Charlotte Axen, François Lieffrig, Neil M. Ruane, Patrick Martin, Timothy Sheehan, Tine M. Iburg, Espen Rimstad, and Niels J. Olesen. Presence and genetic variability of Piscine orthoreovirus genotype 1 (PRV-1) in wild salmonids in Northern Europe and North Atlantic Ocean. Journal of Fish Diseases, 42(8):1107–1118, August 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

vanDyk:2014:CMT

[vD14] J. C. van Dyk. Cholangioma in the Mozambique tilapia, Oreochromis mossambicus (Peters). Journal of Fish Diseases, 37 (9):847–851, September 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Vercauteren:2018:FIV

[VDD^{+}18] M. Vercauteren, E. De Swaef, A. Declercq, L. Bosseler, S. Gulla, S. Balboa, J. L. Romalde, L. Devriese, H. Polet, F. Boyen, K. Chiers, and A. Decostere. First isolation of Vibrio tapetis and an atypical strain of Aeromonas salmonicida from skin ulcerations in common dab (Limanda limanda) in the North Sea. Journal of Fish Diseases, 41(2):329–335, February 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Viljamaa-Dirks:2019:TNS

[VDH19] Satu Viljamaa-Dirks and Sirpa Heinikainen. A tentative new species Aphanomyces fennicus sp. nov. interferes with molecular diagnostic methods for crayfish plague. Journal of Fish Diseases, 42(3):413–422, March 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Viljamaa-Dirks:2016:VHG

[VDHV^{+}16] S. Viljamaa-Dirks, S. Heinikainen, A-M. K. Virtala, H. Torssonen, and S. Pelkonen. Variation in the hyphal growth rate and the virulence of two genotypes of the crayfish plague organism Aphanomyces astaci. Journal of Fish Diseases, 39(6):753–764, June 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Valdes-Donoso:2013:CIP

[VDMJ^{+}13] P. Valdes-Donoso, F. O. Mardones, M. Jarpa, M. Ulloa, T. E. Carpenter, and A. M. Perez. Co-infection patterns of infectious salmon anaemia and sea lice in farmed Atlantic salmon, Salmo salar L., in southern Chile (2007–2009). Journal of Fish Diseases, 36(3):353–360, March 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Valentim:2018:NEO

[VDO^{+}18a] D. S. S. Valentim, J. L. Duarte, A. E. M. F. M. Oliveira, R. A. S. Cruz, J. C. T. Carvalho, E. C. Conceição, C. P. Fernandes, and M. Tavares-Dias. Nanoemulsion from essential oil of Pterodon emarginatus (Fabaceae) shows in vitro efficacy against monogeneans of Colossoma macropomum (Pisces: Serrasalmidae). Journal of Fish Diseases, 41(3): 443–449, March 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Valentim:2018:ENC

[VDO^{+}18b] D. S. S. Valentim, J. L. Duarte, A. E. M. F. M. Oliveira, R. A. S. Cruz, J. C. T. Carvalho, C. Solans, C. P. Fernandes, and M. Tavares-Dias. Effects of a nanoemulsion with Copaifera officinalis oleoresin against monogenean parasites of Colossoma macropomum: a neotropical Serrasalmidae. Journal of Fish Diseases, 41(7):1041–1048, July 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Vilar:2012:MFS

[VFB^{+}12] P. Vilar, L. D. Faílde, R. Bermúdez, F. Vigliano, A. Riaza, R. Silva, Y. Santos, and M. I. Quiroga. Morphopathological features of a severe ulcerative disease outbreak associated with Tenacibaculum maritimum in cultivated sole, Solea senegalensis (L.). Journal of Fish Diseases, 35(6):437– 445, June 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Valladao:2016:EOC

[VGI^{+}16] G. M. R. Valladão, S. U. Gallani, C. V. Ikefuti, C. da Cruz, N. Levy-Pereira, M. V. N. Rodrigues, and F. Pilarski. Essential oils to control ichthyophthiriasis in pacu, Piaractus mesopotamicus (Holmberg): special emphasis on treatment with Melaleuca alternifolia. Journal of Fish Diseases, 39(10): 1143–1152, October 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jorgensen:2011:EED

[vGJHS^{+}11] L. von Gersdorff Jørgensen, R. D. Heinecke, K. Skjødt, K. J. Rasmussen, and K. Buchmann. Experimental evidence for direct in situ binding of IgM and IgT to early trophonts of Ichthyophthirius multifiliis (Fouquet) in the gills of rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 34(10):749–755, October 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jorgensen:2017:RTO

[vGJKR^{+}17] L. von Gersdorff Jørgensen, P. W. Kania, K. J. Rasmussen, A. H. Mattsson, J. Schmidt, A. Al-Jubury, A. Sander, A. Salanti, and K. Buchmann. Rainbow trout (Oncorhynchus mykiss) immune response towards a recombinant vaccine targeting the parasitic ciliate Ichthyophthirius multifiliis. Journal of Fish Diseases, 40(12):1815–1821, December 2017. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Volpe:2018:DMC

[VGP^{+}18] E. Volpe, M. Grodzki, V. Panzarin, A. Guercio, G. Purpari, P. Serratore, and S. Ciulli. Detection and molecular characterization of betanodaviruses retrieved from bivalve molluscs. Journal of Fish Diseases, 41(4):603–611, April 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Verner-Jeffreys:2018:SAM

[VJWC^{+}18] D. W. Verner-Jeffreys, T. J. Wallis, I. Cano Cejas, D. Ryder, D. J. Haydon, J. F. Domazoro, J. Dontwi, T. R. Field, D. Adjei-Boteng, G. Wood, T. Bean, and S. W. Feist. Streptococcus agalactiae Multilocus sequence type 261 is associated with mortalities in the emerging Ghanaian tilapia industry. Journal of Fish Diseases, 41(1):175–179, January 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Vo:2019:CCS

[VKK^{+}19] Nguyen T. K. Vo, Barbara A. Katzenback, Calvin Kellendonk, Tu Duong, Theresa M. Curtis, Brian Dixon, and Niels C. Bols. Characterization of the continuous skin fibroblastoid cell line, WE-skin11f, from walleye (Sander vitreus). Journal of Fish Diseases, 42(11):1587–1599, November 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Vazquez:2016:NMR

[VLVS^{+}16] D. Vázquez, C. López-Vázquez, H. F. Skall, S. S. Mikkelsen, N. J. Olesen, and C. P. Dopazo. A novel multiplex RTqPCR method based on dual-labelled probes suitable for typing all known genotypes of viral haemorrhagic septicaemia virus. Journal of Fish Diseases, 39(4):467–482, April 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Valenzuela-Munoz:2017:UIR

[VMBGE17] V. Valenzuela-Muñoz, S. Boltaña, and C. Gallardo-Escárate. Uncovering iron regulation with species-specific transcriptome patterns in Atlantic and Coho salmon during a Caligus rogercresseyi infestation. Journal of Fish Diseases, 40(9):

1169–1184, September 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Vasquez-Machado:2019:MCA

[VMBGIC19] Gersson Vásquez-Machado, Paola Barato-Gómez, and Carlos Iregui-Castro. Morphological characterization of the adherence and invasion of Streptococcus agalactiae to the intestinal mucosa of tilapia Oreochromis sp.: an in vitro model. Journal of Fish Diseases, 42(9):1223–1231, September 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Vergilio:2014:EVE

[VMCM14] C. S. Vergílio, R. V. Moreira, C. E. V. Carvalho, and E. J. T. Melo. Effects of in vitro exposure to mercury on male gonads and sperm structure of the tropical fish tuvira Gymnotus carapo (L.). Journal of Fish Diseases, 37(6):543–551, June 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Volpe:2019:EFH

[VME^{+}19] Enrico Volpe, Luciana Mandrioli, Francesca Errani, Patrizia Serratore, Emanuele Zavatta, Antonella Rigillo, and Sara Ciulli. Evidence of fish and human pathogens associated with doctor fish (Garra rufa, Heckel, 1843) used for cosmetic treatment. Journal of Fish Diseases, 42(12):1637–1644, December 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Vorbach:2019:MRC

[VPCM19] Bryan S. Vorbach, Lauren B. Peiffer, Leigh A. Clayton, and Lisa M. Mangus. Multiple recurrent cutaneous masses in a cownose ray (Rhinoptera bonasus) with progression from benign lesions to high-grade carcinoma. Journal of Fish Diseases, 42(11):1623–1627, November 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Vendramin:2012:VER

[VPP^{+}12] N. Vendramin, F. Padrós, T. Pretto, E. Cappellozza, V. Panzarin, G. Bovo, A. Toffan, and C. Terregino. Viral encephalopathy and retinopathy outbreak in restocking facilities of the endangered freshwater species, Salaria fluviatilis (Asso). Journal of Fish Diseases, 35(11):867–871, November 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Varello:2014:GLW

[VPS^{+}14] K. Varello, M. Prearo, L. Serracca, D. Meloni, I. Rossini, M. Righetti, M. Pezzolato, M. L. Fioravanti, C. Ercolini, and E. Bozzetta. Granulomatous lesions in a wild mullet population from the eastern Ligurian Sea (Italy): mycobacteriosis vs. pseudotuberculosis. Journal of Fish Diseases, 37 (6):553–558, June 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Valderrama:2019:ASI

[VSDS^{+}19a] Katherinne Valderrama, Manuel Soto-Dávila, Cristopher Segovia, Ignacio Vásquez, My Dang, and Javier Santander. Aeromonas salmonicida infects Atlantic salmon (Salmo salar) erythrocytes. Journal of Fish Diseases, 42 (11):1601–1608, November 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Valderrama:2019:CI

[VSDS^{+}19b] Katherinne Valderrama, Manuel Soto-Dávila, Cristopher Segovia, Ignacio Vásquez, My Dang, and Javier Santander. Cover image. Journal of Fish Diseases, 42(11):i, November 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Valderrama:2017:PAS

[VSS17] K. Valderrama, M. Saravia, and J. Santander. Phenotype of Aeromonas salmonicida sp. salmonicida cyclic adenosine 3’,5’-monophosphate receptor protein (Crp) mutants and its virulence in rainbow trout (Oncorhynchus mykiss). Journal of Fish Diseases, 40(12):1849–1856, December 2017. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Vu:2019:BIR

[VST^{+}19] Nguyen Thanh Vu, Nguyen Van Sang, Trinh Quoc Trong, Nguyen Huynh Duy, Nguyen Thi Dang, and Nguyen Hong Nguyen. Breeding for improved resistance to Edwardsiella ictaluri in striped catfish (Pangasianodon hypophthalmus): Quantitative genetic parameters. Journal of Fish Diseases, 42(10):1409–1417, October 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Vendramin:2014:CPS

[VTM^{+}14] N. Vendramin, A. Toffan, M. Mancin, E. Cappellozza, V. Panzarin, G. Bovo, G. Cattoli, I. Capua, and C. Terregino. Comparative pathogenicity study of ten different betanodavirus strains in experimentally infected European sea bass, Dicentrarchus labrax (L.). Journal of Fish Diseases, 37 (4):371–383, April 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Work:2014:SPH

[WA14] T. M. Work and G. S. Aeby. Skin pathology in Hawaiian goldring surgeonfish, Ctenochaetus strigosus (Bennett). Journal of Fish Diseases, 37(4):357–362, April 2014. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Whittington:2010:IIF

[WBD10a] R. J. Whittington, J. A. Becker, and M. M. Dennis. Iridovirus infections in finfish — critical review with emphasis on ranaviruses. Journal of Fish Diseases, 33(2):95–122, February 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Winton:2010:CLP

[WBD^{+}10b] J. Winton, W. Batts, P. DeKinkelin, M. LeBerre, M. Bremont, and N. Fijan. Current lineages of the Epithelioma papulosum cyprini (EPC) cell line are contaminated with fathead minnow, Pimephales promelas, cells. Journal of Fish Diseases, 33(8):701–704, August 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wunnemann:2017:FRC

[WBE^{+}17] H. Wünnemann, S. M. Bergmann, U. Eskens, A. Scharbert, M. Hundt, and M. Lierz. First report of a cystic malformation on the upper jaw of hatchery-reared allis shad Alosa alosa. Journal of Fish Diseases, 40(1):1–10, January 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2010:IHS

[WBF^{+}10] L. Wang, K. Baruah, T. Fan, M. Yu, and P. Bossier. Influence of heat shock proteins induction in different yeast cell wall mutants on the protection against Vibrio campbellii infection in gnotobiotically grown Artemia franciscana (Kellogg). Journal of Fish Diseases, 33(11):919–923, November 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wu:2012:DEP

[WCL^{+}12] M-S. Wu, C-W. Chen, C-H. Lin, C-S. Tzeng, and C-Y. Chang. Differential expression profiling of orange-spotted grouper larvae, Epinephelus coioides (Hamilton), that survived a betanodavirus outbreak. Journal of Fish Diseases, 35(3):215–225, March 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2019:IFC

[WCL^{+}19] Wenji Wang, Jianlin Chen, Baoshan Liao, Liqun Xia, Suying Hou, Zhiwen Wang, and Yishan Lu. Identification and functional characterization of Histone-like DNA-binding protein in Nocardia seriolae (NsHLP) involved in cell apoptosis. Journal of Fish Diseases, 42(5):657–666, May 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2018:EPA

[WCR^{+}18] H. Wang, Yq Chen, Gm Ru, Yq Xu, and Lq Lu. EGCG: Potential application as a protective agent against grass carp reovirus in aquaculture. Journal of Fish Diseases, 41(8): 1259–1267, August 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2016:DIM

[WCW16a] C-S. Wang, C-Y. Chang, and C-M. Wen. Developing immunological methods for detecting Macrobrachium rosenbergii nodavirus and extra small virus using a recombinant protein preparation. Journal of Fish Diseases, 39(6):715– 727, June 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wen:2016:INP

[WCW^{+}16b] C. M. Wen, M. M. Chen, C. S. Wang, P. C. Liu, and F. H. Nan. Isolation of a novel polyomavirus, related to Japanese eel endothelial cell-infecting virus, from marbled eels, Anguilla marmorata (Quoy & Gaimard). Journal of Fish Diseases, 39(7):889–897, July 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wallace:2015:SWM

[WDM^{+}15] I. S. Wallace, K. Donald, L. A. Munro, W. Murray, C. C. Pert, H. Stagg, M. Hall, and N. Bain. A survey of wild marine fish identifies a potential origin of an outbreak of viral haemorrhagic septicaemia in wrasse, Labridae, used as cleaner fish on marine Atlantic salmon, Salmo salar L., farms. Journal of Fish Diseases, 38(6):515–521, June 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wen:2016:CVS

[Wen16] C. M. Wen. Characterization and viral susceptibility of a brain cell line from brown-marbled grouper Epinephelus fuscoguttatus (Forsskål) with persistent betanodavirus infection. Journal of Fish Diseases, 39(11):1335–1346, November 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wunnemann:2018:LLC

[WEPB^{+}18] H. Wünnemann, U. Eskens, E. Prenger-Berninghoff, C. Ewers, and M. Lierz. Lactococcus lactis, causative agent of an endocarditis valvularis and parietalis thromboticans in the allis shad, Alosa alosa (L.). Journal of Fish Diseases, 41(8): 1207–1215, August 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wohlleben:2018:ESI

[WFH^{+}18] Anika Marie Wohlleben, Frederik Franke, Madeleine Hamley, Joachim Kurtz, and Jörn Peter Scharsack. Early stages of infection of three-spined stickleback (Gasterosteus aculeatus) with the cestode Schistocephalus solidus. Journal of Fish Diseases, 41(11):1701–1708, November 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Welch:2017:ETY

[WGL17] T. J. Welch, T. D. Goodrich, and S. E. LaPatra. Efficacy testing of 35-year-old commercially produced ERM bacterin reveals the remarkable stability of this product. Journal of Fish Diseases, 40(12):1921–1924, December 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wessel:2018:IPO

[WHR^{+}18] Øystein Wessel, Øyvind Haugland, Marit Rode, Børge N. Fredriksen, Maria K. Dahle, and Espen Rimstad. Inactivated Piscine orthoreovirus vaccine protects against heart and skeletal muscle inflammation in Atlantic salmon. Journal of Fish Diseases, 41(9):1411–1419, September 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wei:2019:PCH

[WIC^{+}19] Chang Wei, Hayato Iida, QiuYuan Chuah, Mikio Tanaka, Goshi Kato, and Motohiko Sano. Persistence of cyprinid herpesvirus 2 in asymptomatic goldfish Carassius auratus (L.) that survived an experimental infection. Journal of Fish Diseases, 42(6):913–921, June 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Winger:2012:RTG

[WKK12] A. C. Winger, R. Kristoffersen, and R. Knudsen. Rapid transmission of Gyrodactylus salaris (Malmberg, 1957) between live Arctic charr, Salvelinus alpinus (L.), fry. Journal of Fish Diseases, 35(10):781–784, October 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Weeks:2011:EIS

[WKW^{+}11] C. Weeks, R. Kim, M. Wolgamod, G. Whelan, and M. Faisal. Experimental infection studies demonstrate the high susceptibility of the salmonid, lake herring, Coregonus artedi (Le Sueur), to the Great Lakes strain of viral haemorrhagic septicaemia virus (genotype IVb). Journal of Fish Diseases, 34(11):887–891, November 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wen:2013:VST

[WKW13] C-M. Wen, C-C. Ku, and C-S. Wang. Viral susceptibility, transfection and growth of SPB — a fish neural progenitor cell line from the brain of snubnose pompano, Trachinotus blochii (Lacépède). Journal of Fish Diseases, 36(7):657–667, July 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2010:TTC

[WLX^{+}10] B. Wang, F. Li, J. Xiang, L. Gui, Z. Luo, and H. Yan. Three tetraspanins from Chinese shrimp, Fenneropenaeus chinensis, may play important roles in WSSV infection. Journal of Fish Diseases, 33(1):15–29, January 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2019:GPR

[WM19] Xin-An Wang and Ai-Jun Ma. Genetic parameters for resistance against Vibrio anguillarum in turbot Scophthalmus maximus. Journal of Fish Diseases, 42(5):713–720, May 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wamala:2018:MSA

[WMD^{+}18] Samuel Posian Wamala, Kizito Kahoza Mugimba, Saurabh Dubey, Abayneh Takele, Hetron Mweemba Munang’andu, Øystein Evensen, Stephen Mutoloki, Denis Karuhize Byarugaba, and Henning Sørum. Multilocus sequence analysis revealed a high genotypic diversity of Aeromonas hydrophila infecting fish in Uganda. Journal of Fish Diseases, 41(10): 1589–1600, October 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

White:2014:IIW

[WMF14] V. C. White, J. F. Morado, and C. S. Friedman. Ichthyophonus-infected walleye pollock Theragra chalcogramma (Pallas) in the eastern Bering Sea: a potential reservoir of infections in the North Pacific. Journal of Fish Diseases, 37 (7):641–655, July 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wallace:2016:DAS

[WMM^{+}16] I. S. Wallace, L. A. Munro, A. G. Murray, A. J. Christie, and N. K. G. Salama. A descriptive analysis of Scottish farmed Atlantic salmon, Salmo salar L., movements identifies a potential disease transmission risk from freshwater movements. Journal of Fish Diseases, 39(8):1021–1025, August 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wallace:2017:HRK

[WMM17] I. S. Wallace, P. McKay, and A. G. Murray. A historical review of the key bacterial and viral pathogens of Scottish wild fish. Journal of Fish Diseases, 40(12):1741–1756, December 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wiik-Nielsen:2013:GVN

[WNAF^{+}13] J. Wiik-Nielsen, M. Alarcón, B. Fineid, M. Rode, and Ø. Haugland. Genetic variation in Norwegian piscine myocarditis virus in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 36(2):129–139, February 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wiik-Nielsen:2016:VCI

[WNAJ^{+}16] J. Wiik-Nielsen, M. Alarcón, B. Bang Jensen, Ø. Haugland, and A. B. Mikalsen. Viral co-infections in farmed Atlantic salmon, Salmo salar L., displaying myocarditis. Journal of Fish Diseases, 39(12):1495–1507, December 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wortberg:2012:IYR

[WNCR12] F. Wortberg, E. Nardy, M. Contzen, and J. Rau. Identification of Yersinia ruckeri from diseased salmonid fish by Fourier transform infrared spectroscopy. Journal of Fish Diseases, 35(1):1–10, January 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wiik-Nielsen:2017:CBC

[WNGS^{+}17] J. Wiik-Nielsen, M. Gjessing, H. T. Solheim, A. Litlabø, A-G. Gjevre, A. B. Kristoffersen, M. D. Powell, and D. J. Colquhoun. Ca. Branchiomonas cysticola, Ca. Piscichlamydia salmonis and Salmon Gill Pox Virus transmit horizontally in Atlantic salmon held in fresh water. Journal of Fish Diseases, 40(10):1387–1394, October 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wiik-Nielsen:2012:CML

[WNLF^{+}12] J. Wiik-Nielsen, M. Løvoll, C. Fritsvold, A. B. Kristoffersen, Ø. Haugland, I. Hordvik, M. Aamelfot, E. Jirillo, E. O. Koppang, and S. Grove. Characterization of myocardial lesions associated with cardiomyopathy syndrome in Atlantic salmon, Salmo salar L., using laser capture microdissection. Journal of Fish Diseases, 35(12):907–916, December 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wiik-Nielsen:2016:MDP

[WNMK^{+}16] J. Wiik-Nielsen, T. A. Mo, H. Kolstad, S. N. Mohammad, S. Hytterød, and M. D. Powell. Morphological diversity of Paramoeba perurans trophozoites and their interaction with Atlantic salmon, Salmo salar L., gills. Journal of Fish Diseases, 39(9):1113–1123, September 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wright:2018:RSF

[WNO^{+}18] Daniel William Wright, Barbara Nowak, Frode Oppedal, Phil Crosbie, Lars Helge Stien, and Tim Dempster. Repeated sublethal freshwater exposures reduce the amoebic gill disease parasite, Neoparamoeba perurans, on Atlantic salmon. Journal of Fish Diseases, 41(9):1403–1410, September 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wiik-Nielsen:2012:PVR

[WNSAL12] C. R. Wiik-Nielsen, P-M. R. Ski, A. Aunsmo, and M. Løvoll. Prevalence of viral RNA from piscine reovirus and piscine myocarditis virus in Atlantic salmon, Salmo salar L., broodfish and progeny. Journal of Fish Diseases, 35(2):169–171, February 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wright:2016:ESS

[WOD16] D. W. Wright, F. Oppedal, and T. Dempster. Early-stage sea lice recruits on Atlantic salmon are freshwater sensitive. Journal of Fish Diseases, 39(10):1179–1186, October 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wootten:2012:E

[Woo12] R. Wootten. Editorial. Journal of Fish Diseases, 35(9):621, September 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2013:PSD

[WPH^{+}13] K. Y. Wang, C. Z. Peng, J. L. Huang, Y. D. Huang, M. C. Jin, and Y. Geng. The pathology of selenium deficiency in Cyprinus carpio L. Journal of Fish Diseases, 36(7):609–615, July 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Winton:2010:O

[WR10] Jim Winton and Ronald J. Roberts. Obituary. Journal of Fish Diseases, 33(10):787–788, October 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wonmongkol:2018:GPR

[WSA^{+}18] P. Wonmongkol, S. Sukhavachana, K. Ampolsak, P. Srisapoome, T. Suwanasopee, and S. Poompuang. Genetic parameters for resistance against Flavobacterium columnare in Nile tilapia Oreochromis niloticus (Linnaeus, 1758). Journal of Fish Diseases, 41(2):321–328, February 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2018:RVD

[WSX^{+}18] H. Wang, M. Sun, D. Xu, P. Podok, J. Xie, Ys Jiang, and Lq Lu. Rapid visual detection of cyprinid herpesvirus 2 by recombinase polymerase amplification combined with a lateral flow dipstick. Journal of Fish Diseases, 41(8):1201–1206, August 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Weyl:2014:ALE

[WTM14] O. L. F. Weyl, H. Taraschewski, and F. Moravec. Are African longfin eel imports a potential pathway for parasite invasions? Journal of Fish Diseases, 37(9):843–845, September 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Whyte:2013:FDA

[WWE^{+}13] S. K. Whyte, J. D. Westcott, A. Elmoslemany, K. L. Hammell, and C. W. Revie. A fixed-dose approach to conducting emamectin benzoate tolerance assessments on field-collected sea lice, Lepeophtheirus salmonis. Journal of Fish Diseases, 36(3):283–292, March 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2011:ECV

[WWS^{+}11] N. Wang, X-L. Wang, Z-X. Sha, Y-S. Tian, and S-L. Chen. Establishment, characterization and virus susceptibility of a kidney-derived cell line from southern flounder, Paralichthys lethostigma Jordan & Gilbert. Journal of Fish Diseases, 34 (1):81–85, January 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2018:ECC

[WWZ^{+}18] Yingying Wang, Qing Wang, Weiwei Zeng, Jiyuan Yin, Yingying Li, Yan Ren, Cunbin Shi, Sven M. Bergmann, and Xinping Zhu. Establishment and characterization of a cell line from tilapia brain for detection of tilapia lake virus. Journal of Fish Diseases, 41(12):1803–1809, December 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2016:DCH

[WXL16] H. Wang, Lj Xu, and Lq Lu. Detection of cyprinid herpesvirus 2 in peripheral blood cells of silver crucian carp, Carassius auratus gibelio (Bloch), suggests its potential in viral diagnosis. Journal of Fish Diseases, 39(2):155–162, February 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wu:2015:IEH

[WXWD15] X. Wu, H. Xiong, Y. Wang, and H. Du. Immunomodulatory effects of hyperthermia on resisting WSSV infection in Procambarus clarkii. Journal of Fish Diseases, 38(6):567– 574, June 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2019:RMG

[WYZ^{+}19] Suyun Wang, Qingpi Yan, Meimei Zhang, Lixing Huang, Leilei Mao, Mengmeng Zhang, Xiaojin Xu, Liwei Chen, and Yingxue Qin. The role and mechanism of icmF in Aeromonas hydrophila survival in fish macrophages. Journal of Fish Diseases, 42(6):895–904, June 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2010:CAP

[WZA10] Y. Wang, X. H. Zhang, and B. Austin. Comparative analysis of the phenotypic characteristics of high- and low-virulent strains of Edwardsiella tarda. Journal of Fish Diseases, 33 (12):985–994, December 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2011:OIM

[WZL^{+}11] Q. Wang, W. W. Zeng, K. B. Li, O. Q. Chang, C. Liu, G. H. Wu, C. B. Shi, and S. Q. Wu. Outbreaks of an iridovirus in marbled sleepy goby, Oxyeleotris marmoratus (Bleeker), cultured in southern China. Journal of Fish Diseases, 34 (5):399–402, May 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2013:NIT

[WZL13] J. Wang, L. L. Zou, and A. X. Li. A novel iron transporter in Streptococcus iniae. Journal of Fish Diseases, 36(12):1007– 1015, December 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2016:NPB

[WZQ^{+}16] X. Wang, Y. Zhang, G. Qin, W. Luo, and Q. Lin. A novel pathogenic bacteria (Vibrio fortis) causing enteritis in cultured seahorses, Hippocampus erectus Perry, 1810. Journal of Fish Diseases, 39(6):765–769, June 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2018:ECN

[WZW^{+}18] Y. Wang, W. Zeng, Q. Wang, Y. Li, S. M. Bergmann, S. Zheng, Y. Ren, C. Liu, O. Chang, and P. Lee. Establishment and characterization of a new cell line from koi brain (Cyprinus carpio L.). Journal of Fish Diseases, 41(2):

357–364, February 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Xiao:2015:SNS

[XJL^{+}15] P. Xiao, M. Jiang, Y. Liu, M. Sun, L. Zhang, L. Jie, G. Li, and Z. Mo. Splenic necrosis signs and pathogen detection in cultured half-smooth tongue sole, Cynoglossus semilaevis Günther. Journal of Fish Diseases, 38(1):103–106, January 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Xiang:2019:CTA

[XJL^{+}19] Yangxi Xiang, Peng Jia, Wei Liu, Meisheng Yi, and Kuntong Jia. Comparative transcriptome analysis reveals the role of p53 signalling pathway during red-spotted grouper nervous necrosis virus infection in Lateolabrax japonicus brain cells. Journal of Fish Diseases, 42(4):585–595, April 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Xueqin:2012:CEF

[XKB12] J. Xueqin, P. W. Kania, and K. Buchmann. Comparative effects of four feed types on white spot disease susceptibility and skin immune parameters in rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 35 (2):127–135, February 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Xu:2012:STS

[XKBC12] D-H. Xu, P. H. Klesius, B. G. Bosworth, and N. Chatakondi. Susceptibility of three strains of blue catfish, Ictalurus furcatus (Valenciennes), to Ichthyophthirius multifiliis. Journal of Fish Diseases, 35(12):887–895, December 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Xu:2017:OEM

[XLZ^{+}17] L-W. Xu, X-H. Liu, J-Y. Zhang, G-F. Liu, and J. Feng. Outbreak of enteric microsporidiosis of hatchery-bred juvenile groupers, Epinephelus spp., associated with a new intranuclear microporidian in China. Journal of Fish Diseases, 40 (2):183–189, February 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Xu:2015:MMA

[XQS^{+}15] W. J. Xu, Z. D. Qin, H. Shi, N. Jiang, Y. Zhou, X. L. Liu, J. J. Xie, G. S. Wang, W. M. Wang, M. Asim, L. B. Zeng, and L. Lin. Mass mortality associated with a viral-induced anaemia in cage-reared large yellow croaker, Larimichthys crocea (Richardson). Journal of Fish Diseases, 38(5):499– 502, May 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Xue:2018:ECL

[XWP^{+}18] T. Xue, Y. Z. Wang, Q. H. Pan, Q. Wang, J. F. Yuan, and T. S. Chen. Establishment of a cell line from the kidney of black carp and its susceptibility to spring viremia of carp virus. Journal of Fish Diseases, 41(2):365–374, February 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Xu:2019:VRC

[XXL^{+}19] Jun Xu, Yun-Dan Xie, Ling Liu, Song Guo, You-Lu Su, and An-Xing Li. Virulence regulation of cel-EIIB protein mediated PTS system in Streptococcus agalactiae in Nile tilapia. Journal of Fish Diseases, 42(1):11–19, January 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Xu:2016:TVM

[XZZ16] D-H. Xu, Q-Z. Zhang, and D. Zhang. Two in vitro methods for screening potential parasiticides against Ichthyophthirius multifiliis using Tetrahymena thermophila. Journal of Fish Diseases, 39(3):285–294, March 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yatabe:2011:RFA

[YAHWU11] T. Yatabe, G. Arriagada, C. Hamilton-West, and S. Urcelay. Risk factor analysis for sea lice, Caligus rogercresseyi, levels in farmed salmonids in southern Chile. Journal of Fish Diseases, 34(5):345–354, May 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yemmen:2017:SAO

[YB17] C. Yemmen and S. Bahri. Seasonality of Amyloodinium ocellatum Brown 1931 (Dinophyceae) infesting the Senegalese sole Solea senegalensis from Bizerte lagoon, Tunisia. Journal of Fish Diseases, 40(6):853–857, June 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yeh:2018:CVS

[YCNW18] S. W. Yeh, Y. H. Cheng, F. N. Nan, and C. M. Wen. Characterization and virus susceptibility of a continuous cell line derived from the brain of Aequidens rivulatus (Günther). Journal of Fish Diseases, 41(4):635–641, April 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yardimci:2016:IEE

[YDO^{+}16] B. Yardimci, B. I. Didinen, E. E. Onuk, S. Metin, A. Ciftci, A. Kubilay, G. Z. Pekmezci, and H. Eralp. Immunohistochemical evaluation of experimental Vagococcus salmoninarum infection in rainbow trout (Oncorhynchus mykiss, Walbaum 1792). Journal of Fish Diseases, 39(5):547–553, May 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ye:2017:CAW

[YEK^{+}17] F. Ye, R. H. Easy, S. D. King, D. K. Cone, and P. You. Comparative analyses within Gyrodactylus (Platyhelminthes: Monogenea) mitochondrial genomes and conserved polymerase chain reaction primers for gyrodactylid mitochondrial DNA. Journal of Fish Diseases, 40(4):541–555, April 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yamin:2017:PEH

[YFA^{+}17] G. Yamin, R. Falk, R. R. Avtalion, N. Shoshana, T. Ofek, R. Smirnov, G. Rubenstein, and J. van Rijn. The protective effect of humic-rich substances on atypical Aeromonas salmonicida subsp . salmonicida infection in common carp (Cyprinus carpio L.). Journal of Fish Diseases, 40(12):1783– 1790, December 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yan:2018:IHS

[YGL^{+}18] Wanli Yan, Lide Gu, Shu Liu, Wei Ren, Mingsheng Lyu, and Shujun Wang. Identification of a highly specific DNA aptamer for Vibrio vulnificus using systematic evolution of ligands by exponential enrichment coupled with asymmetric PCR. Journal of Fish Diseases, 41(12):1821–1829, December 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yan:2016:CIH

[YHY^{+}16] D. C. Yan, J. Huang, B. Yang, H. S. Sun, Y. Y. Wang, and X. Liu. Competition of infectious hypodermal and haematopoietic necrosis virus (IHHNV) with white spot syndrome virus (WSSV) for binding to shrimp cellular membrane. Journal of Fish Diseases, 39(10):1225–1229, October 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yeh:2012:CEC

[YK12] H-Y. Yeh and P. H. Klesius. Construction, expression and characterization of 11 putative flagellar apparatus genes of Aeromonas hydrophila AL09-73. Journal of Fish Diseases, 35(11):853–860, November 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yeh:2013:CSM

[YK13] H. Y. Yeh and P. H. Klesius. Changes of serum myeloperoxidase and nitric oxide in the early stage of Edwardsiella ictaluri infection in channel catfish, Ictalurus punctatus (Rafinesque). Journal of Fish Diseases, 36(4):441–446, April 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yoshida:2013:IKH

[YkSKY13] N. Yoshida, R k Sasaki, H. Kasai, and M. Yoshimizu. Inactivation of koi-herpesvirus in water using bacteria isolated from carp intestines and carp habitats. Journal of Fish Diseases, 36(12):997–1005, December 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yang:2018:APT

[YLL^{+}18] Ying-Fei Yang, Tien-Hsuan Lu, Hsing-Chieh Lin, Chi-Yun Chen, and Chung-Min Liao. Assessing the population transmission dynamics of tilapia lake virus in farmed tilapia. Journal of Fish Diseases, 41(9):1439–1448, September 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yu:2019:SCS

[YLS^{+}19] Qing Yu, Mingzhu Liu, Hongfei Su, Hehe Xiao, Siting Wu, Xianling Qin, Siqiao Li, Huizhi Mi, Zijun Lu, Deqiang Shi, and Pengfei Li. Selection and characterization of ssDNA aptamers specifically recognizing pathogenic Vibrio alginolyticus. Journal of Fish Diseases, 42(6):851–858, June 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yan:2014:IPP

[YLSW14] D. C. Yan, H. L. Liu, H. S. Sun, and Y. Y. Wang. Investigation of possible presence of infectious myonecrosis virus in shrimp in China. Journal of Fish Diseases, 37(7):679– 682, July 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ye:2014:UMH

[YLW^{+}14] Lt Ye, Wx Li, Ww Wang, Sg Wu, and Gt Wang. Updated morphology, histopathology and molecular phylogeny of Myxobolus hearti, cardiac myxosporea in gibel carp, Carassius gibelio (Bloch). Journal of Fish Diseases, 37(1): 11–20, January 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yu:2019:DNA

[YLX^{+}19a] Qing Yu, Mingzhu Liu, Hehe Xiao, Siting Wu, Xianling Qin, Ke Ke, Siqiao Li, Huizhi Mi, Deqiang Shi, and Pengfei Li. Development of novel aptamer-based enzyme-linked aptasorbent assay (ELASA) for rapid detection of mariculture pathogen Vibrio alginolyticus. Journal of Fish Diseases, 42 (11):1523–1529, November 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yu:2019:IAA

[YLX^{+}19b] Qing Yu, Mingzhu Liu, Hehe Xiao, Siting Wu, Xianling Qin, Zijun Lu, Deqiang Shi, Siqiao Li, Huizhi Mi, Yibing Wang, Hongfei Su, Taixia Wang, and Pengfei Li. The inhibitory activities and antiviral mechanism of Viola philippica aqueous extracts against grouper iridovirus infection in vitro and in vivo. Journal of Fish Diseases, 42(6):859–868, June 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yemmen:2013:MSP

[YMBE13] C. Yemmen, S. Marton, S. Bahri, and E. Eszterbauer. Morphology, seasonality and phylogeny of Zschokkella soleae sp. n. (Myxozoa, Myxosporea) parasite of Solea solea (L.) (Pleuronectiformes, Soleidae) from Ghar El Melh Lagoon, Tunisia. Journal of Fish Diseases, 36(10):871–879, October 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yingjie:2011:ACC

[YMH^{+}11] S. Yingjie, Z. Min, L. Hong, Y. Zhiqin, Z. Xiaocong, and W. Zhe. Analysis and characterization of the complete genomic sequence of the Chinese strain of hirame rhabdovirus. Journal of Fish Diseases, 34(2):167–171, February 2011. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yan:2011:ECV

[YNL11] W. Yan, P. Nie, and Y. Lu. Establishment, characterization and viral susceptibility of a new cell line derived from goldfish, Carassius auratus (L.), tail fin. Journal of Fish Diseases, 34(10):757–768, October 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yousaf:2017:CBA

[YP17] M. N. Yousaf and T. T. Poppe. Cartilage in the bulbus arteriosus of farmed Atlantic salmon (Salmo salar L.). Journal of Fish Diseases, 40(9):1249–1252, September 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yamamoto:2018:EGD

[YSHG18] Fernando Y. Yamamoto, Fernando J. Sutili, Michael Hume, and Delbert M. Gatlin I.I.I. The effect of β-1,3-glucan derived from Euglena gracilis (Algamune^{T M}) on the innate immunological responses of Nile tilapia (Oreochromis niloticus L.). Journal of Fish Diseases, 41(10):1579–1588, October 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yeh:2013:CAC

[YSK13] H-Y. Yeh, C. A. Shoemaker, and P. H. Klesius. Chemotactic activity of channel catfish, Ictalurus punctatus (Rafinesque), recombinant cyclophilin A. Journal of Fish Diseases, 36(12): 1041–1046, December 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yanez:2013:TNB

[YSV^{+}13] A. J. Yañez, H. Silva, K. Valenzuela, J. P. Pontigo, M. Godoy, J. Troncoso, A. Romero, J. Figueroa, J. G. Carcamo, and R. Avendaño-Herrera. Two novel blood-free solid media for the culture of the salmonid pathogen Piscirickettsia salmonis. Journal of Fish Diseases, 36(6):587–591, June 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yamkasem:2019:EPV

[YTKS19] Jidapa Yamkasem, Puntanat Tattiyapong, Attapon Kamlangdee, and Win Surachetpong. Evidence of potential vertical transmission of tilapia lake virus. Journal of Fish Diseases, 42(9):1293–1300, September 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yan:2010:RTP

[YTL10] D. C. Yan, K. F. J. Tang, and D. V. Lightner. A real-time PCR for the detection of hepatopancreatic parvovirus (HPV) of penaeid shrimp. Journal of Fish Diseases, 33(6):507–511, June 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yanez:2014:BMP

[YVM^{+}14] A. J. Yáñez, K. Valenzuela, C. Matzner, V. Olavarría, J. Figueroa, R. Avendaño-Herrera, and J. G. Carcamo. Broth microdilution protocol for minimum inhibitory concentration (MIC) determinations of the intracellular salmonid pathogen Piscirickettsia salmonis to florfenicol and oxytetracycline. Journal of Fish Diseases, 37(5):505–509, May 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yin:2019:DSR

[YWW^{+}19] Jiyuan Yin, Qing Wang, Yingying Wang, Yingying Li, Weiwei Zeng, Jiexing Wu, Yan Ren, Yafang Tang, Caixia Gao, Huzi Hu, and Sven M. Bergmann. Development of a simple and rapid reverse transcription–loopmediated isothermal amplification (RT-LAMP) assay for sensitive detection of tilapia lake virus. Journal of Fish Diseases, 42(6):817–824, June 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yokoyama:2010:MDM

[YYF^{+}10] H. Yokoyama, T. Yanagida, M. A. Freeman, T. Katagiri, A. Hosokawa, M. Endo, M. Hirai, and S. Takagi. Molecular diagnosis of Myxobolus spirosulcatus associated with encephalomyelitis of cultured yellowtail, Seriola quinqueradiata Temminck & Schlegel. Journal of Fish Diseases, 33(12): 939–946, December 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yu:2014:GMI

[YYW^{+}14] H. Yu, Q. G. Yan, Z. B. Wang, Y. J. Lu, M. J. Xu, H. Li, and X. Q. Zhu. MH II-DAB gene expression in grass carp Ctenopharyngodon idella (Valenciennes) after infection with the ciliate parasite, Ichthyophthirius multifiliis. Journal of Fish Diseases, 37(1):43–50, January 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yang:2019:MLP

[YZW^{+}19] Bintong Yang, Dongxing Zhang, Tonglei Wu, Zhiqiang Zhang, Sayed Haidar Abbas Raza, Nicola Schreurs, Lei Zhang, Guilian Yang, Chunfeng Wang, Aidong Qian, Yuanhuan Kang, and Xiaofeng Shan. Maltoporin (LamB protein) contributes to the virulence and adhesion of Aeromonas veronii TH0426. Journal of Fish Diseases, 42(3):379–389, March 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhang:2012:IPC

[ZB12] S. Zhang and J-R. Bonami. Isolation and partial characterization of a new reovirus in the Chinese mitten crab, Eriocheir sinensis H Milne Edwards. Journal of Fish Diseases, 35(10): 733–739, October 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zainathan:2015:DAM

[ZCC^{+}15] S. C. Zainathan, G. Carlile, J. Carson, K. A. McColl, M. St. J. Crane, L. M. Williams, J. Hoad, N. J. G. Moody, H. M. Aiken, G. F. Browning, and B. F. Nowak. Development and application of molecular methods (PCR) for detection of Tasmanian Atlantic salmon reovirus. Journal of Fish Diseases, 38(8):739–754, August 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhou:2018:FRG

[ZCC18] Q. J. Zhou, F. C. Chai, and J. Chen. First record of Glugea plecoglossi (Takahashi & Egusa, 1977), a microsporidian parasite of ayu (Plecoglossus altivelis altivelis Temminck & Schlegel, 1846) in Mainland China. Journal of Fish Diseases, 41(1):165–169, January 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zainathan:2013:LES

[ZCCN13] S. C. Zainathan, J. Carson, M. St J. Crane, and B. F. Nowak. Laboratory evaluation of sample collection methods (organs vs swabs) for Tasmanian salmon reovirus detection in farmed Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 36 (4):427–436, April 2013. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhao:2018:GMB

[ZCJ^{+}18] Xianliang Zhao, He Chen, Zhaohui Jin, Li Li, Jie Zhang, and Xianghui Kong. GC-MS-based metabolomics analysis reveals L-aspartate enhances the antibiotic sensitivity of neomycin sulfate-resistant Aeromonas hydrophila. Journal of Fish Diseases, 41(12):1831–1841, December 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zerihun:2012:EMA

[ZCP12] M. A. Zerihun, D. J. Colquhoun, and T. T. Poppe. Experimental mycobacteriosis in Atlantic cod, Gadus morhua L. Journal of Fish Diseases, 35(5):365–377, May 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhao:2015:MQS

[ZCQF15] J. Zhao, M. Chen, Cs Quan, and Sd Fan. Mechanisms of quorum sensing and strategies for quorum sensing disruption in aquaculture pathogens. Journal of Fish Diseases, 38(9): 771–786, September 2015. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhu:2018:HPP

[ZDWH18] Z. M. Zhu, C. F. Dong, S. P. Weng, and J. G. He. The high prevalence of pathogenic Vibrio harveyi with multiple antibiotic resistance in scale drop and muscle necrosis disease of the hybrid grouper, Epinephelus fuscoguttatus ([female sign]) × E. lanceolatus ([male sign]), in China. Journal of Fish Diseases, 41(4):589–601, April 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhou:2011:RIS

[ZFZ^{+}11] S. M. Zhou, Y. Fan, X. Q. Zhu, M. Q. Xie, and A. X. Li. Rapid identification of Streptococcus iniae by specific PCR assay utilizing genetic markers in ITS rDNA. Journal of Fish Diseases, 34(4):265–271, April 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhang:2018:NBO

[ZGY^{+}18] Liang Zhang, Zhuying Gao, Li Yu, Bo Zhang, Jing Wang, and Jun Zhou. Nucleotide-binding and oligomerization domain (NOD)-like receptors in teleost fish: Current knowledge and future perspectives. Journal of Fish Diseases, 41 (9):1317–1330, September 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zerihun:2011:ITR

[ZHFC11] M. A. Zerihun, M. J. Hjortaas, K. Falk, and D. J. Colquhoun. Immunohistochemical and Taqman real-time PCR detection of mycobacterial infections in fish. Journal of Fish Diseases, 34(3):235–246, March 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zha:2018:PVF

[ZJDL18] H. Zha, A. Jeffs, Y. Dong, and G. Lewis. Potential virulence factors of bacteria associated with tail fan necrosis in the spiny lobster, Jasus edwardsii. Journal of Fish Diseases, 41 (5):817–828, May 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhang:2018:DTA

[ZJR^{+}18] Qinghua Zhang, Ce Ji, Jianfeng Ren, Qiuyue Zhang, Xuehong Dong, Yao Zu, Liang Jia, and Weiming Li. Differential transcriptome analysis of zebrafish (Danio rerio) larvae challenged by Vibrio parahaemolyticus. Journal of Fish Diseases, 41(7):1049–1062, July 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhang:2019:CPS

[ZKL^{+}19] Defeng Zhang, Xiaoli Ke, Zhigang Liu, Jianmeng Cao, Youlu Su, Maixin Lu, Fengying Gao, Miao Wang, Mengmeng Yi, and Fengling Qin. Capsular polysaccharide of Streptococcus agalactiae is an essential virulence factor for infection in Nile tilapia (Oreochromis niloticus Linn.). Journal of Fish Diseases, 42(2):293–302, February 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhou:2017:RSD

[ZLN^{+}17] L. Zhou, P. Li, S. Ni, Y. Yu, M. Yang, S. Wei, and Q. Qin. Rapid and sensitive detection of redspotted grouper nervous necrosis virus (RGNNV) infection by aptamer–coat protein–aptamer sandwich enzyme-linked apta-sorbent assay (ELASA). Journal of Fish Diseases, 40(12):1831–1838, December 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhang:2019:IIV

[ZLX^{+}19] Wanwan Zhang, Zelin Li, Yangxi Xiang, Peng Jia, Wei Liu, Meisheng Yi, and Kuntong Jia. Isolation and identification of a viral haemorrhagic septicaemia virus (VHSV) isolate from wild largemouth bass Micropterus salmoides in China. Journal of Fish Diseases, 42(11):1563–1572, November 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zahran:2010:ESA

[ZN10] E. Zahran and E. J. Noga. Evidence for synergism of the antimicrobial peptide piscidin 2 with antiparasitic and antioomycete drugs. Journal of Fish Diseases, 33(12):995– 1003, December 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zerihun:2011:MSIa

[ZNHC11a] M. A. Zerihun, H. Nilsen, S. Hodneland, and D. J. Colquhoun. Mycobacterium salmoniphilum infection in farmed Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 34(10):769–781, October 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zerihun:2011:MSIb

[ZNHC11b] M. A. Zerihun, H. Nilsen, S. Hodneland, and D. J. Colquhoun. Mycobacterium salmoniphilum infection in farmed Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 34(12):959–960, December 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zuskova:2018:ETP

[ZPM^{+}18] E. Zuskova, V. Piackova, J. Machova, L. Chupani, C. Steinbach, A. Stara, and J. Velisek. Efficacy and toxicity of praziquantel in helminth-infected barbel (Barbus barbus L.). Journal of Fish Diseases, 41(4):643–649, April 2018. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhang:2011:MPC

[ZQB^{+}11] X-J. Zhang, G-M. Qin, X-W. Bing, B-L. Yan, and L-G. Liang. Molecular and phenotypic characterization of Vibrio aestuarianus, a pathogen of the cultured tongue sole, Cynoglossus semilaevis Günther. Journal of Fish Diseases, 34(1):57–64, January 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhang:2016:IIN

[ZSZF16] S. Zhang, X. Shu, L. Zhou, and B. Fu. Isolation and identification of a new reovirus associated with mortalities in farmed oriental river prawn, Macrobrachium nipponense (de Haan, 1849), in China. Journal of Fish Diseases, 39(3):371– 375, March 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zilberg:2010:DLR

[ZTF^{+}10] D. Zilberg, A. Tal, N. Froyman, S. Abutbul, N. Dudai, and A. Golan-Goldhirsh. Dried leaves of Rosmarinus officinalis as a treatment for streptococcosis in tilapia. Journal of Fish Diseases, 33(4):361–369, April 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhang:2017:ICP

[ZWG^{+}17] J. Zhang, Y. Wang, H. Guo, Z. Mao, and C. Ge. Identification and characterization of a phospholipase A1 activity type three secreted protein, PP ExoU from Pseudomonas plecoglossicida NB2011, the causative agent of visceral granulomas disease in large yellow croaker (Larimichthys crocea). Journal of Fish Diseases, 40(6):831–840, June 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zeng:2018:DVR

[ZWG^{+}18] Weiwei Zeng, Yingying Wang, Yanmin Guo, Sven M. Bergmann, Jiyuan Yin, Yingying Li, Yan Ren, Cunbin Shi, and Qing Wang. Development of a VP38 recombinant protein-based indirect ELISA for detection of antibodies against grass carp reovirus genotype II (iELISA for detection of antibodies against GCRV II). Journal of Fish Diseases, 41(12):1811–1819, December 2018. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhou:2014:CKC

[ZWL^{+}14] J-X. Zhou, H. Wang, X-W. Li, X. Zhu, W-L. Lu, and D-M. Zhang. Construction of KHV-CJ ORF25 DNA vaccine and immune challenge test. Journal of Fish Diseases, 37(4):319– 325, April 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhu:2016:IVP

[ZWL^{+}16] M. Zhu, X. R. Wang, J. Li, G. Y. Li, Z. P. Liu, and Z. L. Mo. Identification and virulence properties of Aeromonas veronii bv. sobria isolates causing an ulcerative syndrome of loach Misgurnus anguillicaudatus. Journal of Fish Diseases, 39 (6):777–781, June 2016. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhang:2010:IMT

[ZWLG10] J. Y. Zhang, J. G. Wang, A. H. Li, and X. N. Gong. Infection of Myxobolus turpisrotundus sp. n. in allogynogenetic gibel carp, Carassius auratus gibelio (Bloch), with revision of Myxobolus rotundus (s. l.) Nemeczek reported from C. auratus auratus (L.). Journal of Fish Diseases, 33(8):625–638, August 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhang:2011:EPH

[ZWY^{+}11] X. Z. Zhang, Z. H. Wu, S. P. Yang, H. Y. Pang, J. C. Jian, and Y. S. Lu. Expression pattern of heat-shock cognate 70 gene of humphead snapper, Lutjanus sanguineus (Cuvier), infected by Vibrio harveyi. Journal of Fish Diseases, 34(10): 719–729, October 2011. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zheng:2012:DAE

[ZXW^{+}12] Y. Zheng, Y. Xiao, H. Wu, Q. Wang, J. Xiao, Y. Zhang, and Q. Liu. Different approaches to expressing Edwardsiella tarda antigen GAPDH in attenuated Vibrio anguillarum for multivalent fish vaccines. Journal of Fish Diseases, 35(8): 569–577, August 2012. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhang:2019:FRI

[ZxYJ^{+}19] Zheng Zhang, Yong xiang Yu, Yong Jiang, Yin geng Wang, Mei jie Liao, Xiao jun Rong, Kai Wang, Hao Zhang, and Jing Chen. First report of isolation and complete genome of Vibrio rotiferianus strain SSVR1601 from cage-cultured black rockfish (Sebastes schlegelii) associated with skin ulcer. Journal of Fish Diseases, 42(5):623–630, May 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhang:2010:UTH

[ZYW^{+}10] J. Y. Zhang, H. Yokoyama, J. G. Wang, A. H. Li, X. N. Gong, A. Ryu-Hasegawa, M. Iwashita, and K. Ogawa. Utilization of tissue habitats by Myxobolus wulii Landsberg & Lom, 1991 in different carp hosts and disease resistance in allogynogenetic gibel carp: redescription of M. wulii from China and Japan. Journal of Fish Diseases, 33(1):57–68, January 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhang:2014:VGF

[ZZA^{+}14] J. T. Zhang, S. M. Zhou, S. W. An, L. Chen, and G. L. Wang. Visceral granulomas in farmed large yellow croaker, Larimichthys crocea (Richardson), caused by a bacterial pathogen, Pseudomonas plecoglossicida. Journal of Fish Diseases, 37 (2):113–121, February 2014. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zago:2017:VAF

[ZZC^{+}17] V. Zago, M. Zambon, M. Civettini, O. Zaltum, and A. Manfrin. Virulence-associated factors in Vibrio cholerae non– O1/ non-O139 and V. mimicus strains isolated in ornamental fish species. Journal of Fish Diseases, 40(12):1857–1868, December 2017. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhang:2019:GTG

[ZZQ^{+}19] Bo Zhang, Huixian Zhang, Geng Qin, Yuhong Liu, Xue Han, Jianping Yin, and Qiang Lin. TLR2 gene in seahorse brood pouch plays key functional roles in LPS-induced antibacterial responses. Journal of Fish Diseases, 42(7):1085–1089, July 2019. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).